A Bibliography of Papers in Lecture Notes in Computer Science (1998), Part 1 of 2

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

22 March 2018 Version 1.24

Title word cross-reference

(1/6) log log n [3127]. (2^{8}) [2682]. + [1385, 2282]. 0 [1742, 1724]. 1 [1423, 1742, 1554, 1724]. 2 [1831, 452, 1729, 1730, 1609, 1554, 1597, 2887, 1610, 989, 2193, 2595, 2258]. 21/2 [1611]. 3 [1617, 490, 1560, 2596, 1609, 2597, 1403, 1597, 387, 1664, 2594, 2317, 1666, 989, 2318, 514, 1661, 1578, 2319, 2595, 990, 1646, 1138, 1648, 1662, 1641, 2174, 3080]. 4 [2166, 1354, 660, 1755]. ^{0} [3128, 2805]. ^{3} [2684]. ^{A} [2102]. _{2} [411]. _{R} [2568]. □ [1210]. c_{0} [2161]. c_{1} [2161]. c_{2} [2161]. ϵ [3115]. F_{−−} [269]. ∀∃^{∗} [3151]. GF(p) [2500]. k [2529, 1494, 2885, 2857, 462, 2541, 436, 2892, 288]. L [2186, 1494]. λ [430, 2101]. λμ [3121]. ≤ [3177]. m [2161]. m^{2} [2161]. m^{3} [2161]. Z_{p} [3177]. L_{π} [686]. μ [761]. N [1344, 779, 3063, 1463]. O(n log^{2}(n)) [3115]. Ω [508]. p [1408, 2818]. π [2859, 2838, 2858]. Π_{1}^{0} [3151]. q [2818]. qAC00[2] [3070]. S [2197]. T [1804, 2888]. T^{A}P [2092]. C⃗_{k} [659]. W^{4}S [1611]. Z_{m} [2860]. Z_{n} [1463].

-Approximation [452]. -approximations [2887]. -Body [1344, 3063]. -Calculus [3121, 2838, 761, 430, 2859, 2858]. -clique-free [2892]. -code [2684]. -Colorable [3080]. -Connectivity [2857]. -Cube [660]. -D [1661, 990, 1646, 1641]. -Edge [1729, 1730]. -Equational [3151]. -Free [3115]. -function [2186]. -Hard [3151]. -integral [2197]. -Learning [1277].

-outconnected [2885]. -power [2193]. -Process [779]. -threads [1408]. -Times [436]. -tree [2529]. -trees [462]. -Variable [2541]. -vertex [2888].

0-balanced [474].

1 [3081]. 1-Hamiltonian [3081]. 1.x [2399]. 16-bit [2500]. 18th [2134]. 18th-century [2134]. 1998 [1298]. 1Bit [204].

2 [3107, 1427, 1440, 288]. 2-Tape [3107]. 2.0 [2093, 2317]. 21st [1321]. 2D [459].

3 [2967, 748, 3064]. 3-Flip [3064]. 3-SAT [2967]. 3-stage [3074].

4 [1355].

5 [1018, 1437, 2322].

6 [1018]. 6-PARITY [1018].

7th [88].

8 [1831]. ’81 [2728]. ’82 [2729, 2730]. ’83 [2731, 2732]. ’84 [2733, 2734]. ’85 [2735, 2736]. ’86 [2737, 2738]. ’87 [2739, 2740]. ’88 [2742, 2741]. ’89 [2744, 2743].

’90 [2745, 2746, 269]. 900 [1419]. ’91 [2747, 2748]. ’92 [2473, 2750, 2749]. ’93 [2751, 2752]. ’94 [2753, 2754]. ’95 [2756, 2755, 1707]. ’96 [2758, 2757]. ’97 [2759, 2760, 1125, 1153, 1132, 1142].

A0* [1347]. Abacus [916]. ABB [2412]. Abbadingo [2540]. ABC02 [1163]. Abelian [2705, 2186, 3118, 2105, 2106]. Ability [3027, 2767]. Able [1808]. Abridged [701]. Absence [568]. Abstract [770, 105, 1504, 2884, 2160, 1478, 1464, 3132, 544, 3124, 106, 239, 2192, 751, 1490, 624, 3133, 1280, 2622, 1068, 2179, 738]. Abstracting [339]. Abstraction [267, 779, 3094]. Abstractions [2381, 2379, 2629, 771, 20, 43, 1704, 25]. Abstractness [1789]. AC [3128, 3110, 2805, 2411, 2845]. AC-Compatible [2845]. AC-Termination [3110]. AC00 [433]. Academic [1688, 1678]. Academy [785]. Access [2695, 2694, 808, 569, 2032, 960, 1038, 2696, 2697, 1364, 1497]. Accessible [2029]. According [1919]. Accountable [2448]. Accounting [2450, 1043]. Accumulations [2143]. Accuracy [958, 1939]. Accurate [2046, 1561]. achieve [1561]. Achievements [1034, 109]. ACL2 [930, 2113, 2354]. Acoi [1536]. Acquiring [1143]. Acquisition [1073, 2962, 2018, 383, 1854, 1241]. Acrobot [824]. Across [2444, 168, 2338]. Action [1780, 2136, 2030, 497, 1156, 3109, 227, 2137, 130, 131]. Actions [1961, 595, 596, 1509, 1508]. Active [68, 1647, 1668, 1653, 787, 1863, 1664, 2496, 1649, 528]. Actively [2624]. Activities [209, 349, 485]. Activity [2127]. Actors [14]. Acyclic [2929]. Ad [601]. Ad-Hoc [601]. Ada [1694, 1715, 1716, 1693, 1696, 1703, 1720, 1701, 1714, 1707, 1702, 1700, 1698, 1710]. Ada95 [1719, 1705, 1706]. Adam [3079]. Adaptable [100, 101, 102, 2457]. Adaptation [917, 3018, 2969, 2910, 2983, 198, 1944, 1306, 103, 1581]. Adaptative [1988]. Adapting [1756, 85, 2688]. Adaptive [2962, 1442, 1027, 2968, 186, 16, 1431, 363, 1996, 15, 992, 310, 956, 2999, 1007, 2622, 2457, 1608, 400, 951, 1645, 1933, 1607, 1620, 1630]. Adaptivity [2621]. Added [789, 1030]. Added-Value [1030]. Adders [2395]. Adding [543, 53]. Addition [84]. Additive [3073, 3177]. Additively [2537]. Additivity [113]. ADF [3031]. Adjacency [435]. Administration [1920]. Admissibility [2116]. Admissible [2168]. Admission [1842]. Adoption [2444]. ADSL [2345]. Advanced [1396, 947, 1531, 1759]. Advances [2926]. Advection [1352]. Adventure [1073]. Advertising [1441]. Advice [3117]. Aerial [987, 1578]. Aerodynamic [1324]. AÉROSPATIALE [1057]. Aesthetics [534, 2129]. AFA [2645]. Affine [1608, 1607, 1556, 1560, 1572]. Affordable [917]. Aflatoxin [1849]. Aftermath [2605]. Against [2281, 2692, 2052, 2496, 1655, 1460]. Age [299]. Agencies [1762]. Agency [2767, 91]. Agenda [21, 709]. Agent [2664, 794, 1152, 1126, 2658, 1037, 2630, 2466, 226, 486, 2976, 1862, 187, 2663, 1236, 2612, 243, 2666, 303, 1861, 2057, 2623, 247, 2618, 1920, 2655, 239, 2665, 1168, 2939, 1863, 241, 2014, 2454, 224, 1127, 238, 2653, 350, 1971, 237, 2662, 212, 382, 1166, 244, 2609, 2587, 1447, 366, 229, 2585, 2765, 234, 2479, 2956, 1975, 227, 2668, 1457, 2775, 2021, 2959, 2846, 1133, 2780]. Agent-Based [794, 2658, 486, 1863, 2662, 1447, 366, 2956, 1975]. Agent-Oriented [1236]. Agent-Supported [2612]. Agents [2626, 2055, 2456, 2656, 1857, 1978, 2043, 2776, 2615, 1103, 1131, 2660, 2455, 2629, 248, 2955, 213, 240, 2667, 2652, 2628, 1976, 487, 488, 2470, 2054, 1960, 2616, 2897, 2610, 1038, 2053, 2773, 1155, 2774, 2608, 2050, 2584, 2052, 2617, 1162, 225, 230, 1400, 1459, 2621, 2627, 2056, 2619, 2961, 484, 2661, 2778, 1456]. Aggregate [268, 586, 345, 2605, 587]. Aggregation [2075, 347, 585]. Aglets [2058]. Agony [2727]. Agreeing [44]. agreement [2496]. AI [2031, 1856, 1124, 1972, 1810, 1964, 1887, 1858]. Aid [1920]. Aided [2494, 507]. AIN [791, 806, 807]. Air [1330, 498]. Airplane [991]. Alarm [1918]. ALBERT [724]. Alerts [1109]. Algebra [1831, 899, 2088, 111, 710, 139, 1792, 473, 545, 2175, 1536, 1964, 3126, 547, 552, 558]. Algebraic [2133, 544, 145, 143, 2196, 2081, 3141, 2188, 440, 471, 1228, 1798, 2801, 457, 140, 2138, 551, 559]. Algebras [644, 2264, 629]. Algorithm [2510, 2967, 853, 2022, 1870, 1076, 2156, 1437, 3085, 3037, 1749, 1730, 2785, 1868, 3165, 2273, 668, 1362, 1818, 2828, 689, 2907, 1337, 1906, 1746, 2514, 780, 681, 1849, 1745, 2196, 779, 3027, 2241, 1174, 1268, 2540, 1891, 575, 1360, 1113, 1114, 402, 2364, 2641, 3034, 592, 878, 2114, 2960, 1276, 1897, 2544, 2973, 3005, 1821, 1326, 2543, 1118, 3082, 1409, 1905, 422, 2256, 3015, 3020, 1546, 2123, 2157, 423, 951, 2177, 2024, 2263, 3063, 2775, 3039, 3036, 2506, 740, 2895]. Algorithmic [2799, 3051, 3101]. Algorithms [193, 665, 282, 2787, 2529, 3059, 3091, 1344, 2832, 2133, 2986, 1846, 2795, 176, 2219, 3058, 1734, 1845, 470, 3066, 900, 459, 1338, 1401, 1433, 3013, 3065, 2243, 1893, 184, 2991, 1575, 2236, 2782, 190, 2977, 2509, 2984, 1869, 2531, 292, 3077, 281, 2687, 2797, 426, 1732, 3054, 3004, 191, 2653, 2803, 2534, 1465, 3096, 3007, 1390, 2023, 2994, 902, 194, 2808, 1473, 188, 3099, 1733, 896, 196, 460, 2972, 2987, 2966, 581, 2539, 1096, 3134, 2980, 1744, 907, 635, 1069, 838, 969, 997, 2817, 3084, 3064, 2965, 2975, 2501, 1574, 127, 2880]. Alias [2904]. Aliases [698]. Alice [528]. ALife [2580]. Aligning [3049, 1754, 3047]. Alignment [535, 277, 1391, 3046, 1621]. Alignments [3047]. All-Optical [2847, 1011, 2891]. Allocation [1674, 2659, 229, 1345, 2891, 740, 739, 729]. allocator [737]. Allocentric [1506]. Allows [1395]. Almost [462, 3065, 666, 550]. Almost-uniform [3065]. along [2766]. Alphabet [508]. Alphabets [3037]. Alternating [751, 2141]. Alternation [430, 1801, 3152]. Alternative [583, 2920, 1959, 1705, 1706]. Always [1251]. Ambient [357]. Ambients [612]. Ambiguous [1265, 1227]. AMInet [314]. Among [539, 2514, 1285, 2776, 2774, 2778, 2781]. amortization [3054]. Analogs [3129].

Analogue [1867]. Analysing [2663]. Analysis [2380, 606, 353, 1838, 2064, 280, 2787, 515, 2018, 927, 1715, 1716, 2115, 1049, 685, 1076, 1674, 176, 775, 688, 2217, 54, 2430, 1105, 2377, 815, 996, 1081, 1814, 2782, 2573, 2087, 2080, 1412, 2923, 2076, 158, 2531, 180, 246, 2784, 2187, 2290, 426, 222, 861, 1764, 1391, 763, 769, 731, 953, 1835, 171, 172, 402, 2360, 1635, 388, 719, 194, 2224, 647, 259, 926, 2709, 718, 2253, 987, 255, 254, 863, 2585, 2250, 954, 988, 2157, 854, 2440, 2269, 266, 1190, 1639, 749, 726, 2252, 1770, 1627, 732, 562, 1634, 733, 1491, 2880, 2270]. Analytic [2169]. Analytical [1903]. Analyzers [767]. Analyzing [704, 728, 2401, 673, 52]. Anaphora [493]. ANFIS [1933]. Angle [2515]. Angry [334]. Angular [3055]. Animal [1863]. Animals [197]. Animated [2567, 2579, 484, 1563]. ANN [849]. Annealing [1678, 886, 1922]. Annex [1712, 1699]. Annotated [2770]. Annotation [2325]. Anomaly [1698]. Anonymity [1448, 330]. Anonymous [2680, 147]. ANSI [1467]. Answers [699]. Ant [2656, 2244, 3030]. Ant-Like [2656]. Antenna [2653]. Antennas [1438, 1326]. Anthropomorphic [382]. any [2705, 159]. Anymore [268]. AORTA [706]. AP1 [2298]. APE [1360]. APE/ [1360]. APE100 [1416]. APE100/ [1416]. Aperture [952, 3024]. API [268, 1708, 2458]. APIs [2326]. Appearance [1592, 1647, 1653, 614, 1594, 1593]. Applicability [1293, 2978]. Application [993, 1324, 1853, 2325, 1846, 1076, 441, 1218, 2898, 912, 2660, 2455, 2383, 9, 2243, 2087, 850, 2984, 326, 787, 1357, 1864, 1925, 1071, 1770, 1388, 179, 1525, 2323, 1242, 806, 807, 2653, 2834, 1184, 1835, 1911, 1358, 3019, 2276, 188, 2861, 2288, 1964, 2413, 2414, 1255, 1848, 2554, 1198, 2927, 1759, 2322, 2296, 2931, 140, 3056, 2134, 1554, 1524, 1585, 2504, 2151]. Applications [2021, 2664, 1931, 1319, 2298, 2314, 2315, 2672, 73, 1368, 663, 1396, 1398, 1313, 1458, 610, 766, 2473, 1383, 765, 1346, 1343, 187, 784, 1413, 650, 210, 597, 1768, 959, 851, 1701, 1322, 1702, 1755, 2690, 1791, 292, 1480, 178, 1179, 602, 950, 966, 824, 960, 1420, 95, 2472, 829, 310, 948, 1907, 873, 2468, 147, 463, 1318, 301, 1987, 136, 1858, 2668, 2537, 2339, 1194, 2522, 1462, 1669, 1671]. Applied [1307, 3175, 3016, 2969, 403, 1135, 2020, 2544, 1092, 1347, 1425, 2295, 881]. Apply [1886]. Applying [1883, 2550, 1957, 1819, 129, 65, 99]. Approach [2386, 1025, 108, 2471, 2672, 2387, 2015, 1368, 104, 276, 2136, 1601, 580, 1397, 486, 96, 892, 231, 94, 1775, 1683, 923, 1856, 1082, 1873, 584, 210, 341, 984, 1438, 1714, 2215, 1877, 2552, 1948, 503, 214, 2937, 2454, 2255, 2411, 1088, 1908, 989, 1769, 2145, 1824, 114, 1112, 827, 1384, 1064, 2235, 167, 2393, 3035, 1690, 876, 885, 1772, 2544, 1859, 2265, 2169, 1595, 2829, 864, 2801, 2599, 1887, 904, 1800, 2337, 251, 1096, 1949, 1834, 1453, 1889, 981, 2821, 2780, 1933, 2883, 2137, 1600, 549, 2138, 1656, 1563, 1663, 732, 2146, 1648]. Approaches [3072, 1989, 509, 1440, 1972, 831, 3051, 2119, 1444, 1750, 118]. Approximability [467, 1747, 284, 290, 289]. Approximate [3103, 2297, 676, 679, 2546, 677, 2196, 242, 120, 2213, 3034, 3040, 680, 2944, 2276, 3038, 682, 3041]. Approximating [3058, 2885, 3146, 1732, 2890, 2891, 2892, 288, 2893, 291, 2883]. Approximation [2510, 2882, 2884, 450, 2832, 452, 2223, 1749, 1730, 1734, 2857, 900, 816, 1746, 292, 2512, 281, 2971, 1822, 3062, 2225, 294, 1748, 2856, 1733, 445, 2247, 3084, 2248, 2232, 2889, 2880, 2895]. Approximations [2303, 2513, 2264, 634, 2246, 2879, 2291, 2887]. AR1 [2297]. Arabic [1877, 509, 520]. Arbitrary [1220, 1221, 873, 718, 1020, 1630]. Arbitrated [1191]. Arbitration [2392, 1970]. Arc [2230, 2891]. Archetypes [934]. Architecting [2439]. Architectural [975, 2433, 2714, 2424, 976, 25, 2415, 356, 2911, 2463, 357]. Architecture [860, 307, 2438, 2427, 365, 1285, 398, 532, 533, 226, 1371, 1862, 2423, 2432, 2412, 755, 1912, 1861, 2655, 882, 2461, 967, 1168, 1758, 1547, 2426, 332, 2605, 1377, 2451, 1331, 2662, 1158, 1294, 1541, 2416, 2413, 2340, 1929, 2414, 2320, 53, 944, 2326, 102, 867, 846, 2479, 1366, 93, 1932, 317, 2855, 2419, 227, 2668, 1446]. Architecture-Centric [2438]. Architectures [704, 1344, 1335, 2431, 2430, 2436, 1713, 16, 2418, 868, 306, 1349, 1971, 905, 856, 2407, 1975, 258, 836, 2437, 1364, 2, 2420, 742]. Arctic [1994]. Area [141, 254, 1602]. Areas [1314]. Arena [1526]. AReVi [2587, 2588]. Argument [2865]. Argumentation [236, 128]. Argumentation-Based [236]. Arguments [2865]. Ariadne [1044]. Arising [3052]. Arithmetic [2812, 2377, 3171, 614, 2175, 1803, 2501]. Arithmetically [2185, 2184]. Arithmetics [2201]. Arms [1907]. Arrangements [2536]. Array [853, 277, 1391, 1326, 266]. Arrays [671, 3174, 2970, 1438, 3043, 2040, 933, 759]. Art [2603, 2602, 988]. Articulated [1850]. Artifacts [68, 62, 70, 63]. Artificial [3000, 1936, 1828, 205, 206, 1886, 880, 3030, 2584, 2579, 1982, 2269, 2270]. Artistic [499, 994]. Asia [300]. Asia-Pacific [300]. ASIS [1709]. Aspect [1521, 97]. Aspect-Oriented [97]. Aspects [1042, 1200, 1291, 1201, 1182, 349, 1969, 440, 212, 557]. Assembly [536, 1774, 491, 1817]. Assertions [1816]. Assessing [360]. Assessment [2434, 1980]. Assigning [1834]. Assignment [199, 1676, 427, 1751, 256, 1900, 888]. Assignment-Type [1676]. Assimilation [1327]. Assist [1538, 2973]. Assistant [1814]. Assistants [2020]. Assisted [207, 2328]. Association [1076, 1077, 1079, 1084, 2207, 2237, 1089, 1111, 992, 1534, 1094, 1122, 2206, 1615]. Associations [1082, 1661]. Associative [3017, 640]. Associative-Commutative [640]. Assume [2391]. Assumption [1237]. assumptions [1485]. Assuring [1808]. ASTL [2638]. Astro [1705, 1706]. Astronomical [1115]. Asymetric [1276]. Asymmetric [607, 157, 1023, 1470, 3022, 1024]. Asymptotic [1743]. Asymptotical [196]. Asymptotically [3105]. Asynchronous [1350, 609, 2634, 2872, 2144, 2815]. Asynchrony [2873, 621]. ATM [2450, 2456, 1825, 805, 2080, 2655, 2458, 2478, 2066, 1842, 1995, 2341, 2657, 2338, 2479, 2339]. ATMAN [2343]. Atmospheric [1327]. ATMS [2770]. Atomic [1469, 3002]. Atoms [3104]. ATR [206]. ATR’s [205]. Attack [153, 406, 1170, 2683, 1171, 1460]. Attacking [420]. Attacks [156, 2680, 2692, 2681, 664, 416, 152, 2496]. Attempt [1886]. Attention [382, 245]. Attitude [2276]. Attractors [438]. Attribute [2925, 2035, 32, 1119, 728, 729]. Attribute-Value [2925]. Attributes [985, 2283, 1084, 330, 2262, 2275]. Audio [2343, 222, 1658, 1323]. Audio-Visual [1658, 1323]. Augmentation [1728]. Augmented [1185, 322]. Augmenting [1051, 1052]. Authenticated [157, 2496]. Authentication [2055, 157, 158, 1191, 1172, 1499, 1491]. Author [296, 1235, 1505, 1551, 1614, 1672, 2083, 2153, 2205, 2445, 2508, 2565, 2762, 2896, 2924]. Authoring [505]. Auto [1461]. Auto-Certifiable [1461]. Auto-Recoverable [1461]. Autoassociative [1953]. Autocalibration [1557]. autoepistemic [134]. AutoLink [170]. AutoMap [170]. Automata [768, 1175, 651, 608, 2633, 2812, 916, 2634, 2841, 3175, 1784, 2377, 1240, 3158, 2840, 204, 757, 429, 2853, 821, 2548, 2637, 2638, 191, 2639, 473, 2834, 845, 438, 3176, 696, 829, 3177, 437, 439, 3035, 454, 1900, 656, 3117, 896, 2644, 457, 463, 883, 1436, 725, 2854, 2649, 2650, 2651, 2547, 726]. Automated [3023, 2998, 780, 1755, 135, 957, 1891, 2120, 718, 2617, 2117, 2118, 1234, 642, 142]. Automatic [73, 535, 931, 802, 501, 1027, 2065, 253, 982, 983, 1571, 984, 502, 389, 277, 1087, 279, 1640, 715, 1835, 1599, 1578, 625, 1958, 1241, 767, 352, 3024, 2562, 3032, 2330, 2018, 1597]. Automatically [1808, 1018, 2381, 1936, 933]. Automating [1696, 2342, 1761]. Automation [789, 827]. automatique [541]. Automaton [2635, 2561, 2549, 1793, 846, 2647, 2503]. Automorphism [442]. Automorphisms [3142]. Automotive [810, 811, 816]. Autonomic [329]. Autonomous [218, 1924, 1525, 2318, 202, 203, 2773, 2774, 1927, 1146, 230, 1149, 1524]. Autotag [529]. Auxiliary [3074]. Availability [1384]. Available [914]. AVANTI [2457]. Avatars [2607]. Avenge [3036]. Average [2278, 3138, 3139, 2541]. Average-Case [3138, 3139]. averaging [1584]. Avionics [1712, 1057]. Avoid [1698, 1273]. Avoidance [2987]. Aware [920, 874]. Awareness [219]. Axiomatisation [2072]. Axiomatization [2878]. Axiomatizations [604]. Axiomatizing [464]. Axioms [2825, 1792].

B [1054, 1061, 1058, 1693, 1694, 1060, 1431, 1071, 1064, 1066, 1059, 1068, 1067, 1069]. B-ISDN [1431]. B-Method [1059, 1069]. Back [302]. Backbone [1446]. Background [996, 2933, 1530]. backgrounds [1644]. Backing [2995]. Backjumping [1224]. Backoff [171, 172]. Backtracking [715]. Backtracking-Free [715]. backwards [1479]. Bahncard [3060]. Balance [2578]. Balanced [1437, 1495, 474, 1496]. Balancing [2789, 1344, 1402, 1343]. BAN [2389]. Bandwidth [1380, 1382, 2852, 1431, 1426, 2335, 2891].

Bar [981]. Bar-Charts [981]. baroque [521]. Barotropic [1328]. Barrier [174, 2599]. Barriers [852]. BARTER [1446]. Base [1538, 80, 2912, 1945]. Based [411, 970, 794, 1931, 386, 1777, 73, 1164, 2438, 1488, 758, 2079, 35, 1966, 2658, 398, 2361, 383, 1875, 610, 706, 1782, 3000, 2043, 2002, 775, 486, 569, 1959, 231, 1131, 1851, 94, 1968, 526, 2331, 1683, 2171, 2624, 1878, 1079, 2047, 1082, 2673, 817, 584, 1814, 341, 404, 984, 645, 2089, 1196, 1136, 1912, 1454, 204, 2659, 2375, 8, 614, 1890, 2945, 502, 1745, 2029, 2173, 385, 325, 691, 67, 753, 1482, 1610, 1156, 1637, 1863, 1773, 1948, 214, 2446, 2008, 2951, 1246, 1268, 1440, 173, 2078, 191, 2459, 114, 1891, 976, 2449, 401, 1064]. Based [1930, 2235, 1980, 953, 308, 1915, 510, 1971, 848, 2670, 2671, 1604, 169, 2994, 3034, 2963, 2770, 1143, 1914, 831, 1916, 1917, 1271, 1195, 2662, 876, 2572, 1942, 2492, 181, 2341, 1821, 2265, 717, 1767, 320, 2576, 1445, 1347, 990, 621, 2570, 2696, 1447, 2682, 1679, 1839, 911, 2698, 1197, 1923, 1369, 331, 1162, 366, 1267, 102, 506, 1937, 1096, 314, 236, 1992, 2006, 1984, 683, 149, 1605, 2469, 2676, 1147, 1949, 211, 2956, 1292, 1710, 1608, 1975, 837, 2274, 1834, 1932, 1750, 400, 1121, 227, 1141, 1453, 1944, 2172, 994, 981, 266, 2781, 1102, 2282, 2232]. Based [1364, 1607, 1650, 1476, 1609, 1856, 39, 557, 1606, 1481, 1636, 127, 1657, 128, 1583, 1651, 1586, 2507, 1321, 1649, 2503, 733, 1501, 1564, 2504, 1485, 909, 1648, 1662, 326, 2461, 1198, 1928, 1613, 1425, 2479, 2495, 262]. Bases [1967, 659, 1813, 135, 2207, 1038, 2266, 2206]. Basic [2107, 1787, 1847, 730, 198, 1685, 1686]. Basic-block [730]. Basins [1963]. Basis [2246, 1816, 2200, 2813, 1633]. Basket [1076, 1086]. Batch [663, 1262, 1477]. Batcher [460]. Battle [3114]. Bayanihan [331]. Bayes [1237, 1246, 1830]. Bayesian [842, 1247, 1666, 1092, 2295, 1584, 2042, 2779, 1261]. BDD [2375]. BDDs [1211, 957, 3155]. Be [1808, 3144, 1464, 1485]. Beget [2428]. Beginning [1606]. Behavior [1150, 231, 1970, 3010, 1162, 4, 1858, 1512]. Behavior-Based [231]. Behavioral [1529, 484, 1919]. Behaviored [2213]. Behaviors [2634, 1135, 1143, 1155, 3028, 1927]. Behaviour [798, 1925, 1863, 2835, 2836, 3030, 719, 2570]. Behaviours [2581]. Behind [852]. Belief [127, 2255, 3007, 2260, 129]. Beliefs [1857]. Bellantoni [1791]. Bellantoni-Cook [1791]. Benchmark [590, 1210]. Benchmarks [1353]. Benefit [863]. Benefits [2242]. Bennion [2164]. Benzenoid [1736]. Best [2344, 2895, 2713]. Best-Effort [2344]. Beta [516, 19]. Beta-Velocity [516]. Better [2003, 3062, 1484]. Between [1826, 2429, 1106, 2807, 3166, 1351, 2844, 439, 2590, 2122, 2952, 1840, 2102, 1750, 1512, 2011, 3039, 1517, 2204, 490, 813, 3013, 1272, 2771, 559, 2768, 560, 2277, 2773, 2482, 698, 2176, 2017, 357, 2779]. Betweenness [1737]. Beyond [754, 2148]. Bi [1335, 3159, 3160]. Bi-Infinite [3159, 3160]. Bi-Parallel [1335]. Bias [1645]. Bias-Variance [1645]. Bibliographic [1038]. Bibliography [295, 1297]. bibliothèque [521]. BiBTEX [2764]. Bidding [229]. Bidirected [1726]. Big [614]. Bijective [2686]. Bilingual [2029]. Billing [2485]. Billion [205, 206]. Bin [1740, 2512]. Binary [2523, 648, 3065, 448, 2127, 201, 2910, 3161, 2261, 2157]. Binder [326]. Binding [685, 154, 1696]. Binding-Time [685]. Biological [602]. Biologically [3003]. Biology [2871, 1737, 2927, 2537]. Biomass [1880]. Biometallurgical [1835]. Biosequences [2564, 2563]. BIP [908]. Bipartite [1723, 1722, 1497]. Biquadrates [2166]. Bisimilarity [2806, 559]. Bisimulation [684, 3125, 2816, 2859, 3100]. Bisimulation-Like [2816]. Bisimulations [822]. Bit [583, 776, 777, 154, 2350, 3034, 3035, 435, 2141, 2500]. Bit-Parallel [3035]. Bit-Vector [3034]. Bit-Vectors [776, 777]. bitos [1159]. Bitrate [2348]. Bits [1969]. Black [1475, 413]. Black-Box [1475, 413]. Blackboard [1871]. Blackbox [2054]. Blind [1195, 1177, 844, 1489, 1181]. Bloat [1016, 1021]. Block [2513, 419, 2677, 2331, 321, 2966, 1340, 449, 2177, 2675, 1186, 1479, 730, 1405]. Block-Reduction [2177]. Block-World [321]. Blocking [2076, 2012, 3075]. Blockwise [3155]. Bluff [1434]. Bluff-Body-Stabilized [1434]. BMS [1685]. Board [491, 1914]. Body [1344, 389, 1434, 3063]. Bolivia [1150]. Boltzmann [1337, 1333]. Books [146, 1049, 1047]. Boolean [2869, 3171, 1495, 3053, 429, 1496, 3068, 467, 2842, 3178]. Boosting [1260]. Both [223]. Bottom [3177]. Bound [3127, 1729, 2998]. Boundaries [2224, 1602]. Boundary [3067]. Bounded [3135, 671, 441, 1731, 1218, 470, 3069, 2822, 1393, 479]. Bounding [2160]. Bounds [1738, 1781, 892, 1740, 3172, 1498, 1465, 1748, 2492, 2520, 2684, 436, 288, 3075, 2886]. Box [1475, 2072, 413]. Boxes [2686]. BP [1282, 1951]. BP-SOM [1282, 1951]. Bracketer [2028]. Brain [205, 206, 205]. Branch [276]. Branching [2822, 436]. Breaking [155, 1464, 2507]. Bremen [1524, 1525]. Bricks [20]. Bridge [698, 559]. Bridges [2837]. Bridging [2140]. bring [738]. Bringing [1047]. British [1689]. Broadband [809, 2446, 2333, 2477, 2339]. Broadcast [1498, 1428, 2321]. Broadcasting [922, 187, 462, 173]. Broker [307, 1452, 1039]. Broker-Centered [1452]. Brokerage [2471, 1450, 2472, 2469]. Browser [2336]. Browsers [2061, 2062]. Browsing [678, 1948, 2465]. Brute [1690]. BSP [872, 887, 929]. Buchberger [2123]. budget [2888]. Buffer [584, 329, 895].

Buffering [1273]. Buffers [671, 2066]. Build [1931, 150, 950, 1588, 404]. BUILD-IT [404]. Builder [1430]. Building [1313, 1161, 248, 1413, 2667, 361, 1743, 1714, 1773, 369, 1030, 1929, 698, 2966, 929, 3008]. Buildings [361, 367, 373, 377, 370, 368, 356]. Bulk [582, 2842, 3178]. Bulk-Synchronous [2842, 3178]. Bundle [1752]. Burns [779]. Burnside [654, 653]. Bus [2392, 960, 2409, 2410]. Buses [3077]. Business [978, 1754, 42, 323, 1088, 350, 1772, 1294, 2469, 2240, 1770]. Byte [1186]. Byte-Oriented [1186].

C [3106, 736, 2425, 102]. C/C [2425]. C1 [2299]. C4.5 [1101]. CA [3177]. Cable [2345]. Cache [1437, 1757, 2354, 935]. Cached [2907]. Caching [1435]. Cacti [1727]. Cactus [1743]. CAD [490]. CAE [2286]. CAESAR [755]. Calculating [723, 2152]. Calculations [1423]. Calculi [1780, 2098, 2872, 2873, 2096]. Calculus [1213, 1216, 2079, 36, 3121, 2918, 74, 75, 1843, 145, 2072, 1791, 241, 2838, 1227, 1795, 2096, 721, 1230, 929, 761, 2839, 140, 430, 2140, 2859, 2858]. calibration [1579, 1554, 1561]. Call [1776]. Call-by-Value [1776]. Called [1896]. Calligraphic [510]. ’Calls [783]. Calm [354]. CAM [206, 205]. CAM-Brain [206, 205]. CAMD [1322]. Camera [1604, 1554, 1571, 1561]. Camera-Based [1604]. Cameras [1565, 1572]. Can [3144, 2003, 1696, 1140, 2585, 1485, 1935, 2568, 2428, 2608, 960]. Cancer [2295]. Canonical [261, 2809, 2127]. Canonization [1788]. Capabilities [2666, 2619]. Capability [2858]. Capacity [1900]. Capture [1766, 1765]. Captures [433]. Capturing [891]. Car [1324, 2355]. Card [2092, 1071, 164, 1184, 2092]. Cardinal [1832]. Cardinality [3085, 2895]. Cards [2505]. Carlo [1335, 1414, 1818]. Carrying [2053]. Cascade [1950]. Cascades [2681].

Case [1075, 810, 271, 811, 2518, 707, 1302, 708, 1078, 1856, 2424, 1713, 1293, 2391, 2531, 1773, 1948, 274, 3138, 3139, 1184, 370, 1281, 1115, 719, 1772, 2482, 700, 1945, 1949, 874, 2992, 1072, 1944, 362, 2781, 1569]. Case- [1856]. Case-Base [1945]. Case-Based [1773, 1948, 1944, 2781]. Cases [3048, 498, 432, 1806]. Cash [1502, 1195, 2492]. CASL [562]. CAST [406]. Categorical [1084]. Categories [617, 1798, 567]. Categorization [1254, 1516]. Catenable [2519]. Causal [94, 306]. caused [1629]. Causes [1016]. CAVE [2599]. CBCM [1467]. CBR [1310, 1307, 1313, 1308, 1309, 1302, 1946, 1312, 1305, 1947, 1311, 2287, 1304]. CBT [313]. cc [814]. CCAIIA [1084]. CCD [1590]. CCS [957]. CCSIMD [1993]. CD [1046]. CD-ROM [1046]. CEC [2608]. Cell [475, 2536]. Cells [2537]. Cellular [1025, 916, 3174, 3175, 2840, 476, 204, 2464, 427, 2997, 2335, 438, 3176, 3177, 437, 883, 2228, 2503]. Cellular-Programming [1025]. Census [3137]. Center [2517, 827]. Center-TRACON [827]. Centered [1452, 2598, 308]. Centers [2443]. Central [1689, 144]. Centralised [1912]. Centrality [2211]. Centre [1317]. Centred [1853]. Centres [1320]. Centric [2438, 1762]. Centrifugal [1333]. Century [1321, 2134]. Certain [2212]. Certifiable [1461]. Certificate [2695]. Certification [2078, 2484]. Certified [2123]. Certifying [2493]. CFMD [1113]. Chain [1743, 1480]. Chains [2078, 673, 2174]. Chaitin [477]. Challenge [1126, 1896, 2616, 1124, 1127, 81]. Challenges [738, 70, 217, 2000, 988, 378]. Challenging [2461]. Chambers [2470, 1333]. Change [2722, 2710, 1764, 700, 2330, 129]. Changes [1629, 3049, 1971]. Changing [1279, 835, 118, 123, 129]. Channel [1848]. Channels [2380, 965, 2350, 2007, 1199]. Chaos [1423, 3175]. ChaosLAN [2010].

Chaotic [2835, 2836, 2010]. Character [2293, 850, 504, 980, 2296, 2179]. Characteristic [135, 1114, 2204]. Characteristics [1133]. Characterization [2635, 1722, 669, 178, 179, 3141, 1268, 3070, 3176, 465, 1796, 1558, 2254, 127]. Characterized [3006, 1086]. Characterizing [2034, 1724]. Characters [971, 973, 979]. Charging [2450]. Charts [618, 2402, 981]. Chase [1542]. Chebyshev [1628]. Checked [2385]. Checker [761, 748]. Checkers [2359, 3097]. Checking [768, 754, 2400, 2370, 663, 2361, 2405, 1203, 2367, 771, 2404, 3130, 2068, 2846, 2371, 2359, 284, 2394, 752, 2859, 2368, 2364, 3096, 716, 2393, 756, 2372, 2373]. Checkpointing [919, 926]. Chemical [3066, 1358, 1821]. Chemistry [2927]. Chess [1956]. Children [1245]. Chinese [1956]. Chip [648, 2996]. Choice [2216, 2622]. Chosen [153, 1460]. Chromatic [666]. Chrominance [2348]. Chromodynamics [1334]. Chronogram [2804]. Chunking [1383]. Chunks [1763]. Church [619, 624, 623]. CIIPS [1527]. Cipher [419, 410, 406, 2682, 1170, 2675, 1186, 417]. Ciphers [2677, 1480, 449, 413, 1479]. ciphertext [1460]. Circuit [3127, 889, 3129, 1071, 868, 667, 491]. Circuits [3128, 2869, 3153, 3171, 2818, 1185, 2279, 2280, 1879]. Circulant [2860, 3079]. circular [2891]. Circumscription [1219]. Cities [367]. Civic [1689]. CKBS [2234]. CL [1212, 2102]. Clairvoyance [2532]. Clam} [2101]. CLARET [1118]. Class [1711, 2189, 1742, 2067, 2190, 111, 2192, 790, 631, 2193, 2834, 80, 2912, 1748, 2201, 833, 2254, 1990, 1757]. Classes [2958, 3088, 444, 3055, 38, 1740, 461, 1468, 690, 3069, 478, 697]. Classical [2958, 1787]. Classification [1025, 1954, 1875, 2935, 2036, 1861, 2215, 2212, 1877, 1806, 1268, 2249, 1257, 2994, 1092, 3024, 1121, 1256]. Classifications [2046, 1262, 1260].

Classifier [1278, 1830, 1261, 1258]. Classifiers [1259, 2238, 1257, 2562]. Classify [2624, 2997, 1243]. Classifying [2922, 381, 1119]. Clause [2126]. Clauses [2944, 2948]. Claw [662, 1726]. Claw-Free [662, 1726]. CLear [364]. ClearCase [2708]. Client [1368, 223]. Client-Server [223]. Client/Server [1368, 221]. Clients [1697, 2678]. Clifford [140, 2088, 139]. Clique [3127, 670, 2892]. Clock [958]. Clocks [1701, 1702]. CLOG [2938]. Clos [3074, 3075]. Close [2011]. Closed [1579, 461, 456, 1571]. Closed-form [1579]. Closure [3119, 1831, 443]. Cloud [1111]. Clouds [1926]. Clueless [2050]. Cluster [912, 1375, 181, 909, 1373]. Clustered [668]. Clustering [1356, 2561, 2035, 1084, 1991, 2951, 1882, 2037, 1273]. Clusters [3002, 873, 2037, 915]. cluttered [1644]. CMIP [2481]. CML [719]. CMM [48]. CMUnited [1153, 1142]. CMUnited-97 [1153, 1142]. CNM [2479]. Co [3031, 209, 611, 638, 1435, 367, 961, 862, 380, 1154, 2023, 2774, 2224, 1803]. Co-continuous [611]. Co-definite [638]. Co-evolution [2023]. Co-evolving [3031, 1154]. Co-Heyting [2224]. Co-operative [1435, 367]. Co-ordination [209, 2774]. Co-processed [961]. Co-processing [862]. Co-speech [380]. co-words [1803]. Coal [1949]. Coalgebra [552]. coalgebraic [553]. Coalition [2617, 234]. Coarse [3059, 972]. Coarse-Grain [3059]. Coastal [1329]. Code [1334, 2049, 2236, 1412, 693, 2425, 264, 2053, 252, 1352, 2583, 1717, 741, 2684]. Coders [380, 2253, 2252]. Codes [1169, 265, 1349, 1895, 1172, 1269, 1499, 2676]. CoDi [204]. CoDi-1Bit [204]. Coding [2348, 2559, 1510, 3045, 2347]. CoEd [2724]. Coevolutionary [1274]. Cognition [1520, 1529]. Cognitive [2039, 2955, 363, 1844, 2855]. Coherence [2354, 935]. Cohort [3019]. Cohort-Driven [3019]. coinductive [2151]. Colimits [561]. Collaboration [207, 215, 363, 324]. Collaborations [222]. Collaborative [1962, 218, 2624, 2623, 321, 149, 362]. Collapses [3117]. Collection [2907, 2611, 2908, 1428, 1546]. Collections [1045, 1958, 529]. Collective [881]. Collinear [1099]. collisions [1485]. Colonoscopy [3083]. Color [2578, 1587, 1590, 1598]. Colorable [3080]. Colorado [511]. Colored [476, 3057, 2514]. Coloring [186, 669, 2794, 737]. Colour [985, 1582, 1580, 1581, 1591]. Coloured [2069, 1648]. Colouring [666, 2891]. Combination [2087, 1123, 1609, 134]. Combinatorial [3102, 3088, 1746, 3052, 3073, 2819, 1498, 132, 1750]. Combinatorics [1736, 1751]. Combinators [644]. Combinatory [644]. Combined [1961, 242, 3004]. Combiner [1171]. Combining [996, 1259, 757, 1664, 1661, 57, 563, 2112, 1128, 2930, 1684, 556]. Come [1502]. Comment [2705]. Commerce [307, 1442, 1443, 1458, 1449, 1450, 2618, 2470, 1455, 2617, 1459, 2469, 2619, 1453]. Commitment [2680, 2773]. Commitments [1549]. Committees [2761, 1261]. Common [2021, 1978, 1374, 1806, 1552, 682]. Common-Pool [1978]. Commonality [2440]. CommonKADS [246]. Communicating [1857, 170]. Communication [676, 910, 2344, 498, 480, 2002, 267, 913, 1959, 3171, 459, 1394, 1041, 1486, 2680, 1993, 1714, 1979, 677, 691, 3172, 178, 2616, 179, 861, 2121, 238, 2566, 1865, 957, 905, 165, 244, 2001, 2856, 2801, 911, 1197, 1926, 1001, 93, 951, 266, 2011]. Communication-Based [691]. Communication-Efficient [676, 459, 677]. Communications [2700, 1427, 941]. Communicative [2765]. Communities [2594]. Commutative [640, 1799]. Compact [2809, 313, 1008]. Company [993]. Comparability [451]. Comparative [178]. Comparator [2523]. Compared [1018]. Comparing [2559, 1973, 1349, 2201]. Comparison [1204, 2195, 3166, 3013, 1678, 2680, 201, 233, 1351, 890, 870, 3045, 673, 2954, 2378, 954, 1352, 1649, 1644]. Comparisons [2932]. Compartmented [2703]. COMPaS [909]. COMPASSION [1383]. Compatibility [2265]. Compatible [2845]. Competence [1316]. Competition [2540, 1158, 245]. Competitions [1685, 1686]. Competitive [2787, 2796, 2782, 2798, 2618, 2784, 2797, 2778]. Compilation [3092, 1397, 259, 262, 742]. Compiler [1398, 273, 267, 311, 278, 265, 1440, 264, 1710, 566]. Compiler-Based [1440]. Compiler-Generated [311, 264]. Compiler-Parallelized [265]. Compiling [251]. Complement [3113]. complet [541]. Complete [271, 2958, 2866, 1576, 3124, 312, 2838, 466, 699, 2867, 2157, 1684]. Completeness [3135, 3136, 3104, 1222, 1223, 471, 3109, 1289, 553, 470]. Completing [2949]. Completion [2813, 1602]. Complex [1150, 569, 2183, 658, 170, 950, 1974, 848, 2589, 252, 1887, 585, 2296, 2843, 2287, 124, 1671]. Complexes [2289]. Complexities [96, 51]. Complexity [2871, 2299, 442, 3170, 444, 687, 432, 2036, 3166, 1219, 446, 674, 2923, 841, 3129, 2853, 3172, 281, 478, 1844, 1797, 2801, 1817, 1950, 675, 2507, 2811]. Complexity-Theoretic [3129]. Complying [2324]. Component [73, 72, 532, 533, 74, 75, 76, 79, 77, 2910, 304, 1547, 80, 2421, 2617, 82, 2422]. Component-Based [73]. Component-Oriented [72, 79, 82]. Components [2447, 2658, 85, 1703, 1934, 709, 1890, 78, 2475, 84, 1834, 1619]. Composing [1279]. Composite [1173, 1189]. Composition [812, 938, 2476, 710, 1059, 507, 3156, 83, 86, 87]. Compositional [1855, 351, 554, 2071, 712, 2364]. Compositionality [763]. Compositionally [2381]. Comprehensive [1582, 1753]. Compression [1383, 470, 434, 787]. Computation [1026, 676, 853, 3090, 1488, 2219, 2691, 2189, 2140, 884, 1993, 1553, 677, 2193, 3095, 1895, 3019, 893, 1031, 1197, 2202, 189, 3008, 2271, 2855, 2648, 2537, 175, 1639]. Computational [1778, 3121, 1737, 139, 2249, 1797, 137, 2956]. Computationally [1342, 3097]. Computationally-Sound [3097]. Computations [2301, 2302, 2529, 2154, 3172, 881, 3152, 2286]. Compute [3088, 952]. Computed [1575, 1574]. Computer [1883, 1035, 480, 912, 1959, 1957, 213, 1044, 7, 395, 1212, 1930, 1526, 872, 513, 1347, 358, 1964, 1369, 507, 722, 1352, 140, 1405]. Computer-Aided [507]. Computer-Generated [513]. Computerized [1200, 1201]. Computers [1359, 1409, 302, 854]. Computing [3127, 2310, 49, 1324, 2381, 2959, 2968, 2848, 2217, 1851, 3055, 2840, 858, 784, 2195, 2186, 877, 1272, 2514, 312, 3115, 1376, 2357, 2674, 2376, 1973, 2897, 2147, 1844, 1547, 1395, 1972, 1377, 1914, 864, 309, 1447, 331, 863, 1318, 941, 2203, 874, 2198, 1472, 1372, 3063]. Concatenable [2824]. Concatenation [2837]. Concentration [325]. Concept [1274, 2336, 1265, 1105, 1814, 2943, 1272, 2026, 785, 786, 2355, 46, 47, 119]. Conception [1037, 1914]. Concepts [1299, 1302, 1515, 1847, 1298, 369, 714, 1816, 1017, 2045, 2276, 123]. Conceptual [1105, 319, 597]. Concerning [2167, 2179, 1584]. Conclusions [1053]. Concrete [687]. Concurrency [1401, 1433]. Concurrent [2849, 2387, 1809, 758, 686, 2922, 2526, 3132, 884, 1993, 1789, 241, 714, 256, 2147, 2915, 1795, 3133, 2417, 1799, 1400, 683, 258, 2839]. CONDENSATION [1613, 1603]. CONDENSATION-based [1613]. Condensed [518]. Condition [606, 1917, 833, 1123]. Conditional [2254]. conditioning [132]. Conditions [193, 966, 826, 719]. cones [1596]. Conference [1691, 1046, 2698]. Conferences [2762, 2764, 2763, 2761]. Confidence [1130]. Configurable [104, 166, 864]. Configuration [2714, 2727, 2460, 2655, 2717, 950, 952, 78, 1938, 2657, 1909, 2716, 1306, 2713, 103]. Configurations [144, 1556, 2811]. Configuring [95]. Confirmer [1490]. Conflict [1113, 986]. Conflict-Free [1113]. Conflicts [2512, 44, 2778]. Confluence [2097]. Confluently [2519]. Conformance [791, 2475]. Conformations [3002]. Congested [2851]. Congestion [1991, 2449]. Congruence [2806, 2859]. Congruences [2838]. conics [1572]. Conjecture [2167, 2810, 3079, 3080]. Conjunctive [1796]. Connected [1729, 1730]. Connecting [372]. Connection [2678, 1222, 1223, 1793, 2125, 898]. Connectionist [1955]. Connections [3003, 1426, 147, 2477]. Connectivity [2857, 1486, 891, 1728, 2017]. CONS [2823]. CONS-Free [2823]. Consecutive [1737]. Consensus [2955, 1298, 960]. Conservative [353, 870]. Consideration [316]. Considerations [1530]. Considered [783]. Considering [1539, 2558]. Consistency [632, 94, 2068]. Consistent [1533]. Consisting [358]. Constancy [1580]. Constant [433, 3088, 469, 256]. Constant-Round [469]. Constants [3128]. Constrained [338, 2530, 2517, 1748, 2978]. Constraint [975, 526, 1683, 3132, 2936, 689, 336, 2853, 1550, 1226, 3133, 115, 1799, 1848, 126, 683, 2040, 1834, 1684, 733, 117, 2895]. Constraint-Based [526, 1683, 1834]. Constraints [1061, 108, 2108, 2109, 814, 638, 1539, 3055, 2030, 2814, 1752, 1357, 2394, 279, 2349, 3150, 1395, 637, 2099, 2477, 1507, 1724, 506, 2839, 2886, 2888, 125, 1626].

Construct [257]. Constructed [1083]. Constructible [1807]. Constructing [1103, 1105, 3043, 2199, 3082, 1950, 3084]. Construction [2632, 2523, 1382, 900, 476, 275, 2704, 1085, 3105, 1294, 1138, 1949, 1685, 1686, 2139]. Constructions [415, 1490, 2684]. Constructive [1259, 2862, 3124, 981]. Constructs [339, 714]. Consumer [2421, 2422]. contact [402]. Contactless [2505]. Containing [1119, 3080]. Contamination [1849]. Contemporary [2000]. Content [2264, 2624, 2342, 1323, 1446, 1589]. Content-Based [2624]. Contents [2085, 2343, 2017, 599]. Contest [999]. Context [3153, 2560, 2641, 388, 619, 1859, 565, 1763, 1929, 2253, 863, 627, 3126, 1858, 3151, 2252]. Context-Free [3153, 2560, 2641, 3126]. Context-Mediated [1858]. Context-Sensitive [619, 1859]. Contexts [705]. Contextual [607, 1857, 2220, 2825, 1015, 2221, 1242, 1093]. Contiguous [3044]. Contingency [2828]. Continued [2180, 2181]. continuity [3072, 1586]. Continuous [583, 916, 2324, 1381, 1275, 818, 820, 241, 1658, 830, 1281, 3019, 2262, 317, 2275, 611]. Continuous-Valued [820]. Contour [983, 1586, 2347, 1609]. Contour-Coding [2347]. contours [1665]. Contracts [76, 77, 86, 87]. Contribution [1066]. Contributors [298]. Control [2712, 2306, 2695, 2694, 1164, 810, 271, 811, 1846, 1270, 498, 1150, 813, 1250, 2968, 1913, 816, 3167, 1720, 2412, 1912, 1886, 329, 2026, 1991, 2376, 1969, 1426, 2131, 1903, 2411, 78, 2838, 2449, 2711, 2120, 1842, 2038, 264, 1135, 848, 1914, 2662, 1911, 1985, 2276, 1066, 718, 2696, 1839, 1922, 1937, 1988, 2006, 1909, 1130, 2697, 93, 2231, 177, 1167, 1933, 749, 732, 2132]. control-flow [749]. Controllability [833]. Controller [817, 769, 1424]. Controllers [2995, 817, 3016, 720, 1927, 832]. Controlling [1162, 2338]. Controls [884]. Convenient [2981, 2982]. Conventional [2694]. Convergence [193, 784, 3009, 1283, 1281, 196, 2250, 2149, 2150]. Convergent [2044, 625]. Conversation [212, 483]. Conversations [350]. Converse [869]. Conversion [999, 2847]. Conversions [1048]. Convert [159]. Converters [3170]. Convex [1436]. convexity [1629]. Cook [604, 1791]. Cool [749]. Cooperating [1149]. Cooperation [490, 498, 482, 2667, 2876, 2091, 1549, 2766]. Cooperative [349, 210, 361, 2666, 2618, 487, 488, 377, 481, 370, 1155, 483, 368, 364, 358, 356, 134, 227, 2961, 2781]. Coordinate [2982]. Coordinated [876, 1907, 2723]. Coordinating [1131, 213, 232, 2621]. Coordination [2776, 231, 1401, 1433, 214, 217, 221, 1154, 350, 208, 234, 230, 211, 2627]. Cope [2585]. Copyright [1053]. Coq [761]. CORBA [2480, 2328, 95, 2481, 2326, 2475]. CORBA-Assisted [2328]. COREMO [10]. Cormas [1978]. Corner [1515]. Corporation [2444, 2357]. Corpus [387]. Correct [3092, 2139, 2491]. Correcting [1257, 1269]. Correction [2578, 2813, 990]. Correctness [2387, 548, 2907, 930, 2365]. Correlation [1875, 1272, 2681, 2683, 1495]. correlation-immunity [1495]. correspondence [1659]. correspondences [1560]. Corresponding [325]. cortical [1640]. Cosmic [1982]. Cost [156, 3091, 853, 2678, 1382, 912, 1851, 2857, 1746, 1057, 887, 1260, 1043, 3063]. Cost-Sensitive [1260]. Costs [1253, 669, 2238]. Cottage [938]. Countable [652]. Counter [2144, 1228]. Counter-Valuations [1228]. Countermeasures [2990]. Counters [2377]. Counting [3128, 2870, 3069, 2196, 3176, 455, 1463]. Coupled [1026, 328, 1651]. Coupling [645, 1257, 82]. Course [2015, 1673, 1678].

Courses [1513]. Couveignes [2194]. Covalently [2998]. Cover [452, 2991, 291, 2888]. Coverage [2333]. Covering [2788, 520, 2971, 2263, 2882, 2883]. Covers [667]. Covert [2674]. CPTHEO [2091]. CQ [327]. Craik [1929]. CRAY [1334, 1410, 1419]. Creaseness [1624]. Creating [529]. Creation [3031, 803, 318]. Creative [212]. Credit [1092]. Credit-Scoring [1092]. CREW [674, 675]. CRIL [2100]. Criteria [1983, 1747, 714]. Criterion [1494]. Critical [705, 707, 708, 1693, 1694, 825, 623, 200, 12, 1067]. Cross [532, 533, 180, 2616]. Cross-media [532, 533]. Cross-Platform [180]. Crossing [954, 979]. Crossover [3029, 1019, 201, 3020]. Cryptanalysis [1467, 1466, 662, 407, 1172, 422, 405, 1474, 1170, 423, 421, 425, 2675]. Cryptanalytic [416]. CRYPTO [2762, 2764, 2763, 2761, 2734, 2744, 2750, 2730, 2732, 2756, 2754, 2748, 2728, 2742, 2760, 2758, 2746, 2738, 2740, 2752, 2736]. Cryptographic [684, 2361, 2687, 1185, 2056, 2174]. cryptographically [1496]. Cryptography [2488, 663, 1469, 1178, 1182, 2693, 2489]. Cryptosystem [1175, 2173, 1482, 161, 1176, 1483, 1180, 1481, 2506, 1501]. Cryptosystems [155, 2486, 2670, 2671, 2676, 2502, 1461, 2500, 2499, 1460]. CS [1427, 417]. CS-2 [1427]. CS-CIPHER [417]. CSP [1899, 716]. CSP-Z [716]. CSPs [2034]. Cube [660, 1085]. Cubes [579]. Cubic [450, 2202]. Cues [1601, 1600]. Cultural [2977, 2616, 2973, 2972, 2975]. Culturing [2976]. Cumulative [2955, 2704]. Cunningham [2174]. Current [2709, 1300]. Curried [119]. Curvature [1624]. Curve [2488, 2195, 2154, 1198, 2204, 2500, 2506, 1463]. Curved [989]. Curves [2194, 3058, 2171, 1622, 143, 2159, 2201, 2197, 2185, 2172, 2184, 2501, 1577]. Customer [1442, 1028]. Customizable [597, 230]. Customized [799]. Cut [676, 810, 2220, 677, 1992]. Cut-and-Paste [2220]. Cut-Off [810]. Cut-Through [1992]. Cuts [1743]. Cutting [3026, 1897]. CVS [592]. CyberGlass [2572]. Cyberspace [2582]. Cycle [450, 3029, 177]. Cycles [1533, 2074]. Cyclic [709, 2193, 2908, 1892]. Cyclotomy [2158]. cylinders [1596]. CyPhone [322]. Czech [2693].

D [2595, 1617, 1831, 1423, 490, 1560, 2596, 1609, 1354, 2597, 1403, 1554, 1597, 387, 1611, 1664, 1610, 2594, 2317, 1666, 989, 2318, 514, 438, 1661, 1578, 2319, 2595, 990, 1646, 2258, 1138, 1648, 1662, 1641]. D-dimensional [438]. D-Object [1610]. D/ [2595]. D0L [3142]. D0L-Systems [3142]. D’Agents [2057]. Daily [1981, 212]. Damaged [980]. Data [1837, 588, 1954, 794, 2783, 108, 1543, 1965, 1075, 1074, 2300, 280, 598, 271, 2281, 1295, 2077, 1356, 2630, 706, 1076, 1687, 1077, 1392, 1435, 2217, 253, 996, 1117, 1099, 1327, 1365, 544, 1288, 577, 1082, 584, 341, 1083, 530, 531, 1293, 1284, 170, 3057, 3052, 1085, 1070, 664, 1086, 1545, 1550, 277, 257, 1087, 1758, 1757, 344, 1339, 2207, 1110, 1973, 279, 949, 1242, 113, 2261, 1112, 1114, 321, 1915, 1038, 1399, 574, 905, 1115, 2544, 3051, 2210, 1871, 2581, 252, 2954, 2541, 579, 700, 1241, 229, 251, 840, 2521, 1429, 1367, 263, 2291, 1119, 1355]. Data [2716, 1120, 1372, 250, 594, 266, 1102, 2239, 736, 2206, 1640, 564]. Data-Computing [1372]. Data-Intensive [598]. Data-Mining [1115]. Data-Parallel [252, 1429]. Data-Processor [251]. Data-Refine [1070]. Data-Security [2630]. Data-Structures [170]. Database [1548, 109, 335, 1539, 590, 110, 646, 584, 1532, 336, 1298, 803, 1769, 343, 328, 1540, 1363, 2570, 2017, 963, 352, 1366, 1367, 599, 2292, 116, 1544, 1297, 1767, 220]. Databases [342, 2695, 589, 1898, 571, 1285, 340, 580, 1286, 570, 1293, 1287, 1538, 602, 1088, 112, 113, 1090, 543, 1116, 1534, 115, 1537, 1546, 1121, 572, 1364, 1583]. Datalog [351, 348, 1542]. Datasets [2240]. Datatypes [2135]. Daubechies [2258]. DBMS [1547]. DDM [1412]. De-amortization [3054]. Deadlock [2077, 2014, 2013, 2012]. Deadlocking [3126]. Debugger [699]. Debugging [1387, 745]. Decades [517]. Decentralised [1872]. Decentralized [2346]. Decidability [636, 2807, 2821, 1213]. Decidable [634, 626, 633]. Deciding [2802, 776, 777, 2816, 618, 2643, 1129]. Decision [2281, 209, 2155, 1253, 2278, 2379, 1131, 1845, 2110, 1365, 1910, 3123, 448, 1156, 1109, 1829, 1303, 1239, 1133, 2277, 2111, 1941, 2280, 2214, 2262, 1130, 1872, 1977, 2275, 1231, 1120, 1963, 2373, 2232, 128]. Decision-Making [1845]. Decision-Mechanism [1131]. Decisions [2695, 2441]. Declarations [1225]. Declarative [699]. Decoder [1440]. Decoding [2349, 1181]. DECOMATE [1053]. Decomposing [2365]. Decomposition [970, 2075, 253, 1362, 1351, 3067, 829, 2288, 1649]. Decompositions [3163]. Decoration [501]. Decorrelation [449]. Decoupling [1619]. Decreasing [2818]. Decryption [1482, 1177, 1481]. DECT [808]. DEDALE [336]. Dedicated [360, 66]. Deduction [685, 2120, 553]. Deductions [2086]. Deductive [828, 1541, 2119]. DEF [257]. DEF/USE [257]. Default [2259, 122, 134]. defaults [554]. Defeat [152]. Defect [3159, 3160, 1888]. Definability [2869, 2223]. Definable [2800]. Define [2568]. Defined [1018, 1058, 455]. Defining [3092, 478, 2952, 2440]. definite [638]. Definition [1833, 802, 2472, 269, 46, 47]. Definitions [1715, 2124, 1716, 3175, 2833, 1506, 2151]. Deflection [1991]. Deformations [824]. degeneracy [1671]. Degree [3164, 1731, 1494, 2818, 325, 443, 666, 1255]. Degree-Decreasing [2818]. Degree-Preserving [3164]. Degrees [337, 2193]. Delaunay [658]. Delay [222]. Delayed [2244, 1991]. Delays [905]. Delegation [2695]. Deleting [1229]. Deletion [1222, 1223]. Deliberation [1167]. Delivery [1053, 2454]. Demand [1451, 2346]. Demosaicing [1590]. DenK [482]. Denotational [2829]. Dense [900, 1576, 2168, 293, 2894]. Density [2935, 2866, 2166, 3161, 2867, 1556]. Department [781]. Dependability [962]. Dependable [945]. Dependence [1517, 1665]. Dependences [698]. Dependencies [2077, 1940, 113, 52, 2701, 2702, 728]. Dependency [639, 1782, 763, 640, 2025]. Dependency-Relations [1782]. Dependent [1019, 885, 1927]. Deployable [928]. Deploying [1371]. Deployment [2654, 2726, 2461]. Depth [3088, 2278, 1019, 2280, 1650, 1570]. Depth-Dependent [1019]. Derandomization [283]. deregulation [1770]. Derinability [479]. Derivation [2824, 2133, 920, 2144, 2134, 2141]. derived [550]. Deriving [2871, 1826, 935, 289, 838]. Dershowitz [2844]. Descendants [634]. descent [727]. Describe [2427]. Describing [1543, 2418, 1165, 1772, 2419, 2420]. Description [2090, 2093, 2103, 2104, 348, 1161, 1520, 2091, 2089, 2092, 2100, 1145, 2099, 2415, 2101, 2102, 2718]. Descriptions [1274, 76, 77, 1156, 2235, 57, 2416, 1609, 124]. descriptors [1602]. Design [19, 536, 598, 1809, 2438, 1324, 758, 365, 589, 109, 910, 1308, 1309, 31, 17, 18, 226, 1027, 2356, 932, 410, 1957, 1539, 2003, 1291, 1117, 60, 1039, 20, 404, 23, 79, 1532, 247, 508, 2355, 32, 2969, 395, 371, 246, 316, 22, 2285, 713, 2441, 2394, 42, 418, 2411, 2901, 90, 2902, 823, 1030, 114, 769, 797, 715, 2915, 2010, 1036, 167, 2487, 510, 2642, 1914, 948, 2341, 2973, 320, 1370, 228, 368, 522, 2351, 499, 2699, 53, 2903, 800, 801, 507, 1065, 506, 995, 356, 2000, 725, 116, 362, 2780, 129, 1496, 564, 2880]. Designand [918]. Designed [197, 908, 413]. Designing [1962, 1760, 512, 946, 2994, 2001, 225, 3015, 1068, 1067, 2579]. Designs [2403, 764, 1723, 2921]. DESIRE [237]. Desktop [1378, 222]. Destage [1437]. destructive [1019]. Detail [588]. Detecting [762, 2953, 1244, 736, 1611, 591, 151, 2701, 2702, 1644]. Detection [975, 2077, 1877, 882, 233, 396, 397, 1923, 2384, 3044, 2295, 1889, 2330, 1640, 1616]. Determinant [2534]. determinate [2933]. Determination [1966]. Determining [1272, 1589]. determinism [727]. Deterministic [459, 2549, 2525, 3004, 1996, 619, 2856, 620, 3118, 898, 2647, 2817]. Deterministic/Adaptive [1996]. Determinization [2841]. Developing [2909, 1860, 597, 489]. Development [2438, 702, 1056, 1151, 2473, 1693, 1694, 2898, 858, 2429, 9, 544, 947, 2666, 2461, 2029, 67, 1109, 1389, 2669, 2121, 950, 1764, 2719, 2474, 65, 1323, 1143, 212, 1057, 2322, 1977, 2435, 2600, 1963]. Developments [1673, 2509, 2158]. Deviation [2206, 2207]. Device [1335]. Devices [156, 155, 512, 2573, 358]. DFA [2558, 2540, 2543]. d’hypertexte [541]. Diagnosing [1884, 2280]. Diagnosis [1953, 1886, 1890, 1303, 1891, 2279, 1821, 189, 1889, 133]. Diagnostic [2431, 1915, 1859, 2274]. Diagonal [3062]. Diagonal-Flip [3062]. Diagonalization [3135]. Diagram [459]. Diagrammatic [838]. Diagrams [475, 448, 2129, 55, 56]. Dialogs [496]. Dialogue [2550, 401]. Dialogues [495]. Diameter [3163, 3081]. Dictionaries [2805]. Dictionary [470, 152, 3036]. Difference [1326, 1956]. Differences [3012, 689, 2716]. Different [447, 1343, 2559, 107, 509, 1352].

Differential [1468, 426, 829, 406, 1386, 422, 1170, 423, 425, 2675, 1638]. Differential-Linear [1468]. Difficult [2811]. Difficulties [3]. Diffie [2155, 2496]. Diffusive [1343]. DIGISAT [2334]. DIGISAT-S3M [2334]. Digital [1504, 1034, 1035, 1049, 2343, 512, 2331, 1046, 2355, 2606, 1195, 159, 1444, 358, 2483, 357, 1477, 2498]. Dijkstra [3165]. Dilemma [2958]. Dimension [3069, 2772, 2257]. Dimensional [679, 3086, 582, 3058, 1365, 1023, 3042, 3057, 891, 3077, 3043, 381, 1816, 1024, 2980, 2931, 2583, 1517, 3087, 438, 124, 1562]. Dimensionality [1110]. Dimensioning [2066]. dinosaurs [730]. Diophantine [1738, 1415]. Direct [1414, 1913, 2195, 404, 1420, 2009, 522, 728]. Directed [1078]. Direction [1507, 400]. Directional [970, 3009, 1902, 971, 1149]. Directions [849, 1832, 217, 1012]. Directives [270]. directly [2490]. Disabled [2457]. Discipline [702, 691]. discontinuities [1566]. discontinuity [1627]. Discontinuous [1350]. Discourse [493]. Discover [2931]. Discovered [1079]. Discovering [1075, 1295, 1898, 3029, 1082, 374, 575, 44]. Discovery [2281, 2564, 2563, 1076, 1104, 1109, 2259, 1088, 2237, 2610, 1116, 1093, 2234, 579, 2262, 1098, 1100]. Discovery-Driven [579]. Discrepant [1199]. Discrete [1639, 903, 2691, 3104, 820, 385, 1262, 1638, 830, 2198, 2204, 1476, 1472]. Discrete-Continuous [830]. Discrete-Event [820]. Discretionary [1100]. Discretization [1107, 1113, 2284, 2275]. Discretizations [2283]. Discretized [1900]. Discriminants [2182, 2183]. Discrimination [1022]. DISCWorld [1376]. Disequations [631]. Disguised [431]. Disjoint [1732, 1568]. Disjoint-Path [1732]. Disjunctive [1741, 2769]. Disk [583, 1437, 2369].

Dispatch [2905]. Dispatching [2905]. Dispersed [1320]. Dispersion [2884]. Display [2567, 512]. Displaying [44]. Displays [357]. dissimilar [1618]. Dissimilarity [1852]. Dissipation [1183]. Distance [2511, 569, 2951, 3176, 2006, 2040, 3084, 3041, 3039, 1585]. Distance-Based [569, 2951, 2006]. Distances [2950, 2952, 2040]. Distant [394]. Distillation [1917]. Distinctions [1506]. Distortion [2569]. distortions [1629]. Distributed [1451, 1274, 2090, 2787, 1344, 1442, 2786, 49, 1853, 2456, 1414, 1407, 932, 1104, 10, 1432, 1387, 1413, 1354, 1327, 1884, 584, 210, 2907, 2068, 361, 2486, 1450, 79, 1707, 961, 1376, 787, 2674, 1550, 1329, 967, 427, 788, 861, 966, 173, 2897, 223, 1865, 1712, 306, 308, 1323, 1342, 2426, 1704, 2611, 310, 1375, 305, 99, 2409, 2410, 948, 2341, 1386, 1326, 1455, 1699, 1363, 2813, 873, 1445, 2417, 926, 2599, 2908, 944, 2569, 720, 219, 2858, 1068, 2384, 874, 1366, 1549, 1762, 93, 1069, 2661, 1372, 1345, 915]. Distributed-Memory [1327]. Distribution [2447, 3003, 146, 2513, 1918, 2331, 1354, 1220, 1221, 788, 313, 2698, 1198, 2539, 12, 1345, 1102]. Distribution-Sensitive [2539]. Distributions [2076, 196]. Distributive [689]. Distributivity [3122]. Disturbances [1971]. Diverse [2021, 2664]. Diversity [2418, 2023, 2419, 2420]. Divertible [1469]. Divide [1428]. Divided [1110]. Divisible [1502, 3078, 2492, 2105, 2106]. Division [2011]. Divisor [2201]. Dixon [135, 136]. dLIMIT [328]. DLP [1205]. DM1 [2300]. dMARS [235]. DNA [2996, 3049, 3045]. DNS [1404]. Do [360, 2166, 1561, 3168, 1935, 2817, 2608, 1405]. DO-loop [1405]. Docking [2998]. Document [536, 515, 523, 1048, 2235, 537, 522, 2556, 529, 1762, 538, 979]. Document-Centric [1762]. Documents [535, 2724, 2725, 539, 2624, 1545, 540, 524, 980, 538, 525]. Does [1015, 2011, 3017]. doesn’t [1251]. DOGMA [2945]. DOL [472]. DoLittle [2016]. DOLMEN [2474]. Domain [1432, 1861, 1351, 3067, 2971, 1163, 2439, 603, 1326, 2625, 1876, 150, 127, 2446, 2484]. Domains [1853, 1940, 1807, 311, 2930, 149, 1759, 1100]. DOME [2599]. Domestic [377]. Dominating [3085]. Domination [2814]. DOTPAR [1432]. Double [431]. Doubling [2651]. Dove [2777]. Down [2681]. Downwards [2143]. Dozen [334]. DPE [2467]. DPFS [1429]. DRAM [865]. Drawing [970, 972, 3056]. Drawings [975, 993, 390, 990, 988, 994]. Dream [184]. Drive [2411]. Driven [2443, 903, 1766, 2470, 2540, 2078, 955, 2544, 3019, 579, 250, 1765, 1770]. Driving [2018]. drop [2008]. DRT [1631]. Drum [1403]. DSM [1380, 2003, 1409]. DSMs [265]. DSP [860]. DTD [523]. Dual [3016]. Duality [1670, 1790, 2844]. Duplication [2072]. Duration [177]. During [1539, 1048, 2001, 1944]. Dwelling [376]. Dynamic [1061, 3031, 704, 3091, 2633, 2511, 671, 2968, 1913, 1402, 2851, 339, 1343, 650, 597, 2652, 842, 851, 2659, 170, 3057, 1979, 1970, 2470, 385, 2804, 257, 173, 78, 1021, 1761, 960, 1384, 2038, 3034, 1534, 542, 1032, 945, 2574, 581, 3020, 599, 2042, 1345, 177, 1581, 1619]. Dynamical [1637, 202, 831, 1836, 1880, 1636, 203]. Dynamically [1083, 856, 2697]. Dynamics [3175, 3033, 1408, 912, 2613, 1410, 1976, 2620, 195, 673, 847, 846, 835].

E-LOTOS [1768]. E-Unification [636, 630]. EAMMO [1165]. Early [1878]. Early-Based [1878]. Earth [1984]. Easy [1502, 3137]. ECOLE [166]. Ecological [1977]. Economics [2956]. Economy [2620, 355]. Ecosystem [3000]. Ecosystems [228]. Ecstasy [2727]. EDF [960]. Edge [1729, 1730, 184, 882, 1728, 907]. Edge-Connectivity [1728]. Edge-Splitting [1728]. EDI [320]. Editing [2722, 1537, 2723, 3039]. Edition [1430, 525]. Editor [1389, 2570]. Education [1674, 1050, 1141]. Educational [1320]. EEG [2269, 2270]. EEPROM [2692]. Effect [534, 1077, 519, 3159, 3160, 2574, 1507, 1880, 2024]. Effective [3072]. Effectiveness [1343, 1057]. Effects [1380, 2985, 1273, 1516]. Efficiency [2073, 2705, 1393, 1364]. Efficient [108, 2632, 3048, 676, 2348, 2832, 2518, 1749, 3058, 2865, 459, 1082, 689, 1023, 3057, 1085, 677, 1107, 1419, 3077, 1948, 2831, 1535, 751, 1997, 573, 1342, 1490, 935, 1503, 875, 1534, 1900, 2875, 2643, 1024, 92, 2698, 2645, 581, 2384, 2128, 1292, 759, 1641, 2177, 2537, 3064, 2821, 3039, 1194, 726, 2501, 1609, 1589]. Efficiently [1334, 2541]. Effort [2344]. EFSM [2378]. Egocentric [1506]. Eigenvalue [1419]. Eilenberg [652]. Elastic [1406, 1942]. Elasticity [2147]. Elections [148]. Electrical [2852, 1851, 1890]. electricity [1770]. Electronic [2471, 1442, 1443, 365, 1452, 1458, 528, 1449, 1450, 2618, 2689, 326, 2470, 2688, 2472, 148, 1455, 2617, 1459, 1879, 2619, 1456, 1446]. Electronics [2421, 2422]. Electrostatic [1337]. ELEM2 [2046]. Element [1350, 809, 1412, 1422, 1421, 1409, 1425]. Elementary [20, 138, 1245, 2063, 2141, 3123]. Elements [1826, 512, 2640, 2988, 981]. ElGamal [1176, 2495]. ElGamal-based [2495]. Eli [746]. Eliminating [911]. Elimination [2108, 2109, 2528, 2091, 2089, 3141, 2813, 116]. Ellipsis [1877]. Elliptic [2488, 2194, 2171, 2195, 2159, 2197, 1198, 2185, 2172, 2204, 2184, 2501, 2500, 2506, 1463]. else [2098]. Elutriation [1333]. Embedded [707, 708, 96, 939, 947, 1692, 709, 943, 946, 1999, 953, 2409, 2410, 948, 2253, 225, 941, 951, 2252]. Embedding [3162]. Emergence [1974, 2579]. Emergent [1812, 2581]. Emission [1838]. Empirical [3027, 1110, 870, 2045, 200]. Employee [1680]. Employing [157]. EMud [2584]. Emulation [868, 897]. EMYCIN [2779]. Enabled [2453]. Encephalomyopathies [2275]. Encoding [1868, 3114]. Encodings [2809, 2861]. Encompassed [333]. Encrypting [152]. Encryption [1478, 2691, 409, 664, 1498, 420, 422, 1183, 423, 2495]. End [1053, 2903, 225]. End-to-End [2903]. End-User [225]. Ends [2798]. Energy [3002]. Enforcement [373]. Engeler [333]. Engine [1778, 810, 3023, 1354, 570, 2606, 1345]. engineer [2972]. Engineering [1310, 1062, 2427, 1774, 724, 2476, 210, 1755, 1329, 1316, 1810, 1358, 2973, 1066, 990, 1753, 988, 2716, 1, 2, 50, 551, 58]. Engines [1106, 2378]. English [2027, 2033]. Enhance [1401, 1433]. Enhanced [520, 2332, 1363]. Enhancement [212]. Enhancing [2404, 2619, 2463]. Enough [1580]. Enqueue [163]. ENS [1315]. Ensembles [1867]. Entailment [2853, 2949, 2929]. Entities [154]. Entity [157]. Entropy [2840, 1179]. Entry [2429]. Enumerable [477, 3173]. Enumeration [2536]. Enumerations [610]. Environment [1931, 1442, 1368, 1407, 2396, 166, 1760, 359, 2358, 858, 577, 494, 2398, 597, 1768, 1979, 1376, 325, 967, 1389, 803, 2121, 950, 394, 2335, 321, 44, 310, 305, 364, 2599, 358, 1929, 2571, 3020, 1429, 1549, 2452, 2723, 854, 354, 2780, 2333, 1564]. Environmental [1911, 2050]. Environments [2471, 914, 1356, 2466, 2959, 2976, 1862, 1381, 150, 2573, 107, 377, 2614, 1426, 788, 403, 2586, 2719, 1761, 1316, 2657, 2574, 2569, 800, 801, 229, 2993, 1530]. environnement [541]. EP-like [2986]. Epipolar [1565, 1566]. Epistasis [200]. Epistemic [243]. Equal [574]. Equality [2108, 2109, 2375, 2832]. Equalization [1952]. Equals [479]. Equation [1738, 1174]. Equational [604, 3144, 2097, 464, 1792, 628, 545, 2099, 3151, 121, 553]. Equations [1833, 3157, 681, 426, 2197, 2861, 1415, 627, 1823, 1471]. Equilibria [194]. Equilibrium [2076]. Equivalence [447, 813, 2072, 1070, 1463, 624, 2826, 3118, 2821, 2204, 559, 1788]. Equivalences [2872, 2816, 2859]. Equivalent [1464, 2185, 2254, 2184]. Erlang [2090]. Erratum [3178]. Error [802, 990, 1269, 2347, 1252, 1657]. Error-Correcting [1269]. Error-Robustness [2347]. Errors [2831]. ERW’97 [2444]. ES/EP [2986]. ES/EP-like [2986]. Escrow [1181, 1180]. Establishing [1538, 2017]. Establishment [2700, 2699]. Esterel [2396]. Estimates [817]. Estimating [3003, 3053, 1648]. Estimation [2935, 1080, 1881, 389, 1652, 2647, 1122, 1982, 1573, 1569, 1620, 1562, 1598, 1570, 1619]. Estimation-Segmentation [1652]. Estimators [1252]. ESTL [2082]. ETA [715]. Ethernet [174, 913]. Ethernets [171, 172]. Euclid [333]. Euclidean [1579, 2857, 1535, 294, 1942, 2176, 1570, 2157, 973, 142]. Eulerian [3089]. EURESCOM [2451]. EUROCRYPT [2762, 2764, 2763, 2761, 2735, 2731, 2729, 2733, 2739, 2745, 2747, 2753, 2759, 2741, 2755, 2751, 2737, 2757, 2743, 2749]. EUROMED [1321]. European [1053, 91]. EV1 [2301]. Evaluate [590, 881]. Evaluating [2018, 2014]. Evaluation [860, 583, 3102, 914, 2115, 2002, 1968, 643, 1983, 2432, 1713, 1427, 79, 2178, 1440, 751, 797, 2915, 1996, 382, 1370, 998, 1210, 2176, 2562, 997, 1667, 1562]. Evaluations [1338]. evaluators [729]. Evalution [2960]. EVEN [1018]. EVEN-5 [1018]. Event [607, 903, 820, 2078, 763, 306, 848, 1916, 1917, 1294]. Event-Based [848, 1916, 1917].

Event-Driven [903, 2078]. Events [2082]. Every [3080]. Evidence [150, 2540]. Evidence-Driven [2540]. Evolution [2962, 59, 1778, 808, 50, 2976, 1851, 205, 1275, 884, 3016, 1023, 367, 206, 3002, 3001, 2441, 2586, 1424, 185, 1024, 2413, 2414, 1848, 1020, 2442, 3101, 2023]. Evolution-Based [1851]. Evolutionarily [2981, 2982]. Evolutionary [3012, 2301, 2967, 2959, 2995, 2970, 3018, 2898, 1868, 1681, 1845, 199, 2955, 186, 884, 2991, 2996, 2984, 1869, 2969, 3027, 2971, 3004, 2983, 3026, 1654, 2963, 885, 1029, 878, 3019, 188, 1031, 1905, 196, 2988, 198, 189, 3015, 3008, 2980, 2956, 2992, 2242, 3014, 2965, 2975]. Evolvable [2968, 592]. Evolve [3000, 1032]. Evolving [3025, 1026, 2995, 884, 877, 111, 3030, 1866, 3028, 2599, 1031, 1020, 1927, 2981, 2982, 2993, 192, 3032, 2961, 3031, 1154]. Exact [614, 3070, 1904, 3134, 3084]. Exam [1679]. Examination [1677, 1683]. Example [1203, 2407, 2843, 1648]. Examples [2871, 1264, 1265, 2048, 192]. Exception [2282]. Exceptions [1760, 687, 2366]. Exchange [1504, 152, 3089, 1197, 1988, 2507]. Exchanging [645]. Exclusion [451, 779]. Executable [686, 795, 796, 4]. Executing [3078]. Execution [1701, 107, 1702, 1388, 232, 1342, 2320, 2366, 2363, 1345]. Exemplar [1436]. Exhibit [371]. Exhibition [1960]. Existence [1794, 445]. Existential [3168, 2826, 1802]. Existentially [2800, 339]. Existing [374, 2482]. Exit [3040]. Exogenous [1510]. Expansion [2180, 2181]. Expected [474]. Expensive [581]. Experience [273, 2898, 361, 1934, 326, 2357, 869, 1242, 778, 909, 1710, 48]. Experiences [2444, 2714, 2909, 1048, 8, 2454, 216, 315, 2709, 2413, 2414, 1047, 1318]. Experiment [2383, 20, 2708, 2426, 1144, 2585, 1982]. Experimental [1954, 1310, 2167, 371, 2192, 222, 2983, 1014].

Experiments [2073, 1407, 1244, 1410, 1258, 1529, 934, 1680, 1904, 2553, 198, 1530]. Expert [1855, 1812, 1915, 1916, 1917, 2781]. Experts [2548, 2625, 2547]. Explain [2583]. Explaining [1245]. Explanation [1267, 3101]. Explanation-Based [1267]. Explanations [1879]. Explicit [2862, 3105, 232, 2096, 1132]. Explicitly [256, 255]. Exploitation [2065]. Exploiting [754, 270, 394, 263]. Exploration [772, 2976, 1970, 579]. Explorative [2986]. Exploratory [2957]. Exploring [519, 602, 1370]. Exponent [627]. Exponentiation [1475, 1477]. Expressed [637]. Expression [3040]. Expressions [3115, 711, 834, 1517, 1613]. Expressive [348, 471, 3109, 1802, 1289]. Expressiveness [1543, 2812]. Extended [770, 1504, 2438, 2884, 2160, 1478, 1464, 1832, 1286, 2192, 734, 751, 1490, 1055, 3035, 131, 2644, 2622, 2179, 136, 1560, 748]. Extending [2899, 1302, 1435, 2888, 1541]. Extensibility [1759]. Extension [2918, 523, 2245, 2230, 32, 51, 2919, 269, 1700]. Extensional [2094, 2095]. Extensions [30, 431, 2804, 628, 1739, 2247, 1292, 263]. External [2518, 2025]. External-Memory [2518]. Extracting [55, 56, 1876, 981]. Extraction [993, 1870, 3023, 502, 2268, 1769, 1599, 1271, 1121, 1231, 979]. Extractors [434, 2522]. Extracts [22]. Extrapersonal [1512]. Extrapolation [1358, 1641]. Extreme [702]. Extrinsic [1624].

Fabric [1994]. Face [2077, 815, 1653, 2566, 1654, 1648]. faces [1655]. Facial [3153]. Facilitate [1545]. Facilitating [244]. Facilities [1378, 1401, 1433]. Facility [2511, 1734, 1709]. Facing [1484]. FaCT [1205, 1232]. Factor [3158, 1081]. Factoring [1464, 1483, 1476, 1463]. Factorization [2170, 1500, 2806, 2161, 2670, 2671, 1595, 2491, 1570]. Factorizing [196]. Factors [1510, 1130]. Factotum [767]. facts [726]. Fail [920]. Fail-Aware [920]. Failure [233, 915]. Failures [2691]. Fair [1504, 2449, 1795]. Faithful [1596]. Families [2433, 2442, 2440, 2437]. Family [2434, 2443, 2436, 2412, 2418, 2411, 1185, 2409, 2410, 2413, 2414, 2419, 3081, 2420, 2504]. FAMOOS [50]. Far [1696]. Fast [2695, 679, 168, 1477, 3037, 174, 410, 409, 1910, 930, 2178, 1482, 180, 2237, 1174, 1420, 3034, 3035, 2114, 1900, 2685, 2537, 2263, 3036, 1639, 1481, 740, 2503, 913]. Faster [1076, 3009, 650, 2805]. Fault [922, 73, 1953, 927, 917, 923, 16, 1890, 964, 2032, 1821, 1700, 925, 944, 928, 888, 2661, 1889, 924]. Fault-Tolerant [922, 927, 923, 16, 964, 1700, 925, 888, 924]. Faults [155, 2279, 2280]. FD [1326]. FD-TD [1326]. FDR [718]. Feature [1870, 1099, 1249, 1248, 637, 1271, 506]. Feature-Based [506]. Features [2962, 1254, 1649]. February [1298]. Federated [1539]. Federation [2476, 2462]. Federations [328]. Feedback [1135, 2682]. Feedforward [1135, 3015]. Feistel [412]. Fermat [2197]. Fermentation [1880]. Few [1904, 2048]. FFT [1360]. FGAL [1821]. Fibring [1215]. Fido [1793]. Field [2162, 2163, 1632, 886, 857, 2203, 3063, 2204]. Fieldbus [94]. Fields [2182, 2183, 2190, 2607, 2191, 2196, 2192, 1903, 2193, 2199, 2169, 2200, 457, 2202, 2185, 2184, 1620]. FIFO [2380]. Figure [476]. Figures [1850]. File [2764, 1429]. Filter [1766, 3042, 1603, 1765]. Filter-Mechanism [1766, 1765]. Filtering [2624, 2623, 2620]. Filters [2258, 2993, 3032]. Finance [1357]. Financial [2796]. Finding [1659, 450, 1602, 3085, 2283, 3002, 2069, 1174, 2038, 1485, 2894, 2174, 1471]. Fine [2719]. Fine-Grained [2719]. Fingertips [396, 397]. Finite [1350, 1175, 2264, 2841, 2379, 1240, 2705, 2876, 1220, 1221, 2549, 1412, 2636, 757, 1422, 2548, 2196, 2816, 1262, 191, 2838, 1421, 2120, 471, 2360, 2642, 2815, 625, 1326, 2169, 3117, 1409, 896, 2541, 2988, 457, 2647, 3173, 2547, 726]. Finite-Control [2838]. Finite-Difference [1326]. Finite-State [2379, 2636, 2816, 2120, 2360, 2642, 2815]. Finite-Valued [1220, 1221]. First [1423, 2932, 2918, 2800, 1049, 1266, 424, 2092, 3131, 2638, 2096, 2952, 2929, 463, 2373]. First-Order [2932, 2918, 2800, 1266, 3131, 2096, 2952, 2929, 463, 2373]. Fitness [2964, 183, 1016, 1021]. Fitting [501, 1596]. Five [2001]. Fixed [776, 777, 3078, 751, 2803, 3046, 2980, 898]. Fixed-Connection [898]. Fixed-Length [2980]. Fixpoint [2116, 689]. Flame [1411, 1434]. Flashover [1828]. Flat [1063, 144]. Fleet [1029, 878]. Flexibility [1910, 2081]. Flexible [923, 2067, 1910, 1622, 2998, 313, 2904, 149, 2467, 3044]. Flight [1720]. Flip [3062, 3064]. Flipping [475]. Floating [2395]. Floating-Point [2395]. Floats [428]. Flocking [2581]. Floorplanning [2243]. Flow [1325, 1746, 2509, 1317, 1745, 1333, 2145, 264, 1652, 1632, 733, 3108, 718, 2006, 1744, 266, 1639, 749, 732, 2146, 1633, 1619, 1372]. Flow-Control [718]. Flowcharts [2830]. Flows [1414, 476]. Fluid [1325, 1333]. Flushing [2363]. Fly [2380, 2370, 197, 2249, 1388, 1491]. Folding [188]. folds [1640]. follow [727]. follow-determinism [727]. Following [1925]. Font [501, 508, 505]. Fonts [518, 506]. Football [1140]. Forbidden [3158]. Force [3063]. Forces [3028]. Forecasting [1894, 1892]. Foreground [996]. Foreign [3057, 2514]. Forest [1070]. Forests [3164, 900]. Form [2076, 2127, 256, 980, 2641, 2161, 116, 979, 1579]. Formal [2530, 1056, 1478, 2194, 2361, 750, 74, 75, 1105, 2383, 2358, 235, 724, 1814, 40, 1520, 158, 239, 2357, 2144, 3094, 769, 825, 1064, 799, 2393, 795, 796, 1066, 1057, 2119, 2180, 718, 781, 2699, 46, 47, 2363, 722, 48, 2181, 556]. Formalisation [1060]. Formalism [797, 800, 801]. Formalisms [49]. Formalization [5, 6, 2140]. Formalizing [36, 1775, 90]. Format [1687]. Formation [240, 1272, 2617, 234]. Formatters [532, 533]. Forms [610, 978, 634, 2126, 463]. Formula [1693, 1694]. Formulae [1216]. Formulas [2370, 3131]. Formulation [1782, 916, 510]. Forte [1690]. FORTRAN [551, 1396, 1336, 269]. Forty [1237]. Forward [919, 1003]. Foundations [2310, 1302, 1487, 243]. Foundry [1888]. Four [144]. Fourier [930, 2045, 1619]. FPGA [856]. FPGAs [942, 852]. Fractal [2250]. Fraction [2180, 2181]. Fractional [614, 1255]. Fragile [80, 2912]. Fragment [2110, 2111]. Frame [715]. Frames [1518, 1520, 1669]. Framework [1441, 1451, 1452, 30, 1423, 1265, 2776, 1940, 37, 1432, 1387, 2408, 2955, 1860, 215, 1961, 1993, 1284, 247, 8, 33, 1552, 1379, 223, 1188, 1030, 328, 489, 545, 2451, 1271, 2112, 29, 1066, 2952, 2903, 366, 236, 2467, 701, 26, 2463, 1639, 1613, 2139, 741, 2888, 1612, 119, 125]. Framework-Specific [30]. Frameworks [31, 17, 18, 2261, 25, 3008]. Fredman [2804]. Free [1216, 2823, 662, 1601, 3153, 653, 1401, 1433, 2560, 3115, 3161, 654, 1113, 715, 1225, 2641, 1726, 2944, 148, 2643, 1923, 3126, 2105, 2106, 1600, 550, 2892]. Freezing [642]. Frege [1781]. French [1144]. Frequency [199, 427, 2013]. Frequent [575]. Frequential [2214]. Friendly [2981, 2982, 2036]. FRIENDS [918]. Frobenius [2501]. Front [1411]. Fruitful [799]. FST [2646]. Full [3119, 1789, 497, 1795, 1438]. Full-JAVA [1438]. Fulltext [1036]. Fully [3006, 346, 3132, 2998, 751, 3133, 3080, 2889]. Fun [2148]. Function [3127, 1756, 1825, 2785, 3029, 268, 1080, 3166, 1786, 2971, 1824, 2199, 2944, 2200, 2202, 2203, 2204, 2186, 2503, 2504]. Function-Free [2944]. Functional [1789, 1209, 692, 2901, 1540, 169, 2288, 1929, 4, 3134, 683, 334]. Functional/Concurrent [1789]. Functionality [1286]. Functionals [1797, 3134]. Functioning [3101]. Functions [3031, 415, 2871, 1018, 3088, 2511, 662, 3072, 2964, 1279, 431, 3053, 2375, 3137, 183, 2376, 3154, 2255, 478, 345, 3070, 1424, 195, 2685, 2374, 1495, 2260, 1496, 1485, 1633]. Functor [617]. Fundamental [1559, 657, 901, 2535]. Fundamentals [514]. Fusion [1550, 986, 2839, 572]. Future [794, 371, 2897, 1012, 362]. fuzziness [2208]. Fuzzy [3106, 1826, 2245, 1924, 1983, 1820, 1873, 1083, 2246, 1864, 1829, 2260, 65, 1424, 1882, 2994, 1821, 1988, 1932, 1823, 1121, 2231, 2975, 1933, 2229]. Fuzzy-Neural [1864].

GA [2964, 2945, 200, 1679, 1425, 1928, 1894, 1852]. GA-Based [1928, 1425, 2945, 1679]. GA-Optimal [2964]. GA1 [2303]. Gaea [1166]. GAGPC [1846]. Game [1073, 2289, 2777, 2846, 1165, 1795, 1267, 2778]. Game-Theoretic [1795, 2778]. Games [1776, 1139, 1779, 2830, 463, 756]. GAMMA [913, 1411, 912, 1166]. Gap [2011, 2140]. Garbage [2907, 2908]. Garden [2582]. Gas [1414, 2412, 1436]. Gate [857]. Gates [3127]. Gateway [2481, 2475]. Gateways [1384]. Gauge [916]. Gaussian [1657, 3141]. Gaze [382, 400]. GCD [2156]. GDMO [3]. Gear [769]. GeM [220]. Gene [1957]. General [2386, 2303, 2304, 2962, 2513, 69, 1169, 1226, 1748, 1281, 2199, 119, 1485]. Generalised [727, 2830, 1055]. Generalising [2130]. Generalization [452, 441, 577, 2946, 1267, 1123, 1102, 119]. Generalizations [2881]. Generalized [2632, 1846, 1676, 919, 3030, 2225, 830, 1726, 1744]. Generalizing [603, 2248]. Generate [1824]. Generated [3144, 311, 2222, 264, 1526, 513]. Generating [1139, 2170, 2379, 2190, 1869, 2126, 2773, 2185, 1909, 1879, 1684, 2184]. Generation [1074, 1214, 267, 3065, 2089, 889, 1979, 803, 2027, 252, 2050, 2263, 1101, 1631, 1491, 541]. Generations [3033]. Generator [2683, 83, 749, 747]. Generators [416, 2819, 1171]. Generic [346, 590, 1976, 1806, 1899, 1465, 1490, 2416, 1827, 701]. Genericity [969]. Genetic [193, 3031, 3006, 2022, 1870, 1846, 1027, 2219, 3033, 1022, 1868, 3013, 2243, 1028, 184, 2979, 1906, 1881, 190, 1015, 183, 1849, 1019, 2241, 1822, 2997, 1021, 191, 1030, 2653, 2974, 1154, 3007, 1424, 2990, 2023, 2994, 1017, 2025, 194, 2960, 1276, 1821, 1866, 1014, 2987, 2966, 1400, 3024, 3020, 2583, 2024]. Genome [3050]. Genuinely [2828]. Geographically [1320]. Geometric [3059, 2884, 2088, 891, 135, 294, 2808, 901, 140, 1659, 1664, 1626]. Geometries [142]. Geometry [678, 1515, 141, 650, 138, 145, 2087, 139, 1904, 137, 1565, 1573, 1656, 1669, 1577]. GEORAL [485]. Gestural [395]. Gesture [398, 383, 392, 393, 403, 388, 384, 399, 378]. Gesture-Based [398, 383]. Gestures [1961, 390, 385, 380, 381, 394, 382, 1613]. Get [1969]. Get-Bits [1969]. Getting [330]. GF} [2682]. GH1 [2304]. Gigabit [2010]. Girault [2497]. Give [452]. Given [2982]. Glauber [673]. GLiB [3031]. Global [1018, 2802, 1348, 1913, 300, 1107, 1757, 81, 1377, 2199, 1328, 2858, 2384, 1656, 741, 1606, 1621, 1633]. Global/Local [2858]. Glory [1527]. Glossary [1299, 1298]. GNAT [1695]. Go [1502, 1696, 1015, 302, 2766]. Goal [1073, 240, 1902, 1156, 1770]. Goal-driven [1770]. Goals [1126, 1034]. God [1251]. Goedel [1804]. GoF [19]. Goggles [1717]. Goldbach [2167]. Good [1580, 1015, 3027, 3105, 2834, 4, 1495]. Governing [2931]. Governmental [1762].

GRADE [1389]. Gradient [837]. Grain [2712, 3059]. Grained [2719]. grains [2208]. Grammar [2302, 2936, 2552, 2545]. Grammar-Rules [2936]. Grammars [2559, 2717, 2222, 2641, 728]. Grammatical [2550, 1020, 2556]. Grand [1896, 2616]. Granular [1403]. Granularities [585]. Granularity [1299]. Granulation [2211]. Graph [665, 442, 2824, 3144, 610, 3166, 186, 3145, 461, 1222, 1223, 712, 788, 255, 2250, 722, 1685, 3083, 3056, 1686, 3080, 2138, 737, 555]. Grapheme [2027]. Graphic [974, 512]. Graphical [1389, 1065, 505]. Graphics [993, 991, 999, 996, 971, 998, 969, 997]. Graphics/Text [997]. Graphs [450, 1731, 3085, 2809, 54, 3153, 668, 2528, 900, 446, 1722, 429, 2879, 451, 669, 426, 1361, 2794, 3147, 479, 667, 3079, 666, 1728, 1726, 1817, 1800, 3148, 3081, 2373, 730, 1615, 2892]. Grassmann [141]. Grayscale [518]. Great [2599]. GRED [1389]. Greedy [1732]. Greenhouse [1922]. Grey [1045]. Grid [2525]. Grids [3162, 1418, 1328, 1801]. Gripen [1720]. Gröbner [135]. Grounding [2593, 2584]. Group [2705, 1192, 2121, 921, 159, 2908, 2502, 128]. Grouping [437, 1595]. Groups [2194, 2691, 2189, 2195, 653, 2703, 2192, 654, 1465, 1700, 2201, 3118, 2105, 2106, 2961]. Groupware [1762]. Growing [2594, 619]. GS1 [2302]. GSPN [2066]. GSPNs [2075]. Guarantee [2391]. Guaranteed [2448, 2491, 568]. Guard [606]. Guarded [2110, 2111]. Guarding [2059]. Guess [3174]. Guidance [1719]. Guided [489, 1772, 987]. Guides [1406].

Hades [1966]. HAL [2398, 164]. Half [1831, 334, 2887, 2888]. Half-Dozen [334]. Half-planes [1831]. Halftoning [514]. Halving [3050]. Hamiltonian [450, 2138, 3081]. Hammer [492]. Hand [382, 396, 397, 1560]. Handed [395].

Handicapped [1959]. Handling [1831, 962, 1666, 2234, 2262, 2437]. Handover [1761, 2066]. Handwritten [516, 2296]. Hao [1743]. Hao-Orlin [1743]. Haptic [383]. Hard [3088, 1742, 1703, 1701, 1788, 1702, 2944, 1750, 3151, 293, 3103]. Hardness [3102, 2804, 2890, 291]. Hardware [866, 2968, 410, 861, 1819, 1995, 865, 2503]. Hardware-Monitor [861]. Hardware-Software [1819]. Harmful [783]. Harmonic [1423]. Hash [415, 662, 1096, 2503, 2504, 1485]. Hashing [411, 409, 2685, 414]. Haulage [2600]. having [1626]. Hawk [2777]. Hawk-and-Dove [2777]. Head [3117]. heads [1563]. Health [3023]. Heap [2523, 2208]. Heat [1245, 1988]. Heat-Exchange [1988]. Hellman [2155, 2496]. Help [1311]. Helping [993]. Helps [2542]. Herbrand [2950]. Heterogeneous [2630, 2327, 1040, 2655, 2461, 1538, 602, 1029, 878, 2657, 1347, 1549]. Heuristic [2015, 2216, 1496, 1897, 1827, 888]. Heuristics [3025, 2259, 2940, 1690, 1909, 889]. Heyting [2224]. Hidden [386, 385, 1821, 384, 1830, 552]. Hierarchical [1837, 1344, 2724, 1912, 1806, 1192, 824, 1800, 507, 725, 2287, 1615]. Hierarchically [1032]. Hierarchies [772, 111, 3149, 456, 1794, 454, 2837]. Hierarchy [605, 433, 430, 2872, 2652, 2418, 1522, 3117, 1801, 2419, 3173, 2420]. High [1319, 2694, 1324, 582, 1718, 1371, 1989, 300, 2909, 1178, 1354, 899, 851, 3057, 1376, 958, 891, 1359, 1439, 1865, 350, 1997, 1719, 1375, 717, 2001, 1370, 1363, 908, 384, 1369, 1183, 1318, 941, 1366, 1546, 1367, 1453, 1405, 2713, 1612, 1578, 1396, 1336]. High-Accuracy [958]. High-Dimensional [582, 891]. High-Level [2909, 350, 717, 1453, 1405, 2713, 1865, 1612].

High-Performance [1989, 1359, 1997, 2001, 1370, 1369, 941]. High-Speed [1178, 1183]. High-Throughput [851]. Higher [2094, 2103, 2095, 2104, 1674, 1784, 643, 1789, 406, 2096, 2101, 1051, 1052, 1170, 1950, 2845, 119]. Higher-Level [1051, 1052]. Higher-Order [2094, 2103, 2095, 2104, 1784, 643, 1789, 2096, 2101, 2845]. Highly [2654, 104, 1495, 1418, 1988]. Highness [3068]. Hilbert [3087, 2112]. Hints [201]. HiPER [1997]. HiPER-P [1997]. Histograms [1610, 1659]. Historian [2712]. historic [2371]. Historical [1203]. histories [130]. History [2722]. HMM [849]. HMM/ANN [849]. Hoc [601]. HOL [1060, 2102]. Holistic [1623]. Home [325, 403, 2351, 302]. homogeneous [659]. Homomorphisms [1169]. Hopfield [3017]. Hopping [2164]. Horizons [2864]. Horn [431, 2944, 2122, 2929]. HOSMII [865]. Host [2673, 2960, 1276]. Host-Parasite [2960, 1276]. Hosts [927, 2054, 926, 2052]. house [1578]. HP [1436]. HP/ [1436]. HPC [1368, 1357]. HPCN [1330, 1317, 1316]. HPF [1396, 1398, 1341, 261, 1397, 1338, 1430, 1399, 251, 1340, 262]. HPF-Builder [1430]. HPF-MPI [1341]. HRS1 [2305]. Human [398, 383, 480, 1250, 1959, 213, 2598, 380, 1526, 396, 397, 2574, 722, 2211, 1640]. Human-Computer [480, 1959]. Human-Machine-Interaction [398]. Humanistic [702]. Humans [2567, 1528]. Humboldt [1151]. Humidity [1922]. Hunted [1106]. Hunter [1106, 2187]. HWW [1318]. Hybrid [1894, 1074, 2304, 2305, 810, 2022, 811, 812, 849, 813, 814, 1845, 1385, 1820, 1856, 821, 1854, 2937, 823, 824, 834, 2044, 825, 828, 830, 1996, 1972, 2963, 1029, 832, 1119, 833, 3014]. Hybridization [1838]. Hydra [3114]. Hyper [3147, 299, 3148, 1214]. Hyper-Graphs [3147, 3148]. Hyperbook [530, 531]. Hypercubes [3162]. Hyperedge [3144]. Hyperelliptic [2201, 2499]. Hypergraph [3091]. Hypergraphs [3146]. Hypermap [541]. Hypermedia [570, 2610]. Hyperplane [2536]. Hypertext [2725, 540, 680, 2930]. Hypothesis [1657, 198, 2966]. HySpirit [570].

I.Total [1982]. I/O [1383, 1385]. IBM [278, 178, 1340]. IC [2568]. IC1 [2306]. ICE [422, 423]. ICONDENSATION [1612]. Iconic [1660]. ID [2507, 1604, 1198]. ID-Based [1198, 2507]. IDEA [1468]. Ideal [2175, 162, 1367]. Ideas [1811]. identical [2886]. Identification [1940, 496, 1888, 2974, 1882, 2706, 2229, 2497]. Identifying [257, 1090]. Identity [2698, 1197, 1501]. Identity-Based [2698, 1197, 1501]. If [492, 2098]. if-then-else [2098]. ifs [1588, 3167]. II [1169, 2041, 1998, 2260, 11]. III [17]. IIR [2993]. illumination [1568]. Illusion [2603]. Illustration [3033]. illustrative [2880]. ILP [2953, 1244, 2939, 2926, 2954, 2927]. ILU [1418]. Image [2307, 2308, 515, 2866, 500, 1582, 1637, 1590, 2598, 2317, 1651, 1938, 2329, 1536, 2576, 2867, 400, 915, 1571, 1631, 1636, 1586, 1638, 1588, 1598, 1585, 1619, 1641]. Imagery [3024, 1578]. Images [970, 389, 1587, 980, 1958, 2257, 972, 987, 1243, 1602, 1563, 1618, 1648]. Imaginary [1482, 2193, 2201, 1481]. Imaging [169]. Immediate [1018]. Immersion [1354]. Immune [1332]. immunity [1495]. Impact [1510, 782, 80, 2025]. Imperative [2906, 2918, 692, 3108, 1233, 2148]. Implement [2638, 601]. Implementation [536, 1335, 589, 1341, 550, 2156, 1039, 20, 1288, 918, 2233, 792, 793, 1337, 2636, 1746, 8, 33, 2029, 1419, 316, 1211, 51, 2637, 595, 596, 1689, 2010, 2364, 1036, 1331, 1810, 1328, 2341, 1326, 2294, 92, 2645, 2479, 1292, 1233, 227, 726, 2500, 2503]. Implementations [509, 664, 2926, 301, 2483]. Implementing [19, 903, 319, 14, 1701, 1702, 344, 165, 99, 2921]. Implication [1826, 2607, 2946]. Implications [938]. Implicit [1616]. Impostors [2567]. Imprimitive [2182]. Improve [1346, 1475, 2561, 1384, 2686, 1852]. Improved [1466, 1378, 1437, 1730, 1734, 3066, 2810, 965, 3149, 2819, 971, 1257, 169, 680, 1473, 2520, 2203, 1098]. Improvement [2979, 2708]. Improvements [1018, 2506]. Improving [582, 1940, 1970, 173, 1682, 1170, 1830, 2600]. Impulse [1874]. IN-Platforms [794]. INAP [791, 806, 807]. Inapproximability [2527]. INCA [2662]. Including [1383, 1836, 1830]. Inclusion [542]. Incomparable [3136]. Incompatible [2558]. Incomplete [2223, 113, 2319, 2234, 2266, 118]. Incompleteness [2224]. Incorporating [1860, 1752, 1626]. Increased [469]. Increasing [1479, 1950, 354]. Incremental [2015, 577, 1196, 788, 344, 345, 2543, 1929, 2363, 593, 122, 747]. Independence [1237]. Independencies [2254]. Independent [267, 1861, 2460, 2879, 2814, 3146, 1492, 2892]. Indeterminate [2948]. Index [296, 1235, 1505, 1551, 1614, 1672, 2083, 2153, 2205, 2445, 2508, 2565, 2762, 2896, 2924, 2763, 297, 582, 1922, 1588]. Indexing [571, 2329, 2019, 1589]. Indexings [3087]. Indian [510]. Indirect [2195]. Indistinguishable [469]. Individual [1156, 549]. Individuals [1018]. indoor [1564]. Induced [2217, 2806]. Inducing [1274, 1250, 437, 124, 2940, 1585]. Induction [2558, 2116, 2936, 1259, 2552, 2545, 1114, 1263, 126]. Inductive [2389, 1265, 2115, 121, 1940, 2925, 2943, 2834, 2940, 1017, 117, 2931, 1943, 555]. Industrial [1319, 1056, 1396, 64, 2358, 1886, 1708, 789, 2285, 1882, 228, 2709, 800, 801, 1318, 1921].

Industrially [2608].

Industrially-Oriented [2608]. Industry [938, 2608]. Inequalities [816]. Inequality [1752]. inertia [131]. Infarction [1119]. Inference [2316, 2557, 687, 2550, 1518, 1240, 2561, 570, 2971, 3007, 2554, 2556, 2858, 2042, 1190, 1933, 1621]. Inferential [1816, 1915]. Infinite [605, 2381, 775, 657, 2376, 3159, 2835, 3160, 2836, 655, 2362]. Infinite-State [775]. Inflation [3033]. Inflows [1894]. Influence [365, 2250]. Influences [1015]. InfoBus [1051, 1052]. Informal [108, 40]. Information [2021, 2664, 1756, 993, 991, 365, 2626, 2864, 534, 2223, 2596, 2615, 1103, 1104, 319, 2624, 2612, 2597, 1044, 2996, 839, 645, 2628, 1979, 1811, 2460, 1755, 2690, 2614, 1550, 257, 2290, 1754, 2620, 1247, 1769, 2610, 2145, 1188, 1237, 114, 1321, 2485, 1509, 2611, 831, 592, 2224, 3051, 1767, 537, 2999, 2581, 2288, 1523, 1051, 1052, 2266, 986, 2555, 2622, 356, 2621, 299, 1771, 2625, 2627, 1876, 1872, 568, 2287, 357, 594, 979, 2211, 118, 1508, 2146]. Information-Based [831]. Information-Filtering [2620]. Information-Rich [2614]. Information-Theoretic [1247]. informatique [521]. InfoSleuth [244, 2609]. Infrastructure [2433, 2217, 150, 244, 2484]. Inherent [3069]. Inheritance [2922, 1698, 86, 87, 552]. Inhibition [1955]. Inhibitors [2998]. Initial [972]. Initialization [1952]. Injection [2013]. Input [2823, 391, 2994, 2288]. Insights [2534]. Inspection [1888]. Inspectorate [373]. Inspired [1088, 1024, 1023]. Instance [2047, 1246, 1268]. Instance-Based [2047, 1246, 1268]. Instances [1687, 879, 293]. Instant [2887]. Instantiating [2124]. Instead [2369]. institutions [560]. Instructing [484]. instruction [741]. Integer [2812, 1741, 2156, 1500, 1732, 2670, 2671, 467, 2491]. Integers [2160, 614, 2188, 428]. Integral [2418, 2419, 2420, 2197]. integrality [2887, 2888]. Integrated [365, 784, 2673, 40, 1990, 371, 1071, 823, 856, 3028, 2340, 865, 2646, 356, 1963]. Integrating [1164, 2361, 706, 1601, 349, 2236, 1660, 819, 277, 2459, 1700, 314, 1600]. Integration [271, 1397, 39, 2004, 2715, 1814, 145, 1811, 1545, 497, 1538, 2973, 1533, 2482, 1848, 593, 2005]. Integrity [589, 1718, 1539, 1719]. Intelligence [1527, 1936, 1896, 1886, 323, 2249, 505, 881, 2211]. Intelligent [1808, 1965, 2306, 2480, 2043, 1825, 798, 2976, 359, 2660, 2455, 2667, 527, 1860, 2652, 1920, 1854, 790, 2454, 2616, 403, 806, 807, 2449, 1972, 804, 1138, 1304, 2600, 1457, 2780]. Intelligently [3028]. Intensive [598, 1342, 208]. Intentional [240]. Inter [2484]. Inter-domain [2484]. Interact [528]. Interacting [1850]. Interaction [398, 383, 1959, 2573, 1133, 481, 2417, 722]. Interactions [790, 153, 249, 2772]. Interactive [2865, 2866, 1044, 1768, 959, 1438, 468, 2598, 381, 2575, 259, 3028, 2599, 499, 525, 2867]. Interactivity [2334]. Interconnect [164]. Interconnected [181]. Interconnection [943, 1997, 1013, 2012]. Interconnections [1506]. Interconnects [911]. Interesting [1084]. Interestingness [1079]. Interface [1037, 261, 2043, 1378, 1959, 76, 2327, 2460, 77, 958, 395, 491, 332, 1375, 2595, 499, 220, 2102, 354, 357]. Interfaces [489, 358, 2457]. Interferometric [1416]. Interior [2963]. Interior-Lagrangian [2963]. Interleaved [2387]. Interlocking [1964]. Intermodular [744]. Internal [1278]. International [999, 2453]. Internationalized [520, 318]. Internet [307, 1378, 1371, 1381, 187, 2448, 1196, 2459, 2688, 574, 332, 804, 1363, 2465, 1051, 1447, 1052, 2338, 149, 1453, 572, 2339]. Internet-Based [2459, 1453]. Interoperability [1285, 2667, 869]. Interoperable [1455]. Interoperation [1548, 335]. Interpolation [1359]. Interpretable [1282, 1951]. Interpretation [1965, 3121, 1966, 982, 983, 984, 1816, 1871, 987, 986, 1919, 132]. Interpretations [3124, 2950]. Interpreted [2768]. Interpreter [537]. Interpreters [83]. Interprocedural [1391]. intersecting [476]. Intersection [1739]. Interval [1217, 814, 2851, 2893]. Intervals [2040]. Interworking [2480, 2482]. INTEX [2646]. Intractability [3138, 3139]. Intractable [3154]. Intranet [1320]. Intrinsic [2224]. Introducing [1061, 2473, 2708, 734]. Introduction [282, 1204, 968, 839, 936, 281, 2406, 1301, 1000, 1003, 1009, 2063]. Intuitionistic [1231]. Intuitive [404, 395]. Invalidate [2002, 2354]. Invalidate-Based [2002]. Invariant [3029, 3042, 3131, 2251, 1583, 2329, 2042, 1588, 1568]. Invariant-based [1583]. Invariants [2397]. Invention [1264, 2941, 2048]. Inverse [191, 2949]. Inversion [2860, 674, 675]. invertible [1479]. InVeSt [2397]. Investigating [371, 391, 2589]. Investigation [1785, 3027, 2983, 1390]. Invisible [354]. Invited [738, 2509, 2527, 2113, 2086, 2119, 2765, 2005, 2766, 2769, 2522]. IP [316, 2340, 175]. IP1 [2307]. IP2 [2308]. Irreducibility [647]. Irregular [1398, 1383, 176, 2008, 2007, 1992, 263]. Irregularity [2191]. Irrelevance [2863]. Isabelle [1060]. Isabelle/HOL [1060]. ISCM [319]. ISDN [1431]. ISO9001 [326]. ISO9001-Based [326]. Isomorphisms [1473]. Issue [31]. Issues [59, 480, 2049, 2505, 743, 2266, 2051, 378]. Issuing [1109]. Itemsets [1086]. Iterated [2958, 2777, 3116]. Iteration [604, 1913, 2865]. Iterative [2075, 2677, 1402, 1948, 1859, 1570]. ITOC [1755]. ITSE [1548]. IUT [791]. IV [3000, 34]. IX [88].

Japanese [401, 402]. JAS [1720]. JAVA [1438, 2906, 2913, 273, 3092, 272, 858, 1696, 743, 303, 33, 2674, 2458, 1379, 1695, 1445, 566, 1369, 331, 301, 1759, 2911, 1364]. Java-Based [1445, 1364]. javar [273]. JavaSoccer [1137]. Java^{T M} [2059]. JEROBOAM [408]. Job [2244, 1899]. Job-Shop [2244, 1899]. Jobs [3078]. John [1425]. Join [1428, 1546]. Joins [581]. Joint [410, 1652, 2773]. Jordan [2538]. Journals [542, 1047]. Joys [3100]. JP [2723]. JPMQ [1531]. Judging [1511]. Junction [2042]. Just [268, 693].

K4 [3080]. KADS [1839]. Kasuga [1159]. Kasuga-bitos [1159]. KBES [1889]. KDD [2312, 2236]. KeLP [1399]. Kerberos [2389]. Kernel [490, 1821, 1827, 2467]. Kernels [1255]. KEY [413, 155, 1175, 157, 2700, 2677, 150, 2486, 1468, 2173, 152, 313, 3089, 2670, 2671, 2484, 1483, 160, 2050, 2699, 2698, 1197, 1198, 1181, 1180, 2701, 2702, 2676, 2502, 1194, 2490, 2496, 2507]. Key-Dependencies [2701, 2702]. Key-Distribution [313]. Key-Recovery [160]. Keyboard [495]. Keyed [1478]. Keys [1179, 3094, 152]. Keyword [2763]. Keywords [2017]. Khepera [202, 203]. KHF [425]. Khoros [947, 949]. KN [1170]. Knapsack [1739, 2889]. Knockout [2080]. Knowledge [1073, 1075, 146, 2018, 1809, 1966, 1875, 1841, 1968, 2865, 469, 1856, 1814, 1967, 785, 786, 2029, 1854, 1813, 1109, 2259, 1969, 1538, 242, 1088, 2933, 2268, 1525, 1815, 1980, 1816, 388, 1116, 1836, 1871, 1093, 990, 2234, 2266, 208, 1241, 1937, 1984, 1805, 1944, 355, 1100, 1524]. Knowledge- [1856]. Knowledge-Based [1966, 1814, 2029, 1980, 990, 1937, 1984]. Knowledge-Intensive [208]. Knowledge-Level [1809]. Known [3027]. Koala [2421, 2422]. Kohonen [1824]. Korean [978, 516]. Kriging [1359]. Kripke [2768, 1228]. Kronos [2405]. Kruskal [2844]. KT4 [1206]. Kullback [1080]. Kumamoto [362]. Kumamoto-Kyoto-MIT [362]. Kyoto [362].

L0kappa [1788]. L0kappa-Equivalence [1788]. Labeled [2075]. labelled [558]. labelling [1640]. Labels [1840]. Labor [2956]. Laboratories [2585]. Labour [373]. Labyrinth [2604]. Ladder [762]. Lagrangian [2963]. Lambda [2098, 1791, 2946, 1795, 1277]. Lambda-Calculi [2098]. lambda-Calculus [1795]. lambda-Subsumption [2946]. LAN [2010, 2351]. Lanczos [1419]. Landcape [1980]. landmarks [1597]. Landscape [3017, 2989, 1980]. LANGAGE [2635]. Language [2874, 342, 108, 386, 1777, 2694, 2530, 346, 30, 351, 31, 17, 18, 272, 2550, 1291, 2218, 724, 379, 494, 2551, 434, 2900, 387, 23, 2057, 1877, 33, 1789, 32, 239, 691, 497, 713, 2637, 1212, 238, 306, 1930, 2915, 1540, 1055, 1541, 29, 2553, 2643, 66, 1020, 1400, 263, 1514, 1233, 227, 2051, 484, 492, 738, 1631, 747, 401, 402]. Language-Action [227]. language-oriented [747]. Languages [3107, 2800, 706, 2635, 2557, 2220, 3153, 2916, 2560, 464, 3167, 2917, 2726, 690, 2850, 2222, 3147, 714, 232, 347, 114, 473, 616, 617, 465, 619, 2820, 3169, 2541, 2554, 3143, 2645, 2555, 455, 1289, 3148, 26, 738, 734, 131]. LAPACK90 [899]. Large [2444, 991, 1356, 2601, 1410, 570, 2080, 1438, 1813, 2411, 578, 1242, 443, 763, 1689, 592, 148, 1534, 2973, 1958, 1355]. Large-Scale [991, 1410, 1242, 592, 2973]. Lateral [1955]. Lattice [1334, 1337, 1333, 1271, 2200, 1032, 2176, 1436]. Lattice-Based [1271]. Lattice-Boltzmann [1333]. Lattices [1026, 2489]. Launch [2664, 950]. Law [373]. Layer [1354, 1374, 1880]. Layered [35, 2138, 241, 1162, 2921, 1168, 2231]. Layering [1069]. Layers [1051, 1052, 2921].

Layout [519, 1847, 279, 981]. Layout-Based [981]. Layouts [1517]. Lazy [1270, 1386, 3134]. Lead [2186]. lean} [1207]. leanK [2093]. Learn [1855, 1265, 360, 1140]. Learned [2029, 703, 1349, 2580]. Learner [2945, 1821]. Learners [1269]. Learning [2046, 3003, 1152, 1164, 1962, 1251, 2795, 1270, 2244, 1264, 2776, 1940, 2546, 2550, 1851, 1266, 2559, 187, 1078, 1236, 2925, 2624, 1279, 2928, 2560, 2047, 1961, 2551, 2549, 2943, 2942, 839, 842, 1278, 2221, 2026, 123, 2548, 1925, 1108, 1258, 1254, 2941, 1246, 1268, 2540, 1938, 2027, 2044, 1939, 2938, 1238, 2235, 2038, 1248, 1283, 1281, 1526, 1143, 1245, 2020, 1155, 1900, 2543, 364, 2553, 1280, 1129, 2929, 1255, 2541, 2972, 1243, 2555, 2930, 1147, 2542, 2992, 1956, 837, 1932, 1256, 3014, 1277, 1943, 2033, 2780, 2948, 2547, 122, 119, 125]. Least [1173, 1635, 1563, 1596, 1634, 119]. least-squares [1596]. Leaves [1092]. Lecture [2509, 2527, 1682, 2522, 738]. LEDA [3096]. Legacy [1755, 1754, 55, 56, 2428]. Legal [1100]. Legged [1136]. Legibility [518]. Legislation [1444]. Leibler [1080]. Lemma [2089, 2818]. Lemmas [608]. Lempel [2861]. Length [3103, 2230, 3116, 1091, 2980, 2174]. Length- [2174]. Length-Preserving [3116]. Lengths [3029, 3036]. LEO [2103, 2104]. Less [1751]. Lessons [361, 2029, 377, 1349, 2580]. Let [197]. Letterforms [507]. Level [1711, 2694, 1809, 910, 2909, 1576, 474, 43, 1624, 350, 617, 2265, 717, 3028, 52, 1051, 1052, 251, 1453, 1405, 2713, 1612, 734, 1865, 24, 249]. Levels [1513, 2472, 2621]. Lexical [1241, 726]. LF [2117, 2118]. LFSC1 [2310]. Liberating [43]. Libraries [1711]. Library [1034, 1042, 1934, 2642, 906, 949]. Life [1533, 2579]. Life-Cycles [1533]. Lifecycle [2342]. Lift [1741]. Lifting [641, 815]. Light [311, 2689]. Light-Weight [2689]. Lighting [2578]. Like [2656, 3017, 2816, 2163, 2474, 2986, 1373]. Limbs [392, 393]. Limitations [2374]. Limited [2678, 2847, 2054, 2520]. Limited-Privilege [2678]. Limiting [2013]. Limits [2950]. Line [970, 2795, 1851, 2429, 390, 1886, 2514, 3149, 2794, 972, 672, 2790, 1928, 1932, 3152, 2789, 2791, 1560, 685, 2788, 646, 2796, 2886, 2793, 2792, 2491, 1671]. Line-Drawing [970]. Lineal [1833]. Linear [1738, 754, 1780, 1870, 1846, 1781, 1875, 1782, 452, 1218, 2524, 1402, 668, 432, 899, 816, 1873, 1468, 614, 1791, 1071, 2799, 2222, 2803, 438, 3176, 2175, 1796, 3082, 2086, 2682, 405, 2176, 2555, 1171, 2070, 1392, 650, 1484, 1357, 1646, 2880, 1671]. Linear-Time [2803, 3082]. Linearised [2690]. Linearly [3103]. Lines [983, 2416, 2428, 1566, 1572]. Linguistic [1822, 1111, 2265]. Link [1222, 1223, 542]. Linked [538]. Links [2331, 2925]. LISA [1777]. List [298, 1691, 2761]. Listeners [382]. Lists [2519]. Literate [778]. Literature [1045]. LITESET [2689]. Live [737, 787]. Living [368, 730]. Load [2789, 1344, 582, 1437, 1402, 1343]. Load-Balanced [1437]. Loadable [2144]. Loading [841]. Local [1527, 166, 1592, 1683, 2531, 753, 222, 751, 739, 3021, 1269, 1905, 463, 2858, 2647, 1872, 1594, 2242, 728, 1656, 1593, 1606, 1621]. Localisation [1668]. Locality [1344, 3131, 879, 263, 250]. localization [1597, 1667]. Localized [1418, 3032]. Locally [2835, 2836, 3105, 2554]. Location [2511, 927, 802, 1734, 459, 1520, 2464, 2997, 2661]. Locations [1751]. Locative [1517]. Locking [353]. Locomotor [1528]. log [1476]. Logarithms [2691, 2198, 2204, 1472]. Logic [762, 754, 2309, 1780, 1295, 1265, 1217, 1782, 686, 1218, 1957, 2925, 1787, 818, 158, 1071, 2799, 2082, 1793, 694, 2267, 65, 1225, 3168, 546, 466, 1816, 637, 2770, 2940, 3108, 2086, 3099, 3109, 2952, 2101, 1799, 695, 1800, 2117, 2118, 463, 126, 211, 2931, 1121, 1231, 1684, 2767, 1345, 2373, 2769, 2718, 117, 134]. Logic-Automaton [1793]. Logical [2310, 580, 690, 595, 596, 2235, 545, 465, 1796, 2826, 2017, 548, 563]. Logics [3106, 1213, 1204, 1208, 348, 432, 1220, 1221, 1211, 797, 1802, 733]. LogLinear [1238]. Logo [977]. Logs [600]. Long [3037]. Look [199, 3017]. Looking [1717]. Lookup [1869, 251]. Loop [273, 2032, 1405]. Loops [275, 260, 2131, 1339, 740, 731, 2132]. Loosely [328]. Loosely-Coupled [328]. Loss [1080]. Lossy [2380, 965]. LOTOS [1768]. Low [860, 156, 2348, 2678, 1382, 912, 2851, 2852, 3002, 1426, 1179, 1908, 3028, 52, 1183, 1405, 1612]. Low-Bandwidth [2852, 1426]. Low-Congested [2851]. Low-Cost [2678, 912]. Low-Entropy [1179]. Low-Level [3028, 52, 1405, 1612]. Low-Power [860]. Low-Risk [1908]. Lower [1738, 3127, 1781, 2886, 1740, 3172, 1465, 2492, 436, 3075]. Lowering [261]. Lowness [3068]. LP [1750]. LP-Based [1750]. LR [2632, 1119]. Lrc [747]. LRS1 [2309]. LSI [1183]. LTL [2372]. LTrL} [2811, 3123, 471]. Luby [1479]. Luby-Rackoff [1479]. Lyapunov [820]. Lyon [1315].

m [3061, 2079, 958, 715, 2102, 2568]. M-Module [958]. M-Net [2079]. M.A.D. [2898]. MAC [2504]. Mach [956]. Machine [105, 2795, 398, 267, 1580, 278, 106, 1939, 1974, 1748, 2020, 2385, 2211, 3092, 1695]. Machine-Checked [2385]. Machines [1749, 1385, 253, 38, 1993, 2636, 1254, 2425, 2875, 859, 1068, 192, 2648, 897, 2882, 2886, 1646]. Macromolecules [1893]. MACs [1480]. Made [1675, 3124, 703]. Made-to-Measure [1675]. Magic [2122]. Magnetic [2369, 1640]. Magnetohydrodynamics [1339]. Magnitude [1833, 1843]. Main [2369]. Maintain [3027]. Maintaining [540, 2023, 594]. Maintenance [598, 1851, 344, 1939, 1945, 593]. Major [1034, 1370]. Makanin [681]. Make [1808].

Makespan [1749]. Making [1845, 1829, 2249, 1133, 368, 2981, 2982, 1872, 2766, 1479]. Malicious [2054, 151, 2052]. Mall [365]. Management [1348, 73, 810, 589, 209, 927, 2456, 2714, 2660, 2455, 2727, 9, 939, 2673, 375, 584, 2652, 1287, 1454, 785, 786, 1329, 1757, 214, 2454, 1812, 1764, 223, 1188, 2478, 1316, 308, 2451, 875, 2484, 2657, 5, 2468, 2481, 945, 1051, 1052, 1369, 2326, 1771, 6, 895, 2716, 2713, 103, 1963, 177]. Manager [2453, 1294]. Managing [104, 1934, 2477]. Manipulation [404, 520, 2808, 522]. Manipulations [700]. Manipulators [1913, 1905, 1933]. Manual [2065]. Manufacturing [1953, 2067, 1910, 1973, 1971, 1972, 1887]. Many [2199, 2240, 1254]. Map [1026, 1110, 987, 986, 2501]. Maple [2635]. Mapping [209, 1099, 190, 946, 3, 546, 1390, 956]. Mappings [1430, 251, 560]. Maps [1931, 2626, 1521, 985, 2039, 486, 982, 983, 984, 3125, 2806, 971, 992, 559]. Maps/Graphics [971]. Marcus [2220]. Market [2443, 2613, 1454, 2659, 2470, 1086, 2459, 2469]. Market-Based [1454, 2659]. Marketing [2292]. Markets [2956, 1456]. Markov [386, 385, 1821, 673, 384]. Marriage [2481]. MAS [246]. MAS-CommonKADS [246]. Massaging [452]. Massive [1406, 3052, 1115]. Massively [1326]. Mastering [1971]. Match [1685, 1686]. Matches [3040]. Matching [338, 3048, 1711, 679, 2281, 3037, 1784, 1623, 1362, 1660, 3042, 1622, 629, 3034, 3035, 680, 3005, 1615, 3038, 672, 3036, 1613, 1577, 1584, 1616]. Mate [3001]. Material [1053]. Materialization [1775, 1085]. Materials [1403, 529]. Mathematical [215, 513, 863]. Mathematics [1200, 1201, 1212, 2134]. Matings [2125]. Matrices [1559, 2860, 429, 2227, 3043, 2842, 3178, 3084, 2226]. Matrix [2513]. Matroid [2813]. Mave [1862]. Max [3067, 1727, 2895, 3064]. MaxClique [290]. MaxE [288]. Maximal [653, 654, 1481, 1482]. Maximize [1677]. Maximizing [1443]. Maximum [2509, 1745, 3146, 575, 666, 1744, 3056, 2892, 1735]. MaxLinEq3 [288]. MaxLinEq3-2 [288]. May [1464, 2558]. Maze [1528]. Mazes [433]. McCulloch [1929]. McCulloch-Craik [1929]. MD4 [424]. MDx [2504]. MDx-family [2504]. Mean [886]. Mean-Field [886]. Meaning [2542]. Means [1851, 1417]. Measure [441, 1675, 2035, 201, 962, 1249, 1852]. Measurement [953]. Measurements [2710]. Measures [3091, 2932, 1826, 1086, 2290, 2288, 200, 61, 1644]. Measuring [539, 943, 955]. Mechanical [1199]. Mechanically [930]. Mechanising [2389]. Mechanism [1711, 339, 1131, 1766, 1765]. Mechanisms [1348, 2002, 2922, 2673, 100, 101, 2332, 718, 92, 1001, 229]. Media [583, 2325, 2324, 1381, 1196, 2351, 532, 533]. Mediated [2680, 2623, 1858]. Mediating [1960]. Medical [1319, 2274]. Medicine [1312]. Medieval [2604]. Medium [1851, 1357, 2999]. Medium-size [1357]. MeDoc [1034, 1037, 1035, 1049, 1042, 1041, 1039, 1048, 1040, 1038, 1036, 1047, 1050, 1043]. Medoid [1082]. Meeting [2594]. Meets [2580]. Meiko [1427]. Melding [2629]. Membership [920, 1824, 545, 921, 1424, 2643, 2148]. Memetic [3005]. Memorizing [2242]. Memory [1348, 2529, 1344, 1407, 2518, 2002, 1392, 2909, 526, 2955, 1327, 2068, 1510, 278, 3017, 2376, 178, 1393, 279, 1528, 1349, 1375, 2393, 99, 1386, 1326, 1511, 911, 2369, 759, 1353, 1405, 2011]. MEMS [939]. Menezes [161]. Menezes-Vanstone [161]. Meninas [2603]. Mental [1519, 1512]. Mereology [2285]. Merge [460]. Merger [2908]. Merging [2540]. Mergings [2558]. MESA [758]. Mesh [501, 1438, 3077, 902, 867].

Meshes [1422, 880, 2535, 1645]. Message [1966, 253, 20, 179, 2013, 1349, 905, 618, 1091, 2402, 1409, 1172, 1531, 2012, 924, 2498]. Message-Passing [253, 179, 1409]. Message-Passing/DSM [1409]. Messages [180, 332]. Messaging [1374]. Meta [889, 2915, 2641, 2696, 2117, 2118, 1537]. Meta-heuristics [889]. Meta-Logic [2117, 2118]. Meta-Normal [2641]. Meta-objects [2915]. Meta-Structure [1537]. Metabolically [3000]. Metabolically-Based [3000]. Metaclass [1775]. Metadata [335]. Metafont [511]. Metaheuristic [1029]. MetaJava [14, 100, 101]. Metaphor [364]. Metaphors [2663]. Metaserver [1377]. MetateM [237]. MÉTÉOR [1056]. Meteorological [1331]. Metering [2450, 1503]. Method [684, 1894, 1350, 2389, 1217, 1774, 1913, 2216, 791, 1766, 1196, 2221, 1085, 1890, 1752, 502, 1107, 2357, 2804, 1087, 1813, 964, 1071, 788, 1351, 2268, 1242, 2044, 976, 1064, 171, 172, 1135, 1281, 2199, 382, 1066, 1059, 1763, 505, 2203, 1428, 2198, 1069, 1602, 551, 1765, 2229, 1651, 1570]. Method-Driven [1766, 1765]. Methodist [1681]. Methodologies [224]. Methodology [1313, 2391, 946, 2474, 2019, 2990, 2451, 1916, 1964, 2992, 1823, 1947]. Methods [2304, 2305, 2311, 2313, 2314, 2315, 49, 1875, 74, 75, 978, 569, 2383, 2358, 2195, 1576, 1806, 1973, 799, 2284, 1358, 2973, 1057, 2930, 1987, 2295, 837, 400, 838]. Metric [2849, 294, 2150, 3041, 2149]. Metrical [2785]. Metrics [611, 1535]. MHEG [2322]. MHEG-5 [2322]. MIB [320]. Micro [202, 203]. Micro-Robot [202, 203]. MICROB [1144]. microcomputer [2500]. Microprocessors [2365]. Microtubule [2999]. Middleware [2667, 328]. Migrating [1458]. Migration [873]. Mill [1864]. Millenium [1805]. Million [205, 206]. MIMD [1385]. Mimetic [1275, 185]. Mimicking [2848]. Mimo [1922]. Min [3067, 1727]. Min-max [1727]. Min-Max-Boundary [3067]. Mind [1129, 3101]. Mines [2600]. Minima [2188]. Minimal [193, 1150, 2278, 3158, 2528, 1486, 3161, 2639, 2803, 2988, 627, 2640]. Minimising [1394]. Minimization [2831, 1283, 2651, 445]. Minimize [3049]. Minimizing [588, 1749, 1197]. Minimum [3103, 676, 853, 3085, 1730, 2857, 1743, 1746, 461, 677, 1091, 2813, 2856, 3084]. Minimum-Cost [1746]. Minimum-Weight [2813]. Mining [600, 1075, 1074, 2300, 1077, 1266, 1117, 1099, 1365, 577, 584, 1083, 1086, 1087, 2207, 1111, 2261, 1112, 1090, 1114, 1115, 1534, 3051, 1095, 1096, 2291, 1122, 1102, 2239, 2240, 2206]. Mining-Based [584]. Minor [610, 461, 259]. Minor-Closed [461]. Minsum [1747]. MIP [286, 287]. Mis [1631]. Misallocation [2985]. Misclassification [1253]. Missile [2990]. Missing [2925, 1094, 699]. Mistakes [703]. Mitochondrial [2275]. Mitologies [2604]. Mix [2680, 1493, 1492]. Mix-Mediated [2680]. mix-net [1492]. mix-servers [1492]. Mixed [1791, 667, 2595, 1733]. Mixes [1448]. Mixin [2921, 547]. Mixing [1325]. Mixture [1081]. Mixtures [1262]. Miya [1156]. MML [1080, 1081]. Mobidyc [1976]. Mobile [2055, 927, 809, 2700, 2466, 612, 1134, 2049, 2660, 2455, 2629, 1668, 1924, 2398, 2628, 2057, 2525, 936, 2054, 1925, 316, 1908, 315, 2897, 2485, 304, 2590, 1143, 322, 926, 147, 781, 1447, 1929, 1926, 1923, 2052, 2228, 2627, 2056, 2619, 2051, 2661, 1456]. Mobile-Agent [2057]. Mobility [808, 2665, 928, 2452, 317, 2463]. MOBY [725]. MOBY/PLC [725]. MOCHA [2400]. MOD [2818]. Modal [1204, 430, 1791, 1224, 496, 1226, 761]. Modal/Linear [1791]. Modalities [1220, 1221]. Mode [1467, 2680, 696]. Mode-Automata [696]. Model [768, 754, 2400, 270, 1809, 2370, 2899, 71, 1846, 1423, 2361, 1857, 2405, 706, 1601, 3000, 1677, 1218, 3033, 2550, 1203, 2367, 2559, 2216, 771, 2404, 319, 3132, 2432, 215, 3130, 2233, 2068, 597, 2091, 2089, 204, 1438, 3073, 2418, 882, 1890, 2846, 2371, 2359, 67, 1969, 2058, 2394, 2545, 78, 752, 2586, 516, 1891, 2368, 321, 350, 2364, 1986, 955, 937, 1795, 716, 2393, 1331, 875, 1328, 2421, 2060, 1821, 3133, 2119, 5, 2417, 945, 887, 1929, 2699, 1839, 61, 507, 4, 1605, 761, 2040, 1759, 756, 881, 1132, 1949, 1067, 6, 2697, 2419, 2568]. Model [2372, 1662, 2373, 2769, 2779, 2422, 2420, 1600, 2142, 1606, 1657, 2229, 1649, 563, 130, 1581, 1584]. Model-Based [706, 1890, 1891, 1662, 1606, 1657]. Model-Checking [2361, 2405, 2404, 716, 2393]. Model-Free [1601, 1600]. Model-Theoretic [2119, 563]. Modeled [2376]. Modelica [2900]. Modeling [2957, 2015, 1442, 62, 35, 1330, 1270, 45, 766, 765, 2039, 37, 1924, 1820, 1676, 40, 2900, 530, 531, 959, 1976, 70, 2720, 1813, 2687, 34, 2255, 43, 578, 306, 826, 1815, 308, 66, 485, 1932, 63, 2721, 759, 854, 2012, 1453, 2282]. Modelling [2725, 2289, 1106, 1081, 1028, 2907, 2068, 1427, 850, 1979, 2460, 711, 237, 2162, 131, 1887, 778, 2337, 1400, 2274, 1626, 1120, 1578]. Models [1833, 600, 386, 3059, 2694, 2387, 843, 49, 36, 2601, 2043, 1250, 2065, 1647, 38, 2516, 784, 1843, 1653, 1520, 1486, 1293, 1450, 850, 1284, 690, 385, 344, 1519, 989, 223, 1111, 2768, 1238, 2066, 2393, 1228, 844, 859, 621, 718, 384, 2469, 1975, 1043, 2718, 1648]. Modems [2345]. Modern [1712, 1349, 2371]. Modes [696, 421]. Modification [1954, 2692]. Modified [1175, 1176]. Modifiers [2028]. Modifying [1128]. Modula [748]. Modula-3 [748]. Modular [442, 3171, 2412, 3123, 674, 1714, 557, 1174, 752, 1985, 2829, 455, 675, 2502, 1477, 1471]. Modularity [2400, 639]. Modularization [3094]. Modulation [1159]. Module [205, 206, 958, 332, 46, 47]. Module/Package [46, 47]. Modules [1936, 547]. Modulo [275, 631, 624, 740]. Molecular [2871, 2302, 2219, 1408, 912, 1410, 881, 1949, 2537]. Moments [2296, 1588, 2335]. MON [413]. Mona [2399, 1793]. Monadic [479, 3168, 637, 1800, 1801, 1803]. Monads [2130]. Mongolian [520]. Monitor [861]. Monitoring [1883, 3023, 1884, 967, 1917, 327]. Monkey [413]. Monoids [625]. Monotone [3053, 3154, 2151]. Monotonic [1249, 2770]. Monotonicity [337]. Monte [1335, 1414, 1818]. Montgomery [2163]. Montgomery-Like [2163]. Morphological [970, 971, 402, 1650]. Morphology [2938]. mosaicing [1621]. Mosaics [500]. Most [3127, 2981, 2982, 907]. Motifs [2997, 682]. Motion [1650, 2516, 1513, 1575, 1906, 1903, 693, 1638, 1904, 1635, 388, 1907, 1905, 1157, 1569, 1574, 741, 1572, 1651, 1634, 1598, 1633, 1584]. Motions [144]. Motor [1164, 1839]. Motorways [1771]. Mouse [2706, 1560]. Movement [387, 380, 396, 397]. Movements [387]. Movie [2575]. Moving [2515, 1296, 1805, 1625, 1609]. MPEG [1440]. MPEG-2 [1440]. MPEG2 [2331]. MPI [914, 1341, 1427, 170, 961, 1374, 178]. MRAI [1407]. MRD [414]. Muir [1425]. Muller [2639]. Multi [2021, 2664, 2712, 3087, 1073, 1488, 1978, 2959, 2356, 1862, 150, 1365, 1597, 1990, 2666, 1976, 247, 2618, 2655, 1902, 2846, 540, 1191, 1669, 1168, 2939, 1757, 2446, 496, 2008, 3043, 224, 2637, 1555, 2653, 2333, 1184, 1133, 24, 1599, 249, 165, 1767, 2587, 566, 2351, 2684, 1499, 1162, 229, 2533, 2585, 2765, 234, 2878, 1405, 2231, 2780, 734, 856]. Multi-Agent [2664, 1862, 2666, 247, 2618, 2655, 2939, 224, 2653, 2587, 229, 2585, 2765, 234, 2021, 2959, 2846, 1133, 2780]. Multi-Agents [1978, 1976]. Multi-Application [1184].

Multi-block [1405]. Multi-class [1990, 1757]. Multi-database [1767]. Multi-Dimensional [1365, 3043, 3087]. Multi-Directional [1902]. Multi-domain [150, 2446]. Multi-drop [2008]. Multi-environment [2333]. Multi-FPGA [856]. Multi-Grain [2712]. Multi-Layered [1162, 1168, 2231]. Multi-level [24, 249, 734]. Multi-Media [2351]. Multi-Modal [496]. Multi-party [1488]. Multi-Processor [2356]. Multi-Receiver [1499]. Multi-scale [1599]. Multi-senders [1191, 2684]. Multi-Stage [2878]. Multi-step [1597]. Multi-tape [2637]. Multi-Threaded [566, 2533]. Multi-User [165, 1073]. Multi-variant [540]. Multi-view [1669]. Multiagent [1912, 2614, 246, 1864, 232, 1865, 2772, 1810, 2588, 1130]. Multiattribute [2215]. Multicast [1382, 2526, 964, 3076, 313, 876]. Multicasting [2008, 1992]. Multichannel [1625]. Multiclass [1258]. Multicommodity [1746]. Multicomputer [1997]. Multicriteria [2215, 1922]. Multicuts [1731]. Multidatabase [1090]. Multidestination [2008, 1992]. Multidimensional [1356, 580, 576]. Multifunctions [2225]. Multigranularity [353]. Multigroups [1192]. Multihead [2876]. Multilanguages [2221]. Multilateral [1445]. Multilective [3172]. Multilevel [1293, 1438, 2703]. Multilingual [535, 869, 274, 2938]. Multimedia [1139, 2325, 809, 1959, 526, 494, 2328, 2342, 1768, 959, 524, 2332, 2323, 2335, 2333, 2349, 481, 491, 1390, 2451, 804, 309, 2320, 2326, 2338, 1537, 2321, 525]. Multimodal [1213, 490, 498, 480, 482, 486, 487, 488, 493, 489, 391, 322, 1618]. Multimodeling [1919]. Multiparadigm [869]. Multipath [2993]. Multiperspective [2898]. Multiple [1045, 1697, 1601, 2809, 2377, 2942, 586, 2057, 508, 1552, 989, 1399, 955, 1116, 358, 585, 871, 2585, 3156, 587, 1549, 421, 2775, 1617, 1600, 1661, 1577, 1570]. Multiple-Alphabet [508]. Multiple-Computer [358]. Multiple-Language [2057]. Multiple-Rate [955]. Multiplexing [2011]. Multiplication [2842, 3178]. Multiprecision [2178]. Multiprocessor [885, 1347, 886, 940]. Multiprocessors [2002, 586, 2393, 587, 1353, 924]. Multiprogramming [937]. Multirate [3076]. multiresolution [1627]. multisignature [2498]. Multistrategy [1245]. Multithreaded [258]. Multivariate [1113, 576, 1471]. Multiversion [342]. Multiview [2319]. Multiway [676, 677, 2373]. Muon [1982]. Murcursive [2369]. MUSE [1768]. Museum [1960, 2323]. Music [2595]. Musical [2559, 2341]. musicale [521]. MUSIST [2336]. Must [1808]. Mutation [3009, 2979, 1016, 2989, 196, 3020]. Mutations [3022]. MUTEX [2073]. MUTEX-Solutions [2073]. Mutual [451, 779]. Myenteric [1986]. Myocardial [1119]. Myrinet [180, 165, 908]. Myths [726].

n02 [1258]. n02-classifier [1258]. Naive [1237, 1830, 1261]. NAMD [274]. Name [2873]. Name-Passing [2873]. Names [992]. Nano [270, 875]. Nano-Threads [270, 875]. Narratives [2575, 2604]. Narrow [2335]. Narrow-Bandwidth [2335]. NAS [1353]. Nationally [2029]. Native [2339]. Natural [108, 387, 1930, 1602, 1631, 1644]. Nature [249]. Navigation [2336, 2626, 2791, 2601, 1044, 1832, 711, 1525, 1138, 2228, 1605, 1524, 1564]. Navigator [2572]. NC [470, 907]. NCSTRL [1045]. Nearest [3057, 2514, 1535, 1941]. Nearest-Foreign-Neighbor [3057]. Nearest-Neighbour [1535]. Nearly [434]. NEC [1355]. Necessity [2862]. Need [764, 1561]. Negation [3127, 2770, 2265]. Negative [3128, 645]. Negligibility [1840, 2208]. Negotiation [2666, 2665, 2617, 236, 128]. Negotiations [2618]. Neighbor [3057]. Neighborhood [1079, 3064]. Neighborhood-Based [1079]. Neighbors [2514]. Neighbour [1535, 1941]. Nerve [2582]. Nervous [1986]. Nested [2135, 344, 345]. Net [2071, 2825, 2076, 785, 615, 2078, 941, 2372, 1492, 1320, 2079, 2004, 2005]. NetAcademy [786]. Netherlands [373]. Nets [607, 1855, 1302, 1924, 2807, 2067, 757, 2081, 1628, 2069, 1824, 948, 717, 2063, 1453, 549]. Network [1380, 975, 1954, 3059, 168, 2480, 2658, 2456, 809, 166, 813, 300, 2455, 1828, 9, 2327, 2230, 2673, 817, 3016, 2549, 2652, 2666, 2655, 1849, 958, 3078, 3074, 967, 2793, 2454, 2032, 1955, 1247, 2792, 2449, 2333, 3007, 1986, 167, 169, 2662, 2020, 2657, 5, 2468, 2477, 887, 1415, 189, 2338, 1880, 2469, 846, 874, 6, 3075, 836, 317, 838, 3014, 1892, 2042, 2779, 127]. Network-Aware [874]. Network-Based [2469]. Networked [1768, 959]. Networking [2021, 2664, 1368, 784, 1384, 908, 302]. Networks [770, 922, 1953, 2654, 866, 808, 2344, 2656, 2523, 1825, 798, 176, 2848, 1989, 2852, 2003, 2218, 2526, 2036, 805, 1023, 977, 1083, 2847, 842, 2659, 7, 2076, 787, 427, 880, 2446, 790, 2008, 3076, 2268, 943, 2478, 2044, 2013, 845, 163, 2066, 489, 1842, 1997, 1680, 1283, 1942, 1767, 2009, 2999, 1867, 1024, 718, 1885, 2972, 904, 1011, 1013, 928, 2228, 2007, 1992, 3015, 2006, 840, 898, 2296, 2479, 1950, 1872, 2012, 1282, 1951, 2661, 1982, 2269, 177, 2339, 1666, 2891, 2270]. Neural [1954, 3003, 1953, 1855, 866, 398, 1828, 205, 2036, 2230, 817, 3016, 1023, 2549, 977, 1083, 206, 1849, 1864, 880, 2032, 2268, 3095, 845, 1986, 1283, 1867, 1024, 1885, 887, 1927, 847, 2228, 3015, 1880, 840, 846, 2296, 836, 1932, 838, 1282, 1951, 1982, 3014, 2269, 1892, 2270].

Neural-Network [817]. Neuro [1933]. Neuro-Fuzzy [1933]. Neurocomputing [1319]. Neurocybernetics [1895]. Neuroidal [2855]. Neuron [205, 206, 204]. Neuronal [843]. Neurons [1950]. Newspapers [1898, 1140]. Newton [2087]. NEXP [3136]. NEXPTIME [3150]. Next [1074, 1214, 205, 2030, 206, 2001, 1805]. NFA [3115]. NICE [1317]. Niching [3007]. Nightwatch [202, 203]. Niken [1156]. Nile [1894]. Ninf [1377]. NL [2545]. NLR [1355]. No [201, 3080]. Node [3082]. Node-Search [3082]. Nodes [1361, 474, 952]. Noise [2273]. Noisy [1241]. Nomadic [1547]. Non [776, 777, 1846, 1392, 476, 650, 3123, 659, 1357, 1747, 284, 1482, 1019, 2933, 402, 290, 2770, 2122, 162, 1434, 289, 1799, 2867, 467, 1171, 3075, 142, 594, 1479, 728, 1481, 2866]. Non-approximability [1747, 284, 290, 289]. Non-blocking [3075]. Non-Boolean [467]. Non-commutative [1799]. Non-contact [402]. Non-destructive [1019]. Non-determinate [2933]. Non-elementary [3123]. Non-Euclidean [142]. Non-fixed [776, 777]. Non-homogeneous [659]. Non-Horn [2122]. Non-interactive-SZK [2867, 2866]. Non-intersecting [476]. non-invertible [1479]. Non-Linear [1846, 1392, 650, 1357]. non-local [728]. Non-maximal [1482, 1481]. Non-monotonic [2770]. Non-perfect [162]. Non-premixed [1434]. Non-temporal [594]. Nondeterministic [1796, 2266, 3152]. Nonelementary [2280]. Noninteger [1724]. Nonlinear [2995, 2216, 1881, 830, 1882, 1988, 832, 1495]. Nonlinearity [2686]. Nonrepudiable [2707]. Nonstandard [2087]. Nonstochastic [3107]. Nonterminal [2222]. Normal [610, 2942, 634, 2641, 2126, 463, 116]. Normalization [1582, 1242, 1804, 633]. Norms [1735]. Norreed [2382]. notation [521]. Notations [39]. Note [1255, 2507].

Notes [2966]. nother [450]. Nothing [586, 587]. Noun [2028]. Nouns [1515]. Novel [2976, 1384]. NP [3103, 2800, 3104, 2865, 3137, 1750]. NP-Completeness [3104]. NP-Hard [3103, 1750]. NT [914, 2004, 2005]. NTI [958]. NTRU [2173]. Nuclear [1884]. Null [2643]. Null-Free [2643]. Number [2538, 1173, 2183, 660, 3078, 3074, 416, 474, 666, 2162, 2163, 457, 1004, 2185, 2165, 1982, 2184, 1492, 1463, 1472]. numbering [734]. Numbers [477, 648, 2178, 2191, 2193, 2161, 1840, 3173]. Numeric [2979]. Numerical [1406, 2831, 1904, 1417, 1987]. Numerous [2567]. Nuprl [2390]. Nyberg [1195]. Nyberg-Rueppel [1195].

O [1383, 1385]. OA&D [35]. OBDD [445]. OBDD-Minimization [445]. OBDDs [447, 446, 3154, 3156]. Object [105, 1756, 1711, 89, 49, 589, 1853, 2899, 36, 1452, 71, 2918, 2136, 85, 45, 226, 50, 10, 54, 37, 1957, 2909, 495, 2922, 590, 404, 40, 2900, 88, 2923, 1284, 7, 106, 711, 70, 882, 2717, 690, 67, 1610, 2687, 34, 1642, 2901, 43, 343, 394, 714, 223, 694, 2238, 2919, 2974, 1064, 65, 304, 308, 955, 2019, 167, 937, 11, 91, 1704, 305, 948, 1541, 1827, 1533, 1598, 66, 695, 61, 98, 507, 102, 1400, 995, 1537, 352, 1705, 1706, 1366, 1546, 1367, 93, 2063, 2911, 86, 87, 103, 1517, 1, 1659, 2137, 1656]. object [556, 1657, 1666, 1663, 1661, 744]. Object-Based [1064]. Object-Centered [308]. Object-Centred [1853]. Object-Orientation [2136, 98, 2137]. Object-Oriented [105, 1711, 89, 49, 589, 36, 71, 85, 45, 50, 10, 54, 37, 590, 40, 2900, 88, 1284, 7, 106, 711, 70, 690, 34, 2901, 43, 714, 694, 65, 955, 167, 11, 91, 305, 66, 695, 507, 102, 1400, 1537, 352, 1366, 1546, 1367, 2911, 103, 1, 744]. Objectives [1053, 225]. Objects [2281, 1458, 490, 686, 94, 1832, 1438, 2059, 107, 1751, 2897, 99, 3108, 2319, 2952, 697, 2696, 92, 871, 1296, 2214, 2063, 2287, 552, 1609, 1664, 2915, 1646, 1662, 1626]. ObjectSpace [303]. Oblivious [1487, 3117, 1006, 1010]. Observation [2859, 2838, 1533, 1665]. Observational [705, 559]. Observations [609, 1679, 783, 130]. Obstacle [1923]. Obtain [1837]. Obtaining [108, 2498]. OCB [590]. Occam [1251]. Occlusions [1566, 1616]. Ocean [1328]. Octic [2182]. Odd [3080]. Odds [2798]. ODMG [2899, 1541]. OF-fields [1620]. Off [810, 685, 1675, 2355, 3149, 3152, 2491]. Off-Line [3149, 3152, 685, 2491]. Off-Road [2355]. Off-the-Peg [1675]. OFFER [1452]. Offering [2452]. Offerings [1050]. Office [373, 325, 370]. Offline [1005]. offs [2009]. Offset [3029]. OI [2267]. OI-Resolution [2267]. Oja [2256]. Okamoto [2507]. Okamoto-Tanaka [2507]. OLAP [578, 579, 576]. Old [2534]. Omega [477, 440, 2649]. omega-words [440]. Omni [1149]. Omni-directional [1149]. Omnidirectional [1148]. OMS [1759]. OMS/Java [1759]. OMT [3]. On-Line [2795, 2794, 672, 2790, 1928, 1932, 2789, 2791, 2788, 646, 2796, 2793, 2792, 2886]. On-the-Fly [2380, 2370, 1388]. One [754, 2883, 3090, 3174, 636, 424, 3111, 3112, 3149, 2222, 2540, 3168, 2257, 1230, 3152, 1405, 1484, 2503, 564, 1485, 1662]. One-Pass [3111, 3112, 1230]. One-Tape [3149]. one-tiered [564]. One-Time [3090]. One-Way [3174, 424, 2503, 1485]. one-wayness [1484]. Ones [1737, 2817]. Online [785, 786, 427, 542, 1047, 2533]. Only [1264, 1786, 279]. Ontological [1549]. Ontology [35]. Ontology-Based [35]. OO [59, 39, 8]. OO-Notations [39]. OODBMS [1759]. OOM [1889]. OOM-KBES [1889]. Open [805, 2327, 755, 247, 2679, 3125, 559, 2806, 152, 244, 378, 2668, 1571, 755]. OPEN/CAESAR [755]. OPENR [1136].

Operating [105, 104, 13, 106, 100, 101, 1184, 98, 102, 3020, 103]. Operating-System [100, 101]. Operation [277, 163, 437, 421, 2600]. Operational [2826]. Operations [970, 582, 3140, 453, 971, 2225, 2213, 876, 2644, 871]. operative [1435, 367]. Operator [670, 2267, 2250, 2024]. Operators [2959, 183, 201, 3019, 2978, 2248]. opolizing} [413]. Opportunities [1692, 738]. Optic [1652, 1632, 1639]. Optical [1838, 166, 2293, 2847, 1011, 1633, 177, 2011, 2891, 1619]. Optikron [2404]. Optimal [2447, 271, 571, 3085, 2283, 913, 892, 3055, 462, 2286, 434, 2514, 2525, 1667, 1246, 952, 1249, 2827, 439, 466, 290, 1911, 1562, 3082, 1347, 3020, 3081, 3063, 740, 2964]. Optimality [2385]. Optimisation [1681, 1394, 1864, 2449, 906, 1921]. Optimise [887]. Optimised [1209, 1171]. Optimistic [1504, 173]. Optimization [3012, 2344, 2970, 2331, 1856, 3013, 650, 1881, 2793, 879, 294, 830, 2990, 2963, 831, 3019, 2903, 255, 2978, 3011, 2980, 258, 3083, 293, 566]. Optimization-based [1856]. Optimizations [2650]. Optimize [277, 225]. Optimized [1324]. Optimizer [1922, 266]. Optimizers [593]. Optimizing [1441, 3154]. Optimum [1470]. Optional [1703]. Options [1288, 1280]. Oracle [3041]. Oracles [2020]. Oral [2550, 485]. ORB [303, 2467]. ORBs [12]. Order [2386, 1833, 2802, 2094, 2103, 2095, 2104, 2932, 2918, 2800, 1494, 2691, 1784, 1266, 643, 1843, 3131, 1789, 773, 774, 479, 626, 3168, 437, 406, 637, 195, 2148, 2973, 2096, 2353, 2952, 2929, 2101, 1800, 463, 1802, 1170, 2363, 1803, 2647, 1950, 2845, 2373, 561, 119]. Order-Invariant [3131]. order-sorted [561]. Ordered [448, 2988, 2374, 3039]. Ordering [641, 2108, 2109, 447, 1240, 637, 581, 1273, 2845, 1650]. Orderings [2809, 3155]. Orders [1482, 3156, 1481].

Ordinal [1076, 1804]. Ordinals [652]. ordination [209, 2774]. Organisation [1423]. Organisational [209, 349]. Organization [69, 361, 1767]. Organizational [59]. Organizations [365, 208]. Organizing [2783, 2596, 2597, 374, 1110, 1942]. Orientation [2136, 1528, 98, 1530, 2137, 1560]. Oriented [105, 1756, 1711, 89, 49, 1187, 589, 36, 71, 85, 72, 45, 226, 50, 10, 54, 37, 495, 2922, 1432, 590, 1236, 319, 27, 40, 2900, 88, 2923, 79, 1284, 7, 106, 711, 70, 2717, 690, 34, 2901, 43, 1769, 1764, 714, 694, 65, 97, 955, 167, 11, 91, 1704, 305, 2608, 1827, 66, 695, 507, 82, 102, 1400, 1537, 352, 1705, 1706, 1366, 1546, 1367, 2911, 28, 103, 1186, 1, 2476, 747, 1598, 744]. Origin [622]. Orlin [1743]. Ornamental [513]. Orthogonal [2906, 2913]. orthogonalizing [1655]. Osaka [1147]. OSAM [2465]. Oscillator [1423]. OSMOSIS [69]. OST [7]. OSTIA [2544]. Other [1691, 783]. OTI [96]. Out-of-Order [2363]. outconnected [2885]. Outperform [2817]. Output [1269, 2536]. Output-Sensitive [2536]. Overhead [265]. Overlap [254]. overlapped [1633]. Overlapping [3033, 1092]. Overload [1885]. Overloading [189]. Overview [2488, 910, 849, 974, 1125, 2608, 2489, 836, 117]. Own [360]. OXSIM [1901]. OZ [305].

P [607, 1997, 1234]. P-SETHEO [1234]. P/T [607]. P610 [2451]. PACE [868]. Pacific [300]. Package [2635, 899, 1421, 46, 47, 2649]. Packages [2723]. Packet [671]. Packing [1721, 2788, 1740, 2512, 1732]. Packings [3105]. Padding [1802]. Padnon [2633]. Page [2623, 2545]. Pages [1106]. Paging [2786, 2464, 2784, 2533]. Pairs [639, 1240, 640, 623, 907]. Pairwise [1257, 1659]. PANAMA [409]. Panel [224]. Panoramic [1565]. Paper [528, 527, 782, 2765, 2766, 2769]. PAPERS [268]. Paradigm [1812, 1395, 1097]. Paradigms [1169, 2909, 544, 2961]. paradox [2208]. Paraflow [1372]. Parallax [1670, 1669]. Parallel [1350, 3059, 676, 853, 1335, 1324, 1332, 1341, 1356, 1414, 1407, 1406, 272, 1383, 2156, 267, 2524, 1077, 1402, 2324, 1381, 1432, 1387, 2218, 1325, 900, 1336, 459, 1327, 1362, 1401, 1433, 884, 1337, 586, 1412, 889, 757, 286, 287, 461, 961, 677, 2969, 277, 1389, 274, 1333, 1393, 1361, 1339, 179, 949, 1379, 667, 256, 870, 473, 1349, 1384, 1420, 953, 171, 172, 264, 167, 1424, 1748, 169, 894, 3035, 1331, 906, 1358, 878, 1328, 1326, 1434, 1409, 1404, 893, 252, 901, 896, 887, 460, 255, 254, 883, 467, 1096, 1416, 1429, 2271, 587, 888]. Parallel [1428, 1366, 1546, 2842, 3178, 263, 895, 897, 951, 2177, 1405, 175, 3063, 3039, 2286]. Parallelisation [1440]. Parallelism [270, 271, 903, 890, 1399, 1395, 1234, 736]. Parallelizable [891]. Parallelization [271, 273, 3165, 668, 965, 260, 934, 1340, 1353, 1405]. Parallelized [265]. Parallelizing [1398, 2383]. Parameter [502, 2653, 2978, 3011, 1932, 2229]. Parameterised [552]. Parameterized [931, 2392]. Parameters [3053, 1849, 2506, 1247, 510]. Parametric [650, 1985, 656]. Parametrization [143]. Parasite [2960, 1276]. ParaStation [910]. parchments [563, 565]. Parentheses [2809]. Parity [669, 1185, 1018, 447]. Parity-OBDDs [447]. Parking [1981]. Parser [1878]. Parsers [2632]. Parsing [848, 492, 727]. Part [1239, 1916, 1917, 2981, 2982]. Part-of-Speech [1239]. Partial [2386, 644, 2802, 2529, 2264, 685, 611, 2116, 643, 462, 942, 773, 774, 2806, 114, 2915, 1116, 3005, 2353, 963, 1428]. Partial-Congruence [2806]. Partial-Order [2802]. Partially [453, 2988, 857]. Partially-Reconfigurable-Field-Programmable [857]. Particle [3012, 3013, 3011]. Particles [3010]. Partioning [857]. Partition [429, 824, 2329, 2994]. Partitioning [889, 1422, 1361, 1819, 856, 2651]. Partitions [2038, 2260]. Partner [2956]. Party [3090, 1488]. Pass [3111, 3112, 1230]. Passing [2073, 253, 20, 1789, 179, 1349, 2873, 905]. Passing/DSM [1409]. Passive [1157]. Past [3098, 2854]. Paste [2220]. Patch [1326]. Path [853, 2530, 1901, 3165, 182, 2655, 1902, 2008, 1732, 1908, 591, 1928, 2138]. Path-Based [2008]. Paths [661, 2510, 428, 907, 973]. Pattern [1025, 3048, 679, 2097, 2564, 2563, 60, 20, 2984, 2251, 1952, 680, 2541, 3014, 1561]. Patterns [19, 1295, 31, 17, 18, 1049, 3037, 154, 21, 23, 32, 22, 178, 179, 2207, 2902, 1030, 24, 513, 368, 1096, 2627, 2206]. Pawlak [2248]. Paying [382]. Payment [1448, 1196, 1447]. Payments [2688]. PC [909]. PC-based [909]. PCI [164, 181]. PCP [285]. PCP-Theorem [285]. PCRC [262]. PCRC-based [262]. PDE [1651]. PDF [542]. PDGP [1023, 1024]. PDGP-inspired [1023, 1024]. PDM [2715]. Peanuts [1849]. Pedagogical [1969, 2268]. Peer [2460, 223]. Peer-to-Peer [2460, 223]. Peg [1675]. Penalties [186]. Penalty [2963]. Pentium [909]. People [1959, 1611, 357]. Perceived [2344]. Perception [1513, 1510]. Perceptual [2273, 1591]. Perceptually [518]. Perceptually-Tuned [518]. perfect [162]. Performance [1380, 3012, 1319, 1324, 1396, 582, 927, 1397, 1371, 1989, 300, 913, 1385, 199, 996, 2171, 1336, 1394, 899, 817, 1427, 1376, 2797, 1388, 1359, 1439, 173, 2345, 2478, 1384, 1874, 953, 1997, 181, 2001, 259, 1370, 2009, 1363, 998, 908, 926, 1852, 2574, 384, 1369, 995, 1318, 941, 1830, 954, 1366, 1546, 1367, 1098, 2562, 854, 997, 2024, 2172, 127, 2499]. Performances [590, 200]. Periodic [3029, 2835, 2836, 1911]. Periodicity [627].

Peripheral [806, 807]. Permutation [3085, 475, 451]. permutations [1484]. Persistence [2913]. Persistent [2519, 453, 1531]. Perso [520]. Perso-Arabic [520]. Personal [2021, 2664, 912, 2665, 322, 2463]. Personalized [1103]. Personalizing [1960]. Perspective [74, 75, 1567, 300, 1203, 207, 2210, 1148, 227, 1570]. perturbations [1655]. Petri [1924, 2067, 2071, 757, 2076, 2072, 2069, 2078, 1827, 2063, 1453]. Phase [1126, 1416]. Phased [2970]. Phases [380]. PhD [88]. Phenomena [1245]. phi [2369]. Philosophical [1200, 1201]. Philosophy [3012]. Phone [781]. photometry [1656]. Photon [1354]. Physical [1126, 852, 1705, 1706]. Physics [1982]. pictorial [1583]. Picture [616, 3169]. Picture-Languages [3169]. Pictures [41, 3168, 1801]. Pieces [1888]. Piecewise [616]. Pigeonhole [1783]. PINCEL [1320]. PINCEL-NET [1320]. Pincer [575]. PINROB [1708]. Pipe [1847]. Pipeline [275]. Pipelined [750, 2365]. Pipelines [946]. Piping [1847]. Pizza [1233]. PKI [2679]. Place [376, 1526, 355]. Placement [693, 1367]. Places [2199]. Plan [487, 488, 232, 2775]. Planar [2809, 3153, 1670, 668, 2528, 459, 476, 1438, 1986, 1728, 1557, 1556]. planes [1831, 1669]. Planet [304, 144]. Planner [1853]. Planning [3025, 2016, 1307, 1901, 1218, 1936, 2516, 1906, 1902, 1806, 1168, 1847, 1903, 1908, 715, 1904, 1827, 2101, 1905, 1817, 928, 1928, 1157]. Plans [976]. Plant [1988]. Plants [1884, 1985]. Plasma [1336, 1337, 1339]. Platform [3092, 1413, 180, 100, 101, 2100, 2409, 2410, 1445, 2465, 2587, 2588, 2482, 2326]. Platforms [794, 2466, 1404, 1352]. Plausible [3003]. Play [2658, 1156, 1149]. Playing [1134, 1159, 1129, 1128, 1267, 1167]. PLC [725]. Plexus [1986]. PLTL [1230].

Plug [2658]. Plug-And-Play [2658]. PMLs [64]. PMMLA [851]. Point [809, 2395, 459, 1080, 954, 1292, 1602, 1617, 1556, 1560, 1597, 1606]. Point-Based [1292]. Pointer [698]. pointers [734]. Pointing [394]. Points [1756, 2429, 143, 2196, 2251, 751, 2803, 2197, 1572, 1463]. Poisson [1341]. Poker [2031]. Polar [2347]. Policies [1343]. Policing [1825]. Policy [373, 2484, 2385]. Pollard [2198]. Polyatomic [3104]. Polygonal [661, 3058]. Polygons [2810]. Polyhedra [1739, 1724]. Polyhedral [1736, 2510, 1735, 1751]. Polylog [2856]. Polymorphic [853, 3120, 3122]. Polynomial [3102, 2857, 2828, 659, 294, 647, 2543, 3117, 1255, 445, 1471, 2889]. Polynomial-Time [2828]. Polynomials [2690, 2162, 2169, 1473]. Polysynaptic [1986]. Polytypic [2143, 2148]. Polyvalent [1816]. Polyvariant [685]. Pool [1978, 2680]. Pool-Mode [2680]. Pop [2577]. Pop-Out [2577]. Populated [2654]. Population [1018]. Populations [1976]. Portable [1708, 745, 1709]. Porting [1411, 912, 1695]. Pose [2982]. Posets [473, 2388]. Position [1296]. Positions [1152]. Positive [3136, 1264, 2166, 645, 655, 2541, 2555, 2174]. Positive/Negative [645]. Possible [2212, 2769, 2895]. Post [2371, 221]. Post-Client [221]. Post-Client/Server [221]. Post-modern [2371]. Postman [1733]. Potential [1847, 1903]. pour [521, 541]. Power [860, 2986, 348, 2852, 1918, 469, 2876, 3073, 3141, 347, 1985, 2009, 2180, 1183, 1802, 1289, 354, 2181, 2193]. Power/Performance [2009]. Powerful [2532, 1676]. Powertrain [811]. Practical [993, 1673, 1338, 1493, 3094, 216, 756, 2650, 1462, 2500, 1491]. Practicality [1808]. Practice [1187, 1151, 2797, 1793, 1001, 2631]. Practice-Oriented [1187]. Practices [2713]. Practises [369]. PRAM [674, 675, 897]. PRCS [2711]. Pre [2371]. Pre-historic [2371]. Preaching [1681]. Precedence [1748, 1817, 2886]. Precise [35, 45, 34, 2902, 2366, 2647]. precomputations [1476]. Preconditioning [1418]. Predicate [1264, 2905, 2942, 2941, 2048]. Predicative [3120]. Predict [1893]. Predicting [1129, 200]. Prediction [1894, 1073, 1828, 1388, 2032, 1973, 2349, 188, 1880, 2292, 1122]. Predictive [1846, 850, 1922]. Predicts [1849]. Preface [2084, 1033]. Preferences [1108]. Preferential [2956]. Prefetching [1392, 1435, 2346, 895]. Prefix [604]. Preimage [2226, 2227]. Preliminary [3003, 914, 231, 539, 1828, 1996, 125, 2992]. premixed [1434]. Preparation [1072]. Presburger [2377, 2378]. Prescriptions [538]. Presence [155, 687, 2985, 2014, 979]. Presentation [599, 2005]. Presentations [1139, 1691, 487, 488, 625, 2320]. Presenting [2321]. Preserving [3164, 3116, 717]. Preventing [2680]. Prevention [790]. price [2883]. Primality [2158, 2159]. Prime [2191, 2165]. Primes [2170]. Primitive [1480, 1128]. Primitives [387, 1643, 691, 2025]. Principal [456]. Principals [2696]. Principia [2087]. Principle [1423, 1783]. Principles [1811, 2]. Print [1451]. Printed [529]. Prior [2935]. Priori [201, 388]. Priority [2518, 2521]. Priors [1247]. Prisoner [2958]. Privacy [1181]. Private [2676]. Private-Key [2676]. Privilege [2678]. Pro [1863, 909]. Pro-Active [1863]. Probabilistic [843, 3132, 2561, 1961, 138, 570, 426, 2044, 1055, 3133, 133, 2683, 2521, 2817, 2254, 1613, 1656, 1664, 1663, 2260]. Probabilities [892, 2935, 2024]. Problem [1894, 2832, 2022, 2155, 2244, 452, 1729, 1737, 2785, 1411, 1785, 199, 648, 646, 1818, 3060, 841, 183, 3113, 143, 1745, 451, 669, 2241, 1124, 3030, 1899, 113, 2653, 202, 203, 1819, 80, 2912, 848, 1726, 2284, 1029, 1942, 878, 1897, 1900, 2774, 1859, 1733, 2554, 2982, 3118, 3156, 2893, 2179, 2881, 907, 3083, 1944, 2888, 2889, 133, 1735]. Problem-Solving [2774]. Problems [1736, 2871, 1018, 2530, 2884, 2885, 2660, 2788, 3066, 319, 2036, 2953, 1244, 2796, 446, 1740, 1676, 3052, 1747, 2804, 1258, 1751, 2793, 2517, 2616, 1732, 890, 191, 626, 3026, 294, 1680, 1748, 2175, 2001, 3019, 1867, 901, 896, 52, 2978, 2981, 2982, 467, 2535, 2927, 137, 888, 263, 1750, 635, 378, 293, 2651, 2883, 2138, 2880, 1633]. ProbRough [2292]. Procedure [261, 2110, 3123, 2111, 2643]. Procedures [2379, 339, 2375, 1231, 1597]. Proceedings [1046]. Process [600, 69, 1846, 2710, 1696, 1886, 1979, 629, 2026, 67, 779, 2078, 2708, 308, 1882, 81, 1821, 1066, 66, 1887, 1753, 1880, 3126, 93, 1921, 2440, 2435, 2240, 1770]. processed [961]. Processes [770, 2824, 365, 62, 71, 609, 1851, 2955, 884, 2398, 2825, 70, 1637, 2816, 241, 1754, 1816, 1772, 1909, 63, 915, 1636]. Processing [108, 2307, 2308, 2864, 1134, 978, 10, 569, 1918, 1432, 494, 586, 2969, 1637, 952, 946, 1938, 171, 172, 169, 2999, 940, 485, 840, 587, 1872, 951, 862, 1636]. Processor [1348, 853, 2356, 868, 2366, 883, 251, 2911]. Processors [750, 2218, 3078, 26]. Producing [504]. Product [2433, 2170, 2443, 2429, 2076, 2418, 3140, 2717, 67, 395, 2421, 2416, 2442, 2676, 2419, 2428, 2437, 2422, 2420]. Production [71, 2342, 142]. Products [2434, 2064, 2409, 2410]. Profile [1718, 385]. Profiles [187, 1873]. profiling [745]. Profit [1443, 2443]. Progol [2936]. Program [605, 2761, 932, 1015, 22, 1389, 879, 1815, 934, 3096, 698, 255, 254, 700, 1263, 883, 2150, 1935, 258, 2051, 1231, 2134, 2139, 2149, 2151]. Programmable [857]. Programmatic [2997]. Programmed [171, 172]. Programming [2874, 1025, 3031, 3006, 2722, 270, 702, 723, 351, 72, 1265, 452, 1027, 272, 2995, 2970, 3033, 2524, 1957, 1022, 2922, 2925, 27, 3132, 1703, 2979, 2996, 1015, 2352, 1970, 239, 2969, 292, 691, 1019, 274, 1212, 1822, 2997, 949, 43, 1021, 1379, 1030, 2983, 3026, 1154, 97, 167, 11, 2990, 3034, 2940, 1017, 875, 1166, 2025, 3133, 1014, 1799, 1848, 778, 581, 1373, 467, 683, 3024, 2878, 211, 2931, 1233, 2723, 28, 334, 2965, 2769, 2975, 738, 2148, 117, 2880, 2894]. Programming-Language [2874]. Programming-Level [43]. Programs [1808, 762, 2090, 2823, 1715, 1716, 931, 1542, 1718, 1741, 54, 1742, 2942, 2923, 757, 277, 2822, 1393, 179, 1732, 256, 694, 2120, 3096, 1055, 2770, 252, 2929, 1848, 695, 1020, 255, 436, 254, 126, 929, 1345, 2821, 550, 551, 736, 744]. Progress [379, 1967]. Progressive [1897]. ProHPC [1315]. Project [1034, 69, 1741, 360, 1432, 205, 1053, 1048, 206, 278, 1850, 2711, 2451, 327, 2457, 1141, 362, 511]. Project-Based [1141]. Projection [1217, 1610, 3032]. Projection-Histograms [1610]. Projections [3107, 819]. Projective [1568, 1554, 1570]. Projects [2113]. Projet [541]. Prolog [2947, 2637, 2272]. Promises [70]. Promote [2901]. Promotion [326]. Proof [3103, 3106, 2863, 2361, 3104, 1785, 1219, 2907, 2810, 2365, 284, 779, 2131, 1212, 2121, 2838, 466, 471, 2354, 2053, 2101, 1229, 1233, 2385, 3080, 2132]. Proof-Based [2361]. Proof-Carrying [2053]. Proof-Irrelevance [2863]. Proof-Theoretical [1785]. Proofs [3102, 705, 632, 469, 1787, 642, 142]. Propagating [689]. Propagation [1494, 1411, 256, 489, 1226]. Proper [2640, 1135]. Properties [2434, 2802, 3088, 2220, 771, 3130, 1272, 2846, 2989, 618, 717, 2384, 933, 84, 2070, 734]. Property [1721, 1272, 623]. Proposal [419, 790]. Proposals [1291]. Propose [1819]. Proposed [982, 421, 997]. Propositional [3103, 432, 1211, 1226, 466, 2954, 1231, 2821].

Propositionalization [2933]. Proteases [2998]. Protected [165]. Protecting [146, 2054, 1181, 2052]. Protection [2874, 13, 2692, 311, 2904]. Protein [3049, 2997, 188, 2931]. Proteins [877]. Protocol [2691, 2324, 791, 1041, 2392, 2390, 2068, 792, 793, 2460, 2689, 964, 1071, 153, 1865, 921, 320, 998, 908, 2346, 911, 1197, 1926, 2339, 2141, 2496]. Protocols [684, 1126, 157, 1469, 2361, 2700, 2408, 3171, 469, 2525, 158, 8, 468, 957, 2494, 2360, 2815, 935, 926, 147, 1051, 1052, 2699, 2698, 1007, 1008, 1005, 1006, 1010, 1011, 314]. Prototype [273, 359, 992, 1072]. Prototyping [2898, 2653]. Provable [1187, 449]. Provable-Security [1187]. Provably [3092, 1462]. Prover [2103, 2104, 2092, 2355, 1206]. Provers [1204, 645]. Provider [1039, 1040, 1038]. Providing [2913, 1448, 2464, 2451, 1459]. Proving [2356, 2088, 2087, 285, 2834, 2158, 2159, 2117, 2118, 933, 1234, 562]. Provis [1850]. Provisioning [2470]. Proxy [1469, 2328, 329, 2707, 1193]. Prune [1251]. Pruning [1253, 2125, 1252]. PSEEs [64]. Pseudo [2640]. Pseudo-minimal [2640]. Pseudoflow [1745]. Pseudoflow-Based [1745]. Pseudorandom [416, 2819, 1171]. PSTNs [804]. Pthreads [1373]. Public [155, 1175, 150, 2582, 2486, 2173, 2446, 152, 2670, 2671, 2484, 1483, 2502, 1194, 2490]. Public-Key [150, 1483, 2502]. Publication [1035]. Publications [1047]. publicly [1462]. Publishing [938, 1046, 785, 786]. PULC [910]. Pullback [615]. Pumping [608]. Punctuated [194]. Purchase [2292]. Pure [3119]. Purely [2202]. Purposes [996]. Purse [2688]. Pursuit [2957, 1362, 1654]. Pushdown [3118, 846]. PVM [961, 1374].

Q [1278, 1277]. Q-Learning [1278]. QCD [916]. QOS [956, 2446, 2340]. Quadratic [2189, 2191, 1482, 2192, 1751, 1174, 2203, 2204, 1481, 1472]. Quadri [1816]. Quadri-Dimensional [1816]. quadric [1626]. Quadrics [1360, 1416]. Quadrifocal [1553]. Qualitative [1837, 1841, 1832, 2255, 1522, 1842, 1915, 1916, 1917, 1836, 1839, 1840, 2040]. Qualities [372]. Quality [1330, 2344, 1989, 1983, 1354, 2350, 2026, 326, 1758, 1773, 568, 1709]. Quality-for-ASIS [1709]. Quantification [1227]. Quantified [339]. Quantifier [1216, 3168, 1801]. Quantifier-Free [1216]. Quantifiers [443, 455, 933]. Quantifying [249]. Quantitative [1837, 1713, 1790, 1267, 2040]. Quantization [1587, 2257]. Quantum [1334, 1423, 2864, 2870, 662, 2693, 1199]. Quasi [881, 1588]. Quasi-Random [881]. Quasideterministic [3107]. quaternions [1617]. Queries [346, 1831, 569, 586, 3057, 345, 573, 2234, 587, 1428, 572]. Query [342, 582, 1293, 1040, 347, 114, 1112, 1536, 1541, 2017, 1289, 593]. Querying [1543, 340, 2615, 341, 1545, 1769, 1296]. Queue [2875]. Queueing [2076]. Queues [2518, 2521]. Queuing [1531]. Quicksort [2827]. Quorum [1488]. Quorum-Based [1488].

R [2645]. R&D [2608]. R-AFA [2645]. R0n [3018]. Rabin [647]. Race [1401, 1433]. Races [762]. Rackoff [1479]. Radar [952, 3024]. Radix [457]. Ragnarok [2714]. RAID [1437]. Railway [1964]. RAIN [1319]. RAM [2805, 458]. Random [3135, 1423, 903, 2546, 448, 891, 426, 620, 893, 435, 881, 2242, 2522]. Randomization [894]. Randomized [282, 900, 3077, 902, 901, 436, 2533, 898, 895, 897]. Randomness [3135, 2868, 890]. Range [954, 737]. Rank [2127, 2202]. Rank/Activity [2127]. Rapid [9, 2998, 944]. Rapide [306]. Rapidly [928]. Rarefied [1414]. Raster [999, 984]. Raster-Based [984]. Rate [583, 3033, 955, 1283, 314]. Rate-Based [314]. Rates [2350, 196, 3020]. Rating [1943]. Rational [651, 613, 2178, 3116, 2642, 2199, 620]. Ravenscar [1718]. Ray [1243, 1982]. Rays [3104, 3061]. Razor [1251]. RC2 [418]. RC5 [1466, 405]. RCC [1831]. RCC-8 [1831]. RCTL [2370]. RDBMS [2236]. Re [50, 2901, 2972, 1935, 58]. Re-engineer [2972]. Re-engineering [50, 58]. Re-Use [1935, 2901]. Reachability [768, 815, 771, 821, 2378]. Reachable [816, 2376]. Reaching [1715, 1716]. Reactions [1128]. Reactive [1423, 490, 775, 2379, 1131, 2352, 712, 1863, 2120, 696, 1158, 1065, 1157, 2766, 1167]. Reactivity [231]. Read [279, 2027, 436]. Read- [436]. Read-Only [279]. Readable [142]. Reading [519]. Real [2567, 2849, 89, 2812, 2405, 3072, 2403, 528, 1718, 1940, 1134, 1104, 10, 154, 2331, 771, 2404, 996, 3130, 1703, 2551, 1713, 1196, 14, 1701, 1702, 1611, 614, 2850, 2191, 2192, 861, 962, 966, 2249, 90, 2566, 960, 797, 953, 2349, 2590, 955, 11, 91, 1704, 848, 956, 2576, 2201, 1907, 864, 945, 384, 92, 925, 1840, 940, 3134, 963, 2467, 2203, 1771, 725, 93, 1355, 3173, 2385, 177, 1167, 2142]. Real-Image-Based [2576]. Real-Time [2567, 89, 2405, 2403, 1718, 1134, 10, 771, 2404, 3130, 1703, 1713, 1196, 14, 1701, 1702, 1611, 2850, 861, 962, 966, 2249, 90, 960, 2349, 955, 11, 91, 1704, 956, 864, 945, 384, 92, 925, 940, 963, 2467, 1771, 725, 93, 1355, 2385, 177, 1167, 2142]. Real-World [1940, 2249, 848]. Realism [2606]. Realistic [904]. Reality [2593, 1529, 2606, 322, 2587, 2588, 2604, 2568]. Realizability [2151]. Realizing [1816]. really [1561]. Reals [477, 468]. Realtime [95]. Rearrangements [942, 3074]. Reasoning [2297, 1202, 1841, 1843, 145, 2771, 690, 1773, 1948, 135, 1519, 242, 1842, 1915, 3098, 1116, 2112, 2210, 2086, 1523, 2040, 2854, 140, 118, 2142, 128]. Receipt [148]. Receipt-Free [148]. Receiver [1499]. Recently [421]. Recently-Proposed [421]. Reciprocity [2190]. Recognition [993, 386, 991, 1556, 849, 523, 974, 999, 1592, 2293, 1653, 379, 977, 392, 393, 850, 385, 2251, 2222, 403, 1654, 401, 1658, 402, 2019, 388, 998, 384, 2556, 995, 3024, 2296, 399, 400, 1594, 3032, 969, 3014, 1659, 1613, 1656, 1593, 1664, 2887, 1657, 1666, 1663, 1661, 1649, 1662]. Recognizability [479]. Recognizable [616, 3169]. Recognize [668, 976]. Recognizing [1873, 1655, 1646]. Recombination [3046]. Recommend [2624]. Recomputation [345]. Reconciliation [1199]. Reconfigurable [860, 866, 855, 2968, 858, 862, 880, 861, 856, 859, 904, 863, 857, 867, 854]. Reconfiguration [176, 1918, 851]. Reconfigure [1936]. Reconstructing [3104, 989]. Reconstruction [983, 658, 2996, 980, 2318, 990, 1229, 1630, 1590, 1627, 1578, 1671, 1570]. Recorded [3009]. Recover [2427]. Recoverable [1386, 1461]. Recovery [2423, 1117, 2424, 919, 233, 2426, 160, 601, 915, 1571, 1638]. Rectangles [2819]. Recurrent [1023, 2549, 1083, 3113, 1024, 836, 1892]. Recursion [1791, 1795]. Recursive [280, 1270, 977, 2833, 3093, 2853, 257, 694, 345, 845, 695, 1263, 1804, 736, 727, 729]. Recursively [477, 3173]. Redesign [1985]. REDLOG [137]. Reduce [2558]. Reducing [1240, 3110, 2810, 265, 2175]. Reduction [2273, 1278, 1890, 1110, 1794, 773, 774, 2013, 829, 624, 2096, 2200, 2353, 2845, 2177, 735]. Reductions [2386, 3071, 2367, 289, 635, 1501]. Reducts [2214, 2271, 2263]. Redundancy [1278, 2035, 391, 1229, 116]. Redundant [1913]. Reef [2599]. Reengineering [2429, 2425]. Reference [1035, 1518, 495, 1520, 2251, 1952, 1517, 1669]. References [1392]. Referent [496]. Referential [595, 596, 1540]. Refine [1070].

Refinement [1152, 1062, 74, 75, 3145, 429, 615, 1112, 934, 935, 717, 1059, 721, 46, 47, 929]. Refinements [2947]. Refining [722]. Reflection [2916, 2917, 15]. Reflections [713]. Reflective [13, 16, 2915, 11, 102, 12]. Region [1931, 2515, 2232, 1586]. Region-Based [1931, 2232]. Regions [2538, 1831, 817, 389]. register [737, 740]. Registration [29, 1659, 1617, 1618]. Regression [2216, 2979, 1095, 1252]. Regular [2538, 2635, 2557, 2559, 464, 3115, 2850, 3147, 834, 3040, 2643, 2645, 3148, 2362, 1563]. Regulated [970]. Regulator [2203]. Rehearsal [2341]. Reinforcement [1152, 1279, 2928, 1925, 1281, 1280]. Reintegration [218]. Relabelling [610]. Related [3106, 1749, 2691, 3140, 1314, 2882, 1501]. Relating [2824, 454, 838]. Relation [2133, 2800, 2429, 2246, 2768, 624, 1823, 1517]. Relation-Algebraic [2133]. Relational [723, 109, 2039, 2935, 2928, 2945, 2951, 113, 347, 2045, 1243, 2930, 1289, 1121, 116, 1585, 567]. Relations [723, 1782, 2227, 443, 2261, 2826, 1511, 1840, 2176, 1516, 1512, 2226]. Relationship [1106, 2122, 1750, 1121]. Relationships [1898, 2290, 2277]. Relative [1560, 2189, 1843, 2193]. Relaxing [1395]. Relay [762]. Relevance [2863, 1272]. Relevancy [2122]. Relevant [1254, 1090]. reliability [1667, 1562]. Reliable [1459, 1561, 1596]. Relies [1939]. Relinking [182]. RELVIEW [723]. Remarks [2218, 1222, 1223, 1177]. Remote [168, 1426, 163, 1364]. Remotely [1478]. Removing [591]. Rendering [1354, 390, 503]. Repair [1683]. Repates [3044]. Repeat [2941]. Repeated [682]. Repetition [494, 286, 287, 3161]. Repetition-Free [3161]. Replacement [3144]. Replicated [882]. Replication [279, 1386]. Report [2444, 999, 231, 131, 1710, 125]. Reporting [3040]. Repository [1532]. Representable [1831]. Representation [2016, 2348, 1521, 261, 3016, 1023, 2317, 2319, 1871, 1093, 1024, 972, 255, 700, 1512, 1919, 1589, 1568]. Representational [1513]. Representations [1743, 183, 1643, 1506, 1522, 457, 2981, 2980, 2374, 122]. Representative [1806, 2237, 1089]. Represented [446, 2287]. Representing [71, 3147, 3148, 973, 563]. Republic [2693]. Requests [494, 496]. Requirements [2529, 1674, 1449, 724, 373, 2726, 2401, 790, 216, 1712, 2437, 1456, 129]. Requires [1394]. Requisitioning [1452]. Resampling [2987]. Rescue [2914]. Research [2031, 849, 300, 1182, 517, 378, 738]. Reservation [1382, 956, 314]. Reset [2807]. Residential [2339]. Resistant [156, 155]. Resolution [2094, 2095, 1783, 2110, 1222, 1223, 2127, 1209, 2267, 313, 2111, 2112, 986, 1578]. resonance [1640]. Resource [3135, 441, 2659, 2655, 967, 1812, 956, 621, 945, 314]. Resource-Bounded [3135, 441]. Resources [1978, 2470, 374, 2477]. Respect [3176, 907, 2208]. Response [1332, 1346]. Responses [494]. Responsibilities [1161]. restoration [1651]. Restricted [1881, 467, 1876]. Restricted-Domain [1876]. Restructuring [273, 60, 343]. Result [469, 553]. Result-Indistinguishable [469]. Resultant [136]. Resultants [135]. Results [1543, 3003, 1213, 539, 2167, 1053, 2527, 1747, 2804, 2192, 1110, 2540, 1014, 1280, 289, 1499, 1267, 405, 2992, 1405]. Resurrected [2720, 2721]. Retail [2618]. Retaining [2279, 1984]. Retention [1028]. Retina [1874, 2467]. Retrenchment [1062]. Retrieval [534, 1302, 2612, 570, 1948, 2598, 1237, 1536, 3051, 1958, 2625, 1583]. Retrofitted [1821]. Retrospective [517]. Reusability [1860]. Reusable [2408].

Reuse [2444, 2443, 76, 77, 1764, 714, 537, 1763, 538, 250]. Reuse-Driven [250]. Reuse-Oriented [1764]. Reveal [1642]. Revenue [2449]. Reverse [2427, 1755, 551]. Review [1878, 2669, 1014]. Revise [1819]. Revised [64]. Revising [2041]. Revision [2711, 129]. Revisited [3091, 109, 816, 3030]. Rewards [2244]. Rewrite [2097, 2088, 631, 722, 2843, 3114, 557]. rewrite-based [557]. Rewriting [613, 643, 3111, 3112, 615, 622, 640, 546, 635]. Rho [2198]. Rice [3129]. Rich [2614, 1803]. rig [1579]. Rigid [636]. Rigidity [1670]. Ring [2173, 2891, 1463]. Ring-Based [2173]. Rings [2188, 1472]. Risk [1908, 1771]. River [1963, 1894]. RLGG’s [2947]. RNA [1893]. Road [2355, 1599]. Roadvehicle [1324]. Robert [2164]. Robin [2152]. RoboCup [1139, 1126, 1136, 1124, 1127, 1163, 1125, 1155, 1158, 1144, 1162, 1138, 1147, 1132]. RoboCup- [1138]. RoboCup-97 [1125, 1132]. RoboCup97 [1148]. Robot [1164, 1134, 1668, 1136, 1902, 1925, 1908, 202, 203, 1904, 1907, 2228, 1147, 1142, 1667]. Robotic [2976, 1896, 2230, 2808, 399, 1933]. Robotics [1931, 1807, 1124, 1133, 1923, 1141]. Robots [1527, 1027, 1924, 2525, 1708, 2590, 1143, 1929, 1926, 1927, 1146, 1149, 1167]. Robust [1931, 1569, 3071, 1906, 980, 2815, 1618, 1621, 12]. Robustness [2347]. Rockies [511]. Rödl [3105]. Rogi [1145]. Rogi-Team [1145]. ROI [1597]. Role [498, 2356, 1131, 111, 2464, 3095, 3007, 391, 1635, 2696, 2452, 122, 1634]. Role-Based [1131, 2696]. Roles [1161]. Roll [1864]. Roll-Mill [1864]. Rollback [873]. ROM [1046]. roofs [1578]. Room [360, 375, 325]. Rooms [360, 363, 364]. Roomware [356]. Root [1174, 2163]. ROSE [2294]. Rosser [619, 624, 623]. Rotating [1403]. Rotation [3042, 3062, 2329, 2685, 1562]. Rotation-Symmetric [2685]. Rotorcraft [2969]. Rough [2298, 2305, 2309, 2311, 2312, 2313, 2314, 2315, 2245, 2293, 2273, 2246, 2215, 1109, 2285, 2255, 1088, 2238, 2267, 2209, 2235, 2272, 2284, 2224, 2210, 2276, 2294, 2247, 2295, 2296, 2274, 2269, 1190, 1102, 2282, 2270]. Rough-Set [2255]. roughness [2208]. Round [469, 2152]. Round-Robin [2152]. Rounds [424, 412]. Route [1847, 1507]. Routed [1991]. Router [923, 1997, 1996]. Routers [2000, 175]. Routine [429]. Routing [2022, 2656, 855, 671, 176, 2851, 2852, 1990, 2659, 964, 3077, 2241, 2446, 2792, 2010, 905, 1029, 878, 672, 1848, 1007, 1008, 1003, 1009, 1005, 1006, 1010, 1004, 1013, 898, 897, 175, 924]. Rowed [2828]. RSA [1464, 2691, 2494, 2487, 160, 1180, 1183]. RSDM [2233]. RSKDD1 [2312]. RSM1 [2311]. RSM2 [2313]. RSMA1 [2314]. RSMA2 [2315]. RT [94]. RT-Objects [94]. RTSP [2324]. Rueppel [1195]. Rule [3023, 2289, 214, 2131, 1268, 1114, 1915, 2277, 2994, 717, 1941, 1098, 2263, 1123, 1101, 2282, 3003, 2132]. Rule-Based [214, 1268, 1915, 2994, 717]. Rules [1837, 2281, 1855, 1076, 1077, 1079, 2936, 1910, 2088, 1084, 2212, 42, 2237, 1089, 1111, 1424, 2272, 1534, 1094, 1095, 2262, 722, 2274, 2931, 1256, 1122, 2033]. Run [73, 1743, 260]. Run-Time [73, 260]. Running [1334, 912, 1352]. Runtime [855, 1383, 858, 311, 870, 1695, 872, 875].

S [2686, 633, 413]. S-Boxes [2686].

S-Terms [633]. S3M [2334]. S4 [630]. SA [1675, 1922]. Saccadic [3003]. Safe [338, 2920, 1791, 2053, 1705, 1706, 549, 2060]. Safe-Tcl [2060]. Safety [707, 708, 1693, 1694, 2377, 825, 826, 831, 719, 717, 720, 941, 2878, 1067, 933, 2600, 1717]. Safety-Critical [707, 708, 825, 1067]. Saks [2804]. Sales [1304]. Salesman [1735, 1942]. Same [1743]. Sample [108]. samples [1590]. Sampling [2828, 891, 2827, 893].

SAR [1416, 3032]. SAT [2967, 2531, 2100, 2895, 3064, 288]. Satellite [2333, 2276]. Satisfiability [1818, 183, 3099, 467]. Satisfying [1449]. Saturation [645]. Saturation-Based [645]. Sauvim [1928]. SC0k [470]. SC0k-Completeness [470]. ScaFiEP [1421]. Scalability [526, 305]. Scalable [2654, 512, 1362, 597, 278, 2332, 1421, 945, 940, 2668, 175]. Scale [991, 1410, 578, 1242, 2329, 592, 148, 2973, 1866, 198, 1649, 1599]. Scales [1105, 2585]. Scaling [3007, 2965, 1101, 2286]. Scan&Track [2571]. Scanned [983, 984, 990, 973]. Scanning [1957]. Scatter [182]. Scenario [758, 2333, 1763]. Scenario-Based [758]. Scenarios [2955, 578]. Scene [788, 1641]. Scenes [1557, 124, 1581]. Scheduled [275, 2477, 740]. Scheduler [1912, 2152]. Schedulers [884]. Schedules [2677, 1685, 1686]. Scheduling [583, 1688, 2244, 1677, 1346, 1675, 1678, 884, 1431, 1747, 451, 15, 905, 1748, 885, 1347, 92, 886, 883, 2790, 2893, 1750, 2385, 2882, 2886, 741]. Schema [1164, 109, 2985, 2726, 803, 2966]. Schema-Based [1164]. Schemata [1014]. Schematizations [3113]. Scheme [411, 927, 1868, 2857, 2705, 275, 1676, 1448, 919, 1191, 2512, 1195, 159, 601, 1367, 1121, 997, 2490, 2498, 1462, 2889, 2507, 2497]. Schemes [605, 2658, 1407, 2851, 2559, 2171, 2680, 2704, 1470, 2707, 294, 1490, 2487, 148, 162, 412, 151, 1263, 445, 898, 2172, 1193, 2882, 1476, 1497]. Schönhage [2156]. Schoof [2506]. Schützenberger [3140]. SCI [1374, 181, 1373]. SCI-cluster [1373]. Science [1035, 1202, 1044, 217]. Scientific [272, 1338, 961, 179]. SCM [2710, 2715, 2709]. SCOOP [906]. Scope [1268]. Scoped [1540]. Score [2871]. Scoring [1092]. SCR0* [2401]. Screen [519]. Screening [1885]. Screens [499]. Script [509]. Scripting [248, 304]. Scripts [520, 516, 510]. SDL [2473, 2071, 757, 781]. Sealing [2059]. Seamless [2021, 2664]. Search [1274, 1875, 1218, 2558, 1240, 199, 648, 1106, 1044, 650, 2771, 182, 3057, 3073, 2984, 1902, 1806, 2531, 2552, 2014, 3004, 453, 575, 3061, 2974, 3022, 195, 1897, 3082, 1866, 1905, 2988, 1400, 1233, 2242, 1684, 3064]. Searches [2187, 1535]. Searching [433, 2791, 610, 458, 2214]. SECOMS [2333]. Second [3129, 479, 626, 3168, 637, 195, 1800, 1802, 1803, 999]. Second-Order [479, 626, 3168, 637, 1800, 1802, 1803]. Secondary [1893]. Secret [2705, 2704, 2703, 3089, 1497, 162, 1462]. Sector [3057]. Secure [415, 146, 1488, 2678, 2700, 1486, 2689, 2357, 1191, 1179, 2145, 1184, 2485, 1490, 1503, 1483, 1447, 2499, 149, 941, 2483, 1457, 1462, 2496, 2506, 2146, 1485]. Securing [2672, 1460]. Securities [1943]. Security [411, 1187, 2630, 1042, 2049, 2505, 319, 2061, 2673, 1293, 2057, 2690, 2054, 316, 1189, 2058, 418, 2332, 1188, 2494, 2360, 2670, 2671, 1501, 2060, 2062, 320, 1445, 1489, 1491, 963, 449, 2051, 1456, 1479, 2497, 2495]. See [2483]. Segment [571]. Segmentation [985, 968, 2578, 2937, 1652, 848, 1091, 1605, 994, 1650, 1651, 1586, 1596, 1620, 1591, 1585, 1584, 1619]. Segments [2514]. Segregation [1403]. Select [1849]. Selecting [2030]. Selection [1039, 2573, 3001, 78, 394, 1249, 3021, 1248, 2994, 2520, 2288, 2253, 3011, 1984, 1987, 2956, 2024, 2252, 1597, 1581]. Selections [581]. Selective [2700, 1085, 2266]. Selectively [2124]. Self [2783, 1423, 1027, 1884, 1554, 2969, 1110, 2983, 1939, 1143, 1942, 925, 254, 1585, 2992, 1667]. Self-Adaptation [2969, 2983]. Self-Adaptive [1027]. Self-calibration [1554]. Self-Diagnosing [1884]. Self-Learning [1143, 2992]. self-localization [1667]. Self-Maintenance [1939]. Self-Organisation [1423]. Self-Organizing [2783, 1110, 1942].

Self-Similar [254]. Self-Testing [925]. Seller [1443]. Selmer [2195]. Semantic [108, 1215, 2626, 37, 2955, 2145, 2112, 2972, 562, 2146]. Semantics [2849, 607, 2079, 35, 45, 110, 39, 2071, 1015, 111, 1790, 239, 41, 34, 112, 238, 114, 595, 596, 543, 44, 2829, 2769, 552, 554]. Semantics-Based [114]. Semi [2018, 3119, 3072, 341, 2114]. Semi-automatic [2018]. Semi-continuity [3072]. Semi-Full [3119]. Semi-structured [341]. Semi-Unification [2114]. Semiconductor [1953]. Semidefinite [292, 2894]. semiformal [556]. Semigroups [653, 654]. Semiqualitative [1835, 1836]. Semistructured [1543, 1545]. senders [1191, 2684]. Sense [2554, 3075]. Sensing [2230, 1872]. Sensitive [2722, 2238, 3176, 619, 1859, 2539, 2536, 1260]. Sensor [2318, 1874, 1871]. Sensors [2591, 2592]. Separate [2443]. Separating [2639]. Separation [971, 844, 997, 1619]. Sequence [340, 2996, 850, 618, 2402, 3038, 1651]. Sequences [3049, 2317, 2997, 396, 397, 620, 1199, 1560, 1571, 1631, 1555, 1638, 1598, 1619]. Sequential [2865, 2047, 859, 1096]. Serial [1375]. Series [1288, 2178, 3141, 667, 473, 1115, 2180, 2181]. Series-Parallel [667, 473]. Serine [2998]. Serpent [419]. Server [1368, 2678, 2785, 2324, 1381, 2328, 329, 221, 2323, 223, 2494, 310, 1544]. Server-Aided [2494]. ServerNet^{T M} [1998]. Servers [1697, 649, 1384, 1492]. Service [2447, 2654, 809, 1035, 2344, 802, 1989, 2455, 2613, 2476, 920, 649, 803, 790, 2454, 2453, 2472, 2451, 1051, 1052, 2326, 2475, 2337, 230, 1531, 568]. Services [2450, 794, 809, 2663, 2448, 784, 2327, 2470, 2461, 789, 949, 2332, 2335, 2459, 2333, 2485, 804, 1363, 720, 568, 1, 2]. SESAME [2672, 2669]. Session [2297, 2298, 2299, 2300, 2301, 2302, 2303, 2304, 2305, 2306, 2307, 2308, 2309, 2310, 2311, 2312, 2313, 2314, 2315, 2316, 2444, 2430, 2423, 2415, 2407, 2435]. Set [2311, 2313, 2314, 2315, 1216, 3085, 638, 775, 37, 1576, 2215, 2461, 1806, 201, 2255, 1088, 575, 1624, 3150, 1726, 2224, 2210, 2294, 2247, 291, 2282, 2651, 2142, 2892]. Set-Based [775]. Set-Inspired [1088]. set-theoretic [2142]. SETHEO [1234]. Sets [3135, 2298, 2305, 2309, 2312, 1356, 2245, 2293, 2273, 816, 634, 3137, 2168, 3052, 2879, 2814, 3146, 2196, 1109, 135, 1439, 2069, 479, 2639, 345, 2209, 2235, 466, 2213, 1115, 2284, 2122, 2988, 2295, 2296, 2274, 1355, 2070, 2269, 1190, 1102, 1617, 1640, 2270]. Sets-Based [2235]. Setting [1764, 2653]. Severely [1959]. Sex [3001]. SG [158]. SGA [195]. SGML [543]. Shading [1643, 1625]. Shadows [1642]. Shallow [628]. Shape [658, 1515, 1643, 692, 1628, 3005, 1625, 1556, 1609, 1583, 1649]. shapes [1568, 1644]. Share [1086]. Shared [2002, 174, 2909, 1354, 942, 2068, 586, 278, 882, 325, 178, 1349, 1375, 99, 1386, 3155, 1032, 911, 587, 1353, 1405, 1199]. Shared-Memory [1375]. Shared-Nothing [586, 587]. shareholders [2491]. Sharing [2705, 2704, 2703, 162, 767, 1462, 2491, 1497]. Shave [1251]. Sheet [526]. Shift [544]. Shifts [2578]. Shimura [2154, 2190]. Ship [1914]. Ship-Board [1914]. Shogi [2940]. Shop [2244, 1899]. Short [661, 1156, 1255, 2966]. Short-Term [1156]. Shortest [2510, 3165, 428, 907]. Shot [2330]. Should [302]. Shrinking [2683]. Shuffle [3140, 440]. Sicily [1160]. SICMA [2323]. Siemens [781]. Sieve [2162, 2163, 2164]. Sieves [2165]. Sign [386, 379, 387, 401, 402, 2483]. Signal [1362, 1873, 392, 393, 946, 940, 2231, 1619]. Signals [2331]. Signature [2691, 2171, 2707, 1490, 2487, 1604, 1195, 159, 151, 1444, 2706, 2172, 1193, 2490, 1491]. Signatures [1504, 1192, 1489, 2483, 1477].

Signcryption [1194, 2490]. Significance [2174]. Signing [2059]. Signs [380]. SIMD [1385, 1993, 1395, 1331]. SimDS [2337]. Similar [254]. Similarities [2716]. Similarity [3086, 2932, 569, 539, 2089, 2246, 61, 2934, 1097]. Similarity-Based [2089]. Simple [3144, 2877, 432, 1015, 2810, 2519, 2803, 194, 2960, 196, 2117, 2118, 2934, 3044, 1744, 1804, 2961]. Simple Com [1714]. Simpler [2805]. Simplex [1745]. Simplification [509, 1224, 1226]. Simplified [892, 204]. Simplifying [430, 264, 2843]. Simulated [2996, 3028, 1922]. Simulating [996, 516, 1245]. Simulation [1335, 1414, 1406, 903, 1325, 1336, 2900, 2080, 755, 1317, 1707, 1890, 1863, 1333, 823, 870, 1326, 1434, 1409, 5, 1436, 1417, 1987, 6, 759, 951]. Simulations [1781, 2382, 1339, 1974, 439, 2875, 2585, 3063]. Simulator [1332, 2591, 791, 1976, 2592, 1827, 867, 1153, 2600]. Simultaneous [1573, 636, 623]. Single [1081, 2014, 256, 428, 907, 3152, 915, 1618, 1648]. Single-Agent [2014]. Single-Tape [3152]. SISCI [1373]. SISCI-Pthreads [1373]. Sites [598]. Situation [1966, 1157]. Six [412]. Size [776, 777, 1730, 470, 1403, 448, 1597, 1357]. Skeleton [504]. Skeuomorphs [2977]. Skewed [979]. Skewness [1077]. Skill [383]. Skills [1250]. Slack [2147]. Slicing [2973, 744]. Small [3163, 3037, 2182, 2183, 1742, 1174, 2126, 1183, 1142, 3081, 1471]. Smalltalk [743]. Smart [2505, 2092, 1874, 1184]. SMATV [2334]. Smog [1109]. Smooth [661, 2160, 2169, 1630]. Smoothed [1995]. Smoothing [1603, 1591]. SMP [1373, 909]. SMP-like [1373]. SN [644]. Snakes [1599]. SNMP [320, 2479]. SNMP-based [2479]. Soccer [1152, 1139, 1164, 1527, 1134, 1168, 1165, 1154, 1133, 1135, 1143, 2589, 1129, 1128, 1146, 1167]. Soccer-Playing [1167]. Social [2289, 1863, 233, 372, 2772, 2774].

Socially [2616]. Societies [247]. Society [2613, 1321]. Socket [911]. SOFM [1824]. Soft [2310, 1973, 1972, 1914, 2239]. Softbot [1154]. Software [1380, 68, 704, 1310, 307, 59, 2438, 2431, 702, 62, 69, 71, 74, 75, 50, 938, 1693, 1694, 1371, 1936, 2714, 410, 2430, 2660, 1413, 2436, 2432, 275, 1713, 2412, 755, 79, 1934, 1322, 1422, 709, 2726, 70, 265, 2470, 2401, 78, 1764, 752, 946, 2719, 1819, 65, 2426, 57, 91, 81, 2409, 2410, 1810, 2421, 2294, 309, 66, 3008, 1067, 63, 2716, 2713, 2437, 1364, 2422]. Software-System [2436]. Sokoban [2014]. Solitons [1406]. Solution [452, 1371, 2877, 1320, 2679, 3027, 191, 1938, 1900, 2861, 1823, 2778, 2781, 1617]. Solutions [2796, 830, 627, 1194, 1579, 1471, 2073]. Solve [2001, 52, 2982]. Solved [202, 203]. Solver [1341, 1402, 1419, 1420, 2099, 1340, 1405]. Solving [1738, 1741, 681, 1258, 3026, 631, 1029, 1942, 2774, 1859, 1415, 2535, 1944, 1633]. SOM [1282, 1951]. Some [1543, 118, 2073, 2194, 2218, 982, 1182, 852, 2527, 1811, 1222, 1223, 2259, 703, 631, 2277, 901, 3169, 1679, 2266, 2992, 1213, 1302, 421]. sorted [561]. Sorting [458, 460, 898, 2651]. Sorts [460]. Sound [1022, 3097]. Source [3065, 2425, 844, 428]. Sources [1104, 2622, 594, 2522]. SP2 [178, 1340]. Space [1674, 3058, 1240, 942, 373, 3149, 374, 1522, 3095, 91, 1767, 2520, 1923, 2176, 2150, 2165, 1975, 1514, 1512, 951, 357, 1629, 2149]. Space-Efficient [3058]. Space-Shared [942]. Space-Time [951]. Spaces [2438, 2223, 2868, 891, 2594, 2225, 294, 2247, 356, 2362, 3041]. Spanish [1851]. Spanners [2890]. Spanning [2133, 1730, 900, 461, 2813, 2856]. Sparse [1241, 2535, 435, 1560]. Sparse-Meshes [2535]. Spase [3038]. Spatial [1521, 985, 1518, 1082, 1520, 1510, 3145, 336, 1085, 1587, 1506, 2207, 1519, 1525, 337, 112, 372, 1528, 1665, 1529, 1508, 1509, 1523, 1511, 1516, 1530, 1512, 1517, 2206, 1524, 124]. Spatialization [2595]. Spatially [3032]. Spatially-Localized [3032]. spatio [1589]. spatio-temporal [1589]. Spatiotemporal [1637, 1636]. Spatiotemporally [1620]. Speaker [494]. Special [3048, 403]. Specialization [2877, 701]. Specialized [1500]. Specific [30, 1987]. Specification [1777, 706, 235, 818, 1768, 1967, 1861, 7, 1071, 713, 42, 2902, 545, 228, 48, 553, 2142, 556]. Specifications [686, 920, 2799, 1063, 3, 795, 796, 722, 548, 551, 557, 561]. Specified [712]. Specify [601]. Specifying [704, 707, 708, 554, 709, 2401, 1189, 1763, 2765]. Spectra [1786]. Spectral [665, 1838, 1752]. specular [1625]. Speculation [276]. Speculations [1904]. Speculative [260, 2366]. Speech [849, 1022, 850, 2637, 1658, 1239, 380, 1649]. Speed [2532, 1178, 1279, 1375, 1839, 1183]. Speeding [1476, 2198, 3152, 1277]. Speeding-Up [3152]. Speeds [2011]. Speedup [1393]. Spelling [682]. Sphere [659, 2978]. spheres [1596]. SPI [64]. Spike [843]. Spiral [882]. Spirit [1150]. Splines [659, 1630]. Splitting [660, 1728, 737]. SPN [2065]. SPNT [260]. Spoofing [2674]. Sports [1685, 1686]. SPP1600 [1436]. SQL [1292]. SQL3 [1290]. Square [2517, 2163]. Square-Center [2517]. Squares [1635, 1563, 1596, 1634]. SSA [734]. SSAPRE [735]. ST1 [2316]. Stability [817, 820, 2256]. Stabilization [835]. Stabilized [1434]. Stable [1418, 1726, 123]. Stack [911, 2339, 739]. Stage [2878, 3074, 1663]. Stakeholders [2436]. Standard [2466, 1687, 2679, 664, 2014, 1364]. Standards [1999]. Stanford [1052, 1051]. Starfree [616]. Start [1917]. Start-Up [1917]. State [651, 2381, 775, 2379, 2558, 38, 1881, 2636, 1278, 2816, 2425, 2540, 763, 2120, 2360, 2642, 2815, 3005, 2709, 1068, 55, 56, 988, 2362]. State/Event [763]. Statecharts [707, 708, 709, 721]. Statemate [2403, 764]. Statements [138, 257]. States [1240, 753, 2376, 2082, 696, 2773, 3126]. Static [353, 280, 589, 2851, 1983, 773, 774, 256, 562, 748]. Statically [2920, 2916, 2917]. Statistical [1894, 2316, 2217, 201, 2037, 2930]. Steady [3005]. Steering [1388]. Steiner [2881]. Step [754, 205, 3009, 206, 3129, 2638, 1597, 828]. Stepping [1407]. Steps [852]. Stepwise [934]. Stereo [1931, 1923, 1564, 1616, 1579, 1566, 1555]. Stereoscopic [2574, 1629]. Stereovision [1576]. Still [3154]. Stochastic [2557, 2985, 2548, 1357, 2552, 2933, 3022, 1827, 2934, 2547, 2286, 1612]. Stochastically [2541]. Stochasticity [847]. Stock [3026, 1897]. Stop [688]. Storage [729, 2520]. Store [2907, 1003]. Store-and-Forward [1003]. Stored [1995]. Storing [1958]. Storytelling [2603]. Strain [1406]. Strategies [2958, 365, 2615, 1161, 2976, 1957, 982, 3002, 2212, 1070, 2969, 1903, 2827, 1013, 2024, 729]. Strategy [2016, 1346, 2449, 3061, 1819, 1942, 3082, 1273, 1234, 1345, 2781]. Stratifications [822]. Stream [410, 409, 1480, 848, 2682]. STREAMER [1995]. Streaming [2325, 1196]. Streams [222, 2338]. Street [992, 1485]. Streets [992]. Strength [1965, 3135, 664, 735]. Strengthened [1489]. Strengths [1946]. Strict [2107, 2554]. String [3037, 3042, 976, 3034, 3035]. Stringology [894]. Strings [3086, 3157, 681, 429, 826, 2037, 435]. Strong [1845, 2527, 2859, 2770, 2262, 1496]. Strong/Weak [2859]. Strongly [2943]. STRPLAN [1853]. Structural [539, 1412, 2833, 2076, 976, 2368, 2442, 2555]. Structurally [2948]. Structure [607, 209, 523, 444, 2424, 1575, 341, 1015, 3057, 448, 1806, 1572, 1642, 1063, 1632, 1985, 1798, 1537, 1949, 2931, 2583, 1932, 1573, 1569, 1574, 1664, 1497]. Structure-Based [341]. Structured [535, 2067, 657, 1660, 839, 691, 332, 529, 599, 548, 341, 1589]. Structures [2783, 280, 607, 582, 3104, 1893, 657, 170, 3093, 257, 840, 2521, 1514, 736, 1615, 567]. Structuring [2441, 1816, 1066]. Stubborn [2069, 2070]. Student [1677]. Students [88]. Studies [2391, 497, 510]. Studio [2341, 2576, 362]. Study [1833, 1954, 271, 811, 707, 1677, 708, 1078, 1828, 2424, 1713, 1293, 1806, 1636, 1637, 274, 2912, 1115, 719, 804, 2045, 2482, 200, 700, 874, 2956, 2992, 362, 2330, 127, 560]. Studying [2076]. Stuff [1015]. Style [2559, 526, 2112, 522]. Sub [2690, 1981, 1546]. Sub-collection [1546]. Sub-linearised [2690]. Subanalytic [822]. Subclass [2560, 2555]. Subgraph [1729, 2885]. Subgraphs [1730, 2894]. Subgroup [2207, 2206]. Subject [297, 1767]. Subject-Based [1767]. Subpaging [168]. Subroutines [1032]. Subscripts [252]. Subsequential [2546]. Subset [2832, 1248]. Subset- [2832]. Subsethood [1826]. Subspace [2957]. Substitution [1055, 1985]. Substitutions [1781, 2862]. Substrate [105, 106]. Substructure [1351]. Subsumption [3113, 2946, 2128]. Subsystems [2001, 630]. Subtype [2853]. Subtyping [3120, 3122, 697, 2858]. Success [1056]. Successive [1954, 2188]. Sufficient [833]. Suffix [3043, 3035, 3038, 682, 3044]. Sugeno [2229]. suitable [2503]. Suite [3006, 2404, 791]. Sum [1480, 2179]. Summaries [1300]. Summary [1204, 200, 1012]. Summation [2831]. Sums [2832, 2166]. Super [651]. Super-State [651]. Supercomputer [1416]. Supercomputers [1406, 1327, 1417]. Supercomputing [1378, 1705, 1706].

Superior [2967]. Superpolynomial [3127]. Superposition [2107, 2105, 2106]. superscalar [742]. Supervised [1954, 1875, 1925]. Supervision [820, 1815, 1916, 1917, 2020, 1935, 1921]. Support [583, 1962, 758, 808, 30, 31, 17, 18, 2976, 360, 1387, 1365, 2327, 2476, 959, 1532, 32, 1109, 1254, 2113, 1303, 2019, 1995, 310, 872, 1455, 2288, 1255, 767, 1290, 1977, 2911, 1921, 1304, 2463, 1963, 128, 1646]. Supported [2612, 395]. Supporting [218, 1285, 1432, 724, 2665, 223, 2719, 12, 1537, 324]. Surface [1627, 3142, 2600, 1645, 1659, 1629]. Surfaces [2510, 1625, 1630, 1626]. Surge [962]. Survey [2700, 890, 901, 115, 2562]. Sustainability [369]. Swaps [3048]. Swarm [3012, 3013, 3011]. Switch [1994, 3074, 2008]. Switch-Based [2008]. Switchbox [2241]. Switched [813, 2478]. Switches [2080, 1992]. Switching [809, 805, 817, 2771, 1902, 3076, 2340, 863]. SWN [2080]. SX [1355]. SX-4 [1355]. Syllabification [2033]. Syllogistic [1816]. Symbiosis [799, 2590]. Symbol [975, 974, 2561, 392, 393, 2222]. Symbolic [2386, 772, 2387, 1715, 1716, 2979, 2371, 2394, 2268, 2384, 759]. Symbols [1786, 976]. Symmetric [1442, 3130, 3070, 2685, 413]. Symmetries [916, 2065]. Symmetry [754, 1567, 2367, 2368]. Synchronisation [688]. Synchronization [174, 265, 958, 524, 966, 173, 1395, 592, 3143]. Synchronized [2064, 2320]. Synchronizing [1697, 871]. Synchronous [2352, 2820, 2842, 3178]. Syngas [1434]. Synonyms [693]. Syntactic [1622]. Syntactical [2264]. Syntax [2542]. Synthesis [429, 277, 964, 2290, 401, 715, 2319, 856, 2781, 2151]. Synthesising [1063]. Synthesizing [2901, 483, 859, 832, 3014]. Synthetic [996, 139, 1127, 952, 228, 3024]. System [2712, 2046, 3003, 1319, 2090, 2018, 991, 2093, 723, 2427, 365, 1831, 2103, 2104, 1966, 1037, 1035, 482, 813, 1397, 1383, 1402, 1959, 2909, 1968, 187, 319, 791, 2436, 1884, 1044, 1410, 918, 375, 1720, 2900, 2091, 2089, 1912, 2666, 2092, 7, 2057, 106, 1707, 1920, 1611, 1755, 326, 2029, 2720, 325, 1232, 1854, 1550, 1864, 316, 1109, 1189, 2545, 100, 101, 1212, 962, 2971, 90, 223, 1865, 824, 1938, 401, 306, 827, 2711, 1930, 1184, 1980, 1835, 304, 308, 350, 2100, 1689, 1036, 1695, 167, 1390, 2611, 1377, 310, 1914, 831, 1245, 2409, 2410, 2609]. System [1386, 856, 1682, 1363, 1871, 2099, 2599, 2101, 2351, 1447, 1181, 1180, 1162, 102, 1138, 485, 2102, 1984, 1429, 2706, 2479, 1771, 2843, 3114, 1876, 1834, 1977, 2292, 2721, 1921, 2275, 1544, 951, 3014, 1372, 1963, 1457, 915, 1943, 2718, 548, 749, 128, 746, 1933]. Systematic [276, 989, 2699, 700, 767, 2440, 560]. Systems [2380, 1380, 1808, 105, 1074, 1451, 2302, 1756, 2064, 1334, 89, 3119, 2388, 758, 365, 2381, 218, 1855, 104, 811, 2630, 1781, 2136, 812, 2244, 2097, 849, 1978, 2405, 707, 2289, 708, 2223, 775, 2049, 13, 2785, 2379, 1918, 94, 526, 2003, 590, 2404, 205, 1983, 939, 2876, 1860, 3130, 1703, 2067, 689, 584, 649, 1910, 657, 818, 206, 1881, 88, 1278, 79, 16, 1714, 2464, 709, 286, 1979, 287, 710, 3145, 1040, 247, 1811, 2352, 820, 712, 2846, 1357, 2614, 1329, 246, 2376, 2939, 2290, 1773, 214, 1754, 323, 224, 2637, 861, 2411, 381]. Systems [2121, 2332, 1812, 943, 823, 824, 1712, 870, 631, 2838, 763, 2653, 825, 2768, 3142, 2474, 1999, 95, 2120, 828, 696, 1974, 953, 2349, 2038, 466, 1719, 80, 830, 391, 955, 11, 1704, 2277, 2772, 2994, 305, 1536, 592, 244, 2492, 160, 1836, 2973, 2096, 1326, 1370, 1985, 1699, 998, 1057, 1347, 2417, 926, 309, 945, 781, 1885, 1964, 2954, 925, 778, 1198, 98, 82, 2982, 1065, 883, 2765, 1937, 767, 963, 1068, 2384, 2621, 55, 56, 601, 1367, 1067, 832, 833, 1311, 2428, 3014, 2287, 103, 2362, 835, 2775, 2781, 2137, 121, 558, 2229]. systems [563, 564]. Systolic [465]. SZK [2866, 2867]. Szymanski [780].

T [607]. T3E [1334, 1338, 1410, 1419].

T3E-900 [1419]. Table

[2085, 2529, 3136, 2183, 503, 251, 175, 1102]. Table-Lookup [251]. Tableau [1213, 1216, 1217, 1224, 1230, 1214]. Tableaux [1215, 1220, 1221, 2125, 1225, 1226, 1227]. Tables [2281, 3090, 2828, 1869, 2214, 1120, 2232]. Tabu [1684]. Tabular [1545, 859]. Tactical [1156]. Tactile [485]. Tagging [1239]. Tail [892]. Tailoring [932, 2989]. Takagi [2229]. Takagi-Sugeno [2229]. Talk [2113, 2086, 2119]. Tally [3137]. Tamagotchis [764]. Tamper [156, 155]. Tanaka [2507]. Tank [1325]. TANKA [2018]. Tape [3149, 2648, 3152, 2637, 3107]. Target [864, 3024, 3032, 1517]. Targeted [2187]. Tarski [333]. Tarskian [3150]. Task [2785, 495, 2573, 1912, 946, 1815, 1347, 1817, 1345]. Task-Oriented [495]. Tasking [1718, 1695]. Tasks [1883, 1806, 1342, 885, 1276, 883, 1834]. Taxonomy [1520, 1525, 1524]. TBSCM [1341]. Tcl [2060]. TCP [2345]. TD [1326]. Teaching [2648]. Team [1164, 1161, 1160, 1156, 1159, 1154, 1166, 1145, 1153, 1142, 1149, 1156]. Teams [2776, 1161, 1131, 360, 219]. Teamwork [360, 1132]. Technical [59, 1050]. Technique [2331, 311, 158, 1361, 591, 1226, 1425, 1119]. Techniques [583, 665, 1838, 1883, 157, 2361, 45, 1774, 1936, 1683, 1117, 1820, 544, 2399, 1678, 40, 1886, 891, 1637, 2014, 34, 1769, 1899, 1888, 65, 99, 1810, 893, 901, 189, 1932, 1944, 1306, 2239, 2651, 1625, 2134, 1569, 556, 1636, 742]. Technologies [2454, 2481, 2617]. Technology [1315, 207, 215, 260, 377, 1370, 1947, 354, 1805, 1]. Technology-Assisted [207]. Telecom [95]. Telecommunication [59, 2466, 2461, 2668]. Telecommunications [2656, 2660, 784, 2659, 8, 1, 2]. Telecontrol [2408]. Telecooperation [366]. TeleEEG [1322]. Telematic [2663]. Telemedical [1321]. Telemedicine [1322]. Telephone [993, 2032, 189]. Telephony [783]. Teleproximity [324]. Telerobotic [1930]. Television [2350]. Temperature [1922]. Tempering [1416]. Template [182]. Temporal [754, 1295, 1898, 1285, 1217, 1218, 94, 432, 1286, 1293, 1287, 1284, 3145, 1211, 1298, 1071, 2799, 2082, 2207, 844, 3109, 2477, 1290, 857, 1289, 1292, 1956, 2070, 838, 1297, 2373, 594, 1613, 1589, 2206, 1661]. Temporally [1280, 1927, 2042]. Ten [2353, 1710]. Tensor [1553, 1637, 1558, 1560, 1636, 1584]. Tensor-Based [1637, 1636]. Tensors [1552]. TENTACLE [1544]. Term [2932, 613, 2186, 3111, 3112, 3113, 1156, 622, 1225, 2492]. Term-Based [2492]. Termination [641, 639, 2115, 3110, 591, 640, 642]. Terminology [1002]. Terms [1079, 633, 2845, 484, 3036]. Ternary [759]. Terrarium [2582]. Tertiary [1690]. Test [3006, 2018, 447, 791, 889, 1792, 260, 2393, 647, 1425, 1072]. Test-Driving [2018]. Testability [2647]. Testable [616, 2554]. Testbed [2021, 2664, 2031, 315]. Testcase [1405]. Testing [768, 3066, 791, 792, 793, 755, 966, 2122, 925, 2475, 1709, 2961]. Text [2018, 528, 1518, 1254, 520, 1248, 1093, 997, 1631]. Texts [1093, 1876]. Textual [534, 2207, 1305, 2206]. Textures [976, 1591]. Their [2064, 2800, 710, 2227, 380, 666, 1358, 2180, 2978, 1292, 2522, 2181, 2220, 509, 2226, 2499, 2174]. Them [703]. Thematic [985, 2350]. Theorem [1204, 652, 2103, 2104, 2356, 1725, 2087, 2985, 2092, 2355, 3129, 285, 2834, 2966, 2117, 2118, 1727, 1234, 562]. Theorems [2194, 2088, 2844, 624, 142]. Theoretic [1216, 3129, 1247, 1795, 2119, 2778, 2142, 563].

Theoretical [1785, 489, 1014, 1280, 2599]. Theories [938, 243, 628, 121, 130]. Theory [606, 3106, 109, 1169, 2043, 1151, 2245, 37, 2905, 464, 1807, 2833, 2375, 3093, 1790, 281, 829, 603, 2926, 2294, 2247, 1685, 2648, 2649, 3151, 1686, 2631, 2766, 2282, 129, 1619]. Theory-Based [2043]. Therapy [1806]. Thermal [939]. Thinking [2074]. Thinned [2970]. Third [1979]. Thread [875]. Threaded [566, 2533]. Threading [1559]. Threads [270, 875, 1408]. Three [2170, 3104, 509, 2411, 1349, 2931, 2961, 1562, 1671]. Three-Dimensional [2931, 1562]. Three-Tier [2411]. Threshold [3066, 1182, 160, 2265, 151, 2554, 435, 1193, 1460]. Throughput [851]. Ticketing [2453]. Tier [2411]. tiered [564]. TIG [1878]. Tile [546]. Tiling [249, 250]. Tim [1940]. Time [754, 2567, 73, 2849, 89, 3090, 2022, 2899, 1299, 1407, 2405, 2403, 685, 1677, 1718, 1134, 1346, 1104, 10, 2857, 668, 1291, 771, 2404, 2828, 1288, 3130, 1703, 1713, 1743, 1196, 14, 1701, 3115, 1702, 1611, 2850, 2054, 958, 260, 3149, 374, 861, 962, 966, 2249, 90, 2803, 960, 797, 3095, 953, 2349, 294, 574, 955, 11, 91, 1704, 848, 956, 1115, 1796, 1326, 2543, 3082, 2276, 864, 2520, 945, 384, 92, 925, 198, 940, 445, 12, 963, 2165, 2467, 1771, 725, 93, 1355, 2070, 951, 2385, 994, 177, 2011, 1167, 2142]. Time-Based [994]. Time-Critical [12]. Time-Division [2011]. Time-Domain [1326]. Time-Series [1288]. Time-Space [2520]. Timed [770, 768, 608, 2388, 3125]. TimeGraph [2041]. TimeGraph-II [2041]. Timeless [1779, 368]. Timeliness [963]. Times [1140, 436]. Timetables [1681, 1684]. Timetabling [2015, 1687, 1673, 1683, 1675, 1689, 1680, 1690, 1682, 1679]. Timewheel [921]. Timing [2434, 664, 2187]. TINA [2471, 2476, 2462, 2461, 2459, 2474, 2482, 2475, 2452, 2463]. TINA-Based [2461]. TINA-like [2474]. TINA-oriented [2476]. TINA-TMN [2482]. Tinkerer [1031]. TKQML [248]. TMN [2473, 2453, 2482].

TMs [3149]. Toft [3080]. Token [2073, 173, 2063, 549]. Token-Based [173]. Token-Passing [2073]. Tolerance [917, 1843, 2265, 944]. Tolerant [922, 927, 923, 16, 964, 1700, 925, 928, 888, 924]. Tomasulo [2364]. Tool [2664, 1837, 2722, 2724, 2397, 30, 2405, 1397, 176, 248, 3029, 2404, 319, 404, 597, 1438, 7, 2355, 540, 950, 1271, 529, 725, 1710]. Toolbox [1031, 2646]. Toolkit [1548, 1545]. Tools [1432, 724, 170, 1322, 690, 602, 221, 2638, 2653, 513, 1827, 2, 747]. Toolset [2401, 306, 2402]. Top [2001]. Topic [212]. Topics [841, 403]. Topographic [982, 984]. Topological [3175, 1924, 1523]. Topologies [718]. Topology [3046, 876, 2007, 2931, 1629, 1621, 1630]. tori [1596]. Toroidal [922]. Torsion [2105, 2106]. Torsion-Free [2105, 2106]. Torus [1991]. Total [1635, 2845, 1634]. Totality [2869]. Totally [2183]. Tour [604]. Trace [1766, 1765]. Traceability [2719]. Traces [2824, 471, 2056]. Tracing [2612, 1470]. Trackies [1147]. Tracking [2612, 1611, 1610, 622, 1604, 382, 864, 2574, 1605, 1608, 1607, 1609, 1606, 1612]. TRACON [827]. Tractable [2034]. Trade [2009, 1457, 1446]. Trade-offs [2009]. Tradeoff [1645]. Tradeoffs [2520]. Trading [2343, 2165]. Traffic [1883, 498, 2953, 1244, 1912, 2611, 2954, 189, 2231, 177]. Trafficopter [2611]. Trail [1425]. Trailer [2995]. Trainable [2028]. Training [1954, 1955, 887, 846]. Trains [843]. Traitor [1470]. Trajectories [3140, 1613]. Tranforming [973]. Transaction [2077, 351, 1291, 1287, 2689]. Transactional [2456, 349]. Transactions [353, 1700, 2468, 149, 352]. Transcript [453]. Transcription [380]. Transducer [2640, 2642, 2553]. Transducer-Learning [2553]. Transducers [340, 2546, 2640, 620, 3118]. Transductions [3116]. Transfer [1018, 1315, 1487, 1245].

Transfinite [3157]. Transform

[930, 2045, 2257, 1097]. Transformation [641, 2108, 2109, 447, 710, 3145, 1040, 712, 3062, 2779, 555, 558]. Transformations [110, 614, 337]. Transformers [2150, 2149]. Transforming [2355, 2425]. Transforms [2258]. Transient [155]. Transition [772, 2064, 657, 820, 558, 1606, 564]. Transitioning [1290]. Transitive [443]. Transitivity [1238]. translating [743]. Translation [1209, 760, 2820]. Translations [2874]. Translator [1768]. Transmission [1995, 1885]. Transparency [2464, 1532, 1540, 873]. Transport [157, 2332, 314]. Transpose [1360]. Transposon [1425]. Traveling [1735, 1942]. Traversal [3091]. traversals [2138]. Traverse [257]. Treatment [1478, 1094]. Tree [641, 2823, 2133, 3163, 1731, 2557, 1784, 2558, 464, 2847, 3057, 479, 2834, 955, 454, 3082, 682, 1992, 1130, 3044, 1119, 2881, 2042, 2529]. Tree-Based [1992]. Tree-Driven [955]. Tree-Width [1731, 479]. Trees [651, 1253, 1382, 2278, 462, 461, 474, 453, 1239, 2277, 637, 1092, 2813, 2856, 3038, 672, 1941, 1095, 2280, 2150, 1260, 1252, 3084, 3039, 2149]. Trend [2217, 1078, 1810]. Trends [2693]. Trials [2985]. Triangles [2810]. Triangulation [3056]. Triangulations [3055]. Tridiagonal [1340]. Trifocal [1558, 1560, 1584]. Trigger [591]. Triggers [2079]. Trios [456]. Triple [420, 2197]. Trouble [2453]. Truck [2995, 2600]. Truck-and-Trailer [2995]. True [2849]. Truemper [1725]. TrueType [505]. Trust [2493]. Truth [3136]. Truth-Table [3136]. TS [1675]. TSQL2 [1290]. TTCN [2475]. TTCN/CORBA [2475]. TTN [1315]. Tuned [518]. Tuning [2273]. Tupling [688]. Turbine [2412]. Turbio [1876]. Turbulence [1337]. Turbulent [1325]. Turing [3136, 192]. Turning [357]. Tutorial [783]. Tutoring [1860, 1311].

Tutorship [1968]. TV [2331, 2321]. Twist [456]. Twist-Closed [456]. Two [1711, 679, 3086, 1335, 3090, 1875, 610, 3058, 2524, 2501, 2283, 424, 2828, 1023, 3042, 1070, 664, 395, 3077, 517, 381, 1440, 617, 1907, 1024, 1905, 883, 251, 2880, 2583, 2854, 1517, 2883, 1663, 124]. Two-Dimensional [679, 3086, 1023, 3042, 3077, 1024, 2583, 1517]. Two-Handed [395]. Two-Level [1711, 617, 251]. Two-Party [3090]. Two-Processor [883]. Two-Rowed [2828]. two-stage [1663]. Two-Variable [2524]. Two-Way [2854]. TWOPRIME [407]. Type [338, 606, 3119, 687, 512, 339, 2877, 27, 1676, 2923, 2833, 691, 510, 1797, 2878, 28]. Type-Inference [687]. Typed [2916, 2943, 3145, 2917, 1795, 2697]. Typeface [507]. Typemaking [511]. Types [2920, 2116, 38, 544, 694, 2826, 695, 564]. Typicality [1516]. Typing [339]. Typographic [502, 503, 517]. Typography [515].

U [2004, 2005]. U-Net [2004, 2005]. Ubiquity [568]. Ultimate [3061]. Ultrafast [900]. Ultrametric [3084]. Ultrasonic [1888]. UltraSparc [181]. UMDL [2613]. UML [2687, 47, 58, 48]. UML’s [46]. UMTS [2485]. Unary [1786, 443, 439]. Unbiased [3065]. Unbounded [2380]. uncalibrated [1569, 1555]. Uncapacitated [1734]. Uncertain [1841, 1296, 2214]. Uncertainty [1846, 2822, 1829, 2990, 1098, 1666, 132]. Unconditionally [1191]. Undecidability [2807]. Undecidable [626]. Underlying [1806]. Understanding [1518, 377, 2235, 2553, 972, 1887, 2240, 1484]. Unexpected [1295]. Unexpectedness [1079]. Unfolding [2069, 656]. Unfoldings [2372]. Unification [2098, 38, 636, 629, 628, 626, 2114, 630, 3151]. Unified [275, 2905, 2900, 503, 2821, 2883]. Uniform [2672, 649, 472, 3146, 2114, 3065].

Uniformly [1749, 478]. Unifying [1612, 2454, 2139]. Uninterpreted [2375, 2376]. Union [2538]. Unit [2202, 1617]. Units [862]. Unity [931, 932, 936, 933]. Universal [201, 1816, 1013]. Universally [1492]. Universe [1975]. University [1689, 1682, 1684, 1147]. University-Wide [1682]. Unix [2672, 1920]. Unknown [2170, 1119, 1564]. Unsolved [378]. Unstructured [1418, 1328]. Unsupervised [1955, 1093]. Untrusted [2053]. Unweighted [1731]. Unwrapping [1416]. Up/Down [2681]. Upcast [2813]. Update [1286, 2159, 327, 1297, 133]. Updates [1540]. upon [2265]. Upper [1738, 392, 393, 2520]. Urban [2601, 1912, 1578]. Urquhart [3106]. Ursa [259, 1370]. URW [517]. Usage [2450, 1049, 2466, 1934, 254]. Use [535, 1560, 2076, 508, 372, 3096, 1719, 388, 1772, 1682, 2468, 883, 1937, 2007, 1050, 1935, 1517, 2901, 257]. Used [2680, 1001, 1944]. Useful [429]. Usefulness [1939]. User [910, 1037, 2043, 1108, 1769, 330, 332, 165, 795, 796, 358, 1181, 225, 2457, 2706, 1073]. User-Level [910]. User-Oriented [1769]. Users [1053, 2334, 2457]. Using [3031, 108, 1075, 3003, 1152, 639, 2672, 386, 59, 515, 2388, 1295, 1855, 2633, 1846, 1356, 528, 1378, 1406, 1901, 985, 1437, 2995, 2970, 1936, 1408, 1161, 54, 1592, 2379, 512, 2909, 2383, 2935, 2273, 1775, 1683, 3029, 1668, 771, 791, 1843, 1080, 939, 2561, 1653, 2230, 2936, 1576, 2088, 1023, 2996, 434, 1083, 1412, 1881, 2190, 2578, 1631, 1701, 170, 757, 3002, 2212, 965, 487, 961, 1970, 488, 1702, 1755, 326, 2470, 32, 2717, 1849, 1869, 1086, 2969, 2674, 3074, 246, 1813, 2939, 1109, 1211, 779, 1071, 178, 1388, 2454, 2008, 2941, 2032, 1732, 2317]. Using [1110, 2997, 1535, 2125, 1824, 957, 763, 1654, 1529, 828, 2329, 1588, 1238, 1239, 934, 1163, 402, 2010, 1915, 15, 1135, 2159, 2940, 2053, 1271, 382, 1900, 1821, 3019, 130, 990, 2581, 1024, 859, 147, 2599, 384, 1885, 1964, 887, 2972, 682, 886, 2645, 1243, 255, 331, 2326, 800, 801, 1922, 506, 2880, 1992, 2363, 929, 2980, 2369, 756, 1130, 3044, 2706, 1132, 1171, 1072, 1685, 400, 1594, 1256, 2372, 86, 87, 1686, 2661, 1306, 2373, 2502, 2231, 177, 2780, 1167, 1659, 1602, 1617, 1556, 1560, 1656, 2501, 551, 1593, 1666, 563, 1615, 1633]. Utility [2043]. Utilizing [1377].

V [49]. Valiant [444]. Valid [1291]. Validation [931, 2331, 959, 2939, 795, 796, 760, 2820, 778, 2037]. Validations [1710]. validity [120]. Valuations [611, 1228]. Value [1776, 2925, 1789, 789, 1030, 848, 1141]. Value-Added [789]. Value-Passing [1789]. Valued [1220, 1221, 820]. Values [1094, 1119]. Valuing [1910]. Vanstone [161]. Variability [2439]. Variable [583, 447, 2524, 636, 1225, 3155, 2541, 3156, 2980]. Variable-Dimensional [2980]. Variables [1738, 1830]. Variance [1394, 1645]. Variant [2437, 540]. Variants [619]. Variation [1062]. Varieties [655]. Varying [2238, 3036, 2229]. VBR [2346]. Vector [1406, 999, 1254, 3034, 1404, 2257, 1255, 1405, 1646]. Vectorization [968, 972]. Vectors [776, 2176, 777]. Vehicle [2022, 2355, 1029, 878, 1606]. Vehicles [1981, 826]. Velocity [385, 516, 1874]. Verbal [1238]. Verfication [825]. Verifiable [686, 1492, 2490, 1462]. Verification [2386, 2090, 2387, 663, 2388, 2397, 2361, 2396, 2403, 764, 1200, 750, 766, 765, 2395, 1201, 791, 2392, 2390, 138, 755, 780, 712, 821, 753, 2939, 2113, 2799, 3094, 957, 763, 827, 828, 2364, 1604, 2393, 831, 948, 151, 2119, 781, 2366, 800, 801, 2363, 352, 2070, 2775, 1492, 1477, 1657]. Verified [761]. Verifier [2369]. Verifiers [151]. Verifying [770, 2398, 930, 566, 2362]. Versatile [1352]. Version [2712, 2722, 1298, 2123, 701]. Versioned [2725, 2723]. Versioning [2724, 2718].

Versus [3012, 625, 1405, 2925]. Vertex [452, 2991, 2888]. Vertices [2538]. Very [2348, 3037]. VFC [1398]. VI [50]. Via [1954, 1882, 768, 2108, 2109, 1476, 2809, 815, 1787, 1426, 735, 173, 2813, 697, 954, 1570]. Video [386, 2344, 2350, 787, 389, 222, 2349, 1995, 2319, 2346, 1589, 1555, 1621]. Video-Based [386]. Videoconference [2317]. Videos [2577]. Videotransmission [965]. Vienna [1398]. View [1552, 344, 1816, 592, 1753, 1549, 1608, 593, 1607, 555, 1606, 1669, 2260, 1662]. View-Based [1608, 1607]. Viewing [2574, 484, 1573]. viewpoint [1555, 1638]. Viewpoints [2585]. Views [989, 343, 573, 1533, 2442, 594, 1661, 1577, 1671, 1570]. VII [62]. VIII [72]. Violating [330]. Violations [1826]. Virtual [2567, 3092, 2920, 1862, 2909, 1354, 2573, 852, 2655, 325, 2593, 1426, 788, 2594, 2598, 363, 2602, 2323, 394, 2566, 1529, 2606, 321, 2590, 1395, 1695, 2605, 2575, 872, 2589, 537, 2576, 364, 2581, 2599, 2587, 2588, 2574, 2604, 2569, 2571, 2585, 865, 2007, 211, 2583, 538, 2568, 3083, 1530]. Visible [2514]. Vision [1931, 1883, 1134, 1668, 2591, 1930, 2592, 1143, 2572, 1347, 1923, 1937, 1147, 140, 1564, 370]. Vision-Based [1143, 2572, 1923, 1147, 1564]. Visual [1356, 2343, 1592, 1593, 1967, 1087, 2249, 2902, 1430, 1658, 797, 1980, 1323, 388, 800, 801, 1474, 26, 1594, 492, 1629, 738, 1665, 1649, 505]. Visualisation [795, 796, 2581, 525]. Visualising [798]. Visualization [2633, 3018, 961, 1087, 799, 953, 1355, 1589, 1641]. Vital [907, 2482]. Vitro [2636]. VLIW [742]. VLSI [2036, 2243]. VLSI-Friendly [2036]. Volunteer [331]. Voronoi [459, 2537]. Voting [1601, 148, 1600]. Voyager [303]. VQL [342]. VR [2603, 2601, 1413, 2599]. VRML [2317, 2570]. VRML2 [2572]. vs [3135, 3171, 23, 2618, 3138, 3139, 1638, 2606, 2119]. Vulnerability [943].

Walks [426]. Wall [1925]. Wallaid [1984]. Walls [1984]. Warehouses [588, 1758, 1757]. Warehousing [577, 594]. Water [1963]. Watermarking [2331]. Wave [1406]. Wavelength [3170, 2847]. Wavelet [2258, 2253, 1097, 1639, 2252]. Wavelets [2269, 2270]. Way [3174, 424, 1981, 368, 2854, 2503, 1485]. wayness [1484]. Weak [1783, 13, 1468, 2859, 2522]. Weakly [1722, 1655]. Weakness [161]. Weaknesses [1946]. WEB [1108, 598, 1346, 1266, 539, 2061, 1106, 1044, 597, 2628, 2623, 329, 2674, 1769, 782, 318, 2453, 573, 2580, 2062, 1370, 720, 331, 2321, 2625, 599, 1544]. Web-Based [331]. Webots [2591, 2592, 2590]. WebSemantics [574]. Webtutor [1968]. WeBUSE [220]. Weight [311, 2689, 1952, 2813, 3056]. Weighted [2510, 2841, 2526, 1361, 1535, 2642, 2537]. Weights [1855]. Welding [2230]. Well [1058, 2067, 657, 2213]. Well-Behaviored [2213]. Well-Structured [657]. Were [197]. wheelchair [1524, 1525]. Where [1015, 302]. Which [2982]. WHILE [605]. White [2578, 782]. Whose [666]. WIA [2631]. Wide [1682, 2625, 3075, 1044]. Wide-Sense [3075]. Width [433, 1731, 479]. Will [3168]. Window [502, 490]. Window-Based [502]. Windows [2022, 2004, 2005]. Windows/NT [2004, 2005]. Wings [197]. Wire [2230]. Wireless [2066, 2351, 928]. Within [1528, 1850, 2982, 1315, 1368, 1402, 2464, 1069]. Without [3122, 2069, 152, 330]. Witness [1173]. Wizard [1847]. Word [458, 3113, 2937, 454, 2861, 3038]. Words [652, 3158, 472, 3159, 2835, 3160, 2836, 2222, 3161, 2027, 655, 440, 1803]. Work [218, 2785, 2573, 2666, 1967, 371, 376, 369, 1245, 1300, 1492]. Workbench [536, 1208, 1755]. Workflow [600, 1454, 214, 216, 363, 1761, 1455, 82, 601, 1453]. Workflows [1760]. Working [373, 3096, 198, 355]. Workloads [1757].

Workplace [371]. Workshop [50, 88]. Workspace [219]. Workspaces [1962]. Workstation [1375, 873]. Workstations [166, 163, 167, 181, 1415, 2007]. World [1940, 154, 96, 996, 1279, 2598, 2249, 2566, 321, 848, 2575, 2583, 123, 1044, 2625]. World-Wide [1044]. Worlds [1236, 2593, 2602, 2606, 2605, 2581, 2584, 211, 2579]. Worldwide [304]. Wormhole [1991, 2013, 904, 1009, 2006]. Wormhole-Routed [1991]. Worms [2008, 1992]. Worst [2518, 2531, 3138, 3139]. Worst-Case [2518, 2531, 3138, 3139]. Wrappers [2914]. Write [2354]. Written [2706]. WS1S [2399]. WS2S [1777, 2399]. WWC [299]. WWW [1441, 326, 1321, 220]. WWW-based [1321]. WWW-Database [220].

X [89, 490, 3104, 1243]. X-Ray [1243]. X-Rays [3104]. X.R.S. [2096]. X9.52 [1467]. Xeve [2396]. XI [97]. XII [98]. Xpect [1451].

Years [2001, 2353, 1710]. Yield [2162].

Z [707, 708, 39, 709, 716]. Z-based [39]. Zantema [1785]. Zernike [2296]. Zero [549, 2865, 469]. Zero-Knowledge [2865, 469]. Zero-safe [549]. Zeuxis [2606]. Zhang [2707]. Ziv [2861]. Zone [1329].

References

Znaty:1998:OOT

[1] Simon Znaty and Jean-Pierre Hubaux. I: Object-oriented technology for telecommunications services engineering. Lecture Notes in Computer Science, 1357:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                              link.springer-ny.com/link/service/

series/0558/bibs/1357/13570001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1357/13570001.pdf.

Znaty:1998:TSE

[2] Simon Znaty and Jean-Pierre Hubaux. Telecommunications services engineering: Principles architectures and tools. Lecture Notes in Computer Science, 1357:3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1357/13570003.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1357/13570003.
pdf.

Meisingset:1998:DMO

[3] Arve Meisingset. Difficulties with mapping OMT specifications into GDMO. Lecture Notes in Computer Science, 1357:12–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1357/13570012.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1357/13570012.
pdf.

Sidou:1998:TGF

[4] D. Sidou. Towards a “good” functional and executable behavior model. Lecture Notes in Computer Science, 1357: 17–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570017.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1357/13570017.
                                        pdf.

Portas:1998:FSN

[5] Jorge L. Tellez Portas, David Watrin, and Tayeb Ben Meriem. The formalization and simulation of a network management model. Lecture Notes in Computer Science, 1357:23– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570023.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570023.
                                        pdf.

TellezPorta:1998:FSN

[6] J. L. Tellez Porta, D. Watrin, and T. Ben Meriem. The formalization and simulation of a network management model. Lecture Notes in Computer Science, 1357:23–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Granville:1998:OOO

[7] Lisandro Granville, Luciano Gaspary, and Janilce Almeida. OST — an object-oriented computer networks system specification tool. Lecture Notes in Computer Science, 1357:29– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570029.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570029.
                                        pdf.

Harju:1998:ETP

[8] Jarmo Harju, Bilhanan Silverajan, and Ilkka Toivanen. Experiences in telecommunications protocols with an OO based implementation framework. Lecture Notes in Computer Science, 1357:34–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570034.
pdf.

Deri:1998:RNM

[9] Luca Deri. Rapid network management application development. Lecture Notes in Computer Science, 1357:39– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570039.
pdf.

Choukair:1998:RTO

[10] Zièd Choukair and Antoine Beugnard. Real-time object-oriented distributed processing with COREMO. Lecture Notes in Computer Science, 1357:45– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570045.
pdf.

Mitchell:1998:IRR

[11] S. E. Mitchell and R. J. Stroud. II: Reflective real-time object-oriented programming and systems. Lecture Notes in Computer Science, 1357:51– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570051.
                                        pdf.

Singhai:1998:ROS

[12] Ashish Singhai, Aamod Sane, and Roy Campbell. Reflective ORBs: Supporting robust, time-critical distribution. Lecture Notes in Computer Science, 1357:55–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1357/13570055.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570055.
                                        pdf.

Chiba:1998:WPR

[13] Shigeru Chiba, Takeshi Nishimura, Kenichi Kourai, Atsushi Ohnoki, and Takashi Masuda. Weak protection for reflective operating systems. Lecture Notes in Computer Science, 1357:62– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570062.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570062.
                                        pdf.

Golm:1998:IRT

[14] Michael Golm and Jürgen Kleinöder. Implementing real-time actors with MetaJava. Lecture Notes in Computer Science, 1357:68–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1357/13570068.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1357/13570068.pdf.

Mitchell:1998:ASU

[15] S. E. Mitchell, A. Burns, and A. J. Wellings. Adaptive scheduling using reflection. Lecture Notes in Computer Science, 1357:74–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570074.
pdf.

Goldberg:1998:AFT

[16] Jack Goldberg and Robert J. Stroud. Adaptive fault-tolerant systems and reflective architectures. Lecture Notes in Computer Science, 1357:80–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570080.
pdf.

Bosch:1998:ILS

[17] Jan Bosch, Görel Hedin, and Kai Koskimies. III: Language support for design patterns and frameworks. Lecture Notes in Computer Science, 1357: 89–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570089.
                                        pdf.

Bosch:1998:LSD

[18] J. Bosch, G. Hedin, and K. Koskimies. Language support for design patterns and frameworks. Lecture Notes in Computer Science, 1357:89–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agerbo:1998:IGD

[19] Ellen Agerbo and Aino Cornils. Implementing GoF design patterns in BETA. Lecture Notes in Computer Science, 1357:92–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1357/13570092.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570092.
                                        pdf.

Ducasse:1998:MPA

[20] Stéphane Ducasse. Message passing abstractions as elementary bricks for design pattern implementation: An experiment. Lecture Notes in Computer Science, 1357:96–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570096.
pdf.

Eden:1998:PA

[21] Amnon H. Eden and Amiram Yehudai. Patterns of the agenda. Lecture Notes in Computer Science, 1357:100– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570100.
pdf.

Jacobsen:1998:DPP

[22] Eyðun Eli Jacobsen. Design patterns as program extracts. Lecture Notes in Computer Science, 1357:105– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570105.
pdf.

Gil:1998:DPV

[23] Joseph Gil and David H. Lorenz. Design patterns vs. language design. Lecture Notes in Computer Science, 1357: 108–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570108.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570108.
                                        pdf.

Matthijs:1998:MLP

[24] Frank Matthijs, Wouter Joosen, Bert Robben, Bart Vanhaute, and Pierre Verbaeten. Multi-level patterns. Lecture Notes in Computer Science, 1357: 112–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570112.
                                        pdf.

Nowack:1998:AAF

[25] Palle Nowack. Architectural abstractions for frameworks. Lecture Notes in Computer Science, 1357:116–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570116.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570116.
                                        pdf.

Tuovinen:1998:FPV

[26] Antti-Pekka Tuovinen. A framework for processors of visual languages. Lecture Notes in Computer Science, 1357: 119–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570119.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570119.
                                        pdf.

DeVolder:1998:TOP

[27] K. De Volder and W. De Meuter. Type oriented programming. Lecture Notes in Computer Science, 1357:123– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volder:1998:TOP

[28] Kris De Volder and Wolfgang De Meuter. Type oriented programming. Lecture Notes in Computer Science, 1357:123–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1357/13570123.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1357/13570123.
pdf.

Peres:1998:FRL

[29] Yardena Peres, Jerry W. Malcolm, Pnina Vortman, and Gabi Zodik. A framework registration language. Lecture Notes in Computer Science, 1357:126– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570126.
pdf.

Bjarnason:1998:TSF

[30] Elizabeth Bjarnason and Görel Hedin. Tool support for framework-specific language extensions. Lecture Notes in Computer Science, 1357:129–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570129.
pdf.

Bosch:1998:DPF

[31] Jan Bosch. Design patterns & frameworks: On the issue of language support. Lecture Notes in Computer Science, 1357:133–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570133.
pdf.

Hedin:1998:LSD

[32] Görel Hedin. Language support for design patterns using attribute extension. Lecture Notes in Computer Science, 1357:137–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1357/13570137.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1357/13570137.
pdf.

Harsu:1998:LIF

[33] Maarit Harsu, Juha Hautamäki, and Kai Koskimies. A language implementation framework in Java. Lecture Notes in Computer Science, 1357:141– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1357/13570141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570141.
pdf.

Kilov:1998:IPS

[34] Haim Kilov and Bernhard Rumpe. IV: Precise semantics for object-oriented modeling techniques. Lecture Notes in Computer Science, 1357:145–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570145.
pdf.

Bezivin:1998:OBL

[35] Jean Bézivin and Richard Lemesle. Ontology-based layered semantics for precise OA&D modeling. Lecture Notes in Computer Science, 1357:151– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570151.
pdf.

Bicarregui:1998:FOO

[36] J. C. Bicarregui, K. C. Lano, and T. S. E. Maibaum. Formalizing object-oriented models in the object calculus. Lecture Notes in Computer Science, 1357:155–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1357/13570155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570155.
                                        pdf.

Cohen:1998:STS

[37] B. Cohen. Set theory as a semantic framework for object-oriented modeling. Lecture Notes in Computer Science, 1357:161–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1357/13570161.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1357/13570161.pdf.

Daugherty:1998:UMT

[38] Gary W. Daugherty. Unification of the models for types, classes and state machines. Lecture Notes in Computer Science, 1357:169–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570169.
                                        pdf.

Ebert:1998:IZB

[39] Jürgen Ebert and Roger Süttenbach. Integration of Z-based semantics of OO-notations. Lecture Notes in Computer Science, 1357:173–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570173.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1357/13570173.
pdf.

France:1998:IIO

[40] Robert B. France and Jean-Michel Bruel. Integrated informal objectoriented and formal modeling techniques. Lecture Notes in Computer Science, 1357:178–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1357/13570178.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1357/13570178.pdf.

Kent:1998:STP

[41] Stuart Kent, Ali Hamie, John Howse, Franco Civello, and Richard Mitchell. Semantics through pictures. Lecture Notes in Computer Science, 1357:182– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570182.
pdf.

Kilov:1998:BRB

[42] Haim Kilov and Ian Simmonds. Business rules: From business specification to design. Lecture Notes in Computer Science, 1357:188–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570188.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1357/13570188.
pdf.

Kurki-Suonio:1998:LOO

[43] Reino Kurki-Suonio and Tommi Mikkonen. Liberating object-oriented modeling from programming-level abstractions. Lecture Notes in Computer Science, 1357:195–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1357/13570195.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1357/13570195.pdf.

Mullery:1998:DDA

[44] Geoff Mullery. Discovering, displaying and agreeing semantics in an environment of conflicts. Lecture Notes in Computer Science, 1357:200–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570200.
pdf.

Breu:1998:TPS

[45] Ruth Breu, Radu Grosu, Franz Huber, Bernhard Rumpe, and Wolfgang Schwerin. Towards a precise semantics for object-oriented modeling techniques. Lecture Notes in Computer Science, 1357:205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570205.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1357/13570205.
pdf.

Schuerr:1998:FDR

[46] A. Schuerr and A. J. Winter. Formal definition and refinement of UML’s module/package concept. Lecture Notes in Computer Science, 1357:211– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schurr:1998:FDR

[47] Andy Schürr and Andreas J. Winter. Formal definition and refinement of UML’s module/package concept. Lecture Notes in Computer Science, 1357: 211–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570211.
pdf.

Warmer:1998:EFS

[48] Jos Warmer, John Hogg, Steve Cook, and Bran Selic. Experience with formal specification of CMM and UML. Lecture Notes in Computer Science, 1357: 216–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570216.
pdf.

Bastide:1998:VMF

[49] Rémi Bastide and Didier Buchs. V: Models, formalisms and methods for object-oriented distributed computing. Lecture Notes in Computer Science, 1357:221–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1357/13570221.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1357/13570221.
pdf.

Casais:1998:VFW

[50] Eduardo Casais, Ari Jaaksi, and Thomas Lindner. VI: FAMOOS workshop on object-oriented software evolution and re-engineering. Lecture Notes in Computer Science, 1357:257– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570257.
pdf.

Kantorowitz:1998:IEC

[51] Eliezer Kantorowitz. On implementation and extension complexities. Lecture Notes in Computer Science, 1357: 264–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570264.
pdf.

Richner:1998:ADS

[52] Tamar Richner and Robb Nebbe. Analyzing dependencies to solve lowlevel problems. Lecture Notes in Computer Science, 1357:266–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570266.
pdf.

Ruonavaara:1998:AAD

[53] Markku Ruonavaara. Adding architecture to design. Lecture Notes in Computer Science, 1357:268–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570268.
pdf.

Ciupke:1998:AOO

[54] Oliver Ciupke. Analysis of objectoriented programs using graphs. Lecture Notes in Computer Science, 1357: 270–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570270.
pdf.

Systa:1998:ESD

[55] Tarja Systä and Kai Koskimies. Extracting state diagrams from legacy systems. Lecture Notes in Computer Science, 1357:272–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1357/13570272.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1357/13570272.pdf.

Systae:1998:ESD

[56] T. Systae and K. Koskimies. Extracting state diagrams from legacy systems. Lecture Notes in Computer Science, 1357:272–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mens:1998:CSD

[57] Kim Mens, Tom Mens, Patrick Steyaert, and Koen De Hondt. Combining software descriptions. Lecture Notes in Computer Science, 1357:274– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570274.
pdf.

Trauter:1998:REU

[58] Roland Trauter. Re-engineering with UML. Lecture Notes in Computer Science, 1357:276–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1357/13570276.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1357/13570276.pdf.

Bachatene:1998:ETS

[59] Hélène Bachatène and Pierre Arnaud. Evolution of telecommunication software using OO: Organizational and technical issues. Lecture Notes in Computer Science, 1357:278– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570278.
pdf.

Demeyer:1998:DPR

[60] Serge Demeyer, Theo-Dirk Meijler, and Matthias Rieger. Design pattern restructuring. Lecture Notes in Computer Science, 1357:280–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570280.
pdf.

Sarirete:1998:SMO

[61] Akila Sarirete and Jean Vaucher. Similarity measures for the object model. Lecture Notes in Computer Science, 1357:283–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1357/13570283.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1357/13570283.
pdf.

Berg:1998:VMS

[62] Klaas van den Berg, Mehmet Aksit, Pim van den Broek, Leon Osterweil, Karl Lieberherr, and Francesco Marcelloni. VII: Modeling software processes and artifacts. Lecture Notes in Computer Science, 1357:285–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570285.
                                        pdf.

VandenBerg:1998:MSP

[63] K. Van den Berg, M. Aksit, P. Van den Broek, and L. Osterwel. Modeling software processes and artifacts. Lecture Notes in Computer Science, 1357:285– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1998:RPP

[64] Reidar Conradi and Chunnian Liu. Revised PMLs and PSEEs for industrial SPI. Lecture Notes in Computer Science, 1357:289–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1357/13570289.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1357/13570289.pdf.

Marcelloni:1998:AFL

[65] Francesco Marcelloni and Mehmet Aksit. Applying fuzzy logic techniques in object-oriented software development.

Lecture Notes in Computer Science, 1357:295–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1357/13570295.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1357/13570295.
pdf.

Reimer:1998:TDO

[66] Wiebke Reimer, Wilhelm Schäfer, and Thomas Schmal. Towards a dedicated object-oriented software process modeling language. Lecture Notes in Computer Science, 1357:299–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570299.
pdf.

Hruby:1998:OMP

[67] Pavel Hruby. The object model for a product based development process. Lecture Notes in Computer Science, 1357:303–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1357/13570303.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1357/13570303.
pdf.

Aksit:1998:ASA

[68] Mehmet Aksit, Francesco Marcelloni, Bedir Tekinerdogan, Klaas van den Berg, and Pim van den Broek. Active software artifacts. Lecture Notes in Computer Science, 1357:307–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570307.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570307.
                                        pdf.

Bezivin:1998:GSP

[69] Jean Bézivin. General software process organization in the OS-MOSIS project. Lecture Notes in Computer Science, 1357:311–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570311.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570311.
                                        pdf.

Han:1998:OOM

[70] Jun Han. Object-oriented modeling of software processes and artifacts: Promises and challenges. Lecture Notes in Computer Science, 1357: 315–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570315.
                                        pdf.

Bider:1998:OOM

[71] I. Bider and M. Khomyakov. Objectoriented model for representing software production processes. Lecture Notes in Computer Science, 1357:319– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570319.
pdf.

Bosch:1998:VCO

[72] Jan Bosch, Clemens Szyperski, and Wolfgang Weck. VIII: Componentoriented programming. Lecture Notes in Computer Science, 1357:323–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570323.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570323.
pdf.

Baggiolini:1998:TAR

[73] Vito Baggiolini and Jürgen Harms. Toward automatic, run-time fault management for component-based applications. Lecture Notes in Computer Science, 1357:327–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1357/13570327.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1357/13570327.pdf.

Buchi:1998:FMC

[74] Martin Büchi and Emil Sekerinski. Formal methods for component software: The refinement calculus perspective. Lecture Notes in Computer Science, 1357:332–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570332.
                                        pdf.

Buechi:1998:FMC

[75] M. Buechi and E. Sekerinski. Formal methods for component software: The refinement calculus perspective. Lecture Notes in Computer Science, 1357: 332–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeHondt:1998:RCC

[76] K. De Hondt, C. Lucas, and P. Steyaert. Reuse contracts as component interface descriptions. Lecture Notes in Computer Science, 1357:338– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hondt:1998:RCC

[77] Koen De Hondt, Carine Lucas, and Patrick Steyaert. Reuse contracts as component interface descriptions. Lecture Notes in Computer Science, 1357: 338–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570338.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570338.
                                        pdf.

Lalanda:1998:CMD

[78] Philippe Lalanda. A control model for the dynamic selection and configuration of software components. Lecture Notes in Computer Science, 1357:343– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570343.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570343.
pdf.

Goedicke:1998:DED

[79] Michael Goedicke and Torsten Meyer. Design and evaluation of distributed component-oriented software systems. Lecture Notes in Computer Science, 1357:348–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1357/13570348.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1357/13570348.
pdf.

Mikhajlov:1998:FBC

[80] Leonid Mikhajlov and Emil Sekerinski. The fragile base class problem and its impact on component systems. Lecture Notes in Computer Science, 1357: 353–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570353.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570353.
pdf.

Murer:1998:CGS

[81] Tobias Murer. The challenge of the global software process. Lecture Notes in Computer Science, 1357:359– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570359.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570359.
                                        pdf.

Schreyjak:1998:CWC

[82] Stefan Schreyjak. Coupling of workflow and component-oriented systems. Lecture Notes in Computer Science, 1357: 364–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570364.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570364.
                                        pdf.

Steensgaard-Madsen:1998:GCI

[83] Jørgen Steensgaard-Madsen. A generator for composition interpreters. Lecture Notes in Computer Science, 1357: 369–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570369.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570369.
                                        pdf.

Troya:1998:APC

[84] Jose M. Troya and Antonio Vallecillo. On the addition of properties to components. Lecture Notes in Computer Science, 1357:374–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570374.
pdf.

Bosch:1998:AOO

[85] Jan Bosch. Adapting objectoriented components. Lecture Notes in Computer Science, 1357:379–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570379.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570379.
pdf.

Weck:1998:IUC

[86] Wolfgang Weck. Inheritance using contracts & object composition. Lecture Notes in Computer Science, 1357:384– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570384.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570384.
pdf.

Week:1998:IUC

[87] W. Week. Inheritance using contracts and object composition. Lecture Notes in Computer Science, 1357:384–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerhardt:1998:IWP

[88] Frank Gerhardt, Lutz Wohlrab, and Erik Ernst. IX: The 7th workshop for PhD students in objectoriented systems. Lecture Notes in Computer Science, 1357:389–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570389.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570389.
                                        pdf.

Barroca:1998:XOO

[89] Leonor Barroca, Eugene Durr, and François Terrier. X: Object-oriented real-time systems. Lecture Notes in Computer Science, 1357:433–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570433.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570433.
                                        pdf.

Lano:1998:FRT

[90] K. C. Lano and S. Goldsack. Formalizing real-time system design. Lecture Notes in Computer Science, 1357:436– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570436.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570436.
                                        pdf.

Monteverde:1998:OOR

[91] Jorge Amador Monteverde. Objectoriented real-time software in the European Space Agency. Lecture Notes in Computer Science, 1357:446–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570446.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570446.
pdf.

Rioux:1998:SME

[92] Laurent Rioux, Patrick Vanuxeem, and François Terrier. Scheduling mechanisms for efficient implementation of real-time objects. Lecture Notes in Computer Science, 1357:450– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570450.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570450.
pdf.

Toinard:1998:OCA

[93] Christian Toinard and Nicolas Chevassus. An object communication architecture for real-time distributed process control. Lecture Notes in Computer Science, 1357:460–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570460.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570460.
pdf.

Cornilleau:1998:ROB

[94] T. Cornilleau and E. Gressier-Soudan. RT-objects based on temporal causal consistency: A new approach for fieldbus systems. Lecture Notes in Computer Science, 1357:464–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570464.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570464.
                                        pdf.

Maisonneuve:1998:CRC

[95] Julien Maisonneuve. Configuring realtime CORBA systems for telecom applications. Lecture Notes in Computer Science, 1357:468–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570468.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570468.
                                        pdf.

Clohessy:1998:NCE

[96] Kim Clohessy, Brian Barry, and Peter Tanner. New complexities in the embedded world — the OTI approach. Lecture Notes in Computer Science, 1357:472–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570472.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570472.
                                        pdf.

Mens:1998:XAO

[97] Kim Mens, Cristina Lopes, Bedir Tekinerdogan, and Gregor Kiczales. XI: Aspect-oriented programming. Lecture Notes in Computer Science, 1357: 483–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570483.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570483.
pdf.

Schmidt:1998:XOO

[98] Henning Schmidt, Frank Schubert, and Lutz Wohlrab. XII: Objectorientation and operating systems. Lecture Notes in Computer Science, 1357: 497–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570497.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570497.
pdf.

Nebro:1998:ADS

[99] Antonio J. Nebro, Ernesto Pimentel, and José M. Troya. Applying distributed shared memory techniques for implementing distributed objects. Lecture Notes in Computer Science, 1357: 499–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570499.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570499.
pdf.

Kleinoder:1998:MPA

[100] Jürgen Kleinöder and Michael Golm. MetaJava — A platform for adaptable operating-system mechanisms. Lecture Notes in Computer Science, 1357:507– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570507.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570507.
                                        pdf.

Kleinoeder:1998:MPA

[101] J. Kleinoeder and M. Golm. MetaJava — A platform for adaptable operatingsystem mechanisms. Lecture Notes in Computer Science, 1357:507–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schubert:1998:RAA

[102] Frank Schubert. A reflective architecture for an adaptable object-oriented operating system based on C++. Lecture Notes in Computer Science, 1357: 515–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570515.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570515.
                                        pdf.

Wohlrab:1998:CAM

[103] Lutz Wohlrab. Configuration and adaptation management for objectoriented operating systems. Lecture Notes in Computer Science, 1357:523– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570523.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570523.
pdf.

Beuche:1998:AMH

[104] Danilo Beuche. An approach for managing highly configurable operating systems. Lecture Notes in Computer Science, 1357:531–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1357/13570531.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1357/13570531.
pdf.

AlvarezGuitierrez:1998:OOA

[105] D. Alvarez Guitierrez, L. Tajes Martinez, F. Alvarez Garcia, and M. A. Diaz Fondon. An object-oriented abstract machine as the substrate for an object-oriented operating systems. Lecture Notes in Computer Science, 1357: 537–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gutierrez:1998:OOA

[106] Darío Álvarez Gutiérrez, Lourdes Tajes Martínez, Fernando Álvarez García, María Ángeles Díaz Fondón, Raúl Izquierdo Castanedo, and Juan Manuel Cueva Lovelle. An object-oriented abstract machine as the substrate for an objectoriented operating system. Lecture Notes in Computer Science, 1357:537– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570537.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570537.
                                        pdf.

Graupner:1998:ODE

[107] Sven Graupner. Objects in different execution environments. Lecture Notes in Computer Science, 1357:545– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1357/13570545.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1357/13570545.
                                        pdf.

Albrecht:1998:IEA

[108] M. Albrecht, E. Buchholz, A. Duesterhoeft, and B. Thalheim. An informal and efficient approach for obtaining semantic constraints using sample data and natural language processing. Lecture Notes in Computer Science, 1358:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biskup:1998:ARD

[109] J. Biskup. Achievements of relational database schema design theory revisited. Lecture Notes in Computer Science, 1358:29–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davidson:1998:SDT

[110] S. Davidson, P. Buneman, and A. Kosky. Semantics of database transformations. Lecture Notes in Computer Science, 1358:55–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gottlob:1998:EAS

[111] G. Gottlob and M. Schrefl. The evolving algebra semantics of class and role hierarchies. Lecture Notes in Computer Science, 1358:92–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuijpers:1998:SSD

[112] B. Kuijpers, J. Paredaens, and L. Vandeurzen. Semantics in spatial databases. Lecture Notes in Computer Science, 1358:114–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levene:1998:APD

[113] M. Levene and G. Loizou. The additivity problem for data dependencies in incomplete relational databases. Lecture Notes in Computer Science, 1358: 136–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Libkin:1998:SBA

[114] L. Libkin. A semantics-based approach to design of query languages for partial information. Lecture Notes in Computer Science, 1358:170–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Revesz:1998:CDS

[115] P. Z. Revesz. Constraint databases: A survey. Lecture Notes in Computer Science, 1358:209–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vincent:1998:REN

[116] M. W. Vincent. Redundancy elimination and a new normal form for relational database design. Lecture Notes in Computer Science, 1358:247– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padmanabhuni:1998:ICL

[117] S. Padmanabhuni and A. K. Ghose. Inductive constraint logic programming: An overview. Lecture Notes in Computer Science, 1359:1–8, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1998:SAR

[118] G. Antoniou and M.-A. Williams. Some approaches to reasoning with incomplete and changing information. Lecture Notes in Computer Science, 1359: 9–44, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padmanabhuni:1998:CLG

[119] S. Padmanabhuni, R. Goebel, and K. Furukawa. Curried least general generalization: A framework for higher order concept learning. Lecture Notes in Computer Science, 1359:45– 60, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kyburg:1998:AV

[120] H. E. Kyburg. Approximate validity. Lecture Notes in Computer Science, 1359:61–77, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bulmer:1998:ITE

[121] M. Bulmer. Inductive theories from equational systems. Lecture Notes in Computer Science, 1359:78–91, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghose:1998:RDR

[122] A. K. Ghose, S. Padmanabhuni, and R. Goebel. The role of default representations in incremental learning. Lecture Notes in Computer Science, 1359: 92–105, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harries:1998:LSC

[123] M. Harries and K. Horn. Learning stable concepts in a changing world. Lecture Notes in Computer Science, 1359:106–122, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCane:1998:ICS

[124] B. McCane, T. Caelli, and O. De Vel. Inducing complex spatial descriptions in two dimensional scenes. Lecture Notes in Computer Science, 1359:123– 132, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padmanabhuni:1998:FLC

[125] S. Padmanabhuni, J.-H. You, and A. K. Ghose. A framework for learning constraints: Preliminary report. Lecture Notes in Computer Science, 1359:133– 147, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebag:1998:ICL

[126] M. Sebag, C. Rouveirol, and J.-F. Puget. Induction of constraint logic programs. Lecture Notes in Computer Science, 1359:148–167, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jitnah:1998:BNA

[127] N. Jitnah and A. E. Nicholson. Belief network algorithms: A study of performance based on domain characterization. Lecture Notes in Computer Science, 1359:168–187, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karacapilidis:1998:GDN

[128] N. Karacapilidis and D. Papadias. A group decision and negotiation support system for argumentation based reasoning. Lecture Notes in Computer Science, 1359:188–205, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacNish:1998:BRD

[129] C. K. MacNish and M.-A. Williams. From belief revision to design revision: Applying theory change to changing requirements. Lecture Notes in Computer Science, 1359:206–220, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinto:1998:UHM

[130] J. A. Pinto. Using histories to model observations in theories of action. Lecture Notes in Computer Science, 1359:221–233, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prokopenko:1998:MIA

[131] M. Prokopenko and P. Peppas. Modelling inertia in action languages (extended report). Lecture Notes in Computer Science, 1359:234–247, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramer:1998:CIU

[132] A. Ramer. Combinatorial interpretation of uncertainty and conditioning. Lecture Notes in Computer Science, 1359:248–255, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Restificar:1998:PDU

[133] A. C. Restificar. Probabilistic diagnosis as an update problem. Lecture Notes in Computer Science, 1359:256–266, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teng:1998:CCD

[134] Choh Man Teng. Cooperative combination of default logic and autoepistemic logic. Lecture Notes in Computer Science, 1359:267–283, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1998:AGR

[135] D. Kapur. Automated geometric reasoning: Dixon resultants, Gröbner bases, and characteristic sets. Lecture Notes in Computer Science, 1360:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tran:1998:EDR

[136] Q.-N Tran. Extended Dixon’s resultant and its applications. Lecture Notes in Computer Science, 1360:37–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sturm:1998:CGP

[137] T. Sturm and V. Weispfenning. Computational geometry problems in RED-LOG. Lecture Notes in Computer Science, 1360:58–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferro:1998:PVE

[138] G. C. Ferro, G. Gallo, and R. Gennaro. Probabilistic verification of elementary geometry statements. Lecture Notes in Computer Science, 1360:87–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havel:1998:CSG

[139] T. F. Havel. Computational synthetic geometry with Clifford algebra. Lecture Notes in Computer Science, 1360:102–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1998:CAC

[140] D. Wang. Clifford algebraic calculus for geometric reasoning with application to computer vision. Lecture Notes in Computer Science, 1360:115–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fearnley-Sander:1998:AGG

[141] D. Fearnley-Sander and T. Stokes. Area in Grassmann geometry. Lecture Notes in Computer Science, 1360:141– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1998:APR

[142] L. Yang, X.-S. Gao, S.-C. Chou, and J.-Z. Zhang. Automated production of readable proofs for theorems in non-Euclidean geometries. Lecture Notes in Computer Science, 1360:171–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hillgarter:1998:PAC

[143] E. Hillgarter and F. Winkler. Points on algebraic curves and the parametrization problem. Lecture Notes in Computer Science, 1360:189–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1998:FCC

[144] H. Shi and F. Zou. Flat central configurations of four planet motions. Lecture Notes in Computer Science, 1360: 208–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fevre:1998:IRA

[145] S. Fevre. Integration of reasoning and algebraic calculus in geometry. Lecture Notes in Computer Science, 1360:218– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1998:SBP

[146] R. J. Anderson, V. Matyas, F. A. Petitcolas, and I. E. Buchan. Secure books: Protecting the distribution of knowledge. Lecture Notes in Computer Science, 1361:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reed:1998:PUA

[147] M. G. Reed, P. F. Syverson, and D. M. Goldschlag. Protocols using anonymous connections: Mobile applications. Lecture Notes in Computer Science, 1361:13–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okamoto:1998:RFE

[148] T. Okamoto. Receipt-free electronic voting schemes for large scale elections. Lecture Notes in Computer Science, 1361:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solana:1998:FIS

[149] E. Solana and J. Harms. Flexible Internet secure transactions based on collaborative domains. Lecture Notes in Computer Science, 1361:37–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crispo:1998:HBE

[150] B. Crispo. How to build evidence in a public-key infrastructure for multidomain environments. Lecture Notes in Computer Science, 1361:53–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersen:1998:SST

[151] H. Petersen and M. Michels. On signature schemes with threshold verification detecting malicious verifiers. Lecture Notes in Computer Science, 1361: 67–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucks:1998:OKE

[152] S. Lucks. Open key exchange: How to defeat dictionary attacks without encrypting public keys. Lecture Notes in Computer Science, 1361:79–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelsey:1998:PIC

[153] J. Kelsey, B. Schneier, and D. Wagner. Protocol interactions and the chosen protocol attack. Lecture Notes in Computer Science, 1361:91–104, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://www.counterpane.com/
chosen_protocol.html.

Christianson:1998:BBP

[154] B. Christianson and J. A. Malcolm. Binding bit patterns to real world entities. Lecture Notes in Computer Science, 1361:105–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bao:1998:BPK

[155] F. Bao, R. H. Deng, Y. F. Han, and A. B. R. Jeng. Breaking public key cryptosystems on tamper resistant devices in the presence of transient faults. Lecture Notes in Computer Science, 1361:115–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1998:LCA

[156] R. J. Anderson and M. Kuhn. Low cost attacks on tamper resistant devices. Lecture Notes in Computer Science, 1361:125–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blake-Wilson:1998:EAA

[157] S. Blake-Wilson and A. Menezes. Entity authentication and authenticated key transport protocols employing asymmetric techniques. Lecture Notes in Computer Science, 1361:137– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guergens:1998:SLF

[158] S. Guergens. SG logic — A formal analysis technique for authentication protocols. Lecture Notes in Computer Science, 1361:159–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersen:1998:HCD

[159] H. Petersen. How to convert any digital signature scheme into a group signature scheme. Lecture Notes in Computer Science, 1361:177–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okamoto:1998:TKR

[160] T. Okamoto. Threshold key-recovery systems for RSA. Lecture Notes in Computer Science, 1361:191–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiefer:1998:WMV

[161] K. Kiefer. A weakness of the Menezes-Vanstone cryptosystem. Lecture Notes in Computer Science, 1361:201–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paillier:1998:INP

[162] P. Paillier. On ideal non-perfect secret sharing schemes. Lecture Notes in Computer Science, 1361:207–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markatos:1998:REO

[163] Evangelos P. Markatos, Manolis G. H. Katevenis, and Penny Vatsolaki. The remote enqueue operation on networks of workstations. Lecture Notes in Computer Science, 1362:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620001.
pdf.

Larson:1998:HIP

[164] Jeff Larson. The HAL interconnect PCI card. Lecture Notes in Computer Science, 1362:15–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620015.
pdf.

Nagar:1998:IPM

[165] Shailabh Nagar, Dale Seed, and Anand Sivasubramaniam. Implementing protected multi-user communication for myrinet. Lecture Notes in Computer Science, 1362:30–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620030.
pdf.

Brebner:1998:ECE

[166] Gordon Brebner and Rob Pooley. ECOLE: A configurable environment for a local optical network of workstations. Lecture Notes in Computer Science, 1362:45–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620045.
pdf.

Ming:1998:DPP

[167] Chan Wai Ming, Samuel Chanson, and Mounir Hamdi. The design of a parallel programming system for a network of workstations: An object-oriented approach. Lecture Notes in Computer Science, 1362:59– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620059.
pdf.

Bangalore:1998:RSA

[168] Manjunath Bangalore and Anand Sivasubramaniam. Remote subpaging across a fast network. Lecture Notes in Computer Science, 1362:74– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620074.
pdf.

Munz:1998:IFI

[169] F. Munz, T. Stephan, U. Maier, T. Ludwig, A. Bode, S. Ziegler, S. Nekolla, P. Bartenstein, and M. Schwaiger. Improved functional imaging through network based parallel processing. Lecture Notes in Computer Science, 1362:88–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620088.
pdf.

Goujon:1998:AAT

[170] Delphine Stéphanie Goujon, Martial Michel, Jasper Peeters, and Judith Ellen Devaney. AutoMap and AutoLink: Tools for communicating complex and dynamic datastructures using MPI. Lecture Notes in Computer Science, 1362:98–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620098.
pdf.

Matloff:1998:APB

[171] Norman Matloff. Analysis of a programmed backoff method for parallel processing on ethernets. Lecture Notes in Computer Science, 1362:110– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1362/13620110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1362/13620110.
                                        pdf.

Matloft:1998:APB

[172] N. Matloft. Analysis of a programmed backoff method for parallel processing on ethernets. Lecture Notes in Computer Science, 1362:110–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1998:IDT

[173] Alexander I-Chi Lai and Chin-Laung Lei. Improving dynamic token-based distributed synchronization performance via optimistic broadcasting. Lecture Notes in Computer Science, 1362:118–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1362/13620118.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1362/13620118.
                                        pdf.

Chiola:1998:FBS

[174] G. Chiola and G. Ciaccio. Fast barrier synchronization on shared fast ethernet. Lecture Notes in Computer Science, 1362:132–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1362/13620132.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1362/13620132.pdf.

Xiao:1998:PRT

[175] Xipeng Xiao and Lionel M. Ni. Parallel routing table computation for scalable IP routers. Lecture Notes in Computer Science, 1362:144–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620144.
pdf.

Casado:1998:TAR

[176] Rafael Casado, Blanca Caminero, Pedro Cuenca, Francisco Quiles, Antonio Garrido, and José Duato. A tool for the analysis of reconfiguration and routing algorithms in irregular networks. Lecture Notes in Computer Science, 1362: 159–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620159.
pdf.

Yu:1998:RTT

[177] Cheng-Chi Yu and Sourav Bhattacharya. Real-time traffic management on optical networks using dynamic control of cycle duration. Lecture Notes in Computer Science, 1362: 174–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1362/13620174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1362/13620174.
                                        pdf.

Karlsson:1998:CCC

[178] Sven Karlsson and Mats Brorsson. A comparative characterization of communication patterns in applications using MPI and shared memory on an IBM SP2. Lecture Notes in Computer Science, 1362:189–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1362/13620189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1362/13620189.
                                        pdf.

Kim:1998:CCP

[179] JunSeong Kim and David J. Lilja. Characterization of communication patterns in message-passing parallel scientific application programs. Lecture Notes in Computer Science, 1362: 202–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1362/13620202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1362/13620202.
                                        pdf.

Iannello:1998:CPA

[180] Giulio Iannello, Mario Lauria, and Stefano Mercolino. Cross-platform analysis of fast messages for myrinet. Lecture Notes in Computer Science, 1362: 217–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620217.
pdf.

Omang:1998:PCP

[181] Knut Omang. Performance of a cluster of PCI based UltraSparc workstations interconnected with SCI. Lecture Notes in Computer Science, 1362:232– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1362/13620232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1362/13620232.
pdf.

Glover:1998:TSS

[182] Fred Glover. A template for scatter search and path relinking. Lecture Notes in Computer Science, 1363:13– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1363/13630003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630003.
pdf.

Gottlieb:1998:RFF

[183] Jens Gottlieb and Nico Voss. Representations, fitness functions and genetic operators for the satisfiability problem. Lecture Notes in Computer Science, 1363:55–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1363/13630055.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1363/13630055.
                                        pdf.

Escazut:1998:GAE

[184] Cathy Escazut and Philippe Collard. Genetic algorithms at the edge of a dream. Lecture Notes in Computer Science, 1363:69–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1363/13630069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1363/13630069.
                                        pdf.

Peyral:1998:ME

[185] Mathieu Peyral, Antoine Ducoulombier, Caroline Ravisé, Marc Schoenauer, and Michèle Sebag. Mimetic evolution. Lecture Notes in Computer Science, 1363:81–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1363/13630081.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1363/13630081.pdf.

Eiben:1998:APE

[186] A. E. Eiben and J. K. Van Der Hauw. Adaptive penalties for evolutionary graph coloring. Lecture Notes in Computer Science, 1363:95– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1363/13630095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630095.
pdf.

Cuenca:1998:ASL

[187] Cristina Cuenca and Jean-Claude Heudin. An agent system for learning profiles in broadcasting applications on the Internet. Lecture Notes in Computer Science, 1363:109–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1363/13630109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630109.
pdf.

Piccolboni:1998:AEA

[188] A. Piccolboni and G. Mauri. Application of evolutionary algorithms to protein folding prediction. Lecture Notes in Computer Science, 1363:123– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1363/13630123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630123.
pdf.

Servet:1998:TNT

[189] Isabelle Servet, Louise Travé-Massuyès, and Daniel Stern. Telephone network traffic overloading diagnosis and evolutionary computation techniques. Lecture Notes in Computer Science, 1363: 137–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1363/13630137.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1363/13630137.
                                        pdf.

Gaspin:1998:GAG

[190] Christine Gaspin and Thomas Schiex. Genetic algorithms for genetic mapping. Lecture Notes in Computer Science, 1363:145–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1363/13630145.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1363/13630145.pdf.

Leblanc:1998:IPF

[191] Benoit Leblanc, Evelyne Lutton, and Jean-Paul Allouche. Inverse problems for finite automata: A solution based on genetic algorithms. Lecture Notes in Computer Science, 1363:157– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1363/13630157.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1363/13630157.
                                        pdf.

Tanomaru:1998:ETM

[192] Julio Tanomaru. Evolving Turing machines from examples. Lecture Notes in Computer Science, 1363:167– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1363/13630167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630167.
pdf.

Agapie:1998:GAM

[193] Alexandru Agapie. Genetic algorithms: Minimal conditions for convergence. Lecture Notes in Computer Science, 1363:183–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1363/13630183.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1363/13630183.
pdf.

Oh:1998:APE

[194] Sangyeop Oh and Hyunsoo Yoon. An analysis of punctuated equilibria in simple genetic algorithms. Lecture Notes in Computer Science, 1363:195– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1363/13630195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630195.
pdf.

Naudts:1998:SSD

[195] Bart Naudts and Alain Verschoren. SGA search dynamics on second order functions. Lecture Notes in Computer Science, 1363:207–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1363/13630207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1363/13630207.
                                        pdf.

Rudolph:1998:ACR

[196] Günter Rudolph. Asymptotical convergence rates of simple evolutionary algorithms under factorizing mutation distributions. Lecture Notes in Computer Science, 1363:223–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1363/13630223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1363/13630223.
                                        pdf.

Dedieu:1998:WWD

[197] Eric Dedieu, Olivier Lebeltel, and Pierre Bessière. Wings were not designed to let animals fly. Lecture Notes in Computer Science, 1363:237– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1363/13630237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1363/13630237.
                                        pdf.

Salomon:1998:AET

[198] Ralf Salomon and Peter Eggenberger. Adaptation on the evolutionary time scale: A working hypothesis and basic experiments. Lecture Notes in Computer Science, 1363:251–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1363/13630251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630251.
pdf.

Crisan:1998:FAP

[199] Christine Crisan and Heinz Mühlenbein. The frequency assignment problem: A look at the performance of evolutionary search. Lecture Notes in Computer Science, 1363:263–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1363/13630263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630263.
pdf.

Rochet:1998:CES

[200] S. Rochet, G. Venturini, M. Slimane, and E. M. El Kharoubi. A critical and empirical study of epistasis measures for predicting GA performances: A summary. Lecture Notes in Computer Science, 1363:275– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1363/13630275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630275.
pdf.

Kallel:1998:PCB

[201] Leila Kallel and Marc Schoenauer. A priori comparison of binary crossover operators: No universal statistical measure, but a set of hints. Lecture Notes in Computer Science, 1363:287– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1363/13630287.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1363/13630287.
                                        pdf.

Loeffler:1998:DNP

[202] A. Loeffler, J. Klahold, and U. Rueckert. The dynamical nightwatch’s problem solved by the autonomous microrobot khepera. Lecture Notes in Computer Science, 1363:303–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loffler:1998:DNP

[203] A. Löffler, J. Klahold, and U. Rückert. The Dynamical Nightwatch’s problem solved by the autonomous microrobot Khepera. Lecture Notes in Computer Science, 1363:303–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1363/13630303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1363/13630303.
                                        pdf.

Gers:1998:CSC

[204] Felix Gers, Hugo de Garis, and Michael Korkin. CoDi-1Bit: A simplified cellular automata based neuron model. Lecture Notes in Computer Science, 1363: 315–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1363/13630315.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630315.
pdf.

DeGaris:1998:MMN

[205] H. De Garis, L. Kang, Q. He, and Z. Pan. Million module neural systems evolution — the next step in ATR’s billion neuron artificial brain (“CAM-Brain”) project. Lecture Notes in Computer Science, 1363:335–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garis:1998:MMN

[206] Hugo de Garis, Lishan Kang, Qiming He, Zhengjun Pan, Masahiro Ootani, and Edmund Ronald. Million module neural systems evolution — the next step in ATR’s billion neuron artificial brain (“CAM-Brain”) project. Lecture Notes in Computer Science, 1363:335– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1363/13630335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1363/13630335.
pdf.

Conen:1998:PTA

[207] W. Conen and G. Neumann. A perspective on technology-assisted collaboration. Lecture Notes in Computer Science, 1364:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiefloe:1998:CKI

[208] P. M. Schiefloe and T. G. Syvertsen. Coordination in knowledgeintensive organizations. Lecture Notes in Computer Science, 1364:9–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernus:1998:COM

[209] P. Bernus and G. Uppington. Coordination of management activities — mapping organisational structure to the decision structure. Lecture Notes in Computer Science, 1364:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1998:CAD

[210] G. Fernandez and I. Wijegunaratne. A cooperative approach to distributed applications engineering. Lecture Notes in Computer Science, 1364:39–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarau:1998:TLP

[211] P. Tarau, V. Dahl, and K. De Bosschere. Towards logic programming based coordination in virtual worlds. Lecture Notes in Computer Science, 1364:49–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishimoto:1998:ECA

[212] K. Nishimoto, Y. Sumi, and K. Mase. Enhancement of creative aspects of a daily conversation with a topic development agent. Lecture Notes in Computer Science, 1364:63–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Decker:1998:CHC

[213] K. S. Decker. Coordinating human and computer agents. Lecture Notes in Computer Science, 1364:77–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappel:1998:CWM

[214] G. Kappel, S. Rausch-Schott, and W. Retschitzegger. Coordination in workflow management systems — A rule-based approach. Lecture Notes in Computer Science, 1364:99–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellis:1998:FMM

[215] C. A. Ellis. A framework and mathematical model for collaboration technology. Lecture Notes in Computer Science, 1364:121–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1998:PER

[216] K.-H. Kim and S.-K Pak. Practical experiences and requirements on workflow. Lecture Notes in Computer Science, 1364:145–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:1998:CSC

[217] M. Klein. Coordination science: Challenges and directions. Lecture Notes in Computer Science, 1364:161–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1998:SAW

[218] M. Berger, G. Voelksen, and A. Schill. Supporting autonomous work and reintegration in collaborative systems. Lecture Notes in Computer Science, 1364: 177–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlichter:1998:WAD

[219] J. Schlichter, M. Koch, and M. Buerger. Workspace awareness for distributed teams. Lecture Notes in Computer Science, 1364:199–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sivadas:1998:GWT

[220] M. Sivadas and G. Fernandez. GeM and WeBUSE: Towards a WWW-Database interface. Lecture Notes in Computer Science, 1364:219–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuehn:1998:PCS

[221] E. Kuehn and G. Nozicka. Postclient/server coordination tools. Lecture Notes in Computer Science, 1364: 231–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1998:EDA

[222] D.-H. Kim, K.-H. Lee, M.-G. Kang, and G.-H. Han. An experimental delay analysis for local audio video streams for desktop collaborations. Lecture Notes in Computer Science, 1364:255– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1998:SBC

[223] T. Lee, J. Park, Seok Hwan Yoon, and Pyeong Jung Kim. Supporting both client-server and peer-to-peer models in a framework of a distributed object management system. Lecture Notes in Computer Science, 1364:263–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kinny:1998:PMM

[224] D. Kinny, J. Treur, L. Gasser, and S. Clark. Panel: Methodologies for multi-agent systems. Lecture Notes in Computer Science, 1365:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoppers:1998:DEA

[225] M. Schoppers and D. Shapiro. Designing embedded agents to optimize end-user objectives. Lecture Notes in Computer Science, 1365:3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryson:1998:AAO

[226] J. Bryson and B. McGonigle. Agent architecture as object oriented design. Lecture Notes in Computer Science, 1365:15–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verharen:1998:ICA

[227] E. M. Verharen, F. Dignum, and S. Bos. Implementation of a cooperative agent architecture based on the language-action perspective. Lecture Notes in Computer Science, 1365:31– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parunak:1998:TSD

[228] V. Parunak, J. Sauter, and S. Clark. Toward the specification and design of industrial synthetic ecosystems. Lecture Notes in Computer Science, 1365: 45–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwartz:1998:BMD

[229] R. Schwartz and S. Kraus. Bidding mechanisms for data allocation in multi-agent environments. Lecture Notes in Computer Science, 1365:61– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singh:1998:CCS

[230] M. P. Singh. A customizable coordination service for autonomous agents. Lecture Notes in Computer Science, 1365:93–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coradeschi:1998:BBA

[231] S. Coradeschi and L. Karlsson. A behavior-based approach to reactivity and coordination: A preliminary report. Lecture Notes in Computer Science, 1365:107–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1998:EPL

[232] J. Lee and E. H. Durfee. On explicit plan languages for coordinating multiagent plan execution. Lecture Notes in Computer Science, 1365:113–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaminka:1998:SCF

[233] G. A. Kaminka and M. Tambe. Social comparison for failure detection and recovery. Lecture Notes in Computer Science, 1365:127–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shehory:1998:MAC

[234] O. M. Shehory, K. Sycara, and S. Jha. Multi-agent coordination through coalition formation. Lecture Notes in Computer Science, 1365:143– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DInverno:1998:FSD

[235] M. D’Inverno, D. Kinny, M. Luck, and M. Wooldridge. A formal specification of dMARS. Lecture Notes in Computer Science, 1365:155–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sierra:1998:FAB

[236] C. Sierra, N. R. Jennings, P. Noriega, and S. Parsons. A framework for argumentation-based negotiation. Lecture Notes in Computer Science, 1365:177–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mulder:1998:AMM

[237] M. Mulder, J. Treur, and M. Fisher. Agent modelling in MetateM and DE-SIRE. Lecture Notes in Computer Science, 1365:193–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Labrou:1998:SAC

[238] Y. Labrou and T. Finin. Semantics for an agent communication language. Lecture Notes in Computer Science, 1365:209–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hindriks:1998:FSA

[239] K. V. Hindriks, F. S. De Boer, W. Van der Hoek, and J.-J. C. Meyer. Formal semantics for an abstract agent programming language. Lecture Notes in Computer Science, 1365:215–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dignum:1998:IAG

[240] F. Dignum and R. Conte. Intentional agents and goal formation. Lecture Notes in Computer Science, 1365:231– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jung:1998:LAC

[241] C. G. Jung and K. Fischer. A layered agent calculus with concurrent, continuous processes. Lecture Notes in Computer Science, 1365:245–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koriche:1998:ARA

[242] F. Koriche. Approximate reasoning about combined knowledge. Lecture Notes in Computer Science, 1365:259– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duc:1998:EFA

[243] Ho Ngoc Duc. On the epistemic foundations of agent theories. Lecture Notes in Computer Science, 1365:275– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nodine:1998:FOC

[244] M. H. Nodine and A. Unruh. Facilitating open communication in agent systems: The InfoSleuth infrastructure. Lecture Notes in Computer Science, 1365:281–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandeVelde:1998:CA

[245] W. Van de Velde, S. Geldof, and R. Schrooten. Competition for attention. Lecture Notes in Computer Science, 1365:297–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iglesias:1998:ADM

[246] C. A. Iglesias, M. Garijo, J. C. Gonzalez, and J. R. Velasco. Analysis and design of multiagent systems using MAS-CommonKADS. Lecture Notes in Computer Science, 1365:313–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustavsson:1998:MAS

[247] R. E. Gustavsson. Multi-agent systems as open societies — A design framework. Lecture Notes in Computer Science, 1365:329–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cost:1998:TST

[248] R. S. Cost, I. Soboroff, J. Lakhani, and T. Finin. TKQML: A scripting tool for building agents. Lecture Notes in Computer Science, 1365:339–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1998:QML

[249] N. Mitchell, L. Carter, J. Ferrante, and K. Hoegstedt. Quantifying the multi-level nature of tiling interactions. Lecture Notes in Computer Science, 1366:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xue:1998:RDT

[250] J. Xue and C.-H. Huang. Reuse-driven tiling for data locality. Lecture Notes in Computer Science, 1366:16–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shih:1998:TLA

[251] K.-P. Shih, J.-P. Sheu, and C.- H. Huang. Table-lookup approach for compiling two-level data-processor mappings in HPF. Lecture Notes in Computer Science, 1366:34–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramanujam:1998:CGC

[252] J. Ramanujam, S. Dutta, and A. Venkatachar. Code generation for complex subscripts in data-parallel programs. Lecture Notes in Computer Science, 1366:49–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damian-Iordache:1998:ADD

[253] M. Damian-Iordache and S. V. Pemmaraju. Automatic data decomposition for message-passing machines. Lecture Notes in Computer Science, 1366: 64–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sawdey:1998:PAO

[254] A. Sawdey and M. O’Keefe. Program analysis of overlap area usage in self-similar parallel programs. Lecture Notes in Computer Science, 1366:79– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarkar:1998:AOE

[255] V. Sarkar. Analysis and optimization of explicitly parallel programs using the parallel program graph representation. Lecture Notes in Computer Science, 1366:94–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1998:CSS

[256] J. Lee, S. P. Midkiff, and D. A. Padua. Concurrent static single assignment form and constant propagation for explicitly parallel programs. Lecture Notes in Computer Science, 1366: 114–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1998:IDU

[257] Y.-S. Hwang and J. Saltz. Identifying DEF/USE information of statements that construct and traverse dynamic recursive data structures. Lecture Notes in Computer Science, 1366:

131–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsai:1998:POC

[258] J.-Y. Tsai, Z. Jiang, and P.-C. Yew. Program optimization for concurrent multithreaded architectures. Lecture Notes in Computer Science, 1366:146– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1998:ICP

[259] I. Park, M. Voss, B. Armstrong, and R. Eigenmann. Interactive compilation and performance analysis with ursa minor. Lecture Notes in Computer Science, 1366:163–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1998:STN

[260] T.-C. Huang and P.-H. Hsu. The SPNT test: A new technology for run-time speculative parallelization of loops. Lecture Notes in Computer Science, 1366:177–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borowiec:1998:LHP

[261] J. Borowiec and A. Veen. Lowering HPF procedure interface to a canonical representation. Lecture Notes in Computer Science, 1366:192–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1998:PBH

[262] G. Zhang, B. Carpenter, G. Fox, and X. Li. PCRC-based HPF compilation. Lecture Notes in Computer Science, 1366:204–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trabado:1998:DPL

[263] G. P. Trabado and E. L. Zapata. Data parallel language extensions for exploiting locality in irregular problems. Lecture Notes in Computer Science, 1366:218–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mellor-Crummey:1998:SCF

[264] J. Mellor-Crummey and V. Adve. Simplifying control flow in compilergenerated parallel code. Lecture Notes in Computer Science, 1366:235–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1998:RSO

[265] H. Han, C.-W. Tseng, and P. Keleher. Reducing synchronization overhead for compiler-parallelized codes on software DSMs. Lecture Notes in Computer Science, 1366:240–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuan:1998:ADF

[266] X. Yuan, R. Gupta, and R. Melhem. An array data flow analysis based communication optimizer. Lecture Notes in Computer Science, 1366:246–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chamberlain:1998:CAM

[267] B. L. Chamberlain, S.-E. Choi, and L. Snyder. A compiler abstraction for machine independent parallel communication generation. Lecture Notes in Computer Science, 1366:261–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1998:AFA

[268] H. G. Dietz, T. I. Mattoz, and G. Krishnamurthy. The aggregate function API: It’s not just for PAPERS anymore. Lecture Notes in Computer Science, 1366:277–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Numrich:1998:DEF

[269] R. W. Numrich, J. L. Steidel, B. H. Johnson, and B. Dupont de Dinechin. Definition of the F_{−−} extension to Fortran 90. Lecture Notes in Computer Science, 1366:292–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayguade:1998:EPT

[270] E. Ayguade, X. Martorell, J. Labarta, and M. Gonzalez. Exploiting parallelism through directives on the nanothreads programming model. Lecture Notes in Computer Science, 1366:307– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banerjee:1998:OPT

[271] D. Banerjee and Y. C. Browne. “optimal” parallelism through integration of data and control parallelism: A case study in complete parallelization. Lecture Notes in Computer Science, 1366:322–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpenter:1998:JLS

[272] B. Carpenter, Y.-J. Chang, G. Fox, and X. Li. Java as a language for scientific parallel programming. Lecture Notes in Computer Science, 1366:340–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bik:1998:ELP

[273] A. J. C. Bik, J. E. Villacis, and D. B. Gannon. Experience with loop parallelization in javar (A prototype restructuring compiler for Java). Lecture Notes in Computer Science, 1366:355– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kale:1998:NCS

[274] L. V. Kale, M. Bhandarkar, R. Brunner, and N. Krawetz. NAMD: A case study in multilingual parallel programming. Lecture Notes in Computer Science, 1366:367–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DupontdeDinechin:1998:USP

[275] B. Dupont de Dinechin. A unified software pipeline construction scheme for modulo scheduled loops. Lecture Notes in Computer Science, 1366:382– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bilardi:1998:SAB

[276] G. Bilardi, A. A. Nicolau, and J. Hummel. A systematic approach to branch speculation. Lecture Notes in Computer Science, 1366:394–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1998:IAD

[277] G.-H. Hwang, J. K. Lee, and D.- C. R. Ju. Integrating automatic data alignment and array operation synthesis to optimize data parallel programs. Lecture Notes in Computer Science, 1366:412–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1998:CIS

[278] M. Gupta and S. P. Midkiff. A compiler for the IBM scalable shared memory project machine. Lecture Notes in Computer Science, 1366:416–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kremer:1998:ADL

[279] U. Kremer. Automatic data layout with read-only replication and memory constraints. Lecture Notes in Computer Science, 1366:419–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1998:SAR

[280] D. K. Arvind and T. A. Lewis. Static analysis of recursive data structures. Lecture Notes in Computer Science, 1366:423–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1998:ITC

[281] Thomas Jansen. 1. introduction to the theory of complexity and approximation algorithms. Lecture Notes in Computer Science, 1367:5– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1367/13670005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1367/13670005.
pdf.

Andrzejak:1998:IRA

[282] Artur Andrzejak. 2. introduction to randomized algorithms. Lecture Notes in Computer Science, 1367:29– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1367/13670029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1367/13670029.
pdf.

Sieling:1998:D

[283] Detlef Sieling. 3. derandomization. Lecture Notes in Computer Science, 1367: 41–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1367/13670041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1367/13670041.
pdf.

Hougardy:1998:PCN

[284] Stefan Hougardy. 4. proof checking and non-approximability. Lecture Notes in Computer Science, 1367:63– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1367/13670063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1367/13670063.
pdf.

Heun:1998:PPT

[285] Volker Heun, Wolfgang Merkle, and Ulrich Weigand. 5. proving the PCP-theorem. Lecture Notes in Computer Science, 1367:83–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1367/13670083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1367/13670083.
                                        pdf.

Groepl:1998:PRM

[286] C. Groepl and M. Skutella. Parallel repetition of MIP(2,1) systems. Lecture Notes in Computer Science, 1367: 161–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gropl:1998:PRM

[287] Clemens Gröpl and Martin Skutella. 6. parallel repetition of MIP(2,1) systems. Lecture Notes in Computer Science, 1367:161–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1367/13670161.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1367/13670161.
                                        pdf.

Seibert:1998:BAM

[288] Sebastian Seibert and Thomas Wilke. 7. bounds for approximating MaxLinEq3-2 and MaxEkSat. Lecture Notes in Computer Science, 1367:179–211, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1367/13670179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1367/13670179.
pdf.

Rick:1998:DNA

[289] Claus Rick and Hein Röhrig. 8. deriving non-approximability results by reductions. Lecture Notes in Computer Science, 1367:213–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1367/13670213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1367/13670213.
pdf.

Mundhenk:1998:ONA

[290] Martin Mundhenk and Anna Slobodová. 9. optimal non-approximability of MaxClique. Lecture Notes in Computer Science, 1367:235–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1367/13670235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1367/13670235.
pdf.

Wolff:1998:HAS

[291] Alexander Wolff. 10. the hardness of approximating set cover. Lecture Notes in Computer Science, 1367:249– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1367/13670249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1367/13670249.
                                        pdf.

Hofmeister:1998:SPA

[292] Thomas Hofmeister and Martin Hühne. 11. semidefinite programming and its applications to approximation algorithms. Lecture Notes in Computer Science, 1367:263–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1367/13670263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1367/13670263.
                                        pdf.

Wolf:1998:DIH

[293] Katja Wolf. 12. dense instances of hard optimization problems. Lecture Notes in Computer Science, 1367:299– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1367/13670299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1367/13670299.
                                        pdf.

Mayr:1998:PTA

[294] Richard Mayr and Annette Schelten. 13. polynomial time approximation schemes for geometric optimization problems in Euclidean metric spaces. Lecture Notes in Computer Science, 1367:313–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1367/13670313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1367/13670313.
pdf.

Anonymous:1998:B

[295] Anonymous. Bibliography. Lecture Notes in Computer Science, 1367:325– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1367/13670325.pdf.

Anonymous:1998:AIa

[296] Anonymous. Author index. Lecture Notes in Computer Science, 1367:335– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1367/13670335.pdf.

Anonymous:1998:SI

[297] Anonymous. Subject index. Lecture Notes in Computer Science, 1367:337– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1367/13670337.pdf.

Anonymous:1998:LC

[298] Anonymous. List of contributors. Lecture Notes in Computer Science, 1367: 343–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1367/13670343.pdf.

Tamura:1998:WHI

[299] K. Tamura. WWC and the hyper information age. Lecture Notes in Computer Science, 1368:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chon:1998:GHP

[300] K. Chon. Global high performance research network: An Asia-Pacific perspective. Lecture Notes in Computer Science, 1368:6–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steele:1998:JAI

[301] G. L. Steele. Java applications and implementations. Lecture Notes in Computer Science, 1368:18–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tokoro:1998:BHW

[302] M. Tokoro. Back to home: Where computers and networking should go. Lecture Notes in Computer Science, 1368: 32–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glass:1998:OVA

[303] G. Glass. ObjectSpace voyager — the agent ORB for Java. Lecture Notes in Computer Science, 1368:38–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsubara:1998:WCS

[304] K. Matsubara, T. Maekawa, and K. Kato. Worldwide component scripting with the planet mobile object system. Lecture Notes in Computer Science, 1368:56–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1998:SOO

[305] A. Nakamura, T. Nishioka, Y. Hamazaki, and M. Tsukamoto. Scalability in object-oriented distributed systems environment OZ. Lecture Notes in Computer Science, 1368:72–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luckham:1998:RLT

[306] D. C. Luckham. Rapide: A language and toolset for causal event modeling of distributed system architectures. Lecture Notes in Computer Science, 1368:88–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aoyama:1998:ASC

[307] M. Aoyama, T. Yamashita, and S. Kobori. An architecture of software commerce broker over the Internet. Lecture Notes in Computer Science, 1368: 97–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsushita:1998:DPM

[308] M. Matsushita, M. Oshita, H. Iida, and K. Inoue. Distributed process management system based on object-centered process modeling. Lecture Notes in Computer Science, 1368:108–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajkumar:1998:SSM

[309] R. Rajkumar. Systems software for multimedia computing. Lecture Notes in Computer Science, 1368:120–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakajima:1998:ESS

[310] T. Nakajima, H. Aizu, M. Kobayashi, and K. Shimamoto. Environment server: A system support for adaptive distributed applications. Lecture Notes in Computer Science, 1368:142– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furukawa:1998:CGP

[311] Y. Furukawa and E. Shibayama. Compiler-generated protection domains and a light weight runtime protection technique. Lecture Notes in Computer Science, 1368:158–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1998:CC

[312] V. Gupta, G. Montenegro, and J. Rulifson. Complete computing. Lecture Notes in Computer Science, 1368:174– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsuura:1998:CFR

[313] K. Matsuura, Y. Zheng, and H. Imai. Compact and flexible resolution of CBT multicast key-distribution. Lecture Notes in Computer Science, 1368: 190–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shionozaki:1998:IRR

[314] A. Shionozaki, K. Yamashita, S. Utsumi, and K. Cho. Integrating resource reservation with rate-based transport protocols in AMInet. Lecture Notes in Computer Science, 1368:206–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1998:EMT

[315] K. Lai, M. Roussopoulos, D. Tang, and X. Zhao. Experiences with a mobile testbed. Lecture Notes in Computer Science, 1368:222–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishiyama:1998:DIM

[316] M. Ishiyama, A. Inoue, A. Fukumoto, and T. Okamoto. Design and implementation of mobile IP system with security consideration. Lecture Notes in Computer Science, 1368:238–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uehara:1998:NAC

[317] K. Uehara, T. Tateoka, Y. Watanabe, and H. Sunahara. A network architecture for continuous mobility. Lecture Notes in Computer Science, 1368:254– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leong:1998:TIW

[318] Kok Yong Leong, H. Liu, and O. Wu. Towards internationalized Web creation. Lecture Notes in Computer Science, 1368:270–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMaria:1998:IIS

[319] P. De Maria and C. De Mei. ISCM, information system conceptual model oriented to security problems and a tool implementing it. Lecture Notes in Computer Science, 1368:282–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1998:DES

[320] T.-K Park. Design of EDI security MIB based on SNMP protocol. Lecture Notes in Computer Science, 1368: 293–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Masunaga:1998:BWD

[321] Y. Masunaga. The block-world data model for a collaborative virtual environment. Lecture Notes in Computer Science, 1368:309–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pulli:1998:CMM

[322] P. Pulli, T. Pyssysalo, K. Kuutti, and J. Similae. CyPhone — mobile multimodal personal augmented reality. Lecture Notes in Computer Science, 1368:325–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1998:BIS

[323] W. Kim. On business intelligence systems. Lecture Notes in Computer Science, 1368:337–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tang:1998:SCT

[324] J. C. Tang. Supporting collaboration through teleproximity. Lecture Notes in Computer Science, 1368:349–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honda:1998:HOS

[325] S. Honda, K.-I. Okada, and Y. Matsushita. A home office system based on a virtual shared room: An environment corresponding to degree of concentration. Lecture Notes in Computer Science, 1368:364–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hazeyama:1998:EBS

[326] A. Hazeyama and M. Hanawa. Electronic binder system: Promotion of an ISO9001-Based quality system using the WWW and experience from its application. Lecture Notes in Computer Science, 1368:381–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pu:1998:UMC

[327] C. Pu and L. Liu. Update monitoring: The CQ project. Lecture Notes in Computer Science, 1368:396–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Looser:1998:DMF

[328] H. Looser and T. Haerder. dLIMIT — A middleware framework for looselycoupled database federations. Lecture Notes in Computer Science, 1368:412– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guo:1998:ABC

[329] Y. Guo, Y. Hiranaka, and T. Akatsuka. Autonomic buffer control of Web proxy server. Lecture Notes in Computer Science, 1368:428–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsumoto:1998:GUA

[330] T. Matsumoto. Getting users’ attributes without violating anonymity. Lecture Notes in Computer Science, 1368:439–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarmenta:1998:BWB

[331] L. F. G. Sarmenta. Bayanihan: Web-based volunteer computing using Java. Lecture Notes in Computer Science, 1368:444–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morioka:1998:AUI

[332] T. Morioka. Architecture of a user interface module for structured Internet messages. Lecture Notes in Computer Science, 1368:462–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paredaens:1998:ETE

[333] J. Paredaens, B. Kuijpers, G. Kuper, and L. Vandeurzen. Euclid, tarski, and engeler encompassed. Lecture Notes in Computer Science, 1369:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wadler:1998:FPA

[334] P. Wadler. Functional programming: An angry half-dozen. Lecture Notes in Computer Science, 1369:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cluet:1998:MDI

[335] S. Cluet, R. Connor, R. Hull, and D. Maier. Metadata for database interoperation. Lecture Notes in Computer Science, 1369:35–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grumbach:1998:DSC

[336] S. Grumbach, P. Rigaux, M. Scholl, and L. Segoufin. DEDALE, a spatial constraint database. Lecture Notes in Computer Science, 1369:38–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuijpers:1998:DMS

[337] B. Kuijpers. Degrees of monotonicity of spatial transformations. Lecture Notes in Computer Science, 1369:60– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alagic:1998:CMT

[338] S. Alagic. Constrained matching is type safe. Lecture Notes in Computer Science, 1369:78–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Connor:1998:EQP

[339] R. Connor. Existentially quantified procedures: A mechanism for abstracting type in dynamic typing constructs. Lecture Notes in Computer Science, 1369:97–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonner:1998:QSD

[340] A. J. Bonner and G. Mecca. Querying sequence databases with transducers. Lecture Notes in Computer Science, 1369:118–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1998:SBA

[341] M. Fernandez, L. Popa, and D. Suciu. A structure-based approach to querying semi-structured data. Lecture Notes in Computer Science, 1369:136– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdessalem:1998:VQL

[342] T. Abdessalem and G. Jomier. VQL: A query language for multiversion databases. Lecture Notes in Computer Science, 1369:160–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lacroix:1998:OVD

[343] Z. Lacroix, C. Delobel, and P. Breche. Object views and database restructuring. Lecture Notes in Computer Science, 1369:180–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawaguchi:1998:IIV

[344] A. Kawaguchi, D. Lieuwen, I. Mumick, and K. Ross. Implementing incremental view maintenance in nested data models. Lecture Notes in Computer Science, 1369:202–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Libkin:1998:IRR

[345] L. Libkin and L. Wong. Incremental recomputation of recursive queries with nested sets and aggregate functions. Lecture Notes in Computer Science, 1369:222–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beeri:1998:TLF

[346] C. Beeri, T. Milo, and P. Ta-Shma. Towards a language for the fully generic queries. Lecture Notes in Computer Science, 1369:239–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Libkin:1998:PAR

[347] L. Libkin and L. Wong. On the power of aggregation in relational query languages. Lecture Notes in Computer Science, 1369:260–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cadoli:1998:DDL

[348] M. Cadoli, L. Palopoli, and M. Lenzerini. Datalog and description logics: Expressive power. Lecture Notes in Computer Science, 1369:281–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faase:1998:IOT

[349] F. J. Faase, S. J. Even, R. A. De By, and P. M. G. Apers. Integrating organisational and transactional aspects of cooperative activities. Lecture Notes in Computer Science, 1369:336–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matthes:1998:BCH

[350] F. Matthes. Business conversations: A high-level system model for agent coordination. Lecture Notes in Computer Science, 1369:355–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonner:1998:TDC

[351] A. J. Bonner. Transaction datalog: A compositional language for transaction programming. Lecture Notes in Computer Science, 1369:373–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spelt:1998:AVT

[352] D. Spelt and H. Balsters. Automatic verification of transactions on an object-oriented database. Lecture Notes in Computer Science, 1369:396– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amaro:1998:SAT

[353] G. Amaro, F. Giannotti, and G. Mainetto. Static analysis of transactions for conservative multigranularity locking. Lecture Notes in Computer Science, 1369: 413–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weiser:1998:III

[354] Mark Weiser. The invisible interface: Increasing the power of the environment through calm technology. Lecture Notes in Computer Science, 1370: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1370/13700001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700001.
pdf.

Worthington:1998:WPK

[355] John Worthington. Working place for the knowledge economy. Lecture Notes in Computer Science, 1370:2– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700002.
pdf.

Streitz:1998:RCB

[356] Norbert A. Streitz, Jörg Geißler, and Torsten Holmer. Roomware for cooperative buildings: Integrated design of architectural spaces and information spaces. Lecture Notes in Computer Science, 1370:4–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700004.
pdf.

Wisneski:1998:ADT

[357] Craig Wisneski, Hiroshi Ishii, Andrew Dahley, Matt Gorbet, Scott Brave, Brygg Ullmer, and Paul Yarin. Ambient displays: Turning architectural space into an interface between people and digital information. Lecture Notes in Computer Science, 1370:22– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1370/13700022.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700022.
                                        pdf.

Rekimoto:1998:MCU

[358] Jun Rekimoto. Multiple-computer user interfaces: A cooperative environment consisting of multiple digital devices. Lecture Notes in Computer Science, 1370:33–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1370/13700033.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700033.
                                        pdf.

Coen:1998:PIE

[359] Michael H. Coen. A prototype intelligent environment. Lecture Notes in Computer Science, 1370:41–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1370/13700041.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700041.
                                        pdf.

Covi:1998:RYO

[360] Lisa M. Covi, Judith S. Olson, Elena Rocco, William J. Miller, and Paul Allie. A room of your own: What do we learn about support of teamwork from assessing teams in dedicated project rooms? Lecture Notes in Computer Science, 1370:53– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700053.
pdf.

Fitzpatrick:1998:EBC

[361] Geraldine Fitzpatrick, Simon Kaplan, and Sara Parsowith. Experience in building a cooperative distributed organization: Lessons for cooperative buildings. Lecture Notes in Computer Science, 1370:66–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700066.
pdf.

Yee:1998:KKC

[362] Susan Yee, William J. Mitchell, Ryusuke Naka, Mitsuo Morozumi, and Sigeyuki Yamaguchi. The Kumamoto-Kyoto-MIT collaborative project: A case study of the design studio of the future. Lecture Notes in Computer Science, 1370:80– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700080.
pdf.

Kirsh:1998:ARV

[363] David Kirsh. Adaptive rooms, virtual collaboration and cognitive workflow. Lecture Notes in Computer Science, 1370:94–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1370/13700094.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700094.
                                        pdf.

Pfister:1998:MVR

[364] Hans-Rüdiger Pfister, Christian Schuckmann, Jennifer Beck-Wilson, and Martin Wessner. The metaphor of virtual rooms in the cooperative learning environment CLear. Lecture Notes in Computer Science, 1370:107–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1370/13700107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700107.
                                        pdf.

Bender:1998:IAE

[365] Kai Bender. Integrated architecture of electronic mall systems — how strategies, processes and organizations influence information system design. Lecture Notes in Computer Science, 1370: 114–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1370/13700114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700114.
                                        pdf.

Schmidt:1998:ABT

[366] Albrecht Schmidt, Alexander Specker, Gerhard Partsch, Michael Weber, and Siegfried Höck. An agent-based telecooperation framework. Lecture Notes in Computer Science, 1370:122–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700122.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700122.
pdf.

Frazer:1998:COE

[367] John Frazer. The co-operative evolution of buildings and cities. Lecture Notes in Computer Science, 1370:130– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700130.
pdf.

Pemberton:1998:TWM

[368] Lyn Pemberton and Richard N. Griffiths. The timeless way: Making living cooperative buildings with design patterns. Lecture Notes in Computer Science, 1370:142–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700142.
pdf.

Kohler:1998:SNW

[369] Niklaus Kohler. Sustainability of new work practises and building concepts. Lecture Notes in Computer Science, 1370:154–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1370/13700154.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700154.
                                        pdf.

Moltke:1998:CBC

[370] Ivar Moltke and Hans H. K. Andersen. Cooperative buildings — the case of office VISION. Lecture Notes in Computer Science, 1370:163– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1370/13700163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700163.
                                        pdf.

Hunt:1998:FEE

[371] Robert Hunt, Andrea Vanecko, and Steven Poltrock. Future@work: An experimental exhibit investigating integrated workplace design. Lecture Notes in Computer Science, 1370:177– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1370/13700177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700177.
                                        pdf.

Lainer:1998:CQS

[372] Rüdiger Lainer and Ina Wagner. Connecting qualities of social use with spatial qualities. Lecture Notes in Computer Science, 1370:191–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700191.
pdf.

Hagen:1998:LEW

[373] Rob H. Hagen. Law enforcement of working space requirements in office buildings — the policy of the labour inspectorate in the netherlands. Lecture Notes in Computer Science, 1370: 204–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1370/13700204.
pdf.

Joeckle:1998:OST

[374] Karin Joeckle, Monika Schneiders, Thomas Sieverts, Marianne Koch, and Hartmut Chodura. Organizing space in time — discovering existing resources. Lecture Notes in Computer Science, 1370:208–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1370/13700208.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1370/13700208.
                                        pdf.

Falkowski:1998:RMS

[375] Christian Falkowski. A room management system. Lecture Notes in Computer Science, 1370:219–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1370/13700219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700219.
                                        pdf.

Junestrand:1998:DPW

[376] Stefan Junestrand and Konrad Tollmar. The dwelling as a place for work. Lecture Notes in Computer Science, 1370:230–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1370/13700230.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700230.
                                        pdf.

Hughes:1998:UTD

[377] John Hughes, Jon O’Brien, and Tom Rodden. Understanding technology in domestic environments: Lessons for cooperative buildings. Lecture Notes in Computer Science, 1370:248– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1370/13700248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1370/13700248.
                                        pdf.

Wexelblat:1998:RCG

[378] Alan Wexelblat. Research challenges in gesture: Open issues and unsolved problems. Lecture Notes in Computer Science, 1371:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710001.
pdf.

Edwards:1998:PSL

[379] Alistair D. N. Edwards. Progress in sign language recognition. Lecture Notes in Computer Science, 1371:13– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710013.
pdf.

Kita:1998:MPS

[380] Sotaro Kita, Ingeborg van Gijn, and Harry van der Hulst. Movement phases in signs and co-speech gestures, and their transcription by human coders. Lecture Notes in Computer Science, 1371:23–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710023.
pdf.

Kramer:1998:CTD

[381] Axel Kramer. Classifying two dimensional gestures in interactive systems. Lecture Notes in Computer Science, 1371:37–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1371/13710037.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1371/13710037.
pdf.

Nobe:1998:LPA

[382] Shuichi Nobe, Satoru Hayamizu, Osamu Hasegawa, and Hideaki Takahashi. Are listeners paying attention to the hand gestures of an anthropomorphic agent? an evaluation using a gaze tracking method. Lecture Notes in Computer Science, 1371:49– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710049.
pdf.

Bordegoni:1998:GBH

[383] Monica Bordegoni and Franco De Angelis. Gesture-based and haptic interaction for human skill acquisition. Lecture Notes in Computer Science, 1371: 61–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710061.
pdf.

Rigoll:1998:HPR

[384] Gerhard Rigoll, Andreas Kosmala, and Stefan Eickeler. High performance real-time gesture recognition using hidden Markov models. Lecture Notes in Computer Science, 1371:69– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710069.
pdf.

Hofmann:1998:VPB

[385] Frank G. Hofmann, Peter Heyer, and Günter Hommel. Velocity profile based recognition of dynamic gestures with discrete hidden Markov models. Lecture Notes in Computer Science, 1371: 81–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710081.
pdf.

Assan:1998:VBS

[386] Marcell Assan and Kirsti Grobel. Video-based sign language recognition using hidden Markov models. Lecture Notes in Computer Science, 1371:97– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710097.
pdf.

Gibet:1998:CNM

[387] S. Gibet, J. Richardson, T. Lebourque, and A. Braffort. Corpus of 3D natural movements and sign language primitives of movement. Lecture Notes in Computer Science, 1371:111–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1371/13710111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1371/13710111.
                                        pdf.

Munk:1998:UCP

[388] Karin Husballe Munk and Erik Granum. On the use of context and A priori knowledge in motion analysis for visual gesture recognition. Lecture Notes in Computer Science, 1371: 123–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1371/13710123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1371/13710123.
                                        pdf.

Hienz:1998:AEB

[389] Hermann Hienz and Kirsti Grobel. Automatic estimation of body regions from video images. Lecture Notes in Computer Science, 1371:135– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1371/13710135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1371/13710135.
                                        pdf.

Godenschweger:1998:RGL

[390] Frank Godenschweger, Thomas Strothotte, and Hubert Wagener. Rendering gestures as line drawings. Lecture Notes in Computer Science, 1371:147– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710147.
pdf.

Mills:1998:IRR

[391] Karen McKenzie Mills and James L. Alty. Investigating the role of redundancy in multimodal input systems. Lecture Notes in Computer Science, 1371:159–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1371/13710159.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1371/13710159.
pdf.

Froehlich:1998:GRU

[392] M. Froehlich and I. Wachsmuth. Gesture recognition of the upper limbs — from signal to symbol. Lecture Notes in Computer Science, 1371:173–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frohlich:1998:GRU

[393] Martin Fröhlich and Ipke Wachsmuth. Gesture recognition of the upper limbs — from signal to symbol. Lecture Notes in Computer Science, 1371:173– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710173.
pdf.

Latoschik:1998:EDP

[394] Marc Erich Latoschik and Ipke Wachsmuth. Exploiting distant pointing gestures for object selection in a virtual environment. Lecture Notes in Computer Science, 1371:185–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710185.
pdf.

Hummels:1998:ITH

[395] Caroline Hummels, Gerda Smets, and Kees Overbeeke. An intuitive twohanded gestural interface for computer supported product design. Lecture Notes in Computer Science, 1371:197– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710197.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710197.
pdf.

Noelker:1998:DFH

[396] C. Noelker and H. Ritter. Detection of fingertips in human hand movement sequences. Lecture Notes in Computer Science, 1371:209–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolker:1998:DFH

[397] Claudia Nölker and Helge Ritter. Detection of fingertips in human hand movement sequences. Lecture Notes in Computer Science, 1371:209–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710209.
pdf.

Boehme:1998:NAG

[398] Hans-Joachim Boehme, Anja Brakensiek, Ulf-Dietrich Braumann, Markus Krabbes, and Horst-Michael Gross. Neural architecture for gesture-based human-machine-interaction. Lecture Notes in Computer Science, 1371:219– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710219.
pdf.

Triesch:1998:RGR

[399] Jochen Triesch and Christoph von der Malsburg. Robotic gesture recognition. Lecture Notes in Computer Science, 1371:233–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1371/13710233.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/1371/13710233.
                                        pdf.

Varchmin:1998:IBR

[400] Axel Christian Varchmin, Robert Rae, and Helge Ritter. Image based recognition of gaze direction using adaptive methods. Lecture Notes in Computer Science, 1371:245–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1371/13710245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1371/13710245.
                                        pdf.

Lu:1998:TDS

[401] Shan Lu, Seiji Igi, Hideaki Matsuo, and Yuji Nagashima. Towards a dialogue system based on recognition and synthesis of Japanese Sign Language. Lecture Notes in Computer Science, 1371: 259–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1371/13710259.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1371/13710259.
                                        pdf.

Matsuo:1998:RAN

[402] Hideaki Matsuo, Seiji Igi, Shan Lu, Yuji Nagashima, Yuji Takata, and Terutaka Teshima. The recognition algorithm with non-contact for Japanese Sign Language using morphological analysis. Lecture Notes in Computer Science, 1371:273–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710273.
pdf.

Kohler:1998:STG

[403] Markus Kohler. Special topics of gesture recognition applied in intelligent home environments. Lecture Notes in Computer Science, 1371:285– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710285.
pdf.

Fjeld:1998:BII

[404] Morten Fjeld, Martin Bichsel, and Matthias Rauterberg. BUILD-IT: An intuitive design tool based on direct object manipulation. Lecture Notes in Computer Science, 1371:297– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1371/13710297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1371/13710297.
pdf.

Selcuk:1998:NRL

[405] Ali Aydin Selçuk. New results in linear cryptanalysis of RC5. Lecture Notes in Computer Science, 1372:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1372/13720001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720001.
                                        pdf.

Moriai:1998:HOD

[406] Shiho Moriai, Takeshi Shimoyama, and Toshinobu Kaneko. Higher order differential attack of a CAST cipher. Lecture Notes in Computer Science, 1372: 17–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1372/13720017.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720017.
                                        pdf.

Coppersmith:1998:CT

[407] Don Coppersmith, David Wagner, Bruce Schneier, and John Kelsey. Cryptanalysis of TWOPRIME. Lecture Notes in Computer Science, 1372: 32–48, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1372/13720032.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720032.
                                        pdf;   http://www.counterpane.com/
                                        twoprime.html.

Chabanne:1998:J

[408] Hervé Chabanne and Emmanuel Michon. JEROBOAM. Lecture Notes in Computer Science, 1372:49–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1372/13720049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1372/13720049.
pdf.

Daemen:1998:FHS

[409] J. Daemen and C. S. K. Clapp. Fast hashing and stream encryption with PANAMA. Lecture Notes in Computer Science, 1372:60–74, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1372/13720060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1372/13720060.
pdf.

Clapp:1998:JHS

[410] Craig S. K. Clapp. Joint hardware / software design of a fast stream cipher. Lecture Notes in Computer Science, 1372:75–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1372/13720075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1372/13720075.
pdf.

Abdukhalikov:1998:SHS

[411] Kanat S. Abdukhalikov and Chul Kim. On the security of the hashing scheme based on SL_{2}. Lecture Notes in Computer Science, 1372:93– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1372/13720093.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720093.
                                        pdf.

Patarin:1998:AFS

[412] Jacques Patarin. About Feistel schemes with six (or more) rounds. Lecture Notes in Computer Science, 1372:103– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1372/13720103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720103.
                                        pdf.

Young:1998:MBB

[413] Adam Young and Moti Yung. Monkey: Black-box symmetric ciphers designed for MONopolizing KEYs. Lecture Notes in Computer Science, 1372:122– 133, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1372/13720122.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720122.
                                        pdf.

Safavi-Naini:1998:MH

[414] Rei Safavi-Naini, Shahram Bakhtiari, and Chris Charnes. MRD hashing. Lecture Notes in Computer Science, 1372: 134–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1372/13720134.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720134.
                                        pdf.

Aiello:1998:NCS

[415] William Aiello, Stuart Haber, and Ramarathnam Venkatesan. New constructions for secure hash functions. Lecture Notes in Computer Science, 1372:150– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1372/13720150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1372/13720150.
pdf.

Kelsey:1998:CAP

[416] John Kelsey, Bruce Schneier, David Wagner, and Chris Hall. Cryptanalytic attacks on pseudorandom number generators. Lecture Notes in Computer Science, 1372:168–188, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1372/13720168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1372/13720168.
pdf;   http://www.counterpane.com/
pseudorandom_number.html.

Stern:1998:CC

[417] Jacques Stern and Serge Vaudenay. CS-CIPHER. Lecture Notes in Computer Science, 1372:189–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1372/13720189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1372/13720189.
pdf.

Knudsen:1998:DSR

[418] Lars R. Knudsen, Vincent Rijmen, Ronald L. Rivest, and Matthew J. B. Robshaw. On the design and security of RC2. Lecture Notes in Computer Science, 1372:206–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1372/13720206.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720206.
                                        pdf.

Biham:1998:SNB

[419] Eli Biham, Ross Anderson, and Lars R. Knudsen. Serpent: A new block cipher proposal. Lecture Notes in Computer Science, 1372:222–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1372/13720222.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720222.
                                        pdf.

Lucks:1998:ATE

[420] Stefan Lucks. Attacking triple encryption. Lecture Notes in Computer Science, 1372:239–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1372/13720239.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1372/13720239.pdf.

Wagner:1998:CSR

[421] David Wagner. Cryptanalysis of some recently-proposed multiple modes of operation. Lecture Notes in Computer Science, 1372:254–269, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1372/13720254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1372/13720254.
pdf.

Rompay:1998:DCI

[422] Bart Van Rompay, Lars R. Knudsen, and Vincent Rijmen. Differential cryptanalysis of the ICE encryption algorithm. Lecture Notes in Computer Science, 1372:270–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1372/13720270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1372/13720270.
pdf.

VanRompay:1998:DCI

[423] B. Van Rompay, L. R. Knudsen, and V. Rijmen. Differential cryptanalysis of the ICE encryption algorithm. Lecture Notes in Computer Science, 1372:270–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobbertin:1998:FTR

[424] Hans Dobbertin. The first two rounds of MD4 are not one-way. Lecture Notes in Computer Science, 1372:284– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1372/13720284.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720284.
                                        pdf.

Wagner:1998:DCK

[425] David Wagner. Differential cryptanalysis of KHF. Lecture Notes in Computer Science, 1372:293–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1372/13720293.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1372/13720293.
                                        pdf.

Karp:1998:RGR

[426] R. M. Karp. Random graphs, random walks, differential equations and the probabilistic analysis of algorithms. Lecture Notes in Computer Science, 1373:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssen:1998:DOF

[427] J. Janssen, D. Krizanc, L. Narayanan, and S. Shende. Distributed online frequency assignment in cellular networks. Lecture Notes in Computer Science, 1373:3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thorup:1998:FIS

[428] M. Thorup. Floats, integers, and single source shortest paths. Lecture Notes in Computer Science, 1373:14–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habib:1998:SPR

[429] M. Habib, C. Paul, and L. Viennot. A synthesis on partition refinement: A useful routine for strings, graphs, Boolean matrices and automata. Lecture Notes in Computer Science, 1373: 25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradfield:1998:SMC

[430] J. C. Bradfield. Simplifying the modal λ-calculus alternation hierarchy. Lecture Notes in Computer Science, 1373: 39–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1998:DDH

[431] T. Eiter, T. Ibaraki, and K. Makino. On disguised double Horn functions and extensions. Lecture Notes in Computer Science, 1373:50–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demri:1998:CPL

[432] S. Demri and P. Schnoebelen. The complexity of propositional linear temporal logics in simple cases. Lecture Notes in Computer Science, 1373:61–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrington:1998:SCW

[433] D. A. M. Barrington, C.-J. Lu, P. B. Miltersen, and S. Skyum. Searching constant width mazes captures the AC00 hierarchy. Lecture Notes in Computer Science, 1373:73–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fortnow:1998:NOL

[434] L. Fortnow and S. Laplante. Nearly optimal language compression using extractors. Lecture Notes in Computer Science, 1373:84–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spencer:1998:RSB

[435] J. H. Spencer and K. St. John. Random sparse bit strings at the threshold of adjacency. Lecture Notes in Computer Science, 1373:94–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sauerhoff:1998:LBR

[436] M. Sauerhoff. Lower bounds for randomized read-k-times branching programs. Lecture Notes in Computer Science, 1373:105–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mazoyer:1998:IOC

[437] J. Mazoyer and I. Rapaport. Inducing an order on cellular automata by a grouping operation. Lecture Notes in Computer Science, 1373:116–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manzini:1998:ADD

[438] G. Manzini and L. Margara. Attractors of D-dimensional linear cellular automata. Lecture Notes in Computer Science, 1373:128–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mereghetti:1998:OSB

[439] C. Mereghetti and G. Pighizzini. Optimal simulations between unary automata. Lecture Notes in Computer Science, 1373:139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mateescu:1998:SOW

[440] A. Mateescu. Shuffle of omega-words: Algebraic aspects. Lecture Notes in Computer Science, 1373:150–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhrman:1998:GRB

[441] H. Buhrman, D. Van Melkebeek, K. W. Regan, and D. Sivakumar. A generalization of resource-bounded measure, with an application. Lecture Notes in Computer Science, 1373:161–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1998:CMG

[442] V. Arvind, R. Beigel, and A. Lozano. The complexity of modular graph automorphism. Lecture Notes in Computer Science, 1373:172–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Libkin:1998:UQT

[443] L. Libkin and L. Wong. Unary quantifiers, transitive closure, and relations of large degree. Lecture Notes in Computer Science, 1373:183–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buergisser:1998:SVC

[444] P. Buergisser. On the structure of Valiant’s complexity classes. Lecture Notes in Computer Science, 1373:194– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sieling:1998:EPT

[445] D. Sieling. On the existence of polynomial time approximation schemes for OBDD-Minimization. Lecture Notes in Computer Science, 1373:205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feigenbaum:1998:CPG

[446] J. Feigenbaum, S. Kannan, M. Y. Vardi, and M. Viswanathan. Complexity of problems on graphs represented as OBDDs. Lecture Notes in Computer Science, 1373:216–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Behrens:1998:ETO

[447] J. Behrens and S. Waack. Equivalence test and ordering transformation for Parity-OBDDs of different variable ordering. Lecture Notes in Computer Science, 1373:227–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groepl:1998:SSR

[448] C. Groepl, H. J. Proemel, and A. Srivastav. Size and structure of random ordered binary decision diagrams. Lecture Notes in Computer Science, 1373: 238–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaudenay:1998:PSB

[449] S. Vaudenay. Provable security for block ciphers by decorrelation. Lecture Notes in Computer Science, 1373:249– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bazgan:1998:AFN

[450] C. Bazgan, M. Santha, and Z. Tuza. On the approximation of finding a(nother) Hamiltonian cycle in cubic Hamiltonian graphs. Lecture Notes in Computer Science, 1373:276–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1998:MES

[451] K. Jansen. The mutual exclusion scheduling problem for permutation and comparability graphs. Lecture Notes in Computer Science, 1373:287– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bshouty:1998:MLP

[452] N. H. Bshouty and L. J. Burroughs. Massaging a linear programming solution to give a 2-approximation for a generalization of the vertex cover problem. Lecture Notes in Computer Science, 1373:298–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1998:PPS

[453] K. S. Larsen. Partially persistent search trees with transcript operations. Lecture Notes in Computer Science, 1373:309–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niwinski:1998:RHW

[454] D. Niwinski and I. Walukiewicz. Relating hierarchies of word and tree automata. Lecture Notes in Computer Science, 1373:320–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Straubing:1998:LDM

[455] H. Straubing. Languages defined with modular counting quantifiers. Lecture Notes in Computer Science, 1373:332– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jantzen:1998:HPT

[456] M. Jantzen. Hierarchies of principal twist-closed trios. Lecture Notes in Computer Science, 1373:344–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Safer:1998:RRA

[457] T. Safer. Radix representations of algebraic number fields and finite automata. Lecture Notes in Computer Science, 1373:356–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagerup:1998:SSW

[458] T. Hagerup. Sorting and searching on the word RAM. Lecture Notes in Computer Science, 1373:366–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diallo:1998:CED

[459] M. Diallo, A. Ferreira, and A. Rau-Chaplin. Communication-efficient deterministic parallel algorithms for planar point location and 2D Voronoi diagram. Lecture Notes in Computer Science, 1373:399–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rueb:1998:BMS

[460] C. Rueb. On Batcher’s merge sorts as parallel sorting algorithms. Lecture Notes in Computer Science, 1373:410– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustedt:1998:MST

[461] J. Gustedt. Minimum spanning trees for minor-closed graph classes in parallel. Lecture Notes in Computer Science, 1373:421–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dessmark:1998:OBA

[462] A. Dessmark, A. Lingas, H. Olsson, and H. Yamamoto. Optimal broadcasting in almost trees and partial k-trees. Lecture Notes in Computer Science, 1373:432–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwentick:1998:LNF

[463] T. Schwentick and K. Barthelmann. Local normal forms for first-order logic with applications to games and automata. Lecture Notes in Computer Science, 1373:444–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esik:1998:AET

[464] Z. Esik. Axiomatizing the equational theory of regular tree languages. Lecture Notes in Computer Science, 1373: 455–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monti:1998:LCS

[465] A. Monti and A. Peron. A logical characterization of systolic languages. Lecture Notes in Computer Science, 1373:466–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Messner:1998:OPS

[466] J. Messner and J. Toran. Optimal proof systems for propositional logic and complete sets. Lecture Notes in Computer Science, 1373:477–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serna:1998:PAN

[467] M. Serna, L. Trevisan, and F. Xhafa. The (parallel) approximability of non-Boolean satisfiability problems and restricted integer programming. Lecture Notes in Computer Science, 1373:488– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ivanov:1998:IPR

[468] S. Ivanov and M. De Rougemont. Interactive protocols on the reals. Lecture Notes in Computer Science, 1373: 499–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiCrescenzo:1998:RIZ

[469] G. Di Crescenzo, K. Sakurai, and M. Yung. Result-indistinguishable zero-knowledge proofs: Increased power and constant-round protocols. Lecture Notes in Computer Science, 1373:511–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAgostino:1998:BSD

[470] S. De Agostino and R. Silvestri. Bounded size dictionary compression: SC0k-Completeness and NC algorithms. Lecture Notes in Computer Science, 1373:522–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1998:ECL

[471] R. Meyer and A. Petit. Expressive completeness of LTrL on finite traces: An algebraic proof. Lecture Notes in Computer Science, 1373:533–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frid:1998:UDW

[472] A. E. Frid. On the uniform DOL words. Lecture Notes in Computer Science, 1373:544–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lodaya:1998:SPP

[473] K. Lodaya and P. Weil. Series-parallel posets: Algebra, automata and languages. Lecture Notes in Computer Science, 1373:555–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemp:1998:ENN

[474] R. Kemp. On the expected number of nodes at level k in 0-balanced trees. Lecture Notes in Computer Science, 1373:566–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golumbic:1998:CFP

[475] M. C. Golumbic and H. Kaplan. Cell flipping in permutation diagrams. Lecture Notes in Computer Science, 1373: 577–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorkenoo:1998:CNI

[476] M. Dorkenoo, M.-C. Eglin-Leclerc, and E. Remila. Construction of nonintersecting colored flows through a planar cellular figure. Lecture Notes in Computer Science, 1373:587–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calude:1998:RER

[477] C. S. Calude, P. H. Hertling, B. Khoussainov, and Y. Wang. Recursively enumerable reals and Chaitin omega numbers. Lecture Notes in Computer Science, 1373:596–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosub:1998:UDC

[478] S. Kosub, H. Schmitz, and H. Vollmer. Uniformly defining complexity classes of functions. Lecture Notes in Computer Science, 1373:607–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lapoire:1998:REM

[479] D. Lapoire. Recognizability equals monadic second-order derinability, for sets of graphs of bounded tree-width. Lecture Notes in Computer Science, 1373:618–??, 1998. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Bunt:1998:IMH

[480] Harry Bunt. Issues in multimodal human-computer communication. Lecture Notes in Computer Science, 1374: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1374/13740001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1374/13740001.
pdf.

Maybury:1998:TCM

[481] Mark T. Maybury. Toward cooperative multimedia interaction. Lecture Notes in Computer Science, 1374:13– 38, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1374/13740013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1374/13740013.
pdf.

Bunt:1998:MCD

[482] Harry Bunt, René Ahn, Robbert-Jan Beun, Tijn Borghuis, and Kees van Overveld. Multimodal cooperation with the DenK system. Lecture Notes in Computer Science, 1374:39– 67, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1374/13740039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1374/13740039.
pdf.

Pelachaud:1998:SCC

[483] Catherine Pelachaud, Justine Cassell, Norman Badler, Mark Steedman, Scott Prevost, and Matthew Stone. Synthesizing cooperative conversation. Lecture Notes in Computer Science, 1374: 68–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1374/13740068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1374/13740068.
                                        pdf.

Webber:1998:IAA

[484] Bonnie Webber. Instructing animated agents: Viewing language in behavioral terms. Lecture Notes in Computer Science, 1374:89–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1374/13740089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1374/13740089.
                                        pdf.

Siroux:1998:MPO

[485] Jacques Siroux, Marc Guyomard, Franck Multon, and Christophe Remondeau. Modeling and processing of oral and tactile activities in the GEORAL system. Lecture Notes in Computer Science, 1374:101–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1374/13740101.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1374/13740101.
pdf.

Cheyer:1998:MMA

[486] Adam Cheyer and Luc Julia. Multimodal maps: An agent-based approach. Lecture Notes in Computer Science, 1374:111–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1374/13740111.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1374/13740111.pdf.

Hah:1998:UCA

[487] Y. Hah and I. Zukerman. Using cooperative agents to plan multimodal presentations. Lecture Notes in Computer Science, 1374:122–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1998:UCA

[488] Yi Han and Ingrid Zukerman. Using cooperative agents to plan multimodal presentations. Lecture Notes in Computer Science, 1374:122–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1374/13740122.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1374/13740122.
pdf.

Martin:1998:DMI

[489] J. C. Martin, R. Veldman, and D. Béroule. Developing multimodal interfaces: A theoretical framework and guided propagation networks. Lecture Notes in Computer Science, 1374:158– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1374/13740158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1374/13740158.
pdf.

Bourdot:1998:CBR

[490] Patrick Bourdot, Mike Krus, and Rachid Gherbi. Cooperation between reactive 3D objects and a multimodal X Window kernel for CAD. Lecture Notes in Computer Science, 1374:188– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1374/13740188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1374/13740188.
pdf.

McCaffery:1998:MIC

[491] Fergal McCaffery, Michael McTear, and Maureen Murphy. A multimedia interface for circuit board assembly. Lecture Notes in Computer Science, 1374:213–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1374/13740213.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1374/13740213.
pdf.

Wittenburg:1998:VLP

[492] Kent Wittenburg. Visual language parsing: If I had a hammer. … Lecture Notes in Computer Science, 1374:

231–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1374/13740231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1374/13740231.
pdf.

Lee:1998:AMD

[493] John Lee and Keith Stenning. Anaphora in multimodal discourse. Lecture Notes in Computer Science, 1374:250– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1374/13740250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1374/13740250.
pdf.

Fais:1998:SRR

[494] Laurel Fais, Kyung ho Loken-Kim, and Young-Duk Park. Speakers’ responses to requests for repetition in a multimedia language processing environment. Lecture Notes in Computer Science, 1374:264–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1374/13740264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1374/13740264.
pdf.

Cremers:1998:ORT

[495] Anita Cremers. Object reference in task-oriented keyboard dialogues. Lecture Notes in Computer Science, 1374: 279–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1374/13740279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1374/13740279.
                                        pdf.

Kato:1998:RIR

[496] Tsuneaki Kato and Yukiko I. Nakano. Referent identification requests in multi-modal dialogs. Lecture Notes in Computer Science, 1374:294–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1374/13740294.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1374/13740294.
                                        pdf.

Huls:1998:SFI

[497] Carla Huls and Edwin Bos. Studies into full integration of language and action. Lecture Notes in Computer Science, 1374:312–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1374/13740312.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1374/13740312.
                                        pdf.

Bressolle:1998:RMC

[498] Marie-Christine Bressolle, Bernard Pavard, and Marcel Leroux. The role of multimodal communication in cooperation: The cases of air traffic control. Lecture Notes in Computer Science, 1374:326–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1374/13740326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1374/13740326.
pdf.

Rudaz:1998:IID

[499] N. Rudaz, R. D. Hersch, and V. Ostromoukhov. An interface for the interactive design of artistic screens. Lecture Notes in Computer Science, 1375: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750001.
pdf.

Finkelstein:1998:IM

[500] Adam Finkelstein and Marisa Range. Image mosaics. Lecture Notes in Computer Science, 1375:11–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750011.
pdf.

Browne:1998:FDA

[501] Cameron Browne. Font decoration by automatic mesh fitting. Lecture Notes in Computer Science, 1375:23– 43, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://diwww.epfl.

                                        ch/w3lsp/ridt98.html;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1375/13750023.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1375/13750023.pdf.

Hertz:1998:WBM

[502] Jacky Hertz, Changyuan Hu, Jakob Gonczarowski, and Roger D. Hersch. A window-based method for automatic typographic parameter extraction. Lecture Notes in Computer Science, 1375: 44–54, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750044.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750044.
                                        pdf.

Kang:1998:UTA

[503] ByungHoon Kang. Unified table approach for typographic rendering. Lecture Notes in Computer Science, 1375: 55–65, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750055.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750055.
                                        pdf.

Gonczarowski:1998:PSC

[504] Jakob Gonczarowski. Producing the skeleton of a character. Lecture Notes in Computer Science, 1375:66– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1375/13750066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750066.
pdf.

Stamm:1998:VTG

[505] Beat Stamm. Visual TrueType: A graphical method for authoring font intelligence. Lecture Notes in Computer Science, 1375:77–92, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750077.
pdf.

Shamir:1998:FBD

[506] Ariel Shamir and Ari Rappoport. Feature-based design of fonts using constraints. Lecture Notes in Computer Science, 1375:93–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750093.
pdf.

Schneider:1998:OOM

[507] Uwe Schneider. An object-oriented model for the hierarchical composition of letterforms in computeraided typeface design. Lecture Notes in Computer Science, 1375:109–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1375/13750109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750109.
                                        pdf.

Haralambous:1998:DUM

[508] Yannis Haralambous and John Plaice. The design and use of a multiplealphabet font with Ω. Lecture Notes in Computer Science, 1375:126–137, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750126.
                                        pdf.

Haralambous:1998:SAS

[509] Yannis Haralambous. Simplification of the Arabic script: Three different approaches and their implementations. Lecture Notes in Computer Science, 1375:138–156, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750138.
                                        pdf.

Mohanty:1998:FPT

[510] S. K. Mohanty. The formulation of parameters for type design of Indian scripts based on calligraphic studies. Lecture Notes in Computer Science, 1375:157–166, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1375/13750157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750157.
pdf.

Southall:1998:MRC

[511] Richard Southall. Metafont in the Rockies: The Colorado Typemaking Project. Lecture Notes in Computer Science, 1375:167–180, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750167.
pdf.

Connare:1998:DUS

[512] Vincent Connare. Designing and using scalable type for display as graphic elements on digital devices. Lecture Notes in Computer Science, 1375:181– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750181.
pdf.

Ostromoukhov:1998:MTC

[513] Victor Ostromoukhov. Mathematical tools for computer-generated ornamental patterns. Lecture Notes in Computer Science, 1375:193–223, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1375/13750193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750193.
                                        pdf.

Lou:1998:FH

[514] Qun Lou and Peter Stucki. Fundamentals of 3D halftoning. Lecture Notes in Computer Science, 1375:224– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750224.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750224.
                                        pdf.

Bapst:1998:UTD

[515] Frédéric Bapst and Rolf Ingold. Using typography in document image analysis. Lecture Notes in Computer Science, 1375:240–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1375/13750240.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750240.
                                        pdf.

Lee:1998:BVM

[516] Do-Hoon Lee and Hwan-Gue Cho. The beta-velocity model for simulating handwritten Korean scripts. Lecture Notes in Computer Science, 1375: 252–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750252.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1375/13750252.
pdf.

Karow:1998:TDT

[517] Peter Karow. Two decades of typographic research at URW: A retrospective. Lecture Notes in Computer Science, 1375:265–280, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750265.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750265.
pdf.

Morris:1998:LCP

[518] R. A. Morris, R. D. Hersch, and A. Coimbra. Legibility of condensed perceptually-tuned grayscale fonts. Lecture Notes in Computer Science, 1375:281–293, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1375/13750281.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1375/13750281.pdf.

Dyson:1998:EEL

[519] Mary C. Dyson and Gary J. Kipping. Exploring the effect of layout on reading from screen. Lecture Notes in Computer Science, 1375:294– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750294.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1375/13750294.
pdf.

Kataoka:1998:ITM

[520] Tomoko I. Kataoka, Yutaka Kataoka, Kazutomo Uezono, and Hiroyoshi Ohara. Internationalized text manipulation covering Perso-Arabic enhanced for Mongolian scripts. Lecture Notes in Computer Science, 1375:305– 318, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750305.
pdf.

Bouzaiene:1998:BIP

[521] Nabil Bouzaiene, Loïc Le Gall, and Emmanuel Saint-James. Une bibliothèque informatique pour la notation musicale baroque. Lecture Notes in Computer Science, 1375:319– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750319.
pdf.

Richy:1998:DSD

[522] Hélène Richy. Document style design by direct manipulation. Lecture Notes in Computer Science, 1375:331– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750331.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1375/13750331.
pdf.

Brugger:1998:DED

[523] Rolf Brugger, Frédéric Bapst, and Rolf Ingold. A DTD extension for document structure recognition. Lecture Notes in Computer Science, 1375:343– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750343.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750343.
pdf.

King:1998:SMD

[524] Peter King, Helen Cameron, Howard Bowman, and Simon Thompson. Synchronization in multimedia documents. Lecture Notes in Computer Science, 1375:355–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1375/13750355.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1375/13750355.
pdf.

Jourdan:1998:EVI

[525] Muriel Jourdan, Cécile Roisin, and Laurent Tardif. Édition et visualisation interactive de documents multimédia. Lecture Notes in Computer Science, 1375:370–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1375/13750370.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/1375/13750370.
                                        pdf.

Cumaranatunge:1998:MSC

[526] Terry Cumaranatunge and Ethan V. Munson. Memory scalability in constraint-based multimedia style sheet systems. Lecture Notes in Computer Science, 1375:381–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750381.
                                        pdf.

Dymetman:1998:IP

[527] Marc Dymetman and Max Copperman. Intelligent paper. Lecture Notes in Computer Science, 1375:392– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750392.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750392.
                                        pdf.

Brown:1998:AAU

[528] Heather Brown, Robert Harding, Steven Lay, Peter Robinson, Dan Sheppard, and Richard Watts. Active Alice: Using real paper to interact with electronic text. Lecture Notes in Computer Science, 1375:407– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750407.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1375/13750407.
pdf.

Taghva:1998:ATC

[529] Kazem Taghva, Allen Condit, and Julie Borsack. Autotag: A tool for creating structured document collections from printed materials. Lecture Notes in Computer Science, 1375:420– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750420.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750420.
pdf.

Froehitch:1998:HDM

[530] P. Froehitch, N. Henze, and W. Nejdl. Hyperbook data modeling. Lecture Notes in Computer Science, 1375: 432–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frohlich:1998:HDM

[531] Peter Fröhlich, Nicola Henze, and Wolfgang Nejdl. Hyperbook data modeling. Lecture Notes in Computer Science, 1375:432–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1375/13750432.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1375/13750432.
pdf.

Brettbauer:1998:CAC

[532] T. Brettbauer, A. Brüggemann-Klein, S. Hermann, and R. Klein. A component architecture for cross-media formatters. Lecture Notes in Computer Science, 1375:444–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bretthauer:1998:CAC

[533] Thomas Bretthauer, Anne Brüggemann-Klein, Stefan Hermann, and Rolf Klein. A component architecture for cross-media formatters. Lecture Notes in Computer Science, 1375:444– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750444.
pdf.

Brooks:1998:ETA

[534] Terrence A. Brooks. The effect of textual aesthetics on information retrieval. Lecture Notes in Computer Science, 1375:454–463, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750454.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750454.
pdf.

Ballim:1998:UAA

[535] A. Ballim, G. Coray, A. Linden, and C. Vanoirbeek. The use of automatic alignment on structured multilingual documents. Lecture Notes in Computer Science, 1375:464–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750464.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750464.
pdf.

Ahonen:1998:DID

[536] Helena Ahonen, Barbara Heikkinen, Oskari Heinonen, Jani Jaakkola, Pekka Kilpeläinen, and Greger Lindén. Design and implementation of a document assembly workbench. Lecture Notes in Computer Science, 1375:476– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750476.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750476.
pdf.

Paradis:1998:VDI

[537] François Paradis, Anne-Marie Vercoustre, and Brendan Hills. A virtual document interpreter for reuse of information. Lecture Notes in Computer Science, 1375:487–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750487.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750487.
pdf.

Vercoustre:1998:RLD

[538] Anne-Marie Vercoustre and François Paradis. Reuse of linked documents through virtual document prescriptions. Lecture Notes in Computer Science, 1375:499–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1375/13750499.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1375/13750499.pdf.

Cruz:1998:MSS

[539] Isabel F. Cruz, Slava Borisov, Michael A. Marks, and Timothy R. Webb. Measuring structural similarity among Web documents: Preliminary results. Lecture Notes in Computer Science, 1375:513–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750513.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750513.
                                        pdf.

Hu:1998:TMM

[540] Shueh-Cheng Hu and Richard Furuta. A tool for maintaining multivariant hypertext documents. Lecture Notes in Computer Science, 1375:525– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1375/13750525.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1375/13750525.
                                        pdf.

Rhissassi:1998:PHP

[541] Habib Rhissassi and Alain Lelu. Projet Hypermap: pour un environnement complet de génération automatique d’hypertexte. Lecture Notes in Computer Science, 1375:537–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750537.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750537.
pdf.

Probets:1998:DLI

[542] Steve Probets, David F. Brailsford, Les Carr, and Wendy Hall. Dynamic link inclusion in online PDF journals. Lecture Notes in Computer Science, 1375: 550–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1375/13750550.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1375/13750550.
pdf.

Mazumdar:1998:ASS

[543] Subhasish Mazumdar, Gary Yuan, Weifeng Bao, and Jonathan Price. Adding semantics to SGML databases. Lecture Notes in Computer Science, 1375:563–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1375/13750563.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1375/13750563.
pdf.

Ehrig:1998:ADT

[544] H. Ehrig, M. Gajewsky, and U. Wolter. From abstract data types to algebraic development techniques: A shift of paradigms. Lecture Notes in Computer Science, 1376:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meseguer:1998:MAL

[545] J. Meseguer. Membership algebra as a logical framework for equational specification. Lecture Notes in Computer Science, 1376:18–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meseguer:1998:MTL

[546] J. Meseguer and U. Montanari. Mapping tile logic into rewriting logic. Lecture Notes in Computer Science, 1376: 62–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ancona:1998:AMM

[547] D. Ancona and E. Zucca. An algebra of mixin modules. Lecture Notes in Computer Science, 1376:92–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borzyszkowski:1998:CLS

[548] T. Borzyszkowski. Correctness of a logical system for structured specifications. Lecture Notes in Computer Science, 1376:107–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruni:1998:ZSN

[549] R. Bruni and U. Montanari. Zerosafe nets: The individual token approach. Lecture Notes in Computer Science, 1376:122–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerioli:1998:IDP

[550] M. Cerioli and E. Zucca. Implementation of derived programs (almost) for free. Lecture Notes in Computer Science, 1376:141–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherki:1998:MFP

[551] S. Cherki and C. Choppy. A method for FORTRAN programs reverse engineering using algebraic specifications. Lecture Notes in Computer Science, 1376: 156–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cirstea:1998:CSH

[552] C. Cirstea. Coalgebra semantics for hidden algebra: Parameterised objects and inheritance. Lecture Notes in Computer Science, 1376:174–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1998:CRE

[553] A. Corradini. A completeness result for equational deduction in coalgebraic specification. Lecture Notes in Computer Science, 1376:190–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dionisio:1998:SDC

[554] F. M. Dionisio and U. W. Lipeck. Specifying with defaults: Compositional semantics. Lecture Notes in Computer Science, 1376:206–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gadducci:1998:IVG

[555] F. Gadducci and R. Heckel. An inductive view of graph transformation. Lecture Notes in Computer Science, 1376:223–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1998:CSF

[556] M. Gogolla and M. Richters. On combining semiformal and formal object specification techniques. Lecture Notes in Computer Science, 1376:238– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gramlich:1998:MAR

[557] B. Gramlich. Modular aspects of rewrite-based specifications. Lecture Notes in Computer Science, 1376:253– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosse-Rhode:1998:ATL

[558] M. Grosse-Rhode. From algebra transformation to labelled transition systems. Lecture Notes in Computer Science, 1376:269–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lasota:1998:OMB

[559] S. Lasota. Open maps as a bridge between algebraic observational equivalence and bisimilarity. Lecture Notes in Computer Science, 1376:285–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martini:1998:SSM

[560] A. Martini and U. Welter. A systematic study of mappings between institutions. Lecture Notes in Computer Science, 1376:300–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossakowski:1998:COS

[561] T. Mossakowski. Colimits of ordersorted specifications. Lecture Notes in Computer Science, 1376:316–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossakowski:1998:SSA

[562] T. Mossakowski and B. Krieg-Brueckner. Static semantic analysis and theorem proving for CASL. Lecture Notes in Computer Science, 1376: 333–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossakowski:1998:CRL

[563] T. Mossakowski, A. Tarlecki, and W. Pawtowski. Combining and representing logical systems using modeltheoretic parchments. Lecture Notes in Computer Science, 1376:349–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padawitz:1998:TOT

[564] P. Padawitz. Towards the one-tiered design of data types and transition systems. Lecture Notes in Computer Science, 1376:365–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pawtowski:1998:CP

[565] W. Pawtowski. Context parchments. Lecture Notes in Computer Science, 1376:381–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reus:1998:VCO

[566] B. Reus, A. Knapp, P. Cenciarelli, and M. Wirsing. Verifying a compiler optimization for multi-threaded Java. Lecture Notes in Computer Science, 1376:402–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walicki:1998:CRS

[567] M. Walicki and M. Bialasik. Categories of relational structures. Lecture Notes in Computer Science, 1376:418– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weikum:1998:UIS

[568] Gerhard Weikum. On the ubiquity of information services and the absence of guaranteed service quality. Lecture Notes in Computer Science, 1377:3– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770003.
pdf.

Ciaccia:1998:PCS

[569] Paolo Ciaccia, Marco Patella, and Pavel Zezula. Processing complex similarity queries with distancebased access methods. Lecture Notes in Computer Science, 1377:9– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770009.
pdf.

Fuhr:1998:HPI

[570] Norbert Fuhr and Thomas Rölleke. HySpirit — A probabilistic inference engine for hypermedia retrieval in large databases. Lecture Notes in Computer Science, 1377:24–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770024.
pdf.

Bertino:1998:TOI

[571] E. Bertino, B. Catania, and B. Shidlovsky. Towards optimal indexing for segment databases. Lecture Notes in Computer Science, 1377:39–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770039.
pdf.

Yerneni:1998:FQI

[572] Ramana Yerneni, Yannis Papakonstantinou, Serge Abiteboul, and Hector Garcia-Molina. Fusion queries over Internet databases. Lecture Notes in Computer Science, 1377:57– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770057.
                                        pdf.

Mecca:1998:EQW

[573] Giansalvatore Mecca, Alberto O. Mendelzon, and Paolo Merialdo. Efficient queries over Web views. Lecture Notes in Computer Science, 1377:72– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770072.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770072.
                                        pdf.

Mihaila:1998:ETD

[574] George A. Mihaila, Louiqa Raschid, and Anthony Tomasic. Equal time for data on the Internet with WebSemantics. Lecture Notes in Computer Science, 1377:87–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770087.
                                        pdf.

Lin:1998:PSN

[575] Dao-I Lin and Zvi M. Kedem. Pincer search: A new algorithm for discovering the maximum frequent set. Lecture Notes in Computer Science, 1377: 105–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770105.
pdf.

Shao:1998:MMO

[576] Shin-Chung Shao. Multivariate and multidimensional OLAP. Lecture Notes in Computer Science, 1377:120–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770120.
pdf.

Ester:1998:IGM

[577] Martin Ester and Rüdiger Wittmann. Incremental generalization for mining in a data warehousing environment. Lecture Notes in Computer Science, 1377:135–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1377/13770135.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1377/13770135.
pdf.

Lehner:1998:MLS

[578] Wolfgang Lehner. Modeling large scale OLAP scenarios. Lecture Notes in Computer Science, 1377:153–167, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770153.
                                        pdf.

Sarawagi:1998:DDE

[579] Sunita Sarawagi, Rakesh Agrawal, and Nimrod Megiddo. Discovery-driven exploration of OLAP data cubes. Lecture Notes in Computer Science, 1377: 168–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770168.
                                        pdf.

Cabibbo:1998:LAM

[580] Luca Cabibbo and Riccardo Torlone. A logical approach to multidimensional databases. Lecture Notes in Computer Science, 1377:183–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770183.
                                        pdf.

Scheufele:1998:EDP

[581] Wolfgang Scheufele and Guido Moerkotte. Efficient dynamic programming algorithms for ordering expensive joins and selections. Lecture Notes in Computer Science, 1377:201– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770201.
pdf.

Berchtold:1998:IQP

[582] Stefan Berchtold, Christian Böhm, and Hans-Peter Kriegel. Improving the query performance of highdimensional index structures by bulk load operations. Lecture Notes in Computer Science, 1377:216–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770216.
pdf.

Al-Marri:1998:EAD

[583] Jaber A. Al-Marri and Shahram Ghandeharizadeh. An evaluation of alternative disk scheduling techniques in support of variable bit rate continuous media. Lecture Notes in Computer Science, 1377:231–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770231.
pdf.

Feng:1998:BMD

[584] L. Feng, H. Lu, Y. C. Tay, and K. H. Tung. Buffer management in distributed database systems: A data mining-based approach. Lecture Notes in Computer Science, 1377:246– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770246.
                                        pdf.

Ross:1998:CAM

[585] Kenneth A. Ross, Divesh Srivastava, and Damianos Chatziantoniou. Complex aggregation at multiple granularities. Lecture Notes in Computer Science, 1377:263–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770263.
                                        pdf.

Fukuda:1998:PPM

[586] T. Fukuda and H. Matsuzawa. Parallel processing of multiple aggregate queries on shared-nothing multiprocessors. Lecture Notes in Computer Science, 1377:278–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takeshi:1998:PPM

[587] Fukuda Takeshi and Hirofumi Matsuzawa. Parallel processing of multiple aggregate queries on sharednothing multiprocessors. Lecture Notes in Computer Science, 1377:278–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770278.
pdf.

Akinde:1998:MDD

[588] M. O. Akinde, O. G. Jensen, and M. H. Böhlen. Minimizing detail data in data warehouses. Lecture Notes in Computer Science, 1377:293– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770293.
pdf.

Benzaken:1998:SMI

[589] Véronique Benzaken and Xavier Schaefer. Static management of integrity in object-oriented databases: Design and implementation. Lecture Notes in Computer Science, 1377:311–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770311.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770311.
pdf.

Darmont:1998:OGB

[590] Jérôme Darmont, Bertrand Petit, and Michel Schneider. OCB: A generic benchmark to evaluate the performances of object-oriented database systems. Lecture Notes in Computer Science, 1377:326–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770326.
                                        pdf.

Lee:1998:PRT

[591] Sin Yeung Lee and Tok Wang Ling. A path removing technique for detecting trigger termination. Lecture Notes in Computer Science, 1377:341– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770341.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770341.
                                        pdf.

Nica:1998:CAV

[592] Anisoara Nica, Amy J. Lee, and Elke A. Rundensteiner. The CVS algorithm for view synchronization in evolvable large-scale information systems. Lecture Notes in Computer Science, 1377: 359–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770359.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770359.
                                        pdf.

Vista:1998:IIV

[593] Dimitra Vista. Integration of incremental view maintenance into query optimizers. Lecture Notes in Computer Science, 1377:374–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770374.
pdf.

Yang:1998:MTV

[594] Jun Yang and Jennifer Widom. Maintaining temporal views over nontemporal information sources for data warehousing. Lecture Notes in Computer Science, 1377:389–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770389.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770389.
pdf.

Ludaescher:1998:RAL

[595] B. Ludaescher and W. May. Referential actions: From logical semantics to implementation. Lecture Notes in Computer Science, 1377:404–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludascher:1998:RAL

[596] Bertram Ludäscher and Wolfgang May. Referential actions: From logical semantics to implementation. Lecture Notes in Computer Science, 1377:404– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770404.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770404.
                                        pdf.

Fraternali:1998:CMT

[597] Piero Fraternali and Paolo Paolini. A conceptual model and a tool environment for developing more scalable, dynamic, and customizable Web applications. Lecture Notes in Computer Science, 1377:421–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770421.
                                        pdf.

Atzeni:1998:DMD

[598] Paolo Atzeni, Giansalvatore Mecca, and Paolo Merialdo. Design and maintenance of data-intensive Web sites. Lecture Notes in Computer Science, 1377:436–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1377/13770436.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770436.
                                        pdf.

Toyama:1998:DSP

[599] Motomichi Toyama and Takuhiro Nagafuji. Dynamic and structured presentation of database contents on the Web. Lecture Notes in Computer Science, 1377:451–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1377/13770451.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770451.
pdf.

Agrawal:1998:MPM

[600] Rakesh Agrawal, Dimitrios Gunopulos, and Frank Leymann. Mining process models from workflow logs. Lecture Notes in Computer Science, 1377:469– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770469.
pdf.

Tang:1998:SSI

[601] Jian Tang and San-Yih Hwang. A scheme to specify and implement adhoc recovery in workflow systems. Lecture Notes in Computer Science, 1377: 484–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1377/13770484.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1377/13770484.
pdf.

Kosky:1998:EHB

[602] Anthony S. Kosky, I-Min A. Chen, Victor M. Markowitz, and Ernest Szeto. Exploring heterogeneous biological databases: Tools and applications. Lecture Notes in Computer Science, 1377: 499–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1377/13770499.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1377/13770499.
                                        pdf.

Mislove:1998:GDT

[603] Michael Mislove. Generalizing domain theory. Lecture Notes in Computer Science, 1378:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1378/13780001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1378/13780001.
                                        pdf.

Aceto:1998:CTE

[604] Luca Aceto, Wan Fokkink, and Anna Ingólfsdóttir. A cook’s tour of equational axiomatizations for prefix iteration. Lecture Notes in Computer Science, 1378:20–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1378/13780020.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1378/13780020.
                                        pdf.

Albayrak:1998:WHP

[605] Can Adam Albayrak and Thomas Noll. The WHILE hierarchy of program schemes is infinite. Lecture Notes in Computer Science, 1378:35– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1378/13780035.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1378/13780035.
pdf.

Amadio:1998:AGC

[606] Roberto M. Amadio and Solange Coupet-Grimal. Analysis of a guard condition in type theory. Lecture Notes in Computer Science, 1378:48– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1378/13780048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1378/13780048.
pdf.

Baldan:1998:ESS

[607] Paolo Baldan, Andrea Corradini, and Ugo Montanari. An event structure semantics for P/T contextual nets: Asymmetric event structures. Lecture Notes in Computer Science, 1378:63– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1378/13780063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1378/13780063.
pdf.

Beauquier:1998:PLT

[608] Danièle Beauquier. Pumping lemmas for timed automata. Lecture Notes in Computer Science, 1378:81– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1378/13780081.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1378/13780081.
                                        pdf.

Boreale:1998:AOP

[609] Michele Boreale, Rocco De Nicola, and Rosario Pugliese. Asynchronous observations of processes. Lecture Notes in Computer Science, 1378:95– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1378/13780095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1378/13780095.
                                        pdf.

Bottreau:1998:MSN

[610] Anne Bottreau and Yves Métivier. Minor searching, normal forms of graph relabelling: Two applications based on enumerations by graph relabelling. Lecture Notes in Computer Science, 1378: 110–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1378/13780110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1378/13780110.
                                        pdf.

Bukatin:1998:PMC

[611] Michael A. Bukatin and Svetlana Yu. Shorina. Partial metrics and cocontinuous valuations. Lecture Notes in Computer Science, 1378:125–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1378/13780125.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1378/13780125.
pdf.

Cardelli:1998:MA

[612] Luca Cardelli and Andrew D. Gordon. Mobile ambients. Lecture Notes in Computer Science, 1378:140– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1378/13780140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1378/13780140.
pdf.

Corradini:1998:RTR

[613] Andrea Corradini and Fabio Gadducci. Rational term rewriting. Lecture Notes in Computer Science, 1378: 156–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1378/13780156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1378/13780156.
pdf.

Heckmann:1998:ABI

[614] Reinhold Heckmann. The appearance of big integers in exact real arithmetic based on linear fractional transformations. Lecture Notes in Computer Science, 1378:172–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1378/13780172.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1378/13780172.
                                        pdf.

Klempien-Hinrichs:1998:NRP

[615] Renate Klempien-Hinrichs. Net refinement by pullback rewriting. Lecture Notes in Computer Science, 1378:189– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1378/13780189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1378/13780189.
                                        pdf.

Matz:1998:PTS

[616] Oliver Matz. On piecewise testable, starfree, and recognizable picture languages. Lecture Notes in Computer Science, 1378:203–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1378/13780203.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1378/13780203.pdf.

Moggi:1998:FCT

[617] Eugenio Moggi. Functor categories and two-level languages. Lecture Notes in Computer Science, 1378:211– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1378/13780211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1378/13780211.
                                        pdf.

Muscholl:1998:DPM

[618] Anca Muscholl, Doron Peled, and Zhendong Su. Deciding properties for message sequence charts. Lecture Notes in Computer Science, 1378:226– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1378/13780226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1378/13780226.
pdf.

Niemann:1998:CRL

[619] Gundula Niemann and Friedrich Otto. The Church–Rosser languages are the deterministic variants of the growing context-sensitive languages. Lecture Notes in Computer Science, 1378:243– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1378/13780243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1378/13780243.
pdf.

Porrot:1998:DRT

[620] Sylvain Porrot, Max Dauchet, Bruno Durand, and Nikolai K. Vereshchagin. Deterministic rational transducers and random sequences. Lecture Notes in Computer Science, 1378:258– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1378/13780258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1378/13780258.
pdf.

Rathke:1998:RBM

[621] Julian Rathke. Resource based models for asynchrony. Lecture Notes in Computer Science, 1378:273–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1378/13780273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1378/13780273.
                                        pdf.

Klop:1998:OTT

[622] Jan Willem Klop. Origin tracking in term rewriting. Lecture Notes in Computer Science, 1379:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790001.
                                        pdf.

Okui:1998:SCP

[623] Satoshi Okui. Simultaneous critical pairs and Church–Rosser property. Lecture Notes in Computer Science, 1379: 2–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790002.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790002.
                                        pdf.

Ohlebusch:1998:CRT

[624] Enno Ohlebusch. Church–Rosser theorems for abstract reduction modulo an equivalence relation. Lecture Notes in Computer Science, 1379:17– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790017.
pdf.

Otto:1998:AMV

[625] Friedrich Otto, Andrea Sattler-Klein, and Klaus Madlener. Automatic monoids versus monoids with finite convergent presentations. Lecture Notes in Computer Science, 1379:32– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790032.
pdf.

Levy:1998:DUS

[626] Jordi Levy. Decidable and undecidable second-order unification problems. Lecture Notes in Computer Science, 1379: 47–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790047.
pdf.

Schmidt-Schauss:1998:EPM

[627] Manfred Schmidt-Schauß and Klaus U. Schulz. On the exponent of periodicity of minimal solutions of context equations. Lecture Notes in Computer Science, 1379:61–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790061.
                                        pdf.

Jacquemard:1998:UES

[628] Florent Jacquemard, Christoph Meyer, and Christoph Weidenbach. Unification in extensions of shallow equational theories. Lecture Notes in Computer Science, 1379:76–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790076.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790076.
                                        pdf.

Guo:1998:UMP

[629] Qing Guo, Paliath Narendran, and Sandeep K. Shukla. Unification and matching in process algebras. Lecture Notes in Computer Science, 1379:91– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790091.
                                        pdf.

Schmidt:1998:USS

[630] Renate A. Schmidt. E-unification for subsystems of S4. Lecture Notes in Computer Science, 1379:106–120, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790106.
pdf.

Limet:1998:SDM

[631] Sébastien Limet and Pierre Réty. Solving disequations modulo some class of rewrite systems. Lecture Notes in Computer Science, 1379:121–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790121.
pdf.

Comon:1998:APC

[632] Hubert Comon. About proofs by consistency. Lecture Notes in Computer Science, 1379:136–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790136.
pdf.

Waldmann:1998:NTD

[633] Johannes Waldmann. Normalization of S-terms is decidable. Lecture Notes in Computer Science, 1379:138– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790138.
                                        pdf.

Genet:1998:DAS

[634] Thomas Genet. Decidable approximations of sets of descendants and sets of normal forms. Lecture Notes in Computer Science, 1379:151–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790151.
                                        pdf.

Verma:1998:ARR

[635] Rakesh M. Verma, Michael Rusinowitch, and Denis Lugiez. Algorithms and reductions for rewriting problems. Lecture Notes in Computer Science, 1379: 166–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790166.
                                        pdf.

Degtyarev:1998:DSR

[636] Anatoli Degtyarev, Yuri Gurevich, Paliath Narendran, Margus Veanes, and Andrei Voronkov. The decidability of simultaneous rigid E-unification with one variable. Lecture Notes in Computer Science, 1379:181–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790181.
pdf.

Muller:1998:OCF

[637] Martin Müller and Joachim Niehren. Ordering constraints over feature trees expressed in second-order monadic logic. Lecture Notes in Computer Science, 1379:196–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1379/13790196.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1379/13790196.pdf.

Charatonik:1998:CDS

[638] Witold Charatonik and Andreas Podelski. Co-definite set constraints. Lecture Notes in Computer Science, 1379:211– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790211.
pdf.

Arts:1998:MTU

[639] Thomas Arts and Jürgen Giesl. Modularity of termination using dependency pairs. Lecture Notes in Computer Science, 1379:226–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1379/13790226.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790226.
                                        pdf.

Marche:1998:TAC

[640] Claude Marché and Xavier Urbain. Termination of associativecommutative rewriting by dependency pairs. Lecture Notes in Computer Science, 1379:241–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1379/13790241.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1379/13790241.pdf.

Aoto:1998:TTT

[641] Takahito Aoto and Yoshihito Toyama. Termination transformation by tree lifting ordering. Lecture Notes in Computer Science, 1379:256–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790256.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790256.
                                        pdf.

Xi:1998:TAT

[642] Hongwei Xi. Towards automated termination proofs through “freezing”. Lecture Notes in Computer Science, 1379: 271–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790271.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1379/13790271.
pdf.

Danvy:1998:HOR

[643] Olivier Danvy and Kristoffer Høgsbro Rose. Higher-order rewriting and partial evaluation. Lecture Notes in Computer Science, 1379:286–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790286.
pdf.

Akama:1998:SCP

[644] Yohji Akama. SN combinators and partial combinatory algebras. Lecture Notes in Computer Science, 1379:302– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790302.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790302.
pdf.

Fuchs:1998:CSB

[645] Dirk Fuchs. Coupling saturationbased provers by exchanging positive/negative information. Lecture Notes in Computer Science, 1379:317– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1379/13790317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1379/13790317.
pdf.

Dershowitz:1998:LPD

[646] Nachum Dershowitz and Ralf Treinen. An on-line problem database. Lecture Notes in Computer Science, 1379:332– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1379/13790332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1379/13790332.
                                        pdf.

Panario:1998:ARP

[647] Daniel Panario and Alfredo Viola. Analysis of Rabin’s polynomial irreducibility test. Lecture Notes in Computer Science, 1380:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1380/13800001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1380/13800001.
                                        pdf.

Damaschke:1998:CSP

[648] Peter Damaschke. A chip search problem on binary numbers. Lecture Notes in Computer Science, 1380:11– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1380/13800011.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1380/13800011.
                                        pdf.

Feuerstein:1998:USS

[649] Esteban Feuerstein. Uniform service systems with k servers. Lecture Notes in Computer Science, 1380:23– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800023.
pdf.

Fernandez-Baca:1998:FNL

[650] David Fernández-Baca. Faster nonlinear parametric search with applications to optimization and dynamic geometry. Lecture Notes in Computer Science, 1380:33–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800033.
pdf.

Bassino:1998:SSA

[651] Frédérique Bassino, Marie-Pierre Béal, and Dominique Perrin. Super-state automata and rational trees. Lecture Notes in Computer Science, 1380:42– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800042.
pdf.

Bedon:1998:ETW

[652] Nicolas Bedon and Olivier Carton. An Eilenberg theorem for words on countable ordinals. Lecture Notes in Computer Science, 1380:53–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1380/13800053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1380/13800053.
                                        pdf.

DoLago:1998:MGF

[653] A. P. Do Lago. Maximal groups in free Burnside semigroups. Lecture Notes in Computer Science, 1380:65–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lago:1998:MGF

[654] Alair Pereira do Lago. Maximal groups in free burnside semigroups. Lecture Notes in Computer Science, 1380:65– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1380/13800065.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1380/13800065.
                                        pdf.

Pin:1998:PVI

[655] Jean-Éric Pin. Positive varieties and infinite words. Lecture Notes in Computer Science, 1380:76–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1380/13800076.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1380/13800076.
                                        pdf.

Peixoto:1998:UPA

[656] Marcos Veloso Peixoto and Laurent Fribourg. Unfolding parametric automata. Lecture Notes in Computer Science, 1380:88–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800088.
pdf.

Finkel:1998:FSW

[657] Alain Finkel and Philippe Schnoebelen. Fundamental structures in wellstructured infinite transition systems. Lecture Notes in Computer Science, 1380:102–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1380/13800102.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1380/13800102.
pdf.

Edelsbrunner:1998:SRD

[658] Herbert Edelsbrunner. Shape reconstruction with Delaunay complex. Lecture Notes in Computer Science, 1380: 119–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800119.
pdf.

Gomide:1998:BNH

[659] Anamaria Gomide and Jorge Stolfi. Bases for non-homogeneous polynomial C⃗_{k} splines on the sphere. Lecture Notes in Computer Science, 1380: 133–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800133.
pdf.

Faria:1998:SNC

[660] Luerbio Faria, Celina Miraglia Herrera de Figueiredo, and Candido Ferreira Xavier de Mendonça Neto. The splitting number of the 4-cube. Lecture Notes in Computer Science, 1380: 141–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800141.
pdf.

Abello:1998:SSP

[661] James Abello and Emden Gansner. Short and smooth polygonal paths. Lecture Notes in Computer Science, 1380:151–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1380/13800151.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1380/13800151.
pdf.

Brassard:1998:QCH

[662] Gilles Brassard, Peter Høyer, and Alain Tapp. Quantum cryptanalysis of hash and claw-free functions. Lecture Notes in Computer Science, 1380: 163–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800163.
pdf.

Bellare:1998:BVA

[663] Mihir Bellare, Juan A. Garay, and Tal Rabin. Batch verification with applications to cryptography and checking. Lecture Notes in Computer Science, 1380:170–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800170.
pdf.

Hevia:1998:STD

[664] Alejandro Hevia and Marcos Kiwi. Strength of two data encryption standard implementations under timing attacks. Lecture Notes in Computer Science, 1380:192–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1380/13800192.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1380/13800192.pdf.

Alon:1998:STG

[665] Noga Alon. Spectral techniques in graph algorithms. Lecture Notes in Computer Science, 1380:206–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1380/13800206.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1380/13800206.
                                        pdf.

Molloy:1998:CGW

[666] Michael Molloy and Bruce Reed. Colouring graphs whose chromatic number is almost their maximum degree. Lecture Notes in Computer Science, 1380:216–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1380/13800216.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1380/13800216.pdf.

Lee:1998:CCS

[667] Orlando Lee and Yoshiko Wakabayashi. Circuit covers in seriesparallel mixed graphs. Lecture Notes in Computer Science, 1380:226–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1380/13800226.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1380/13800226.
                                        pdf.

Dahlhaus:1998:LTA

[668] Elias Dahlhaus. A linear time algorithm to recognize clustered planar graphs and its parallelization. Lecture Notes in Computer Science, 1380:239– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800239.
pdf.

Jansen:1998:NCP

[669] Klaus Jansen. A new characterization for parity graphs and a coloring problem with costs. Lecture Notes in Computer Science, 1380:249– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800249.
pdf.

Gutierrez:1998:CO

[670] Marisa Gutierrez and João Meidanis. On the clique operator. Lecture Notes in Computer Science, 1380:261– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800261.
pdf.

Broder:1998:DPR

[671] Andrei Z. Broder, Alan M. Frieze, and Eli Upfal. Dynamic packet routing on arrays with bounded buffers. Lecture Notes in Computer Science, 1380:273– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1380/13800273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1380/13800273.
                                        pdf.

Roberts:1998:LMR

[672] Alan Roberts and Antonios Symvonis. On-line matching routing on trees. Lecture Notes in Computer Science, 1380: 282–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1380/13800282.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1380/13800282.
                                        pdf.

Randall:1998:AGD

[673] Dana Randall and Prasad Tetali. Analyzing Glauber dynamics by comparison of Markov chains. Lecture Notes in Computer Science, 1380:292– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1380/13800292.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1380/13800292.
                                        pdf.

Gathen:1998:CPC

[674] Joachim von zur Gathen and Igor Shparlinski. The CREW PRAM complexity of modular inversion. Lecture Notes in Computer Science, 1380:305– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800305.
pdf.

VonZurGathen:1998:CPC

[675] J. Von Zur Gathen and I. Shparlinski. The CREW PRAM complexity of modular inversion. Lecture Notes in Computer Science, 1380:305–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AufDerHeide:1998:CEP

[676] F. M. Auf Der Heide and G. T. Martinez. Communication-efficient parallel multiway and approximate minimum cut computation. Lecture Notes in Computer Science, 1380:316–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heide:1998:CEP

[677] Friedhelm Meyer auf der Heide and Gabriel Terán Martinez. Communicationefficient parallel multiway and approximate minimum cut computation. Lecture Notes in Computer Science, 1380: 316–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800316.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1380/13800316.
pdf.

Beigel:1998:GB

[678] Richard Beigel and Egemen Tanin. The geometry of browsing. Lecture Notes in Computer Science, 1380:331– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800331.
pdf.

Baeza-Yates:1998:FTD

[679] Ricardo Baeza-Yates and Gonzalo Navarro. Fast two-dimensional approximate pattern matching. Lecture Notes in Computer Science, 1380:341– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800341.
pdf.

Navarro:1998:IAP

[680] Gonzalo Navarro. Improved approximate pattern matching on hypertext. Lecture Notes in Computer Science, 1380:352–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1380/13800352.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1380/13800352.
pdf.

Gutierrez:1998:SES

[681] Claudio Gutiérrez. Solving equations in strings: On Makanin’s algorithm. Lecture Notes in Computer Science, 1380: 358–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800358.
pdf.

Sagot:1998:SAR

[682] Marie-France Sagot. Spelling approximate repeated or common motifs using a suffix tree. Lecture Notes in Computer Science, 1380:374– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1380/13800374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1380/13800374.
pdf.

Smolka:1998:CCP

[683] Gert Smolka. Concurrent constraint programming based on functional programming. Lecture Notes in Computer Science, 1381:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810001.
pdf.

Abadi:1998:BMC

[684] Martín Abadi and Andrew D. Gordon. A bisimulation method for cryptographic protocols. Lecture Notes in Computer Science, 1381:12–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810012.
pdf.

Bruynooghe:1998:PBT

[685] Maurice Bruynooghe, Michael Leuschel, and Konstantinos Sagonas. A polyvariant binding-time analysis for offline partial deduction. Lecture Notes in Computer Science, 1381:27–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810027.
pdf.

Caires:1998:VEL

[686] Luís Caires and Luís Monteiro. Verifiable and executable logic specifications of concurrent objects in L_{π}. Lecture Notes in Computer Science, 1381:42– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810042.
pdf.

Chatterjee:1998:CCT

[687] Ramkrishna Chatterjee, Barbara G. Ryder, and William A. Landi. Complexity of concrete type-inference in the presence of exceptions. Lecture Notes in Computer Science, 1381:57– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810057.
pdf.

Chin:1998:SAS

[688] Wei-Ngan Chin, Siau-Cheng Khoo, and Tat-Wee Lee. Synchronisation analysis to stop tupling. Lecture Notes in Computer Science, 1381:75– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810075.
pdf.

Fecht:1998:PDE

[689] Christian Fecht and Helmut Seidl. Propagating differences: An efficient new fixpoint algorithm for distributive constraint systems. Lecture Notes in Computer Science, 1381:90– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810090.
pdf.

Hensel:1998:RAC

[690] Ulrich Hensel, Marieke Huisman, Bart Jacobs, and Hendrik Tews. Reasoning about classes in object-oriented languages: Logical models and tools. Lecture Notes in Computer Science, 1381: 105–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1381/13810105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1381/13810105.
                                        pdf.

Honda:1998:LPT

[691] Kohei Honda, Vasco T. Vasconcelos, and Makoto Kubo. Language primitives and type discipline for structured communicationbased programming. Lecture Notes in Computer Science, 1381:122–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1381/13810122.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1381/13810122.
                                        pdf.

Jay:1998:FIS

[692] C. B. Jay and P. A. Steckler. The functional imperative: Shape! Lecture Notes in Computer Science, 1381: 139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1381/13810139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1381/13810139.
                                        pdf.

Knoop:1998:CMC

[693] Jens Knoop, Oliver Rüthing, and Bernhard Steffen. Code motion and code placement: Just synonyms? Lecture Notes in Computer Science, 1381:154– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810154.
pdf.

Leino:1998:ROT

[694] K. R. M. Leino. Recursive object types in a logic of object-oriented programs. Lecture Notes in Computer Science, 1381:170–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rustan:1998:ROT

[695] K. Rustan and M. Leino. Recursive object types in a logic of objectoriented programs. Lecture Notes in Computer Science, 1381:170–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810170.
pdf.

Maraninchi:1998:MAA

[696] Florence Maraninchi and Yann Rémond. Mode-automata: About modes and states for reactive systems. Lecture Notes in Computer Science, 1381:185– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810185.
pdf.

Remy:1998:COS

[697] Didier Rémy. From classes to objects via subtyping. Lecture Notes in Computer Science, 1381:200–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810200.
pdf.

Ross:1998:BBB

[698] John L. Ross and Mooly Sagiv. Building a bridge between pointer aliases and program dependences. Lecture Notes in Computer Science, 1381:221– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810221.
pdf.

Ruggieri:1998:CDD

[699] Salvatore Ruggieri. A complete declarative debugger of missing answers. Lecture Notes in Computer Science, 1381: 236–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1381/13810236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810236.
pdf.

Scherlis:1998:SCD

[700] William L. Scherlis. Systematic change of data representation: Program manipulations and a case study. Lecture Notes in Computer Science, 1381:252– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1381/13810252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1381/13810252.
pdf.

Thiemann:1998:GFS

[701] Peter Thiemann. A generic framework for specialization (abridged version). Lecture Notes in Computer Science, 1381:267–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1381/13810267.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1381/13810267.
pdf.

Beck:1998:EPH

[702] Kent Beck. Extreme programming: A humanistic discipline of software development. Lecture Notes in Computer Science, 1382:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820001.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1382/13820001.
                                        pdf.

Jones:1998:SMM

[703] Cliff B. Jones. Some mistakes I have made and what I have learned from them. Lecture Notes in Computer Science, 1382:7–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1382/13820007.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1382/13820007.
                                        pdf.

Allen:1998:SAD

[704] Robert Allen, Rémi Douence, and David Garlan. Specifying and analyzing dynamic software architectures. Lecture Notes in Computer Science, 1382:21–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1382/13820021.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1382/13820021.
                                        pdf.

Berregeb:1998:OPC

[705] Narjes Berregeb, Adel Bouhoula, and Michaël Rusinowitch. Observational proofs with critical contexts. Lecture Notes in Computer Science, 1382:38– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1382/13820038.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1382/13820038.
pdf.

Bradley:1998:IAM

[706] Steven Bradley, William Henderson, David Kendall, and Adrian Robson. Integrating AORTA with model-based data specification languages. Lecture Notes in Computer Science, 1382:54– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820054.
pdf.

Buessow:1998:SSC

[707] R. Buessow, R. Geisler, and M. Klar. Specifying safety-critical embedded systems with statecharts and Z: A case study. Lecture Notes in Computer Science, 1382:71–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bussow:1998:SSC

[708] Robert Büssow, Robert Geisler, and Marcus Klar. Specifying safetycritical embedded systems with statecharts and Z: A case study. Lecture Notes in Computer Science, 1382:71– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820071.
pdf.

Grieskamp:1998:SES

[709] Wolfgang Grieskamp, Maritta Heisel, and Heiko Dörr. Specifying embedded systems with statecharts and Z: An agenda for cyclic software components. Lecture Notes in Computer Science, 1382:88–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820088.
pdf.

Grosse-Rhode:1998:ATS

[710] Martin Große-Rhode. Algebra transformation systems and their composition. Lecture Notes in Computer Science, 1382:107–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1382/13820107.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1382/13820107.pdf.

Hamie:1998:NEO

[711] Ali Hamie, John Howse, and Stuart Kent. Navigation expressions in object-oriented modelling. Lecture Notes in Computer Science, 1382:123– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820123.
pdf.

Heckel:1998:CVR

[712] Reiko Heckel. Compositional verification of reactive systems specified by graph transformation. Lecture Notes in Computer Science, 1382:138– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820138.
pdf.

Kahrs:1998:RDS

[713] Stefan Kahrs and Donald Sannella. Reflections on the design of a specification language. Lecture Notes in Computer Science, 1382:154–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820154.
pdf.

Lechner:1998:CCC

[714] Ulrike Lechner. Constructs, concepts, and criteria for reuse in concurrent object-oriented languages. Lecture Notes in Computer Science, 1382:171– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820171.
pdf.

Margaria:1998:BFD

[715] Tiziana Margaria and Bernhard Steffen. Backtracking-free design planning by automatic synthesis in M ETA frame. Lecture Notes in Computer Science, 1382:188–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820188.
pdf.

Mota:1998:MCC

[716] Alexandre Mota and Augusto Sampaio. Model-checking CSP-Z. Lecture Notes in Computer Science, 1382:205– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820205.
pdf.

Padberg:1998:RBR

[717] J. Padberg, M. Gajewsky, and C. Ermel. Rule-based refinement of high-level nets preserving safety properties. Lecture Notes in Computer Science, 1382:221–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820221.
pdf.

Reed:1998:AFA

[718] J. N. Reed, D. M. Jackson, B. Deianov, and G. M. Reed. Automated formal analysis of networks: FDR models of arbitrary topologies and flowcontrol mechanisms. Lecture Notes in Computer Science, 1382:239–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820239.
pdf.

Nielson:1998:BAS

[719] Hanne Riis Nielson, Torben Amtoft, and Flemming Nielson. Behaviour analysis and safety conditions: A case study in CML. Lecture Notes in Computer Science, 1382:255–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820255.
pdf.

Sandholm:1998:DSC

[720] Anders Sandholm and Michael I. Schwartzbach. Distributed safety controllers for Web services. Lecture Notes in Computer Science, 1382:270– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820270.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1382/13820270.
                                        pdf.

Scholz:1998:RCS

[721] Peter Scholz. A refinement calculus for statecharts. Lecture Notes in Computer Science, 1382:285–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1382/13820285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1382/13820285.
                                        pdf.

Sucrow:1998:RFS

[722] Bettina Eva Sucrow. Refining formal specifications of human computer interaction by graph rewrite rules. Lecture Notes in Computer Science, 1382: 302–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1382/13820302.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1382/13820302.
                                        pdf.

Behnke:1998:RSC

[723] Ralf Behnke, Rudolf Berghammer, Erich Meyer, and Peter Schneider. RELVIEW — A system for calculating with relations and relational programming. Lecture Notes in Computer Science, 1382:318–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1382/13820318.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1382/13820318.
pdf.

Dubois:1998:AFL

[724] Eric Dubois. ALBERT: A formal language and its supporting tools for requirements engineering. Lecture Notes in Computer Science, 1382:322– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820322.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820322.
pdf.

Tapken:1998:MOP

[725] Josef Tapken. MOBY/PLC — A design tool for hierarchical realtime automata. Lecture Notes in Computer Science, 1382:326–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1382/13820326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1382/13820326.
pdf.

Brouwer:1998:MFA

[726] K. Brouwer, W. Gellerich, and E. Ploedereder. Myths and facts about the efficient implementation of finite automata and lexical analysis. Lecture Notes in Computer Science, 1383:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnstone:1998:GRD

[727] A. Johnstone and E. Scott. Generalised recursive descent parsing and follow-determinism. Lecture Notes in Computer Science, 1383:16–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyland:1998:ADN

[728] J. T. Boyland. Analyzing direct nonlocal dependencies in attribute grammars. Lecture Notes in Computer Science, 1383:31–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizushima:1998:SAS

[729] K. Mizushima and T. Katayama. Storage allocation strategies for recursive attribute evaluators. Lecture Notes in Computer Science, 1383:50–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knoop:1998:BBG

[730] J. Knoop, D. Koschuetzki, and B. Steffen. Basic-block graphs: living dinosaurs? Lecture Notes in Computer Science, 1383:65–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1998:AL

[731] F. Martin, M. Alt, R. Wilhelm, and C. Ferdinand. Analysis of loops. Lecture Notes in Computer Science, 1383: 80–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malacaria:1998:NAC

[732] P. Malacaria and C. Hankin. A new approach to control flow analysis. Lecture Notes in Computer Science, 1383: 95–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielson:1998:FLC

[733] H. R. Nielson and F. Nielson. Flow logics for constraint based analysis. Lecture Notes in Computer Science, 1383: 109–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lapkowski:1998:ESN

[734] C. Lapkowski and L. J. Hendren. Extended SSA numbering: introducing SSA properties to languages with multi-level pointers. Lecture Notes in Computer Science, 1383:128–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennedy:1998:SRS

[735] R. Kennedy, F. Chow, P. Dahl, and S.-M. Liu. Strength reduction via SS-APRE. Lecture Notes in Computer Science, 1383:144–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghiya:1998:DPC

[736] R. Ghiya, L. J. Hendren, and Y. Zhu. Detecting parallelism in C programs with recursive data structures. Lecture Notes in Computer Science, 1383:159– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:1998:LRS

[737] K. D. Cooper and L. T. Simpson. Live range splitting in a graph coloring register allocator. Lecture Notes in Computer Science, 1383:174–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burnett:1998:ILA

[738] M. M. Burnett. Invited lecture (abstract): Challenges and opportunities visual programming languages bring to programming language research. Lecture Notes in Computer Science, 1383: 188–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maierhofer:1998:LSA

[739] M. Maierhofer and M. A. Ertl. Local stack allocation. Lecture Notes in Computer Science, 1383:189–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lelait:1998:NFA

[740] S. Lelait, G. R. Gao, and C. Eisenbeis. A new fast algorithm for optimal register allocation in modulo scheduled loops. Lecture Notes in Computer Science, 1383:204–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1998:CMF

[741] R. Gupta. A code motion framework for global instruction scheduling. Lecture Notes in Computer Science, 1383:219–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stuempel:1998:VCT

[742] E. Stuempel, M. Thies, and U. Kastens. VLIW compilation techniques for superscalar architectures. Lecture Notes in Computer Science, 1383:234– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelbrecht:1998:ITS

[743] R. L. Engelbrecht and D. Kourie. Issues in translating Smalltalk to Java. Lecture Notes in Computer Science, 1383: 249–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steindl:1998:ISO

[744] C. Steindl. Intermodular slicing of object-oriented programs. Lecture Notes in Computer Science, 1383:264– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pettersson:1998:PDP

[745] M. Pettersson. Portable debugging and profiling. Lecture Notes in Computer Science, 1383:279–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kastens:1998:ES

[746] U. Kastens, P. Pfahler, and M. Jung. The eli system. Lecture Notes in Computer Science, 1383:294–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuiper:1998:LGI

[747] M. Kuiper and J. Saraiva. Lrc — a generator for incremental languageoriented tools. Lecture Notes in Computer Science, 1383:298–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leino:1998:ESC

[748] K. R. M. Leino and G. Nelson. An extended static checker for Modula-3. Lecture Notes in Computer Science, 1383:302–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braun:1998:CCF

[749] V. Braun, J. Knoop, and D. Koschuetzki. Cool: a control-flow generator for system analysis. Lecture Notes in Computer Science, 1383:306–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryant:1998:FVP

[750] Randal E. Bryant. Formal verification of pipelined processors. Lecture Notes in Computer Science, 1384:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840001.
pdf.

Liu:1998:FLE

[751] Xinxin Liu, C. R. Ramakrishnan, and Scott A. Smolka. Fully local and efficient evaluation of alternating fixed points (extended abstract). Lecture Notes in Computer Science, 1384:5– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840005.
pdf.

Laster:1998:MMC

[752] Karen Laster and Orna Grumberg. Modular model checking of software. Lecture Notes in Computer Science, 1384:20–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1384/13840020.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1384/13840020.
pdf.

Huhn:1998:VBL

[753] Michaela Huhn, Peter Niebert, and Frank Wallner. Verification based on local states. Lecture Notes in Computer Science, 1384:36–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840036.
pdf.

Ajami:1998:ESL

[754] K. Ajami, S. Haddad, and J-M. Ilié. Exploiting symmetry in linear time temporal logic model checking: One step beyond. Lecture Notes in Computer Science, 1384:52– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840052.
pdf.

Garavel:1998:OCO

[755] Hubert Garavel. OPEN/CAESAR: An open software architecture for verification, simulation, and testing. Lecture Notes in Computer Science, 1384:68– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1384/13840068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1384/13840068.
                                        pdf.

Stevens:1998:PMC

[756] Perdita Stevens and Colin Stirling. Practical model checking using games. Lecture Notes in Computer Science, 1384:85–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1384/13840085.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1384/13840085.
                                        pdf.

Grahlmann:1998:CFA

[757] Bernd Grahlmann. Combining finite automata, parallel programs and SDL using Petri nets. Lecture Notes in Computer Science, 1384:102–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1384/13840102.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1384/13840102.
                                        pdf.

Ben-Abdallah:1998:MES

[758] Hanêne Ben-Abdallah and Stefan Leue. MESA: Support for scenario-based design of concurrent systems. Lecture Notes in Computer Science, 1384:118– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840118.
pdf.

Velev:1998:EMM

[759] Miroslav N. Velev and Randal E. Bryant. Efficient modeling of memory arrays in symbolic ternary simulation. Lecture Notes in Computer Science, 1384:136–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1384/13840136.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1384/13840136.
pdf.

Pnueli:1998:TV

[760] A. Pnueli, M. Siegel, and E. Singerman. Translation validation. Lecture Notes in Computer Science, 1384:151– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840151.
pdf.

Sprenger:1998:VMC

[761] Christoph Sprenger. A verified model checker for the modal μcalculus in Coq. Lecture Notes in Computer Science, 1384:167–183, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840167.
pdf.

Aiken:1998:DRR

[762] Alexander Aiken, Manuel Fähndrich, and Zhendong Su. Detecting races in relay ladder logic programs. Lecture Notes in Computer Science, 1384:184– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840184.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840184.
pdf.

Lind-Nielsen:1998:VLS

[763] Jørn Lind-Nielsen, Henrik Reif Andersen, Gerd Behrmann, Henrik Hulgaard, Kåre Kristoffersen, and Kim G. Larsen. Verification of large state/ event systems using compositionality and dependency analysis. Lecture Notes in Computer Science, 1384:201– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840201.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1384/13840201.
pdf.

Brockmeyer:1998:TNV

[764] Udo Brockmeyer and Gunnar Wittich. Tamagotchis need not die — verification of Statemate designs. Lecture Notes in Computer Science, 1384:217– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840217.
pdf.

Cartel:1998:MVS

[765] T. Cartel. Modeling and verification of sC++ applications. Lecture Notes in Computer Science, 1384:232–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CATTEL:1998:MVS

[766] Thierry CATTEL. Modeling and verification of sC++ applications. Lecture Notes in Computer Science, 1384: 232–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840232.
pdf.

Sherman:1998:FAS

[767] David James Sherman and Nicolas Magnier. Factotum: Automatic and systematic sharing support for systems analyzers. Lecture Notes in Computer Science, 1384:249–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840249.
pdf.

Aceto:1998:MCR

[768] Luca Aceto, Augusto Burgueño, and Kim G. Larsen. Model checking via reachability testing for timed automata. Lecture Notes in Computer Science, 1384:263–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840263.
pdf.

Lindahl:1998:FDA

[769] Magnus Lindahl, Paul Pettersson, and Wang Yi. Formal design and analysis of a gear controller. Lecture Notes in Computer Science, 1384:281– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840281.
pdf.

Abdulla:1998:VNT

[770] Parosh Aziz Abdulla and Bengt Jonsson. Verifying networks of timed processes (extended abstract). Lecture Notes in Computer Science, 1384:298– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840298.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840298.
pdf.

Daws:1998:MCR

[771] Conrado Daws and Stavros Tripakis. Model checking of real-time reachability properties using abstractions. Lecture Notes in Computer Science, 1384: 313–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840313.
pdf.

Alur:1998:SET

[772] Rajeev Alur, Thomas A. Henzinger, and Sriram K. Rajamani. Symbolic exploration of transition hierarchies. Lecture Notes in Computer Science, 1384: 330–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840330.
pdf.

Kurshan:1998:SPO

[773] R. Kurshan, V. Levin, M. Minea, D. Peled, and H. Yenigün. Static partial order reduction. Lecture Notes in Computer Science, 1384:345–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1384/13840345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1384/13840345.
                                        pdf.

Kutshah:1998:SPO

[774] R. Kutshah, V. Levin, M. Minea, and D. Peled. Static partial order reduction. Lecture Notes in Computer Science, 1384:345–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charatonik:1998:SBA

[775] Witold Charatonik and Andreas Podelski. Set-based analysis of reactive infinite-state systems. Lecture Notes in Computer Science, 1384:358–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1384/13840358.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1384/13840358.
                                        pdf.

Bjoerner:1998:DFN

[776] N. S. Bjoerner and M. C. Pichora. Deciding fixed and non-fixed size bitvectors. Lecture Notes in Computer Science, 1384:376–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjorner:1998:DFN

[777] Nikolaj S. Bjørner and Mark C. Pichora. Deciding fixed and nonfixed size bit-vectors. Lecture Notes in Computer Science, 1384:376–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840376.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840376.
pdf.

Ruys:1998:ELP

[778] Theo C. Ruys and Ed Brinksma. Experience with literate programming in the modelling and validation of systems. Lecture Notes in Computer Science, 1384:393–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840393.
pdf.

Jensen:1998:PBP

[779] Henrik E. Jensen and Nancy A. Lynch. A proof of Burns Nprocess mutual exclusion algorithm using abstraction. Lecture Notes in Computer Science, 1384:409–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1384/13840409.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1384/13840409.
pdf.

Gribomont:1998:AVS

[780] E. Pascal Gribomont and Guy Zenner. Automated verification of Szymanski’s algorithm. Lecture Notes in Computer Science, 1384:424–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1384/13840424.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1384/13840424.
                                        pdf.

Regensburger:1998:FVS

[781] Franz Regensburger and Aenne Barnard. Formal verification of SDL systems at the Siemens Mobile Phone Department. Lecture Notes in Computer Science, 1384:439–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1384/13840439.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1384/13840439.pdf.

Lardenois:1998:WIW

[782] Alain Lardenois. The Web impact: A white paper. Lecture Notes in Computer Science, 1385:8– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1385/13850001.
                                        pdf.

Zave:1998:CCH

[783] Pamela Zave. ’calls considered harmful’ and other observations: A tutorial on telephony. Lecture Notes in Computer Science, 1385:8–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850008.
pdf.

Decina:1998:CTC

[784] Maurizio Decina. Convergence of telecommunications and computing on networking models for integrated services and applications. Lecture Notes in Computer Science, 1385:28– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850028.
pdf.

Handschub:1998:NAN

[785] S. Handschub, U. Lechner, D.-M. Lincke, and B. Schmid. The net academy — A new concept for online publishing and knowledge management. Lecture Notes in Computer Science, 1385:29–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Handschuh:1998:NNC

[786] Siegfried Handschuh, Ulrike Lechner, David-Michael Lincke, Beat Schmid, Petra Schubert, Dorian Selz, and Katarina Stanoevska-Slabeva. The NetAcademy — A new concept for online publishing and knowledge management. Lecture Notes in Computer Science, 1385:29–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850029.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1385/13850029.
                                        pdf.

Hess:1998:DCL

[787] Robert Hess, Dagmar Geske, Sascha Kuemmel, and Henrik Thuermer. Distributed compression of live video — an application for active networks. Lecture Notes in Computer Science, 1385: 44–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850044.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1385/13850044.
                                        pdf.

Kakizaki:1998:ISG

[788] Ken’ichi Kakizaki. Incremental scene graph distribution method for distributed virtual environments. Lecture Notes in Computer Science, 1385:61– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1385/13850061.
                                        pdf.

Hergenhan:1998:VAS

[789] André Hergenhan, Christoph Weiler, Karlheinz Weiß, and Wolfgang Rosenstiel. Value-added services in industrial automation. Lecture Notes in Computer Science, 1385:75–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850075.
pdf.

Keck:1998:RPP

[790] Dirk O. Keck. Requirements and a proposal for the prevention of a class of service interactions in intelligent networks. Lecture Notes in Computer Science, 1385:90–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850090.
pdf.

Do:1998:IPT

[791] Hyunsook Do, Seongyong Bae, and Sangki Kim. INAP protocol test suite verification method using the IUT simulator for AIN system conformance testing. Lecture Notes in Computer Science, 1385:106–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850106.
pdf.

Florschuetz:1998:TPI

[792] P. Florschuetz. Testing IN protocol implementation. Lecture Notes in Computer Science, 1385:117–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Florschutz:1998:TPI

[793] P. Florschütz. Testing IN protocol implementation. Lecture Notes in Computer Science, 1385:117–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1385/13850117.
                                        pdf.

Albayrak:1998:ABD

[794] Sahin Albayrak and Jens Meinköhn. Agent-based data services in future IN-platforms. Lecture Notes in Computer Science, 1385:128–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1385/13850128.
                                        pdf.

Oezcan:1998:VEF

[795] M. B. Oezcan, P. W. Parry, I. C. Morrey, and J. I. Siddiqi. Visualisation of executable formal specifications for user validation. Lecture Notes in Computer Science, 1385:142–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ozcan:1998:VEF

[796] M. B. Özcan, P. W. Parry, I. C. Morrey, and J. I. Siddiqi. Visualisation of executable formal specifications for user validation. Lecture Notes in Computer Science, 1385:142– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850142.
pdf.

Lusini:1998:DEV

[797] M. Lusini and E. Vicario. Design and evaluation of a visual formalism for real time logics. Lecture Notes in Computer Science, 1385:158– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850158.
pdf.

Capellmann:1998:VBI

[798] Carla Capellmann, Søren Christensen, and Uwe Herzog. Visualising the behaviour of intelligent networks. Lecture Notes in Computer Science, 1385: 174–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850174.
pdf.

Margaria:1998:FMC

[799] Tiziana Margaria and Volker Braun. Formal methods and customized visualization: A fruitful symbiosis. Lecture Notes in Computer Science, 1385: 190–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1385/13850190.
                                        pdf.

Schloer:1998:UVF

[800] R. Schloer, B. Josko, and D. Werth. Using a visual formalism for design verification in industrial environments. Lecture Notes in Computer Science, 1385: 208–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlor:1998:UVF

[801] Rainer Schlör, Bernhard Josko, and Dieter Werth. Using a visual formalism for design verification in industrial environments. Lecture Notes in Computer Science, 1385:208–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1385/13850208.
                                        pdf.

Braun:1998:AEL

[802] Volker Braun, Tiziana Margaria, Bernhard Steffen, and Haiseung Yoo. Automatic error location for IN service definition. Lecture Notes in Computer Science, 1385:222–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850222.
pdf.

Kang:1998:GSD

[803] Jeomja Kang, Jeonghun Choi, and Sehyeong Cho. The generation of service database schema through service creation environment. Lecture Notes in Computer Science, 1385:238– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850238.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850238.
pdf.

Orozco-Barbosa:1998:SIM

[804] L. Orozco-Barbosa, D. Makrakis, and N. D. Georganas. A study of intelligent multimedia services over PSTNs and the Internet. Lecture Notes in Computer Science, 1385:250–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850250.
pdf.

Duque-Anton:1998:OSA

[805] Manuel Duque-Antón, Ralf Günther, Raschid Karabek, Thomas Meuser, and Josef Wasel. Open switching for ATM networks. Lecture Notes in Computer Science, 1385:265–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850265.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1385/13850265.
                                        pdf.

Lim:1998:AIA

[806] Hee Jin Lim and Go Bong Choi. Application of INAP to AIN intelligent peripheral. Lecture Notes in Computer Science, 1385:278–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850278.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1385/13850278.
                                        pdf.

Lira:1998:AIA

[807] H. J. Lira and G. B. Choi. Application of INAP to AIN intelligent peripheral. Lecture Notes in Computer Science, 1385:278–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biacchi:1998:ESM

[808] S. Biacchi, G. Ferrari, and R. Gobbi. IN evolution to support mobility in DECT access networks. Lecture Notes in Computer Science, 1385:291–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1385/13850291.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1385/13850291.
pdf.

Blaiotta:1998:MBS

[809] D. Blaiotta, L. Faglia, M. Varisco, and L. Vezzoli. A mobile broadband service switching point: A new network element for multimedia and mobile services. Lecture Notes in Computer Science, 1385:306– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1385/13850306.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1385/13850306.
pdf.

Balluchi:1998:HCA

[810] A. Balluchi, M. Di Benedetto, C. Pinello, and C. Rossi. Hybrid control for automotive engine management: The cut-off case. Lecture Notes in Computer Science, 1386:13– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beydoun:1998:HCA

[811] A. Beydoun, L. Y. Wang, J. Sun, and S. Sivashankar. Hybrid control of automotive powertrain systems: A case study. Lecture Notes in Computer Science, 1386:33–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bornot:1998:CHS

[812] S. Bornot and J. Sifakis. On the composition of hybrid systems. Lecture Notes in Computer Science, 1386:49– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bushnell:1998:EBC

[813] L. Bushnell, O. Beldiman, and G. Walsh. An equivalence between a control network and a switched hybrid system. Lecture Notes in Computer Science, 1386:64–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlson:1998:HCI

[814] B. Carlson and V. Gupta. Hybrid cc with interval constraints. Lecture Notes in Computer Science, 1386:80– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dang:1998:RAF

[815] T. Dang and O. Maler. Reachability analysis via face lifting. Lecture Notes in Computer Science, 1386:96– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fehnker:1998:ACR

[816] A. Fehnker. Automotive control revisited: Linear inequalities as approximation of reachable sets. Lecture Notes in Computer Science, 1386:110–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farreira:1998:SCB

[817] E. D. Farreira and B. H. Krogh. Switching controllers based on neural-network estimates of stability regions and controller performance. Lecture Notes in Computer Science, 1386:126–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Friesen:1998:LSC

[818] V. Friesen. A logic for the specification of continuous systems. Lecture Notes in Computer Science, 1386:143–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenstreet:1998:IP

[819] M. R. Greenstreet and I. Mitchell. Integrating projections. Lecture Notes in Computer Science, 1386:159–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1998:LSC

[820] K. X. He and M. D. Lemmon. Lyapunov stability of continuous-valued systems under the supervision of discrete-event transition systems. Lecture Notes in Computer Science, 1386: 175–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1998:RVH

[821] T. A. Henzinger and V. Rusu. Reachability verification for hybrid automata. Lecture Notes in Computer Science, 1386:190–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lafferriere:1998:SSB

[822] G. Lafferriere, G. J. Pappas, and S. Sastry. Subanalytic stratifications and bisimulations. Lecture Notes in Computer Science, 1386:205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehrenfeld:1998:IDS

[823] G. Lehrenfeld, R. Naumann, R. Rasche, and C. Rust. Integrated design and simulation of hybrid systems. Lecture Notes in Computer Science, 1386:221– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lemch:1998:HHS

[824] E. S. Lemch and P. E. Caines. Hierarchical hybrid systems: Partition deformations and applications to the acrobot system. Lecture Notes in Computer Science, 1386:237–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Livadas:1998:FVS

[825] C. Livadas and N. A. Lynch. Formal verfication of safety-critical hybrid systems. Lecture Notes in Computer Science, 1386:253–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lygeros:1998:SVM

[826] J. Lygeros and N. A. Lynch. Strings of vehicles: Modeling and safety conditions. Lecture Notes in Computer Science, 1386:273–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lygeros:1998:AVC

[827] J. Lygeros, G. J. Pappas, and S. Sastry. An approach to the verification of the Center-TRACON automation system. Lecture Notes in Computer Science, 1386:289–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manna:1998:DVH

[828] Z. Manna and H. B. Sipma. Deductive verification of hybrid systems using STeP. Lecture Notes in Computer Science, 1386:305–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matveev:1998:RDD

[829] A. S. Matveev and A. V. Savkin. Reduction and decomposition of differential automata: Theory and applications. Lecture Notes in Computer Science, 1386:319–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1998:OGS

[830] B. M. Miller. Optimization of generalized solutions on nonlinear hybrid (discrete-continuous) systems. Lecture Notes in Computer Science, 1386:334– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neller:1998:IBO

[831] T. W. Neller. Information-based optimization approaches to dynamical system safety verification. Lecture Notes in Computer Science, 1386:346–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomlin:1998:SCN

[832] C. Tomlin, J. Lygeros, and S. Sastry. Synthesizing controllers for nonlinear hybrid systems. Lecture Notes in Computer Science, 1386:360–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanSchuppen:1998:SCC

[833] J. H. Van Schuppen. A sufficient condition for controllability of a class of hybrid systems. Lecture Notes in Computer Science, 1386:374–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1998:HRE

[834] X. Li, T. Zheng, J. Hou, and J. Zhao. Hybrid regular expressions. Lecture Notes in Computer Science, 1386:384– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zefran:1998:SSC

[835] M. Zefran and J. W. Burdick. Stabilization of systems with changing dynamics. Lecture Notes in Computer Science, 1386:400–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsoi:1998:RNN

[836] Ah Chung Tsoi. Recurrent neural network architectures: An overview. Lecture Notes in Computer Science, 1387: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1387/13870001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1387/13870001.
pdf.

Tsoi:1998:GBL

[837] Ah Chung Tsoi. Gradient based learning methods. Lecture Notes in Computer Science, 1387:27–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1387/13870027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1387/13870027.
pdf.

Wan:1998:DMD

[838] Eric A. Wan and Françoise Beaufays. Diagrammatic methods for deriving and relating temporal neural network algorithms. Lecture Notes in Computer Science, 1387:63–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1387/13870063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1387/13870063.
pdf.

Frasconi:1998:ILS

[839] Paolo Frasconi. An introduction to learning structured information. Lecture Notes in Computer Science, 1387: 99–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1387/13870099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1387/13870099.
pdf.

Sperduti:1998:NNP

[840] Alessandro Sperduti. Neural networks for processing data structures. Lecture Notes in Computer Science, 1387:121– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1387/13870121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1387/13870121.
                                        pdf.

Gori:1998:LPT

[841] Marco Gori. The loading problem: Topics in complexity. Lecture Notes in Computer Science, 1387:145–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1387/13870145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1387/13870145.
                                        pdf.

Ghahramani:1998:LDB

[842] Zoubin Ghahramani. Learning dynamic Bayesian networks. Lecture Notes in Computer Science, 1387:168– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1387/13870168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1387/13870168.
                                        pdf.

Baldi:1998:PMN

[843] Pierre Baldi. Probabilistic models of neuronal spike trains. Lecture Notes in Computer Science, 1387:198– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1387/13870198.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1387/13870198.
pdf.

Parra:1998:TMB

[844] Lucas C. Parra. Temporal models in blind source separation. Lecture Notes in Computer Science, 1387:229– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1387/13870229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1387/13870229.
pdf.

Maggini:1998:RNN

[845] Marco Maggini. Recursive neural networks and automata. Lecture Notes in Computer Science, 1387:248–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1387/13870248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1387/13870248.
pdf.

Sun:1998:NNP

[846] G. Z. Sun, C. L. Giles, and H. H. Chen. The neural network pushdown automaton: Architecture, dynamics and training. Lecture Notes in Computer Science, 1387:296–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1387/13870296.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1387/13870296.
pdf.

Siegelmann:1998:NDS

[847] Hava T. Siegelmann. Neural dynamics with stochasticity. Lecture Notes in Computer Science, 1387:346– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1387/13870346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1387/13870346.
                                        pdf.

Mozer:1998:PST

[848] Michael C. Mozer and Debra Miller. Parsing the stream of time: The value of event-based segmentation in a complex real-world control problem. Lecture Notes in Computer Science, 1387: 370–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1387/13870370.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1387/13870370.
                                        pdf.

Bourlard:1998:HHA

[849] Hervé Bourlard and Nelson Morgan. Hybrid HMM/ANN systems for speech recognition: Overview and new research directions. Lecture Notes in Computer Science, 1387:389–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1387/13870389.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1387/13870389.
                                        pdf.

Gallinari:1998:PMS

[850] P. Gallinari. Predictive models for sequence modelling, application to speech and character recognition. Lecture Notes in Computer Science, 1387: 418–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1387/13870418.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1387/13870418.
pdf.

Ghare:1998:DRP

[851] G. Ghare and S.-Y. Lee. Dynamic reconfiguration of a PMMLA for highthroughput applications. Lecture Notes in Computer Science, 1388:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fornaciari:1998:VFS

[852] W. Fornaciari and V. Piuri. Virtual FPGAs: Some steps behind the physical barriers. Lecture Notes in Computer Science, 1388:7–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baglietto:1998:PAM

[853] P. Baglietto, M. Maresca, and M. Migliardi. A parallel algorithm for minimum cost path computation on polymorphic processor array. Lecture Notes in Computer Science, 1388:13– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walrath:1998:PMA

[854] J. Walrath and R. Vemuri. A performance modeling and analysis environment for reconfigurable computers. Lecture Notes in Computer Science, 1388:19–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brebner:1998:RRR

[855] G. Brebner and A. Donlin. Runtime reconfigurable routing. Lecture Notes in Computer Science, 1388:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ouaiss:1998:IPS

[856] I. Ouaiss, S. Govindarajan, V. Srinivasan, and M. Kaul. An integrated partitioning and synthesis system for dynamically reconfigurable Multi-FPGA architectures. Lecture Notes in Computer Science, 1388:31–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spillane:1998:TPP

[857] J. Spillane and H. Owen. Temporal partioning for partially-reconfigurable-field-programmable gate. Lecture Notes in Computer Science, 1388:37– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:1998:JDR

[858] D. Davis, M. Bart, T. Bennett, and S. Edwards. A Java development and runtime environment for reconfigurable computing. Lecture Notes in Computer Science, 1388:43–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rath:1998:SRS

[859] K. Rath and J. Li. Synthesizing reconfigurable sequential machines using tabular models. Lecture Notes in Computer Science, 1388:49–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abnous:1998:ELP

[860] A. Abnous, K. Seno, Y. Ichikawa, and M. Wan. Evaluation of a low-power reconfigurable DSP architecture. Lecture Notes in Computer Science, 1388: 55–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirschbaum:1998:RHM

[861] A. Kirschbaum, J. Becker, and M. Glesner. A reconfigurable hardwaremonitor for communication analysis in distributed real-time systems. Lecture Notes in Computer Science, 1388:61– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartenstein:1998:RCP

[862] R. W. Hartenstein, M. Herz, T. Hoffmann, and U. Nageldinger. On reconfigurable co-processing units. Lecture Notes in Computer Science, 1388:67– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scalera:1998:MBA

[863] S. M. Scalera, J. J. Murray, and S. Lease. A mathematical benefit analysis of context switching reconfigurable computing. Lecture Notes in Computer Science, 1388:73–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pudipeddi:1998:CCA

[864] B. Pudipeddi, A. L. Abbot, and P. M. Athanas. A configurable computing approach towards real-time target tracking. Lecture Notes in Computer Science, 1388:79–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shibata:1998:HVH

[865] Y. Shibata, H. Miyazaki, X. Ling, and H. Amano. HOSMII: A virtual hardware integrated with DRAM. Lecture Notes in Computer Science, 1388:85– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beuchat:1998:HRN

[866] J. L. Beuchat, J. O. Haenni, and E. Sanchez. Hardware reconfigurable neural networks. Lecture Notes in Computer Science, 1388:91–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steckel:1998:SRM

[867] C. Steckel, M. Middendorf, H. El-Gindy, and H. Schmeck. A simulator for the reconfigurable mesh architecture. Lecture Notes in Computer Science, 1388:99–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolla:1998:PPA

[868] R. Kolla and O. Springauf. PACE: Processor architectures for circuit emulation. Lecture Notes in Computer Science, 1388:105–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kale:1998:MMI

[869] L. V. Kale, M. Bhandarkar, R. Brunner, and J. Yelon. Multiparadigm, multilingual interoperability: Experience with converse. Lecture Notes in Computer Science, 1388:111–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1998:ECR

[870] C. C. Lim, Y. H. Low, W. Cai, and W. J. Hsu. An empirical comparison of runtime systems for conservative parallel simulation. Lecture Notes in Computer Science, 1388:123–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruehl:1998:SOM

[871] T. Ruehl and H. E. Bal. Synchronizing operations on multiple objects. Lecture Notes in Computer Science, 1388:135–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nibhanupudi:1998:RSV

[872] M. V. Nibhanupudi and B. K. Szymanski. Runtime support for virtual BSP computer. Lecture Notes in Computer Science, 1388:147–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petri:1998:MRT

[873] S. Petri, M. Bolz, and H. Langendoerfer. Migration and rollback transparency for arbitrary distributed applications in workstation clusters. Lecture Notes in Computer Science, 1388:

159–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tangmunarunkit:1998:NAD

[874] H. Tangmunarunkit and P. Steenkiste. Network-aware distributed computing: A case study. Lecture Notes in Computer Science, 1388:171–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikolopoulos:1998:ERT

[875] D. S. Nikolopoulos, E. D. Polychronopoulos, and T. S. Papatheodorou. Efficient runtime thread management for the nano-threads programming model. Lecture Notes in Computer Science, 1388:183–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolte:1998:TBA

[876] J. Nolte. A topology based approach to coordinated multicast operations. Lecture Notes in Computer Science, 1388:195–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Villacanas:1998:CEP

[877] J. L. Fernandez-Villacañas, J. M. Fatah, and S. Amin. Computing with evolving proteins. Lecture Notes in Computer Science, 1388:207–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ochi:1998:PEA

[878] L. S. Ochi, D. S. Vianna, L. M. A. Drummond, and A. O. Victor. A parallel evolutionary algorithm for the vehicle routing problem with heterogeneous fleet. Lecture Notes in Computer Science, 1388:216–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leopold:1998:LOP

[879] C. Leopold. Locality optimization for program instances. Lecture Notes in Computer Science, 1388:225–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jenq:1998:ANN

[880] J.-F. Jenq and W. N. Li. Artificial neural networks on reconfigurable meshes. Lecture Notes in Computer Science, 1388:234–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szuba:1998:MQR

[881] T. Szuba. A molecular quasi-random model of computations applied to evaluate collective intelligence. Lecture Notes in Computer Science, 1388:243– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1998:RSO

[882] X. He, T. Hintz, and U. Szewcow. Replicated shared object model for edge detection with spiral architecture. Lecture Notes in Computer Science, 1388:252–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seredynski:1998:STP

[883] F. Seredynski. Scheduling tasks of a parallel program in two-processor systems with use of cellular automata. Lecture Notes in Computer Science, 1388:261–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elrad:1998:EPE

[884] T. Elrad and J. Lin. Evolving processes and evolution schedulers for concurrent scheduling controls and parallel evolutionary computation. Lecture Notes in Computer Science, 1388:270–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nossal:1998:EAM

[885] R. Nossal. An evolutionary approach to multiprocessor scheduling of dependent tasks. Lecture Notes in Computer Science, 1388:279–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salleh:1998:MSU

[886] S. Salleh and A. Y. Zomaya. Multiprocessor scheduling using mean-field annealing. Lecture Notes in Computer Science, 1388:288–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rogers:1998:UBC

[887] R. O. Rogers and D. B. Skillicorn. Using the BSP cost model to optimise parallel neural network training. Lecture Notes in Computer Science, 1388:297–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talbi:1998:FTP

[888] E.-G. Talbi, J.-M. Geib, Z. Hafidi, and D. Kebbal. A fault-tolerant parallel heuristic for assignment problems. Lecture Notes in Computer Science, 1388:306–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gil:1998:MHC

[889] C. Gil, J. Ortega, A. F. Diaz, and M. G. Montoya. Meta-heuristics for circuit partitioning in parallel test generation. Lecture Notes in Computer Science, 1388:315–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krizanc:1998:SRP

[890] D. Krizanc. A survey of randomness and parallelism in comparison problems. Lecture Notes in Computer Science, 1388:324–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1998:CCH

[891] D. Hsu, L. E. Kavraki, J. C. Latombe, and R. Motwani. Capturing the connectivity of high-dimensional geometric spaces by parallelizable random sampling techniques. Lecture Notes in Computer Science, 1388:330–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1998:OBT

[892] A. Cohen, Y. Rabinovich, A. Schuster, and H. Shachnai. Optimal bounds on tail probabilities — A simplified approach. Lecture Notes in Computer Science, 1388:341–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raman:1998:RST

[893] R. Raman. Random sampling techniques in parallel computation. Lecture Notes in Computer Science, 1388: 351–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muthukrishnan:1998:RPS

[894] S. Muthukrishnan. Randomization in parallel stringology. Lecture Notes in Computer Science, 1388:361–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varman:1998:RPP

[895] P. J. Varman. Randomized parallel prefetching and buffer management. Lecture Notes in Computer Science, 1388:363–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravikumar:1998:PAF

[896] B. Ravikumar. Parallel algorithms for finite automata problems. Lecture Notes in Computer Science, 1388:373– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wei:1998:RRP

[897] D. S. L. Wei. Randomized routing and PRAM emulation on parallel machines. Lecture Notes in Computer Science, 1388:374–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suel:1998:RDS

[898] T. Suel. On randomized and deterministic schemes for routing and sorting on fixed-connection networks. Lecture Notes in Computer Science, 1388: 384–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1998:HPL

[899] J. Dongarra and J. Wasniewski. High performance linear algebra package LAPACK90. Lecture Notes in Computer Science, 1388:387–391, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dessmark:1998:URP

[900] A. Dessmark, C. Dorgerloh, A. Lingas, and J. Wirtgen. Ultrafast randomized parallel construction and approximation algorithms for spanning forests in dense graphs. Lecture Notes in Computer Science, 1388:392–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramaswami:1998:PRT

[901] S. Ramaswami. Parallel randomized techniques for some fundamental geometric problems: A survey. Lecture Notes in Computer Science, 1388:400– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narayanan:1998:RAM

[902] L. Narayanan. Randomized algorithms on the mesh. Lecture Notes in Computer Science, 1388:408–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bumble:1998:IPR

[903] M. Bumble and L. Coraor. Implementing parallelism in random discrete event-driven simulation. Lecture Notes in Computer Science, 1388:418– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez:1998:RWN

[904] J. L. Sanchez, J. M. Garcia, and F. J. Alfaro. Reconfigurable wormhole networks: A realistic approach. Lecture Notes in Computer Science, 1388:428– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moukrim:1998:SCD

[905] A. Moukrim and A. Quilliot. Scheduling with communication delays and data routing in message passing architectures. Lecture Notes in Computer Science, 1388:438–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsen:1998:POS

[906] P. K. Nilsen and N. Prcovic. Parallel optimisation in the SCOOP library. Lecture Notes in Computer Science, 1388:452–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Venema:1998:NAS

[907] S. Venema, H. Shen, and F. Suraweera. NC algorithms for the single most vital edge problem with respect to all pairs shortest paths. Lecture Notes in Computer Science, 1388:464–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prylli:1998:BNP

[908] L. Prylli and B. Tourancheau. BIP: A new protocol designed for high performance networking on myrinet. Lecture Notes in Computer Science, 1388:

472–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1998:CPP

[909] Y. Tanaka, M. Matsuda, M. Ando, and K. Kubota. COMPaS: A Pentium Pro PC-based SMP cluster and its experience. Lecture Notes in Computer Science, 1388:486–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1998:PPU

[910] J. M. Blum, T. M. Warschko, and W. F. Tichy. PULC: ParaStation user-level communication. design and overview. Lecture Notes in Computer Science, 1388:498–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryan:1998:EPS

[911] S. J. Ryan and H. Bryhni. Eliminating the protocol stack for socket based communication in shared memory interconnects. Lecture Notes in Computer Science, 1388:510–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciaccio:1998:PMD

[912] G. Ciaccio and V. Di Martino. Porting a molecular dynamics application on a low-cost cluster of personal computer running GAMMA. Lecture Notes in Computer Science, 1388:524–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciaccio:1998:OCP

[913] G. Ciaccio. Optimal communication performance on Fast Ethernet with GAMMA. Lecture Notes in Computer Science, 1388:534–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1998:MNP

[914] M. Baker and G. Fox. MPI on NT: A preliminary evaluation of the available environments. Lecture Notes in Computer Science, 1388:549–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zabarsky:1998:FRD

[915] J. Zabarsky. Failure recovery for distributed processes in single system image clusters. Lecture Notes in Computer Science, 1388:564–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brower:1998:QAC

[916] R. C. Brower. The QCD abacus: A cellular automata formulation for continuous gauge symmetries. Lecture Notes in Computer Science, 1388:584–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1998:AFT

[917] I. Chang, M. A. Hiltunen, and R. D. Schlichting. Affordable fault tolerance through adaptation. Lecture Notes in Computer Science, 1388:585–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fabre:1998:DIF

[918] J.-C Fabre. Designand implementation of the FRIENDS system. Lecture Notes in Computer Science, 1388:604– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1998:GFR

[919] K. Huang, J. Wu, and E. B. Fernandez. A generalized forward recovery checkpointing scheme. Lecture Notes in Computer Science, 1388:623–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fetzer:1998:DFA

[920] C. Fetzer and F. Cristian. Derivation of fail-aware membership service specifications. Lecture Notes in Computer Science, 1388:644–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mishra:1998:TGM

[921] S. Mishra, C. Fetzer, and F. Cristian. The timewheel group membership protocol. Lecture Notes in Computer Science, 1388:664–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Almohammad:1998:FTB

[922] B. Almohammad and B. Bose. Faulttolerant broadcasting in toroidal networks. Lecture Notes in Computer Science, 1388:681–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doering:1998:FAF

[923] A. C. Doering, W. Obeloeer, G. Lustig, and E. Maehle. A flexible approach for a fault-tolerant router. Lecture Notes in Computer Science, 1388:693–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zakrevski:1998:FTM

[924] L. Zakrevski and M. Karpovsky. Faulttolerant message routing for multiprocessors. Lecture Notes in Computer Science, 1388:714–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rooholamini:1998:STF

[925] M. Rooholamini and S. H. Hosseini. Self-testing fault-tolerant real-time systems. Lecture Notes in Computer Science, 1388:731–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quaglia:1998:CPD

[926] F. Quaglia, B. Ciciani, and R. Baldoni. Checkpointing protocols in distributed systems with mobile hosts: A performance analysis. Lecture Notes in Computer Science, 1388:742–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biaz:1998:PAF

[927] S. Biaz and N. H. Vaidya. Performance analysis of a fault-tolerant scheme for location management of mobile hosts. Lecture Notes in Computer Science, 1388:756–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shields:1998:FTM

[928] C. Shields, V. Jain, S. Ntafos, and R. Prakash. Fault tolerant mobility planning for rapidly deployable wireless networks. Lecture Notes in Computer Science, 1388:770–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skillicorn:1998:BBP

[929] D. B. Skillicorn. Building BSP programs using the refinement calculus. Lecture Notes in Computer Science, 1388:790–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gamboa:1998:MVC

[930] R. A. Gamboa. Mechanically verifying the correctness of the fast Fourier transform in ACL2. Lecture Notes in Computer Science, 1388:796–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodeveix:1998:AVP

[931] J. P. Bodeveix and M. Filali. On the automatic validation of parameterized unity programs. Lecture Notes in Computer Science, 1388:807–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charpentier:1998:TUD

[932] M. Charpentier, M. Filali, P. Mauran, and G. Padiou. Tailoring UNITY to distributed program design. Lecture Notes in Computer Science, 1388:820– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thirioux:1998:APU

[933] X. Thirioux. Automatically proving UNITY safety properties with arrays and quantifiers. Lecture Notes in Computer Science, 1388:833–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massingill:1998:EPP

[934] B. L. Massingill. Experiments with program parallelization using archetypes and stepwise refinement. Lecture Notes in Computer Science, 1388:844–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nalumasu:1998:DEC

[935] R. Nalumasu and G. Copalakrishnan. Deriving efficient cache coherence protocols through refinement. Lecture Notes in Computer Science, 1388:857– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruia-Catalin:1998:IMU

[936] R. Gruia-Catalin and P. J. McCann. An introduction to mobile UNITY. Lecture Notes in Computer Science, 1388:871–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Misra:1998:OMM

[937] J. Misra. An object model for multiprogramming. Lecture Notes in Computer Science, 1388:881–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandy:1998:CIS

[938] K. M. Chandy, P. A. G. Sivilotti, and J. R. Kiniry. A cottage industry of software publishing: Implications for theories of composition. Lecture Notes in Computer Science, 1388:890–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Draper:1998:TME

[939] J. Draper, J. Block, J. Koller, and C. Steele. Thermal management in embedded systems using MEMS. Lecture Notes in Computer Science, 1388:900– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scherrer:1998:SMR

[940] D. Scherrer and H. Eberle. A scalable multiprocessor for real-time signal processing. Lecture Notes in Computer Science, 1388:902–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steele:1998:SNS

[941] C. Steele, J. Draper, and J. Koller. Safety net: Secure communications for embedded high-performance computing. Lecture Notes in Computer Science, 1388:908–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diessei:1998:PRS

[942] O. Diessei and H. Elgindy. Partial rearrangements of space-shared FP-GAs. Lecture Notes in Computer Science, 1388:913–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakamraju:1998:MVI

[943] V. Lakamraju, Z. Koren, I. Koren, and C. M. Krishna. Measuring the vulnerability of interconnection networks in embedded systems. Lecture Notes in Computer Science, 1388:919–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Russ:1998:AAR

[944] S. H. Russ. All architecture for rapid distributed fault tolerance. Lecture Notes in Computer Science, 1388:925– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravindran:1998:RMM

[945] B. Ravindran, L. R. Welch, C. Bruggerman, and B. A. Shirazi. A resource management model for dynamic, scalable, dependable, real-time systems. Lecture Notes in Computer Science, 1388:931–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1998:MMD

[946] M. Lee, W. Liu, and V. K. Prasanna. A mapping methodology for designing software task pipelines for embedded signal processing. Lecture Notes in Computer Science, 1388:937–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogler:1998:TED

[947] J. Fogler, T. Robey, and M. Young. Toward embedded development from advanced khoros. Lecture Notes in Computer Science, 1388:945–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nuetzel:1998:OND

[948] J. Nuetzel, B. Daene, and W. Fengler. Object nets for the design and verification of distributed and embedded applications. Lecture Notes in Computer Science, 1388:953–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kubica:1998:DPP

[949] S. Kubica, T. Robey, and C. Moorman. Data parallel programming with the Khoros Data Services Library. Lecture Notes in Computer Science, 1388:963–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krueger:1998:DTE

[950] M. Krueger. A development tool environment for configuration, build, and launch of complex applications. Lecture Notes in Computer Science, 1388: 969–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

West:1998:SCT

[951] J. M. West and J. K. Antonio. Simulation of the communication time for a space-time adaptive processing algorithm on a parallel embedded system. Lecture Notes in Computer Science, 1388:979–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kubica:1998:OCC

[952] S. Kubica, T. Robey, and C. Moorman. Optimal configuration of compute nodes for synthetic aperture radar processing. Lecture Notes in Computer Science, 1388:987–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinez:1998:PAP

[953] J. G. Martinez, J. L. D. De Arriba, J. E. Castaño, and D. F. G. Martinez. Performance analysis of parallel embedded real time systems based on measurement and visualization. Lecture Notes in Computer Science, 1388:1015–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sun:1998:PRC

[954] X.-H. Sun. Performance range comparison via crossing point analysis. Lecture Notes in Computer Science, 1388:1025– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mincev:1998:TDM

[955] V. Mincev and D. Milicev. A treedriven multiple-rate model of time measuring in object-oriented real-time systems. Lecture Notes in Computer Science, 1388:1037–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakajima:1998:RRA

[956] T. Nakajima. Resource reservation for adaptive QOS mapping in real-time mach. Lecture Notes in Computer Science, 1388:1047–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lichtenecker:1998:AVC

[957] R. Lichtenecker, K. Gotthart, and J. Zalewski. Automated verification of communication protocols using CCS and BDDs. Lecture Notes in Computer Science, 1388:1057–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horauer:1998:NNT

[958] M. Horauer, U. Schmid, and K. Schossmaier. NTI: A network time interface M-module for high-accuracy clock synchronization. Lecture Notes in Computer Science, 1388:1067–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaspary:1998:MVS

[959] L. P. Gaspary and M. J. B. Almeida. Modeling and validation support for interactive networked multimedia applications. Lecture Notes in Computer Science, 1388:1077–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Livani:1998:ECC

[960] M. A. Livani and J. Kaiser. EDF consensus on CAN bus access for dynamic real-time applications. Lecture Notes in Computer Science, 1388:1088–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haimes:1998:UPM

[961] R. Haimes and K. E. Jordan. Using PVM and MPI for co-processed, distributed and parallel scientific visualization. Lecture Notes in Computer Science, 1388:1098–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koren:1998:SHM

[962] Z. Koren, I. Koren, and C. M. Krishna. Surge handling as a measure of real-time system dependability. Lecture Notes in Computer Science, 1388:1106–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Son:1998:PST

[963] S. H. Son. Partial security and timeliness in real-time database systems. Lecture Notes in Computer Science, 1388:1117–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishida:1998:SMF

[964] K. Ishida and K. Amano. A synthesis method for fault-tolerant multicast routing protocol. Lecture Notes in Computer Science, 1388:1121–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guenzel:1998:IVL

[965] C. Guenzel, F. Riemenschneider, and J. Wirtgen. Improved videotransmission over lossy channels using parallelization. Lecture Notes in Computer Science, 1388:1131–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kshemkalyani:1998:TSC

[966] A. D. Kshemkalyani. Testing of synchronization conditions for distributed real-time applications. Lecture Notes in Computer Science, 1388:1140–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Irey:1998:ANR

[967] P. M. Irey, R. W. Hott, and D. T. Marlow. An architecture for network resource monitoring in a distributed environment. Lecture Notes in Computer Science, 1388:1153–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doermann:1998:IVS

[968] D. S. Doermann. An introduction to vectorization and segmentation. Lecture Notes in Computer Science, 1389: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wenyin:1998:GGR

[969] L. Wenyin and D. Dori. Genericity in graphics recognition algorithms. Lecture Notes in Computer Science, 1389: 9–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agam:1998:DDL

[970] G. Agam and I. Dinstein. Directional decomposition of line-drawing images based on regulated morphological operations. Lecture Notes in Computer Science, 1389:21–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luo:1998:IDM

[971] H. Luo and R. Kasturi. Improved directional morphological operations for separation of characters from maps/ graphics. Lecture Notes in Computer Science, 1389:35–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramel:1998:CVI

[972] J. Y. Ramel, N. Vincent, and H. Emptoz. A coarse vectorization as an initial representation for the understanding of line drawing images. Lecture Notes in Computer Science, 1389:48–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vialard:1998:EPR

[973] A. Vialard. Euclidean paths for representing and tranforming scanned characters. Lecture Notes in Computer Science, 1389:58–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chhabra:1998:GSR

[974] A. K. Chhabra. Graphic symbol recognition: An overview. Lecture Notes in Computer Science, 1389:68–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ah-Soon:1998:CNS

[975] C. Ah-Soon. A constraint network for symbol detection in architectural drawings. Lecture Notes in Computer Science, 1389:80–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Llados:1998:SBM

[976] J. Llados, G. Sanchez, and E. Marti. A string based method to recognize symbols and structural textures in architectural plans. Lecture Notes in Computer Science, 1389:91–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Francesconi:1998:LRR

[977] E. Francesconi, P. Frasconi, M. Gori, and S. Marinai. Logo recognition by recursive neural networks. Lecture Notes in Computer Science, 1389:104– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1998:MKB

[978] Y.-W. Choi and Y.-B Kwon. Methods for Korean business forms processing. Lecture Notes in Computer Science, 1389:118–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoo:1998:IES

[979] J.-Y. Yoo, M.-K. Kim, Sang Yong Hang, and Y.-B. Kwon. Information extraction from a skewed form document in the presence of crossing characters. Lecture Notes in Computer Science, 1389:139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1998:RRD

[980] K. Lee, H. Byun, and Y. Lee. Robust reconstruction of damaged character images on the form documents. Lecture Notes in Computer Science, 1389: 149–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokokura:1998:LBA

[981] N. Yokokura and T. Watanabe. Layout-based approach for extracting constructive elements of bar-charts. Lecture Notes in Computer Science, 1389:163–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deseilligny:1998:TMA

[982] M. P. Deseilligny, R. Mariani, J. Labiche, and R. Mullot. Topographic maps automatic interpretation: Some proposed strategies. Lecture Notes in Computer Science, 1389:175–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dupont:1998:AIS

[983] F. Dupont, M. P. Deseilligny, and M. Gondran. Automatic interpretation of scanned maps: Reconstruction of contour lines. Lecture Notes in Computer Science, 1389:194–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frischknecht:1998:AIS

[984] S. Frischknecht and E. Kanani. Automatic interpretation of scanned topographic maps: A raster-based approach. Lecture Notes in Computer Science, 1389:207–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Centeno:1998:STM

[985] J. S. Centeno. Segmentation of thematic maps using colour and spatial attributes. Lecture Notes in Computer Science, 1389:221–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schavemaker:1998:IFC

[986] J. G. M. Schavemaker and M. J. T. Reinders. Information fusion for conflict resolution in map interpretation. Lecture Notes in Computer Science, 1389:231–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roux:1998:MAG

[987] M. Roux and H. Maitre. Map analysis for guided interpretation of aerial images. Lecture Notes in Computer Science, 1389:243–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tombre:1998:AED

[988] K. Tombre. Analysis of engineering drawings: State of the art and challenges. Lecture Notes in Computer Science, 1389:257–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuo:1998:SAT

[989] M. H. Kuo. A systematic approach towards reconstructing 3D curved models from multiple 2D views. Lecture Notes in Computer Science, 1389:265– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poliakoff:1998:ECS

[990] J. F. Poliakoff and P. D. Thomas. Error correction in scanned engineering drawings using 3-D knowledge-based reconstruction. Lecture Notes in Computer Science, 1389:280–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baum:1998:GRL

[991] L. S. Baum, J. H. Boose, and R. J. Kelley. Graphics recognition for a large-scale airplane information system. Lecture Notes in Computer Science, 1389:291–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagy:1998:PAA

[992] G. Nagy, A. Samal, S. Seth, and T. Fisher. A prototype for adaptive association of street names with streets on maps. Lecture Notes in Computer Science, 1389:302–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arias:1998:PAG

[993] J. F. Arias, A. Chhabra, and V. Misra. A practical application of graphics recognition: Helping with the extraction of information from telephone company drawings. Lecture Notes in Computer Science, 1389:314–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yagi:1998:TBS

[994] K. Yagi, M. Minoh, and K. Ikeda. Time-based segmentation of artistic drawings. Lecture Notes in Computer Science, 1389:322–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smeulders:1998:DPO

[995] A. W. M. Smeulders and C. De Boer. Design and performance in object recognition. Lecture Notes in Computer Science, 1389:335–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBoer:1998:SGP

[996] C. De Boer and A. W. M. Smeulders. Simulating graphics for performance analysis purposes by combining real world background and synthetic foreground data. Lecture Notes in Computer Science, 1389:347–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wenyin:1998:PSP

[997] L. Wenyin and D. Dori. A proposed scheme for performance evaluation of graphics/text separation algorithms. Lecture Notes in Computer Science, 1389:359–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phillips:1998:PEP

[998] I. T. Phillips, J. Liang, A. K. Chhabra, and R. Haralick. A performance evaluation protocol for graphics recognition systems. Lecture Notes in Computer Science, 1389:372–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chhabra:1998:SIG

[999] A. K. Chhabra and I. T. Phillips. The Second International Graphics Recognition Contest — raster to vector conversion: A report. Lecture Notes in Computer Science, 1389:390–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheideler:1998:I

[1000] Christian Scheideler. 1. introduction.

Lecture Notes in Computer Science, 1390:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1390/13900001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1390/13900001.
pdf.

Scheideler:1998:CMU

[1001] Christian Scheideler. 2. communication mechanisms used in practice. Lecture Notes in Computer Science, 1390:15– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1390/13900015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1390/13900015.
pdf.

Scheideler:1998:T

[1002] Christian Scheideler. 3. terminology. Lecture Notes in Computer Science, 1390:27–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1390/13900027.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1390/13900027.
pdf.

Scheideler:1998:ISF

[1003] Christian Scheideler. 4. introduction to store-and-forward routing. Lecture Notes in Computer Science, 1390:41– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1390/13900041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1390/13900041.
pdf.

Scheideler:1998:RN

[1004] Christian Scheideler. 5. the routing number. Lecture Notes in Computer Science, 1390:47–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1390/13900047.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1390/13900047.pdf.

Scheideler:1998:ORPa

[1005] Christian Scheideler. 6. offline routing protocols. Lecture Notes in Computer Science, 1390:57–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1390/13900057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1390/13900057.
pdf.

Scheideler:1998:ORPb

[1006] Christian Scheideler. 7. oblivious routing protocols. Lecture Notes in Computer Science, 1390:73–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1390/13900073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1390/13900073.
pdf.

Scheideler:1998:ARP

[1007] Christian Scheideler. 8. adaptive routing protocols. Lecture Notes in Computer Science, 1390:115–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1390/13900115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1390/13900115.
pdf.

Scheideler:1998:CRP

[1008] Christian Scheideler. 9. compact routing protocols. Lecture Notes in Computer Science, 1390:139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1390/13900139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1390/13900139.
pdf.

Scheideler:1998:IWR

[1009] Christian Scheideler. 10. introduction to wormhole routing. Lecture Notes in Computer Science, 1390:163– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1390/13900163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1390/13900163.
pdf.

Scheideler:1998:ORPc

[1010] Christian Scheideler. 11. oblivious routing protocols. Lecture Notes in Computer Science, 1390:167–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1390/13900167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1390/13900167.
pdf.

Scheideler:1998:PAO

[1011] Christian Scheideler. 12. protocols for all-optical networks. Lecture Notes in Computer Science, 1390:179–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1390/13900179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1390/13900179.
pdf.

Scheideler:1998:SFD

[1012] Christian Scheideler. 13. summary and future directions. Lecture Notes in Computer Science, 1390:209–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1390/13900209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1390/13900209.
pdf.

Scheideler:1998:URS

[1013] C. Scheideler. Universal routing strategies for interconnection networks. Lecture Notes in Computer Science, 1390: ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poli:1998:RTE

[1014] Riccardo Poli and W. B. Langdon. A review of theoretical and experimental results on schemata in genetic programming. Lecture Notes in Computer Science, 1391:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910001.
pdf.

Goldberg:1998:WDG

[1015] David E. Goldberg and Una-May O’Reilly. Where does the good stuff go, and why? how contextual semantics influences program structure in simple genetic programming. Lecture Notes in Computer Science, 1391:16– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910016.
pdf.

Langdon:1998:FCB

[1016] W. B. Langdon and R. Poli. Fitness causes bloat: Mutation. Lecture Notes in Computer Science, 1391:37– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910037.
pdf.

Nikolaev:1998:CIG

[1017] Nikolay I. Nikolaev and Vanio Slavov. Concepts of inductive genetic programming. Lecture Notes in Computer Science, 1391:49–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910049.
pdf.

Aler:1998:ITG

[1018] Ricardo Aler. Immediate transfer of global improvements to all individuals in a population compared to automatically defined functions for the EVEN-5,6-PARITY problems. Lecture Notes in Computer Science, 1391:60– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910060.
pdf.

Ito:1998:NDD

[1019] Takuya Ito, Hitoshi Iba, and Satoshi Sato. Non-destructive depth-dependent crossover for genetic programming. Lecture Notes in Computer Science, 1391:71–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1391/13910071.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1391/13910071.
pdf.

Ryan:1998:GEE

[1020] Conor Ryan, J. J. Collins, and Michael O. Neill. Grammatical evolution: Evolving programs for an arbitrary language. Lecture Notes in Computer Science, 1391:83–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910083.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1391/13910083.
                                        pdf.

Langdon:1998:GPB

[1021] W. B. Langdon and R. Poli. Genetic programming bloat with dynamic fitness. Lecture Notes in Computer Science, 1391:97–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1391/13910097.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1391/13910097.pdf.

Conrads:1998:SSD

[1022] Markus Conrads, Peter Nordin, and Wolfgang Banzhaf. Speech sound discrimination with genetic programming. Lecture Notes in Computer Science, 1391:113–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1391/13910113.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1391/13910113.
                                        pdf.

FigueiraPujol:1998:EEA

[1023] J. C. Figueira Pujol and R. Poli. Efficient evolution of asymmetric recurrent neural networks using a PDGPinspired two-dimensional representation. Lecture Notes in Computer Science, 1391:130–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pujol:1998:EEA

[1024] João Carlos Figueira Pujol and Riccardo Poli. Efficient evolution of asymmetric recurrent neural networks using a PDGP-inspired twodimensional representation. Lecture Notes in Computer Science, 1391:130– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910130.
pdf.

Adorni:1998:CPA

[1025] Giovanni Adorni, Federico Bergenti, and Stefano Cagnoni. A cellularprogramming approach to pattern classification. Lecture Notes in Computer Science, 1391:142–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1391/13910142.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1391/13910142.pdf.

Andersson:1998:ECM

[1026] Claes Andersson and Mats G. Nordahl. Evolving coupled map lattices for computation. Lecture Notes in Computer Science, 1391:151–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910151.
pdf.

Calderoni:1998:GPA

[1027] Stéphane Calderoni and Pierre Marcenac.

Genetic programming for automatic design of self-adaptive robots. Lecture Notes in Computer Science, 1391:163– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910163.
pdf.

Eiben:1998:GMC

[1028] A. E. Eiben, A. E. Koudijs, and F. Slisser. Genetic modelling of customer retention. Lecture Notes in Computer Science, 1391:178–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910178.
pdf.

Ochi:1998:EHM

[1029] Luiz S. Ochi, Dalessandro S. Vianna, Lúcia M. A. Drummond, and André O. Victor. An evolutionary hybrid metaheuristic for solving the vehicle routing problem with heterogeneous fleet. Lecture Notes in Computer Science, 1391: 187–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910187.
pdf.

Lenaerts:1998:BGP

[1030] Tom Lenaerts and Bernard Manderick. Building a genetic programming framework: The added-value of design patterns. Lecture Notes in Computer Science, 1391:196–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910196.
pdf.

Reiser:1998:ECT

[1031] Philip G. K. Reiser. Evolutionary computation and the tinkerer’s evolving toolbox. Lecture Notes in Computer Science, 1391:209–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910209.
pdf.

Racine:1998:DLE

[1032] Alain Racine, Marc Schoenauer, and Philippe Dague. A dynamic lattice to evolve hierarchically shared subroutines. Lecture Notes in Computer Science, 1391:220–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1391/13910220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1391/13910220.
pdf.

Barth:1998:P

[1033] Andreas Barth, Michael Breu, Albert Endres, and Arnoud de Kemp. Preface. Lecture Notes in Computer Science, 1392:0–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1392/13920000.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1392/13920000.
                                        pdf.

Barth:1998:MDL

[1034] A. Barth, M. Breu, A. Brüggemann-Klein, A. Endres, and A. de Kemp. The MeDoc digital library project — its goals and major achievements. Lecture Notes in Computer Science, 1392: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1392/13920001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1392/13920001.
                                        pdf.

Boles:1998:MSD

[1035] D. Boles, M. Dreger, K. Großjohann, C. Haber, A. Kusserow, S. Lohrum, D. Menke, J. Meyer, G. Möller, and R. Weber. The MeDoc system — A digital publication and reference service for computer science. Lecture Notes in Computer Science, 1392:13– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1392/13920013.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1392/13920013.
pdf.

Meyer:1998:DIM

[1036] J. Meyer and H.-J. Appelrath. Design and implementation of the MeDoc fulltext system. Lecture Notes in Computer Science, 1392:21– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1392/13920021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1392/13920021.
pdf.

Boles:1998:CUA

[1037] D. Boles, A. Kusserow, G. Möller, H.- J. Appelrath, and D. Fellner. Conception of the user agent — the user interface of the MeDoc system. Lecture Notes in Computer Science, 1392:35– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1392/13920035.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1392/13920035.
pdf.

Menke:1998:ABD

[1038] D. Menke, C. Gnörlich, and H. Helbig. Access to bibliographic data bases — the provider agents in MeDoc. Lecture Notes in Computer Science, 1392: 51–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1392/13920051.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1392/13920051.
                                        pdf.

Dreger:1998:PSD

[1039] M. Dreger, N. Fuhr, K. Großjohann, and S. Lohrum. Provider selection — design and implementation of the MeDoc broker. Lecture Notes in Computer Science, 1392:67– 78, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1392/13920067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1392/13920067.
                                        pdf.

Grossjohann:1998:QTH

[1040] K. Großjohann and D. Menke. Query transformation for heterogeneous provider systems in MeDoc. Lecture Notes in Computer Science, 1392: 79–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1392/13920079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1392/13920079.
                                        pdf.

Dreger:1998:MCP

[1041] Markus Dreger, Stefan Lohrum, and Peter Müller. The MeDoc communication protocol. Lecture Notes in Computer Science, 1392:89–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1392/13920089.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1392/13920089.
pdf.

Bruggemann-Klein:1998:MLS

[1042] A. Brüggemann-Klein and C. Haber.

The MeDoc library: Security aspects. Lecture Notes in Computer Science, 1392:103–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1392/13920103.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1392/13920103.
pdf.

Weber:1998:CMA

[1043] R. Weber, M. Breu, and A. Endres. Cost models and accounting in MeDoc. Lecture Notes in Computer Science, 1392:115–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1392/13920115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1392/13920115.
pdf.

Dreger:1998:AIN

[1044] M. Dreger, S. Lohrum, H. Schweppe, and C. D. Ziegler. Ariadne, an interactive navigation and search system for computer — science information on the World-Wide Web. Lecture Notes in Computer Science, 1392:127– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1392/13920127.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1392/13920127.
                                        pdf.

Adler:1998:GLM

[1045] S. Adler, U. Berger, A. Brüggemann-Klein, C. Haber, W. Lamersdorf, M. Münke, S. Rücker, and H. Spahn. Grey literature and multiple collections in NCSTRL. Lecture Notes in Computer Science, 1392:145–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1392/13920145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1392/13920145.
                                        pdf.

Fellner:1998:DPC

[1046] Dieter W. Fellner. Digital publishing for conference proceedings on CD-ROM. Lecture Notes in Computer Science, 1392:171–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1392/13920171.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1392/13920171.pdf.

Schwab:1998:BJM

[1047] U. Schwab. Books and journals in MeDoc — experiences on bringing publications online. Lecture Notes in Computer Science, 1392:179–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1392/13920179.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1392/13920179.
pdf.

Feyrer:1998:EDC

[1048] Hubert Feyrer. Experiences in document conversions during the MeDoc project. Lecture Notes in Computer Science, 1392:187–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1392/13920187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1392/13920187.
pdf.

Breu:1998:DBM

[1049] Michael Breu. Digital books in MeDoc — A first analysis of usage patterns. Lecture Notes in Computer Science, 1392:193–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1392/13920193.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1392/13920193.
pdf.

Spillner:1998:UMO

[1050] Andreas Spillner. Use of MeDoc offerings in technical education. Lecture Notes in Computer Science, 1392:207– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1392/13920207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1392/13920207.
pdf.

Roescheisen:1998:SIS

[1051] M. Roescheisen, M. Baldonado, K. Chang, and L. Gravano. The Stanford InfoBus and its service layers: Augmenting the Internet with higherlevel information management protocols. Lecture Notes in Computer Science, 1392:213–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roscheisen:1998:SIS

[1052] Martin Röscheisen, Michelle Baldonado, Kevin Chang, Luis Gravano, Steven Ketchpel, and Andreas Paepcke. The stanford InfoBus and its service layers: Augmenting the Internet with higher-level information management protocols. Lecture Notes in Computer Science, 1392:213– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1392/13920213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1392/13920213.
pdf.

Dijkstra:1998:ORC

[1053] Joost Dijkstra. Objectives, results and conclusions of the european DE-COMATE project: Delivery of copyright material to end users. Lecture Notes in Computer Science, 1392:231– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1392/13920231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1392/13920231.
pdf.

Abrial:1998:B

[1054] Jean-Raymond Abrial. On B. Lecture Notes in Computer Science, 1393: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1393/13930001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1393/13930001.
pdf.

Morgan:1998:GSL

[1055] Carroll Morgan. The generalised substitution language extended to probabilistic programs. Lecture Notes in Computer Science, 1393:9–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1393/13930009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1393/13930009.
pdf.

Behm:1998:MIS

[1056] Patrick Behm, Pierre Desforges, and Jean-Marc Meynadier. MÉTÉOR : An industrial success in formal development. Lecture Notes in Computer Science, 1393:26–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1393/13930026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1393/13930026.
pdf.

Pilarski:1998:CEF

[1057] François Pilarski. Cost effectiveness of formal methods in the development of avionics systems at AÉROSPATIALE. Lecture Notes in Computer Science, 1393:27–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1393/13930027.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1393/13930027.
                                        pdf.

Behm:1998:WDB

[1058] Patrick Behm, Lilian Burdy, and Jean-Marc Meynadier. Well defined B. Lecture Notes in Computer Science, 1393: 29–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1393/13930029.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1393/13930029.
                                        pdf.

Potet:1998:CRB

[1059] Marie-Laure Potet and Yann Rouzaud. Composition and refinement in the Bmethod. Lecture Notes in Computer Science, 1393:46–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1393/13930046.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1393/13930046.pdf.

Chartier:1998:FBI

[1060] Pierre Chartier. Formalisation of B in Isabelle/HOL. Lecture Notes in Computer Science, 1393:66–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1393/13930066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1393/13930066.
pdf.

Abrial:1998:IDC

[1061] Jean-Raymond Abrial and Louis Mussat. Introducing dynamic constraints in B. Lecture Notes in Computer Science, 1393:83–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1393/13930083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1393/13930083.
pdf.

Banach:1998:REV

[1062] R. H. Banach and M. R. Poppleton. Retrenchment: An engineering variation on refinement. Lecture Notes in Computer Science, 1393:129– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1393/13930129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1393/13930129.
pdf.

Matthews:1998:SSF

[1063] Brian Matthews, Brian Ritchie, and Juan Bicarregui. Synthesising structure from flat specifications. Lecture Notes in Computer Science, 1393:148– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1393/13930148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1393/13930148.
                                        pdf.

Malioukov:1998:OBA

[1064] Alexander Malioukov. An object-based approach to the B formal method. Lecture Notes in Computer Science, 1393: 162–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1393/13930162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1393/13930162.
                                        pdf.

Sekerinski:1998:GDR

[1065] Emil Sekerinski. Graphical design of reactive systems. Lecture Notes in Computer Science, 1393:182–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1393/13930182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1393/13930182.
                                        pdf.

Petin:1998:PCE

[1066] Jean-François Pétin, Gérard Morel, Dominique Méry, and Patrick Lamboley. Process control engineering: Contribution to a formal structuring framework with the B method. Lecture Notes in Computer Science, 1393:198– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1393/13930198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1393/13930198.
pdf.

Taouil-Traverson:1998:DBM

[1067] Souâd Taouil-Traverson and Sylvie Vignes. Designing a B model for safetycritical software systems. Lecture Notes in Computer Science, 1393:210– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1393/13930210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1393/13930210.
pdf.

Stoddart:1998:ASM

[1068] Bill Stoddart, Steve Dunne, Andy Galloway, and Richard Shore. Abstract state machines: Designing distributed systems with state machines and B. Lecture Notes in Computer Science, 1393:226–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1393/13930226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1393/13930226.
pdf.

Walden:1998:LDA

[1069] M. Waldén. Layering distributed algorithms within the B-method. Lecture Notes in Computer Science, 1393: 243–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1393/13930243.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1393/13930243.
                                        pdf.

Heuberger:1998:TSD

[1070] Philipp Heuberger. Two strategies to data-refine an equivalence to a forest. Lecture Notes in Computer Science, 1393:261–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1393/13930261.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1393/13930261.
                                        pdf.

Julliand:1998:SIC

[1071] Jacques Julliand, Bruno Legeard, Thierry Machicoane, Benoit Parreaux, and Bruno Tatibouët. Specification of an integrated circuit card protocol application using the B method and linear temporal logic. Lecture Notes in Computer Science, 1393:273– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1393/13930273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1393/13930273.
                                        pdf.

Treharne:1998:TCP

[1072] H. Treharne, J. Draper, and S. Schneider. Test case preparation using a prototype. Lecture Notes in Computer Science, 1393:293–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1393/13930293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1393/13930293.
pdf.

Albrecht:1998:KAG

[1073] D. W. Albrecht, A. E. Nicholson, and I. Zukerman. Knowledge acquisition for goal prediction in a multi-user adventure game. Lecture Notes in Computer Science, 1394:1–12, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anand:1998:HDM

[1074] S. S. Anand and J. G. Hughes. Hybrid data mining systems: The next generation. Lecture Notes in Computer Science, 1394:13–24, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anand:1998:DCK

[1075] S. S. Anand, D. Patterson, J. G.

Hughes, and D. A. Bell. Discovering case knowledge using data mining. Lecture Notes in Computer Science, 1394:25–35, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buechter:1998:DAR

[1076] O. Buechter and R. Wirth. Discovery of association rules over ordinal data: A new and faster algorithm and its application to basket analysis. Lecture Notes in Computer Science, 1394:36–47, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheung:1998:EDS

[1077] D. W. Cheung and Y. Xiao. Effect of data skewness in parallel mining of association rules. Lecture Notes in Computer Science, 1394:48–60, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dai:1998:TDL

[1078] H. Dai. Trend directed learning: A case study. Lecture Notes in Computer Science, 1394:61–71, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dong:1998:IDA

[1079] G. Dong and J. Li. Interestingness of discovered association rules in terms of neighborhood-based unexpectedness. Lecture Notes in Computer Science, 1394:72–86, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dowe:1998:PEU

[1080] D. L. Dowe, R. A. Baxter, J. J. Oliver, and C. S. Wallace. Point estimation using the Kullback–Leibler loss function and MML. Lecture Notes in Computer Science, 1394:87–95, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edwards:1998:SFA

[1081] R. T. Edwards and D. L. Dowe. Single factor analysis in MML mixture modelling. Lecture Notes in Computer Science, 1394:96–100, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estivill-Castro:1998:DAS

[1082] V. Estivill-Castro and A. T. Murray. Discovering associations in spatial data — an efficient medoid based approach. Lecture Notes in Computer Science, 1394:110–121, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frayman:1998:DMU

[1083] Y. Frayman and L. Wang. Data mining using dynamically constructed recurrent fuzzy neural networks. Lecture Notes in Computer Science, 1394:122– 131, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gray:1998:CCC

[1084] B. Gray and M. E. Orlowska. CCAIIA:

Clustering categorical attributes into interesting association rules. Lecture Notes in Computer Science, 1394:132– 143, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1998:SME

[1085] J. Han, N. Stefanovic, and K. Koperski. Selective materialization: An efficient method for spatial data cube construction. Lecture Notes in Computer Science, 1394:144–158, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hilderman:1998:MMB

[1086] R. J. Hilderman, C. L. Carter, H. J. Hamilton, and N. Cercone. Mining market basket data using share measures and characterized itemsets. Lecture Notes in Computer Science, 1394: 159–170, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iizuka:1998:AVM

[1087] Y. Iizuka, H. Shiohara, T. Iizuka, and S. Isobe. Automatic visualization method for visual data mining. Lecture Notes in Computer Science, 1394:171–185, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kowalczyk:1998:RSI

[1088] W. Kowalczyk and Z. Piasta. Rough set-inspired approach to knowledge discovery in business databases. Lecture Notes in Computer Science, 1394:186– 197, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kryszkiewicz:1998:RAR

[1089] M. Kryszkiewicz. Representative association rules. Lecture Notes in Computer Science, 1394:198–209, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1998:IRD

[1090] H. Liu, H. Lu, and J. Yao. Identifying relevant databases for multidatabase mining. Lecture Notes in Computer Science, 1394:210–221, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliver:1998:MML

[1091] J. J. Oliver, R. A. Baxter, and C. S.

Wallace. Minimum message length segmentation. Lecture Notes in Computer Science, 1394:222–233, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paass:1998:BCT

[1092] G. Paass and J. Kindermann. Bayesian classification trees with overlapping leaves applied to credit-scoring. Lecture Notes in Computer Science, 1394:234– 245, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perrin:1998:CTR

[1093] P. Perrin and F. Petry. Contextual text representation for unsupervised knowledge discovery in texts. Lecture Notes in Computer Science, 1394:246–257, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ragel:1998:TMV

[1094] A. Ragel and B. Cremilleux. Treatment of missing values for association rules. Lecture Notes in Computer Science, 1394:258–270, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sher:1998:MRR

[1095] B.-Y. Sher, S.-C. Shao, and W.-S.

Hsieh. Mining regression rules and regression trees. Lecture Notes in Computer Science, 1394:271–282, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shintani:1998:MAS

[1096] T. Shintani and M. Kitsuregawa. Mining algorithms for sequential patterns in parallel: Hash based approach. Lecture Notes in Computer Science, 1394: 283–294, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Struzik:1998:WTS

[1097] Z. R. Struzik and A. Siebes. Wavelet transform in similarity paradigm. Lecture Notes in Computer Science, 1394: 295–309, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tolun:1998:IRD

[1098] M. R. Tolun, H. Sever, and M. Uludag. Improved rule discovery performance on uncertainty. Lecture Notes in Computer Science, 1394:310–321, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeVel:1998:FMM

[1099] O. De Vel, D. Coomans, and S. Patrick. Feature mining and mapping of collinear data. Lecture Notes in Computer Science, 1394:322–335, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeleznikow:1998:KDD

[1100] J. Zeleznikow and A. Stranieri. Knowledge discovery in discretionary legal domains. Lecture Notes in Computer Science, 1394:336–347, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1998:SRG

[1101] Z. Zheng. Scaling up the rule generation of C4.5. Lecture Notes in Computer Science, 1394:348–359, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1998:DMB

[1102] N. Zhong, J. Dong, and S. Ohsuga.

Data mining based on the generalization distribution table and rough sets. Lecture Notes in Computer Science, 1394:360–373, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1998:CPI

[1103] C.-H. Chang and C.-C. Hsu. Constructing personalized information agents. Lecture Notes in Computer Science, 1394:374–375, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1998:TRT

[1104] V. Cho and B. Wuethrich. Towards real time discovery from distributed information sources. Lecture Notes in Computer Science, 1394:376–377, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cole:1998:CCS

[1105] R. Cole, P. Eklund, and D. Walker.

Constructing conceptual scales in formal concept analysis. Lecture Notes in Computer Science, 1394:378–379, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dowe:1998:HHM

[1106] D. L. Dowe, L. Allison, and G. Pringle. The hunter and the hunted — modelling the relationship between Web pages and search engines. Lecture Notes in Computer Science, 1394:380–382, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ho:1998:EGD

[1107] K. M. Ho and P. D. Scott. An efficient global discretization method. Lecture Notes in Computer Science, 1394:383– 384, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquenet:1998:LUP

[1108] F. Jacquenet and P. Brenot. Learning user preferences on the WEB. Lecture Notes in Computer Science, 1394:385– 387, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jagielska:1998:URS

[1109] I. Jagielska. Using rough sets for knowledge discovery in the development of a decision support system for issuing smog alerts. Lecture Notes in Computer Science, 1394:388–389, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koshizen:1998:ERD

[1110] T. Koshizen, H. Ogawa, and J. Fulcher. Empirical results on data dimensionality reduction using the divided selforganizing map. Lecture Notes in Computer Science, 1394:390–391, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1998:MAR

[1111] Deyi Li, Kaichang Di, Deren Li, and Xuemei Shi. Mining association rules with linguistic cloud models. Lecture Notes in Computer Science, 1394:392– 393, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1998:DMA

[1112] Y. Liu, H. Chen, J. X. Yu, and N. Ohbo. A data mining approach for query refinement. Lecture Notes in Computer Science, 1394:394–396, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1998:CCF

[1113] Y. Lu, H. Liu, and C. L. Tan. CFMD: A conflict-free multivariate discretization algorithm. Lecture Notes in Computer Science, 1394:397–398, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maeda:1998:CRI

[1114] A. Maeda, H. Maki, and H. Akimori.

Characteristic rule induction algorithm for data mining. Lecture Notes in Computer Science, 1394:399–400, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ng:1998:DMM

[1115] M. K. Ng, Z. Huang, and M. Hegland. Data-mining massive time series astronomical data sets — A case study. Lecture Notes in Computer Science, 1394:401–402, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nowak:1998:MDP

[1116] C. Nowak. Multiple databases, partial reasoning, and knowledge discovery. Lecture Notes in Computer Science, 1394:403–404, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeOca:1998:DRD

[1117] C. M. De Oca and D. L. Carver. Design recovery with data mining techniques. Lecture Notes in Computer Science, 1394:405–406, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pearce:1998:CA

[1118] A. R. Pearce and T. Caelli. The CLARET algorithm. Lecture Notes in Computer Science, 1394:407–408, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsien:1998:LTH

[1119] C. L. Tsien, S. M. Hamish, S. F. Fraser, and I. S. Kohane. LR tree: A hybrid technique for classifying myocardial infarction data containing unknown attribute values. Lecture Notes in Computer Science, 1394:409–411, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wets:1998:MDT

[1120] G. Wets, J. Vanthienen, and H. Timmermans. Modelling decision tables from data. Lecture Notes in Computer Science, 1394:412–413, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vazirgiannis:1998:CRE

[1121] M. Vazirgiannis. A classification and relationship extraction scheme for relational databases based on fuzzy logic. Lecture Notes in Computer Science, 1394:414–416, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Washio:1998:MAR

[1122] T. Washio, H. Matsuura, and H. Motoda. Mining association rules for estimation and prediction. Lecture Notes in Computer Science, 1394:417–419, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xue:1998:RGC

[1123] H. Xue and Q. Cai. Rule generalization by condition combination. Lecture Notes in Computer Science, 1394:420– 422, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitano:1998:RCP

[1124] H. Kitano, M. Asada, Y. Kuniyoshi, and I. Noda. RoboCup: A challenge problem for AI and robotics. Lecture Notes in Computer Science, 1395:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noda:1998:OR

[1125] I. Noda, S. Suzuki, H. Matsubara, and M. Asada. Overview of RoboCup-97. Lecture Notes in Computer Science, 1395:20–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asada:1998:RPA

[1126] M. Asada, P. Stone, H. Kitano, and A. Drogoul. The RoboCup physical agent challenge: Goals and protocols for phase I. Lecture Notes in Computer Science, 1395:42–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitano:1998:RSA

[1127] H. Kitano, M. Tambe, P. Stone, and M. Veloso. The RoboCup synthetic agent challenge 97. Lecture Notes in Computer Science, 1395:62–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riekki:1998:PSM

[1128] J. Riekki and J. Roening. Playing soccer by modifying and combining primitive reactions. Lecture Notes in Computer Science, 1395:74–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Price:1998:LDP

[1129] A. Price, A. Jennings, J. Kneen, and E. Kendall. Learning, deciding, predicting: The soccer playing mind. Lecture Notes in Computer Science, 1395: 88–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stone:1998:UDT

[1130] P. Stone and M. Veloso. Using decision tree confidence factors for multiagent control. Lecture Notes in Computer Science, 1395:99–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coradeschi:1998:RBD

[1131] S. Coradeschi and L. Karlsson. A rolebased decision-mechanism for teams of reactive and coordinating agents. Lecture Notes in Computer Science, 1395: 112–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tambe:1998:UEM

[1132] M. Tambe, J. Adibi, Y. Al-Onaizan, and A. Erdem. Using an explicit model of teamwork in RoboCup-97. Lecture Notes in Computer Science, 1395:123– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsumoto:1998:DMC

[1133] A. Matsumoto and H. Nagai. Decision making by the characteristics and the interaction in multi-agent robotics soccer. Lecture Notes in Computer Science, 1395:132–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1998:RTV

[1134] G. Cheng and A. Zelinsky. Real-time vision processing for a soccer playing mobile robot. Lecture Notes in Computer Science, 1395:144–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizuno:1998:MAS

[1135] H. Mizuno, M. Kourogi, Y. Kawamoto, and Y. Muraoka. A method applied for soccer’s behaviors using proper feedback and feedforward control. Lecture Notes in Computer Science, 1395:156– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1998:LRR

[1136] M. Fujita, H. Kitano, and K. Kageyama. A legged robot for RoboCup based on “OPENR”. Lecture Notes in Computer Science, 1395:168–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balch:1998:J

[1137] T. Balch. JavaSoccer. Lecture Notes in Computer Science, 1395:181–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shinjoh:1998:RCI

[1138] A. Shinjoh. RoboCup-3D: The construction of intelligent navigation system. Lecture Notes in Computer Science, 1395:188–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andre:1998:GMP

[1139] E. Andre, G. Herzog, and T. Rist. Generating multimedia presentations for RoboCup soccer games. Lecture Notes in Computer Science, 1395:200– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frank:1998:FRT

[1140] I. Frank. Football in recent times: What we can learn from the newspapers. Lecture Notes in Computer Science, 1395:216–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verner:1998:VPB

[1141] I. M. Verner. The value of projectbased education in robotics. Lecture Notes in Computer Science, 1395:231– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veloso:1998:CSR

[1142] M. Veloso, P. Stone, K. Han, and S. Achim. The CMUnited-97 small robot team. Lecture Notes in Computer Science, 1395:242–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1998:DSL

[1143] T. Nakamura. Development of selflearning vision-based mobile robots for acquiring soccer robots behaviors. Lecture Notes in Computer Science, 1395: 257–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rocher:1998:MFE

[1144] S. Rocher, V. Idasiak, S. Doncker, and A. Drogoul. MICROB: The French experiment in RoboCup. Lecture Notes in Computer Science, 1395:277–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oller:1998:DRT

[1145] A. Oller, R. Garcia, J. A. Ramon, and A. Figueras. Description of rogiteam. Lecture Notes in Computer Science, 1395:286–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shen:1998:ASR

[1146] W.-M. Shen, J. Adibi, R. Adobbati, and B. Cho. Autonomous soccer robots. Lecture Notes in Computer Science, 1395:295–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1998:VBR

[1147] S. Suzuki, Y. Takahashi, E. Uchibe, and M. Nakamura. Vision-based robot learning towards RoboCup: Osaka University “trackies”. Lecture Notes in Computer Science, 1395:305–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Price:1998:ROP

[1148] A. Price, A. Jennings, and J. Kneen. RoboCup97: An omnidirectional perspective. Lecture Notes in Computer Science, 1395:320–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokota:1998:ODA

[1149] K. Yokota, K. Ozaki, A. Matsumoto, and K. Kawabata. Omni-directional autonomous robots cooperating for team play. Lecture Notes in Computer Science, 1395:333–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brian:1998:SBC

[1150] B. Brian, M. Schneider-Fontan, R. Sargent, and C. Witty. The spirit of bolivia: Complex behavior through minimal control. Lecture Notes in Computer Science, 1395:348–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkhard:1998:HDP

[1151] H.-D. Burkhard, M. Hannebauer, and J. Wendler. AT Humboldt — development, practice and theory. Lecture Notes in Computer Science, 1395:357– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andou:1998:RSA

[1152] T. Andou. Refinement of soccer agents’ positions using reinforcement learning. Lecture Notes in Computer Science, 1395:373–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stone:1998:CST

[1153] P. Stone and M. Veloso. The CMUnited-97 simulator team. Lecture Notes in Computer Science, 1395:389– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luke:1998:CES

[1154] S. Luke, C. Hohn, J. Farris, and G. Jackson. Co-evolving soccer softbot team coordination with genetic programming. Lecture Notes in Computer Science, 1395:398–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohta:1998:LCB

[1155] M. Ohta. Learning cooperative behaviors in RoboCup agents. Lecture Notes in Computer Science, 1395:412– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Igarashi:1998:ITP

[1156] H. Igarashi, S. Kosue, M. Miyahara, and T. Umaba. Individual tactical play and action decision based on a shortterm goal — team descriptions of Team Miya and Team Niken. Lecture Notes in Computer Science, 1395:420–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takaki:1998:RMP

[1157] S. Takaki. The reactive motion planning in the passive situation. Lecture Notes in Computer Science, 1395:428– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pagello:1998:RAR

[1158] E. Pagello, F. Montesello, A. D’Angelo, and C. Ferrari. A reactive architecture for RoboCup competition. Lecture Notes in Computer Science, 1395: 434–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inden:1998:TKB

[1159] T. Inden and T. Takahashi. Team: Kasuga-bitos with modulation of playing. Lecture Notes in Computer Science, 1395:443–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fry:1998:TS

[1160] J. Fry, L. Huang, and S. Peters. Team sicily. Lecture Notes in Computer Science, 1395:450–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chng:1998:TDB

[1161] S. Ch’ng and L. Padgham. Team description: Building teams using roles, responsibilities, and strategies. Lecture Notes in Computer Science, 1395:458– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scerri:1998:MLB

[1162] P. Scerri. A multi-layered behavior based system for controlling RoboCup agents. Lecture Notes in Computer Science, 1395:467–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matellan:1998:UAR

[1163] V. Matellan, D. Borrajo, and C. Fernandez. Using ABC02 in the RoboCup domain. Lecture Notes in Computer Science, 1395:475–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balch:1998:ILM

[1164] T. Balch. Integrating learning with motor schema-based control for a robot soccer team. Lecture Notes in Computer Science, 1395:483–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1998:DSG

[1165] N. Ito, T. Hotta, and N. Ishii. Describing soccer game in EAMMO. Lecture Notes in Computer Science, 1395:492– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noda:1998:TGA

[1166] I. Noda. Team GAMMA: Agent programming on gaea. Lecture Notes in Computer Science, 1395:500–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1998:URD

[1167] Y. Zhang and A. K. Mackworth. Using reactive deliberation for real-time control of soccer-playing robots. Lecture Notes in Computer Science, 1395: 508–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iso:1998:MLP

[1168] R. Iso and H. Inazumi. A multilayered planning architecture for soccer agent. Lecture Notes in Computer Science, 1395:513–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blakley:1998:GTC

[1169] G. R. Blakley and I. Borosh. A general theory of codes, II: Paradigms and homomorphisms. Lecture Notes in Computer Science, 1396:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimoyama:1998:IHO

[1170] T. Shimoyama, S. Moriai, and T. Kaneko. Improving the higher order differential attack and cryptanalysis of the KN cipher. Lecture Notes in Computer Science, 1396:32–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1998:OLA

[1171] H. Tanaka, T. Ohishi, and T. Kaneko.

An optimised linear attack on pseudorandom generators using a non linear combiner. Lecture Notes in Computer Science, 1396:43–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preneel:1998:CMA

[1172] B. Preneel. Cryptanalysis of message authentication codes. Lecture Notes in Computer Science, 1396:55–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balasubramanian:1998:LWC

[1173] R. Balasubramanian and S. V. Nagaraj. The least witness of a composite number. Lecture Notes in Computer Science, 1396:66–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuwakado:1998:FAF

[1174] H. Kuwakado and H. Tanaka. Fast algorithm for finding a small root of a quadratic modular equation. Lecture Notes in Computer Science, 1396:75– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bao:1998:MFA

[1175] F. Bao, R. H. Deng, X. Gao, and Y. Igarashi. Modified finite automata public key cryptosystem. Lecture Notes in Computer Science, 1396:82– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1998:MEC

[1176] D. Nakamura and K. Kobayashi. Modified ElGamal cryptosystem. Lecture Notes in Computer Science, 1396:96– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohta:1998:RBD

[1177] K. Ohta. Remarks on blind decryption. Lecture Notes in Computer Science, 1396:109–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davida:1998:HSC

[1178] G. Davida and R. Peralta. Highspeed cryptography. Lecture Notes in Computer Science, 1396:116–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelsey:1998:SAL

[1179] J. Kelsey, B. Schneier, C. Hall, and D. Wagner. Secure applications of low-entropy keys. Lecture Notes in Computer Science, 1396:121–134, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://www.counterpane.com/
                                        low-entropy.html.

Sameshima:1998:KES

[1180] Y. Sameshima. A key escrow system of the RSA cryptosystem. Lecture Notes in Computer Science, 1396: 135–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakurai:1998:KES

[1181] K. Sakurai, Y. Yamane, S. Miyazaki, and T. Inoue. A key escrow system with protecting user’s privacy by blind decoding. Lecture Notes in Computer Science, 1396:147–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desmedt:1998:SRR

[1182] Y. Desmedt. Some recent research aspects of threshold cryptography. Lecture Notes in Computer Science, 1396: 158–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Satoh:1998:HSS

[1183] A. Satoh, Y. Kobayashi, H. Niijima, and N. Ooba. A high-speed small RSA encryption LSI with low power dissipation. Lecture Notes in Computer Science, 1396:174–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markantonakis:1998:CSM

[1184] C. Markantonakis. The case for a secure multi-application smart card operating system. Lecture Notes in Computer Science, 1396:188–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koyama:1998:AFC

[1185] K. Koyama and R. Terada. An augmented family of cryptographic parity circuits. Lecture Notes in Computer Science, 1396:198–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1998:NBO

[1186] X. Yi, Kwok Yan Lam, Shi Xin Cheng, and Xiao Hu You. A new byteoriented block cipher. Lecture Notes in Computer Science, 1396:209–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellare:1998:POP

[1187] M. Bellare. Practice-oriented provablesecurity. Lecture Notes in Computer Science, 1396:221–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leiwo:1998:FMI

[1188] J. Leiwo and Y. Zheng. A framework for the management of information security. Lecture Notes in Computer Science, 1396:232–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kabasele-Tenday:1998:SSC

[1189] J.-M Kabasele-Tenday. Specifying security in a composite system. Lecture Notes in Computer Science, 1396:246– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1998:RSI

[1190] K. Zhang. On rough sets and inference analysis. Lecture Notes in Computer Science, 1396:256–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1998:AUS

[1191] T. Hwang and C.-H. Wang. Arbitrated unconditionally secure authentication scheme with multi-senders. Lecture Notes in Computer Science, 1396: 266–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1998:GSH

[1192] S. Kim, S. Park, and D. Won. Group signatures for hierarchical multigroups. Lecture Notes in Computer Science, 1396:273–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1998:TPS

[1193] K. Zhang. Threshold proxy signature schemes. Lecture Notes in Computer Science, 1396:282–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1998:SAE

[1194] Y. Zheng. Signcryption and its applications in efficient public key solutions. Lecture Notes in Computer Science, 1396:291–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1998:NDC

[1195] Khanh Quoc Nguyen, Y. Mu, and V. Varadharajan. A new digital cash scheme based on blind Nyberg-Rueppel digital signature. Lecture Notes in Computer Science, 1396:313–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchsberger:1998:IPM

[1196] A. Fuchsberger. An incremental payment method for Internet based streaming real-time media. Lecture Notes in Computer Science, 1396:321– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saeednia:1998:NIB

[1197] S. Saeednia and R. Safavi-Naini. A new identity-based key exchange protocol minimizing computation and communication. Lecture Notes in Computer Science, 1396:328–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakazaki:1998:AIB

[1198] H. Sakazaki, E. Okamoto, and M. Mambo. The application of ID-Based key distribution systems to an elliptic curve. Lecture Notes in Computer Science, 1396:335–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamazaki:1998:RDS

[1199] K. Yamazaki, M. Osaki, and O. Hirota. On reconciliation of discrepant sequences shared through quantum mechanical channels. Lecture Notes in Computer Science, 1396:345–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruijn:1998:PAC

[1200] N. G. de Bruijn. Philosophical aspects of computerized verification of mathematics. Lecture Notes in Computer Science, 1397:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970001.
pdf.

DeBruijn:1998:PAC

[1201] N. G. De Bruijn. Philosophical aspects of computerized verification of mathematics. Lecture Notes in Computer Science, 1397:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bundy:1998:SR

[1202] Alan Bundy. A science of reasoning.

Lecture Notes in Computer Science, 1397:10–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1397/13970010.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1397/13970010.
pdf.

Clarke:1998:MCH

[1203] Edmund M. Clarke and Sergey Berezin.

Model checking: Historical perspective and example. Lecture Notes in Computer Science, 1397:18–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970018.
pdf.

Balsiger:1998:CTP

[1204] Peter Balsiger and Alain Heuerding. Comparison of theorem provers for modal logics — introduction and summary. Lecture Notes in Computer Science, 1397:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970025.
pdf.

Horrocks:1998:FD

[1205] Ian Horrocks and Peter F. Patel-Schneider. FaCT and DLP. Lecture Notes in Computer Science, 1397:27– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970027.
pdf.

Levy:1998:PK

[1206] Michel Levy. Prover KT4. Lecture Notes in Computer Science, 1397:31– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970031.
pdf.

Beckert:1998:LK

[1207] Bernhard Beckert and Rajeev Goré. lean K 2.0. Lecture Notes in Computer Science, 1397:33–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1397/13970033.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1397/13970033.pdf.

Balsiger:1998:LW

[1208] Peter Balsiger, Alain Heuerding, and Stefan Schwendimann. Logics workbench 1.0. Lecture Notes in Computer Science, 1397:35–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970035.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970035.
pdf.

Hustadt:1998:OFT

[1209] Ullrich Hustadt, Renate A. Schmidt, and Christoph Weidenbach. Optimised functional translation and resolution. Lecture Notes in Computer Science, 1397:36–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1397/13970036.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1397/13970036.
pdf.

Pitt:1998:BE

[1210] Jeremy Pitt. Benchmark evaluation of □KE. Lecture Notes in Computer Science, 1397:38–39, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970038.
pdf.

Janssen:1998:IPT

[1211] G. L. J. M. Janssen. Implementation of propositional temporal logics using BDDs. Lecture Notes in Computer Science, 1397:40–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970040.
pdf.

Komara:1998:CPM

[1212] Ján Komara and Paul J. Voda. Computer programming as mathematics in a programming language and proof system CL. Lecture Notes in Computer Science, 1397:42–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970042.
pdf.

Baldoni:1998:TCM

[1213] Matteo Baldoni, Laura Giordano, and Alberto Martelli. A tableau calculus for multimodal logics and some (un)decidability results. Lecture Notes in Computer Science, 1397:44– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970044.
pdf.

Baumgartner:1998:HTN

[1214] Peter Baumgartner. Hyper Tableau — the next generation. Lecture Notes in Computer Science, 1397:60– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970060.
pdf.

Beckert:1998:FST

[1215] Bernhard Beckert and Dov Gabbay.

Fibring semantic tableaux. Lecture Notes in Computer Science, 1397:77– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970077.
pdf.

Beckert:1998:TCQ

[1216] Bernhard Beckert and Ulrike Hartmer. A tableau calculus for quantifierfree set theoretic formulae. Lecture Notes in Computer Science, 1397:93– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970093.
pdf.

Bowman:1998:TMI

[1217] Howard Bowman and Simon Thompson. A tableau method for interval temporal logic with projection. Lecture Notes in Computer Science, 1397: 108–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970108.
pdf.

Cerrito:1998:BMS

[1218] Serenella Cerrito and Marta Cialdea Mayer. Bounded model search in linear temporal logic and its application to planning. Lecture Notes in Computer Science, 1397:124–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970124.
pdf.

Egly:1998:PCC

[1219] Uwe Egly and Hans Tompits. On proof complexity of circumscription. Lecture Notes in Computer Science, 1397:141– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970141.
pdf.

Fermueller:1998:TFV

[1220] C. G. Fermueller and H. Langsteiner. Tableaux for finite-valued logics with arbitrary distribution modalities. Lecture Notes in Computer Science, 1397: 156–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fermuller:1998:TFV

[1221] Christian G. Fermüller and Herbert Langsteiner. Tableaux for finitevalued logics with arbitrary distribution modalities. Lecture Notes in Computer Science, 1397:156–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970156.
pdf.

Haehnle:1998:SRC

[1222] R. Haehnle, N. V. Murray, and E. Rosenthal. Some remarks on completeness, connection graph resolution and link deletion. Lecture Notes in Computer Science, 1397:172–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahnle:1998:SRC

[1223] Reiner Hähnle, Neil V. Murray, and Erik Rosenthal. Some remarks on completeness, connection graph resolution, and link deletion. Lecture Notes in Computer Science, 1397:172– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970172.
pdf.

Hustadt:1998:SBM

[1224] Ullrich Hustadt and Renate A. Schmidt. Simplification and backjumping in modal tableau. Lecture Notes in Computer Science, 1397:187– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1397/13970187.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1397/13970187.
                                        pdf.

Martin:1998:FVT

[1225] P. J. Martín, A. Gavilanes, and J. Leach. Free variable tableaux for a logic with term declarations. Lecture Notes in Computer Science, 1397: 202–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1397/13970202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1397/13970202.
                                        pdf.

Massacci:1998:SGC

[1226] Fabio Massacci. Simplification: A general constraint propagation technique for propositional and modal tableaux. Lecture Notes in Computer Science, 1397:217–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1397/13970217.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1397/13970217.
                                        pdf.

Monz:1998:TCA

[1227] Christof Monz and Maarten de Rijke. A tableaux calculus for ambiguous quantification. Lecture Notes in Computer Science, 1397:232–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970232.
pdf.

Negri:1998:KMA

[1228] Sara Negri and Jan von Plato.

From Kripke models to algebraic counter-valuations. Lecture Notes in Computer Science, 1397:247–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970247.
pdf.

Schmitt:1998:DRP

[1229] Stephan Schmitt and Christoph Kreitz. Deleting redundancy in proof reconstruction. Lecture Notes in Computer Science, 1397:262–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970262.
pdf.

Schwendimann:1998:NOP

[1230] Stefan Schwendimann. A new onepass tableau calculus for PLTL. Lecture Notes in Computer Science, 1397:277– 291, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1397/13970277.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1397/13970277.
                                        pdf.

Weich:1998:DPI

[1231] Klaus Weich. Decision procedures for intuitionistic propositional logic by program extraction. Lecture Notes in Computer Science, 1397:292–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1397/13970292.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1397/13970292.
                                        pdf.

Horrocks:1998:FS

[1232] Ian Horrocks. The FaCT system. Lecture Notes in Computer Science, 1397: 307–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1397/13970307.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1397/13970307.
                                        pdf.

Urban:1998:IPS

[1233] Christian Urban. Implementation of proof search in the imperative programming language pizza. Lecture Notes in Computer Science, 1397:313– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1397/13970313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1397/13970313.
pdf.

Wolf:1998:PSS

[1234] Andreas Wolf. P-SETHEO: Strategy parallelism in automated theorem proving. Lecture Notes in Computer Science, 1397:320–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1397/13970320.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1397/13970320.pdf.

Anonymous:1998:AIb

[1235] Anonymous. Author index. Lecture Notes in Computer Science, 1397:325– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1397/1397auth.pdf.

DeJong:1998:LAO

[1236] K. De Jong. Learning in agent-oriented worlds. Lecture Notes in Computer Science, 1398:3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewis:1998:NBF

[1237] D. D. Lewis. Naive (bayes) at forty:

The independence assumption in information retrieval. Lecture Notes in Computer Science, 1398:4–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marques:1998:LVT

[1238] N. M. Marques, G. Pereira Lopes, and C. A. Coelho. Learning verbal transitivity using LogLinear models. Lecture Notes in Computer Science, 1398:19–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marquez:1998:PST

[1239] L. Marquez and H. Rodriguez. Partof-speech tagging using decision trees. Lecture Notes in Computer Science, 1398:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coste:1998:IFA

[1240] F. Coste and J. Nicolas. Inference of finite automata: Reducing the search space with an ordering of pairs of states. Lecture Notes in Computer Science, 1398:37–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1998:AAL

[1241] R. Schneider. Automatic acquisition of lexical knowledge from sparse and noisy data. Lecture Notes in Computer Science, 1398:43–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Letourneau:1998:NMC

[1242] S. Letourneau, S. Matwin, and F. Famili. A normalization method for contextual data: Experience from a large-scale application. Lecture Notes in Computer Science, 1398:49– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sammut:1998:LCX

[1243] C. Sammut and T. Zrimec. Learning to classify X-ray images using relational learning. Lecture Notes in Computer Science, 1398:55–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzeroski:1998:IED

[1244] S. Dzeroski, N. Jacobs, M. Molina, and C. Moure. ILP experiments in detecting traffic problems. Lecture Notes in Computer Science, 1398:61–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neri:1998:SCL

[1245] F. Neri. Simulating children learning and explaining elementary heat transfer phenomena: A multistrategy system at work. Lecture Notes in Computer Science, 1398:67–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kontkanen:1998:BOI

[1246] P. Kontkanen, P. Myllymaeki, T. Silander, and H. Tirri. Bayes optimal instance-based learning. Lecture Notes in Computer Science, 1398:77– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kontkanen:1998:BIT

[1247] P. Kontkanen, P. Myllymaeki, T. Silander, and H. Tirri. Bayesian and information-theoretic priors for Bayesian network parameters. Lecture Notes in Computer Science, 1398:89– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mladenic:1998:FSS

[1248] D. Mladenic. Feature subset selection in text learning. Lecture Notes in Computer Science, 1398:95–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1998:MMO

[1249] H. Liu, H. Motoda, and M. Dash. A monotonic measure for optimal feature selection. Lecture Notes in Computer Science, 1398:101–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camacho:1998:IMH

[1250] R. Camacho. Inducing models of human control skills. Lecture Notes in Computer Science, 1398:107–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bensusan:1998:GDA

[1251] H. Bensusan. God doesn’t always shave with Occam’s razor — learning when and how to prune. Lecture Notes in Computer Science, 1398:119–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torgo:1998:EEP

[1252] L. Torgo. Error estimators for pruning regression trees. Lecture Notes in Computer Science, 1398:125–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradford:1998:PDT

[1253] J. P. Bradford, C. Kunz, R. Kohavi, and C. Brunk. Pruning decision trees with misclassification costs. Lecture Notes in Computer Science, 1398:131– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joachims:1998:TCS

[1254] T. Joachims. Text categorization with support vector machines: Learning with many relevant features. Lecture Notes in Computer Science, 1398:137– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossius:1998:SNA

[1255] R. Rossius, G. Zenker, A. Ittner, and W. Dilger. A short note about the application of polynomial kernels with fractional degree in support vector learning. Lecture Notes in Computer Science, 1398:143–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viswanathan:1998:CLU

[1256] M. Viswanathan and G. I. Webb. Classification learning using all rules. Lecture Notes in Computer Science, 1398: 149–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreira:1998:IPC

[1257] M. Moreira and E. Mayoraz. Improved pairwise coupling classification with correcting classifiers. Lecture Notes in Computer Science, 1398:160–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jelonek:1998:ESM

[1258] J. Jelonek and J. Stefanowski. Experiments on solving multiclass learning problems by n02-classifier. Lecture Notes in Computer Science, 1398:172– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gama:1998:CCC

[1259] J. Gama. Combining classifiers with constructive induction. Lecture Notes in Computer Science, 1398:178–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ting:1998:BTC

[1260] K. M. Ting and Z. Zheng. Boosting trees for cost-sensitive classifications. Lecture Notes in Computer Science, 1398:190–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1998:NBC

[1261] Z. Zheng. Naive Bayesian classifier committees. Lecture Notes in Computer Science, 1398:196–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kontkanen:1998:BCD

[1262] P. Kontkanen, P. Myllymaeki, T. Silander, and H. Tirri. Batch classifications with discrete finite mixtures. Lecture Notes in Computer Science, 1398:208–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmid:1998:IRP

[1263] U. Schmid and F. Wysotzki. Induction of recursive program schemes. Lecture Notes in Computer Science, 1398: 214–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bostroem:1998:PIL

[1264] H. Bostroem. Predicate invention and learning from positive examples only. Lecture Notes in Computer Science, 1398:226–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouthinon:1998:ILP

[1265] D. Bouthinon and H. Soldano. An inductive logic programming framework to learn a concept from ambiguous examples. Lecture Notes in Computer Science, 1398:238–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Craven:1998:FOL

[1266] M. Craven, S. Slattery, and K. Nigam.

First-order learning for Web mining. Lecture Notes in Computer Science, 1398:250–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroedl:1998:EBG

[1267] S. Schroedl. Explanation-based generalization in game playing: Quantitative results. Lecture Notes in Computer Science, 1398:256–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lachiche:1998:SCI

[1268] N. Lachiche and P. Marquis. Scope classification: An instance-based learning algorithm with a rule-based characterization. Lecture Notes in Computer Science, 1398:268–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ricci:1998:ECO

[1269] F. Ricci and D. W. Aha. Errorcorrecting output codes for local learners. Lecture Notes in Computer Science, 1398:280–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bontempi:1998:RLL

[1270] G. Bontempi, M. Birattari, and H. Bersini. Recursive lazy learning for modeling and control. Lecture Notes in Computer Science, 1398:292–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguifo:1998:ULB

[1271] E. Mephu Nguifo and P. Njiwoua. Using lattice-based framework as a tool for feature extraction. Lecture Notes in Computer Science, 1398:304–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FurtadoVasco:1998:DPR

[1272] J. J. Furtado Vasco. Determining property relevance in concept formation by computing correlation between properties. Lecture Notes in Computer Science, 1398:310–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talavera:1998:BSA

[1273] L. Talavera and J. Roure. A buffering strategy to avoid ordering effects in clustering. Lecture Notes in Computer Science, 1398:316–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anglano:1998:CDS

[1274] C. Anglano, A. Giordana, G. Lo Bello, and L. Saitta. Coevolutionary, distributed search for inducing concept descriptions. Lecture Notes in Computer Science, 1398:322–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ducoulombier:1998:CME

[1275] A. Ducoulombier and M. Sebag. Continuous mimetic evolution. Lecture Notes in Computer Science, 1398:334– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olsson:1998:HPG

[1276] B. Olsson. A host-parasite genetic algorithm for asymetric tasks. Lecture Notes in Computer Science, 1398:346– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiering:1998:SQL

[1277] M. Wiering and J. Schmidhuber.

Speeding up Q(lambda)-learning. Lecture Notes in Computer Science, 1398: 352–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giani:1998:QLR

[1278] A. Giani, A. Sticca, F. Baiardi, and A. Starita. Q-learning and redundancy reduction in classifier systems with internal state. Lecture Notes in Computer Science, 1398:364–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drummond:1998:CFS

[1279] C. Drummond. Composing functions to speed up reinforcement learning in a changing world. Lecture Notes in Computer Science, 1398:370–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Precup:1998:TRR

[1280] D. Precup, R. S. Sutton, and S. Singh. Theoretical results on reinforcement learning with temporally abstract options. Lecture Notes in Computer Science, 1398:382–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munos:1998:GCM

[1281] R. Munos. A general convergence method for reinforcement learning in the continuous case. Lecture Notes in Computer Science, 1398:394–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weijters:1998:INNa

[1282] T. Weijters, A. Van den Bosch, and J. Van den Herik. Interpretable neural networks with BP-SOM. Lecture Notes in Computer Science, 1398:406– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohamed:1998:CRM

[1283] M. H. Mohamed, T. Minamoto, and K. Niijima. Convergence rate of minimization learning for neural networks. Lecture Notes in Computer Science, 1398:412–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goralwalla:1998:OOF

[1284] Iqbal A. Goralwalla, M. Tamer Özsu, and Duane Szafron. An object-oriented framework for temporal data models. Lecture Notes in Computer Science, 1399:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990001.
pdf.

Bettini:1998:ASI

[1285] Claudio Bettini, X. Sean Wang, and Sushil Jajodia. An architecture for supporting interoperability among temporal databases. Lecture Notes in Computer Science, 1399:36– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990036.
pdf.

Etzion:1998:EUF

[1286] Opher Etzion, Avigdor Gal, and Arie Segev. Extended update functionality in temporal databases. Lecture Notes in Computer Science, 1399:56– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990056.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990056.
pdf.

Gal:1998:TMT

[1287] Avigdor Gal. On transaction management in temporal databases. Lecture Notes in Computer Science, 1399:96– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990096.
pdf.

Elmasri:1998:IOT

[1288] Ramez Elmasri and Jae Young Lee. Implementation options for time-series data. Lecture Notes in Computer Science, 1399:115–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990115.
pdf.

Tansel:1998:EPT

[1289] Abdullah Uz Tansel and Erkan Tin. Expressive power of temporal relational query languages and temporal completeness. Lecture Notes in Computer Science, 1399:129–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990129.
pdf.

Snodgrass:1998:TTS

[1290] Richard T. Snodgrass, Michael H.

Böhlen, Christian S. Jensen, and Andreas Steiner. Transitioning temporal support in TSQL2 to SQL3. Lecture Notes in Computer Science, 1399:150– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990150.
pdf.

Darwen:1998:VTT

[1291] Hugh Darwen. Valid time and transaction time proposals: Language design aspects. Lecture Notes in Computer Science, 1399:195–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990195.
pdf.

Toman:1998:PBT

[1292] David Toman. Point-based temporal extensions of SQL and their efficient implementation. Lecture Notes in Computer Science, 1399:211–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1399/13990211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1399/13990211.
                                        pdf.

Gadia:1998:ATD

[1293] Shashi K. Gadia. Applicability of temporal data models to query multilevel security databases: A case study. Lecture Notes in Computer Science, 1399: 238–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1399/13990238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1399/13990238.
                                        pdf.

Patankar:1998:ACB

[1294] Ajit K. Patankar and Arie Segev. An architecture and construction of a business event manager. Lecture Notes in Computer Science, 1399:257– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1399/13990257.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1399/13990257.
                                        pdf.

Berger:1998:DUP

[1295] Gideon Berger and Alexander Tuzhilin. Discovering unexpected patterns in temporal data using temporal logic. Lecture Notes in Computer Science, 1399:281–??, 1998. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1399/13990281.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1399/13990281.
pdf.

Sistla:1998:QUP

[1296] A. Prasad Sistla, Ouri Wolfson, Sam Chamberlain, and Son Dao. Querying the uncertain position of moving objects. Lecture Notes in Computer Science, 1399:310–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990310.
pdf.

Wu:1998:TDB

[1297] Yu Wu, Sushil Jajodia, and X. Sean Wang. Temporal database bibliography update. Lecture Notes in Computer Science, 1399:338–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990338.
pdf.

Jensen:1998:CGT

[1298] Christian S. Jensen, Curtis E. Dyreson, Michael Böhlen, James Clifford, Ramez Elmasri, Shashi K. Gadia, Fabio Grandi, Pat Hayes, Sushil Jajodia, Wolfgang Käfer, Nick Kline, Nikos Lorentzos, Yannis Mitsopoulos, Angelo Montanari, Daniel Nonen, Elisa Peressi, Barbara Pernici, John F. Roddick, Nandlal L. Sarda, Maria Rita Scalas, Arie Segev, Richard T. Snodgrass, Mike D. Soo, Abdullah Tansel, Paolo Tiberio, and Gio Wiederhold. The consensus glossary of temporal database concepts — February 1998 version. Lecture Notes in Computer Science, 1399:367– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990367.
pdf.

Bettini:1998:GTG

[1299] Claudio Bettini, Curtis E. Dyreson, William S. Evans, Richard T. Snodgrass, and X. Sean Wang. A glossary of time granularity concepts. Lecture Notes in Computer Science, 1399:406– 413, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1399/13990406.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990406.
pdf.

Researchers:1998:SCW

[1300] The Dagstuhl Seminar Researchers. Summaries of current work. Lecture Notes in Computer Science, 1399:414– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1399/13990414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1399/13990414.
pdf.

Richter:1998:I

[1301] Michael M. Richter. 1. introduction.

Lecture Notes in Computer Science, 1400:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1400/14000001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1400/14000001.
pdf.

Burkhard:1998:ESC

[1302] Hans-Dieter Burkhard. 2. extending some concepts of CBR — foundations of case retrieval nets. Lecture Notes in Computer Science, 1400:17– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1400/14000017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1400/14000017.
pdf.

Lenz:1998:DDS

[1303] Mario Lenz, Eric Auriol, and Michel Manago. 3. diagnosis and decision support. Lecture Notes in Computer Science, 1400:51–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1400/14000051.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1400/14000051.
pdf.

Wilke:1998:ISS

[1304] Wolfgang Wilke, Mario Lenz, and Stefan Wess. 4. intelligent sales support with CBR. Lecture Notes in Computer Science, 1400:91–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1400/14000091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1400/14000091.
pdf.

Lenz:1998:TC

[1305] Mario Lenz, André Hübner, and Mirjam Kunze. 5. textual CBR. Lecture Notes in Computer Science, 1400:115– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1400/14000115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1400/14000115.
pdf.

Wilke:1998:UCT

[1306] Wolfgang Wilke, Barry Smyth, and Pádraig Cunningham. 6. using configuration techniques for adaptation. Lecture Notes in Computer Science, 1400: 139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1400/14000139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1400/14000139.
pdf.

Bergmann:1998:CAP

[1307] Ralph Bergmann, Héctor Muñoz-Avila, Manuela Veloso, and Erica Melis. 7. CBR applied to planning. Lecture Notes in Computer Science, 1400:169– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1400/14000169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1400/14000169.
pdf.

Boerner:1998:CD

[1308] K. Boerner. CBR for design. Lecture Notes in Computer Science, 1400:201– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borner:1998:CD

[1309] Katy Börner. 8. CBR for design. Lecture Notes in Computer Science, 1400: 201–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1400/14000201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1400/14000201.
pdf.

Althoff:1998:CES

[1310] Klaus-Dieter Althoff, Andreas Birk, Christiane Gresse von Wangenheim, and Carsten Tautz. 9. CBR for experimental software engineering. Lecture Notes in Computer Science, 1400: 235–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1400/14000235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1400/14000235.
pdf.

Weber:1998:CTH

[1311] Gerhard Weber and Thomas J. Schult. 10. CBR for tutoring and help systems. Lecture Notes in Computer Science, 1400:255–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1400/14000255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1400/14000255.
pdf.

Gierl:1998:CM

[1312] Lothar Gierl, Mathias Bull, and Rainer Schmidt. 11. CBR in medicine. Lecture Notes in Computer Science, 1400: 273–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1400/14000273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1400/14000273.
pdf.

Bergmann:1998:MBC

[1313] Ralph Bergmann and Klaus-Dieter Althoff. 12. methodology for building CBR applications. Lecture Notes in Computer Science, 1400:299–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1400/14000299.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1400/14000299.
pdf.

Kamp:1998:RA

[1314] Gerd Kamp, Steffen Lange, and Christoph Globig. 13. related areas. Lecture Notes in Computer Science, 1400:327–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1400/14000327.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1400/14000327.
pdf.

Barberet:1998:TTW

[1315] C. Barberet, L. Brunie, F. Desprez, and G. Lebourgeois. Technology transfer within the ProHPC TTN at ENS Lyon. Lecture Notes in Computer Science, 1401:3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loeve:1998:CME

[1316] W. Loeve and M. E. S. Vogels. Competence management in engineering environments with HPCN. Lecture Notes in Computer Science, 1401:13–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groothuizen:1998:NHC

[1317] R. Groothuizen and H. Van der Ven.

A NICE HPCN centre for flow simulation. Lecture Notes in Computer Science, 1401:23–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stadler:1998:IAH

[1318] J. Stadler. Industrial applications of high performance computing — the experiences from HWW. Lecture Notes in Computer Science, 1401:30–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anguita:1998:HPN

[1319] D. Anguita, A. Boni, M. Chirico, and F. Giudici. High performance neurocomputing: Industrial and medical applications of the RAIN system. Lecture Notes in Computer Science, 1401: 34–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:1998:ISG

[1320] J. C. Gonzalez, J. Herrera, S. Garcia, and F. H. Priano. An intranet solution for geographically dispersed educational centres: PINCEL-NET. Lecture Notes in Computer Science, 1401:44–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marsh:1998:ECW

[1321] A. Marsh. EUROMED — A 21st century WWW-based telemedical information society. Lecture Notes in Computer Science, 1401:54–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grandinetti:1998:CTS

[1322] L. Grandinetti, D. Conforti, and L. De Luca. CAMD and TeleEEG: Software tools for telemedicine applications. Lecture Notes in Computer Science, 1401:64–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meliones:1998:DAV

[1323] A. Meliones, A. Karidis, S. Perrakis, and V. Siganos. Distributed audiovisual content development. Lecture Notes in Computer Science, 1401:74– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beccaria:1998:HPR

[1324] M. Beccaria, G. Buresti, A. Ciampa, and G. Curci. High performance roadvehicle optimized aerodynamic design: Application of parallel computing to car design. Lecture Notes in Computer Science, 1401:86–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derksen:1998:PST

[1325] J. Derksen and H. Van den Akker. Parallel simulation of turbulent fluid flow in a mixing tank. Lecture Notes in Computer Science, 1401:96–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palazzari:1998:SPA

[1326] P. Palazzari, P. D’Atanasio, and F. Ragusini. Simulation of patch array antennas through the implementation of finite-difference time-domain (FD-TD) algorithm on distributed memory massively parallel systems. Lecture Notes in Computer Science, 1401:105–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:1998:ADA

[1327] C. H. Q. Ding, P. M. Lyster, J. W. Larson, and J. Guo. Atmospheric data assimilation on distributed-memory parallel supercomputers. Lecture Notes in Computer Science, 1401:115–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oksuzoglu:1998:BGO

[1328] H. Öksüzoğlu and A. G. M. Van Hees. A barotropic global ocean model and its parallel implementation on unstructured grids. Lecture Notes in Computer Science, 1401:125–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hummel:1998:DES

[1329] S. Hummel, I. J. P. Elshoff, and A. E. Mynett. Distributed engineering systems in coastal zone management. Lecture Notes in Computer Science, 1401:133–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blom:1998:HAQ

[1330] J. G. Blom, W. M. Lioen, and J. G. Verwer. HPCN and air quality modeling. Lecture Notes in Computer Science, 1401:141–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicastro:1998:PIM

[1331] S. Nicastro and F. Valentinotti. Parallel implementation of a meteorological model on a SIMD architecture. Lecture Notes in Computer Science, 1401:151–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernaschi:1998:PSI

[1332] M. Bernaschi, F. Castiglione, and S. Succi. A parallel simulator of the immune response. Lecture Notes in Computer Science, 1401:163–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kandhai:1998:PLB

[1333] D. Kandhai, D. Dubbeldam, A. G.

Hoekstra, and P. M. A. Sloot. Parallel lattice-Boltzmann simulation of fluid flow in centrifugal elutriation chambers. Lecture Notes in Computer Science, 1401:173–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attig:1998:RCL

[1334] N. Attig, S. Guesken, P. Lacock, and T. Lippert. Running a code for lattice quantum chromodynamics efficiently on CRAY T3E systems. Lecture Notes in Computer Science, 1401: 183–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banse:1998:IBP

[1335] F. Banse, J.-L. Dekeyser, R. Fauquembergue, and F. Dessenne. Implementation of a bi-parallel Monte Carlo device simulation on two architectures. Lecture Notes in Computer Science, 1401: 193–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1998:PPS

[1336] B. Di Martino, S. Briguglio, G. Vlad, and P. Sguazzero. Parallel plasma simulation in High Performance Fortran. Lecture Notes in Computer Science, 1401:203–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogaccia:1998:PIL

[1337] G. Fogaccia. Parallel implementation of a lattice Boltzmann algorithm for the electrostatic plasma turbulence. Lecture Notes in Computer Science, 1401:213–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:1998:EHP

[1338] C. H. Q. Ding. Evaluations of HPF for practical scientific algorithms on T3E. Lecture Notes in Computer Science, 1401:223–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keppens:1998:DPS

[1339] R. Keppens, S. Poedts, and J. P. Goedbloed. Data parallel simulations of the magnetohydrodynamics of plasma loops. Lecture Notes in Computer Science, 1401:233–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderPloeg:1998:PBT

[1340] A. Van der Ploeg. Parallelization of a block tridiagonal solver in HPF on an IBM SP2. Lecture Notes in Computer Science, 1401:242–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthou:1998:PHM

[1341] J.-Y. Berthou and L. Plagne. Parallel HPF-MPI implementation of the TB-SCM Poisson solver. Lecture Notes in Computer Science, 1401:252–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meliones:1998:EDE

[1342] A. N. Meliones, I. Barosan, and T. T.

Varvarigou. Efficient distributed execution of computationally intensive tasks. Lecture Notes in Computer Science, 1401:262–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradi:1998:EDD

[1343] A. Corradi, L. Leonardi, and F. Zambonelli. On the effectiveness of different diffusive load balancing policies in dynamic applications. Lecture Notes in Computer Science, 1401:274–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baiardi:1998:LBL

[1344] F. Baiardi, P. Becuzzi, P. Mori, and M. Paoli. Load balancing and locality in hierarchical N-body algorithms on distributed memory architectures. Lecture Notes in Computer Science, 1401:284–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xirogiannis:1998:DTD

[1345] G. Xirogiannis and H. Taylor. A dynamic task distribution and engine allocation strategy for distributed execution of logic programs. Lecture Notes in Computer Science, 1401:294–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherkasova:1998:SSI

[1346] L. Cherkasova. Scheduling strategy to improve response time for Web applications. Lecture Notes in Computer Science, 1401:305–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piriyakumar:1998:NAB

[1347] D. A. L. Piriyakumar, C. S. R. Murthy, and P. Levi. A new A0* based optimal task scheduling in heterogeneous multiprocessor systems applied to computer vision. Lecture Notes in Computer Science, 1401:315–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andre:1998:MGP

[1348] F. Andre, C. Morin, and M.-T. Segarra. Mechanisms for global processor and memory management on a NoW. Lecture Notes in Computer Science, 1401: 324–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacLaren:1998:LLW

[1349] J. M. MacLaren and J. M. Bull. Lessons learned when comparing shared memory and message passing codes on three modern parallel architectures. Lecture Notes in Computer Science, 1401:337–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aharoni:1998:APD

[1350] D. Aharoni and A. Barak. Asynchronous parallel discontinuous finite element method. Lecture Notes in Computer Science, 1401:347–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitagawa:1998:CBS

[1351] K. Kitagawa, H. Nakamura, and G. Yagawa. Comparison between substructure method and domain decomposition method. Lecture Notes in Computer Science, 1401:358–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toth:1998:CDC

[1352] G. Toth and R. Keppens. Comparison of different computer platforms for running the versatile advection code. Lecture Notes in Computer Science, 1401:368–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waheed:1998:PNB

[1353] A. Waheed and J. Yan. Parallelization of NAS benchmarks for shared memory multiprocessors. Lecture Notes in Computer Science, 1401:377–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diard:1998:PDE

[1354] F. Diard. The Photon4D distributed engine: A distribution layer for high quality rendering in shared virtual immersion. Lecture Notes in Computer Science, 1401:387–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderVen:1998:RTV

[1355] H. Van der Ven, B. Schultheiss, S. Doi, and H. Matsumoto. Real-time visualization of large data sets on NLR’s NEC SX-4. Lecture Notes in Computer Science, 1401:397–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blasiak:1998:VCM

[1356] J. Blasiak and W. Dzwinel. Visual clustering of multidimensional and large data sets using parallel environments. Lecture Notes in Computer Science, 1401:403–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hodgson:1998:AHM

[1357] G. S. Hodgson, P. Dzwig, H. M. Liddell, and D. Parkinson. Application of HPC to medium-size stochastic systems with non-linear constraints in finance. Lecture Notes in Computer Science, 1401:411–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nowak:1998:PEM

[1358] U. Nowak, R. Ehrig, and L. Oeverdieck. Parallel extrapolation methods and their application in chemical engineering. Lecture Notes in Computer Science, 1401:419–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerry:1998:KIH

[1359] K. E. Kerry and K. A. Hawick. Kriging interpolation on high-performance computers. Lecture Notes in Computer Science, 1401:429–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lippert:1998:TAF

[1360] T. Lippert, K. Schilling, F. Toschi, and S. Trentmann. Transpose algorithm for FFT on APE/ quadrics. Lecture Notes in Computer Science, 1401:439– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kechadi:1998:PTP

[1361] M.-T. Kechadi and D. F. Hegarty.

A parallel technique for partitioning nodes of weighted graphs. Lecture Notes in Computer Science, 1401:449– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dodero:1998:SPA

[1362] G. Dodero, V. Gianuzzi, M. Moscati, and M. Corvi. A scalable parallel algorithm for matching pursuit signal decomposition. Lecture Notes in Computer Science, 1401:458–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peddemors:1998:HPD

[1363] A. J. H. Peddemors and L. O.

Hertzberger. A high performance distributed database system for enhanced Internet services. Lecture Notes in Computer Science, 1401:469–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zingirian:1998:ESS

[1364] N. Zingirian, M. Maresca, and S. Nalin.

Efficiency of standard software architectures for Java-based access to remote databases. Lecture Notes in Computer Science, 1401:479–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donato:1998:MMD

[1365] J. M. Donato, J. C. Schryver, G. C. Hinkel, and R. L. Schmoyer. Mining multi-dimensional data for decision support. Lecture Notes in Computer Science, 1401:489–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taniar:1998:HPO

[1366] D. Taniar and Y. Jiang. A high performance object-oriented distributed parallel database architecture. Lecture Notes in Computer Science, 1401:498– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taniar:1998:TID

[1367] D. Taniar. Towards an ideal data placement scheme for high performance object-oriented database systems. Lecture Notes in Computer Science, 1401: 508–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaugendre:1998:CSA

[1368] P. Beaugendre, T. Priol, G. Allion, and D. Delavaux. A client/server approach for HPC applications within a networking environment. Lecture Notes in Computer Science, 1401:518–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sander:1998:HPC

[1369] V. Sander, D. Erwin, and V. Huber. High-performance computer management based on Java. Lecture Notes in Computer Science, 1401:526–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1998:UME

[1370] I. Park and R. Eigenmann. URSA MAJOR: Exploring Web technology for design and evaluation of highperformance systems. Lecture Notes in Computer Science, 1401:535–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chattratichat:1998:SAD

[1371] J. Chattratichat, J. Darlington, Y. Guo, and S. Hedvall. A software architecture for deploying high performance solution on the Internet. Lecture Notes in Computer Science, 1401: 545–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1998:PDF

[1372] R. Williams and B. Sears. Paraflow: A data flow distributed data-computing system. Lecture Notes in Computer Science, 1401:556–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1998:SPS

[1373] M. Schulz. SISCI-Pthreads SMP-like programming on an SCI-cluster. Lecture Notes in Computer Science, 1401: 566–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herland:1998:CML

[1374] B. G. Herland, M. Eberl, and H. Hellwagner. A common messaging layer for MPI and PVM over SCI. Lecture Notes in Computer Science, 1401:576– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakajo:1998:DSM

[1375] H. Nakajo, H. Tanaka, Y. Nakanishi, and M. Kohata. Distributed shared-memory for a workstation cluster with a high speed serial interface. Lecture Notes in Computer Science, 1401:588–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hawick:1998:DDH

[1376] K. A. Hawick, H. A. James, C. J. Patten, and F. A. Vaughan. DISCWorld: A distributed high performance computing environment. Lecture Notes in Computer Science, 1401:598–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakada:1998:UMA

[1377] H. Nakada, H. Takagi, S. Matsuoka, and U. Nagashima. Utilizing the metaserver architecture in the ninf global computing system. Lecture Notes in Computer Science, 1401:607– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buehlmann:1998:SDI

[1378] B. Buehlmann and H. Bieri. Supercomputing at the desktop: An improved interface using Internet facilities. Lecture Notes in Computer Science, 1401:617–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Launay:1998:FPP

[1379] P. Launay and J.-L Pazat. A framework for parallel programming in Java. Lecture Notes in Computer Science, 1401:628–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1998:ENB

[1380] A. Aggarwal and D. Grunwald. Effects of network bandwidth on performance in software DSM systems. Lecture Notes in Computer Science, 1401: 638–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CortesGomez:1998:PCM

[1381] F. Cortes Gomez and R. Lueling. A parallel continuous media server for Internet environments. Lecture Notes in Computer Science, 1401:648–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cavendish:1998:CLC

[1382] D. Cavendish, A. Fei, M. Gerla, and R. Rom. On the construction of low cost multicast trees with bandwidth reservation. Lecture Notes in Computer Science, 1401:658–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carretero:1998:CPR

[1383] J. Carretero, J. No, S.-S. Park, and A. Choudhary. COMPASSION: A parallel I/O runtime system including chunking and compression for irregular applications. Lecture Notes in Computer Science, 1401:668–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maciel:1998:DGN

[1384] F. B. Maciel, N. Sagawa, and T. Tanaka. Dynamic gateways: A novel approach to improve networking performance and availability on parallel servers. Lecture Notes in Computer Science, 1401:678–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cremonesi:1998:PHM

[1385] P. Cremonesi, C. Gennaro, and R. Marega. I/O performance in hybrid MIMD + SIMD machines. Lecture Notes in Computer Science, 1401: 688–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Osawa:1998:LDR

[1386] N. Osawa and T. Yuba. Lazy and differential replication in a recoverable distributed shared memory system. Lecture Notes in Computer Science, 1401:698–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cunha:1998:FSP

[1387] J. C. Cunha, J. Lourenco, J. Vieira, and B. Moscao. A framework to support parallel and distributed debugging. Lecture Notes in Computer Science, 1401:708–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerbyson:1998:AES

[1388] D. J. Kerbyson, E. Papaefstathiou, and G. R. Nudd. Application execution steering using on-the-fly performance prediction. Lecture Notes in Computer Science, 1401:718–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kacsuk:1998:GGE

[1389] P. Kacsuk, G. Dozsa, T. Fadgyas, and R. Lovas. The GRED graphical editor for the GRADE parallel program development environment. Lecture Notes in Computer Science, 1401:728–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mirenkov:1998:MSI

[1390] N. Mirenkov, O. G. Monakhov, and O. J. Chunikhin. A multimedia system for the investigation of mapping algorithms. Lecture Notes in Computer Science, 1401:738–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laure:1998:IAA

[1391] E. Laure and B. Chapman. Interprocedural array alignment analysis. Lecture Notes in Computer Science, 1401:747–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chi:1998:DPN

[1392] C.-H. Chi and C. M. Cheung. Data prefetching for non-linear memory references. Lecture Notes in Computer Science, 1401:757–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kartawidjaja:1998:MEP

[1393] M. A. Kartawidjaja and A. G. Hoekstra. Memory efficiency of parallel programs and memory bounded speedup. Lecture Notes in Computer Science, 1401:766–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donaldson:1998:CPO

[1394] S. R. Donaldson, J. M. D. Hill, and D. B. Skillicorn. Communication performance optimisation requires minimising variance. Lecture Notes in Computer Science, 1401:776–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Migliardi:1998:VPA

[1395] M. Migliardi, P. Baglietto, and M. Maresca. Virtual parallelism allows relaxing the synchronization constraints of SIMD computing paradigm. Lecture Notes in Computer Science, 1401:784–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:1998:HHP

[1396] S. Benkner. HPF+: High Performance Fortran for advanced industrial applications. Lecture Notes in Computer Science, 1401:797–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calzarossa:1998:ICS

[1397] M. Calzarossa, L. Massari, A. Merlo, and M. Pantano. Integration of a compilation system and a performance tool: The HPF+ approach. Lecture Notes in Computer Science, 1401:809–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:1998:PIA

[1398] S. Benkner, K. Sanjari, V. Sipkova, and B. Velkov. Parallelizing irregular applications with the Vienna HPF+ compiler VFC. Lecture Notes in Computer Science, 1401:816–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merlin:1998:MDP

[1399] J. H. Merlin, S. B. Baden, S. J.

Fink, and B. M. Chapman. Multiple data parallelism with HPF and KeLP. Lecture Notes in Computer Science, 1401:828–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slootmaekers:1998:MGS

[1400] R. Slootmaekers, H. Van Wulpen, and W. Joosen. Modelling genetic search agents with a concurrent objectoriented language. Lecture Notes in Computer Science, 1401:843–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doroshenko:1998:CFEa

[1401] A. E. Doroshenko. Coordination facilities to enhance concurrency of race free parallel algorithms. Lecture Notes in Computer Science, 1401:854–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christen:1998:DLB

[1402] P. Christen. Dynamic load balancing within a parallel iterative linear system solver. Lecture Notes in Computer Science, 1401:857–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dury:1998:SSG

[1403] C. M. Dury, R. Knecht, and G. H. Ristow. Size segregation of granular materials in a 3D rotating drum. Lecture Notes in Computer Science, 1401:860– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pourquie:1998:DPV

[1404] M. Pourquie. LES and DNS on parallel and vector platforms. Lecture Notes in Computer Science, 1401:863–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wijnandts:1998:HLV

[1405] P. Wijnandts and M. E. S. Vogels. High-level versus low-level DO-loop parallelization: Results for one testcase of a multi-block solver on a shared memory parallel vector computer. Lecture Notes in Computer Science, 1401:866–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bukhanovsky:1998:NSS

[1406] A. V. Bukhanovsky and A. M. Samsonov. Numerical simulation of strain solitons in elastic wave guides using vector and massive parallel supercomputers. Lecture Notes in Computer Science, 1401:869–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botchev:1998:EMT

[1407] M. A. Botchev. Experiments with MRAI time stepping schemes on a distributed memory parallel environment. Lecture Notes in Computer Science, 1401:872–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chillemi:1998:MDU

[1408] G. Chillemi and N. Sanna. Molecular dynamics using p-threads. Lecture Notes in Computer Science, 1401:875– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plazek:1998:FEM

[1409] J. Plazek, K. Banas, and J. Kitowski.

Finite element message-passing/DSM simulation algorithm for parallel computers. Lecture Notes in Computer Science, 1401:878–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzwinel:1998:LSM

[1410] W. Dzwinel, W. Alda, J. Kitowski, and J. Moscinski. Large-scale molecular dynamics experiments on Cray T3E system. Lecture Notes in Computer Science, 1401:881–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciaccio:1998:PFF

[1411] G. Ciaccio, V. DiMartino, and P. Lanucara. Porting the flame front propagation problem on GAMMA. Lecture Notes in Computer Science, 1401:884– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garatani:1998:PFE

[1412] K. Garatani, H. Nakamura, and G. Yagawa. Parallel finite element structural analysis code using DDM. Lecture Notes in Computer Science, 1401: 887–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demuynck:1998:BSP

[1413] K. Demuynck, J. Broeckhove, and F. Arickx. Building a software platform for distributed VR applications. Lecture Notes in Computer Science, 1401: 890–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bogdanov:1998:DPD

[1414] A. V. Bogdanov, N. Y. Bykov, and G. A. Lukianov. Distributed and parallel direct simulation Monte Carlo of rarefied gas flows. Lecture Notes in Computer Science, 1401:893–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakellariou:1998:SDE

[1415] R. Sakellariou. Solving Diophantine equations on a network of workstations. Lecture Notes in Computer Science, 1401:896–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stramaglia:1998:ISP

[1416] S. Stramaglia, G. Pasquariello, L. Guerriero, and A. Distante. Interferometric SAR phase unwrapping by parallel tempering on a APE100/ quadrics supercomputer. Lecture Notes in Computer Science, 1401:898–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stankova:1998:NSM

[1417] E. N. Stankova and E. V. Zudilova. Numerical simulation by means of supercomputers. Lecture Notes in Computer Science, 1401:901–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakajima:1998:HSL

[1418] K. Nakajima, H. Nakamura, and H. Okuda. Highly stable localized ILU preconditioning for unstructured grids. Lecture Notes in Computer Science, 1401:904–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horoi:1998:EIL

[1419] M. Horoi and R. Enbody. Efficient implementation of a Lanczos eigenvalue solver on a Cray T3E-900. Lecture Notes in Computer Science, 1401:907– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maekinen:1998:PFD

[1420] R. A. E. Maekinen, T. Rossi, and J. Toivanen. A parallel fast direct solver with applications. Lecture Notes in Computer Science, 1401:910–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lingen:1998:SSF

[1421] F. J. Lingen. ScaFiEP: The scalable finite element package. Lecture Notes in Computer Science, 1401:913–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenough:1998:SPF

[1422] C. Greenough and R. Fowler. Software for partitioning finite element meshes. Lecture Notes in Computer Science, 1401:916–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bogdanov:1998:FPC

[1423] A. V. Bogdanov, A. Gevorkyan, and A. Grigoryan. First principle calculations of quantum chaos and its selforganisation in the framework of 1D model of random quantum reactive harmonic oscillator. Lecture Notes in Computer Science, 1401:919–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mondelli:1998:PGE

[1424] G. Mondelli, G. Castellano, G. Attolico, and C. Distante. Parallel genetic evolution of membership functions and rules for a fuzzy controller. Lecture Notes in Computer Science, 1401:922– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spirov:1998:TET

[1425] A. V. Spirov and A. S. Kadyrov. Transposon element technique applied to GA-based John Muir’s trail test. Lecture Notes in Computer Science, 1401: 925–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jung:1998:RCV

[1426] T. Jung. Remote control of virtual environments via low-bandwidth connections. Lecture Notes in Computer Science, 1401:929–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Folino:1998:PEM

[1427] G. Folino, G. Spezzano, and D. Talia. Performance evaluation and modelling of MPI communications on the Meiko CS-2. Lecture Notes in Computer Science, 1401:932–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taniar:1998:DPB

[1428] D. Taniar and J. Wenny Rahayu. Divide and partial broadcast method for parallel collection join queries. Lecture Notes in Computer Science, 1401:937– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sueur:1998:DDP

[1429] D. Sueur, J.-L. Dekeyser, and P. Marquet. DPFS: A data-parallel file system environment. Lecture Notes in Computer Science, 1401:940–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lefebvre:1998:VEH

[1430] C. Lefebvre. Visual edition of HPF mappings with HPF-Builder. Lecture Notes in Computer Science, 1401:943– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoang:1998:ABS

[1431] H. D. Hoang. Adaptive bandwidth scheduling in B-ISDN. Lecture Notes in Computer Science, 1401:946–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cunha:1998:DPT

[1432] J. C. Cunha, P. Medeiros, J. Lourenco, and V. Duarte. The DOTPAR project: Towards a framework supporting domain oriented tools for parallel and distributed processing. Lecture Notes in Computer Science, 1401:952–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doroshenko:1998:CFEb

[1433] A. E. Doroshenko. Coordination facilities to enhance concurrency of race free parallel algorithms. Lecture Notes in Computer Science, 1401:955–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peeters:1998:PSB

[1434] T. W. J. Peeters and P. P. J. Stroomer. Parallel simulation of a bluff-body-stabilized non-premixed syngas flame. Lecture Notes in Computer Science, 1401:958–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chi:1998:EDP

[1435] C.-H. Chi and S. C. Lau. Extending data prefetching to co-operative caching. Lecture Notes in Computer Science, 1401:961–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slota:1998:LGA

[1436] R. Slota and J. Moscinski. Lattice gas automata simulation on HP/ convex exemplar SPP1600. Lecture Notes in Computer Science, 1401:963–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1998:IRD

[1437] Y.-S. Chang and B.-Y. Kim. The improved RAID 5 with the disk cache using the load-balanced destage algorithm. Lecture Notes in Computer Science, 1401:966–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gheri:1998:MIO

[1438] G. Gheri and J. R. Mosig. The multilevel interactive objects approach to model and mesh large planar antennas and arrays: A Full-JAVA tool. Lecture Notes in Computer Science, 1401: 969–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosch:1998:HPS

[1439] H. Kosch and L. Boeszoermenyi. High performance sets. Lecture Notes in Computer Science, 1401:972–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laffitte:1998:ETC

[1440] A. Laffitte, R. Sakellariou, and J. R.

Gurd. Evaluation of two compilerbased approaches for the parallelisation of an MPEG-2 decoder. Lecture Notes in Computer Science, 1401:975– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1998:FOW

[1441] Charu C. Aggarwal, Joel L. Wolf, and Philip S. Yu. A framework for the optimizing of WWW advertising. Lecture Notes in Computer Science, 1402: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1402/14020001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1402/14020001.
pdf.

Barra:1998:SAC

[1442] Maria Barra, Giuseppe Cattaneo, Mario Izzo, Alberto Negro, and Vittorio Scarano. Symmetric adaptive customer modeling for electronic commerce in a distributed environment. Lecture Notes in Computer Science, 1402:11–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1402/14020011.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1402/14020011.
pdf.

Belegradek:1998:MSP

[1443] Bella Belegradek, Konstantinos Kalpakis, and Yelena Yesha. Maximizing seller’s profit for electronic commerce. Lecture Notes in Computer Science, 1402:26– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1402/14020026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1402/14020026.
pdf.

Rebel:1998:ADS

[1444] Thomas F. Rebel, Olaf Darge, and Wolfgang Koenig. Approaches of digital signature legislation. Lecture Notes in Computer Science, 1402:39– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1402/14020039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1402/14020039.
pdf.

Pfitzmann:1998:JBD

[1445] A. Pfitzmann, A. Schill, A. Westfeld, G. Wicke, G. Wolf, and J. Zöllner. A Java-based distributed platform for multilateral security. Lecture Notes in Computer Science, 1402:52– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1402/14020052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1402/14020052.
pdf.

Shamir:1998:BBA

[1446] Gadi Shamir, Michael Ben-Or, and Danny Dolev. BARTER: A Backbone ARchitecture for Trade of ElectRonic content. Lecture Notes in Computer Science, 1402:65–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1402/14020065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1402/14020065.
pdf.

Romao:1998:ABS

[1447] Artur Romão and Miguel Mira da Silva. An agent-based secure Internet payment system for mobile computing. Lecture Notes in Computer Science, 1402:80–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1402/14020080.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1402/14020080.
                                        pdf.

Franz:1998:PSM

[1448] Elke Franz, Anja Jerichow, and Guntram Wicke. A payment scheme for mixes providing anonymity. Lecture Notes in Computer Science, 1402:94– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1402/14020094.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1402/14020094.
                                        pdf.

Cunningham:1998:SRE

[1449] Jim Cunningham, Shamimabi Paurobally, Athanassios Diacakis, Lorenz Lorenzen, Gabriel Gross, and Stephen McConnell. Satisfying requirements for electronic commerce. Lecture Notes in Computer Science, 1402:109– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1402/14020109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1402/14020109.
                                        pdf.

Gallego:1998:DMB

[1450] Isabel Gallego, Jaime Delgado, and José J. Acebrón. Distributed models for brokerage on electronic commerce. Lecture Notes in Computer Science, 1402: 129–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1402/14020129.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1402/14020129.
pdf.

Andreoli:1998:DPD

[1451] Jean-Marc Andreoli and François Pacull. Distributed print on demand systems in the xpect framework. Lecture Notes in Computer Science, 1402:141– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1402/14020141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1402/14020141.
pdf.

Bichler:1998:OBC

[1452] Martin Bichler, Carrie Beam, and Arie Segev. OFFER: A brokercentered object framework for electronic requisitioning. Lecture Notes in Computer Science, 1402:154–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1402/14020154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1402/14020154.
pdf.

Weitz:1998:WMI

[1453] Wolfgang Weitz. Workflow modeling for Internet-based commerce: An approach based on high-level Petri nets. Lecture Notes in Computer Science, 1402:166–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1402/14020166.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1402/14020166.
                                        pdf.

Geppert:1998:MBW

[1454] Andreas Geppert, Markus Kradolfer, and Dimitrios Tombros. Marketbased workflow management. Lecture Notes in Computer Science, 1402:179– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1402/14020179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1402/14020179.
                                        pdf.

Papazoglou:1998:DIW

[1455] Mike P. Papazoglou, Manfred A. Jeusfeld, Hans Weigand, and Matthias Jarke. Distributed, interoperable workflow support for electronic commerce. Lecture Notes in Computer Science, 1402:192–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1402/14020192.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1402/14020192.
                                        pdf.

Zapf:1998:SRM

[1456] Michael Zapf, Helge Müller, and Kurt Geihs. Security requirements for mobile agents in electronic markets. Lecture Notes in Computer Science, 1402: 205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1402/14020205.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1402/14020205.
pdf.

Yi:1998:SIT

[1457] X. Yi, X. F. Wang, and K. Y. Lam. A secure intelligent trade agent system. Lecture Notes in Computer Science, 1402:218–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1402/14020218.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1402/14020218.
pdf.

Boger:1998:MOE

[1458] Marko Boger. Migrating objects in electronic commerce applications. Lecture Notes in Computer Science, 1402: 229–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1402/14020229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1402/14020229.
pdf.

Strasser:1998:PRA

[1459] Markus Straßer, Kurt Rothermel, and Christian Maihöfer. Providing reliable agents for electronic commerce. Lecture Notes in Computer Science, 1402: 241–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1402/14020241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1402/14020241.
pdf.

Shoup:1998:STC

[1460] Victor Shoup and Rosario Gennaro. Securing threshold cryptosystems against chosen ciphertext attack. Lecture Notes in Computer Science, 1403:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030001.
                                        pdf.

Young:1998:ARA

[1461] Adam Young and Moti Yung. Autorecoverable auto-certifiable cryptosystems. Lecture Notes in Computer Science, 1403:17–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1403/14030017.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1403/14030017.pdf.

Fujisaki:1998:PPS

[1462] Eiichiro Fujisaki and Tatsuaki Okamoto. A practical and provably secure scheme for publicly verifiable secret sharing and its applications. Lecture Notes in Computer Science, 1403:32– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030032.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030032.
                                        pdf.

Kunihiro:1998:ECN

[1463] Noboru Kunihiro and Kenji Koyama.

Equivalence of counting the number of points on elliptic curve over the ring Z_{n} and factoring n. Lecture Notes in Computer Science, 1403:47– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030047.
pdf.

Boneh:1998:BRM

[1464] Dan Boneh and Ramarathnam Venkatesan. Breaking RSA may not be equivalent to factoring (extended abstract). Lecture Notes in Computer Science, 1403:59–71, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030059.
pdf;   http://theory.stanford.edu/
~dabo/papers/no_rsa_red.ps.gz.

Maurer:1998:LBG

[1465] Ueli M. Maurer and Stefan Wolf.

Lower bounds on generic algorithms in groups. Lecture Notes in Computer Science, 1403:72–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030072.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1403/14030072.
                                        pdf.

Biryukov:1998:ICR

[1466] Alex Biryukov and Eyal Kushilevitz. Improved cryptanalysis of RC5. Lecture Notes in Computer Science, 1403: 85–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030085.
                                        pdf.

Biham:1998:CAX

[1467] Eli Biham and Lars R. Knudsen. Cryptanalysis of the ANSI X9.52 CBCM mode. Lecture Notes in Computer Science, 1403:100–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030100.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030100.
                                        pdf.

Hawkes:1998:DLW

[1468] Philip Hawkes. Differential-linear weak key classes of IDEA. Lecture Notes in Computer Science, 1403:112– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030112.
                                        pdf.

Blaze:1998:DPA

[1469] Matt Blaze, Gerrit Bleumer, and Martin Strauss. Divertible protocols and atomic proxy cryptography. Lecture Notes in Computer Science, 1403:127– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030127.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030127.
pdf.

Kurosawa:1998:OTT

[1470] Kaoru Kurosawa and Yvo Desmedt.

Optimum traitor tracing and asymmetric schemes. Lecture Notes in Computer Science, 1403:145–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030145.
pdf.

Jutla:1998:FSS

[1471] Charanjit S. Jutla. On finding small solutions of modular multivariate polynomial equations. Lecture Notes in Computer Science, 1403:158–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030158.
pdf.

Weber:1998:CDL

[1472] Damian Weber. Computing discrete logarithms with quadratic number rings. Lecture Notes in Computer Science, 1403:171–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1403/14030171.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1403/14030171.pdf.

Patarin:1998:IAI

[1473] Jacques Patarin, Louis Goubin, and Nicolas Courtois. Improved algorithms for isomorphisms of polynomials. Lecture Notes in Computer Science, 1403: 184–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030184.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030184.
                                        pdf.

Shamir:1998:VC

[1474] Adi Shamir. Visual cryptanalysis. Lecture Notes in Computer Science, 1403: 201–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030201.
                                        pdf.

Cohen:1998:HIE

[1475] Gérard D. Cohen, Antoine Lobstein, David Naccache, and Gilles Zémor. How to improve an exponentiation black-box. Lecture Notes in Computer Science, 1403:211–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030211.
pdf.

Boyko:1998:SDL

[1476] Victor Boyko, Marcus Peinado, and Ramarathnam Venkatesan. Speeding up discrete log and factoring based schemes via precomputations. Lecture Notes in Computer Science, 1403:221– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030221.
pdf.

Bellare:1998:FBV

[1477] Mihir Bellare, Juan A. Garay, and Tal Rabin. Fast batch verification for modular exponentiation and digital signatures. Lecture Notes in Computer Science, 1403:236–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030236.
pdf.

Blaze:1998:FTR

[1478] Matt Blaze, Joan Feigenbaum, and Moni Naor. A formal treatment of remotely keyed encryption (extended abstract). Lecture Notes in Computer Science, 1403:251–265, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030251.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030251.
                                        pdf.

Bellare:1998:LRB

[1479] Mihir Bellare, Ted Krovetz, and Phillip Rogaway. Luby-Rackoff backwards: Increasing security by making block ciphers non-invertible. Lecture Notes in Computer Science, 1403:266–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030266.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030266.
                                        pdf.

Jakubowski:1998:CSP

[1480] Mariusz H. Jakubowski and Ramarathnam Venkatesan. The chain & sum primitive and its applications to MACs and stream ciphers. Lecture Notes in Computer Science, 1403:281– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030281.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1403/14030281.
pdf.

Huehnlein:1998:CBN

[1481] D. Huehnlein, M. J. Jacobson, S. Paulus, and T. Takagi. A cryptosystem based on non-maximal imaginary quadratic orders with fast decryption. Lecture Notes in Computer Science, 1403:294–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huhnlein:1998:CBN

[1482] Detlef Hühnlein, Michael J. Jacobson, Jr., Sachar Paulus, and Tsuyoshi Takagi. A cryptosystem based on non-maximal imaginary quadratic orders with fast decryption. Lecture Notes in Computer Science, 1403:294– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030294.
pdf.

Okamoto:1998:NPK

[1483] Tatsuaki Okamoto and Shigenori Uchiyama. A new public-key cryptosystem as secure as factoring. Lecture Notes in Computer Science, 1403: 308–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030308.
pdf.

Hiltgen:1998:TBU

[1484] Alain P. Hiltgen. Towards a better understanding of one-wayness: Facing linear permutations. Lecture Notes in Computer Science, 1403:319–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030319.
pdf.

Simon:1998:FCO

[1485] Daniel R. Simon. Finding collisions on a one-way street: Can secure hash functions be based on general assumptions? Lecture Notes in Computer Science, 1403:334–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030334.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030334.
pdf.

Franklin:1998:SCM

[1486] Matthew Franklin and Rebecca N. Wright. Secure communication in minimal connectivity models. Lecture Notes in Computer Science, 1403:346– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030346.
pdf.

Cachin:1998:FOT

[1487] Christian Cachin. On the foundations of oblivious transfer. Lecture Notes in Computer Science, 1403:361– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030361.
pdf.

Beaver:1998:QBS

[1488] Donald Beaver and Avishai Wool.

Quorum-based secure multi-party computation. Lecture Notes in Computer Science, 1403:375–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1403/14030375.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1403/14030375.pdf.

Pointcheval:1998:SSB

[1489] David Pointcheval. Strengthened security for blind signatures. Lecture Notes in Computer Science, 1403:391– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030391.
pdf.

Michels:1998:GCS

[1490] Markus Michels and Markus Stadler.

Generic constructions for secure and efficient confirmer signature schemes (extended abstract). Lecture Notes in Computer Science, 1403:406–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030406.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030406.
pdf.

Poupard:1998:SAP

[1491] Guillaume Poupard and Jacques Stern. Security analysis of a practical “on the fly” authentication and signature generation. Lecture Notes in Computer Science, 1403:422–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030422.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030422.
pdf.

Abe:1998:UVM

[1492] Masayuki Abe. Universally verifiable mix-net with verification work independent of the number of mix-servers. Lecture Notes in Computer Science, 1403: 437–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030437.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030437.
pdf.

Jakobsson:1998:PM

[1493] Markus Jakobsson. A practical mix.

Lecture Notes in Computer Science, 1403:448–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1403/14030448.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1403/14030448.
pdf.

Carlet:1998:PCD

[1494] Claude Carlet. On the propagation criterion of degree l and order k. Lecture Notes in Computer Science, 1403: 462–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030462.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030462.
pdf.

Filiol:1998:HNB

[1495] Eric Filiol and Caroline Fontaine.

Highly nonlinear balanced Boolean functions with a good correlationimmunity. Lecture Notes in Computer Science, 1403:475–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1403/14030475.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1403/14030475.pdf.

Millan:1998:HDC

[1496] William Millan, Andrew Clark, and Ed Dawson. Heuristic design of cryptographically strong balanced Boolean functions. Lecture Notes in Computer Science, 1403:489–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030489.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030489.
                                        pdf.

Padro:1998:SSS

[1497] Carles Padró and Germán Sáez. Secret sharing schemes with bipartite access structure. Lecture Notes in Computer Science, 1403:500–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030500.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030500.
                                        pdf.

Luby:1998:CBB

[1498] Michael Luby and Jessica Staddon. Combinatorial bounds for broadcast encryption. Lecture Notes in Computer Science, 1403:512–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030512.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030512.
                                        pdf.

Safavi-Naini:1998:NRM

[1499] Reihaneh Safavi-Naini and Huaxiong Wang. New results on multi-receiver authentication codes. Lecture Notes in Computer Science, 1403:527–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030527.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030527.
pdf.

Coppersmith:1998:SIF

[1500] Don Coppersmith. Specialized integer factorization. Lecture Notes in Computer Science, 1403:542–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030542.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030542.
pdf.

Okamoto:1998:SIB

[1501] Tatsuaki Okamoto and Shigenori Uchiyama. Security of an identitybased cryptosystem and the related reductions. Lecture Notes in Computer Science, 1403:546–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1403/14030546.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1403/14030546.
pdf.

Chan:1998:ECE

[1502] Agnes Chan, Yair Frankel, and Yiannis Tsiounis. Easy come — easy go divisible cash. Lecture Notes in Computer Science, 1403:561–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030561.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030561.
                                        pdf.

Naor:1998:SEM

[1503] Moni Naor and Benny Pinkas. Secure and efficient metering. Lecture Notes in Computer Science, 1403:576– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030576.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030576.
                                        pdf.

Asokan:1998:OFE

[1504] N. Asokan, Victor Shoup, and Michael Waidner. Optimistic fair exchange of digital signatures (extended abstract). Lecture Notes in Computer Science, 1403:591–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1403/14030591.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1403/14030591.
                                        pdf.

Anonymous:1998:AIc

[1505] Anonymous. Author index. Lecture Notes in Computer Science, 1403:607– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1403/1403auth.pdf.

Klatzky:1998:AES

[1506] Roberta L. Klatzky. Allocentric and egocentric spatial representations: Definitions, distinctions, and interconnections. Lecture Notes in Computer Science, 1404:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040001.
pdf.

Schweizer:1998:RDE

[1507] Karin Schweizer, Theo Herrmann, Gabriele Janzen, and Steffi Katz. The route direction effect and its constraints. Lecture Notes in Computer Science, 1404:19–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1404/14040019.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1404/14040019.pdf.

Mecklenboeuker:1998:SIA

[1508] S. Mecklenboeuker, W. Wippich, M. Wageher, and J. E. Saathoff. Spatial information and actions. Lecture Notes in Computer Science, 1404:39– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mecklenbrauker:1998:SIA

[1509] Silvia Mecklenbräuker, Werner Wippich, Monika Wagener, and Jörg E. Saathoff. Spatial information and actions. Lecture Notes in Computer Science, 1404:39–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1404/14040039.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1404/14040039.pdf.

Gehrke:1998:IEF

[1510] Jörg Gehrke and Bernhard Hommel. The impact of exogenous factors on spatial coding in perception and memory. Lecture Notes in Computer Science, 1404:63–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040063.
pdf.

Rothkegel:1998:JSR

[1511] Rainer Rothkegel, Karl F. Wender, and Sabine Schumacher. Judging spatial relations from memory. Lecture Notes in Computer Science, 1404:79– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040079.
pdf.

Werner:1998:RBM

[1512] Steffen Werner, Christina Saade, and Gerd Lüer. Relations between the mental representation of extrapersonal space and spatial behavior. Lecture Notes in Computer Science, 1404:107– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040107.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040107.
pdf.

Eisenkolb:1998:RLP

[1513] A. Eisenkolb, A. Musto, K. Schill, D. Hernández, and W. Brauer. Representational levels for the perception of the courses of motion. Lecture Notes in Computer Science, 1404:129– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040129.
pdf.

Tversky:1998:HSS

[1514] Barbara Tversky and Paul U. Lee.

How space structures language. Lecture Notes in Computer Science, 1404: 157–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040157.
pdf.

Eschenbach:1998:SNS

[1515] Carola Eschenbach, Christopher Habel, Lars Kulik, and Annette Leßmöllmann. Shape nouns and shape concepts: A geometry for ‘corner’. Lecture Notes in Computer Science, 1404:177–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1404/14040177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1404/14040177.
                                        pdf.

Vorwerg:1998:TEC

[1516] Constanze Vorwerg and Gert Rickheit. Typicality effects in the categorization of spatial relations. Lecture Notes in Computer Science, 1404:203– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1404/14040203.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1404/14040203.
                                        pdf.

Zimmer:1998:ULE

[1517] Hubert D. Zimmer, Harry R. Speiser, Jörg Baus, Anselm Blocher, and Eva Stopp. The use of locative expressions in dependence of the spatial relation between target and reference object in two-dimensional layouts. Lecture Notes in Computer Science, 1404: 223–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1404/14040223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1404/14040223.
                                        pdf.

Claus:1998:RFS

[1518] Berry Claus, Klaus Eyferth, Carsten Gips, Robin Hörnig, Ute Schmid, Sylvia Wiebrock, and Fritz Wysotzki. Reference frames for spatial inference in text understanding. Lecture Notes in Computer Science, 1404:241– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040241.
pdf.

Knauff:1998:MMS

[1519] Markus Knauff, Reinhold Rauh, Christoph Schlieder, and Gerhard Strube. Mental models in spatial reasoning. Lecture Notes in Computer Science, 1404:267–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040267.
pdf.

Frank:1998:FMC

[1520] Andrew U. Frank. Formal models for cognition — taxonomy of spatial location description and frames of reference. Lecture Notes in Computer Science, 1404:293–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1404/14040293.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1404/14040293.pdf.

Berendt:1998:SRA

[1521] Bettina Berendt, Thomas Barkowsky, Christian Freksa, and Stephanie Kelter. Spatial representation with aspect maps. Lecture Notes in Computer Science, 1404:313–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1404/14040313.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1404/14040313.pdf.

Kuipers:1998:HQR

[1522] Benjamin Kuipers. A hierarchy of qualitative representations for space. Lecture Notes in Computer Science, 1404: 337–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1404/14040337.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1404/14040337.
                                        pdf.

Renz:1998:SRT

[1523] Jochen Renz and Bernhard Nebel. Spatial reasoning with topological information. Lecture Notes in Computer Science, 1404:351–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1404/14040351.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1404/14040351.
                                        pdf.

Krieg-Broeckner:1998:TSK

[1524] B. Krieg-Broeckner, T. Roefer, H.-O.

Carmesin, and R. Mueller. A taxonomy of spatial knowledge for navigation and its application to the bremen autonomous wheelchair. Lecture Notes in Computer Science, 1404:373–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krieg-Bruckner:1998:TSK

[1525] Bernd Krieg-Brückner, Thomas Röfer, Hans-Otto Carmesin, and Rolf Müller. A taxonomy of spatial knowledge for navigation and its application to the Bremen Autonomous Wheelchair. Lecture Notes in Computer Science, 1404: 373–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040373.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040373.
pdf.

Nadel:1998:HPL

[1526] L. Nadel, K. G. F. Thomas, H. E. Laurance, R. Skelton, T. Tal, and W. J. Jacobs. Human place learning in a computer generated arena. Lecture Notes in Computer Science, 1404:399– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040399.
pdf.

Braunl:1998:CGS

[1527] Thomas Bräunl. CIIPS glory soccer robots with local intelligence. Lecture Notes in Computer Science, 1404:416– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040416.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040416.
pdf.

Leplow:1998:SOS

[1528] Bernd Leplow, Doris Höll, Lingju Zeng, and Maximilian Mehdorn. Spatial orientation and spatial memory within a ‘locomotor maze’ for humans. Lecture Notes in Computer Science, 1404:429– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040429.
pdf.

Mallot:1998:BES

[1529] Hanspeter A. Mallot, Sabine Gillner, Hendrik A. H. C. van Veen, and Heinrich H. Bülthoff. Behavioral experiments in spatial cognition using virtual reality. Lecture Notes in Computer Science, 1404:447– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040447.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040447.
pdf.

Wartenberg:1998:SOV

[1530] Fredrik Wartenberg, Mark May, and Patrick Péruch. Spatial orientation in virtual environments: Background considerations and experiments. Lecture Notes in Computer Science, 1404:469– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1404/14040469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1404/14040469.
pdf.

Steiert:1998:JAP

[1531] Hans-Peter Steiert and Jürgen Zimmermann. JPMQ — an advanced persistent message queuing service. Lecture Notes in Computer Science, 1405:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050001.
pdf.

Griebel:1998:RST

[1532] Gerrit Griebel, Brian Lings, and Björn Lundell. A repository to support transparency in database design. Lecture Notes in Computer Science, 1405:19– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050019.
pdf.

Preuner:1998:OCI

[1533] Günter Preuner and Michael Schrefl. Observation consistent integration of views of object life-cycles. Lecture Notes in Computer Science, 1405:32– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1405/14050032.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1405/14050032.
                                        pdf.

Omiecinski:1998:EMA

[1534] Edward Omiecinski and Ashok Savasere. Efficient mining of association rules in large dynamic databases. Lecture Notes in Computer Science, 1405:49– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1405/14050049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1405/14050049.
                                        pdf.

Kurniawati:1998:ENN

[1535] Ruth Kurniawati, Jesse S. Jin, and John A. Shepherd. Efficient nearestneighbour searches using weighted Euclidean metrics. Lecture Notes in Computer Science, 1405:64–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1405/14050064.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1405/14050064.
                                        pdf.

Nes:1998:AAQ

[1536] Niels Nes and Martin Kersten. The acoi algebra: A query algebra for image retrieval systems. Lecture Notes in Computer Science, 1405:77– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050077.
pdf.

Speegle:1998:MSS

[1537] Greg Speegle, Xiaojun Wang, and Le Gruenwald. A meta-structure for supporting multimedia editing in object-oriented databases. Lecture Notes in Computer Science, 1405:89– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050089.
pdf.

Karunaratna:1998:EKB

[1538] D. D. Karunaratna, W. A. Gray, and N. J. Fiddian. Establishing a knowledge base to assist integration of heterogeneous databases. Lecture Notes in Computer Science, 1405:103– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050103.
pdf.

Conrad:1998:CIC

[1539] Stefan Conrad, Ingo Schmitt, and Can Türker. Considering integrity constraints during federated database design. Lecture Notes in Computer Science, 1405:119–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1405/14050119.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1405/14050119.pdf.

Meredith:1998:SRT

[1540] P. F. Meredith and P. J. H. King. Scoped referential transparency in a functional database language with updates. Lecture Notes in Computer Science, 1405:134–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1405/14050134.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1405/14050134.
                                        pdf.

Paton:1998:EOA

[1541] Norman W. Paton and Pedro R. Falcone Sampaio. Extending the ODMG architecture with a deductive object query language. Lecture Notes in Computer Science, 1405:149–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1405/14050149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1405/14050149.
                                        pdf.

Brisaboa:1998:CDP

[1542] Nieves R. Brisaboa, Agustin González, Héctor J. Hernández, and José R. Paramá. The chase of Datalog programs. Lecture Notes in Computer Science, 1405:165–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050165.
pdf.

Alechina:1998:DQS

[1543] Natasha Alechina and Maarten de Rijke. Describing and querying semistructured data: Some expressiveness results. Lecture Notes in Computer Science, 1405:167–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050167.
pdf.

Welz:1998:TDS

[1544] Marc Welz and Peter Wood. The TENTACLE database system as a Web server. Lecture Notes in Computer Science, 1405:169–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050169.
pdf.

Hodge:1998:TFQ

[1545] L. E. Hodge, W. A. Gray, and N. J. Fiddian. A toolkit to facilitate the querying and integration of tabular data from semistructured documents. Lecture Notes in Computer Science, 1405: 171–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1405/14050171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1405/14050171.
                                        pdf.

Taniar:1998:PSC

[1546] David Taniar and J. Wenny Rahayu. Parallel sub-collection join algorithm for high performance objectoriented databases. Lecture Notes in Computer Science, 1405:173–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1405/14050173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1405/14050173.
                                        pdf.

McCann:1998:CDA

[1547] J. A. McCann and J. S. Crane. Component DBMS architecture for nomadic computing. Lecture Notes in Computer Science, 1405:175–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1405/14050175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1405/14050175.
                                        pdf.

Behrendt:1998:IDI

[1548] W. Behrendt, N. J. Fiddian, W. A.

Gray, and A. P. Madurapperuma. ITSE database interoperation toolkit. Lecture Notes in Computer Science, 1405: 177–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050177.
pdf.

Tawil:1998:OCM

[1549] A.-R. H. Tawil, W. A. Gray, and N. J. Fiddian. Ontological commitments for multiple view cooperation in a distributed heterogeneous environment. Lecture Notes in Computer Science, 1405:179–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050179.
pdf.

Hui:1998:CDF

[1550] Kit ying Hui and Peter M. D. Gray.

Constraint and data fusion in a distributed information system. Lecture Notes in Computer Science, 1405:181– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1405/14050181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1405/14050181.
pdf.

Anonymous:1998:AId

[1551] Anonymous. Author index. Lecture Notes in Computer Science, 1405:183– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1405/14050183.pdf.

Heyden:1998:CFM

[1552] Anders Heyden. A common framework for multiple view tensors. Lecture Notes in Computer Science, 1406: 3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060003.
                                        pdf.

Hartley:1998:CQT

[1553] Richard I. Hartley. Computation of the quadrifocal tensor. Lecture Notes in Computer Science, 1406:20– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060020.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060020.
                                        pdf.

Faugeras:1998:SCP

[1554] Olivier Faugeras, Long Quan, and Peter Sturm. Self-calibration of a 1D projective camera and its application to the self-calibration of a 2D projective camera. Lecture Notes in Computer Science, 1406:36– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060036.
pdf.

Koch:1998:MVS

[1555] Reinhard Koch, Marc Pollefeys, and Luc Van Gool. Multi viewpoint stereo from uncalibrated video sequences. Lecture Notes in Computer Science, 1406:55–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1406/14060055.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1406/14060055.
pdf.

Berthilsson:1998:RPP

[1556] Rikard Berthilsson and Anders Heyden. Recognition of planar point configurations using the density of affine shape. Lecture Notes in Computer Science, 1406:72–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060072.
pdf.

Triggs:1998:APS

[1557] Bill Triggs. Autocalibration from planar scenes. Lecture Notes in Computer Science, 1406:89–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1406/14060089.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1406/14060089.pdf.

Papadopoulo:1998:NCT

[1558] Thèodore Papadopoulo and Olivier Faugeras. A new characterization of the trifocal tensor. Lecture Notes in Computer Science, 1406:109–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060109.
                                        pdf.

Avidan:1998:TFM

[1559] Shai Avidan and Amnon Shashua. Threading fundamental matrices. Lecture Notes in Computer Science, 1406: 124–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060124.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060124.
                                        pdf.

Bretzner:1998:UYH

[1560] Lars Bretzner and Tony Lindeberg. Use your hand as a 3D mouse, or, relative orientation from extended sequences of sparse point and line correspondences using the affine trifocal tensor. Lecture Notes in Computer Science, 1406: 141–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1406/14060141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060141.
pdf.

Lavest:1998:DWR

[1561] Jean-Marc Lavest, Marc Viala, and M. Dhome. Do we really need an accurate calibration pattern to achieve a reliable camera calibration? Lecture Notes in Computer Science, 1406:158– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060158.
pdf.

Ohta:1998:OET

[1562] Naoya Ohta and Kenichi Kanatani.

Optimal estimation of three-dimensional rotation and reliability evaluation. Lecture Notes in Computer Science, 1406: 175–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060175.
pdf.

Fua:1998:RIA

[1563] P. Fua and C. Miccio. From regular images to animated heads: A least squares approach. Lecture Notes in Computer Science, 1406:188–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060188.
                                        pdf.

Schreer:1998:SVB

[1564] Oliver Schreer. Stereo visionbased navigation in unknown indoor environment. Lecture Notes in Computer Science, 1406:203–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060203.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060203.
                                        pdf.

Svoboda:1998:EGP

[1565] Tomás Svoboda, Tomás Pajdla, and Václav Hlavác. Epipolar geometry for panoramic cameras. Lecture Notes in Computer Science, 1406:218– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060218.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060218.
                                        pdf.

Ishikawa:1998:ODE

[1566] Hiroshi Ishikawa and Davi Geiger. Occlusions, discontinuities, and epipolar lines in stereo. Lecture Notes in Computer Science, 1406:232–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1406/14060232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060232.
pdf.

Carlsson:1998:SP

[1567] Stefan Carlsson. Symmetry in perspective. Lecture Notes in Computer Science, 1406:249–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1406/14060249.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1406/14060249.pdf.

Startchik:1998:PII

[1568] Sergei Startchik, Ruggero Milanese, and Thierry Pun. Projective and illumination invariant representation of disjoint shapes. Lecture Notes in Computer Science, 1406:264–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060264.
pdf.

Brooks:1998:RTE

[1569] Michael J. Brooks, Wojciech Chojnacki, Anton van den Hengel, and Luis Baumela. Robust techniques for the estimation of structure from motion in the uncalibrated case. Lecture Notes in Computer Science, 1406:281– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1406/14060281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060281.
                                        pdf.

Ueshiba:1998:FMP

[1570] Toshio Ueshiba and Fumiaki Tomita. A factorization method for projective and Euclidean reconstruction from multiple perspective views via iterative depth estimation. Lecture Notes in Computer Science, 1406:296–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060296.
                                        pdf.

Fitzgibbon:1998:ACR

[1571] Andrew W. Fitzgibbon and Andrew Zisserman. Automatic camera recovery for closed or open image sequences. Lecture Notes in Computer Science, 1406:311–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060311.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060311.
                                        pdf.

Kahl:1998:SMP

[1572] Fredrik Kahl and Anders Heyden. Structure and motion from points, lines and conics with affine cameras. Lecture Notes in Computer Science, 1406: 327–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060327.
pdf.

Brodsky:1998:SEV

[1573] Tomás Brodský, Cornelia Fermüller, and Yiannis Aloimonos. Simultaneous estimation of viewing geometry and structure. Lecture Notes in Computer Science, 1406:342–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060342.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060342.
pdf.

Fermueller:1998:WCS

[1574] C. Fermueller and Y. Aloimonos. What is computed by structure from motion algorithms? Lecture Notes in Computer Science, 1406:359–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fermuller:1998:WCS

[1575] Cornelia Fermüller and Yiannis Aloimonos. What is computed by structure from motion algorithms? Lecture Notes in Computer Science, 1406:359– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060359.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060359.
pdf.

Faugeras:1998:CDS

[1576] Olivier Faugeras and Renaud Keriven. Complete dense stereovision using level set methods. Lecture Notes in Computer Science, 1406:379–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060379.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060379.
pdf.

Schmid:1998:GMC

[1577] Cordelia Schmid and Andrew Zisserman. The geometry and matching of curves in multiple views. Lecture Notes in Computer Science, 1406:394– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060394.
pdf.

Moons:1998:AMR

[1578] Theo Moons, David Frère, Jan Vandekerckhove, and Luc Van Gool. Automatic modelling and 3D reconstruction of urban house roofs from high resolution aerial imagery. Lecture Notes in Computer Science, 1406:410– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060410.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060410.
pdf.

Csurka:1998:CFS

[1579] G. Csurka, D. Demirdjian, A. Ruf, and R. Horaud. Closed-form solutions for the Euclidean calibration of a stereo rig. Lecture Notes in Computer Science, 1406:426–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060426.
pdf.

Funt:1998:MCC

[1580] Brian Funt, Kobus Barnard, and Lindsay Martin. Is machine colour constancy good enough? Lecture Notes in Computer Science, 1406:445–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060445.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060445.
pdf.

Raja:1998:CMS

[1581] Yogesh Raja, Stephen J. McKenna, and Shaogang Gong. Colour model selection and adaptation in dynamic scenes. Lecture Notes in Computer Science, 1406:460–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1406/14060460.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1406/14060460.pdf.

Finlayson:1998:CCI

[1582] Graham D. Finlayson, Bernt Schiele, and James L. Crowley. Comprehensive colour image normalization. Lecture Notes in Computer Science, 1406: 475–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060475.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060475.
                                        pdf.

Kliot:1998:IBS

[1583] Michael Kliot and Ehud Rivlin. Invariant-based shape retrieval in pictorial databases. Lecture Notes in Computer Science, 1406:491–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060491.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060491.
                                        pdf.

Torr:1998:CBM

[1584] P. H. S. Torr and A. Zisserman. Concerning Bayesian motion segmentation, model averaging, matching and the trifocal tensor. Lecture Notes in Computer Science, 1406:511–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060511.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060511.
                                        pdf.

Shi:1998:SIR

[1585] Jianbo Shi and Jitendra Malik. Self inducing relational distance and its application to image segmentation. Lecture Notes in Computer Science, 1406: 528–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060528.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060528.
pdf.

Leung:1998:CCR

[1586] Thomas Leung and Jitendra Malik.

Contour continuity in region based image segmentation. Lecture Notes in Computer Science, 1406:544–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060544.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060544.
pdf.

Ketterer:1998:SQC

[1587] Jens Ketterer, Jan Puzicha, Marcus Held, Martin Fischer, Joachim M. Buhmann, and Dieter Fellner. On spatial quantization of color images. Lecture Notes in Computer Science, 1406:563– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060563.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060563.
pdf.

Marie-Julie:1998:UIM

[1588] Jean Michel Marie-Julie and Hassane Essafi. Using ifs and moments to build a quasi invariant image index. Lecture Notes in Computer Science, 1406:578– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060578.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060578.
                                        pdf.

Gelgon:1998:DSS

[1589] Marc Gelgon and Patrick Bouthemy. Determining a structured spatiotemporal representation of video content for efficient visualization and indexing. Lecture Notes in Computer Science, 1406:595–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1406/14060595.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1406/14060595.pdf.

Kimmel:1998:DIR

[1590] Ron Kimmel. Demosaicing: Image reconstruction from color CCD samples. Lecture Notes in Computer Science, 1406:610–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1406/14060610.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060610.
                                        pdf.

Petrou:1998:PSS

[1591] M. Petrou, M. Mirmehdi, and M. Coors. Perceptual smoothing and segmentation of colour textures. Lecture Notes in Computer Science, 1406: 623–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060623.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060623.
pdf.

ColindeVerdiere:1998:VRU

[1592] Vincent Colin de Verdière and James L.

Crowley. Visual recognition using local appearance. Lecture Notes in Computer Science, 1406:640–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060640.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060640.
pdf.

DeVerdiere:1998:VRU

[1593] V. C. De Verdiere and J. L. Crowley. Visual recognition using local appearance. Lecture Notes in Computer Science, 1406:640–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verdiere:1998:VRU

[1594] Vincent Colin de Verdière and James L.

Crowley. Visual recognition using local appearance. Lecture Notes in Computer Science, 1406:640–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060640.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060640.
pdf.

Perona:1998:FAG

[1595] P. Perona and W. Freeman. A factorization approach to grouping. Lecture Notes in Computer Science, 1406:655– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060655.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060655.
pdf.

Lukacs:1998:FLS

[1596] Gabor Lukács, Ralph Martin, and Dave Marshall. Faithful least-squares fitting of spheres, cylinders, cones and tori for reliable segmentation. Lecture Notes in Computer Science, 1406:671– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060671.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060671.
pdf.

Frantz:1998:MSP

[1597] Sönke Frantz, Karl Rohr, and H. Siegfried Stiehl. Multi-step procedures for the localization of 2D and 3D point landmarks and automatic ROI size selection. Lecture Notes in Computer Science, 1406:687– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060687.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060687.
pdf.

Rehrmann:1998:OOM

[1598] Volker Rehrmann. Object oriented motion estimation in color image sequences. Lecture Notes in Computer Science, 1406:704–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060704.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060704.
pdf.

Mayer:1998:MSS

[1599] Helmut Mayer, Ivan Laptev, and Albert Baumgartner. Multi-scale and snakes for automatic road extraction. Lecture Notes in Computer Science, 1406:720–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1406/14060720.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1406/14060720.
pdf.

Braeutigam:1998:MFV

[1600] C. G. Braeutigam, J.-O. Eklundh, and H. I. Christensen. A model-free voting approach for integrating multiple cues. Lecture Notes in Computer Science, 1406:734–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brautigam:1998:MFV

[1601] Carsten G. Bräutigam, Jan-Olof Eklundh, and Henrik I. Christensen. A model-free voting approach for integrating multiple cues. Lecture Notes in Computer Science, 1406:734–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060734.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060734.
                                        pdf.

Belongie:1998:FBN

[1602] Serge Belongie and Jitendra Malik. Finding boundaries in natural images: A new method using point descriptors and area completion. Lecture Notes in Computer Science, 1406:751– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060751.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060751.
                                        pdf.

Isard:1998:SFC

[1603] Michael Isard and Andrew Blake. A smoothing filter for CONDENSA-TION. Lecture Notes in Computer Science, 1406:767–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1406/14060767.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
1406/14060767.pdf.

Munich:1998:CBI

[1604] Mario E. Munich and Pietro Perona. Camera-based ID verification by signature tracking. Lecture Notes in Computer Science, 1406:782–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060782.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060782.
pdf.

Southall:1998:MBT

[1605] B. Southall, J. A. Marchant, T. Hague, and B. F. Buxton. Model based tracking for navigation and segmentation. Lecture Notes in Computer Science, 1406:797–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1406/14060797.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1406/14060797.
pdf.

Haag:1998:BTL

[1606] Michael Haag and Hans-Hellmut Nagel.

Beginning a transition from a local to a more global point of view in model-based vehicle tracking. Lecture Notes in Computer Science, 1406:812– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060812.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1406/14060812.
                                        pdf.

delaTorre:1998:VBA

[1607] Fernando de la Torre, Shaogang Gong, and Stephen McKenna. View-based adaptive affine tracking. Lecture Notes in Computer Science, 1406:828–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060828.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060828.
                                        pdf.

Torre:1998:VBA

[1608] Fernando de la Torre, Shaogang Gong, and Stephen McKenna. View-based adaptive affine tracking. Lecture Notes in Computer Science, 1406:828–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060828.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060828.
                                        pdf.

Denzler:1998:ECS

[1609] J. Denzler, B. Heigl, and H. Niemann. An efficient combination of 2D and 3D shape descriptions for contour based tracking of moving objects. Lecture Notes in Computer Science, 1406:843– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060843.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1406/14060843.
pdf.

Huwer:1998:OTB

[1610] S. Huwer and H. Niemann. 2Dobject tracking based on projectionhistograms. Lecture Notes in Computer Science, 1406:861–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1406/14060861.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1406/14060861.pdf.

Haritaoglu:1998:WRT

[1611] Ismail Haritaoglu, David Harwood, and Larry S. Davis. W^{4}S: A realtime system for detecting and tracking people in 21/2D. Lecture Notes in Computer Science, 1406:877–892, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060877.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1406/14060877.
pdf.

Isard:1998:IUL

[1612] Michael Isard and Andrew Blake.

ICONDENSATION: Unifying lowlevel and high-level tracking in a stochastic framework. Lecture Notes in Computer Science, 1406:893–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1406/14060893.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1406/14060893.
                                        pdf.

Black:1998:PFM

[1613] Michael J. Black and Allan D. Jepson. A probabilistic framework for matching temporal trajectories: CONDENSATION-based recognition of gestures and expressions. Lecture Notes in Computer Science, 1406:909– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1406/14060909.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1406/14060909.
                                        pdf.

Anonymous:1998:AIe

[1614] Anonymous. Author index. Lecture Notes in Computer Science, 1406:925– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1406/14060925.pdf.

Pelillo:1998:MHS

[1615] Marcello Pelillo, Kaleem Siddiqi, and Steven W. Zucker. Matching hierarchical structures using association graphs. Lecture Notes in Computer Science, 1407:3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070003.
                                        pdf.

Trapp:1998:SMI

[1616] Ralph Trapp, Siegbert Drüe, and Georg Hartmann. Stereo matching with implicit detection of occlusions. Lecture Notes in Computer Science, 1407:17–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1407/14070017.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1407/14070017.
pdf.

Benjemaa:1998:SRM

[1617] Raouf Benjemaa and Francis Schmitt.

A solution for the registration of multiple 3D point sets using unit quaternions. Lecture Notes in Computer Science, 1407:34–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1407/14070034.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1407/14070034.pdf.

Nikou:1998:RRD

[1618] Christophoros Nikou, Fabrice Heitz, and Jean-Paul Armspach. Robust registration of dissimilar single and multimodal images. Lecture Notes in Computer Science, 1407:51–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070051.
pdf.

Vernon:1998:DFC

[1619] David Vernon. Decoupling Fourier components of dynamic image sequences: A theory of signal separation, image segmentation, and optical flow estimation. Lecture Notes in Computer Science, 1407:69–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070069.
                                        pdf.

Nagel:1998:SAE

[1620] H.-H. Nagel and A. Gehrke. Spatiotemporally adaptive estimation and segmentation of OF-fields. Lecture Notes in Computer Science, 1407:86– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070086.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070086.
                                        pdf.

Sawhney:1998:RVM

[1621] Harpreet S. Sawhney, Steve Hsu, and R. Kumar. Robust video mosaicing through topology inference and local to global alignment. Lecture Notes in Computer Science, 1407:103– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070103.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1407/14070103.
pdf.

Gdalyahu:1998:FSM

[1622] Yoram Gdalyahu and Daphna Weinshall. Flexible syntactic matching of curves. Lecture Notes in Computer Science, 1407:123–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070123.
pdf.

Cross:1998:HM

[1623] Andrew D. J. Cross and Edwin R.

Hancock. Holistic matching. Lecture Notes in Computer Science, 1407:140– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070140.
pdf.

Lopez:1998:CLS

[1624] Antonio M. López, Felipe Lumbreras, and Joan Serrat. Creaseness from level set extrinsic curvature. Lecture Notes in Computer Science, 1407:156– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070156.
pdf.

Balschbach:1998:MSS

[1625] Günther Balschbach, Jochen Klinke, and Bernd Jähne. Multichannel shape from shading techniques for moving specular surfaces. Lecture Notes in Computer Science, 1407:170–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070170.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070170.
                                        pdf.

Werghi:1998:MOH

[1626] Naoufel Werghi, Robert Fisher, Craig Robertson, and Anthony Ashbrook. Modelling objects having quadric surfaces incorporating geometric constraints. Lecture Notes in Computer Science, 1407:185–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070185.
                                        pdf.

Law:1998:SRM

[1627] N. F. Law and R. Chung. Surface reconstruction with multiresolution discontinuity analysis. Lecture Notes in Computer Science, 1407:202– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070202.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1407/14070202.
pdf.

Koenderink:1998:SCN

[1628] Jan Koenderink and Andrea van Doorn. Shape from Chebyshev nets. Lecture Notes in Computer Science, 1407:215–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1407/14070215.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1407/14070215.
pdf.

Baratoff:1998:CSC

[1629] Gregory Baratoff and Yiannis Aloimonos. Changes in surface convexity and topology caused by distortions of stereoscopic visual space. Lecture Notes in Computer Science, 1407:226– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070226.
pdf.

Stoddart:1998:RSS

[1630] A. J. Stoddart and M. Baker. Reconstruction of smooth surfaces with arbitrary topology adaptive splines. Lecture Notes in Computer Science, 1407: 241–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070241.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1407/14070241.
pdf.

Gerber:1998:MUD

[1631] Ralf Gerber and Hans-Hellmut Nagel. (mis?-) using DRT for generation of natural language text from image sequences. Lecture Notes in Computer Science, 1407:255–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070255.
pdf.

Nielsen:1998:SOF

[1632] Mads Nielsen and Ole Fogh Olsen. The structure of the optic flow field. Lecture Notes in Computer Science, 1407: 271–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070271.
pdf.

Srinivasan:1998:OFU

[1633] Sridhar Srinivasan and Rama Chellappa. Optical flow using overlapped basis functions for solving global motion problems. Lecture Notes in Computer Science, 1407:288– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070288.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1407/14070288.
pdf.

Muehlich:1998:RTL

[1634] M. Muehlich and R. Mester. The role of total least squares in motion analysis. Lecture Notes in Computer Science, 1407:305–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muhlich:1998:RTL

[1635] Matthias Mühlich and Rudolf Mester.

The role of total least squares in motion analysis. Lecture Notes in Computer Science, 1407:305–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070305.
pdf.

Jaehne:1998:SDP

[1636] B. Jaehne, H. Haussecker, H. Scharr, and H. Spies. Study of dynamical processes with tensor-based spatiotemporal image processing techniques. Lecture Notes in Computer Science, 1407: 322–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jahne:1998:SDP

[1637] B. Jähne, H. Haußecker, H. Scharr, H. Spies, D. Schmundt, and U. Schurr. Study of dynamical processes with tensor-based spatiotemporal image processing techniques. Lecture Notes in Computer Science, 1407:322–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070322.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070322.
pdf.

Ma:1998:MRI

[1638] Yi Ma, Jana Kosecká, and Shankar Sastry. Motion recovery from image sequences: Discrete viewpoint vs. differential viewpoint. Lecture Notes in Computer Science, 1407:337– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070337.
pdf.

Bernard:1998:DWA

[1639] Christophe P. Bernard. Discrete wavelet analysis: A new framework for fast optic flow computation. Lecture Notes in Computer Science, 1407:354– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070354.
pdf.

Lohmann:1998:ADL

[1640] Gabriele Lohmann and D. Yves von Cramon. Automatic detection and labelling of the human cortical folds in magnetic resonance data sets. Lecture Notes in Computer Science, 1407:369– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070369.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070369.
pdf.

Werner:1998:EDS

[1641] Tomás Werner, Tomás Pajdla, and Václav Hlavác. Efficient 3-D scene visualization by image extrapolation. Lecture Notes in Computer Science, 1407: 382–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070382.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070382.
pdf.

Kriegman:1998:WSR

[1642] David J. Kriegman and Peter N.

Belhumeur. What shadows reveal about object structure. Lecture Notes in Computer Science, 1407:399–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070399.
pdf.

Haddon:1998:SRS

[1643] John Haddon and David Forsyth.

Shape representations from shading primitives. Lecture Notes in Computer Science, 1407:415–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070415.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070415.
                                        pdf.

Williams:1998:CMD

[1644] Lance R. Williams and Karvel K. Thornber. A comparison of measures for detecting natural shapes in cluttered backgrounds. Lecture Notes in Computer Science, 1407:432– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070432.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070432.
                                        pdf.

Wilson:1998:BVT

[1645] Richard C. Wilson and Edwin R. Hancock. Bias-variance tradeoff for adaptive surface meshes. Lecture Notes in Computer Science, 1407:449–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070449.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070449.
                                        pdf.

Pontil:1998:RDO

[1646] Massimiliano Pontil, Stefano Rogai, and Alessandro Verri. Recognizing 3-D objects with linear support vector machines. Lecture Notes in Computer Science, 1407:469–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070469.
pdf.

Cootes:1998:AAM

[1647] T. F. Cootes, G. J. Edwards, and C. J.

Taylor. Active appearance models. Lecture Notes in Computer Science, 1407: 484–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070484.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070484.
pdf.

Vetter:1998:ECF

[1648] Thomas Vetter and Volker Blanz.

Estimating coloured 3D face models from single images: An example based approach. Lecture Notes in Computer Science, 1407:499–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070499.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070499.
pdf.

Matthews:1998:CAS

[1649] Iain Matthews, J. Andrew Bangham, Richard Harvey, and Stephen Cox. A comparison of active shape model and scale decomposition based features for visual speech recognition. Lecture Notes in Computer Science, 1407:514– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070514.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070514.
                                        pdf.

Bergen:1998:MSD

[1650] Lothar Bergen and Fernand Meyer. Motion segmentation and depth ordering based on morphological segmentation. Lecture Notes in Computer Science, 1407:531–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1407/14070531.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1407/14070531.pdf.

Kornprobst:1998:ISR

[1651] Pierre Kornprobst, Rachid Deriche, and Gilles Aubert. Image sequence restoration: A PDE based coupled method for image restoration and motion segmentation. Lecture Notes in Computer Science, 1407:548–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070548.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070548.
                                        pdf.

Memin:1998:JES

[1652] Étienne Mémin and Patrick Pérez. Joint estimation-segmentation of optic flow. Lecture Notes in Computer Science, 1407:563–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1407/14070563.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1407/14070563.pdf.

Edwards:1998:FRU

[1653] G. J. Edwards, T. F. Cootes, and C. J. Taylor. Face recognition using active appearance models. Lecture Notes in Computer Science, 1407:581– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070581.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070581.
pdf.

Liu:1998:FRU

[1654] Chengjun Liu and Harry Wechsler. Face recognition using evolutionary pursuit. Lecture Notes in Computer Science, 1407:596–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070596.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070596.
pdf.

Nagao:1998:RFW

[1655] Kenji Nagao and Masaki Sohma. Recognizing faces by weakly orthogonalizing against perturbations. Lecture Notes in Computer Science, 1407:613– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070613.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070613.
                                        pdf.

Burl:1998:PAO

[1656] Michael C. Burl, Markus Weber, and Pietro Perona. A probabilistic approach to object recognition using local photometry and global geometry. Lecture Notes in Computer Science, 1407: 628–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070628.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070628.
                                        pdf.

Jurie:1998:HVM

[1657] Frédéric Jurie. Hypothesis verification in model-based object recognition with a Gaussian error model. Lecture Notes in Computer Science, 1407:642– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070642.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070642.
                                        pdf.

Luettin:1998:CAV

[1658] Juergen Luettin and Stéphane Dupont. Continuous audio-visual speech recognition. Lecture Notes in Computer Science, 1407:657–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1407/14070657.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1407/14070657.pdf.

Ashbrook:1998:FSC

[1659] A. P. Ashbrook, R. B. Fisher, C. Robertson, and N. Werghi. Finding surface correspondence for object recognition and registration using pairwise geometric histograms. Lecture Notes in Computer Science, 1407:674– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070674.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070674.
pdf.

Fisher:1998:IIS

[1660] R. B. Fisher and A. MacKirdy. Integrating iconic and structured matching. Lecture Notes in Computer Science, 1407:687–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1407/14070687.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1407/14070687.
pdf.

Massad:1998:CMV

[1661] Amin Massad, Bärbel Mertsching, and Steffen Schmalz. Combining multiple views and temporal associations for 3-D object recognition. Lecture Notes in Computer Science, 1407:699– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070699.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070699.
                                        pdf.

Weiss:1998:MBR

[1662] Isaac Weiss and Manjit Ray. Modelbased recognition of 3D objects from one view. Lecture Notes in Computer Science, 1407:716–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070716.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070716.
                                        pdf.

Li:1998:TSP

[1663] Stan Z. Li and Joachim Hornegger. A two-stage probabilistic approach for object recognition. Lecture Notes in Computer Science, 1407:733– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070733.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070733.
                                        pdf.

Herbin:1998:CGP

[1664] Stéphane Herbin. Combining geometric and probabilistic structure for active recognition of 3D objects. Lecture Notes in Computer Science, 1407:748– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070748.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070748.
pdf.

MacCormick:1998:SDO

[1665] John MacCormick and Andrew Blake.

Spatial dependence in the observation of visual contours. Lecture Notes in Computer Science, 1407:765–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070765.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070765.
pdf.

Krebs:1998:HUO

[1666] B. Krebs, M. Burkhardt, and B. Korn.

Handling uncertainty in 3D object recognition using Bayesian networks. Lecture Notes in Computer Science, 1407:782–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1407/14070782.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1407/14070782.
pdf.

Kanatani:1998:ORS

[1667] Kenichi Kanatani and Naoya Ohta.

Optimal robot self-localization and reliability evaluation. Lecture Notes in Computer Science, 1407:796–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070796.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070796.
                                        pdf.

Davison:1998:MRL

[1668] Andrew J. Davison and David W. Murray. Mobile robot localisation using active vision. Lecture Notes in Computer Science, 1407:809–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070809.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070809.
                                        pdf.

Irani:1998:RFR

[1669] M. Irani, P. Anandan, and D. Weinshall. From reference frames to reference planes: Multi-view parallax geometry and applications. Lecture Notes in Computer Science, 1407:829– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1407/14070829.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1407/14070829.
                                        pdf.

Criminisi:1998:DRP

[1670] A. Criminisi, I. Reid, and A. Zisserman. Duality, rigidity and planar parallax. Lecture Notes in Computer Science, 1407:846–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070846.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070846.
pdf.

Stein:1998:DLR

[1671] Gideon P. Stein and Amnon Shashua.

On degeneracy of linear reconstruction from three views: Linear line complex and applications. Lecture Notes in Computer Science, 1407:862– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1407/14070862.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1407/14070862.
pdf.

Anonymous:1998:AIf

[1672] Anonymous. Author index. Lecture Notes in Computer Science, 1407:879– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1407/14070879.pdf.

Carter:1998:RDP

[1673] Michael W. Carter and Gilbert Laporte. Recent developments in practical course timetabling. Lecture Notes in Computer Science, 1408:3– 19, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1408/14080003.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1408/14080003.
                                        pdf.

Burke:1998:SAA

[1674] E. K. Burke and D. B. Varley. Space allocation: An analysis of higher education requirements. Lecture Notes in Computer Science, 1408:20–36, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1408/14080020.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1408/14080020.
                                        pdf.

Dowsland:1998:PMM

[1675] Kathryn A. Dowsland. Off-the-peg or made-to-measure? timetabling and scheduling with SA and TS. Lecture Notes in Computer Science, 1408:37– 52, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1408/14080037.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1408/14080037.
                                        pdf.

Ferland:1998:GAT

[1676] Jacques A. Ferland. Generalized assignment-type problems: A powerful modeling scheme. Lecture Notes in Computer Science, 1408:53–77, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1408/14080053.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1408/14080053.
pdf.

Bullnheimer:1998:ESM

[1677] Bernd Bullnheimer. An examination scheduling model to maximize students’ study time. Lecture Notes in Computer Science, 1408:78–91, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1408/14080078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1408/14080078.
pdf.

Elmohamed:1998:CAT

[1678] M. A. Saleh Elmohamed, Paul Coddington, and Geoffrey Fox. A comparison of annealing techniques for academic course scheduling. Lecture Notes in Computer Science, 1408:92– 114, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1408/14080092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1408/14080092.
pdf.

Ross:1998:SOA

[1679] Peter Ross, Emma Hart, and Dave Corne. Some observations about GAbased exam timetabling. Lecture Notes in Computer Science, 1408:115–129, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1408/14080115.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1408/14080115.
                                        pdf.

Meisels:1998:ENE

[1680] Amnon Meisels and Natalia Lusternik. Experiments on networks of employee timetabling problems. Lecture Notes in Computer Science, 1408:130–141, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1408/14080130.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1408/14080130.
                                        pdf.

Corne:1998:EOM

[1681] David Corne and John Ogden. Evolutionary optimisation of Methodist preaching timetables. Lecture Notes in Computer Science, 1408:142–155, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1408/14080142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1408/14080142.
                                        pdf.

Paechter:1998:ILT

[1682] Ben Paechter, R. C. Rankin, and Andrew Cumming. Improving a lecture timetabling system for university-wide use. Lecture Notes in Computer Science, 1408:156–168, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1408/14080156.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1408/14080156.
pdf.

David:1998:CBA

[1683] Philippe David. A constraint-based approach for examination timetabling using local repair techniques. Lecture Notes in Computer Science, 1408:169– 186, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1408/14080169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1408/14080169.
pdf.

White:1998:GCU

[1684] George M. White and Junhan Zhang.

Generating complete university timetables by combining tabu search with constraint logic. Lecture Notes in Computer Science, 1408:187–200, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1408/14080187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1408/14080187.
pdf.

VanWeert:1998:CBM

[1685] A. Van Weert and J. A. M. Schreuder.

Construction of basic match schedules (BMS) for sports competitions by using graph theory. Lecture Notes in Computer Science, 1408:201–212, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weert:1998:CBM

[1686] Arjen van Weert and Jan A. M. Schreuder. Construction of basic match schedules for sports competitions by using graph theory. Lecture Notes in Computer Science, 1408:201– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1408/14080201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1408/14080201.
pdf.

Burke:1998:SDF

[1687] Edmund K. Burke, Jeffrey H. Kingston, and Paul A. Pepper. A standard data format for timetabling instances. Lecture Notes in Computer Science, 1408: 213–222, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1408/14080213.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1408/14080213.
pdf.

Blakesley:1998:AS

[1688] James F. Blakesley, Keith S. Murray, Frederick H. Wolf, and Dagmar Murray. Academic scheduling. Lecture Notes in Computer Science, 1408:223– 236, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1408/14080223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1408/14080223.
pdf.

McCollum:1998:ICT

[1689] Barry McCollum. The implementation of a central timetabling system in a large British civic university. Lecture Notes in Computer Science, 1408:237– 253, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1408/14080237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1408/14080237.
pdf.

Nepal:1998:BFH

[1690] T. Nepal, S. W. Melville, and M. I.

Ally. A brute forte and heuristics approach to tertiary timetabling. Lecture Notes in Computer Science, 1408:254– 265, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1408/14080254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1408/14080254.
pdf.

Anonymous:1998:LOC

[1691] Anonymous. A list of other conference presentations. Lecture Notes in Computer Science, 1408:269–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1408/14080269.pdf.

Gasperoni:1998:EO

[1692] Franco Gasperoni. Embedded opportunities. Lecture Notes in Computer Science, 1411:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1411/14110001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1411/14110001.pdf.

Chapfont:1998:AFS

[1693] P. Chapfont. Ada + B: The formula for safety critical software development. Lecture Notes in Computer Science, 1411:14–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapront:1998:AFS

[1694] Pierre Chapront. Ada+B the formula for safety critical software development. Lecture Notes in Computer Science, 1411:14–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110014.
pdf.

Millet:1998:PGT

[1695] Laurent Millet and Ted Baker. Porting the GNAT tasking runtime system to the Java Virtual Machine. Lecture Notes in Computer Science, 1411:19– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110019.
pdf.

Emery:1998:AAB

[1696] David E. Emery, Robert F. Mathis, and Karl A. Nyberg. Automating the Ada binding process for Java — how far can we go? Lecture Notes in Computer Science, 1411:29– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110029.
pdf.

Ben-Ari:1998:SMC

[1697] Mordechai Ben-Ari. Synchronizing multiple clients and servers. Lecture Notes in Computer Science, 1411:41– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110041.
pdf.

Schumacher:1998:HAI

[1698] Guido Schumacher and Wolfgang Nebel. How to avoid the inheritance anomaly in Ada. Lecture Notes in Computer Science, 1411:53– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110053.
pdf.

Pautet:1998:IDS

[1699] Laurent Pautet and Samuel Tardieu. Inside the distributed systems annex. Lecture Notes in Computer Science, 1411:65–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1411/14110065.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1411/14110065.
                                        pdf.

Patino-Martinez:1998:IGT

[1700] Marta Patiño-Martínez, Ricardo Jiménez-Peris, and Sergio Arévalo. Integrating groups and transactions: A faulttolerant extension of Ada. Lecture Notes in Computer Science, 1411:78– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1411/14110078.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1411/14110078.
                                        pdf.

GonzalesHarbour:1998:IUE

[1701] M. Gonzales Harbour, M. A. Rivas, J. J. Gutierrez Garcia, and J. C. P. Gutierrez. Implementing and using execution time clocks in Ada hard realtime applications. Lecture Notes in Computer Science, 1411:90–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harbour:1998:IUE

[1702] M. González Harbour, M. Aldea Rivas, J. J. Gutiérrez García, and J. C. Palencia Gutiérrez. Implementing and using execution time clocks in Ada hard real-time applications. Lecture Notes in Computer Science, 1411:90– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110090.
pdf.

Espinosa:1998:PHR

[1703] Agustín Espinosa, Vicente Julián, Carlos Carrascosa, Andrés Terrasa, and Ana García-Fornes. Programming hard real-time systems with optional components in Ada. Lecture Notes in Computer Science, 1411:102–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110102.
pdf.

Moody:1998:OOA

[1704] Scott Arthur Moody. Object oriented abstractions for real-time distributed systems. Lecture Notes in Computer Science, 1411:112–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110112.
pdf.

Stiff:1998:APS

[1705] M. J. Stiff. (astro)physical supercomputing: Ada95 as a safe, object oriented alternative. Lecture Notes in Computer Science, 1411:128–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stift:1998:APS

[1706] Martin J. Stift. (astro)physical supercomputing: Ada95 as a safe, object oriented alternative. Lecture Notes in Computer Science, 1411:128– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1411/14110128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1411/14110128.
                                        pdf.

Hagenauer:1998:ADS

[1707] Helge Hagenauer and Werner Pohlmann. Ada 95 for a distributed simulation system. Lecture Notes in Computer Science, 1411:140–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1411/14110140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1411/14110140.
                                        pdf.

Harbour:1998:PPA

[1708] M. González Harbour, R. Gómez Somarriba, A. Strohmeier, and J. Jacot. PINROB: A portable API for industrial robots. Lecture Notes in Computer Science, 1411:151–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110151.
pdf.

Strohmeier:1998:QAP

[1709] Alfred Strohmeier, Vasiliy Fofanov, Sergey Rybin, and Stéphane Barbey. Quality-for-ASIS: A portable testing facility for ASIS. Lecture Notes in Computer Science, 1411:163– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110163.
pdf.

Tonndorf:1998:TYT

[1710] Michael Tonndorf. Ten years of tool based Ada compiler validations. an experience report. Lecture Notes in Computer Science, 1411:176–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110176.
pdf.

Araban:1998:TLM

[1711] S. Araban and A. S. M. Sajeev.

A two-level matching mechanism for object-oriented class libraries. Lecture Notes in Computer Science, 1411:188– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1411/14110188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1411/14110188.
                                        pdf.

Lewis:1998:MAR

[1712] Bruce Lewis, Steve Vestal, and David McConnell. Modern avionics requirements for the distributed systems annex. Lecture Notes in Computer Science, 1411:201–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1411/14110201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1411/14110201.
                                        pdf.

Fernandez:1998:CSQ

[1713] José L. Fernández, Bárbara Álvarez, Francisco García, Ángel Pérez, and Juan A. de la Puente. A case study in quantitative evaluation of realtime software architectures. Lecture Notes in Computer Science, 1411:213– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1411/14110213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1411/14110213.
                                        pdf.

Gonzalez-Barahona:1998:BMC

[1714] Jesús M. González-Barahona, Pedro de-las Heras-Quirós, José Centeno-González, and Francisco Ballesteros. Building modular communication systems in Ada: The Simple Com approach. Lecture Notes in Computer Science, 1411:225–237, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110225.
pdf.

Biieberger:1998:SRD

[1715] J. Biieberger and B. Burgstaller. Symbolic reaching definitions analysis of Ada programs. Lecture Notes in Computer Science, 1411:238–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blieberger:1998:SRD

[1716] Johann Blieberger and Bernd Burgstaller.

Symbolic reaching definitions analysis of Ada programs. Lecture Notes in Computer Science, 1411:238–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1411/14110238.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1411/14110238.
pdf.

Wong:1998:LCY

[1717] Ken Wong. Looking at code with your safety goggles on. Lecture Notes in Computer Science, 1411:251–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1411/14110251.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1411/14110251.
                                        pdf.

Burns:1998:RTP

[1718] A. Burns, B. Dobbing, and G. Romanski. The ravenscar tasking profile for high integrity real-time programs. Lecture Notes in Computer Science, 1411: 263–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1411/14110263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1411/14110263.
                                        pdf.

Michell:1998:GUA

[1719] Steve Michell and Mark Saaltink. Guidance on the use of Ada95 in high integrity systems. Lecture Notes in Computer Science, 1411:276–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1411/14110276.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1411/14110276.
                                        pdf.

Frisberg:1998:AJG

[1720] Bo Frisberg. Ada in the JAS 39 Gripen flight control system. Lecture Notes in Computer Science, 1411:288– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1411/14110288.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1411/14110288.
pdf.

Cornuejols:1998:PP

[1721] Gérard Cornuéjols, Bertrand Guenin, and François Margot. The packing property. Lecture Notes in Computer Science, 1412:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1412/14120001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1412/14120001.pdf.

Guenin:1998:CWB

[1722] Bertrand Guenin. A characterization of weakly bipartite graphs. Lecture Notes in Computer Science, 1412:9– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120009.
pdf.

Gasparyan:1998:BD

[1723] Grigor Gasparyan. Bipartite designs.

Lecture Notes in Computer Science, 1412:23–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1412/14120023.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1412/14120023.
pdf.

Sebo:1998:CNP

[1724] András Sebo. Characterizing noninteger polyhedra with 0-1 constraints. Lecture Notes in Computer Science, 1412:37–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1412/14120037.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120037.
                                        pdf.

Conforti:1998:TT

[1725] Michele Conforti and Ajai Kapoor. A theorem of truemper. Lecture Notes in Computer Science, 1412:53– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120053.
                                        pdf.

Nakamura:1998:GSS

[1726] Daishin Nakamura and Akihisa Tamura. The generalized stable set problem for claw-free bidirected graphs. Lecture Notes in Computer Science, 1412:69– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120069.
                                        pdf.

Szigeti:1998:MMT

[1727] Zoltán Szigeti. On a min-max theorem of cacti. Lecture Notes in Computer Science, 1412:84–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120084.
pdf.

Nagamochi:1998:ESE

[1728] Hiroshi Nagamochi and Peter Eades.

Edge-splitting and edge-connectivity augmentation in planar graphs. Lecture Notes in Computer Science, 1412: 96–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120096.
pdf.

Carr:1998:NBE

[1729] Robert Carr and R. Ravi. A new bound for the 2-edge connected subgraph problem. Lecture Notes in Computer Science, 1412:112–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120112.
pdf.

Cheriyan:1998:IAA

[1730] Joseph Cheriyan, András Sebo, and Zoltán Szigeti. An improved approximation algorithm for minimum size 2-edge connected spanning subgraphs. Lecture Notes in Computer Science, 1412:126–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120126.
                                        pdf.

Calinescu:1998:MUG

[1731] Gruia Calinescu, Cristina G. Fernandes, and Bruce Reed. Multicuts in unweighted graphs with bounded degree and bounded tree-width. Lecture Notes in Computer Science, 1412:137– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120137.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120137.
                                        pdf.

Kolliopoulos:1998:ADP

[1732] Stavros G. Kolliopoulos and Clifford Stein. Approximating disjoint-path problems using greedy algorithms and packing integer programs. Lecture Notes in Computer Science, 1412:153– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120153.
                                        pdf.

Raghavachari:1998:AAM

[1733] Balaji Raghavachari and Jeyakesavan Veerasamy. Approximation algorithms for the mixed postman problem. Lecture Notes in Computer Science, 1412: 169–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120169.
pdf.

Chudak:1998:IAA

[1734] Fabián A. Chudak. Improved approximation algorithms for uncapacitated facility location. Lecture Notes in Computer Science, 1412:180–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120180.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120180.
pdf.

Barvinok:1998:MTS

[1735] Alexander Barvinok, David S. Johnson, Gerhard J. Woeginger, and Russell Woodroofe. The Maximum Traveling Salesman Problem under polyhedral norms. Lecture Notes in Computer Science, 1412:195–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120195.
pdf.

Abeledo:1998:PCB

[1736] Hernán Abeledo and Gary Atkinson. Polyhedral combinatorics of benzenoid problems. Lecture Notes in Computer Science, 1412:202–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120202.
                                        pdf.

Christof:1998:COB

[1737] Thomas Christof, Marcus Oswald, and Gerhard Reinelt. Consecutive ones and a betweenness problem in computational biology. Lecture Notes in Computer Science, 1412:213–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120213.
                                        pdf.

Aardal:1998:SLD

[1738] Karen Aardal, Cor Hurkens, and Arjen K. Lenstra. Solving a linear Diophantine equation with lower and upper bounds on the variables. Lecture Notes in Computer Science, 1412:229– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120229.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120229.
                                        pdf.

Martin:1998:IKP

[1739] Alexander Martin and Robert Weismantel. The intersection of knapsack polyhedra and extensions. Lecture Notes in Computer Science, 1412: 243–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120243.
pdf.

Fekete:1998:NCL

[1740] Sándor P. Fekete and Jörg Schepers. New classes of lower bounds for bin packing problems. Lecture Notes in Computer Science, 1412:257– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120257.
pdf.

Ceria:1998:SID

[1741] Sebastián Ceria and Gábor Pataki.

Solving integer and disjunctive programs by lift and project. Lecture Notes in Computer Science, 1412:271– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120271.
pdf.

Cornuejols:1998:CHS

[1742] Gérard Cornuéjols and Milind Dawande. A class of hard small 0-1 programs. Lecture Notes in Computer Science, 1412: 284–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120284.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120284.
                                        pdf.

Fleischer:1998:BCC

[1743] Lisa Fleischer. Building chain and cactus representations of all minimum cuts from Hao-Orlin in the same asymptotic run time. Lecture Notes in Computer Science, 1412:294– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120294.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120294.
                                        pdf.

Tardos:1998:SGM

[1744] Éva Tardos and Kevin D. Wayne. Simple generalized maximum flow algorithms. Lecture Notes in Computer Science, 1412:310–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1412/14120310.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1412/14120310.pdf.

Hochbaum:1998:PAP

[1745] Dorit S. Hochbaum. The pseudoflow algorithm and the pseudoflow-based simplex for the maximum flow problem. Lecture Notes in Computer Science, 1412:325–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1412/14120325.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1412/14120325.
pdf.

Goldberg:1998:ICA

[1746] Andrew V. Goldberg, Jeffrey D. Oldham, Serge Plotkin, and Cliff Stein. An implementation of a combinatorial approximation algorithm for minimumcost multicommodity flow. Lecture Notes in Computer Science, 1412:338– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120338.
pdf.

Hoogeveen:1998:NAR

[1747] Han Hoogeveen, Petra Schuurman, and Gerhard J. Woeginger. Nonapproximability results for scheduling problems with minsum criteria. Lecture Notes in Computer Science, 1412: 353–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120353.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1412/14120353.
                                        pdf.

Munier:1998:ABG

[1748] Alix Munier, Maurice Queyranne, and Andreas S. Schulz. Approximation bounds for a general class of precedence constrained parallel machine scheduling problems. Lecture Notes in Computer Science, 1412:367–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120367.
                                        pdf.

Chekuri:1998:EAA

[1749] Chandra Chekuri and Michael Bender. An efficient approximation algorithm for minimizing makespan on uniformly related machines. Lecture Notes in Computer Science, 1412:383– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1412/14120383.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1412/14120383.
                                        pdf.

Uma:1998:RBC

[1750] R. N. Uma and Joel Wein. On the relationship between combinatorial and LP-based approaches to NPhard scheduling problems. Lecture Notes in Computer Science, 1412:394– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1412/14120394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120394.
pdf.

Kaibel:1998:PCQ

[1751] Volker Kaibel. Polyhedral combinatorics of quadratic assignment problems with less objects than locations. Lecture Notes in Computer Science, 1412:409–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1412/14120409.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1412/14120409.
pdf.

Helmberg:1998:IIC

[1752] Christoph Helmberg, Krzysztof C. Kiwiel, and Franz Rendl. Incorporating inequality constraints in the spectral bundle method. Lecture Notes in Computer Science, 1412:423–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1412/14120423.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1412/14120423.
pdf.

Rolland:1998:CVP

[1753] Colette Rolland. A comprehensive view of process engineering. Lecture Notes in Computer Science, 1413:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1413/14130001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1413/14130001.
                                        pdf.

Kardasis:1998:ALI

[1754] Panos Kardasis and Peri Loucopoulos. Aligning legacy information systems to business processes. Lecture Notes in Computer Science, 1413:25– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1413/14130025.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1413/14130025.
                                        pdf.

Harrison:1998:ARE

[1755] John V. Harrison and Wie Ming Lim. Automated reverse engineering of legacy 4GL information system applications using the ITOC workbench. Lecture Notes in Computer Science, 1413:41–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1413/14130041.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1413/14130041.
                                        pdf.

Antoniol:1998:AFP

[1756] G. Antoniol, F. Calzolari, L. Cristoforetti, R. Fiutem, and G. Caldiera. Adapting function points to object oriented information systems. Lecture Notes in Computer Science, 1413:59– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130059.
pdf.

Jin:1998:GCM

[1757] Shudong Jin and Xiaowei Sun. Global cache management for multi-class workloads in data warehouses. Lecture Notes in Computer Science, 1413:77– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130077.
pdf.

Jarke:1998:AQD

[1758] Matthias Jarke, Manfred A. Jeusfeld, Christoph Quix, and Panos Vassiliadis. Architecture and quality in data warehouses. Lecture Notes in Computer Science, 1413:93–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130093.
pdf.

Steiner:1998:OJM

[1759] Andreas Steiner, Adrian Kobler, and Moira C. Norrie. OMS/Java : Model extensibility of OODBMS for advanced application domains. Lecture Notes in Computer Science, 1413:115–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1413/14130115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1413/14130115.
                                        pdf.

Casati:1998:EDE

[1760] F. Casati, M. G. Fugini, and I. Mirbel. An environment for designing exceptions in workflows. Lecture Notes in Computer Science, 1413:139– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1413/14130139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1413/14130139.
                                        pdf.

Liu:1998:AHD

[1761] Chengfei Liu, Maria E. Orlowska, and Hui Li. Automating handover in dynamic workflow environments. Lecture Notes in Computer Science, 1413:159– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1413/14130159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1413/14130159.
                                        pdf.

Tietze:1998:DCG

[1762] Daniel A. Tietze, Ajit Bapat, and Rolf Reinema. Document-centric groupware for distributed governmental agencies. Lecture Notes in Computer Science, 1413:173–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130173.
pdf.

Rolland:1998:SRC

[1763] Colette Rolland, Véronique Plihon, and Jolita Ralyté. Specifying the reuse context of scenario method chunks. Lecture Notes in Computer Science, 1413:191–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130191.
pdf.

Lam:1998:CAM

[1764] Wing Lam. Change analysis and management in a reuse-oriented software development setting. Lecture Notes in Computer Science, 1413:219– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130219.
pdf.

Doemges:1998:FMM

[1765] R. Doemges, K. Pohl, and K. Schreck.

A filter-mechanism for method-driven trace capture. Lecture Notes in Computer Science, 1413:237–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domges:1998:FMM

[1766] Ralf Dömges, Klaus Pohl, and Klaus Schreck. A filter-mechanism for method-driven trace capture. Lecture Notes in Computer Science, 1413:237– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130237.
pdf.

Papazoglou:1998:SBO

[1767] Michael P. Papazoglou and Steven Milliner. Subject-based organization of the information space in multidatabase networks. Lecture Notes in Computer Science, 1413:251–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130251.
pdf.

Gaspary:1998:MIN

[1768] Luciano Paschoal Gaspary and Maria Janilce B. Almeida. MUSE — an interactive networked multimedia applications specification environment with E-LOTOS translator. Lecture Notes in Computer Science, 1413:273– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1413/14130273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130273.
pdf.

Lacroix:1998:IED

[1769] Zoé Lacroix, Arnaud Sahuguet, and Raman Chandrasekar. Information extraction and database techniques: A user-oriented approach to querying the Web. Lecture Notes in Computer Science, 1413:289–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130289.
pdf.

Kavakli:1998:GDB

[1770] V. Kavakli and P. Loucopoulos. Goaldriven business process analysis — application in electricity deregulation. Lecture Notes in Computer Science, 1413:305–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1413/14130305.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1413/14130305.
pdf.

Tanzi:1998:RTI

[1771] Tullio Tanzi, Sylvie Servigne, and Régis Guiol. Real-time information system for risk management on motorways. Lecture Notes in Computer Science, 1413:325–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1413/14130325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1413/14130325.
                                        pdf.

Nurcan:1998:DBP

[1772] Selmin Nurcan, Georges Grosz, and Carine Souveyet. Describing business processes with a guided use case approach. Lecture Notes in Computer Science, 1413:339–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1413/14130339.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1413/14130339.
                                        pdf.

Jurisica:1998:BQC

[1773] Igor Jurisica and Brian A. Nixon. Building quality into case-based reasoning systems. Lecture Notes in Computer Science, 1413:363–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1413/14130363.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1413/14130363.
                                        pdf.

Brinkkemper:1998:ATM

[1774] Sjaak Brinkkemper, Motoshi Saeki, and Frank Harmsen. Assembly techniques for method engineering. Lecture Notes in Computer Science, 1413: 381–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130381.
pdf.

Dahchour:1998:FMU

[1775] Mohamed Dahchour. Formalizing materialization using a metaclass approach. Lecture Notes in Computer Science, 1413:401–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1413/14130401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1413/14130401.
pdf.

Abramsky:1998:CVG

[1776] S. Abramsky and G. McCusker. Callby-value games. Lecture Notes in Computer Science, 1414:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayari:1998:LSL

[1777] A. Ayari, D. Basin, and A. Podelski. LISA: A specification language based on WS2S. Lecture Notes in Computer Science, 1414:18–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Backofen:1998:ECE

[1778] R. Backofen and P. Clote. Evolution as a computational engine. Lecture Notes in Computer Science, 1414:35– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baillot:1998:TG

[1779] P. Baillot, V. Danos, T. Erhard, and L. Regnier. Timeless games. Lecture Notes in Computer Science, 1414:56– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barber:1998:ACL

[1780] A. Barber, P. Gardner, M. Hasegawa, and G. Plotkin. From action calculi to linear logic. Lecture Notes in Computer Science, 1414:78–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonet:1998:LLB

[1781] M. L. Bonet and N. Galesi. Linear lower bounds and simulations in frege systems with substitutions. Lecture Notes in Computer Science, 1414:115– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauener:1998:FLL

[1782] T. Brauener and V. De Paiva. A formulation of linear logic based on dependency-relations. Lecture Notes in Computer Science, 1414:129–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buss:1998:RWP

[1783] S. Buss and T. Pitassi. Resolution and the weak pigeonhole principle. Lecture Notes in Computer Science, 1414:149–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comon:1998:HOM

[1784] H. Comon and Y. Jurski. Higher-order matching and tree automata. Lecture Notes in Computer Science, 1414:157– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coquand:1998:PTI

[1785] T. Coquand and H. Persson. A prooftheoretical investigation of Zantema’s problem. Lecture Notes in Computer Science, 1414:177–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand:1998:SOU

[1786] A. Durand, R. Fagin, and B. Loescher.

Spectra with only unary function symbols. Lecture Notes in Computer Science, 1414:189–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faggian:1998:CPB

[1787] C. Faggian. Classical proofs via basic logic. Lecture Notes in Computer Science, 1414:203–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grohe:1998:CLE

[1788] M. Grohe. Canonization for L0kappa-Equivalence is hard. Lecture Notes in Computer Science, 1414:220–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartonas:1998:FAF

[1789] C. Hartonas and M. Hennessy. Full abstractness for a functional/concurrent language with higher-order valuepassing. Lecture Notes in Computer Science, 1414:239–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heckmann:1998:DTQ

[1790] R. Heckmann and M. Huth. A duality theory for quantitative semantics. Lecture Notes in Computer Science, 1414:255–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1998:MML

[1791] M. Hofmann. A mixed modal/linear lambda calculus with applications to Bellantoni-Cook safe recursion. Lecture Notes in Computer Science, 1414: 275–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hollenberg:1998:EAT

[1792] M. Hollenberg. Equational axioms of test algebra. Lecture Notes in Computer Science, 1414:295–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klarlund:1998:MFL

[1793] N. Klarlund. Mona and fido: The logic-automaton connection in practice. Lecture Notes in Computer Science, 1414:311–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kupferman:1998:ERH

[1794] O. Kupferman, R. P. Kurshan, and M. Yannakakis. Existence of reduction hierarchies. Lecture Notes in Computer Science, 1414:327–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moschovakis:1998:GTC

[1795] Y. N. Moschovakis. A game-theoretic, concurrent and fair model of the typed lambda-calculus, with full recursion. Lecture Notes in Computer Science, 1414:341–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olive:1998:CLC

[1796] F. Olive. A conjunctive logical characterization of nondeterministic linear time. Lecture Notes in Computer Science, 1414:360–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pezzoli:1998:CCT

[1797] E. Pezzoli. On the computational complexity of type 2 functionals. Lecture Notes in Computer Science, 1414:373– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Power:1998:CAS

[1798] J. Power. Categories with algebraic structure. Lecture Notes in Computer Science, 1414:389–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruet:1998:CCP

[1799] P. Ruet and F. Fages. Concurrent constraint programming and noncommutative logic. Lecture Notes in Computer Science, 1414:406–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiering:1998:HAM

[1800] I. Schiering. A hierarchical approach to monadic second-order logic over graphs. Lecture Notes in Computer Science, 1414:424–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schweikardt:1998:MQA

[1801] N. Schweikardt. The monadic quantifier alternation hierarchy over grids and pictures. Lecture Notes in Computer Science, 1414:441–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwentick:1998:PEP

[1802] T. Schwentick. Padding and the expressive power of existential second-order logics. Lecture Notes in Computer Science, 1414:461–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staiger:1998:RCW

[1803] L. Staiger. Rich co-words and monadic second-order arithmetic. Lecture Notes in Computer Science, 1414:478–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voda:1998:SOR

[1804] P. J. Voda. A simple ordinal recursive normalization of Goedel’s T. Lecture Notes in Computer Science, 1414: 491–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wielinga:1998:KTM

[1805] B. J. Wielinga and A. T. Schreiber. Knowledge technology: Moving into the next millenium. Lecture Notes in Computer Science, 1415:1–20, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrero:1998:SCS

[1806] J. C. Herrero and J. Mira. In search of a common structure underlying a representative set of generic tasks and methods: The hierarchical classification and therapy planning cases study. Lecture Notes in Computer Science, 1415:21–36, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franova:1998:TCD

[1807] M. Franova and M. Kooli. Theory of constructible domains for robotics: Why? Lecture Notes in Computer Science, 1415:37–46, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aida:1998:ISM

[1808] T. Aida and S. Ohsuga. Intelligent systems must be able to make programs automatically for assuring the practicality. Lecture Notes in Computer Science, 1415:47–56, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barker:1998:TKL

[1809] R. Barker, A. Meehan, and I. Tranter. Towards a knowledge-level model for concurrent design. Lecture Notes in Computer Science, 1415:57–67, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nourani:1998:MAI

[1810] C. F. Nourani. Multiagent AI implementation techniques: A new software engineering trend. Lecture Notes in Computer Science, 1415:68–78, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haddar:1998:ISI

[1811] N. Haddar, F. Gargouri, A. B. Hamadou, and C. F. Ducateau. Information systems integration: Some principles and ideas. Lecture Notes in Computer Science, 1415:79–88, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kutti:1998:EPE

[1812] S. Kutti and B. Garner. An emergent paradigm for expert resource management systems. Lecture Notes in Computer Science, 1415:89–101, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ikeda:1998:NMM

[1813] W. Ikeda and S. Ohsuga. New modeling method for using large knowledge bases. Lecture Notes in Computer Science, 1415:102–111, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Manjon:1998:IFC

[1814] B. Fernandez-Manjon, A. Navarro, J. M. Cigarran, and A. Fernandez-Valmayor. Integration of formal concept analysis in a knowledge-based assistant. Lecture Notes in Computer Science, 1415:112–123, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcos:1998:KMP

[1815] M. Marcos, S. Moisan, and A. P. Del Pobil. Knowledge modeling of program supervision task. Lecture Notes in Computer Science, 1415:124–133, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mirita:1998:QDI

[1816] I. I. Mirita. Quadri-dimensional interpretation of syllogistic inferential processes in polyvalent logic, with a view to structuring concepts and assertions for realizing the universal knowledge basis. Lecture Notes in Computer Science, 1415:134–148, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rocha:1998:CPG

[1817] J. Rocha, C. Ramos, and Z. Vale. Complexity of precedence graphs for assembly and task planning. Lecture Notes in Computer Science, 1415:149–158, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drias:1998:MCA

[1818] H. Drias. A Monte Carlo algorithm for the satisfiability problem. Lecture Notes in Computer Science, 1415:159– 168, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez-Vallejo:1998:APR

[1819] M. L. Lopez-Vallejo, C. A. Iglesias, and J. C. Lopez. Applying the propose and revise strategy to the hardwaresoftware partitioning problem. Lecture Notes in Computer Science, 1415:169– 179, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delgado:1998:FHT

[1820] M. Delgado, A. F. Gomez-Skarmeta, J. Gomez Marin-Blazquez, and H. Martinez Barbera. Fuzzy hybrid techniques in modeling. Lecture Notes in Computer Science, 1415:180–189, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ozyurt:1998:CPF

[1821] I. Burak Özyurt, A. K. Sunol, and L. O. Hall. Chemical process fault diagnosis using kernel retrofitted fuzzy genetic algorithm based learner (FGAL) with a hidden Markov model. Lecture Notes in Computer Science, 1415:190–199, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kowalczyk:1998:LAG

[1822] R. Kowalczyk. On linguistic approximation with genetic programming. Lecture Notes in Computer Science, 1415: 200–224, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tzafestas:1998:NSM

[1823] S. G. Tzafestas and G. B. Stamou. A new solution methodology for fuzzy relation equations. Lecture Notes in Computer Science, 1415:210–224, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1998:AGM

[1824] M. Li, Y. Wang, and X. Zhang. An approach to generate membership function by using Kohonen’s SOFM nets. Lecture Notes in Computer Science, 1415:220–224, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calisti:1998:IPF

[1825] A. R. Calisti, F. D. Trujillo Aguilera, A. Diaz-Estrella, and F. Sandoval Hernandez. Intelligent policing function for ATM networks. Lecture Notes in Computer Science, 1415:225–233, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botana:1998:DFS

[1826] F. Botana. Deriving fuzzy subsethood measures from violations of the implication between elements. Lecture Notes in Computer Science, 1415:234–243, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piera:1998:OOS

[1827] M. A. Piera, A. J. Gambin, and R. Vilanova. Object oriented stochastic Petri new simulator: A generic kernel for heuristic planning tools. Lecture Notes in Computer Science, 1415:244–253, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawson:1998:ANN

[1828] C. W. Dawson, P. D. Wilson, and A. N.

Beard. An artificial neural network for flashover prediction. A preliminary study. Lecture Notes in Computer Science, 1415:254–262, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahraman:1998:FDM

[1829] C. Kahraman and E. Tolga. Fuzzy decision making under uncertainty. Lecture Notes in Computer Science, 1415:263– 271, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stewart:1998:IPN

[1830] B. Stewart. Improving performance of naive Bayes classifier by including hidden variables. Lecture Notes in Computer Science, 1415:272–280, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennett:1998:SHR

[1831] B. Bennett, A. Isli, and A. G. Cohn. A system handling RCC-8 queries on 2D regions representable in the closure algebra of half-planes. Lecture Notes in Computer Science, 1415:281–290, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Escrig:1998:CDE

[1832] M. T. Escrig and F. Toledo. Cardinal directions on extended objects for qualitative navigation. Lecture Notes in Computer Science, 1415:291–301, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agell:1998:DSL

[1833] N. Agell, F. Febles, and N. Piera. Definition and study of lineal equations on order of magnitude models. Lecture Notes in Computer Science, 1415:302–311, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turner:1998:CBA

[1834] E. H. Turner and R. M. Turner. A constraint-based approach to assigning system components to tasks. Lecture Notes in Computer Science, 1415:312– 320, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MartinezGasca:1998:ASA

[1835] R. Martinez Gasca, J. A. Ortega, and M. Toro. Automatic semiqualitative analysis: Application to a biometallurgical system. Lecture Notes in Computer Science, 1415:321–328, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ortega:1998:IQK

[1836] J. A. Ortega, R. Martinez Gasca, and M. Toro. Including qualitative knowledge in semiqualitative dynamical systems. Lecture Notes in Computer Science, 1415:329–335, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguilar:1998:TOH

[1837] J. Aguilar, J. Riquelme, and M. Toro.

A tool to obtain a hierarchical qualitative rules from quantitative data. Lecture Notes in Computer Science, 1415: 336–346, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ampratwum:1998:HTO

[1838] C. S. Ampratwum, P. D. Picton, A. A.

Hopgood, and A. Browne. Hybridization techniques in optical emission spectral analysis. Lecture Notes in Computer Science, 1415:347–356, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RuizGomez:1998:KQM

[1839] J. Ruiz Gomez. KADS qualitative model based motor speed control. Lecture Notes in Computer Science, 1415:357–366, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez:1998:NRB

[1840] M. Sanchez, F. Prats, and N. Piera. Negligibility relations between real numbers and qualitative labels. Lecture Notes in Computer Science, 1415:367– 376, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chachoua:1998:QRU

[1841] M. Chachoua and D. Pacholczyk. Qualitative reasoning under uncertain knowledge. Lecture Notes in Computer Science, 1415:377–386, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marzo:1998:QRA

[1842] J. L. Marzo, A. Bueno, R. Fabregat, and T. Jove. Qualitative reasoning for admission control in ATM networks. Lecture Notes in Computer Science, 1415:387–398, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dollinger:1998:UTC

[1843] R. Dollinger and I. A. Letia. Using tolerance calculus for reasoning in relative order of magnitude models. Lecture Notes in Computer Science, 1415:399–407, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MattosBrasil:1998:CCC

[1844] L. Mattos Brasil, F. Mendes De Azevedo, ?. Muniz, J. Barreto, and M. Noirhomme-Fraiture. Complexity and cognitive computing. Lecture Notes in Computer Science, 1415:408– 417, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cotta:1998:DMS

[1845] C. Cotta and J. M. Troya. On decisionmaking in strong hybrid evolutionary algorithms. Lecture Notes in Computer Science, 1415:418–427, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blasco:1998:GPC

[1846] X. Blasco, M. Martinez, J. Senent, and J. Sanchis. Generalized predictive control using genetic algorithms (GAGPC). an application to control of a non-linear process with model uncertainty. Lecture Notes in Computer Science, 1415:428–437, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1998:PLW

[1847] T. Ito. Piping layout wizard: Basic concepts and its potential for pipe route planning. Lecture Notes in Computer Science, 1415:438–447, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruiz-Andino:1998:ICP

[1848] A. Ruiz-Andino and J. J. Ruz. Integration of constraint programming and evolution programs: Application to channel routing. Lecture Notes in Computer Science, 1415:448–459, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henderson:1998:UGA

[1849] C. E. Henderson, W. D. Potter, R. W. McClendon, and G. Hoogenboom. Using a genetic algorithm to select parameters for a neural network that predicts aflatoxin contamination in peanuts. Lecture Notes in Computer Science, 1415:460–469, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luga:1998:IAF

[1850] H. Luga, O. Balet, Y. Duthen, and R. Caubet. Interacting with articulated figures within the provis project. Lecture Notes in Computer Science, 1415: 470–477, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordon:1998:CSM

[1851] O. Cordon, F. Herrera, and L. Sanchez. Computing the Spanish medium electrical line maintenance cost by means of evolution-based learning processes. Lecture Notes in Computer Science, 1415:478–486, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RaghavendraRao:1998:NDM

[1852] G. Raghavendra Rao and K. Chidananda Gowda. A new dissimilarity measure to improve the Ga performance. Lecture Notes in Computer Science, 1415:487–492, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BerlangaLlavori:1998:SDP

[1853] R. Berlanga Llavori. STRPLAN: A distributed planner for object-centred application domains. Lecture Notes in Computer Science, 1415:493–502, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1998:IHS

[1854] C.-C. Hsu and C.-S. Ho. An intelligent hybrid system for knowledge acquisition. Lecture Notes in Computer Science, 1415:503–510, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berka:1998:UNN

[1855] P. Berka and M. Slama. Using neural nets to learn weights of rules for compositional expert systems. Lecture Notes in Computer Science, 1415:511–519, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donskoy:1998:CKO

[1856] V. I. Donskoy. Case-, knowledge-, and optimization-based hybrid approach in AI. Lecture Notes in Computer Science, 1415:520–527, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonzon:1998:CMB

[1857] P. E. Bonzon. A contextual model of beliefs for communicating agents. Lecture Notes in Computer Science, 1415: 528–537, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turner:1998:CMB

[1858] R. M. Turner. Context-mediated behavior for AI applications. Lecture Notes in Computer Science, 1415:538– 546, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ozturk:1998:CSI

[1859] P. Öztürk. A context-sensitive, iterative approach to diagnostic problem solving. Lecture Notes in Computer Science, 1415:546–557, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Sheikh:1998:FDI

[1860] E. El-Sheikh and J. Sticklen. A framework for developing intelligent tutoring systems incorporating reusability. Lecture Notes in Computer Science, 1415:558–567, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goel:1998:CSD

[1861] A. Goel and K. S. Barber. The classification and specification of a domain independent agent architecture. Lecture Notes in Computer Science, 1415:568–576, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coble:1998:MMA

[1862] J. Coble and K. Harbison. Mave: a multi-agent architecture for virtual environments. Lecture Notes in Computer Science, 1415:577–583, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonker:1998:ABS

[1863] C. M. Jonker and J. Treur. Agentbased simulation of reactive, pro-active and social animal behaviour. Lecture Notes in Computer Science, 1415:584– 595, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iglesias:1998:FNM

[1864] C. A. Iglesias, J. C. Gonzalez, and J. R. Velasco. A fuzzy-neural multiagent system for optimisation of a roll-mill application. Lecture Notes in Computer Science, 1415:596–605, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lejouad-Chaari:1998:HLC

[1865] W. Lejouad-Chaari and F. Mouria-Beji. High-level communication protocol in a distributed multiagent system. Lecture Notes in Computer Science, 1415:606–616, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Podlena:1998:ESG

[1866] J. R. Podlena. Evolving the scale of genetic search. Lecture Notes in Computer Science, 1415:617–624, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Philpot:1998:ENN

[1867] D. Philpot and T. Hendtlass. Ensembles of neural networks for analogue problems. Lecture Notes in Computer Science, 1415:625–631, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Copland:1998:EAG

[1868] H. Copland and T. Hendtlass. An evolutionary algorithm with a genetic encoding scheme. Lecture Notes in Computer Science, 1415:632–639, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hendtlass:1998:GLT

[1869] T. Hendtlass. Generating lookup tables using evolutionary algorithms. Lecture Notes in Computer Science, 1415:640– 646, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhandarkar:1998:GAL

[1870] S. M. Bhandarkar, J. Zeppen, and W. D. Potter. A genetic algorithm for linear feature extraction. Lecture Notes in Computer Science, 1415:647–656, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peers:1998:KRB

[1871] S. M. C. Peers. Knowledge representation in a blackboard system for sensor data interpretation. Lecture Notes in Computer Science, 1415:657–666, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Utete:1998:LIP

[1872] S. W. Utete. Local information processing for decision making in decentralised sensing networks. Lecture Notes in Computer Science, 1415:667–676, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felix:1998:FLA

[1873] P. Felix, S. Fraga, R. Marin, and S. Barro. A fuzzy linear approach to recognizing signal profiles. Lecture Notes in Computer Science, 1415:677– 686, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marilly:1998:PSV

[1874] E. Marilly, C. Coroyer, O. Cachard, and A. Faure. Performance of a smart velocity sensor: The impulse retina. Lecture Notes in Computer Science, 1415:687–695, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgi:1998:TML

[1875] A. Borgi, J.-M. Bazin, and H. Akdag.

Two methods of linear correlation search for a knowledge based supervised classification. Lecture Notes in Computer Science, 1415:696–707, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turmo:1998:TSE

[1876] J. Turmo, N. Catala, and H. Rodriguez. Turbio: A system for extracting information from restricted-domain texts. Lecture Notes in Computer Science, 1415:708–721, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haddar:1998:AED

[1877] K. Haddar and A. B. Hamadou. An approach to ellipsis detection and classification for the Arabic language. Lecture Notes in Computer Science, 1415:722– 738, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz:1998:REB

[1878] V. J. Diaz, V. Carrillo, and M. Toro. A review of early-based parser for TIG. Lecture Notes in Computer Science, 1415:732–738, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1998:GEE

[1879] T. Tanaka. Generating explanations from electronic circuits. Lecture Notes in Computer Science, 1415:739–748, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soufian:1998:EDL

[1880] Majeed Soufian, Mustapha Soufian, and M. J. Dempsey. The effect of a dynamical layer in neural network prediction of biomass in a fermentation process. Lecture Notes in Computer Science, 1415:749–757, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garrido:1998:SEN

[1881] S. Garrido, L. Moreno, and C. Balaguer. State estimation for nonlinear systems using restricted genetic optimization. Lecture Notes in Computer Science, 1415:758–767, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moshiri:1998:INI

[1882] B. Moshiri and S. C. Maleki. Identification of a nonlinear industrial process via fuzzy clustering. Lecture Notes in Computer Science, 1415:768–775, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badenas:1998:ACV

[1883] J. Badenas and F. Pla. Applying computer vision techniques to traffic monitoring tasks. Lecture Notes in Computer Science, 1415:776–785, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragoni:1998:SDD

[1884] A. F. Dragoni, P. Giorgini, and M. Panti. A self-diagnosing distributed monitoring system for nuclear plants. Lecture Notes in Computer Science, 1415:786–795, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riquelme:1998:OST

[1885] J. M. Riquelme, A. Gomez, J. L. Martinez, and J. A. Pecas Lopez. Overload screening of transmission systems using neural networks. Lecture Notes in Computer Science, 1415:796–803, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:1998:LID

[1886] C. A. Gonzalez, B. Pulido Junquera, and G. Acosta. On line industrial diagnosis: An attempt to apply artificial intelligence techniques to process control. Lecture Notes in Computer Science, 1415:804–813, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodrigues:1998:AMA

[1887] M. A. Rodrigues and L. Bottaci. An AI modelling approach to understanding a complex manufacturing process. Lecture Notes in Computer Science, 1415:814–821, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lazaro:1998:IDI

[1888] A. Lazaro, I. Serrano, J. P. Oria, and C. De Miguel. Identification and defect inspection with ultrasonic techniques in foundry pieces. Lecture Notes in Computer Science, 1415:822–830, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xiao:1998:OKA

[1889] Y. Xiao and C. Y. Han. An OOM-KBES approach for fault detection and diagnosis. Lecture Notes in Computer Science, 1415:831–839, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heinzelmann:1998:MBF

[1890] A. Heinzelmann and P. Chammas. Model-based fault simulation: A reduction method for the diagnosis of electrical components. Lecture Notes in Computer Science, 1415:840–847, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1998:MBA

[1891] A. Lin. A model-based automated diagnosis algorithm. Lecture Notes in Computer Science, 1415:848–856, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1998:CFR

[1892] S.-I. Wu. Cyclic forecasting with recurrent neural network. Lecture Notes in Computer Science, 1415:857–863, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elloumi:1998:NAP

[1893] M. Elloumi. New algorithms to predict secondary structures of RNA macromolecules. Lecture Notes in Computer Science, 1415:864–875, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdel-Wahab:1998:HGS

[1894] A. H. Abdel-Wahab, M. E. El-Telbany, and S. I. Shaheen. A hybrid Ga statistical method for the forecasting problem: The prediction of the River Nile inflows. Lecture Notes in Computer Science, 1415:876–882, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MorenoDiaz:1998:NCC

[1895] R. Moreno Diaz. Neurocybernetics, codes and computation. Lecture Notes in Computer Science, 1416:1–14, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelPobil:1998:GCC

[1896] A. P. Del Pobil. The grand challenge is called: Robotic intelligence. Lecture Notes in Computer Science, 1416:15– 24, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Onaindia:1998:PHS

[1897] E. Onaindia, F. Barber, V. Botti, and C. Carrascosa. A progressive heuristic search algorithm for the cutting stock problem. Lecture Notes in Computer Science, 1416:25–35, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BerlangaLlavori:1998:DTR

[1898] R. Berlanga Llavori, M. J. Aramburu, and F. Barber. Discovering temporal relationships in databases of newspapers. Lecture Notes in Computer Science, 1416:36–45, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larrosa:1998:GCT

[1899] J. Larrosa and P. Meseguer. Generic CSP techniques for the job-shop problem. Lecture Notes in Computer Science, 1416:46–55, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oommen:1998:FES

[1900] B. J. Oommen and T. Dale Roberts. A fast and efficient solution to the capacity assignment problem using discretized learning automata. Lecture Notes in Computer Science, 1416:56– 65, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cameron:1998:UOP

[1901] S. Cameron. Using OXSIM for path planning. Lecture Notes in Computer Science, 1416:66–74, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henrich:1998:MDS

[1902] D. Henrich, C. Wurll, and H. Woern. Multi-directional search with goal switching for robot path planning. Lecture Notes in Computer Science, 1416:75–84, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1998:APF

[1903] S.-W. Kim and D. Boley. Analytical potential fields and control strategies for motion planning. Lecture Notes in Computer Science, 1416:85–94, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mirolo:1998:EGR

[1904] C. Mirolo and E. Pagello. Exact geometry and robot motion planning: Speculations on a few numerical experiments. Lecture Notes in Computer Science, 1416:95–104, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ridao:1998:ELS

[1905] M. A. Ridao, J. Riquelme, E. F. Camacho, and M. Toro. An evolutionary and local search algorithm for planning two manipulators motion. Lecture Notes in Computer Science, 1416:105–114, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gallardo:1998:GAR

[1906] D. Gallardo, O. Colomina, F. Florez, and R. Rizo. A genetic algorithm for robust motion planning. Lecture Notes in Computer Science, 1416:115–121, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perez-Francisco:1998:CMT

[1907] M. Perez-Francisco, A. P. Del Pobil, and B. Martinez-Salvador. Coordinated motion of two robot arms for real applications. Lecture Notes in Computer Science, 1416:122–131, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kruusmaa:1998:LRA

[1908] M. Kruusmaa and B. Svensson. A low-risk approach to mobile robot path planning. Lecture Notes in Computer Science, 1416:132–141, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stein:1998:GHC

[1909] B. Stein. Generating heuristics to control configuration processes. Lecture Notes in Computer Science, 1416:142– 152, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feurstein:1998:VFF

[1910] M. Feurstein and M. Natter. Valuing the flexibility of flexible manufacturing systems with fast decision rules. Lecture Notes in Computer Science, 1416:153–162, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikulin:1998:OPC

[1911] V. Nikulin. Optimal periodic control with environmental application. Lecture Notes in Computer Science, 1416: 163–172, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1998:CHT

[1912] L. A. Garcia and F. Toledo. A centralised hierarchical task scheduler for an urban traffic control system based on a multiagent architecture. Lecture Notes in Computer Science, 1416:173– 182, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1998:DIM

[1913] W. Chen, Z. Wu, Q. Zhang, and J. Li.

A direct iteration method for global dynamic control of redundant manipulators. Lecture Notes in Computer Science, 1416:183–191, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nechaev:1998:DSB

[1914] Y. I. Nechaev and Y. L. Siek. Design of ship-board control system based on the soft computing conception. Lecture Notes in Computer Science, 1416:192– 199, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melendez-Frigola:1998:EDU

[1915] J. Melendez-Frigola, J. Colomer-Llinas, J. Lluis De La Rosa-Esteva, and O. Contreras. Expert diagnostic using qualitative data and rule-based inferential reasoning. Lecture Notes in Computer Science, 1416:200–209, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

NevesJunior:1998:QEBa

[1916] F. Neves Junior and J. Aguilar-Martin.

Qualitative event-based expert supervision. part 1: Methodology. Lecture Notes in Computer Science, 1416:210– 219, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

NevesJunior:1998:QEBb

[1917] F. Neves Junior and J. Aguilar-Martin.

Qualitative event-based expert supervision. part 2: Distillation start-up condition monitoring. Lecture Notes in Computer Science, 1416:220–229, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordier:1998:APR

[1918] M.-O. Cordier, J.-P. Krivine, P. Laborie, and S. Thiebaux. Alarm processing and reconfiguration in power distribution systems. Lecture Notes in Computer Science, 1416:230–241, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zouaoui:1998:BIA

[1919] F. Zouaoui, R. Thetiot, and M. Dumas. Behavioral interpretation according to multimodeling representation. Lecture Notes in Computer Science, 1416:242–151, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamlin:1998:IAA

[1920] J. H. Hamlin and W. D. Potter. An intelligent agent to aid in Unix system administration. Lecture Notes in Computer Science, 1416:252–260, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Villanueva:1998:POI

[1921] H. Villanueva and H. Lamba. Process optimisation in an industrial supervision support system. Lecture Notes in Computer Science, 1416:261–270, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Senent:1998:MPC

[1922] J. S. Senent, M. A. Martinez, X. Blasco, and J. Sanchis. Mimo predictive control of temperature and humidity inside a greenhouse using simulated annealing (SA) as optimizer of a multicriteria index. Lecture Notes in Computer Science, 1416:271–279, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchiz:1998:SVB

[1923] J. M. Sanchiz, A. Broggi, and F. Pla.

Stereo vision-based obstacle and free space detection in mobile robotics. Lecture Notes in Computer Science, 1416: 280–289, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLope:1998:TMF

[1924] J. De Lope, D. Maravall, and J. G.

Zato. Topological modeling with fuzzy Petri nets for autonomous mobile robots. Lecture Notes in Computer Science, 1416:290–299, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iglesias:1998:SRL

[1925] R. Iglesias, C. V. Regueiro, J. Correa, and S. Barro. Supervised reinforcement learning: Application to a wall following behaviour in a mobile robot. Lecture Notes in Computer Science, 1416:300–309, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SanchezLopez:1998:CPC

[1926] M. Sanchez Lopez and P. Manzoni.

A communication protocol for clouds of mobile robots. Lecture Notes in Computer Science, 1416:310–318, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1998:ENC

[1927] J. Santos and R. J. Duro. Evolving neural controllers for temporally dependent behaviors in autonomous robots. Lecture Notes in Computer Science, 1416:319–328, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sugihara:1998:GBL

[1928] K. Sugihara. GA-Based on-line path planning for sauvim. Lecture Notes in Computer Science, 1416:329–338, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Romo:1998:IBM

[1929] J. Romo, F. De La Paz, and J. Mira. Incremental building of a model of environment in the context of the McCulloch-Craik’s functional architecture for mobile robots. Lecture Notes in Computer Science, 1416:339–352, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marin:1998:TSB

[1930] R. Marin, G. Recatala, P. J. Sanz, and J. M. Iñesta. Telerobotic system based on natural language and computer vision. Lecture Notes in Computer Science, 1416:353–364, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AngelesLopez:1998:RRB

[1931] M. Angeles Lopez and F. Pla. Robust region-based stereo vision to build environment maps for robotics applications. Lecture Notes in Computer Science, 1416:365–374, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tzafestas:1998:NLS

[1932] S. G. Tzafestas and K. C. Zikidis. A new on-line structure and parameter learning architecture for fuzzy modeling, based on neural and fuzzy techniques. Lecture Notes in Computer Science, 1416:375–382, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zilouchian:1998:ANF

[1933] A. Zilouchian, D. W. Howard, and T. Jordanides. An adaptive neurofuzzy inference System(ANFIS) approach to control of robotic manipulators. Lecture Notes in Computer Science, 1416:383–392, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez-Calero:1998:MUE

[1934] P. A. Gonzalez-Calero, M. Gomez-Albarran, and C. Fernandez-Chamizo. Managing the usage experience in a library of software components. Lecture Notes in Computer Science, 1416:393– 402, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thonnat:1998:WCP

[1935] M. Thonnat and S. Moisan. What can program supervision do for program re-use? Lecture Notes in Computer Science, 1416:403–414, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chien:1998:UAI

[1936] S. Chien, F. Fisher, H. Mortensen, and E. Lo. Using artificial intelligence planning techniques to automatically reconfigure software modules. Lecture Notes in Computer Science, 1416:415–426, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shekhar:1998:UKB

[1937] C. Shekhar, S. Moisan, R. Vincent, and P. Burlina. Use of knowledge-based control for vision systems. Lecture Notes in Computer Science, 1416:427– 436, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liedtke:1998:SLC

[1938] C.-E. Liedtke, H. Muenkel, and U. Rost. Solution for a learning configuration system for image processing. Lecture Notes in Computer Science, 1416:437–447, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luaces:1998:MLU

[1939] O. Luaces, J. Alonso, E. A. De La Cal, and J. Ranilla. Machine learning usefulness relies on accuracy and self-maintenance. Lecture Notes in Computer Science, 1416:448–457, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caraca-Valente:1998:IIL

[1940] J. P. Caraca-Valente and C. Montes. Improving inductive learning in realworld domains through the identification of dependencies: The tim framework. Lecture Notes in Computer Science, 1416:458–468, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez:1998:NNR

[1941] J. S. Sanchez, F. Pla, and F. J. Ferri.

From the nearest neighbour rule to decision trees. Lecture Notes in Computer Science, 1416:469–478, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ochi:1998:NSO

[1942] L. Satoru Ochi, N. Maculan, and R. M. Videira Figueiredo. A new self-organizing strategy based on elastic networks for solving the Euclidean traveling salesman problem. Lecture Notes in Computer Science, 1416:479– 487, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zargham:1998:ILS

[1943] M. R. Zargham. An inductive learning system for rating securities. Lecture Notes in Computer Science, 1416:488– 496, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilke:1998:TKU

[1944] W. Wilke and R. Bergmann. Techniques and knowledge used for adaptation during case-based problem solving. Lecture Notes in Computer Science, 1416:497–506, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smyth:1998:CBM

[1945] B. Smyth. Case-base maintenance. Lecture Notes in Computer Science, 1416: 507–516, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cunningham:1998:CSW

[1946] P. Cunningham. CBR: Strengths and weaknesses. Lecture Notes in Computer Science, 1416:517–524, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1998:CTM

[1947] I. Watson. Is CBR a technology or a methodology? Lecture Notes in Computer Science, 1416:525–534, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jurisica:1998:EAI

[1948] I. Jurisica and J. Glasgow. An efficient approach to iterative browsing and retrieval for case-based reasoning. Lecture Notes in Computer Science, 1416:535–546, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1998:CBA

[1949] K. Tanaka, T. Ohkawa, and N. Komoda. Case based approach to the construction of a coal molecular structure model. Lecture Notes in Computer Science, 1416:547–556, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treadgold:1998:CHO

[1950] N. K. Treadgold and T. D. Gedeon. Constructing higher order neurons of increasing complexity in cascade networks. Lecture Notes in Computer Science, 1416:557–563, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weijters:1998:INNb

[1951] T. Weijters and A. Van Den Bosch. Interpretable neural networks with BP-SOM. Lecture Notes in Computer Science, 1416:564–573, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehtokangas:1998:RPW

[1952] M. Lehtokangas. Reference pattern weight initialization for equalization. Lecture Notes in Computer Science, 1416:574–581, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrios:1998:ANN

[1953] L. J. Barrios and L. Lemus. Autoassociative neural networks for fault diagnosis in semiconductor manufacturing. Lecture Notes in Computer Science, 1416:582–592, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aladjem:1998:STN

[1954] M. Aladjem. Supervised training of a neural network for classification via successive modification of the training data — an experimental study. Lecture Notes in Computer Science, 1416:593– 602, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kocsis:1998:UTC

[1955] L. Kocsis and N. B. Szirbik. An unsupervised training connectionist network with lateral inhibition. Lecture Notes in Computer Science, 1416:603–611, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trinh:1998:TDL

[1956] T. B. Trinh, A. S. Bashi, and N. Dashpande. Temporal difference learning in Chinese chess. Lecture Notes in Computer Science, 1416:612–618, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coltell:1998:AOL

[1957] O. Coltell and J. M. Ordovas. Applying object logic programming to design computer strategies in gene scanning. Lecture Notes in Computer Science, 1416:619–627, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poggi:1998:ASR

[1958] A. Poggi and G. Golinelli. Automatic storing and retrieval of large collections of images. Lecture Notes in Computer Science, 1416:628–637, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coltell:1998:ACI

[1959] O. Coltell, J. Llach, P. Sanz, and C. Ruiz. Alternative communication interface for severely handicapped people based on a multimedia humancomputer interaction system. Lecture Notes in Computer Science, 1416:638– 647, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kadobayashi:1998:PME

[1960] R. Kadobayashi, K. Nishimoto, Y. Sumi, and K. Mase. Personalizing museum exhibition by mediating agents. Lecture Notes in Computer Science, 1416:648–667, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

EscolanoRuiz:1998:CPF

[1961] F. Escolano Ruiz, M. Cazorla, D. Gallardo, and F. Llorens. A combined probabilistic framework for learning gestures and actions. Lecture Notes in Computer Science, 1416:668–667, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barros:1998:DWS

[1962] B. Barros and F. Verdejo. Designing workspaces to support collaborative learning. Lecture Notes in Computer Science, 1416:668–677, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1998:DDS

[1963] Z. Xu, K. Ito, K. Jinno, and T. Kojiri. Development of decision support system for integrated water management in river basins. Lecture Notes in Computer Science, 1416:678–686, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roanes-Lozano:1998:AAM

[1964] E. Roanes-Lozano, L. M. Laita, and E. Roanes-Macias. An application of an AI methodology to railway interlocking systems using computer algebra. Lecture Notes in Computer Science, 1416:687–696, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AlonsoAmo:1998:IIS

[1965] F. Alonso Amo, J. M. Barreiro, J. P. Caraca-Valente, and C. Montes. Intelligent interpretation of strength data. Lecture Notes in Computer Science, 1416:697–706, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biermann:1998:HKB

[1966] J. Biermann. Hades — A knowledgebased system for message interpretation and situation determination. Lecture Notes in Computer Science, 1416: 707–716, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gavrilova:1998:WPV

[1967] T. Gavrilova and A. Voinov. Work in progress: Visual specification of knowledge bases. Lecture Notes in Computer Science, 1416:717–726, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coronado:1998:WKB

[1968] S. Coronado, A. Garcia-Beltran, J. A. Jaen, and R. Martinez. Webtutor, A knowledge based system for evaluation and tutorship. Lecture Notes in Computer Science, 1416:727–734, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jerinic:1998:CKP

[1969] L. Jerinic, V. Devedzic, and D. Radovic. Control knowledge and pedagogical aspects of the get-bits model. Lecture Notes in Computer Science, 1416:735– 744, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamdi:1998:IBA

[1970] Mohamed Salah Hamdi and K. Kaiser. Improving behavior arbitration using exploration and dynamic programming. Lecture Notes in Computer Science, 1416:745–754, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monostori:1998:ABA

[1971] L. Monostori and B. Kadar. Agent based architectures for mastering changes and disturbances in manufacturing. Lecture Notes in Computer Science, 1416:755–764, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monostori:1998:SCH

[1972] L. Monostori, J. Hornyak, C. Egresits, and Z. J. Viharos. Soft computing and hybrid AI approaches to intelligent manufacturing. Lecture Notes in Computer Science, 1416:765–774, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koskimaeki:1998:CSC

[1973] E. Koskimaeki, J. Goeoes, P. Kontkanen, and P. Myllymaeki. Comparing soft computing methods in prediction of manufacturing data. Lecture Notes in Computer Science, 1416:775–784, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcenac:1998:TEM

[1974] P. Marcenac, R. Courdier, S. Calderoni, and J. C. Soulie. Towards an emergence machine for complex systems simulations. Lecture Notes in Computer Science, 1416:785–794, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treuil:1998:SMA

[1975] J.-P. Treuil. Space models and agentbased universe architectures. Lecture Notes in Computer Science, 1416:795– 804, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ginot:1998:MGM

[1976] V. Ginot and C. Le Page. Mobidyc, a generic multi-agents simulator for modeling populations dynamics. Lecture Notes in Computer Science, 1416:805–814, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBeusekom:1998:DED

[1977] F. Van Beusekom, F. Brazier, P. Schipper, and J. Treur. Development of an ecological decision support system. Lecture Notes in Computer Science, 1416:815–825, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bousquet:1998:CCP

[1978] F. Bousquet, I. Bakam, H. Proton, and C. Le Page. Cormas: Commonpool resources and multi-agents systems. Lecture Notes in Computer Science, 1416:826–837, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grohmann:1998:DPM

[1979] A. Grohmann and R. Kopetzky. Dynamic process modelling and communication in environment information systems of the third generation. Lecture Notes in Computer Science, 1416:838– 848, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinez-Bejar:1998:LKB

[1980] R. Martinez-Bejar and F. Martin-Rubio. Landcape: A knowledge-based system for visual landscape assessment. Lecture Notes in Computer Science, 1416:849–856, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lassoued:1998:DPS

[1981] B. Lassoued and R. M’Hallah. Daily parking of sub way vehicles. Lecture Notes in Computer Science, 1416:857– 866, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wibig:1998:ANN

[1982] T. Wibig. The artificial neural networks in cosmic ray physics experiment; I.Total muon number estimation. Lecture Notes in Computer Science, 1416:867–876, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelAcebo:1998:SCF

[1983] E. Del Acebo, A. Oller, J. Lluis De La Rosa, and A. Ligeza. Static criteria for fuzzy systems quality evaluation. Lecture Notes in Computer Science, 1416:877–887, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1998:WKB

[1984] I. G. N. Smith, A. Oliver, and J. Oliphant. Wallaid: A knowledgebased system for the selection of earth retaining walls. Lecture Notes in Computer Science, 1416:888–895, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parra:1998:MPS

[1985] A. Parra, M. Rincon, J. R. Alvarez, and J. Mira. A modular and parametric structure for the substitution redesign of power plants control systems. Lecture Notes in Computer Science, 1416:896–906, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miftakov:1998:PPN

[1986] R. N. Miftakov and J. Christensen. A polysynaptic planar neural network as a model of the myenteric nervous plexus. Lecture Notes in Computer Science, 1416:907–917, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stein:1998:SNM

[1987] B. Stein and D. Curatolo. Selection of numerical methods in specific simulation applications. Lecture Notes in Computer Science, 1416:918–927, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skrjanc:1998:FAC

[1988] I. Skrjanc and D. Matko. Fuzzy adaptative control of the highly nonlinear heat-exchange plant. Lecture Notes in Computer Science, 1416:928–937, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chien:1998:AQS

[1989] Andrew A. Chien and Jae H. Kim. Approaches to quality of service in high-performance networks. Lecture Notes in Computer Science, 1417:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1417/14170001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1417/14170001.
pdf.

Fulgham:1998:IMC

[1990] Melanie L. Fulgham and Lawrence Snyder. Integrated multi-class routing. Lecture Notes in Computer Science, 1417:21–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1417/14170021.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1417/14170021.
pdf.

Hyatt:1998:CCW

[1991] Craig Hyatt and Dharma P. Agrawal.

Congestion control in the wormholerouted torus with clustering and delayed deflection. Lecture Notes in Computer Science, 1417:33–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1417/14170033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1417/14170033.
pdf.

Sivaram:1998:MIN

[1992] Rajeev Sivaram, Dhabaleswar K.

Panda, and Craig B. Stunkel. Multicasting in irregular networks with cutthrough switches using tree-based multidestination worms. Lecture Notes in Computer Science, 1417:39–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1417/14170039.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1417/14170039.
                                        pdf.

Garg:1998:CCC

[1993] Vivek Garg and David E. Schimmel. CCSIMD: A concurrent communication and computation framework for SIMD machines. Lecture Notes in Computer Science, 1417:55– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1417/14170055.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1417/14170055.
                                        pdf.

Boughton:1998:ASF

[1994] G. Andrew Boughton. Arctic switch fabric. Lecture Notes in Computer Science, 1417:65–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1417/14170065.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1417/14170065.pdf.

Moon:1998:SHS

[1995] Sung-Whan Moon, Padmanabhan Pillai, and Kang G. Shin. STREAMER: Hardware support for smoothed transmission of stored video over ATM. Lecture Notes in Computer Science, 1417: 75–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1417/14170075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1417/14170075.
pdf.

Miller:1998:PEH

[1996] Dianne Miller and Walid A. Najjar.

Preliminary evaluation of a hybrid deterministic/adaptive router. Lecture Notes in Computer Science, 1417:89– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1417/14170089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1417/14170089.
pdf.

May:1998:HPE

[1997] Phil May, Sek M. Chai, and D. Scott Wills. HiPER-P: An efficient, highperformance router for multicomputer interconnection networks. Lecture Notes in Computer Science, 1417:103– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1417/14170103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1417/14170103.
pdf.

Garcia:1998:SI

[1998] David Garcia and William Watson. ServerNet^{T M} II. Lecture Notes in Computer Science, 1417:119–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1417/14170119.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1417/14170119.
                                        pdf.

Lund:1998:ESS

[1999] Craig Lund. Embedded systems standards. Lecture Notes in Computer Science, 1417:137–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1417/14170137.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1417/14170137.pdf.

Stunkel:1998:CDC

[2000] Craig B. Stunkel. Challenges in the design of contemporary routers. Lecture Notes in Computer Science, 1417: 139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1417/14170139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1417/14170139.
                                        pdf.

Panda:1998:DHP

[2001] D. Panda. Designing high-performance communication subsystems: Top five problems to solve and five problems not to solve during the next five years. Lecture Notes in Computer Science, 1417: 153–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1417/14170153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1417/14170153.
pdf.

Byrd:1998:ECM

[2002] Gregory T. Byrd and Michael J. Flynn.

Evaluation of communication mechanisms in invalidate-based shared memory multiprocessors. Lecture Notes in Computer Science, 1417:159–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1417/14170159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1417/14170159.
pdf.

Dai:1998:HCW

[2003] Donglai Dai and Dhabaleswar K.

Panda. How can we design better networks for DSM systems? Lecture Notes in Computer Science, 1417:171– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1417/14170171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1417/14170171.
pdf.

Eicken:1998:IUN

[2004] Thorsten von Eicken. Integration of U-Net into Windows/NT. Lecture Notes in Computer Science, 1417:185– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1417/14170185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1417/14170185.
                                        pdf.

VonEicken:1998:IUN

[2005] T. Von Eicken. Integration of U-Net into Windows/NT (invited presentation). Lecture Notes in Computer Science, 1417:185–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smai:1998:DBF

[2006] A.-H. Smai and L.-E. Thorelli. Distance-based flow control in wormhole networks. Lecture Notes in Computer Science, 1417:189–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1417/14170189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1417/14170189.
                                        pdf.

Silla:1998:UVC

[2007] F. Silla and J. Duato. On the use of virtual channels in networks of workstations with irregular topology. Lecture Notes in Computer Science, 1417: 203–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1417/14170203.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1417/14170203.
                                        pdf.

Kesavan:1998:MSB

[2008] Ram Kesavan and Dhabaleswar K.

Panda. Multicasting on switch-based irregular networks using multi-drop path-based multidestination worms. Lecture Notes in Computer Science, 1417:217–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1417/14170217.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1417/14170217.
pdf.

Patel:1998:PPT

[2009] Chirag S. Patel, Sek M. Chai, Sudhakar Yalamanchili, and David E. Schimmel. Power/performance tradeoffs for direct networks. Lecture Notes in Computer Science, 1417:231– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1417/14170231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1417/14170231.
pdf.

McKenzie:1998:CDI

[2010] Neil R. McKenzie, Kevin Bolding, Carl Ebeling, and Lawrence Snyder. ChaosLAN: Design and implementation of a gigabit LAN using chaotic routing. Lecture Notes in Computer Science, 1417:247–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1417/14170247.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1417/14170247.
                                        pdf.

Yuan:1998:DTD

[2011] X. Yuan, R. Gupta, and R. Melhem. Does time-division multiplexing close the gap between memory and optical communication speeds? Lecture Notes in Computer Science, 1417:261– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1417/14170261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1417/14170261.
                                        pdf.

Warnakulasuriya:1998:MMB

[2012] Sugath Warnakulasuriya and Timothy Mark Pinkston. Modeling message blocking and deadlock in interconnection networks. Lecture Notes in Computer Science, 1417:275–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1417/14170275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1417/14170275.
                                        pdf.

Lopez:1998:RDF

[2013] P. López, J. M. Martínez, J. Duato, and F. Petrini. On the reduction of deadlock frequency by limiting message injection in wormhole networks. Lecture Notes in Computer Science, 1417: 295–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1417/14170295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1417/14170295.
pdf.

Junghanns:1998:SES

[2014] A. Junghanns and J. Schaeffer.

Sokoban: Evaluating standard singleagent search techniques in the presence of deadlock. Lecture Notes in Computer Science, 1418:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banks:1998:HIM

[2015] D. Banks, P. Van Beek, and A. Meisels.

A heuristic incremental modeling approach to course timetabling. Lecture Notes in Computer Science, 1418:16– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baltes:1998:PSR

[2016] J. Baltes. Planning strategy representation in DoLittle. Lecture Notes in Computer Science, 1418:30–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shin:1998:ELC

[2017] D.-G. Shin and L.-Y. Chu. Establishing logical connectivity between query keywords and database contents. Lecture Notes in Computer Science, 1418: 45–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barker:1998:TDT

[2018] K. Barker, S. Delisle, and S. Szpakowicz. Test-driving TANKA: Evaluating a semi-automatic system of text analysis for knowledge acquisition. Lecture Notes in Computer Science, 1418:60– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mineau:1998:OIM

[2019] G. W. Mineau, M. Lahboub, and J.-M. Beaulieu. An object indexing methodology as support to object recognition. Lecture Notes in Computer Science, 1418:72–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nock:1998:OAM

[2020] R. Nock and B. Esfandiari. Oracles and assistants: Machine learning applied to network supervision. Lecture Notes in Computer Science, 1418:86–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abu-Hakima:1998:CMA

[2021] S. Abu-Hakima, R. Liseano, and R. Impey. A common multi-agent testbed for diverse seamless personal information networking applications. Lecture Notes in Computer Science, 1418:99– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1998:HGA

[2022] J. Berger, M. Salois, and R. Begin. A hybrid genetic algorithm for the vehicle routing problem with time windows. Lecture Notes in Computer Science, 1418:114–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morrison:1998:MGD

[2023] J. Morrison and F. Oppacher. Maintaining genetic diversity in genetic algorithms through co-evolution. Lecture Notes in Computer Science, 1418:128– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiese:1998:EGO

[2024] K. Wiese and S. D. Goodwin. The effect of genetic operator probabilities and selection strategies on the performance of a genetic algorithm. Lecture Notes in Computer Science, 1418:139– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OReilly:1998:IED

[2025] U.-M. O’Reilly. The impact of external dependency in genetic programming primitives. Lecture Notes in Computer Science, 1418:154–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahn:1998:QCC

[2026] U. Hahn and K. Schnattinger. Quality control in the concept learning process. Lecture Notes in Computer Science, 1418:169–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ling:1998:GGL

[2027] C. X. Ling and B. Zhang. Grapheme generation in learning to read English words. Lecture Notes in Computer Science, 1418:184–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barker:1998:TBN

[2028] K. Barker. A trainable bracketer for noun modifiers. Lecture Notes in Computer Science, 1418:196–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hodges:1998:LLD

[2029] B. Hodges and L. Hodgson. Lessons learned in the development and implementation of a bilingual nationally accessible knowledge-based system. Lecture Notes in Computer Science, 1418:211–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donaldson:1998:SNA

[2030] T. Donaldson and R. Cohen. Selecting the next action with constraints. Lecture Notes in Computer Science, 1418: 220–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billings:1998:PTA

[2031] D. Billings, D. Papp, J. Schaeffer, and D. Szafron. Poker as a testbed for AI research. Lecture Notes in Computer Science, 1418:228–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knudsen:1998:FPT

[2032] B. R. Knudsen. Fault prediction in the telephone access loop using a neural network. Lecture Notes in Computer Science, 1418:239–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1998:LES

[2033] J. Zhang and H. J. Hamilton. Learning English syllabification rules. Lecture Notes in Computer Science, 1418: 246–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pang:1998:CTC

[2034] W. Pang and S. D. Goodwin. Characterizing tractable CSPs. Lecture Notes in Computer Science, 1418:259– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goncalves:1998:ARM

[2035] T. Goncalves and F. Moura-Pires. An attribute redundancy measure for clustering. Lecture Notes in Computer Science, 1418:273–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Draghici:1998:CVF

[2036] S. Draghici. On the complexity of VLSI-Friendly neural networks for classification problems. Lecture Notes in Computer Science, 1418:285–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebban:1998:SCS

[2037] M. Sebban and A. M. Landraud-Lamole. Strings clustering and statistical validation of clusters. Lecture Notes in Computer Science, 1418:298– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McGarity:1998:FPL

[2038] M. McGarity. Finding partitions for learning control of dynamic systems. Lecture Notes in Computer Science, 1418:310–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaib-Draa:1998:RMC

[2039] B. Chaib-Draa. A relational modeling of cognitive maps. Lecture Notes in Computer Science, 1418:322–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staab:1998:DCA

[2040] S. Staab and U. Hahn. Distance constraint arrays: A model for reasoning on intervals with qualitative and quantitative distances. Lecture Notes in Computer Science, 1418:334–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delgrande:1998:RTI

[2041] J. P. Delgrande and A. Gupta. Revising TimeGraph-II. Lecture Notes in Computer Science, 1418:349–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xiang:1998:TIJ

[2042] Y. Xiang. Temporally invariant junction tree for inference in dynamic Bayesian network. Lecture Notes in Computer Science, 1418:363–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1998:UTB

[2043] S. M. Brown, E. Santos, and S. B. Banks. Utility theory-based user models for intelligent interface agents. Lecture Notes in Computer Science, 1418: 378–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1998:HCM

[2044] J. Liu, K.-C. Chang, and J. Zhou. A hybrid convergent method for learning probabilistic networks. Lecture Notes in Computer Science, 1418:393– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perez:1998:RCF

[2045] E. Perez and L. Rendell. Relational concepts and the Fourier transform: An empirical study. Lecture Notes in Computer Science, 1418:411–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

An:1998:ELS

[2046] A. An and N. Cercone. ELEM2: A learning system for more accurate classifications. Lecture Notes in Computer Science, 1418:426–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Epstein:1998:SIB

[2047] S. L. Epstein and J. Shih. Sequential instance-based learning. Lecture Notes in Computer Science, 1418:442– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rios:1998:PIF

[2048] R. Rios and S. Matwin. Predicate invention from a few examples. Lecture Notes in Computer Science, 1418:455– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chess:1998:SIM

[2049] David M. Chess. Security issues in mobile code systems. Lecture Notes in Computer Science, 1419:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1419/14190001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1419/14190001.
pdf.

Riordan:1998:EKG

[2050] James Riordan and Bruce Schneier. Environmental key generation towards clueless agents. Lecture Notes in Computer Science, 1419:15–24, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1419/14190015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1419/14190015.
pdf;   http://www.counterpane.com/
clueless-agents.html.

Volpano:1998:LIM

[2051] Dennis Volpano and Geoffrey Smith. Language issues in mobile program security. Lecture Notes in Computer Science, 1419:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1419/14190025.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
1419/14190025.pdf.

Sander:1998:PMA

[2052] Tomas Sander and Christian F.

Tschudin. Protecting mobile agents against malicious hosts. Lecture Notes in Computer Science, 1419:44– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1419/14190044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1419/14190044.
pdf.

Necula:1998:SUA

[2053] George C. Necula and Peter Lee.

Safe, untrusted agents using proofcarrying code. Lecture Notes in Computer Science, 1419:61–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1419/14190061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1419/14190061.
pdf.

Hohl:1998:TLB

[2054] Fritz Hohl. Time limited blackbox security: Protecting mobile agents from malicious hosts. Lecture Notes in Computer Science, 1419:92–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1419/14190092.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1419/14190092.
                                        pdf.

Berkovits:1998:AMA

[2055] Shimshon Berkovits, Joshua D. Guttman, and Vipin Swarup. Authentication for mobile agents. Lecture Notes in Computer Science, 1419:114–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1419/14190114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1419/14190114.
                                        pdf.

Vigna:1998:CTM

[2056] Giovanni Vigna. Cryptographic traces for mobile agents. Lecture Notes in Computer Science, 1419:137–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1419/14190137.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1419/14190137.
                                        pdf.

Gray:1998:DSM

[2057] Robert S. Gray, David Kotz, George Cybenko, and Daniela Rus. D’agents: Security in a multiple-language, mobile-agent system. Lecture Notes in Computer Science, 1419:154–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1419/14190154.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1419/14190154.
pdf.

Karjoth:1998:SMA

[2058] Günter Karjoth, Danny B. Lange, and Mitsuru Oshima. A security model for aglets. Lecture Notes in Computer Science, 1419:188–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1419/14190188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1419/14190188.
pdf.

Gong:1998:SSG

[2059] Li Gong and Roland Schemers. Signing, sealing, and guarding Java^{T M} objects. Lecture Notes in Computer Science, 1419:206–216, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1419/14190206.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1419/14190206.pdf.

Ousterhout:1998:STS

[2060] John K. Ousterhout, Jacob Y. Levy, and Brent B. Welch. The Safe-Tcl security model. Lecture Notes in Computer Science, 1419:217–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1419/14190217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1419/14190217.
pdf.

DePaoli:1998:WBS

[2061] F. De Paoli, A. L. Dos Santos, and R. A. Kemmerer. Web browsers and security. Lecture Notes in Computer Science, 1419:235–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paoli:1998:WBS

[2062] Flavio De Paoli, Andre L. Dos Santos, and Richard A. Kemmerer. Web browsers and security. Lecture Notes in Computer Science, 1419:235–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1419/14190235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1419/14190235.
                                        pdf.

Valk:1998:PNT

[2063] Rüdiger Valk. Petri nets as token objects — an introduction to elementary object nets. Lecture Notes in Computer Science, 1420:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1420/14200001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1420/14200001.
                                        pdf.

Arnold:1998:SPT

[2064] André Arnold. Synchronized products of transition systems and their analysis. Lecture Notes in Computer Science, 1420:26–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1420/14200026.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1420/14200026.
pdf.

Chiola:1998:MAE

[2065] Giovanni Chiola. Manual and automatic exploitation of symmetries in SPN models. Lecture Notes in Computer Science, 1420:28–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1420/14200028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1420/14200028.
pdf.

Marsan:1998:DHB

[2066] Marco Ajmone Marsan, Carla-Fabiana Chiasserini, and Andrea Fumagalli. Dimensioning handover buffers in wireless ATM networks with GSPN models. Lecture Notes in Computer Science, 1420:44–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1420/14200044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1420/14200044.
pdf.

Ezpeleta:1998:CWS

[2067] J. Ezpeleta, F. Garcia-Vallés, and J. M. Colom. A class of well structured Petri nets for flexible manufacturing systems. Lecture Notes in Computer Science, 1420:64–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1420/14200064.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1420/14200064.
                                        pdf.

Fisler:1998:MMC

[2068] Kathi Fisler and Claude Girault. Modelling and model checking a distributed shared memory consistency protocol. Lecture Notes in Computer Science, 1420:84–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1420/14200084.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1420/14200084.
                                        pdf.

Kristensen:1998:FSS

[2069] Lars Michael Kristensen and Antti Valmari. Finding stubborn sets of coloured Petri nets without unfolding. Lecture Notes in Computer Science, 1420:104– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1420/14200104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1420/14200104.
                                        pdf.

Varpaaniemi:1998:SSV

[2070] Kimmo Varpaaniemi. On stubborn sets in the verification of linear time temporal properties. Lecture Notes in Computer Science, 1420:124–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1420/14200124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1420/14200124.
pdf.

Fleischhack:1998:CPN

[2071] Hans Fleischhack and Bernd Grahlmann.

A compositional Petri net semantics for SDL. Lecture Notes in Computer Science, 1420:144–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1420/14200144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1420/14200144.
pdf.

Hesketh:1998:ADE

[2072] Martin Hesketh and Maciej Koutny. An axiomatisation of duplication equivalence in the Petri box calculus. Lecture Notes in Computer Science, 1420: 165–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1420/14200165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1420/14200165.
pdf.

Bihler:1998:ETP

[2073] Elmar Bihler and Walter Vogler. Efficiency of token-passing MUTEX-Solutions — some experiments. Lecture Notes in Computer Science, 1420: 185–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1420/14200185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1420/14200185.
                                        pdf.

Stehr:1998:TC

[2074] Mark-Oliver Stehr. Thinking in cycles. Lecture Notes in Computer Science, 1420:205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1420/14200205.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1420/14200205.
                                        pdf.

Buchholz:1998:IDA

[2075] Peter Buchholz. Iterative decomposition and aggregation of labeled GSPNs. Lecture Notes in Computer Science, 1420:226–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1420/14200226.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1420/14200226.
                                        pdf.

Gribaudo:1998:USP

[2076] Marco Gribaudo and Matteo Sereno. On the use of structural Petri net analysis for studying product form equilibrium distributions of queueing networks with blocking. Lecture Notes in Computer Science, 1420:246– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1420/14200246.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1420/14200246.
pdf.

Bertino:1998:DDF

[2077] E. Bertino, G. Chiola, and L. V.

Mancini. Deadlock detection in the face of transaction and data dependencies. Lecture Notes in Computer Science, 1420:266–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1420/14200266.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1420/14200266.
pdf.

Langner:1998:PNB

[2078] Peter Langner, Christoph Schneider, and Joachim Wehler. Petri net based certification of eventdriven process chains. Lecture Notes in Computer Science, 1420:286–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1420/14200286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1420/14200286.
pdf.

Benzaken:1998:MNC

[2079] Véronique Benzaken, Nicolas Hugon, Hanna Klaudel, Elisabeth Pelz, and Robert-C. Riemann. M-Net calculus based semantics for triggers. Lecture Notes in Computer Science, 1420:306– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1420/14200306.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1420/14200306.
                                        pdf.

Gaeta:1998:SAS

[2080] Rossano Gaeta and Marco Ajmone Marsan. SWN analysis and simulation of large knockout ATM switches. Lecture Notes in Computer Science, 1420: 326–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1420/14200326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1420/14200326.
                                        pdf.

Kindler:1998:FAN

[2081] Ekkart Kindler and Hagen Völzer. Flexibility in algebraic nets. Lecture Notes in Computer Science, 1420:345– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1420/14200345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1420/14200345.
                                        pdf.

Kindler:1998:ETL

[2082] Ekkart Kindler and Tobias Vesper. ESTL: A temporal logic for events and states. Lecture Notes in Computer Science, 1420:365–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1420/14200365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1420/14200365.
pdf.

Anonymous:1998:AIg

[2083] Anonymous. Author index. Lecture Notes in Computer Science, 1420:385– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1420/14200385.pdf.

Anonymous:1998:P

[2084] Anonymous.  Preface.  Lecture

Notes in Computer Science, 1420: V, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1420/1420000v.pdf.

Anonymous:1998:TC

[2085] Anonymous. Table of contents. Lecture Notes in Computer Science, 1420: VII, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1420/14200vii.pdf.

Pfenning:1998:ITR

[2086] Frank Pfenning. Invited talk — reasoning about deductions in linear logic. Lecture Notes in Computer Science, 1421:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1421/14210001.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/1421/14210001.
pdf.

Fleuriot:1998:CNA

[2087] Jacques D. Fleuriot and Lawrence C. Paulson. A combination of nonstandard analysis and geometry theorem proving, with application to Newton’s Principia. Lecture Notes in Computer Science, 1421:3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210003.
pdf.

Fevre:1998:PGT

[2088] Stéphane Fèvre and Dongming Wang. Proving geometric theorems using Clifford algebra and rewrite rules. Lecture Notes in Computer Science, 1421:17– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210017.
pdf.

Fuchs:1998:SDS

[2089] Marc Fuchs. System description: Similarity-based lemma generation for model elimination. Lecture Notes in Computer Science, 1421:33–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210033.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1421/14210033.
pdf.

Arts:1998:SDV

[2090] Thomas Arts, Mads Dam, Lars åke Fredlund, and Dilian Gurov. System description: Verification of distributed Erlang programs. Lecture Notes in Computer Science, 1421:38– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210038.
pdf.

Fuchs:1998:SDC

[2091] Marc Fuchs and Andreas Wolf. System description: Cooperation in model elimination: CPTHEO. Lecture Notes in Computer Science, 1421:42– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210042.
pdf.

Gore:1998:SDC

[2092] Rajeev Goré, Joachim Posegga, Andrew Slater, and Harald Vogt. System description: card T^{A}P: The first theorem prover on a smart card. Lecture Notes in Computer Science, 1421:47– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1421/14210047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1421/14210047.
                                        pdf.

Beckert:1998:SDL

[2093] Bernhard Beckert and Rajeev Goré. System description: leanK 2.0. Lecture Notes in Computer Science, 1421: 51–55, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1421/14210051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1421/14210051.
                                        pdf.

Benzmueller:1998:EHO

[2094] C. Benzmueller and M. Kohlhase. Extensional higher-order resolution. Lecture Notes in Computer Science, 1421: 56–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benzmuller:1998:EHO

[2095] Christoph Benzmüller and Michael Kohlhase. Extensional higherorder resolution. Lecture Notes in Computer Science, 1421:56–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1421/14210056.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1421/14210056.
                                        pdf.

Pagano:1998:XRE

[2096] Bruno Pagano. X.R.S.: Explicit reduction systems — A first-order calculus for higher-order calculi. Lecture Notes in Computer Science, 1421:72– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210072.
pdf.

Boudet:1998:ACE

[2097] Alexandre Boudet and Evelyne Contejean. About the confluence of equational pattern rewrite systems. Lecture Notes in Computer Science, 1421:88– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210088.
pdf.

Beeson:1998:ULC

[2098] Michael Beeson. Unification in lambdacalculi with if-then-else. Lecture Notes in Computer Science, 1421:103–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210103.
pdf.

Peltier:1998:SDE

[2099] Nicolas Peltier. System description: An equational constraints solver. Lecture Notes in Computer Science, 1421:119– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1421/14210119.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1421/14210119.
                                        pdf.

Mazure:1998:SDC

[2100] Bertrand Mazure, Lakhdar Saïs, and Éric Grégoire. System description: CRIL platform for SAT. Lecture Notes in Computer Science, 1421:124– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1421/14210124.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1421/14210124.
                                        pdf.

Richardson:1998:SDP

[2101] Julian Richardson, Alan Smaill, and Ian Green. System description: Proof planning in higher-order logic with λClam. Lecture Notes in Computer Science, 1421:129–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1421/14210129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1421/14210129.
                                        pdf.

Slind:1998:SDI

[2102] Konrad Slind, Mike Gordon, Richard Boulton, and Alan Bundy. System description: An interface between CL^{A}M and HOL. Lecture Notes in Computer Science, 1421:134– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210134.
pdf.

Benzmueller:1998:SDL

[2103] C. Benzmueller and M. Kohlhase. System description: LEO — A higherorder theorem prover. Lecture Notes in Computer Science, 1421:139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benzmuller:1998:SDE

[2104] Christoph Benzmüller and Michael Kohlhase. System description: Leo — A higher-order theorem prover. Lecture Notes in Computer Science, 1421: 139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210139.
pdf.

Waldmann:1998:SDT

[2105] Uwe Waldmann. Superposition for divisible torsion-free Abelian groups. Lecture Notes in Computer Science, 1421:144–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1421/14210144.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1421/14210144.
pdf.

Waldrnann:1998:SDT

[2106] U. Waldrnann. Superposition for divisible torsion-free Abelian groups. Lecture Notes in Computer Science, 1421: 144–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1998:SBS

[2107] Leo Bachmair and Harald Ganzinger. Strict basic superposition. Lecture Notes in Computer Science, 1421:160– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1421/14210160.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1421/14210160.
                                        pdf.

Bachmair:1998:EET

[2108] Leo Bachmair, Harald Ganzinger, and Andrei Voronkov. Elimination of equality via transformation with ordering constraints. Lecture Notes in Computer Science, 1421:175–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1421/14210175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1421/14210175.
                                        pdf.

Bachrnair:1998:EET

[2109] L. Bachrnair, H. Ganzinger, and A. Voronkov. Elimination of equality via transformation with ordering constraints. Lecture Notes in Computer Science, 1421:175–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNivelle:1998:RDP

[2110] H. De Nivelle. A resolution decision procedure for the guarded fragment. Lecture Notes in Computer Science, 1421:191–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nivelle:1998:RDP

[2111] Hans de Nivelle. A resolution decision procedure for the guarded fragment. Lecture Notes in Computer Science, 1421:191–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1421/14210191.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1421/14210191.
pdf.

Ohlbach:1998:CHS

[2112] Hans Jürgen Ohlbach. Combining Hilbert style and semantic reasoning in a resolution framework. Lecture Notes in Computer Science, 1421:205– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210205.
pdf.

Kaufmann:1998:ITA

[2113] Matt Kaufmann. Invited talk — ACL2 support for verification projects. Lecture Notes in Computer Science, 1421: 220–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210220.
pdf.

Oliart:1998:FAU

[2114] Alberto Oliart and Wayne Snyder. A fast algorithm for uniform semi-unification. Lecture Notes in Computer Science, 1421:239–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210239.
pdf.

Brauburger:1998:TAI

[2115] Jürgen Brauburger and Jürgen Giesl. Termination analysis by inductive evaluation. Lecture Notes in Computer Science, 1421:254–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1421/14210254.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1421/14210254.pdf.

Crary:1998:AFI

[2116] Karl Crary. Admissibility of fixpoint induction over partial types. Lecture Notes in Computer Science, 1421:270– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1421/14210270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210270.
pdf.

Schuermann:1998:ATP

[2117] C. Schuermann and F. Pfenning. Automated theorem proving in a simple meta-logic for LF. Lecture Notes in Computer Science, 1421:286–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schurmann:1998:ATP

[2118] Carsten Schürmann and Frank Pfenning. Automated theorem proving in a simple meta-logic for LF. Lecture Notes in Computer Science, 1421:286– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210286.
pdf.

Pnueli:1998:ITD

[2119] Amir Pnueli. Invited talk — deductive vs. model-theoretic approaches to formal verification. Lecture Notes in Computer Science, 1421:301–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210301.
pdf.

Malik:1998:ADF

[2120] Robi Malik. Automated deduction of finite-state control programs for reactive systems. Lecture Notes in Computer Science, 1421:302–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210302.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210302.
pdf.

Kreitz:1998:PED

[2121] Christoph Kreitz, Mark Hayden, and Jason Hickey. A proof environment for the development of group communication systems. Lecture Notes in Computer Science, 1421:317–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210317.
pdf.

Ohta:1998:RBN

[2122] Yoshihiko Ohta, Katsumi Inoue, and Ryuzo Hasegawa. On the relationship between non-Horn magic sets and relevancy testing. Lecture Notes in Computer Science, 1421:333–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210333.
pdf.

Thery:1998:CVB

[2123] Laurent Théry. A certified version of Buchberger’s algorithm. Lecture Notes in Computer Science, 1421:349– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210349.
pdf.

Bishop:1998:SID

[2124] Matthew Bishop and Peter B.

Andrews. Selectively instantiating definitions. Lecture Notes in Computer Science, 1421:365–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210365.
pdf.

Letz:1998:UMP

[2125] Reinhold Letz. Using matings for pruning connection tableaux. Lecture Notes in Computer Science, 1421:381– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210381.
pdf.

Nonnengart:1998:GSC

[2126] Andreas Nonnengart, Georg Rock, and Christoph Weidenbach. On generating small clause normal forms. Lecture Notes in Computer Science, 1421: 397–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210397.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210397.
pdf.

Horton:1998:RAC

[2127] J. D. Horton and Bruce Spencer. Rank/activity: A canonical form for binary resolution. Lecture Notes in Computer Science, 1421:412–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210412.
pdf.

Tammet:1998:TES

[2128] Tanel Tammet. Towards efficient subsumption. Lecture Notes in Computer Science, 1421:427–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1421/14210427.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1421/14210427.
pdf.

Harel:1998:AD

[2129] David Harel. On the aesthetics of diagrams. Lecture Notes in Computer Science, 1422:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220001.
pdf.

Hughes:1998:GM

[2130] John Hughes. Generalising monads.

Lecture Notes in Computer Science, 1422:6–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1422/14220006.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1422/14220006.
pdf.

Karger:1998:PRC

[2131] Burghard von Karger. A proof rule for control loops. Lecture Notes in Computer Science, 1422:7– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220007.
pdf.

VonKarger:1998:PRC

[2132] B. Von Karger. A proof rule for control loops. Lecture Notes in Computer Science, 1422:7–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berghammer:1998:RAD

[2133] Rudolf Berghammer, Burghard von Karger, and Andreas Wolf. Relationalgebraic derivation of spanning tree algorithms. Lecture Notes in Computer Science, 1422:23–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220023.
pdf.

Bijlsma:1998:APD

[2134] A. Bijlsma. An application of program derivation techniques to 18thcentury mathematics. Lecture Notes in Computer Science, 1422:44–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220044.
pdf.

Bird:1998:ND

[2135] Richard Bird and Lambert Meertens. Nested datatypes. Lecture Notes in Computer Science, 1422:52–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220052.
pdf.

Bonsangue:1998:AOO

[2136] Marcello M. Bonsangue, Joost N. Kok, and Kaisa Sere. An approach to objectorientation in action systems. Lecture Notes in Computer Science, 1422:68– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220068.
pdf.

Bosangue:1998:AOO

[2137] M. M. Bosangue, J. N. Kok, and K. Sere. An approach to objectorientation in action systems. Lecture Notes in Computer Science, 1422:68– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brunn:1998:LGT

[2138] Thomas Brunn, Bernhard Möller, and Martin Russling. Layered graph traversals and Hamiltonian path problems — an algebraic approach. Lecture Notes in Computer Science, 1422:96– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220096.
pdf.

Dierks:1998:UFC

[2139] Henning Dierks and Michael Schenke.

A unifying framework for correct program construction. Lecture Notes in Computer Science, 1422:122–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1422/14220122.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1422/14220122.
                                        pdf.

Dijkstra:1998:CCB

[2140] Rutger M. Dijkstra. Computation calculus — bridging a formalization gap. Lecture Notes in Computer Science, 1422:151–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1422/14220151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1422/14220151.
                                        pdf.

Feijen:1998:EDA

[2141] W. H. J. Feijen, A. J. M. Van Gasteren, and B. Schieder. An elementary derivation of the alternating bit protocol. Lecture Notes in Computer Science, 1422:175–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1422/14220175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1422/14220175.
                                        pdf.

Fidge:1998:STM

[2142] C. J. Fidge, I. J. Hayes, A. P. Martin, and A. K. Wabenhorst. A settheoretic model for real-time specification and reasoning. Lecture Notes in Computer Science, 1422:188–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1422/14220188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220188.
pdf.

Gibbons:1998:PDA

[2143] Jeremy Gibbons. Polytypic downwards accumulations. Lecture Notes in Computer Science, 1422:207–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220207.
pdf.

Josephs:1998:FDL

[2144] Mark B. Josephs. Formal derivation of a loadable asynchronous counter. Lecture Notes in Computer Science, 1422: 234–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220234.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220234.
pdf.

Leino:1998:SAS

[2145] K. Rustan M. Leino and Rajeev Joshi. A semantic approach to secure information flow. Lecture Notes in Computer Science, 1422:254–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220254.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1422/14220254.
                                        pdf.

Rustan:1998:SAS

[2146] K. Rustan, M. Leino, and R. Joshi. A semantic approach to secure information flow. Lecture Notes in Computer Science, 1422:254–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manohar:1998:SEC

[2147] Rajit Manohar and Alain J. Martin. Slack elasticity in concurrent computing. Lecture Notes in Computer Science, 1422:272–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1422/14220272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1422/14220272.
                                        pdf.

Naumann:1998:BFO

[2148] David A. Naumann. Beyond fun: Order and membership in polytypic imperative programming. Lecture Notes in Computer Science, 1422:286– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1422/14220286.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1422/14220286.
                                        pdf.

Soerensen:1998:CPT

[2149] M. H. B. Soerensen. Convergence of program transformers in the metric space of trees. Lecture Notes in Computer Science, 1422:315–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sorensen:1998:CPT

[2150] Morten Heine B. Sørensen. Convergence of program transformers in the metric space of trees. Lecture Notes in Computer Science, 1422:315– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220315.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220315.
pdf.

Tatsuta:1998:RMC

[2151] Makoto Tatsuta. Realizability of monotone coinductive definitions and its application to program synthesis. Lecture Notes in Computer Science, 1422: 338–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1422/14220338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1422/14220338.
pdf.

Vaccari:1998:CRR

[2152] Matteo Vaccari and Roland Backhouse.

Calculating a round-robin scheduler. Lecture Notes in Computer Science, 1422:365–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1422/14220365.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/1422/14220365.
                                        pdf.

Anonymous:1998:AIh

[2153] Anonymous. Author index. Lecture Notes in Computer Science, 1422:383– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1422/1422auth.pdf.

Elkies:1998:SCC

[2154] Noam D. Elkies. Shimura curve computations. Lecture Notes in Computer Science, 1423:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1423/14230001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1423/14230001.pdf.

Boneh:1998:DDH

[2155] Dan Boneh. The decision Diffie– Hellman problem. Lecture Notes in Computer Science, 1423:48–63, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230048.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230048.
                                        pdf;   http://theory.stanford.edu/
                                        ~dabo/papers/DDH.ps.gz.

Cesari:1998:PIS

[2156] Giovanni Cesari. Parallel implementation of Schönhage’s integer GCD algorithm. Lecture Notes in Computer Science, 1423:64–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230064.
pdf.

Vallee:1998:CAB

[2157] Brigitte Vallée. The complete analysis of the binary Euclidean algorithm. Lecture Notes in Computer Science, 1423: 77–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230077.
pdf.

Mihailescu:1998:CPP

[2158] Preda Mihailescu. Cyclotomy primality proving — recent developments. Lecture Notes in Computer Science, 1423: 95–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230095.
pdf.

Morain:1998:PPU

[2159] F. Morain. Primality proving using elliptic curves: An update. Lecture Notes in Computer Science, 1423:111– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1423/14230111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230111.
                                        pdf.

Bernstein:1998:BSI

[2160] Daniel J. Bernstein. Bounding smooth integers (extended abstract). Lecture Notes in Computer Science, 1423:128– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230128.
                                        pdf.

Mingzhi:1998:FNF

[2161] Zhang Mingzhi. Factorization of the numbers of the form m^{3} + c_{2} m^{2} + c_{1} m + c_{0}. Lecture Notes in Computer Science, 1423:131–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230131.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230131.
                                        pdf.

Murphy:1998:MYN

[2162] Brian Murphy. Modelling the yield of number field sieve polynomials. Lecture Notes in Computer Science, 1423: 137–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230137.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1423/14230137.
pdf.

Nguyen:1998:MLS

[2163] Phong Nguyen. A Montgomerylike square root for the number field sieve. Lecture Notes in Computer Science, 1423:151–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230151.
pdf.

Galway:1998:RBH

[2164] William F. Galway. Robert Bennion’s “Hopping Sieve”. Lecture Notes in Computer Science, 1423:169–178, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230169.
pdf.

Sorenson:1998:TTS

[2165] Jonathan P. Sorenson. Trading time for space in prime number sieves. Lecture Notes in Computer Science, 1423: 179–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230179.
pdf.

Deshouillers:1998:DSB

[2166] Jean-Marc Deshouillers, François Hennecart, and Bernard Landreau. Do sums of 4 biquadrates have a positive density? Lecture Notes in Computer Science, 1423:196–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230196.
pdf.

Deshouillers:1998:NER

[2167] J.-M. Deshouillers, H. J. J. Te Riele, and Y. Saouter. New experimental results concerning the Goldbach conjecture. Lecture Notes in Computer Science, 1423:204–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230204.
pdf.

Gordon:1998:DAS

[2168] Daniel M. Gordon and Gene Rodemich. Dense admissible sets. Lecture Notes in Computer Science, 1423: 216–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230216.
pdf.

Panario:1998:AAS

[2169] Daniel Panario, Xavier Gourdon, and Philippe Flajolet. An analytic approach to smooth polynomials over finite fields. Lecture Notes in Computer Science, 1423:226–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230226.
pdf.

Boneh:1998:GPT

[2170] Dan Boneh and Jeremy Horwitz. Generating a product of three primes with an unknown factorization. Lecture Notes in Computer Science, 1423:237– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230237.
pdf.

DeWin:1998:PSS

[2171] E. De Win, S. Mister, B. Preneel, and M. Wiener. On the performance of signature schemes based on elliptic curves. Lecture Notes in Computer Science, 1423:252–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Win:1998:PSS

[2172] Erik De Win, Serge Mister, Bart Preneel, and Michael Wiener. On the performance of signature schemes based on elliptic curves. Lecture Notes in Computer Science, 1423:252– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230252.
pdf.

Hoffstein:1998:NRB

[2173] Jeffrey Hoffstein, Jill Pipher, and Joseph H. Silverman. NTRU: A ringbased public key cryptosystem. Lecture Notes in Computer Science, 1423: 267–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230267.
pdf.

Young:1998:FLP

[2174] Adam Young and Moti Yung. Finding length-3 positive Cunningham chains and their cryptographic significance. Lecture Notes in Computer Science, 1423:289–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1423/14230289.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1423/14230289.
pdf.

Neis:1998:RIA

[2175] Stefan Neis. Reducing ideal arithmetic to linear algebra problems. Lecture Notes in Computer Science, 1423:299– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230299.
pdf.

Semaev:1998:ELR

[2176] I. A. Semaev. Evaluation of linear relations between vectors of a lattice in Euclidean space. Lecture Notes in Computer Science, 1423:311– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230311.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230311.
pdf.

Wetzel:1998:EPB

[2177] Susanne Wetzel. An efficient parallel block-reduction algorithm. Lecture Notes in Computer Science, 1423:323– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230323.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230323.
pdf.

Haible:1998:FME

[2178] Bruno Haible and Thomas Papanikolaou. Fast multiprecision evaluation of series of rational numbers. Lecture Notes in Computer Science, 1423:338– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230338.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230338.
                                        pdf.

Teske:1998:PCC

[2179] E. Teske and H. C. Williams. A problem concerning a character sum — extended abstract. Lecture Notes in Computer Science, 1423:351–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230351.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230351.
                                        pdf.

Poorten:1998:FPS

[2180] Alf van der Poorten. Formal power series and their continued fraction expansion. Lecture Notes in Computer Science, 1423:358–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230358.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230358.
                                        pdf.

vanderPoorten:1998:FPS

[2181] Alf van der Poorten. Formal power series and their continued fraction expansion. Lecture Notes in Computer Science, 1423:358–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1423/14230358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230358.
pdf.

Cohen:1998:IOF

[2182] Henri Cohen, Francisco Diaz y Diaz, and Michel Olivier. Imprimitive octic fields with small discriminants. Lecture Notes in Computer Science, 1423: 372–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230372.
pdf.

Cohen:1998:TTC

[2183] Henri Cohen, Francisco Diaz y Diaz, and Michel Olivier. A table of totally complex number fields of small discriminants. Lecture Notes in Computer Science, 1423:381–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230381.
pdf.

deSmit:1998:GAE

[2184] Bart de Smit. Generating arithmetically equivalent number fields with elliptic curves. Lecture Notes in Computer Science, 1423:392–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230392.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230392.
                                        pdf.

Smit:1998:GAE

[2185] Bart de Smit. Generating arithmetically equivalent number fields with elliptic curves. Lecture Notes in Computer Science, 1423:392–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230392.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230392.
                                        pdf.

Dummit:1998:CLT

[2186] David S. Dummit and Brett A. Tangedal. Computing the lead term of an Abelian L-function. Lecture Notes in Computer Science, 1423:400– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230400.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230400.
                                        pdf.

Jones:1998:TAT

[2187] John W. Jones and David P. Roberts. Timing analysis of targeted hunter searches. Lecture Notes in Computer Science, 1423:412–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1423/14230412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230412.
pdf.

Martinet:1998:SMR

[2188] Jacques Martinet. On successive minima of rings of algebraic integers. Lecture Notes in Computer Science, 1423: 424–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230424.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230424.
pdf.

Cohen:1998:CRQ

[2189] Henri Cohen, Francisco Diaz y Diaz, and Michel Olivier. Computation of relative quadratic class groups. Lecture Notes in Computer Science, 1423: 433–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230433.
pdf.

Gee:1998:GCF

[2190] Alice Gee and Peter Stevenhagen. Generating class fields using Shimura reciprocity. Lecture Notes in Computer Science, 1423:441–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1423/14230441.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
1423/14230441.pdf.

Holden:1998:IPN

[2191] Joshua Holden. Irregularity of prime numbers over real quadratic fields. Lecture Notes in Computer Science, 1423: 454–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230454.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230454.
pdf.

Jacobson:1998:ERC

[2192] Michael J. Jacobson, Jr. Experimental results on class groups of real quadratic fields (extended abstract). Lecture Notes in Computer Science, 1423:463– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230463.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230463.
pdf.

Louboutin:1998:CRC

[2193] Stéphane Louboutin. Computation of relative class numbers of imaginary cyclic fields of 2-power degrees. Lecture Notes in Computer Science, 1423: 475–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230475.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230475.
pdf.

Bluher:1998:FGE

[2194] Antonia W. Bluher. Formal groups, elliptic curves, and some theorems of couveignes. Lecture Notes in Computer Science, 1423:482–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230482.
pdf.

Djabri:1998:CDI

[2195] Z. Djabri and N. P. Smart. A comparison of direct and indirect methods for computing selmer groups of an elliptic curve. Lecture Notes in Computer Science, 1423:502–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230502.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230502.
pdf.

Huang:1998:AAC

[2196] Ming-Deh Huang and Yiu-Chung Wong. An algorithm for approximate counting of points on algebraic sets over finite fields. Lecture Notes in Computer Science, 1423:514– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230514.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230514.
pdf.

Petho:1998:IPE

[2197] Attila Pethö, Emanuel Herrmann, and Horst G. Zimmer. S-integral points on elliptic curves and Fermat’s triple equations. Lecture Notes in Computer Science, 1423:528–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230528.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230528.
                                        pdf.

Teske:1998:SPR

[2198] Edlyn Teske. Speeding up Pollard’s rho method for computing discrete logarithms. Lecture Notes in Computer Science, 1423:541–554, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230541.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230541.
                                        pdf.

Niederreiter:1998:GMC

[2199] Harald Niederreiter and Chaoping Xing. A general method of constructing global function fields with many rational places. Lecture Notes in Computer Science, 1423:555–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230555.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230555.
                                        pdf.

Paulus:1998:LBR

[2200] Sachar Paulus. Lattice basis reduction in function fields. Lecture Notes in Computer Science, 1423:567– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230567.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230567.
pdf.

Paulus:1998:CRI

[2201] Sachar Paulus and Andreas Stein.

Comparing real and imaginary arithmetics for divisor class groups of hyperelliptic curves. Lecture Notes in Computer Science, 1423:576–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230576.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230576.
pdf.

Scheidler:1998:UCP

[2202] Renate Scheidler and Andreas Stein.

Unit computation in purely cubic function fields of unit rank 1. Lecture Notes in Computer Science, 1423:592– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1423/14230592.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1423/14230592.
pdf.

Stein:1998:IMC

[2203] Andreas Stein and Hugh C. Williams. An improved method of computing the regulator of a real quadratic function field. Lecture Notes in Computer Science, 1423:607–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1423/14230607.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1423/14230607.pdf.

Zuccherato:1998:EBE

[2204] Robert J. Zuccherato. The equivalence between elliptic curve and quadratic function field discrete logarithms in characteristic 2. Lecture Notes in Computer Science, 1423:621– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1423/14230621.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1423/14230621.
                                        pdf.

Anonymous:1998:AIi

[2205] Anonymous. Author index. Lecture Notes in Computer Science, 1423:639– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1423/1423auth.pdf.

Kloesgen:1998:DAP

[2206] W. Kloesgen. Deviation and association patterns for subgroup mining in temporal, spatial, and textual data bases. Lecture Notes in Computer Science, 1424:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klosgen:1998:DAP

[2207] Willi Klösgen. Deviation and association patterns for subgroup mining in temporal, spatial, and textual data bases. Lecture Notes in Computer Science, 1424:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240001.
pdf.

Marcus:1998:PHG

[2208] Solomon Marcus. The paradox of the heap of grains in respect to roughness, fuzziness, and negligibility. Lecture Notes in Computer Science, 1424: 19–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240019.
pdf.

Marek:1998:RSW

[2209] V. Wiktor Marek and Mirosław Truszczynski. Rough sets — what are they about? Lecture Notes in Computer Science, 1424:24–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1424/14240024.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240024.
                                        pdf.

Pawlak:1998:RAD

[2210] Zdzisław Pawlak. Reasoning about data — A rough set perspective. Lecture Notes in Computer Science, 1424: 25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240025.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240025.
                                        pdf.

Zadeh:1998:IGC

[2211] Lotfi A. Zadeh. Information granulation and its centrality in human and machine intelligence. Lecture Notes in Computer Science, 1424:35– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240035.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240035.
                                        pdf.

Grzymala-Busse:1998:CSU

[2212] Jerzy W. Grzymala-Busse and Xihong Zou. Classification strategies using certain and possible rules. Lecture Notes in Computer Science, 1424:37– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240037.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1424/14240037.
pdf.

Miro-Julia:1998:WBO

[2213] Jose Miro-Julia and José Miró.

Well-behaviored operations for approximate sets. Lecture Notes in Computer Science, 1424:45–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240045.
pdf.

Slezak:1998:SFR

[2214] Dominik Slezak. Searching for frequential reducts in decision tables with uncertain objects. Lecture Notes in Computer Science, 1424:52–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240052.
pdf.

Greco:1998:NRS

[2215] Salvatore Greco, Benedetto Matarazzo, and Roman Slowinski. A new rough set approach to multicriteria and multiattribute classification. Lecture Notes in Computer Science, 1424:60– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240060.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1424/14240060.
                                        pdf.

Cwik:1998:HMM

[2216] J. Cwik and J. Koronacki. A heuristic method of model choice for nonlinear regression. Lecture Notes in Computer Science, 1424:68–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240068.
                                        pdf.

Ciok:1998:HNS

[2217] A. Ciok, T. Kowalczyk, and E. Pleszczynska. How a new statistical infrastructure induced a new computing trend in data analysis. Lecture Notes in Computer Science, 1424:75– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240075.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240075.
                                        pdf.

Dassow:1998:SRN

[2218] Jürgen Dassow. Some remarks on networks of parallel language processors. Lecture Notes in Computer Science, 1424:83–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1424/14240083.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/1424/14240083.
pdf.

Castellanos:1998:MCG

[2219] J. Castellanos, S. Leiva, J. Rodrigo, and A. Rodríguez-Patón. Molecular computation for genetic algorithms. Lecture Notes in Computer Science, 1424:91–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1424/14240091.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1424/14240091.
pdf.

Ceterchi:1998:MCL

[2220] Rodica Ceterchi. Marcus contextual languages and their cut-and-paste properties. Lecture Notes in Computer Science, 1424:99–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240099.
pdf.

Gramatovici:1998:CML

[2221] Radu Gramatovici. Contextual multilanguages: A learning method. Lecture Notes in Computer Science, 1424: 107–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240107.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240107.
pdf.

Knyazev:1998:RWL

[2222] Alexander Knyazev. On recognition of words from languages generated by linear grammars with one nonterminal symbol. Lecture Notes in Computer Science, 1424:111–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240111.
                                        pdf.

Buszkowski:1998:ASD

[2223] Wojciech Buszkowski. Approximation spaces and definability for incomplete information systems. Lecture Notes in Computer Science, 1424:115– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240115.
                                        pdf.

Pagliani:1998:ICH

[2224] Piero Pagliani. Intrinsic co-Heyting boundaries and information incompleteness in rough set analysis. Lecture Notes in Computer Science, 1424: 123–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240123.
                                        pdf.

Maritz:1998:MAO

[2225] P. Maritz. Multifunctions as approximation operations in generalized approximation spaces. Lecture Notes in Computer Science, 1424:131– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240131.
pdf.

Jaervinen:1998:PRT

[2226] J. Jaervinen. Preimage relations and their matrices. Lecture Notes in Computer Science, 1424:139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarvinen:1998:PRT

[2227] J. Järvinen. Preimage relations and their matrices. Lecture Notes in Computer Science, 1424:139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240139.
pdf.

Siemiatkowska:1998:CNN

[2228] Barbara Siemiatkowska and Artur Dubrawski. Cellular neural networks for navigation of a mobile robot. Lecture Notes in Computer Science, 1424: 147–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1424/14240147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240147.
pdf.

Keming:1998:TSF

[2229] Xie Keming, T. Y. Lin, and Zhang Jianwei. The Takagi-Sugeno fuzzy model identification method of parameter varying systems. Lecture Notes in Computer Science, 1424:155– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240155.
pdf.

Eguchi:1998:SAL

[2230] Kazuhiko Eguchi, Satoshi Yamane, Hideo Sugi, and Kenji Oshima. Sensing of arc length and wire extension using neural network in robotic welding. Lecture Notes in Computer Science, 1424: 163–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240163.
pdf.

Yamane:1998:TSC

[2231] Satoshi Yamane, Kazuhiro Okada, Kenji Shinoda, and Kenji Oshima. Traffic signal control using multilayered fuzzy control. Lecture Notes in Computer Science, 1424:171–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240171.
pdf.

Ziarko:1998:ARB

[2232] Wojciech Ziarko. Approximation region-based decision tables. Lecture Notes in Computer Science, 1424:178– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240178.
pdf.

Fernandez-Baizan:1998:MRI

[2233] María C. Fernández-Baizán, Ernestina Menasalvas Ruiz, and Anita Wasilewska. A model of RSDM implementation. Lecture Notes in Computer Science, 1424: 186–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240186.
pdf.

Ras:1998:HQI

[2234] Zbigniew W. Rás. Handling queries in incomplete CKBS through knowledge discovery. Lecture Notes in Computer Science, 1424:194–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240194.
                                        pdf.

Martienne:1998:LLD

[2235] Emmanuelle Martienne and Mohamed Quafafou. Learning logical descriptions for document understanding: A rough sets-based approach. Lecture Notes in Computer Science, 1424:202– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240202.
                                        pdf.

Fernandez-Baizan:1998:IKA

[2236] María C. Fernández-Baizán, Ernestina Menasalvas Ruiz, José M. Peña Sánchez, and Borja Pardo Pastrana. Integrating KDD algorithms and RDBMS code. Lecture Notes in Computer Science, 1424:210–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1424/14240210.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240210.
                                        pdf.

Kryszkiewicz:1998:FDR

[2237] Marzena Kryszkiewicz. Fast discovery of representative association rules. Lecture Notes in Computer Science, 1424: 214–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1424/14240214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240214.
pdf.

Lenarcik:1998:RCS

[2238] Andrzej Lenarcik and Zdzisław Piasta. Rough classifiers sensitive to costs varying from object to object. Lecture Notes in Computer Science, 1424:222– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240222.
pdf.

Zhong:1998:STD

[2239] Ning Zhong, Juzhen Dong, and Setsuo Ohsuga. Soft techniques to data mining. Lecture Notes in Computer Science, 1424:231–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240231.
pdf.

Zytkow:1998:BPU

[2240] Jan M. Zytkow and Arun P. Sanjeev. Business process understanding: Mining many datasets. Lecture Notes in Computer Science, 1424:239– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1424/14240239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240239.
                                        pdf.

Kanasugi:1998:GAS

[2241] Akinori Kanasugi, Takashi Shimayama, Naoshi Nakaya, and Takeshi Iizuka. A genetic algorithm for switchbox routing problem. Lecture Notes in Computer Science, 1424:247–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240247.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240247.
                                        pdf.

Voigt:1998:BRM

[2242] Hans-Michael Voigt and Jan Matti Lange. On the benefits of random memorizing in local evolutionary search. Lecture Notes in Computer Science, 1424:255–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240255.
                                        pdf.

Eguchi:1998:AGA

[2243] Kazuhiko Eguchi, Junya Suzuki, Satoshi Yamane, and Kenji Oshima. An application of genetic algorithms to floorplanning of VLSI. Lecture Notes in Computer Science, 1424:263– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240263.
pdf.

Boryczka:1998:LDR

[2244] Urszula Boryczka. Learning with delayed rewards in ant systems for the job-shop scheduling problem. Lecture Notes in Computer Science, 1424:271– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240271.
pdf.

Cattaneo:1998:FER

[2245] G. Cattaneo. Fuzzy extension of rough sets theory. Lecture Notes in Computer Science, 1424:275–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240275.
pdf.

Greco:1998:FSR

[2246] Salvatore Greco, Benedetto Matarazzo, and Roman Slowinski. Fuzzy similarity relation as a basis for rough approximations. Lecture Notes in Computer Science, 1424:283–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240283.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240283.
                                        pdf.

Stepaniuk:1998:ASE

[2247] Jarosław Stepaniuk. Approximation spaces in extensions of rough set theory. Lecture Notes in Computer Science, 1424:290–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1424/14240290.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240290.
                                        pdf.

Yao:1998:GPA

[2248] Y. Y. Yao. On generalizing Pawlak approximation operators. Lecture Notes in Computer Science, 1424:298– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240298.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240298.
                                        pdf.

Lange:1998:RTR

[2249] Jan Matti Lange, Hans-Michael Voigt, Steffen Burkhardt, and Ralph Göbel. Real-time real-world visual classification — making computational intelligence fly. Lecture Notes in Computer Science, 1424:308–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1424/14240308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240308.
pdf.

Skarbek:1998:FOC

[2250] Wladysław Skarbek. Fractal operator convergence by analysis of influence graph. Lecture Notes in Computer Science, 1424:316–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240316.
pdf.

Ignasiak:1998:PRI

[2251] Krystian Ignasiak and Wladysław Skarbek. Pattern recognition by invariant reference points. Lecture Notes in Computer Science, 1424:322–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240322.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240322.
pdf.

DelRosario:1998:ACS

[2252] J. M. Del Rosario and C. Jedrzejek. An analysis of context selection in embedded wavelet coders. Lecture Notes in Computer Science, 1424:330–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosario:1998:ACS

[2253] Juan Miguel del Rosario and Czesław Jedrzejek. An analysis of context selection in embedded wavelet coders. Lecture Notes in Computer Science, 1424: 330–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240330.
pdf.

Wong:1998:ECC

[2254] S. K. M. Wong and C. J. Butz. Equivalent characterization of a class of conditional probabilistic independencies. Lecture Notes in Computer Science, 1424:338–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1424/14240338.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1424/14240338.
pdf.

Klopotek:1998:NQR

[2255] Mieczysław A. Klopotek and Slawomir T. Wierzchon. A new qualitative rough-set approach to modeling belief functions. Lecture Notes in Computer Science, 1424:346–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240346.
pdf.

Sikora:1998:SOA

[2256] Radosław Sikora and Wladysław Skarbek. On stability of Oja algorithm. Lecture Notes in Computer Science, 1424: 354–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240354.
pdf.

Rak:1998:TVQ

[2257] Remigiusz J. Rak. Transform vector quantization of images in one dimension. Lecture Notes in Computer Science, 1424:361–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240361.
pdf.

Rakowski:1998:DFW

[2258] Waldemar Rakowski and Zbigniew Bartosiewicz. Daubechies filters for 2D wavelet transforms. Lecture Notes in Computer Science, 1424:369–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240369.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240369.
pdf.

Jenssen:1998:SHD

[2259] Tor-Kristian Jenssen, Jan Komorowski, and Aleksander Ohrn. Some heuristics for default knowledge discovery. Lecture Notes in Computer Science, 1424: 373–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240373.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240373.
                                        pdf.

Lin:1998:FPI

[2260] T. Y. Lin. Fuzzy partitions II: belief functions. A probabilistic view. Lecture Notes in Computer Science, 1424: 381–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240381.
                                        pdf.

Lin:1998:FMB

[2261] T. Y. Lin, N. Zhong, J. J. Dong, and S. Ohsuga. Frameworks for mining binary relations in data. Lecture Notes in Computer Science, 1424:387– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240387.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240387.
                                        pdf.

Stefanowski:1998:HCA

[2262] Jerzy Stefanowski. Handling continuous attributes in discovery of strong decision rules. Lecture Notes in Computer Science, 1424:394–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240394.
pdf.

Wroblewski:1998:CRF

[2263] Jakub Wróblewski. Covering with reducts — A fast algorithm for rule generation. Lecture Notes in Computer Science, 1424:402–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240402.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240402.
pdf.

Bartol:1998:SCF

[2264] Wiktor Bartol, Xavier Caicedo, and Francesc Rosselló. Syntactical content of finite approximations of partial algebras. Lecture Notes in Computer Science, 1424:408–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240408.
pdf.

Pacholczyk:1998:NAL

[2265] Daniel Pacholczyk. A new approach to linguistic negation based upon compatibility level and tolerance threshold. Lecture Notes in Computer Science, 1424:416–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240416.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240416.
                                        pdf.

Sakai:1998:SIN

[2266] Hiroshi Sakai. Some issues on nondeterministic knowledge bases with incomplete and selective information. Lecture Notes in Computer Science, 1424: 424–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240424.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240424.
                                        pdf.

Liu:1998:ORO

[2267] Qing Liu. The OI-resolution of operator rough logic. Lecture Notes in Computer Science, 1424:432–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240432.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240432.
                                        pdf.

Krawiec:1998:PME

[2268] Krzysztof Krawiec, Roman Slowinski, and Irmina Szczesniak. Pedagogical method for extraction of symbolic knowledge from neural networks. Lecture Notes in Computer Science, 1424: 436–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240436.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240436.
pdf.

Wojdyllo:1998:WRS

[2269] Piotr Wojdyllo. Wavelets, rough sets and artificial neural networks in EEG analysis. Lecture Notes in Computer Science, 1424:444–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240444.
pdf.

Wojdytto:1998:WRS

[2270] P. Wojdytto. Wavelets, rough sets and artificial neural networks in EEG analysis. Lecture Notes in Computer Science, 1424:444–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Susmaga:1998:PCR

[2271] Robert Susmaga. Parallel computation of reducts. Lecture Notes in Computer Science, 1424:450–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240450.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240450.
pdf.

Mrozek:1998:RRP

[2272] Adam Mrózek and Krzysztof Skabek. Rough rules in Prolog. Lecture Notes in Computer Science, 1424:458–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240458.
pdf.

Czyzewski:1998:TPN

[2273] Andrzej Czyzewski and Bozena Kostek. Tuning the perceptual noise reduction algorithm using rough sets. Lecture Notes in Computer Science, 1424:467– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240467.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240467.
pdf.

Tsumoto:1998:MMD

[2274] Shusaku Tsumoto. Modelling medical diagnostic rules based on rough sets. Lecture Notes in Computer Science, 1424:475–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

0558/bibs/1424/14240475.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1424/14240475.
pdf.

Wakulicz-Deja:1998:DCA

[2275] Alicja Wakulicz-Deja, Mariusz Boryczka, and Piotr Paszek. Discretization of continuous attributes on decision system in mitochondrial encephalomyopathies. Lecture Notes in Computer Science, 1424:483–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240483.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240483.
pdf.

Peters:1998:ATR

[2276] J. F. Peters, K. Ziaei, and S. Ramanna. Approximate time rough control: Concepts and application to satellite attitude control. Lecture Notes in Computer Science, 1424:491– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240491.
pdf.

Moshkov:1998:SRB

[2277] Mikhail Moshkov. Some relationships between decision trees and decision rule systems. Lecture Notes in Computer Science, 1424:499–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240499.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240499.
                                        pdf.

Chikalov:1998:DTM

[2278] Igor Chikalov. On decision trees with minimal average depth. Lecture Notes in Computer Science, 1424:506– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240506.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240506.
                                        pdf.

Moshkova:1998:DRF

[2279] Albina Moshkova. On diagnosis of retaining faults in circuits. Lecture Notes in Computer Science, 1424:513– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240513.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240513.
                                        pdf.

Shevtchenko:1998:DDT

[2280] Vladimir Shevtchenko. On the depth of decision trees for diagnosing of nonelementary faults in circuits. Lecture Notes in Computer Science, 1424:517– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240517.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1424/14240517.
pdf.

Bazan:1998:DDR

[2281] Jan G. Bazan. Discovery of decision rules by matching new objects against data tables. Lecture Notes in Computer Science, 1424:521–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240521.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240521.
pdf.

Zhou:1998:REM

[2282] Yujian Zhou and Jue Wang. Rule + exception modeling based on rough set theory. Lecture Notes in Computer Science, 1424:529–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240529.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240529.
pdf.

Chlebus:1998:FOD

[2283] Bogdan S. Chlebus and Sinh Hoa Nguyen. On finding optimal discretizations for two attributes. Lecture Notes in Computer Science, 1424:537– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240537.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1424/14240537.
                                        pdf.

Nguyen:1998:DPR

[2284] Hung Son Nguyen. Discretization problem for rough sets methods. Lecture Notes in Computer Science, 1424:545– 552, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240545.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240545.
                                        pdf.

Johnson:1998:RMI

[2285] Julia Johnson. Rough mereology for industrial design. Lecture Notes in Computer Science, 1424:553–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240553.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240553.
                                        pdf.

Flasinski:1998:OSS

[2286] Mariusz Flasinski, Robert Schaefer, and Wojciech Toporkiewicz. Optimal stochastic scaling of CAE parallel computations. Lecture Notes in Computer Science, 1424:557–564, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240557.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240557.
                                        pdf.

Wierzbicki:1998:CCO

[2287] Jan Wierzbicki. CBR for complex objects represented in hierarchical information systems. Lecture Notes in Computer Science, 1424:565–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240565.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240565.
pdf.

Rawski:1998:AIM

[2288] M. Rawski, L. Józwiak, and A. Chojnacki. Application of the information measures to input support selection in functional decomposition. Lecture Notes in Computer Science, 1424: 573–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240573.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240573.
pdf.

Burns:1998:MSG

[2289] Tom R. Burns and Anna Gomolinska. Modelling social game systems by rule complexes. Lecture Notes in Computer Science, 1424:581–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240581.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240581.
pdf.

Jozwiak:1998:ASI

[2290] Lech Józwiak. Analysis and synthesis of information systems with information relationships and measures. Lecture Notes in Computer Science, 1424: 585–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240585.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240585.
                                        pdf.

Traczyk:1998:ADM

[2291] Wiesław Traczyk. Approximations in data mining. Lecture Notes in Computer Science, 1424:589–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240589.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240589.
                                        pdf.

VandenPoel:1998:PPD

[2292] Dirk Van den Poel and Zdzisław Piasta. Purchase prediction in database marketing with the ProbRough system. Lecture Notes in Computer Science, 1424:593–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240593.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240593.
                                        pdf.

Czajewski:1998:RSO

[2293] Witold Czajewski. Rough sets in optical character recognition. Lecture Notes in Computer Science, 1424:601– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240601.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240601.
pdf.

Predki:1998:RSI

[2294] Bartlomiej Predki, Roman Slowinski, Jerzy Stefanowski, Robert Susmaga, and Szymon Wilk. ROSE — software implementation of the rough set theory. Lecture Notes in Computer Science, 1424:605–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240605.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240605.
pdf.

Swiniarski:1998:RSB

[2295] Roman W. Swiniarski. Rough sets and Bayesian methods applied to cancer detection. Lecture Notes in Computer Science, 1424:609–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1424/14240609.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1424/14240609.
pdf.

Swiniarski:1998:RSN

[2296] Roman W. Swiniarski. Rough sets and neural networks application to handwritten character recognition by complex Zernike moments. Lecture Notes in Computer Science, 1424:617– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1424/14240617.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1424/14240617.
                                        pdf.

Anonymous:1998:SAAa

[2297] Anonymous. Session AR1 approximate reasoning. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SAAb

[2298] Anonymous. Session AP1 applications of rough sets. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SCC

[2299] Anonymous. Session C1 complexity. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SDD

[2300] Anonymous. Session DM1 data mining. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SEE

[2301] Anonymous. Session EV1 evolutionary computations. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SGGa

[2302] Anonymous. Session GS1 grammar systems and molecular computations. Lecture Notes in Computer Science, 1424: ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SGGb

[2303] Anonymous. Session GA1 general approximations. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SGGc

[2304] Anonymous. Session GH1 general hybrid methods. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SHH

[2305] Anonymous. Session HRS1 hybrid methods in rough sets. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SIIa

[2306] Anonymous. Session IC1 intelligent control. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SIIb

[2307] Anonymous. Session IP1 image processing. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SIIc

[2308] Anonymous. Session IP2 image processing. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SLLa

[2309] Anonymous. Session LRS1 logic in rough sets. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SLLb

[2310] Anonymous. Session LFSC1 logical foundations of soft computing. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SRRa

[2311] Anonymous. Session RSM1 rough set methods. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SRRb

[2312] Anonymous. Session RSKDD1 rough sets in KDD. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SRRc

[2313] Anonymous. Session RSM2 rough set methods. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SRRd

[2314] Anonymous. Session RSMA1 rough set methods and applications. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SRRe

[2315] Anonymous. Session RSMA2 rough set methods and applications. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:SSS

[2316] Anonymous. Session ST1 statistical inference. Lecture Notes in Computer Science, 1424:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kompatsiaris:1998:RVI

[2317] I. Kompatsiaris and M. G. Strintzis. 3D representation of videoconference image sequences using VRML 2.0. Lecture Notes in Computer Science, 1425:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leevers:1998:ASR

[2318] D. Leevers, P. Gil, F. M. Lopes, and J. Pereira. An autonomous sensor for 3D reconstruction. Lecture Notes in Computer Science, 1425:13–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohm:1998:IMR

[2319] J.-R. Ohm and K. Mueller. Incomplete 3D for multiview representation and synthesis of video objects. Lecture Notes in Computer Science, 1425: 26–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rousseau:1998:EAS

[2320] F. Rousseau and A. Duda. An execution architecture for synchronized multimedia presentations. Lecture Notes in Computer Science, 1425:42–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

TenKate:1998:PMW

[2321] W. Ten Kate, P. Deunhouwer, D. Bulterman, and L. Hardman. Presenting multimedia on the Web and in TV broadcast. Lecture Notes in Computer Science, 1425:56–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stolp:1998:MAD

[2322] R. Stolp, A. Scheller, and A. Kraft. MHEG-5 application development. Lecture Notes in Computer Science, 1425:70–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kyriakaki:1998:SMS

[2323] G. Kyriakaki, Y. Maragoudakis, Y. Mavraganis, and N. Pappas. The SICMA multimedia server and virtual museum application. Lecture Notes in Computer Science, 1425:83–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cortes:1998:PCM

[2324] F. Cortes and R. Lueling. A parallel continuous media server complying to the RTSP protocol. Lecture Notes in Computer Science, 1425:97–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bessler:1998:MSM

[2325] S. Bessler and M. Hager. Media streaming in a multimedia annotation application. Lecture Notes in Computer Science, 1425:111–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saydam:1998:MSM

[2326] T. Saydam, S. Yucel, and T. Kusano.

A multimedia service and management architecture using APIs over CORBA platform. Lecture Notes in Computer Science, 1425:121–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edwards:1998:OIS

[2327] C. Edwards, D. Hutchison, and D. Waddington. Open interface support for heterogeneous network services. Lecture Notes in Computer Science, 1425:135–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fallows:1998:CAM

[2328] J. R. Fallows and I. W. Marshall.

A CORBA-Assisted multimedia proxy server. Lecture Notes in Computer Science, 1425:149–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marie-Julie:1998:IIU

[2329] J. M. Marie-Julie and H. Essafi. Image indexing by using a rotation and scale invariant partition. Lecture Notes in Computer Science, 1425:163–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yusoff:1998:SAS

[2330] Y. Yusoff, W. Christmas, and J. Kittler. A study on automatic shot change detection. Lecture Notes in Computer Science, 1425:177–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darmstaedter:1998:BBW

[2331] V. Darmstaedter, J.-F. Delaigle, D. Nicholson, and B. Macq. A block based watermarking technique for MPEG2 signals: Optimization and validation on real digital TV distribution links. Lecture Notes in Computer Science, 1425:190–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kunkelmann:1998:SSM

[2332] T. Kunkelmann, H. Vogler, M.-L. Moschgath, and L. Wolf. Scalable security mechanisms in transport systems for enhanced multimedia services. Lecture Notes in Computer Science, 1425:207–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Losquadro:1998:SBS

[2333] G. Losquadro, M. Barbieri, M. Luglio, and F. Vatalaro. The SECOMS broadband satellite network in the multimedia services and multi-environment coverage scenario. Lecture Notes in Computer Science, 1425:221–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molina:1998:DSI

[2334] A. Molina and J. Seseña. DIGISAT-S3M: The interactivity for SMATV users. Lecture Notes in Computer Science, 1425:233–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leisenberg:1998:MMS

[2335] M. Leisenberg and T. Lindgren. MO-MENTS — multimedia services in a narrow-bandwidth cellular environment. Lecture Notes in Computer Science, 1425:246–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniazzi:1998:MBN

[2336] S. Antoniazzi, H. Marmolin, G. Schapeler, and B. Weickert. MUSIST browser and navigation concept. Lecture Notes in Computer Science, 1425:260–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sellek:1998:SAS

[2337] P. A. Sellek and D. O. Beaumont.

SimDS — an approach to service modelling. Lecture Notes in Computer Science, 1425:274–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shin:1998:CMS

[2338] Y. M. Shin and S. Cho. Controlling multimedia streams across Internet and ATM network. Lecture Notes in Computer Science, 1425:288–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zahariadis:1998:NAP

[2339] T. Zahariadis, C. Georgopoulos, V. Nellas, and T. Arvanitis. Native ATM protocol stack for Internet applications in residential broadband networks. Lecture Notes in Computer Science, 1425:300–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Romano:1998:IQA

[2340] S. P. Romano, C. Deleuze, J. Rezende, and S. Fdida. Integrated QoS architecture for IP switching. Lecture Notes in Computer Science, 1425:312–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlary:1998:DIA

[2341] Y. Orlary, O. Carbonel, S. Gibbs, and D. Konstantas. Design and implementation of an ATM based distributed musical rehearsal studio. Lecture Notes in Computer Science, 1425: 326–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foster:1998:AMC

[2342] P. Foster, S. Banthorpe, and R. Gepp. Automating the multimedia content production lifecycle. Lecture Notes in Computer Science, 1425:340–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caire:1998:ATD

[2343] G. Caire. ATMAN: Trading of digital audio visual contents. Lecture Notes in Computer Science, 1425:352–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolla:1998:PQS

[2344] R. Bolla, A. Iscra, M. Marchese, and S. Zappatore. A perceived quality of service optimization for video communication in ‘best-effort’ networks. Lecture Notes in Computer Science, 1425: 366–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lampard:1998:PTC

[2345] G. J. Lampard. Performance of TCP over cable modems and ADSL. Lecture Notes in Computer Science, 1425: 380–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reisslein:1998:DPP

[2346] M. Reisslein, K. W. Ross, and V. Verillotte. A decentralized prefetching protocol for VBR video on demand. Lecture Notes in Computer Science, 1425: 388–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spaan:1998:ERP

[2347] F. H. P. Spaan, R. L. Lagendijk, and J. Biemond. Error-robustness of polar contour-coding. Lecture Notes in Computer Science, 1425:402–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bartkowiak:1998:ERC

[2348] M. Bartkowiak and M. Domanski. Efficient representation of chrominance for very low bitrate coding. Lecture Notes in Computer Science, 1425:415– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mattavelli:1998:RTC

[2349] M. Mattavelli and S. Brunetton. Realtime constraints and prediction of video decoding time for multimedia systems. Lecture Notes in Computer Science, 1425:425–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunetti:1998:VQB

[2350] M. Gunetti, P. Sunna, and M. Visca. Video quality and bit rates in thematic television channels. Lecture Notes in Computer Science, 1425:439–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robertson:1998:SDW

[2351] P. Robertson, H.-P. Huth, K. Fazel, and O. Klank. A system design for a wireless home multi-media LAN. Lecture Notes in Computer Science, 1425: 453–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halbwachs:1998:SPR

[2352] N. Halbwachs. Synchronous programming of reactive systems. Lecture Notes in Computer Science, 1427:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peled:1998:TYP

[2353] D. Peled. Ten years of partial order reduction. Lecture Notes in Computer Science, 1427:17–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moore:1998:APW

[2354] J. S. Moore. An ACL2 proof of write invalidate cache coherence. Lecture Notes in Computer Science, 1427:29– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hardin:1998:TTP

[2355] D. Hardin, M. Wilding, and D. Greve.

Transforming the theorem prover into a digital design tool: From concept car to off-road vehicle. Lecture Notes in Computer Science, 1427:39–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camilleri:1998:RTP

[2356] A. J. Camilleri. A role for theorem proving in multi-processor design. Lecture Notes in Computer Science, 1427:45–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffman:1998:FME

[2357] J. Hoffman and C. Payne. A formal method experience at secure computing corporation. Lecture Notes in Computer Science, 1427:49–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuellar:1998:FMI

[2358] J. R. Cuellar. Formal methods in an industrial environment. Lecture Notes in Computer Science, 1427:57–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzmann:1998:CMC

[2359] G. J. Holzmann. On checking model checkers. Lecture Notes in Computer Science, 1427:61–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1998:FSA

[2360] J. C. Mitchell. Finite-state analysis of security protocols. Lecture Notes in Computer Science, 1427:71–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolignano:1998:IPB

[2361] D. Bolignano. Integrating proof-based and model-checking techniques for the formal verification of cryptographic protocols. Lecture Notes in Computer Science, 1427:77–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolper:1998:VSI

[2362] P. Wolper and B. Boigelot. Verifying systems with infinite but regular state spaces. Lecture Notes in Computer Science, 1427:88–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skakkebaek:1998:FVO

[2363] J. U. Skakkebaek, R. B. Jones, and D. L. Dill. Formal verification of outof-order execution using incremental flushing. Lecture Notes in Computer Science, 1427:98–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McMillan:1998:VIT

[2364] K. L. McMillan. Verification of an implementation of Tomasulo’s algorithm by compositional model checking. Lecture Notes in Computer Science, 1427:110–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hosabettu:1998:DPC

[2365] R. Hosabettu, M. Srivas, and G. Gopalakrishnan. Decomposing the proof of correctness of pipelined microprocessors. Lecture Notes in Computer Science, 1427:122–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sawada:1998:PVP

[2366] J. Sawada and W. A. Hunt. Processor verification with precise exceptions and speculative execution. Lecture Notes in Computer Science, 1427:135–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1998:SRM

[2367] E. M. Clarke, E. A. Emerson, S. Jha, and A. P. Sistla. Symmetry reductions in model checking. Lecture Notes in Computer Science, 1427:147–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manku:1998:SSM

[2368] G. S. Manku, R. Hojati, and R. Brayton. Structural symmetry and model checking. Lecture Notes in Computer Science, 1427:159–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stern:1998:UMD

[2369] U. Stern and D. L. Dill. Using magnetic disk instead of main memory in the murcursive phi verifier. Lecture Notes in Computer Science, 1427:172– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beer:1998:FMC

[2370] I. Beer, S. Ben-David, and A. Landver. On-the-fly model checking of RCTL formulas. Lecture Notes in Computer Science, 1427:184–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1998:PHP

[2371] T. A. Henzinger, O. Kupferman, and S. Qadeer. From pre-historic to postmodern symbolic model checking. Lecture Notes in Computer Science, 1427: 195–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallner:1998:MCL

[2372] F. Wallner. Model checking LTL using net unfoldings. Lecture Notes in Computer Science, 1427:207–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1998:MCF

[2373] Y. Xu, E. Cerny, X. Song, and F. Cotella. Model checking for a first-order temporal logic using multiway decision graphs. Lecture Notes in Computer Science, 1427:219–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thathachar:1998:LOR

[2374] J. S. Thathachar. On the limitations of ordered representations of functions. Lecture Notes in Computer Science, 1427:232–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goel:1998:BBP

[2375] A. Goel, K. Sajid, H. Zhou, and A. Aziz. BDD based procedures for a theory of equality with uninterpreted functions. Lecture Notes in Computer Science, 1427:244–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Isles:1998:CRC

[2376] A. J. Isles, R. Hojati, and R. K.

Brayton. Computing reachable control states of systems modeled with uninterpreted functions and infinite memory. Lecture Notes in Computer Science, 1427:256–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comon:1998:MCA

[2377] H. Comon and Y. Jurski. Multiple counters automata, safety analysis, and Presburger arithmetic. Lecture Notes in Computer Science, 1427: 268–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shiple:1998:CPE

[2378] T. R. Shiple, J. H. Kukula, and R. K. Ranjan. A comparison of Presburger engines for EFSM reachability. Lecture Notes in Computer Science, 1427:280–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colon:1998:GFS

[2379] M. A. Colon and T. E. Uribe. Generating finite-state abstractions of reactive systems using decision procedures. Lecture Notes in Computer Science, 1427:293–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdulla:1998:FAS

[2380] P. A. Abdulla, A. Bouajjani, and B. Jonsson. On-the-fly analysis of systems with unbounded, lossy FIFO channels. Lecture Notes in Computer Science, 1427:305–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bensalem:1998:CAI

[2381] S. Bensalem, Y. Lakhnech, and S. Owre. Computing abstractions of infinite state systems compositionally and automatically. Lecture Notes in Computer Science, 1427:319–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griffioen:1998:NS

[2382] D. Griffioen and F. Vaandrager. Norreed simulations. Lecture Notes in Computer Science, 1427:332–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Couturier:1998:EPA

[2383] R. Couturier and D. Mery. An experiment in parallelizing an application using formal methods. Lecture Notes in Computer Science, 1427:345–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoller:1998:ESD

[2384] S. D. Stoller and Y. A. Liu. Efficient symbolic detection of global properties in distributed systems. Lecture Notes in Computer Science, 1427:357– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilding:1998:MCP

[2385] M. Wilding. A machine-checked proof of the optimality of a real-time scheduling policy. Lecture Notes in Computer Science, 1427:369–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdulla:1998:GAP

[2386] P. A. Abdulla, B. Jonsson, M. Kindahl, and D. Peled. A general approach to partial order reductions in symbolic verification. Lecture Notes in Computer Science, 1427:379–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balarin:1998:CCA

[2387] F. Balarin. Correctness of the concurrent approach to symbolic verification of interleaved models. Lecture Notes in Computer Science, 1427:391–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belluomini:1998:VTS

[2388] W. Belluomini and C. J. Myers. Verification of timed systems using POSETs. Lecture Notes in Computer Science, 1427:403–??, 1998. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Bella:1998:MBK

[2389] G. Bella and L. C. Paulson. Mechanising BAN Kerberos by the inductive method. Lecture Notes in Computer Science, 1427:416–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felty:1998:PVN

[2390] A. P. Felty, D. J. Howe, and F. A. Stomp. Protocol verification in Nuprl. Lecture Notes in Computer Science, 1427:428–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1998:YAW

[2391] T. A. Henzinger, S. Qadeer, and S. K. Rajamani. You assume, we guarantee: Methodology and case studies. Lecture Notes in Computer Science, 1427:440–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerson:1998:VPB

[2392] E. A. Emerson and K. S. Namjoshi. Verification of a parameterized bus arbitration protocol. Lecture Notes in Computer Science, 1427:452–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nalumasu:1998:TMC

[2393] R. Nalumasu, R. Ghughal, A. Mokkedem, and G. Gopalakrishnan. The ‘test model-checking’ approach to the verification of formal memory models of multiprocessors. Lecture Notes in Computer Science, 1427:464–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufmann:1998:DCS

[2394] M. Kaufmann, A. Martin, and C. Pixley. Design constraints in symbolic model checking. Lecture Notes in Computer Science, 1427:477–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1998:VFP

[2395] Y.-A. Chen and R. E. Bryant. Verification of floating-point adders. Lecture Notes in Computer Science, 1427: 488–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouali:1998:XEV

[2396] A. Bouali. Xeve: An Esterel verification environment. Lecture Notes in Computer Science, 1427:500–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bensalem:1998:ITV

[2397] S. Bensalem, Y. Lakhnech, and S. Owre. InVeSt: A tool for the verification of invariants. Lecture Notes in Computer Science, 1427:505–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1998:VMP

[2398] G. Ferrari, S. Gnesi, U. Montanari, and M. Pistore. Verifying mobile processes in the HAL environment. Lecture Notes in Computer Science, 1427: 511–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elgaard:1998:MXN

[2399] J. Elgaard, N. Klarlund, and A. Moeller. Mona 1.x: New techniques for WS1S and WS2S. Lecture Notes in Computer Science, 1427:516–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1998:MMM

[2400] R. Alur, T. A. Henzinger, F. Y. C. Mang, and S. Qadeer. MOCHA: Modularity in model checking. Lecture Notes in Computer Science, 1427:521– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heitmeyer:1998:STS

[2401] C. Heitmeyer, J. Kirby, B. Labaw, and R. Bharadwaj. SCR0*: A toolset for specifying and analyzing software requirements. Lecture Notes in Computer Science, 1427:526–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peled:1998:TMS

[2402] D. A. Peled. A toolset for message sequence charts. Lecture Notes in Computer Science, 1427:532–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brockmeyer:1998:RTV

[2403] U. Brockmeyer and G. Wittich. Realtime verification of statemate designs. Lecture Notes in Computer Science, 1427:537–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daws:1998:OTS

[2404] C. Daws. Optikron: A tool suite for enhancing model-checking of real-time systems. Lecture Notes in Computer Science, 1427:542–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozga:1998:KMC

[2405] M. Bozga, C. Daws, O. Maler, and A. Olivero. Kronos: A model-checking tool for real-time systems. Lecture Notes in Computer Science, 1427:546– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Obbink:1998:I

[2406] Henk Obbink, Paul C. Clements, and Frank van der Linden. Introduction. Lecture Notes in Computer Science, 1429:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1429/14290001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1429/14290001.
pdf.

Scherlis:1998:SEA

[2407] William Scherlis and Juha Kuusela.

Session 1: Example architectures. Lecture Notes in Computer Science, 1429: 4–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290004.
pdf.

Cysewski:1998:RFT

[2408] G. Cysewski, T. Gromadzki, H. Lyskawa, M. Piechowka, S. Szejko, W. E. Kozlowski, and O. Vahamaki. Reusable framework for telecontrol protocols. Lecture Notes in Computer Science, 1429:6–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1429/14290006.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290006.
                                        pdf.

Niemela:1998:SBP

[2409] Eila Niemelä, Harri Perunka, and Tomi Korpipää. A software bus as a platform for a family of distributed embedded system products. Lecture Notes in Computer Science, 1429:14– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290014.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290014.
                                        pdf.

Niemelae:1998:SBP

[2410] E. Niemelae, H. Perunka, and T. Korpipaeae. A software bus as a platform for a family of distributed embedded system products. Lecture Notes in Computer Science, 1429:14–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolb:1998:TTD

[2411] Peter Kolb and Beat Huber. A threetier design approach for a family of large AC drive control systems. Lecture Notes in Computer Science, 1429: 24–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290024.
pdf.

Ganz:1998:MTC

[2412] Christopher Ganz and Michael Layes.

Modular turbine control software: A control software architecture for the ABB gas turbine family. Lecture Notes in Computer Science, 1429:32– 38, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290032.
pdf.

Roesel:1998:EEA

[2413] A. Roesel. Experiences with the evolution of an application family architecture. Lecture Notes in Computer Science, 1429:39–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosel:1998:EEA

[2414] Andreas Rösel. Experiences with the evolution of an application family architecture. Lecture Notes in Computer Science, 1429:39–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1429/14290039.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290039.
                                        pdf.

Perry:1998:SAD

[2415] Dewayne E. Perry and Jeff Kramer. Session 2: Architectural description. Lecture Notes in Computer Science, 1429:49–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1429/14290049.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290049.
                                        pdf.

Perry:1998:GAD

[2416] Dewayne E. Perry. Generic architecture descriptions for product lines. Lecture Notes in Computer Science, 1429:51– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290051.
                                        pdf.

Pryce:1998:MIC

[2417] Nat Pryce and Steve Crane. A model of interaction in concurrent and distributed systems. Lecture Notes in Computer Science, 1429:57– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290057.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1429/14290057.
pdf.

Hamer:1998:IHD

[2418] Peter van den Hamer, Frank van der Linden, Alison Saunders, and Henk te Sligte. An integral hierarchy and diversity model for describing product family architectures. Lecture Notes in Computer Science, 1429:66– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290066.
pdf.

VanDeHamer:1998:IHD

[2419] P. Van De Hamer, F. J. Van Der Linden, A. Saunders, and H. Te Sligte. An integral hierarchy and diversity model for describing product family architecture. Lecture Notes in Computer Science, 1429:66–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vandenHamer:1998:IHD

[2420] Peter van den Hamer, Frank van der Linden, Alison Saunders, and Henk te Sligte. An integral hierarchy and diversity model for describing product family architectures. Lecture Notes in Computer Science, 1429:66– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290066.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1429/14290066.
pdf.

Ommering:1998:KCM

[2421] Rob van Ommering. Koala, a component model for consumer electronics product software. Lecture Notes in Computer Science, 1429:76–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290076.
pdf.

vanOmmering:1998:KCM

[2422] Rob van Ommering. Koala, a component model for consumer electronics product software. Lecture Notes in Computer Science, 1429:76–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290076.
pdf.

DeBaud:1998:SAR

[2423] Jean-Marc DeBaud and Mehdi Jazayeri. Session 3: Architecture recovery. Lecture Notes in Computer Science, 1429:87–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1429/14290087.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1429/14290087.
pdf.

Eixelsberger:1998:RAS

[2424] Wolfgang Eixelsberger, Manfred Kalan, Michaela Ogris, Håkon Beckman, Berndt Bellay, and Harald Gall. Recovery of architectural structure: A case study. Lecture Notes in Computer Science, 1429:89–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290089.
pdf.

Knor:1998:RCC

[2425] Roland Knor, Georg Trausmuth, and Johannes Weidl. Reengineering C/C++ source code by transforming state machines. Lecture Notes in Computer Science, 1429:97– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290097.
pdf.

Mendonca:1998:EDS

[2426] Nabor C. Mendonça and Jeff Kramer.

An experiment in distributed software architecture recovery. Lecture Notes in Computer Science, 1429:106– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290106.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1429/14290106.
                                        pdf.

Bellay:1998:RER

[2427] Berndt Bellay and Harald Gall. Reverse engineering to recover and describe a system’s architecture. Lecture Notes in Computer Science, 1429:115– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290115.
                                        pdf.

Weiderman:1998:CLS

[2428] Nelson Weiderman, John Bergey, Dennis Smith, and Scott Tilley. Can legacy systems beget product lines? Lecture Notes in Computer Science, 1429:123– 131, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290123.
                                        pdf.

DeBaud:1998:RBP

[2429] Jean-Marc DeBaud and Jean-François Girard. The relation between the product line development entry points and reengineering. Lecture Notes in Computer Science, 1429:132–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290132.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1429/14290132.
pdf.

Clements:1998:SAS

[2430] Paul C. Clements and Juan Antonio de la Puente. Session 4: Analysis of software architectures. Lecture Notes in Computer Science, 1429:140– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290140.
pdf.

Bechtold:1998:DSA

[2431] Richard T. Bechtold. Diagnostic software architectures. Lecture Notes in Computer Science, 1429:143–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290143.
pdf.

Duenas:1998:SAE

[2432] Juan C. Dueñas, William L. de Oliveira, and Juan A. de la Puente. A software architecture evaluation model. Lecture Notes in Computer Science, 1429:148– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290148.
pdf.

Balzer:1998:AIP

[2433] Robert Balzer. An architectural infrastructure for product families. Lecture Notes in Computer Science, 1429:158– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290158.
                                        pdf.

Alonso:1998:ATP

[2434] Alejandro Alonso, Marisol García-Valls, and Juan A. de la Puente. Assessment of timing properties of family products. Lecture Notes in Computer Science, 1429:161–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290161.
                                        pdf.

Weiss:1998:SDP

[2435] David M. Weiss and Frank J. van der Linden. Session 5: Development process. Lecture Notes in Computer Science, 1429:170–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290170.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290170.
                                        pdf.

Dolan:1998:SSS

[2436] Tom Dolan, Ruud Weterings, and J. C.

Wortmann. Stakeholders in softwaresystem family architectures. Lecture Notes in Computer Science, 1429:172– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290172.
pdf.

Yu:1998:HVR

[2437] Chye Cheong Yu, Ananda L. Akkihebbal, and Stan Jarzabek. Handling variant requirements in software architectures for product families. Lecture Notes in Computer Science, 1429:188– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290188.
pdf.

Baum:1998:ACS

[2438] Lothar Baum, Lars Geyer, Georg Molter, Steffen Rothkugel, and Peter Sturm. Architecture-centric software development based on extended design spaces. Lecture Notes in Computer Science, 1429:197–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290197.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1429/14290197.
                                        pdf.

Meekel:1998:ADV

[2439] Jacques Meekel, Thomas B. Horton, and Charlie Mellone. Architecting for domain variability. Lecture Notes in Computer Science, 1429:205– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290205.
                                        pdf.

Weiss:1998:CAS

[2440] David M. Weiss. Commonality analysis: A systematic process for defining families. Lecture Notes in Computer Science, 1429:214–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290214.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1429/14290214.
                                        pdf.

Karhinen:1998:SDD

[2441] Anssi Karhinen and Juha Kuusela. Structuring design decisions for evolution. Lecture Notes in Computer Science, 1429:223–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1429/14290223.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1429/14290223.
pdf.

Scherlis:1998:SVS

[2442] William L. Scherlis. Structural views, structural evolution, and product families. Lecture Notes in Computer Science, 1429:235–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290235.
pdf.

Brouwer:1998:PFR

[2443] Jeroen Brouwer, Ad Jurriens, Henk van Kessel, and Alef Schippers. Product family and reuse in separate market driven profit centers. Lecture Notes in Computer Science, 1429:241– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1429/14290241.
pdf.

Bandinelli:1998:ESR

[2444] Sergio Bandinelli. ERW’97 session report: Reuse adoption experiences across a large corporation. Lecture Notes in Computer Science, 1429:249– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1429/14290249.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1429/14290249.
                                        pdf.

Anonymous:1998:AIj

[2445] Anonymous. Author index. Lecture Notes in Computer Science, 1429:257– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1429/1429auth.pdf.

Karali:1998:QBM

[2446] D. Karali, F. Karayannis, K. Berdekas, and J. Reilly. QoS based multidomain routing in public broadband networks. Lecture Notes in Computer Science, 1430:3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anagnostou:1998:ODS

[2447] M. E. Anagnostou. Optimal distribution of service components. Lecture Notes in Computer Science, 1430:17– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeZen:1998:AGS

[2448] G. De Zen, M. A. Marsiglia, G. Ricagni, and L. Vezzoli. Accountable and guaranteed services in Internet. Lecture Notes in Computer Science, 1430:31– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lodge:1998:FIN

[2449] F. Lodge, D. Botvich, and T. Curran. A fair intelligent network congestion control strategy based on revenue optimisation. Lecture Notes in Computer Science, 1430:43–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Acuna:1998:ASU

[2450] C. M. Acuna, J. Sassi, R. Salminen, and A. Casaca. ATM services usage metering for accounting and charging. Lecture Notes in Computer Science, 1430:61–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nesbitt:1998:EPP

[2451] F. Nesbitt, T. Counihan, and J. Hickie.

The EURESCOM P610 project: Providing a framework, architecture and methodology for multimedia service management. Lecture Notes in Computer Science, 1430:73–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiropanis:1998:ORM

[2452] T. Tiropanis. Offering role mobility in a TINA environment. Lecture Notes in Computer Science, 1430:89–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lynch:1998:WET

[2453] N. Lynch and K. Hyland. Web enabled TMN manager for an international trouble ticketing service. Lecture Notes in Computer Science, 1430: 101–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerr:1998:EUI

[2454] D. Kerr, D. O’Sullivan, R. Evans, and R. Richardson. Experiences using intelligent agent technologies as a unifying approach to network management, service management and service delivery. Lecture Notes in Computer Science, 1430:115–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corley:1998:AIMa

[2455] S. Corley, M. Tesselaar, J. Cooley, and J. Meinkoehn. The application of intelligent and mobile agents to network and service management. Lecture Notes in Computer Science, 1430: 127–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjanger:1998:ANM

[2456] B. W. Bjanger and A. Solhaug. ATM network management with distributed transactional agents. Lecture Notes in Computer Science, 1430:139–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stephanidis:1998:AAU

[2457] C. Stephanidis, A. Paramythis, M. Sfyrakis, and A. Stergiou. Adaptable and adaptive user interfaces for disabled users in the AVANTI project. Lecture Notes in Computer Science, 1430:153–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jepsen:1998:AAJ

[2458] T. C. Jepsen and Z. Zhang. An ATM API for Java. Lecture Notes in Computer Science, 1430:167–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewis:1998:ITI

[2459] D. Lewis and T. Tiropanis. Integrating TINA into an Internet-based services market. Lecture Notes in Computer Science, 1430:183–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1998:PII

[2460] J. Hall, M. Best, R. Ferry, and S. Fratini. Protocol independent information modelling for a peer-to-peer configuration interface. Lecture Notes in Computer Science, 1430:193–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hellemans:1998:DDH

[2461] P. Hellemans, M. Mampaey, H. Vanderstraeten, and H. Zandbelt. Development and deployment of a heterogeneous set of services challenging a TINA-Based telecommunication architecture. Lecture Notes in Computer Science, 1430:205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1998:FT

[2462] J. C. Garcia and P. F. Hansen. Federation in TINA. Lecture Notes in Computer Science, 1430:219–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wind:1998:ETA

[2463] B. Wind, M. Samarotto, P. Nicosia, and M. Lambrou. Enhancing the TINA architectural framework for personal mobility support. Lecture Notes in Computer Science, 1430:233–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gray:1998:RPP

[2464] W. T. Gray and J. Nelson. The role of paging in providing location transparency within cellular systems. Lecture Notes in Computer Science, 1430: 249–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raatikainen:1998:IBO

[2465] K. Raatikainen, A. Dede, and O. Koskimies. Internet browsing on OSAM platform. Lecture Notes in Computer Science, 1430:261–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breugst:1998:USM

[2466] M. Breugst and T. Magedanz. On the usage of standard mobile agent platforms in telecommunication environments. Lecture Notes in Computer Science, 1430:275–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stefani:1998:RDK

[2467] J.-B. Stefani, B. Dumant, F. Dang Tran, and F. Horn. The RETINA DPE kernel: A flexible, real-time ORB framework. Lecture Notes in Computer Science, 1430:287–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranc:1998:UTN

[2468] D. Ranc, I. P. Van der Bijl, and S. Sedillot. Use of transactions in network management applications. Lecture Notes in Computer Science, 1430: 297–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strens:1998:BMM

[2469] R. Strens, M. Martin, J. Dobson, and S. Plagemann. Business and market models of brokerage in network-based commerce. Lecture Notes in Computer Science, 1430:315–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Healstroem:1998:DMD

[2470] A. Healstroem, J. Sallros, S. Martin, and R. Smith. Dynamic market driven provisioning of services and resources using software agents and electronic chambers of commerce. Lecture Notes in Computer Science, 1430:327– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asensio:1998:AEB

[2471] J. I. Asensio, V. A. Villagra, J. I.

Moreno, and J. Berrocal. An approach to electronic brokerage in TINA environments. Lecture Notes in Computer Science, 1430:339–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marti:1998:DSL

[2472] R. Marti and J. Delgado. Definition of service levels for electronic brokerage applications. Lecture Notes in Computer Science, 1430:351–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carls:1998:ISD

[2473] G. Carls and B. Frohnhoff. Introducing SDL ’92 in the development of TMN applications. Lecture Notes in Computer Science, 1430:365–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucidi:1998:DTL

[2474] F. Lucidi, H. Idzenga, and S. Batistatos. Development of TINA-like systems: The DOLMEN methodology. Lecture Notes in Computer Science, 1430:379–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schieferdecker:1998:CTT

[2475] L. Schieferdecker, M. Li, and A. Hoffmann. Conformance testing of TINA service components — the TTCN/ CORBA gateway. Lecture Notes in Computer Science, 1430:393–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Efremidis:1998:TOS

[2476] S. Efremidis, D. Prevedourou, L. Demounem, and K. Milsted. TINAoriented service engineering support to service composition and federation. Lecture Notes in Computer Science, 1430:409–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reilly:1998:SCM

[2477] J. Reilly and M. Abate. Scheduled connections: Managing temporal constraints on broadband network resources. Lecture Notes in Computer Science, 1430:425–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindh:1998:PMS

[2478] T. Lindh. Performance management in switched ATM networks. Lecture Notes in Computer Science, 1430:439– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1998:SBC

[2479] T. Tanaka, S. Chida, Y. Tsuchiya, and T. Ikegawa. An SNMP-based CNM agent for ATM networks: System architecture and implementation. Lecture Notes in Computer Science, 1430: 451–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berg:1998:CIN

[2480] H. A. Berg and S. Brennan. CORBA and intelligent network (IN) interworking. Lecture Notes in Computer Science, 1430:463–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rasmussen:1998:CCG

[2481] S. Rasmussen and C. Baeumer. A CORBA to CMIP gateway: A marriage of management technologies. Lecture Notes in Computer Science, 1430: 477–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reynders:1998:TTI

[2482] J.-M. Reynders, G. Bilsen, F. Vandermeulen, and T. Mota. TINA-TMN interworking: Case study of an integration between the VITAL TINA platform and an existing TMN platform. Lecture Notes in Computer Science, 1430:493–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1998:SWY

[2483] A. Weber. See what you sign: Secure implementations of digital signatures. Lecture Notes in Computer Science, 1430:509–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oelnes:1998:PKI

[2484] J. Oelnes, M. Verdier, N. Ganivet, and D. Maillot. Public key infrastructure and certification policy for interdomain management. Lecture Notes in Computer Science, 1430:521–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1998:SBM

[2485] K. M. Martin, B. Preneel, C. J. Mitchell, and H. J. Hitz. Secure billing for mobile information services in UMTS. Lecture Notes in Computer Science, 1430:535–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frankel:1998:DPK

[2486] Yair Frankel and Moti Yung. Distributed public key cryptosystems. Lecture Notes in Computer Science, 1431: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1431/14310001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1431/14310001.
                                        pdf.

Misarsky:1998:HDR

[2487] Jean-François Misarsky. How (not) to design RSA signature schemes. Lecture Notes in Computer Science, 1431:

14–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1431/14310014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1431/14310014.
pdf.

Araki:1998:OEC

[2488] Kiyomichi Araki, Takakazu Satoh, and Shinji Miura. Overview of elliptic curve cryptography. Lecture Notes in Computer Science, 1431:29– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1431/14310029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1431/14310029.
pdf.

Stern:1998:LCO

[2489] Jacques Stern. Lattices and cryptography: An overview. Lecture Notes in Computer Science, 1431:50– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1431/14310050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1431/14310050.
pdf.

Bao:1998:SSS

[2490] Feng Bao and Robert H. Deng. A signcryption scheme with signature directly verifiable by public key. Lecture Notes in Computer Science, 1431:55– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1431/14310055.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1431/14310055.
                                        pdf.

Mao:1998:GCS

[2491] Wenbo Mao. Guaranteed correct sharing of integer factorization with off-line shareholders. Lecture Notes in Computer Science, 1431:60–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1431/14310060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1431/14310060.
                                        pdf.

Okamoto:1998:LBT

[2492] Tatsuaki Okamoto and Moti Yung. Lower bounds on term-based divisible cash systems. Lecture Notes in Computer Science, 1431:72–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1431/14310072.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1431/14310072.
                                        pdf.

Lehti:1998:CT

[2493] Ilari Lehti and Pekka Nikander. Certifying trust. Lecture Notes in Computer Science, 1431:83–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1431/14310083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1431/14310083.
pdf.

Merkle:1998:SSA

[2494] Johannes Merkle and Ralph Werchner. On the security of serveraided RSA protocols. Lecture Notes in Computer Science, 1431:99–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1431/14310099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1431/14310099.
pdf.

Tsiounis:1998:SEB

[2495] Yiannis Tsiounis and Moti Yung. On the security of ElGamal-based encryption. Lecture Notes in Computer Science, 1431:117–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1431/14310117.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1431/14310117.pdf.

Hirose:1998:ADH

[2496] Shouichi Hirose and Susumu Yoshida.

An authenticated Diffie–Hellman key agreement protocol secure against active attacks. Lecture Notes in Computer Science, 1431:135–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1431/14310135.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1431/14310135.
                                        pdf.

Saeednia:1998:SGI

[2497] Shahrokh Saeednia and Rei Safavi-Naini. On the security of Girault’s identification scheme. Lecture Notes in Computer Science, 1431:149–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1431/14310149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1431/14310149.
                                        pdf.

Chang:1998:SOM

[2498] Chin-Chen Chang, Jyh-Jong Leu, Pai-Cheng Huang, and Wei-Bin Lee. A scheme for obtaining a message from the digital multisignature. Lecture Notes in Computer Science, 1431:154– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1431/14310154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1431/14310154.
                                        pdf.

Sakai:1998:SHC

[2499] Yasuyuki Sakai, Kouichi Sakurai, and Hirokazu Ishizuka. Secure hyperelliptic cryptosystems and their performance. Lecture Notes in Computer Science, 1431:164–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1431/14310164.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/1431/14310164.
pdf.

Hasegawa:1998:PIE

[2500] Toshio Hasegawa, Junko Nakajima, and Mitsuru Matsui. A practical implementation of elliptic curve cryptosystems over GF(p) on a 16bit microcomputer. Lecture Notes in Computer Science, 1431:182–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1431/14310182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1431/14310182.
pdf.

Cheon:1998:TEA

[2501] Jung Hee Cheon, Sungmo Park, Sangwoo Park, and Daeho Kim. Two efficient algorithms for arithmetic of elliptic curves using Frobenius map. Lecture Notes in Computer Science, 1431: 195–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1431/14310195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1431/14310195.
pdf.

Yamamura:1998:PKC

[2502] Akihiro Yamamura. Public-key cryptosystems using the modular group. Lecture Notes in Computer Science, 1431:203–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/1431/14310203.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1431/14310203.
                                        pdf.

Mihaljevic:1998:CAB

[2503] Miodrag Mihaljevic, Yuliang Zheng, and Hideki Imai. A cellular automaton based fast one-way hash function suitable for hardware implementation. Lecture Notes in Computer Science, 1431: 217–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1431/14310217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1431/14310217.
                                        pdf.

Shin:1998:NHF

[2504] Sang Uk Shin, Kyung Hyune Rhee, Dae Hyun Ryu, and Sang Jin Lee. A new hash function based on MDxfamily and its application to MAC. Lecture Notes in Computer Science, 1431: 234–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1431/14310234.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1431/14310234.
                                        pdf.

David:1998:SIC

[2505] Michael W. David and Kouichi Sakurai. Security issues for contactless smart cards. Lecture Notes in Computer Science, 1431:247–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1431/14310247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1431/14310247.
pdf.

Izu:1998:PSE

[2506] Tetsuya Izu, Jun Kogure, Masayuki Noro, and Kazuhiro Yokoyama. Parameters for secure elliptic curve cryptosystem — improvements on Schoof’s algorithm. Lecture Notes in Computer Science, 1431:253–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1431/14310253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1431/14310253.
pdf.

Mambo:1998:NCB

[2507] Masahiro Mambo and Hiroki Shizuya.

A note on the complexity of breaking Okamoto-Tanaka ID-based key exchange scheme. Lecture Notes in Computer Science, 1431:258–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1431/14310258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1431/14310258.
pdf.

Anonymous:1998:AIk

[2508] Anonymous. Author index. Lecture Notes in Computer Science, 1431:263– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        papers/1431/1431indx.pdf.

Goldberg:1998:RDM

[2509] A. V. Goldberg. Recent developments in maximum flow algorithms (invited lecture). Lecture Notes in Computer Science, 1432:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aleksandrov:1998:AAW

[2510] L. Aleksandrov, M. Lanthier, A. Maheshwari, and J.-R. Sack. An — approximation algorithm for weighted shortest paths on polyhedral surfaces. Lecture Notes in Computer Science, 1432:11–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhatia:1998:FLD

[2511] R. Bhatia, S. Guha, S. Khuller, and Y. J. Sussmann. Facility location with dynamic distance functions. Lecture Notes in Computer Science, 1432:23– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1998:ASP

[2512] K. Jansen. An approximation scheme for bin packing with conflicts. Lecture Notes in Computer Science, 1432:35– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aspvall:1998:AGB

[2513] B. Aspvall, M. M. Halldorsson, and F. Manne. Approximations for the general block distribution of a matrix. Lecture Notes in Computer Science, 1432:47–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1998:OAC

[2514] T. Graf and K. Veezhinathan. An optimal algorithm for computing visible nearest foreign neighbors among colored line segments. Lecture Notes in Computer Science, 1432:59–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1998:MAA

[2515] F. Hoffmann, C. Icking, R. Klein, and K. Kriegel. Moving an angle around a region. Lecture Notes in Computer Science, 1432:71–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBerg:1998:MMP

[2516] M. De Berg, M. J. Katz, M. Overmars, and A. F. Van Der Stappen. Models and motion planning. Lecture Notes in Computer Science, 1432:83–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katz:1998:CSC

[2517] M. J. Katz, K. Kedem, and M. Segal. Constrained square-center problems. Lecture Notes in Computer Science, 1432:95–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodal:1998:WCE

[2518] G. S. Brodal and J. Katajainen. Worstcase efficient external-memory priority queues. Lecture Notes in Computer Science, 1432:107–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaplan:1998:SCP

[2519] H. Kaplan, C. Okasaki, and R. E. Tarjan. Simple confluently persistent catenable lists. Lecture Notes in Computer Science, 1432:119–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raman:1998:IUB

[2520] V. Raman and S. RamNath. Improved upper bounds for time-space tradeoffs for selection with limited storage. Lecture Notes in Computer Science, 1432:131–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sridhar:1998:PDS

[2521] R. Sridhar, K. Rajasekar, and C. P. Rangan. Probabilistic data structures for priority queues. Lecture Notes in Computer Science, 1432:143–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zuckerman:1998:EWR

[2522] D. Zuckerman. Extractors for weak random sources and their applications (invited lecture). Lecture Notes in Computer Science, 1432:155–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodal:1998:CNB

[2523] G. S. Brodal and M. C. Pinotti. Comparator networks for binary heap construction. Lecture Notes in Computer Science, 1432:158–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1998:TVL

[2524] D. Z. Chen and J. Xu. Two-variable linear programming in parallel. Lecture Notes in Computer Science, 1432: 169–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grossi:1998:ODP

[2525] R. Grossi, A. Pietracaprina, and G. Pucci. Optimal deterministic protocols for mobile robots on a grid. Lecture Notes in Computer Science, 1432: 181–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMarco:1998:CMW

[2526] G. De Marco, L. Gargano, and U. Vaccaro. Concurrent multicast in weighted networks. Lecture Notes in Computer Science, 1432:193–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haastad:1998:SRS

[2527] J. Haastad. Some recent strong inapproximability results (invited lecture). Lecture Notes in Computer Science, 1432:205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahlhaus:1998:MEP

[2528] E. Dahlhaus. Minimal elimination of planar graphs. Lecture Notes in Computer Science, 1432:210–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aspvall:1998:MRT

[2529] B. Aspvall and A. J. A. Proskurowski Telle. Memory requirements for table computations in partial k-tree algorithms. Lecture Notes in Computer Science, 1432:222–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrett:1998:FLC

[2530] C. Barrett, R. Jacob, and M. Marathe. Formal language constrained path problems. Lecture Notes in Computer Science, 1432:234–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirsch:1998:LSA

[2531] E. A. Hirsch. Local search algorithms for SAT: Worst-case analysis. Lecture Notes in Computer Science, 1432:246– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1998:SMP

[2532] P. Berman and C. Coulston. Speed is more powerful than clairvoyance. Lecture Notes in Computer Science, 1432: 255–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seiden:1998:ROM

[2533] S. S. Seiden. Randomized online multithreaded paging. Lecture Notes in Computer Science, 1432:264–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mahajan:1998:DOA

[2534] M. Mahajan and V. Vinay. Determinant: Old algorithms, new insights. Lecture Notes in Computer Science, 1432:276–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1998:SFP

[2535] J. F. Sibeyn. Solving fundamental problems on sparse-meshes. Lecture Notes in Computer Science, 1432:288– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sleumer:1998:OSC

[2536] N. Sleumer. Output-sensitive cell enumeration in hyperplane arrangements. Lecture Notes in Computer Science, 1432:300–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Will:1998:FEC

[2537] H.-M Will. Fast and efficient computation of additively weighted Voronoi cells for applications in molecular biology. Lecture Notes in Computer Science, 1432:310–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aronov:1998:NRV

[2538] B. Aronov, A. Eftrat, D. Halperin, and M. Sharir. On the number of regular vertices of the union of Jordan regions. Lecture Notes in Computer Science, 1432:322–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sen:1998:DSA

[2539] S. Sen and N. Gupta. Distributionsensitive algorithms. Lecture Notes in Computer Science, 1432:335–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1998:RAO

[2540] Kevin J. Lang, Barak A. Pearlmutter, and Rodney A. Price. Results of the abbadingo one DFA learning competition and a new evidencedriven state merging algorithm. Lecture Notes in Computer Science, 1433: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330001.
pdf.

Rossmanith:1998:LVP

[2541] Peter Rossmanith and Thomas Zeugmann. Learning k-variable pattern languages efficiently stochastically finite on average from positive data. Lecture Notes in Computer Science, 1433:13– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330013.
pdf.

Tellier:1998:MHL

[2542] Isabelle Tellier. Meaning helps learning syntax. Lecture Notes in Computer Science, 1433:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330025.
pdf.

Parekh:1998:PTI

[2543] Rajesh Parekh, Codrin Nichitiu, and Vasant Honavar. A polynomial time incremental algorithm for learning DFA. Lecture Notes in Computer Science, 1433:37–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330037.
pdf.

Oncina:1998:DDA

[2544] José Oncina. The data driven approach applied to the OSTIA algorithm. Lecture Notes in Computer Science, 1433: 50–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330050.
pdf.

Keselj:1998:GMG

[2545] Vlado Keselj. Grammar model and grammar induction in the system NL PAGE. Lecture Notes in Computer Science, 1433:57–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1433/14330057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1433/14330057.
                                        pdf.

Castellanos:1998:ALR

[2546] Antonio Castellanos. Approximate learning of random subsequential transducers. Lecture Notes in Computer Science, 1433:67–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1433/14330067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1433/14330067.
                                        pdf.

delaHiguera:1998:LSF

[2547] Colin de la Higuera. Learning stochastic finite automata from experts. Lecture Notes in Computer Science, 1433: 79–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1433/14330079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1433/14330079.
                                        pdf.

Higuera:1998:LSF

[2548] Colin de la Higuera. Learning stochastic finite automata from experts. Lecture Notes in Computer Science, 1433: 79–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1433/14330079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330079.
pdf.

Firoiu:1998:LDF

[2549] Laura Firoiu, Tim Oates, and Paul R.

Cohen. Learning a deterministic finite automaton with a recurrent neural network. Lecture Notes in Computer Science, 1433:90–101, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330090.
pdf.

Chodorowski:1998:AGI

[2550] Jacques Chodorowski and Laurent Miclet. Applying grammatical inference in learning a language model for oral dialogue. Lecture Notes in Computer Science, 1433:102–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330102.
pdf.

Feldman:1998:RLL

[2551] Jerome A. Feldman. Real language learning. Lecture Notes in Computer Science, 1433:114–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1433/14330114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1433/14330114.
                                        pdf.

Juille:1998:SSA

[2552] Hugues Juillé and Jordan B. Pollack. A stochastic search approach to grammar induction. Lecture Notes in Computer Science, 1433:126–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1433/14330126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1433/14330126.
                                        pdf.

Pico:1998:TLE

[2553] David Picó and Enrique Vidal. Transducer-learning experiments on language understanding. Lecture Notes in Computer Science, 1433:138–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1433/14330138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1433/14330138.
                                        pdf.

Ruiz:1998:LTT

[2554] José Ruiz, Salvador España, and Pedro García. Locally threshold testable languages in strict sense: Application to the inference problem. Lecture Notes in Computer Science, 1433:150– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1433/14330150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330150.
pdf.

Sempere:1998:LSL

[2555] José M. Sempere and G. Nagaraja. Learning a subclass of linear languages from positive structural information. Lecture Notes in Computer Science, 1433:162–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330162.
pdf.

Saidi:1998:GID

[2556] Alexander S. Saidi and Souad Tayebbey. Grammatical inference in document recognition. Lecture Notes in Computer Science, 1433:175–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330175.
pdf.

Carrasco:1998:SIR

[2557] Rafael C. Carrasco, Jose Oncina, and Jorge Calera. Stochastic inference of regular tree languages. Lecture Notes in Computer Science, 1433:187– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1433/14330187.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1433/14330187.
                                        pdf.

Coste:1998:HCI

[2558] François Coste and Jacques Nicolas. How considering incompatible state mergings may reduce the DFA induction search tree. Lecture Notes in Computer Science, 1433:199–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1433/14330199.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1433/14330199.
                                        pdf.

Cruz-Alcazar:1998:LRG

[2559] Pedro P. Cruz-Alcázar and Enrique Vidal-Ruiz. Learning regular grammars to model musical style: Comparing different coding schemes. Lecture Notes in Computer Science, 1433:211– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1433/14330211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1433/14330211.
                                        pdf.

Emerald:1998:LSC

[2560] J. D. Emerald, K. G. Subramanian, and D. G. Thomas. Learning a subclass of context-free languages. Lecture Notes in Computer Science, 1433: 223–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1433/14330223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330223.
pdf.

Dupont:1998:USC

[2561] Pierre Dupont and Lin Chase. Using symbol clustering to improve probabilistic automaton inference. Lecture Notes in Computer Science, 1433:232– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1433/14330232.
pdf.

Viechnicki:1998:PEA

[2562] Peter Viechnicki. A performance evaluation of automatic survey classifiers. Lecture Notes in Computer Science, 1433:244–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1433/14330244.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1433/14330244.
pdf.

Brzma:1998:PDB

[2563] Alvis Brāzma, Inge Jonassen, Jaak Vilo, and Esko Ukkonen. Pattern discovery in biosequences. Lecture Notes in Computer Science, 1433:257– 270, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1433/14330257.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1433/14330257.
                                        pdf.

Brazma:1998:PDB

[2564] A. Brazma, I. Jonassen, J. Vilo, and E. Ukkonen. Pattern discovery in biosequences. Lecture Notes in Computer Science, 1433:257–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:AIl

[2565] Anonymous. Author index. Lecture Notes in Computer Science, 1433:271– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1433/1433auth.pdf.

Lee:1998:RFC

[2566] Won-Sook Lee, Elwin Lee, and Nadia Magnenat Thalmann. Real face communication in a virtual world. Lecture Notes in Computer Science, 1434: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340001.
                                        pdf.

Aubel:1998:AIR

[2567] Amaury Aubel, Ronan Boulic, and Daniel Thalmann. Animated impostors for real-time display of numerous virtual humans. Lecture Notes in Computer Science, 1434:14– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340014.
pdf.

Verna:1998:CWD

[2568] Didier Verna and Alain Grumbach.

Can we define virtual reality? the M_{R} IC model. Lecture Notes in Computer Science, 1434:29–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340029.
pdf.

Ryan:1998:DDV

[2569] Matthew D. Ryan and Paul M.

Sharkey. Distortion in distributed virtual environments. Lecture Notes in Computer Science, 1434:42–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340042.
pdf.

Richardson:1998:VBB

[2570] John F. Richardson. VRML based behaviour database editor. Lecture Notes in Computer Science, 1434:49– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340049.
                                        pdf.

Semwal:1998:SVE

[2571] Sudhanshu Kumar Semwal and Jun Ohya. The scan&track virtual environment. Lecture Notes in Computer Science, 1434:63–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1434/14340063.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1434/14340063.pdf.

Numaoka:1998:CVB

[2572] Chisato Numaoka. CyberGlass: Vision-based VRML2 navigator. Lecture Notes in Computer Science, 1434: 81–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340081.
                                        pdf.

Flaig:1998:WTA

[2573] Thomas Flaig. Work task analysis and selection of interaction devices in virtual environments. Lecture Notes in Computer Science, 1434:88– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340088.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1434/14340088.
pdf.

Richard:1998:ESV

[2574] Paul Richard, Philippe Hareux, Philippe Coiffet, and Grigore Burdea. Effect of stereoscopic viewing on human tracking performance in dynamic virtual environments. Lecture Notes in Computer Science, 1434:97– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340097.
pdf.

Nakatsu:1998:IMV

[2575] Ryohei Nakatsu, Naoko Tosa, and Takeshi Ochi. Interactive movie: A virtual world with narratives. Lecture Notes in Computer Science, 1434:107– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340107.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340107.
pdf.

Park:1998:RIB

[2576] Jong-Il I. Park and Seiki Inoue. Realimage-based virtual studio. Lecture Notes in Computer Science, 1434:117– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340117.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1434/14340117.
                                        pdf.

Carraro:1998:PV

[2577] Gianpaolo U. Carraro, John T. Edmark, and J. Robert Ensor. Pop-out videos. Lecture Notes in Computer Science, 1434:123–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340123.
                                        pdf.

Gerard:1998:CSC

[2578] Philippe Gérard, Clifton Phillips, and Ramesh Jain. Color segmentation and color correction using lighting and white balance shifts. Lecture Notes in Computer Science, 1434:129– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340129.
                                        pdf.

Ventrella:1998:DEA

[2579] Jeffrey Ventrella. Designing emergence in animated artificial life worlds. Lecture Notes in Computer Science, 1434: 143–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340143.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1434/14340143.
pdf.

Pagliarini:1998:AMW

[2580] Luigi Pagliarini, Ariel Dolan, Filippo Menczer, and Henrik Hautop Lund. ALife meets Web: Lessons learned. Lecture Notes in Computer Science, 1434:156–167, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340156.
pdf.

Proctor:1998:IFD

[2581] Glenn Proctor and Chris Winter. Information flocking: Data visualisation in virtual worlds using emergent behaviours. Lecture Notes in Computer Science, 1434:168–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340168.
pdf.

Damer:1998:NGP

[2582] Bruce Damer, Karen Marcelo, and Frank Revi. Nerve garden: A public terrarium in cyberspace. Lecture Notes in Computer Science, 1434:177– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340177.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340177.
                                        pdf.

Tyran:1998:TDV

[2583] Jean-Luc L. Tyran. A two-dimensional virtual world to explain the genetic code structure ? Lecture Notes in Computer Science, 1434:186–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340186.
                                        pdf.

Robert:1998:GAE

[2584] Antony Robert, Fabrice Chantemargue, and Michèle Courant. Grounding agents in EMud artificial worlds. Lecture Notes in Computer Science, 1434: 193–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340193.
                                        pdf.

Servat:1998:TVE

[2585] David Servat, Edith Perrier, Jean-Pierre Treuil, and Alexis Drogoul. Towards virtual experiment laboratories: How multi-agent simulations can cope with multiple scales of analysis and viewpoints. Lecture Notes in Computer Science, 1434:205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1434/14340205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340205.
pdf.

Lattaud:1998:MEE

[2586] Claude Lattaud and Cristina Cuenca.

A model for the evolution of environments. Lecture Notes in Computer Science, 1434:218–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1434/14340218.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1434/14340218.pdf.

Reigner:1998:AVR

[2587] P. Reigner, F. Harrouet, S. Morvan, and J. Tisseau. AReVi: A virtual reality multi-agent platform. Lecture Notes in Computer Science, 1434:229– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reignier:1998:AVR

[2588] P. Reignier, F. Harrouet, S. Morvan, J. Tisseau, and T. Duval. ARéVi: A virtual reality multiagent platform. Lecture Notes in Computer Science, 1434:229–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340229.
pdf.

Noda:1998:ICV

[2589] Itsuki Noda and Ian Frank. Investigating the complex with virtual soccer. Lecture Notes in Computer Science, 1434:241–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1434/14340241.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340241.
                                        pdf.

Michel:1998:WSB

[2590] Olivier Michel. Webots: Symbiosis between virtual and real mobile robots. Lecture Notes in Computer Science, 1434:254–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340254.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340254.
                                        pdf.

DeMeneses:1998:VSW

[2591] Y. L. De Meneses and O. Michel. Vision sensors on the webots simulator. Lecture Notes in Computer Science, 1434:264–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meneses:1998:VSW

[2592] Yuri López de Meneses and Olivier Michel. Vision sensors on the webots simulator. Lecture Notes in Computer Science, 1434:264–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340264.
pdf.

Hutzler:1998:GVW

[2593] Guillaume Hutzler, Bernard Gortais, and Alexis Drogoul. Grounding virtual worlds in reality. Lecture Notes in Computer Science, 1434:274–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340274.
pdf.

Kaplan:1998:GVC

[2594] Frédéric Kaplan, Angus McIntyre, Chisato Numaoka, and Silvère Tajan. Growing virtual communities in 3D meeting spaces. Lecture Notes in Computer Science, 1434:286–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340286.
pdf.

Pachet:1998:MIM

[2595] François Pachet and Olivier Delerue.

A mixed 2D/3D interface for music spatialization. Lecture Notes in Computer Science, 1434:298–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340298.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340298.
                                        pdf.

Cavallar:1998:OI

[2596] Claudia Cavallar and Daniel Dögl. Organizing information in 3D. Lecture Notes in Computer Science, 1434:308– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340308.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340308.
                                        pdf.

Doegl:1998:OI

[2597] D. Doegl and C. Cavallar. Organizing information in 3D. Lecture Notes in Computer Science, 1434:308–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kimoto:1998:HCV

[2598] Haruo Kimoto. Human centered virtual interactive image world for image retrieval. Lecture Notes in Computer Science, 1434:315–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340315.
                                        pdf.

Refsland:1998:VGB

[2599] Scot Thrane Refsland, Takeo Ojika, Tom Defanti, Andy Johnson, Jason Leigh, Carl Loeffler, and Xiaoyuan Tu. Virtual great barrier reef: A theoretical approach towards an evolving, interactive VR environment using a distributed DOME and CAVE system. Lecture Notes in Computer Science, 1434:323–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340323.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340323.
pdf.

Williams:1998:DIH

[2600] Matthew Williams, Damian Schofield, and Bryan Denby. The development of an intelligent haulage truck simulator for improving the safety of operation in surface mines. Lecture Notes in Computer Science, 1434:337– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340337.
pdf.

Bourdakis:1998:NLV

[2601] Vassilis Bourdakis. Navigation in large VR urban models. Lecture Notes in Computer Science, 1434:345– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1434/14340345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340345.
                                        pdf.

Kisseleva:1998:AVW

[2602] Olga Kisseleva. Art and virtual worlds. Lecture Notes in Computer Science, 1434:357–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1434/14340357.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340357.
                                        pdf.

Bizri:1998:MVS

[2603] Hisham Bizri, Andrew Johnson, and Christina Vasilakis. Las meninas in VR: Storytelling and the illusion in art. Lecture Notes in Computer Science, 1434:360–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340360.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1434/14340360.
                                        pdf.

Roussos:1998:MML

[2604] Maria Roussos and Hisham Bizri. Mitologies: Medieval labyrinth narratives in virtual reality. Lecture Notes in Computer Science, 1434:373–383, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1434/14340373.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1434/14340373.
pdf.

Muzhesky:1998:AWV

[2605] Vladimir Muzhesky. Aggregate worlds:

Virtual architecture aftermath. Lecture Notes in Computer Science, 1434: 384–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340384.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340384.
pdf.

Manovich:1998:ZVR

[2606] Lev Manovich. Zeuxis vs. reality engine: Digital realism and virtual worlds. Lecture Notes in Computer Science, 1434:394–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1434/14340394.
pdf.

Hayem:1998:ANF

[2607] Raphael Hayem, Tancrede Fourmaintraux, Keran Petit, Nicolas Rauber, and Olga Kisseleva. Avatars: New fields of implication. Lecture Notes in Computer Science, 1434:406–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1434/14340406.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1434/14340406.
pdf.

Parunak:1998:WCA

[2608] H. Van Dyke Parunak. What can agents do in industry, and why? an overview of industriallyoriented R&D at CEC. Lecture Notes in Computer Science, 1435:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1435/14350001.
pdf.

Nodine:1998:IAS

[2609] Marian Nodine. The InfoSleuth agent system. Lecture Notes in Computer Science, 1435:19–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1435/14350019.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1435/14350019.pdf.

Lazarou:1998:AHI

[2610] Vasilios S. Lazarou and Keith L. Clark. Agents for hypermedia information discovery. Lecture Notes in Computer Science, 1435:21–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1435/14350021.
pdf.

Moukas:1998:TDC

[2611] Alexandros Moukas, Konstantinos Chandrinos, and Pattie Maes. Trafficopter: A distributed collection system for traffic information. Lecture Notes in Computer Science, 1435:33– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1435/14350033.
pdf.

Deschner:1998:ASI

[2612] Dominik Deschner, Oliver Hofmann, Stefan Reinheimer, and Freimut Bodendorf. Agent-supported information retrieval for tracking and tracing. Lecture Notes in Computer Science, 1435: 44–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1435/14350044.
pdf.

Durfee:1998:DUS

[2613] Edmund H. Durfee, Tracy Mullen, Sunju Park, José M. Vidal, and Peter Weinstein. The dynamics of the UMDL service market society. Lecture Notes in Computer Science, 1435:55– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350055.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1435/14350055.
                                        pdf.

Huhns:1998:MSI

[2614] Michael N. Huhns and Munindar P. Singh. Multiagent systems in information-rich environments. Lecture Notes in Computer Science, 1435:79– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1435/14350079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1435/14350079.
                                        pdf.

Chalasani:1998:SQI

[2615] Prasad Chalasani, Somesh Jha, Onn Shehory, and Katia P. Sycara. Strategies for querying information agents. Lecture Notes in Computer Science, 1435:94–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1435/14350094.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1435/14350094.
                                        pdf.

Kido:1998:GCP

[2616] Takashi Kido. Grand challenge problems on cross cultural communication — toward socially intelligent agents. Lecture Notes in Computer Science, 1435:108–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1435/14350108.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1435/14350108.
pdf.

Sandholm:1998:AEC

[2617] Tuomas Sandholm. Agents in electronic commerce: Component technologies for automated negotiation and coalition formation. Lecture Notes in Computer Science, 1435:113–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1435/14350113.
pdf.

Guttman:1998:CVC

[2618] Robert H. Guttman and Pattie Maes.

Cooperative vs. competitive multiagent negotiations in retail electronic commerce. Lecture Notes in Computer Science, 1435:135–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1435/14350135.
pdf.

Vogler:1998:EMA

[2619] Hartmut Vogler, Marie-Luise Moschgath, and Thomas Kunkelmann. Enhancing mobile agents with electronic commerce capabilities. Lecture Notes in Computer Science, 1435:148–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1435/14350148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1435/14350148.
                                        pdf.

Kephart:1998:DIF

[2620] Jeffrey O. Kephart, James E. Hanson, David W. Levine, Benjamin N. Grosof, Jakka Sairamesh, Richard B. Segal, and Steve R. White. Dynamics of an information-filtering economy. Lecture Notes in Computer Science, 1435:160– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1435/14350160.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1435/14350160.
                                        pdf.

Sycara:1998:LAS

[2621] Katia P. Sycara. Levels of adaptivity in systems of coordinating information agents. Lecture Notes in Computer Science, 1435:172–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1435/14350172.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1435/14350172.
                                        pdf.

Sen:1998:ACI

[2622] Sandip Sen. Adaptive choice of information sources (extended abstract). Lecture Notes in Computer Science, 1435:190–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/1435/14350190.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1435/14350190.
pdf.

Green:1998:AMC

[2623] Shaw Green, Pádraig Cunningham, and Fergal Somers. Agent mediated collaborative Web page filtering. Lecture Notes in Computer Science, 1435: 195–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1435/14350195.
pdf.

Delgado:1998:CBC

[2624] Joaquin Delgado, Naohiro Ishii, and Tomoki Ura. Content-based collaborative information filtering: Actively learning to classify and recommend documents. Lecture Notes in Computer Science, 1435:206–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1435/14350206.
pdf.

Theilmann:1998:DEI

[2625] Wolfgang Theilmann and Kurt Rothermel. Domain experts for information retrieval in the World Wide Web. Lecture Notes in Computer Science, 1435: 216–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1435/14350216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1435/14350216.
                                        pdf.

Benn:1998:SNM

[2626] Wolfgang Benn and Otmar Görlitz. Semantic navigation maps for information agents. Lecture Notes in Computer Science, 1435:228–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1435/14350228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1435/14350228.
                                        pdf.

Tolksdorf:1998:CPM

[2627] Robert Tolksdorf. Coordination patterns of mobile information agents. Lecture Notes in Computer Science, 1435: 246–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1435/14350246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1435/14350246.
                                        pdf.

Gehmeyr:1998:MIA

[2628] Andreas Gehmeyr, Jürgen Müller, and Albert Schappert. Mobile information agents on the Web. Lecture Notes in Computer Science, 1435:262– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1435/14350262.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1435/14350262.
pdf.

Corradi:1998:MAM

[2629] Antonio Corradi, Marco Cremonini, and Cesare Stefanelli. Melding abstractions with mobile agents. Lecture Notes in Computer Science, 1435:278– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1435/14350278.
pdf.

Bonatti:1998:DSH

[2630] Piero A. Bonatti, Sarit Kraus, Jose Salinas, and V. S. Subrahmanian. Data-security in heterogeneous agent systems. Lecture Notes in Computer Science, 1435:290–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1435/14350290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1435/14350290.
pdf.

Wood:1998:WTP

[2631] D. Wood. WIA and theory and practice. Lecture Notes in Computer Science, 1436:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1998:CEG

[2632] M. A. Alonso, D. Cabrero, and M. Vilares. Construction of efficient generalized LR parsers. Lecture Notes in Computer Science, 1436:7–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertault:1998:VDA

[2633] F. Bertault and G. Kucherov. Visualization of dynamic automata using padnon. Lecture Notes in Computer Science, 1436:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brzozowski:1998:AAB

[2634] J. A. Brzozowski and R. Negulescu. Automata of asynchronous behaviors. Lecture Notes in Computer Science, 1436:29–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caron:1998:LMP

[2635] P. Caron. LANGAGE: A Maple package for automaton characterization of regular languages. Lecture Notes in Computer Science, 1436:46–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garzon:1998:VIF

[2636] M. Garzon, Y. Gao, J. A. Rose, and R. C. Murphy. In vitro implementation of finite-state machines. Lecture Notes in Computer Science, 1436:56– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiraz:1998:MTA

[2637] G. A. Kiraz and E. Grimley-Evans. Multi-tape automata for speech and language systems: A Prolog implementation. Lecture Notes in Computer Science, 1436:87–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeMaout:1998:TIA

[2638] V. Le Maout. Tools to implement automata, a first step: ASTL. Lecture Notes in Computer Science, 1436:104– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lescow:1998:MSS

[2639] H. Lescow and J. Voege. Minimal separating sets for Muller automata. Lecture Notes in Computer Science, 1436:109–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurel:1998:PMT

[2640] D. Maurel and L. Chauvier. Pseudominimal transducers: A transducer with proper elements. Lecture Notes in Computer Science, 1436:122–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1998:TMN

[2641] C. L. Miller. Towards a meta-normal form algorithm for context-free grammars. Lecture Notes in Computer Science, 1436:133–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohri:1998:RDW

[2642] M. Mohri, F. Pereira, and M. Riley. A rational design for a weighted finite-state transducer library. Lecture Notes in Computer Science, 1436:144– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ponty:1998:ENF

[2643] J.-L. Ponty. An efficient null-free procedure for deciding regular language membership. Lecture Notes in Computer Science, 1436:159–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Revuz:1998:OEA

[2644] D. Revuz. Operations on extended automata. Lecture Notes in Computer Science, 1436:171–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomaa:1998:EIR

[2645] K. Salomaa, X. Wu, and S. Yu. Efficient implementation of regular languages using R-AFA. Lecture Notes in Computer Science, 1436:176–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silberztein:1998:IIF

[2646] M. Silberztein. INTEX: An integrated FST toolbox. Lecture Notes in Computer Science, 1436:185–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trahtman:1998:PEO

[2647] A. N. Trahtman. A precise estimation of the order of local testability of a deterministic finite automaton. Lecture Notes in Computer Science, 1436:198–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Velinov:1998:TTC

[2648] Y. Velinov. Teaching theory of computation with tape machines. Lecture Notes in Computer Science, 1436:213– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voege:1998:ATP

[2649] J. Voege, S. Ulbrand, O. Matz, and N. Buhrke. The automata theory package omega. Lecture Notes in Computer Science, 1436:228–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1998:POA

[2650] B. W. Watson. Practical optimizations for automata. Lecture Notes in Computer Science, 1436:232–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziadi:1998:SDT

[2651] D. Ziadi. Sorting and doubling techniques for set partitioning and automata minimization problems. Lecture Notes in Computer Science, 1436: 241–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frei:1998:DHI

[2652] Christian Frei and Boi Faltings. A dynamic hierarchy of intelligent agents for network management. Lecture Notes in Computer Science, 1437:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370001.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1437/14370001.
pdf.

Lissajoux:1998:GAP

[2653] T. Lissajoux, V. Hilaire, A. Koukam, and A. Caminada. Genetic algorithms as prototyping tools for multi-agent systems: Application to the antenna parameter setting problem. Lecture Notes in Computer Science, 1437:17– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370017.
pdf.

Bernardo:1998:SSDa

[2654] Luis Bernardo and Paulo Pinto. Scalable service deployment on highly populated networks. Lecture Notes in Computer Science, 1437:29–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370029.
pdf.

Hayzelden:1998:HMA

[2655] A. L. G. Hayzelden and J. Bigham. Heterogeneous multi-agent architecture for ATM virtual path network resource configuration. Lecture Notes in Computer Science, 1437:45–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1437/14370045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370045.
pdf.

Bonabeau:1998:RTN

[2656] Eric Bonabeau, Florian Henaux, Sylvain Guérin, Dominique Snyers, Pascale Kuntz, and Guy Theraulaz. Routing in telecommunications networks with ant-like agents. Lecture Notes in Computer Science, 1437:60– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370060.
pdf.

Pagurek:1998:NCM

[2657] Bernard Pagurek, Yanrong Li, Andrzej Bieszczad, and Gatot Susilo. Network configuration management in heterogeneous ATM environments. Lecture Notes in Computer Science, 1437:72– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370072.
pdf.

Bieszczad:1998:ABS

[2658] Andrzej Bieszczad, Syed Kamran Raza, Bernard Pagurek, and Tony White. Agent-based schemes for plug-and-play network components. Lecture Notes in Computer Science, 1437:89– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370089.
pdf.

Gibney:1998:DRA

[2659] M. A. Gibney and N. R. Jennings. Dynamic resource allocation by market-based routing in telecommunications networks. Lecture Notes in Computer Science, 1437:102–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370102.
pdf.

Corley:1998:AIM

[2660] Stephen Corley, Diego Magro, Fabio Malabocchia, Jens Meinköhn, Luisella Sisto, Sahin Albayrak, and Alexander Grosse. The application of intelligent and mobile agents to the management of software problems in telecommunications. Lecture Notes in Computer Science, 1437:118–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370118.
pdf.

White:1998:DFL

[2661] Tony White, Andrzej Bieszczad, and Bernard Pagurek. Distributed fault location in networks using mobile agents. Lecture Notes in Computer Science, 1437:130–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370130.
pdf.

Nicklisch:1998:IAB

[2662] Jan Nicklisch, Jürgen Quittek, Andreas Kind, and Shinya Arao. INCA: An agent-based network control architecture. Lecture Notes in Computer Science, 1437:142–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370142.
pdf.

Davies:1998:AMA

[2663] Guy Davies and Love Ekenberg. Agent metaphors for analysing telematic services. Lecture Notes in Computer Science, 1437:156–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1437/14370156.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1437/14370156.pdf.

Abu-Hakima:1998:MAT

[2664] Suhayya Abu-Hakima, Roger Impey, Ramiro Liscano, and Amir Zeid. A multi-agent testbed and an agent launch tool for diverse seamless personal information networking applications. Lecture Notes in Computer Science, 1437:171–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370171.
pdf.

Hooda:1998:ANS

[2665] A. Hooda, A. Karmouch, and S. Abu-Hakima. Agent negotiation for supporting personal mobility. Lecture Notes in Computer Science, 1437:190– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370190.
pdf.

Garijo:1998:DMA

[2666] Francisco Garijo, Juan Tous, José M. Matias, Stephen Corley, and Marius Tesselaar. Development of a multiagent system for cooperative work with network negotiation capabilities. Lecture Notes in Computer Science, 1437: 204–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370204.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1437/14370204.
pdf.

Dillenseger:1998:ICB

[2667] Bruno Dillenseger. From interoperability to cooperation: Building intelligent agents on middleware. Lecture Notes in Computer Science, 1437:220– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370220.
pdf.

Wieczorek:1998:OSA

[2668] Dirk Wieczorek and Sahin Albayrak.

Open scalable agent architecture for telecommunication applications. Lecture Notes in Computer Science, 1437: 233–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1437/14370233.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1437/14370233.
pdf.

Kaijser:1998:RSD

[2669] Per Kaijser. A review of the SESAME development. Lecture Notes in Computer Science, 1438:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380001.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1438/14380001.
                                        pdf.

Mueller:1998:SPK

[2670] S. Mueller and W. B. Mueller. The security of public key cryptosystems based on integer factorization. Lecture Notes in Computer Science, 1438:9–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1998:SPK

[2671] Siguna Müller and Winfried B. Müller. The security of public key cryptosystems based on integer factorization. Lecture Notes in Computer Science, 1438:9–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380009.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380009.
                                        pdf.

Ashley:1998:UAS

[2672] Paul Ashley, Mark Vandenwauver, and Bradley Broom. A uniform approach to securing Unix applications using SESAME. Lecture Notes in Computer Science, 1438:24–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380024.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380024.
                                        pdf.

Falk:1998:IMN

[2673] Rainer Falk and Markus Trommer. Integrated management of network and host based security mechanisms. Lecture Notes in Computer Science, 1438: 36–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380036.
pdf.

Horton:1998:CDC

[2674] Jeffrey Horton and Jennifer Seberry.

Covert distributed computing using Java through Web spoofing. Lecture Notes in Computer Science, 1438:48– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380048.
pdf.

Yi:1998:DCB

[2675] Xun Yi, Kwok Yan Lam, and Yongfei Han. Differential cryptanalysis of a block cipher. Lecture Notes in Computer Science, 1438:58– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380058.
pdf.

Sun:1998:PKC

[2676] Hung-Min Sun and Shiuh-Pyng Shieh. On private-key cryptosystems based on product codes. Lecture Notes in Computer Science, 1438:68–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380068.
                                        pdf.

Carter:1998:KSI

[2677] G. Carter, E. Dawson, and L. Nielsen. Key schedules of iterative block ciphers. Lecture Notes in Computer Science, 1438:80–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1438/14380080.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1438/14380080.pdf.

Blumenthal:1998:LCS

[2678] Uri Blumenthal, N. C. Hien, and J. H. Rooney. Low-cost secure server connection with limited-privilege clients. Lecture Notes in Computer Science, 1438: 90–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380090.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380090.
                                        pdf.

He:1998:SOS

[2679] Qi He, Katia Sycara, and Zhongmin Su. A solution to open standard of PKI. Lecture Notes in Computer Science, 1438:99–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380099.
pdf.

Franz:1998:CCS

[2680] E. Franz, A. Graubner, A. Jerichow, and A. Pfitzmann. Comparison of commitment schemes used in mix-mediated anonymous communication for preventing pool-mode attacks. Lecture Notes in Computer Science, 1438:111– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380111.
pdf.

Golic:1998:CAC

[2681] Jovan Dj. Golic and Renato Menicocci. Correlation attacks on up/ down cascades. Lecture Notes in Computer Science, 1438:123–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380123.
pdf.

Rose:1998:SCB

[2682] Greg Rose. A stream cipher based on linear feedback over GF (2^{8}). Lecture Notes in Computer Science, 1438:135– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380135.
                                        pdf.

Simpson:1998:PCA

[2683] L. Simpson, J. Dj. Golic, and E. Dawson. A probabilistic correlation attack on the shrinking generator. Lecture Notes in Computer Science, 1438:147– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380147.
                                        pdf.

Safavi-Naini:1998:BCC

[2684] Rei Safavi-Naini and Yejing Wang. Bounds and constructions for A^{3}-code with multi-senders. Lecture Notes in Computer Science, 1438:159–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380159.
                                        pdf.

Pieprzyk:1998:RSF

[2685] Josef Pieprzyk and Cheng Xin Qu. Rotation-symmetric functions and fast hashing. Lecture Notes in Computer Science, 1438:169–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380169.
pdf.

Millan:1998:HIN

[2686] William Millan. How to improve the nonlinearity of bijective S-boxes. Lecture Notes in Computer Science, 1438: 181–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380181.
pdf.

Kaksonen:1998:OMC

[2687] Rauli Kaksonen and Petri Mähönen.

Object modeling of cryptographic algorithms with UML. Lecture Notes in Computer Science, 1438:193–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380193.
pdf.

Manninger:1998:AEP

[2688] Martin Manninger and Robert Schischka. Adapting an electronic purse for Internet payments. Lecture Notes in Computer Science, 1438:205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380205.
pdf.

Hanaoka:1998:LLW

[2689] Goichiro Hanaoka, Yuliang Zheng, and Hideki Imai. LITESET: A light-weight secure electronic transaction protocol. Lecture Notes in Computer Science, 1438:215–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380215.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380215.
pdf.

Henderson:1998:ALS

[2690] Marie Henderson. Applications of linearised and sub-linearised polynomials to information security. Lecture Notes in Computer Science, 1438:227– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380227.
pdf.

Chen:1998:PFR

[2691] Mark Chen and Eric Hughes. Protocol failures related to order of encryption and signature: Computation of discrete logarithms in RSA groups. Lecture Notes in Computer Science, 1438: 238–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380238.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380238.
pdf.

Fung:1998:PAE

[2692] W. W. Fung and J. W. Gray, III.

Protection against EEPROM modification attacks. Lecture Notes in Computer Science, 1438:250–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380250.
pdf.

Hruby:1998:TQC

[2693] Jaroslav Hruby. Trends in quantum cryptography in Czech Republic. Lecture Notes in Computer Science, 1438: 261–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380261.
pdf.

Bai:1998:HLL

[2694] Yun Bai and Vijay Varadharajan. A high level language for conventional access control models. Lecture Notes in Computer Science, 1438:273– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380273.
                                        pdf.

Aura:1998:FAC

[2695] Tuomas Aura. Fast access control decisions from delegation certificate databases. Lecture Notes in Computer Science, 1438:284–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380284.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380284.
                                        pdf.

Riechmann:1998:MOA

[2696] Thomas Riechmann and Jürgen Kleinöder. Meta objects for access control: Role-based principals. Lecture Notes in Computer Science, 1438:296– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380296.
                                        pdf.

Tidswell:1998:DTA

[2697] Jonathon E. Tidswell and John M.

Potter. A dynamically typed access control model. Lecture Notes in Computer Science, 1438:308–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380308.
pdf.

Saeednia:1998:EIB

[2698] Shahrokh Saeednia and Rei Safavi-Naini. Efficient identity-based conference key distribution protocols. Lecture Notes in Computer Science, 1438: 320–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380320.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380320.
pdf.

Rudolph:1998:FMS

[2699] Carsten Rudolph. A formal model for systematic design of key establishment protocols. Lecture Notes in Computer Science, 1438:332–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380332.
pdf.

Boyd:1998:KEP

[2700] Colin Boyd and Anish Mathuria. Key establishment protocols for secure mobile communications: A selective survey. Lecture Notes in Computer Science, 1438:344–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380344.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380344.
                                        pdf.

Stabell-Kulo:1998:DKD

[2701] Tage Stabell-Kulø, Arne Helme, and Gianluca Dini. Detecting key-dependencies. Lecture Notes in Computer Science, 1438:356–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380356.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380356.
                                        pdf.

Stabell-Kuloe:1998:DKD

[2702] T. Stabell-Kuloe, A. Helme, and G. Dini. Detecting key-dependencies. Lecture Notes in Computer Science, 1438:356–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghodosi:1998:SSM

[2703] Hossein Ghodosi, Josef Pieprzyk, and Rei Safavi-Naini. Secret sharing in multilevel and compartmented groups. Lecture Notes in Computer Science, 1438:

367–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380367.
pdf.

Ghodosi:1998:CCS

[2704] Hossein Ghodosi, Josef Pieprzyk, Rei Safavi-Naini, and Huaxiong Wang. On construction of cumulative secret sharing schemes. Lecture Notes in Computer Science, 1438:379–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380379.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380379.
pdf.

Desmedt:1998:CES

[2705] Yvo Desmedt, Brian King, Wataru Kishimoto, and Kaoru Kurosawa. A comment on the efficiency of secret sharing scheme over any finite Abelian group. Lecture Notes in Computer Science, 1438:391–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1438/14380391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1438/14380391.
pdf.

Syukri:1998:UIS

[2706] Agus Fanar Syukri, Eiji Okamoto, and Masahiro Mambo. A user identification system using signature written with mouse. Lecture Notes in Computer Science, 1438:403–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380403.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380403.
                                        pdf.

Lee:1998:ZNP

[2707] Narn-Yih Lee, Tzonelih Hwang, and Chih-Hung Wang. On Zhang’s nonrepudiable proxy signature schemes. Lecture Notes in Computer Science, 1438: 415–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1438/14380415.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1438/14380415.
                                        pdf.

Larsen:1998:ICP

[2708] Jens-Otto Larsen and Helge M. Roald. Introducing ClearCase as a process improvement experiment. Lecture Notes in Computer Science, 1439:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1439/14390001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1439/14390001.
                                        pdf.

Rahikkala:1998:IES

[2709] Tua Rahikkala, Jorma Taramaa, and Antti Välimäki. Industrial experiences from SCM current state analysis. Lecture Notes in Computer Science, 1439: 13–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1439/14390013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1439/14390013.
pdf.

Crnkovic:1998:CMS

[2710] Ivica Crnkovic and Per Willför. Change measurements in an SCM process. Lecture Notes in Computer Science, 1439: 26–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1439/14390026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1439/14390026.
pdf.

MacDonald:1998:PPR

[2711] Josh MacDonald, Paul N. Hilfinger, and Luigi Semenzato. PRCS: The project revision control system. Lecture Notes in Computer Science, 1439: 33–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1439/14390033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1439/14390033.
pdf.

Abu-Shakra:1998:MGV

[2712] Makram Abu-Shakra and Gene L.

Fisher. Multi-grain version control in the historian system. Lecture Notes in Computer Science, 1439:46– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1439/14390046.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1439/14390046.
                                        pdf.

Wingerd:1998:HLB

[2713] Laura Wingerd and Christopher Seiwald. High-level best practices in software configuration management. Lecture Notes in Computer Science, 1439: 57–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1439/14390057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1439/14390057.
                                        pdf.

Christensen:1998:EAS

[2714] Henrik Bærbak Christensen. Experiences with architectural software configuration management in ragnarok. Lecture Notes in Computer Science, 1439:67–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1439/14390067.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1439/14390067.
                                        pdf.

Estublier:1998:TSP

[2715] Jacky Estublier, Jean-Marie Favre, and Philippe Morat. Toward SCM / PDM integration? Lecture Notes in Computer Science, 1439:75–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1439/14390075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1439/14390075.
pdf.

Westfechtel:1998:SCM

[2716] Bernhard Westfechtel and Reidar Conradi. Software configuration management and engineering data management: Differences and similarities. Lecture Notes in Computer Science, 1439: 95–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1439/14390095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1439/14390095.
pdf.

Hedin:1998:PCU

[2717] Görel Hedin, Lennart Ohlsson, and John McKenna. Product configuration using object oriented grammars. Lecture Notes in Computer Science, 1439: 107–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1439/14390107.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1439/14390107.
pdf.

Zeller:1998:VSM

[2718] Andreas Zeller. Versioning system models through description logic. Lecture Notes in Computer Science, 1439: 127–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1439/14390127.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1439/14390127.
                                        pdf.

Lindsay:1998:SFG

[2719] Peter Lindsay and Owen Traynor. Supporting fine-grained traceability in software development environments. Lecture Notes in Computer Science, 1439: 133–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1439/14390133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1439/14390133.
                                        pdf.

Hoek:1998:SMR

[2720] André van der Hoek, Dennis Heimbigner, and Alexander L. Wolf. System modeling resurrected. Lecture Notes in Computer Science, 1439:140– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1439/14390140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1439/14390140.
                                        pdf.

VanderHoek:1998:SMR

[2721] A. Van der Hoek, D. Heimbigner, and A. L. Wolf. System modeling resurrected. Lecture Notes in Computer Science, 1439:140–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkins:1998:VSE

[2722] David L. Atkins. Version sensitive editing: Change history as a programming tool. Lecture Notes in Computer Science, 1439:146–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1439/14390146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1439/14390146.
pdf.

VandeVanter:1998:CEV

[2723] Michael L. Van De Vanter. Coordinated editing of versioned packages in the JP programming environment. Lecture Notes in Computer Science, 1439: 158–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1439/14390158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1439/14390158.
pdf.

Bendix:1998:CTV

[2724] Lars Bendix, Per Nygaard Larsen, Anders Ingemann Nielsen, and Jesper Lai Søndergaard Petersen. CoEd — A tool for versioning of hierarchical documents. Lecture Notes in Computer Science, 1439:174–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1439/14390174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1439/14390174.
pdf.

Bielikova:1998:MVH

[2725] Mária Bieliková and Pavol Návrat. Modelling versioned hypertext documents. Lecture Notes in Computer Science, 1439:188–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1439/14390188.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1439/14390188.pdf.

Hall:1998:RSD

[2726] Richard S. Hall, Dennis Heimbigner, and Alexander L. Wolf. Requirements for software deployment languages and schema. Lecture Notes in Computer Science, 1439:198–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1439/14390198.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1439/14390198.
                                        pdf.

Dart:1998:AEC

[2727] Susan Dart. The agony and ecstasy of configuration management. Lecture Notes in Computer Science, 1439:204– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1439/14390204.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1439/14390204.
                                        pdf.

Gersho:1998:C

[2728] A. Gersho. Crypto ’81. Lecture Notes in Computer Science, 1440:3–8, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beth:1998:Ea

[2729] T. Beth. Eurocrypt ’82. Lecture Notes in Computer Science, 1440:9–12, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaum:1998:Ca

[2730] D. Chaum, R. L. Rivest, and A. T.

Sherman. Crypto ’82. Lecture Notes in Computer Science, 1440:13–20, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:Ea

[2731] Anonymous. Eurocrypt ’83. Lecture Notes in Computer Science, 1440: 21, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1440/14400021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1440/14400021.
pdf.

Chaum:1998:Cb

[2732] D. Chaum. Crypto ’83. Lecture Notes in Computer Science, 1440:23– 28, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beth:1998:Eb

[2733] T. Beth, N. Cot, and I. Ingemarsson. Eurocrypt ’84. Lecture Notes in Computer Science, 1440:29–34, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blakley:1998:C

[2734] G. R. Blakley and D. Chaum. Crypto ’84. Lecture Notes in Computer Science, 1440:35–40, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pichler:1998:E

[2735] F. Pichler. EUROCRYPT ’85. Lecture Notes in Computer Science, 1440: 41–48, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1998:C

[2736] H. C. Williams. Crypto ’85. Lecture Notes in Computer Science, 1440:49– 54, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ingemarsson:1998:E

[2737] I. Ingemarsson. Eurocrypt ’86. Lecture Notes in Computer Science, 1440: 55–60, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Odlyzko:1998:C

[2738] A. M. Odlyzko. Crypto ’86. Lecture Notes in Computer Science, 1440:61– 68, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaum:1998:E

[2739] D. Chaum and W. L. Price. Eurocrypt ’87. Lecture Notes in Computer Science, 1440:69–74, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pomerance:1998:C

[2740] C. Pomerance. Crypto ’87. Lecture Notes in Computer Science, 1440:75– 80, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guenther:1998:E

[2741] C. G. Guenther. Eurocrypt ’88. Lecture Notes in Computer Science, 1440: 81–86, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldwasser:1998:C

[2742] S. Goldwasser. Crypto ’88. Lecture Notes in Computer Science, 1440:87– 92, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quisquater:1998:E

[2743] J.-J. Quisquater and J. Vandewalle.

Eurocrypt ’89. Lecture Notes in Computer Science, 1440:93–100, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brassard:1998:C

[2744] G. Brassard. Crypto ’89. Lecture Notes in Computer Science, 1440:101–110, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damgard:1998:E

[2745] I. B. Damgard. Eurocrypt ’90. Lecture Notes in Computer Science, 1440:111– 118, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menezes:1998:C

[2746] A. J. Menezes and S. A. Vanstone. Crypto ’90. Lecture Notes in Computer Science, 1440:119–126, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davies:1998:E

[2747] D. W. Davies. Eurocrypt ’91. Lecture Notes in Computer Science, 1440:127– 134, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feigenbaum:1998:C

[2748] J. Feigenbaum. Crypto ’91. Lecture Notes in Computer Science, 1440:135– 140, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rueppel:1998:E

[2749] R. A. Rueppel. Eurocrypt ’92. Lecture Notes in Computer Science, 1440:141– 146, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brickell:1998:C

[2750] E. F. Brickell. Crypto ’92. Lecture Notes in Computer Science, 1440:147– 152, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helleseth:1998:E

[2751] T. Helleseth. Eurocrypt ’93. Lecture Notes in Computer Science, 1440:153– 158, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stinson:1998:C

[2752] D. R. Stinson. Crypto ’93. Lecture Notes in Computer Science, 1440:159– 164, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSantis:1998:E

[2753] A. De Santis. Eurocrypt ’94. Lecture Notes in Computer Science, 1440:165– 172, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desmedt:1998:C

[2754] Y. Desmedt. Crypto ’94. Lecture Notes in Computer Science, 1440:173–180, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guillou:1998:E

[2755] L. C. Guillou and J.-J. Quisquater. Eurocrypt ’95. Lecture Notes in Computer Science, 1440:181–190, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coppersmith:1998:C

[2756] D. Coppersmith. Crypto ’95. Lecture Notes in Computer Science, 1440:191– 198, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1998:E

[2757] U. Maurer. Eurocrypt ’96. Lecture Notes in Computer Science, 1440:199– 206, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koblitz:1998:C

[2758] N. Koblitz. Crypto ’96. Lecture Notes in Computer Science, 1440:207–214, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fumy:1998:E

[2759] W. Fumy. Eurocrypt ’97. Lecture Notes in Computer Science, 1440:215–222, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaliski:1998:C

[2760] B. Kaliski. Crypto ’97. Lecture Notes in Computer Science, 1440:223–232, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1998:LPC

[2761] Anonymous. List of program committees of the CRYPTO and EURO-CRYPT conferences 1981–1997. Lecture Notes in Computer Science, 1440: 233–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1440/1440clst.pdf.

Anonymous:1998:AIm

[2762] Anonymous. Author index of the CRYPTO and EUROCRYPT conferences 1981–1997. Lecture Notes in Computer Science, 1440:238–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1440/1440auth.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1440/1440auth.
pdf.

Anonymous:1998:KIC

[2763] Anonymous. Keyword index of the CRYPTO and EUROCRYPT conferences 1981–1997. Lecture Notes in Computer Science, 1440:447–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1440/1440keyw.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1440/1440keyw.
pdf.

Anonymous:1998:BFC

[2764] Anonymous. BiBTEX file of the CRYPTO and EUROCRYPT conferences 1981–1997. Lecture Notes in Computer Science, 1440:??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/supp/1440/1440btex.
bib.

Shapiro:1998:SCM

[2765] Steven Shapiro, Yves Lespérance, and Hector J. Levesque. Specifying communicative multi-agent systems (invited paper). Lecture Notes in Computer Science, 1441:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1441/14410001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1441/14410001.
pdf.

Wooldridge:1998:MIT

[2766] Michael Wooldridge and Afsaneh Haddadi. Making it up as they go along: A theory of reactive cooperation (invited paper). Lecture Notes in Computer Science, 1441:15– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1441/14410015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1441/14410015.
pdf.

Wobcke:1998:ALA

[2767] Wayne Wobcke. Agency and the logic of ability. Lecture Notes in Computer Science, 1441:31–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1441/14410031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1441/14410031.
pdf.

Lomuscio:1998:RBI

[2768] Alessio Lomuscio and Mark Ryan. On the relation between interpreted systems and Kripke models. Lecture Notes in Computer Science, 1441:46– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1441/14410046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1441/14410046.
pdf.

Yuan:1998:DLPa

[2769] Li-Yan Yuan, Jia-Huai You, and Randy Goebel. Disjunctive logic programming and possible model semantics (invited paper). Lecture Notes in Computer Science, 1441:60– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1441/14410060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1441/14410060.
pdf.

Nakamatsu:1998:NMA

[2770] Kazumi Nakamatsu and Atsuyuki Suzuki. A non-monotonic ATMS based on annotated logic programs with strong negation. Lecture Notes in Computer Science, 1441:79–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1441/14410079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1441/14410079.
pdf.

Gibbon:1998:SBR

[2771] Greg Gibbon and Janet Aisbett.

Switching between reasoning and search. Lecture Notes in Computer Science, 1441:94–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1441/14410094.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1441/14410094.pdf.

Moulin:1998:SDI

[2772] Bernard Moulin. The social dimension of interactions in multiagent systems. Lecture Notes in Computer Science, 1441:109–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1441/14410109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1441/14410109.
                                        pdf.

Norman:1998:GSJ

[2773] Timothy J. Norman and Nicholas R. Jennings. Generating states of joint commitment between autonomous agents. Lecture Notes in Computer Science, 1441:123–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1441/14410123.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1441/14410123.pdf.

Ossowski:1998:SCO

[2774] Sascha Ossowski and Ana García-Serrano. Social co-ordination among autonomous problem-solving agents. Lecture Notes in Computer Science, 1441:134–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1441/14410134.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1441/14410134.
                                        pdf.

Zhang:1998:APV

[2775] Chengqi Zhang and Yuefeng Li. An algorithm for plan verification in multiple agent systems. Lecture Notes in Computer Science, 1441:149–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1441/14410149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1441/14410149.
pdf.

Bui:1998:FCL

[2776] Hung Hai Bui, Svetha Venkatesh, and Dorota Kieronska. A framework for coordination and learning among teams of agents. Lecture Notes in Computer Science, 1441:164– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1441/14410164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1441/14410164.
pdf.

Carlsson:1998:IHD

[2777] Bengt Carlsson and Stefan Johansson.

An iterated hawk-and-dove game. Lecture Notes in Computer Science, 1441: 179–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1441/14410179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1441/14410179.
pdf.

Yoshida:1998:GTS

[2778] Satoru Yoshida, Nobuhiro Inuzuka, Tin Tin Naing, Hirohisa Seki, and Hidenori Itoh. A game-theoretic solution of conflicts among competitive agents. Lecture Notes in Computer Science, 1441:193–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1441/14410193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1441/14410193.
                                        pdf.

Zhang:1998:TBE

[2779] Chengqi Zhang and Xudong Luo. Transformation between the EMYCIN model and the Bayesian network. Lecture Notes in Computer Science, 1441: 205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1441/14410205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1441/14410205.
                                        pdf.

Zhang:1998:UMA

[2780] Dong Mei Zhang, Leila Alem, and Kalina Yacef. Using multi-agent approach for the design of an intelligent learning environment. Lecture Notes in Computer Science, 1441:220– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1441/14410220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1441/14410220.
                                        pdf.

Zhang:1998:CBS

[2781] Minjie Zhang. A case-based strategy for solution synthesis among cooperative expert systems. Lecture Notes in Computer Science, 1441:231– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1441/14410231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1441/14410231.
pdf.

Fiat:1998:CAA

[2782] A. Fiat and G. J. Woeginger. Competitive analysis of algorithms. Lecture Notes in Computer Science, 1442:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1998:SOD

[2783] S. Albers and J. Westbrook. Selforganizing data structures. Lecture Notes in Computer Science, 1442:13– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Irani:1998:CAP

[2784] S. Irani. Competitive analysis of paging. Lecture Notes in Computer Science, 1442:52–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chrobak:1998:MTS

[2785] M. Chrobak and L. L. Larmore. Metrical task systems, the server problem and the work function algorithm. Lecture Notes in Computer Science, 1442: 74–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bartal:1998:DP

[2786] Y. Bartal. Distributed paging. Lecture Notes in Computer Science, 1442: 97–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aspnes:1998:CAD

[2787] J. Aspnes. Competitive analysis of distributed algorithms. Lecture Notes in Computer Science, 1442:118–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Csirik:1998:LPC

[2788] J. Csirik and G. J. Woeginger. On-line packing and covering problems. Lecture Notes in Computer Science, 1442: 147–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Azar:1998:LLB

[2789] Y. Azar. On-line load balancing. Lecture Notes in Computer Science, 1442: 178–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sgall:1998:LS

[2790] J. Sgall. On-line scheduling. Lecture Notes in Computer Science, 1442:196– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1998:LSN

[2791] P. Berman. On-line searching and navigation. Lecture Notes in Computer Science, 1442:232–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leonardi:1998:LNR

[2792] S. Leonardi. On-line network routing. Lecture Notes in Computer Science, 1442:242–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalyanasundaram:1998:LNO

[2793] B. Kalyanasundaram and K. Pruhs.

On-line network optimization problems. Lecture Notes in Computer Science, 1442:268–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kierstead:1998:CGL

[2794] H. A. Kierstead. Coloring graphs online. Lecture Notes in Computer Science, 1442:281–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1998:LAM

[2795] A. Blum. On-line algorithms in machine learning. Lecture Notes in Computer Science, 1442:306–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Yaniv:1998:CSL

[2796] R. El-Yaniv. Competitive solutions for on-line financial problems. Lecture Notes in Computer Science, 1442:326– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karlin:1998:PCA

[2797] A. R. Karlin. On the performance of competitive algorithms in practice. Lecture Notes in Computer Science, 1442:373–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiat:1998:COE

[2798] A. Fiat and G. J. Woeginger. Competitive odds and ends. Lecture Notes in Computer Science, 1442:385–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesten:1998:AVL

[2799] Yonit Kesten, Amir Pnueli, and Li on Raviv. Algorithmic verification of linear temporal logic specifications. Lecture Notes in Computer Science, 1443:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1443/14430001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1443/14430001.
pdf.

Borchert:1998:EFO

[2800] Bernd Borchert, Dietrich Kuske, and Frank Stephan. On existentially first-order definable languages and their relation to NP. Lecture Notes in Computer Science, 1443:17–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430017.
pdf.

Raymond:1998:AAC

[2801] Jean-François Raymond, Pascal Tesson, and Denis Thérien. An algebraic approach to communication complexity. Lecture Notes in Computer Science, 1443:29–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430029.
pdf.

Alur:1998:DGP

[2802] Rajeev Alur, Ken McMillan, and Doron Peled. Deciding global partialorder properties. Lecture Notes in Computer Science, 1443:41–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430041.
pdf.

Liu:1998:SLT

[2803] Xinxin Liu and Scott A. Smolka. Simple linear-time algorithms for minimal fixed points. Lecture Notes in Computer Science, 1443:53–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430053.
pdf.

Husfeldt:1998:HRD

[2804] Thore Husfeldt and Theis Rauhe.

Hardness results for dynamic problems by extensions of Fredman and Saks’ chronogram method. Lecture Notes in Computer Science, 1443:67– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430067.
                                        pdf.

Hagerup:1998:SFD

[2805] Torben Hagerup. Simpler and faster dictionaries on the AC^{0} RAM. Lecture Notes in Computer Science, 1443: 79–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430079.
                                        pdf.

Lasota:1998:PCF

[2806] S. Lasota. Partial-congruence factorization of bisimilarity induced by open maps. Lecture Notes in Computer Science, 1443:91–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430091.
                                        pdf.

Dufourd:1998:RNB

[2807] C. Dufourd, A. Finkel, and Ph. Schnoebelen. Reset nets between decidability and undecidability. Lecture Notes in Computer Science, 1443:103– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430103.
pdf.

Overmars:1998:GAR

[2808] Mark H. Overmars. Geometric algorithms for robotic manipulation. Lecture Notes in Computer Science, 1443: 116–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430116.
pdf.

Chuang:1998:CEP

[2809] Richie Chih-Nan Chuang, Ashim Garg, Xin He, Ming-Yang Kao, and Hsueh-I Lu. Compact encodings of planar graphs via canonical orderings and multiple parentheses. Lecture Notes in Computer Science, 1443:118– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430118.
pdf.

Graf:1998:RSP

[2810] Thorsten Graf and Kamakoti Veezhinathan. Reducing simple polygons to triangles — A proof for an improved conjecture. Lecture Notes in Computer Science, 1443:130–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430130.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430130.
                                        pdf.

Walukiewicz:1998:DCC

[2811] Igor Walukiewicz. Difficult configurations — on the complexity of ltrl. Lecture Notes in Computer Science, 1443: 140–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430140.
                                        pdf.

Boigelot:1998:ERI

[2812] Bernard Boigelot, Stéphane Rassart, and Pierre Wolper. On the expressiveness of real and integer arithmetic automata. Lecture Notes in Computer Science, 1443:152–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430152.
                                        pdf.

Peleg:1998:DMB

[2813] David Peleg. Distributed matroid basis completion via elimination upcast and distributed correction of minimumweight spanning trees. Lecture Notes in Computer Science, 1443:164–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430164.
pdf.

Halldorsson:1998:ISD

[2814] Magnús M. Halldórsson, Jan Kratochvíl, and Jan Arne Telle. Independent sets with domination constraints. Lecture Notes in Computer Science, 1443:176–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1443/14430176.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1443/14430176.
pdf.

Mukund:1998:RAP

[2815] Madhavan Mukund, K. Narayan Kumar, Jaikumar Radhakrishnan, and Milind Sohoni. Robust asynchronous protocols are finite-state. Lecture Notes in Computer Science, 1443:188– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430188.
pdf.

Jancar:1998:DBL

[2816] Petr Jancar, Antonín Kucera, and Richard Mayr. Deciding bisimulationlike equivalences with finite-state processes. Lecture Notes in Computer Science, 1443:200–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1443/14430200.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1443/14430200.pdf.

Wigderson:1998:DPA

[2817] Avi Wigderson. Do probabilistic algorithms outperform deterministic ones? Lecture Notes in Computer Science, 1443:212–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430212.
                                        pdf.

Grolmusz:1998:DDL

[2818] Vince Grolmusz. A degree-decreasing lemma for (MOD q, MOD p ) circuits. Lecture Notes in Computer Science, 1443:215–222, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1443/14430215.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1443/14430215.pdf.

Lu:1998:IPG

[2819] Chi-Jen Lu. Improved pseudorandom generators for combinatorial rectangles. Lecture Notes in Computer Science, 1443:223–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430223.
pdf.

Pnueli:1998:TVS

[2820] A. Pnueli, O. Shtrichman, and M. Siegel. Translation validation for synchronous languages. Lecture Notes in Computer Science, 1443:235– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430235.
pdf.

Zakharov:1998:EUA

[2821] Vladimir A. Zakharov. An efficient and unified approach to the decidability of equivalence of propositional programs. Lecture Notes in Computer Science, 1443:247–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430247.
pdf.

Jukna:1998:BPB

[2822] Stasys Jukna and Stanislav Zák. On branching programs with bounded uncertainty. Lecture Notes in Computer Science, 1443:259–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1443/14430259.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1443/14430259.pdf.

Ben-Amram:1998:CFP

[2823] Amir M. Ben-Amram and Holger Petersen. CONS-free programs with tree input. Lecture Notes in Computer Science, 1443:271–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430271.
                                        pdf.

Baldan:1998:CGP

[2824] Paolo Baldan, Andrea Corradini, and Ugo Montanari. Concatenable graph processes: Relating processes and derivation traces. Lecture Notes in Computer Science, 1443:283–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430283.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430283.
                                        pdf.

Gadducci:1998:ACN

[2825] F. Gadducci and U. Montanari. Axioms for contextual net processes. Lecture Notes in Computer Science, 1443: 296–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1443/14430296.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430296.
pdf.

Pitts:1998:ETL

[2826] Andrew M. Pitts. Existential types: Logical relations and operational equivalence. Lecture Notes in Computer Science, 1443:309–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430309.
pdf.

Martinez:1998:OSS

[2827] Conrado Martínez and Salvador Roura. Optimal sampling strategies in Quicksort. Lecture Notes in Computer Science, 1443:327–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430327.
pdf.

Dyer:1998:GPT

[2828] Martin Dyer and Catherine Greenhill. A genuinely polynomial-time algorithm for sampling two-rowed contingency tables. Lecture Notes in Computer Science, 1443:339–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430339.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430339.
                                        pdf.

Power:1998:MAD

[2829] John Power and Giuseppe Rosolini. A modular approach to denotational semantics. Lecture Notes in Computer Science, 1443:351–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430351.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430351.
                                        pdf.

Malacaria:1998:GFG

[2830] Pasquale Malacaria and Chris Hankin. Generalised flowcharts and games. Lecture Notes in Computer Science, 1443: 363–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430363.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430363.
                                        pdf.

Kao:1998:EMN

[2831] Ming-Yang Kao and Jie Wang. Efficient minimization of numerical summation errors. Lecture Notes in Computer Science, 1443:375–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430375.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1443/14430375.
pdf.

Bazgan:1998:EAA

[2832] Cristina Bazgan, Miklos Santha, and Zsolt Tuza. Efficient approximation algorithms for the Subset-Sums Equality problem. Lecture Notes in Computer Science, 1443:387–396, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430387.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430387.
pdf.

Gimenez:1998:SRD

[2833] Eduardo Giménez. Structural recursive definitions in type theory. Lecture Notes in Computer Science, 1443:397– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430397.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430397.
pdf.

Lugiez:1998:GCT

[2834] D. Lugiez. A good class of tree automata. application to inductive theorem proving. Lecture Notes in Computer Science, 1443:409–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430409.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1443/14430409.
                                        pdf.

Karhumaeki:1998:LPI

[2835] J. Karhumaeki, A. Lepistoe, and W. Plandowski. Locally periodic infinite words and a chaotic behaviour. Lecture Notes in Computer Science, 1443:421–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karhumaki:1998:LPI

[2836] Juhani Karhumäki, Arto Lepistö, and Wojciech Plandowski. Locally periodic infinite words and a chaotic behaviour. Lecture Notes in Computer Science, 1443:421–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430421.
                                        pdf.

Pin:1998:BCH

[2837] Jean-Éric Pin. Bridges for concatenation hierarchies. Lecture Notes in Computer Science, 1443:431–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430431.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430431.
                                        pdf.

Lin:1998:CPS

[2838] H. Lin. Complete proof systems for observation congruences in finitecontrol π-calculus. Lecture Notes in Computer Science, 1443:443–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430443.
pdf.

Victor:1998:CCF

[2839] Björn Victor and Joachim Parrow. Concurrent constraints in the fusion calculus. Lecture Notes in Computer Science, 1443:455–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430455.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430455.
pdf.

Damico:1998:CEC

[2840] Michele D’amico, Giovanni Manzini, and Luciano Margara. On computing the entropy of cellular automata. Lecture Notes in Computer Science, 1443: 470–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430470.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430470.
pdf.

Buchsbaum:1998:DWF

[2841] Adam L. Buchsbaum, Raffaele Giancarlo, and Jeffery R. Westbrook. On the determinization of weighted finite automata. Lecture Notes in Computer Science, 1443:482–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430482.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430482.
                                        pdf.

Tiskin:1998:BSP

[2842] A. Tiskin. Bulk-synchronous parallel multiplication of Boolean matrices. Lecture Notes in Computer Science, 1443:494–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1443/14430494.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430494.

pdf. See erratum [3178].

Touzet:1998:CES

[2843] Hélène Touzet. A complex example of a simplifying rewrite system. Lecture Notes in Computer Science, 1443:507– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430507.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430507.
                                        pdf.

Mellies:1998:DBK

[2844] Paul-André Melliès. On a duality between kruskal and dershowitz theorems. Lecture Notes in Computer Science, 1443:518–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430518.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430518.
pdf.

Walukiewicz:1998:TAC

[2845] Daria Walukiewicz. A total ACcompatible reduction ordering on higher-order terms. Lecture Notes in Computer Science, 1443:530–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430530.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430530.
pdf.

Henzinger:1998:MCG

[2846] Thomas A. Henzinger. Model checking game properties of multiagent systems. Lecture Notes in Computer Science, 1443:543–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430543.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430543.
pdf.

Gargano:1998:LWC

[2847] Luisa Gargano. Limited wavelength conversion in all-optical tree networks. Lecture Notes in Computer Science, 1443:544–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1443/14430544.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430544.
                                        pdf.

Chaudhuri:1998:CMN

[2848] Shiva Chaudhuri, K. V. Subrahmanyam, Frank Wagner, and Christos D. Zaroliagis. Computing mimicking networks. Lecture Notes in Computer Science, 1443:556–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430556.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430556.
                                        pdf.

Baier:1998:MST

[2849] Christel Baier, Joost-Pieter Katoen, and Diego Latella. Metric semantics for true concurrent real time. Lecture Notes in Computer Science, 1443:568– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430568.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430568.
                                        pdf.

Henzinger:1998:RRT

[2850] T. A. Henzinger, J.-F. Raskin, and P.-Y. Schobbens. The regular realtime languages. Lecture Notes in Computer Science, 1443:580–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1443/14430580.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430580.
pdf.

Cicerone:1998:SDL

[2851] Serafino Cicerone, Gabriele Di Stefano, and Michele Flammini. Static and dynamic low-congested interval routing schemes. Lecture Notes in Computer Science, 1443:592–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430592.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430592.
pdf.

Cook:1998:LBR

[2852] Doug Cook, Vance Faber, Madhav Marathe, Aravind Srinivasan, and Yoram J. Sussmann. Lowbandwidth routing and electrical power networks. Lecture Notes in Computer Science, 1443:604–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430604.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430604.
pdf.

Henglein:1998:CAC

[2853] Fritz Henglein and Jakob Rehof. Constraint automata and the complexity of recursive subtype entailment. Lecture Notes in Computer Science, 1443: 616–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430616.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430616.
                                        pdf.

Vardi:1998:RAP

[2854] Moshe Y. Vardi. Reasoning about the past with two-way automata. Lecture Notes in Computer Science, 1443:628– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430628.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430628.
                                        pdf.

Valiant:1998:NAC

[2855] Leslie G. Valiant. A neuroidal architecture for cognitive computation. Lecture Notes in Computer Science, 1443: 642–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430642.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430642.
                                        pdf.

Peleg:1998:DPA

[2856] David Peleg and Eilon Reshef. Deterministic polylog approximation for minimum communication spanning trees. Lecture Notes in Computer Science, 1443:670–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1443/14430670.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430670.
pdf.

Czumaj:1998:PTA

[2857] Artur Czumaj and Andrzej Lingas. A polynomial time approximation scheme for Euclidean minimum cost k-connectivity. Lecture Notes in Computer Science, 1443:682– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430682.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430682.
pdf.

Sewell:1998:GLS

[2858] Peter Sewell. Global/local subtyping and capability inference for a distributed π-calculus. Lecture Notes in Computer Science, 1443:695–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430695.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430695.
pdf.

Li:1998:CSW

[2859] Zhoujun Li and Huowang Chen. Checking strong/weak bisimulation equivalences and observation congruence for the π-calculus. Lecture Notes in Computer Science, 1443:707–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430707.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430707.
                                        pdf.

Bini:1998:ICM

[2860] Dario Bini, Gianna M. Del Corso, Giovanni Manzini, and Luciano Margara. Inversion of circulant matrices over Z_{m}. Lecture Notes in Computer Science, 1443:719–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430719.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430719.
                                        pdf.

Plandowski:1998:ALZ

[2861] Wojciech Plandowski and Wojciech Rytter. Application of Lempel– Ziv encodings to the solution of word equations. Lecture Notes in Computer Science, 1443:731–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430731.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430731.
                                        pdf.

Ghani:1998:ESC

[2862] Neil Ghani, Valeria de Paiva, and Eike Ritter. Explicit substitutions for constructive necessity. Lecture Notes in Computer Science, 1443:743– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430743.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430743.
pdf.

Barthe:1998:RPI

[2863] Gilles Barthe. The relevance of proof-irrelevance. Lecture Notes in Computer Science, 1443:755–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430755.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430755.
pdf.

Brassard:1998:NHQ

[2864] Gilles Brassard. New horizons in quantum information processing. Lecture Notes in Computer Science, 1443:769– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430769.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430769.
pdf.

Damgaard:1998:SII

[2865] Ivan Damgård and Birgit Pfitzmann.

Sequential iteration of interactive arguments and an efficient zero-knowledge argument for NP. Lecture Notes in Computer Science, 1443:772–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430772.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430772.
                                        pdf.

DeSantis:1998:IDC

[2866] A. De Santis, G. Di Crescenzo, G. Persiano, and M. Yung. Image density is complete for Non-Interactive-SZK. Lecture Notes in Computer Science, 1443:784–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santis:1998:IDC

[2867] A. De Santis, G. Di Crescenzo, G. Persiano, and M. Yung. Image density is complete for non-interactive-SZK. Lecture Notes in Computer Science, 1443: 784–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430784.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430784.
                                        pdf.

Hertling:1998:RS

[2868] Peter Hertling and Klaus Weihrauch. Randomness spaces. Lecture Notes in Computer Science, 1443:796–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430796.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430796.
                                        pdf.

Bucciarelli:1998:TDB

[2869] Antonio Bucciarelli and Ivano Salvo.

Totality, definability and Boolean circuits. Lecture Notes in Computer Science, 1443:808–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1443/14430808.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1443/14430808.pdf.

Brassard:1998:QC

[2870] Gilles Brassard, Peter Høyer, and Alain Tapp. Quantum counting. Lecture Notes in Computer Science, 1443:820– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430820.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430820.
pdf.

Akutsu:1998:CDS

[2871] Tatsuya Akutsu and Mutsunori Yagiura. On the complexity of deriving score functions from examples for problems in molecular biology. Lecture Notes in Computer Science, 1443:832– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430832.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430832.
pdf.

Fournet:1998:HEA

[2872] Cédric Fournet and Georges Gonthier. A hierarchy of equivalences for asynchronous calculi. Lecture Notes in Computer Science, 1443:844–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430844.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430844.
                                        pdf.

Merro:1998:ANP

[2873] Massimo Merro and Davide Sangiorgi. On asynchrony in namepassing calculi. Lecture Notes in Computer Science, 1443:856–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430856.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430856.
                                        pdf.

Abadi:1998:PPLa

[2874] Martin Abadi. Protection in programminglanguage translations. Lecture Notes in Computer Science, 1443:868–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1443/14430868.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1443/14430868.
                                        pdf.

Petersen:1998:ESQ

[2875] Holger Petersen and John Michael Robson. Efficient simulations by queue machines. Lecture Notes in Computer Science, 1443:884–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430884.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430884.
pdf.

Duris:1998:PCM

[2876] Pavol Duris, Tomasz Jurdzinski, Mirosław Kutyłowski, and Krzysztof Lorys. Power of cooperation and multihead finite systems. Lecture Notes in Computer Science, 1443:896– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430896.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430896.
pdf.

Danvy:1998:SST

[2877] Olivier Danvy. A simple solution to type specialization. Lecture Notes in Computer Science, 1443:908–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430908.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430908.
pdf.

Taha:1998:MSP

[2878] Walid Taha, Zine-El-Abidine Benaissa, and Tim Sheard. Multistage programming: Axiomatization and type safety. Lecture Notes in Computer Science, 1443:918–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1443/14430918.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1443/14430918.
pdf.

Halldorsson:1998:AIS

[2879] Magnús M. Halldórsson. Approximations of independent sets in graphs. Lecture Notes in Computer Science, 1444:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1444/14440001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1444/14440001.
pdf.

Shmoys:1998:ULP

[2880] David B. Shmoys. Using linear programming in the design and analysis of approximation algorithms: Two illustrative problems. Lecture Notes in Computer Science, 1444:15–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440015.
pdf.

Vazirani:1998:STP

[2881] Vijay V. Vazirani. The Steiner tree problem and its generalizations. Lecture Notes in Computer Science, 1444: 33–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440033.
pdf.

Azar:1998:ASC

[2882] Yossi Azar and Leah Epstein. Approximation schemes for covering and scheduling on related machines. Lecture Notes in Computer Science, 1444: 39–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440039.
pdf.

Bar-Yehuda:1998:OPT

[2883] Reuven Bar-Yehuda. One for the price of two: A unified approach for approximating covering problems. Lecture Notes in Computer Science, 1444: 49–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440049.
pdf.

Baur:1998:AGD

[2884] Christoph Baur and Sándor P. Fekete. Approximation of geometric dispersion problems (extended abstract). Lecture Notes in Computer Science, 1444:63– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440063.
pdf.

Cheriyan:1998:AOS

[2885] Joseph Cheriyan, Tibor Jordán, and Zeev Nutov. Approximating koutconnected subgraph problems. Lecture Notes in Computer Science, 1444: 77–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440077.
pdf.

Epstein:1998:LBL

[2886] Leah Epstein. Lower bounds for online scheduling with precedence constraints on identical machines. Lecture Notes in Computer Science, 1444:89– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440089.
pdf.

Hochbaum:1998:IRH

[2887] Dorit S. Hochbaum. Instant recognition of half integrality and 2approximations. Lecture Notes in Computer Science, 1444:99–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440099.
pdf.

Hochbaum:1998:VCP

[2888] Dorit S. Hochbaum. The tvertex cover problem: Extending the half integrality framework with budget constraints. Lecture Notes in Computer Science, 1444:111–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440111.
pdf.

Kellerer:1998:NFP

[2889] Hans Kellerer and Ulrich Pferschy. A new fully polynomial approximation scheme for the knapsack problem. Lecture Notes in Computer Science, 1444: 123–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440123.
pdf.

Kortsarz:1998:HAS

[2890] Guy Kortsarz. On the hardness of approximating spanners. Lecture Notes in Computer Science, 1444:135– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1444/14440135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1444/14440135.
                                        pdf.

Kumar:1998:ACA

[2891] Vijay Kumar. Approximating circular arc colouring and bandwidth allocation in all-optical ring networks. Lecture Notes in Computer Science, 1444:147– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1444/14440147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1444/14440147.
                                        pdf.

Schiermeyer:1998:AMI

[2892] Ingo Schiermeyer. Approximating maximum independent set in kclique-free graphs. Lecture Notes in Computer Science, 1444:159–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1444/14440159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1444/14440159.
                                        pdf.

Spieksma:1998:AIS

[2893] Frits C. R. Spieksma. Approximating an interval scheduling problem. Lecture Notes in Computer Science, 1444: 169–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440169.
pdf.

Srivastav:1998:FDS

[2894] Anand Srivastav and Katja Wolf. Finding dense subgraphs with semidefinite programming. Lecture Notes in Computer Science, 1444:181–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440181.
pdf.

Sviridenko:1998:BPA

[2895] Maxim I. Sviridenko. Best possible approximation algorithm for MAX SAT with cardinality constraint. Lecture Notes in Computer Science, 1444:193– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1444/14440193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1444/14440193.
pdf.

Anonymous:1998:AIn

[2896] Anonymous. Author index. Lecture Notes in Computer Science, 1444:201– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1444/1444auth.pdf.

Lange:1998:MOM

[2897] Danny B. Lange. Mobile objects and mobile agents: The future of distributed computing? Lecture Notes in Computer Science, 1445:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450001.
pdf.

Christensen:1998:MDE

[2898] Michael Christensen, Andy Crabtree, Christian Heide Damm, Klaus Marius Hansen, Ole Lehrmann Madsen, Pernille Marqvardsen, Preben Mogensen, Elmer Sandvad, Lennert Sloth, and Michael Thomsen. The M.A.D. experience: Multiperspective application development in evolutionary prototyping. Lecture Notes in Computer Science, 1445:13–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450013.
pdf.

Bertino:1998:EOO

[2899] Elisa Bertino, Elena Ferrari, Giovanna Guerrini, and Isabella Merlo. Extending the ODMG object model with time. Lecture Notes in Computer Science, 1445:41–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1445/14450041.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1445/14450041.
pdf.

Fritzson:1998:MUO

[2900] Peter Fritzson and Vadim Engelson. Modelica — A unified objectoriented language for system modeling and simulation. Lecture Notes in Computer Science, 1445:67–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450067.
pdf.

Krishnamurthi:1998:SOO

[2901] Shriram Krishnamurthi, Matthias Felleisen, and Daniel P. Friedman. Synthesizing object-oriented and functional design to promote re-use. Lecture Notes in Computer Science, 1445: 91–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450091.
pdf.

Lauder:1998:PVS

[2902] Anthony Lauder and Stuart Kent. Precise visual specification of design patterns. Lecture Notes in Computer Science, 1445:114–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1445/14450114.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1445/14450114.pdf.

Sane:1998:FDE

[2903] Aamod Sane, Ashish Singhai, and Roy H. Campbell. Framework design for end-to-end optimization. Lecture Notes in Computer Science, 1445:135– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1445/14450135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1445/14450135.
                                        pdf.

Noble:1998:FAP

[2904] James Noble, Jan Vitek, and John Potter. Flexible alias protection. Lecture Notes in Computer Science, 1445:158– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1445/14450158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1445/14450158.
                                        pdf.

Ernst:1998:PDU

[2905] Michael Ernst, Craig Kaplan, and Craig Chambers. Predicate dispatching: A unified theory of dispatch. Lecture Notes in Computer Science, 1445: 186–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450186.
pdf.

Alagic:1998:OJI

[2906] Suad Alagic, Jose Solorzano, and David Gitchell. Orthogonal to the Java imperative. Lecture Notes in Computer Science, 1445:212–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450212.
pdf.

Ferreira:1998:MDC

[2907] Paulo Ferreira and Marc Shapiro. Modelling a distributed cached store for garbage collection: The algorithm and its correctness proof. Lecture Notes in Computer Science, 1445:234– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450234.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450234.
pdf.

Rodrigues:1998:CDG

[2908] Helena Rodrigues and Richard Jones.

Cyclic distributed garbage collection with group merger. Lecture Notes in Computer Science, 1445:260–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1445/14450260.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1445/14450260.
                                        pdf.

Cordsen:1998:EDV

[2909] J. Cordsen, J. Nolte, and W. Schröder-Preikschat. Experiences developing a virtual shared memory system using high-level object paradigms. Lecture Notes in Computer Science, 1445:285– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1445/14450285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1445/14450285.
                                        pdf.

Keller:1998:BCA

[2910] Ralph Keller and Urs Hölzle. Binary component adaptation. Lecture Notes in Computer Science, 1445:307– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1445/14450307.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1445/14450307.
                                        pdf.

Vijaykrishnan:1998:OOA

[2911] N. Vijaykrishnan, N. Ranganathan, and R. Gadekarla. Object-oriented architectural support for a Java processor. Lecture Notes in Computer Science, 1445:330–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1445/14450330.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1445/14450330.pdf.

Mikhajlov:1998:SFB

[2912] Leonid Mikhajlov and Emil Sekerinski. A study of the fragile base class problem. Lecture Notes in Computer Science, 1445:355–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450355.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450355.
pdf.

Atkinson:1998:POP

[2913] Malcolm Atkinson and Mick Jordan. Providing orthogonal persistence for Java. Lecture Notes in Computer Science, 1445:383–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450383.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450383.
pdf.

Brant:1998:WR

[2914] John Brant, Brian Foote, Ralph E.

Johnson, and Donald Roberts. Wrappers to the rescue. Lecture Notes in Computer Science, 1445:396–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1445/14450396.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1445/14450396.
                                        pdf.

Masuhara:1998:DPE

[2915] Hidehiko Masuhara and Akinori Yonezawa. Design and partial evaluation of meta-objects for a concurrent reflective language. Lecture Notes in Computer Science, 1445:418– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1445/14450418.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1445/14450418.
                                        pdf.

DeOliveiraGuimaraes:1998:RST

[2916] J. De Oliveira Guimaraes. Reflection for statically typed languages. Lecture Notes in Computer Science, 1445:440– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guimaraes:1998:RST

[2917] José de Oliveira Guimarães. Reflection for statically typed languages. Lecture Notes in Computer Science, 1445:440– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1445/14450440.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1445/14450440.
pdf.

Bono:1998:IFO

[2918] Viviana Bono and Kathleen Fisher.

An imperative, first-order calculus with object extension. Lecture Notes in Computer Science, 1445:462–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450462.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450462.
pdf.

Liquori:1998:OE

[2919] Luigi Liquori. On object extension.

Lecture Notes in Computer Science, 1445:498–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1445/14450498.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1445/14450498.
pdf.

Bruce:1998:SSA

[2920] Kim B. Bruce, Martin Odersky, and Philip Wadler. A statically safe alternative to virtual types. Lecture Notes in Computer Science, 1445:523– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450523.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450523.
pdf.

Smaragdakis:1998:ILD

[2921] Yannis Smaragdakis and Don Batory. Implementing layered designs with mixin layers. Lecture Notes in Computer Science, 1445:550–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450550.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450550.
pdf.

Crnogorac:1998:CIM

[2922] Lobel Crnogorac, Anand S. Rao, and Kotagiri Ramamohanarao. Classifying inheritance mechanisms in concurrent object oriented programming. Lecture Notes in Computer Science, 1445:571– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450571.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450571.
pdf.

Gil:1998:CTA

[2923] Joseph (Yossi) Gil and Alon Itai. The complexity of type analysis of object oriented programs. Lecture Notes in Computer Science, 1445:601– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1445/14450601.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1445/14450601.
pdf.

Anonymous:1998:AIo

[2924] Anonymous. Author index. Lecture Notes in Computer Science, 1445:635– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1445/1445ind.pdf.

DeRaedt:1998:AVL

[2925] L. De Raedt. Attribute-value learning versus inductive logic programming: The missing links. Lecture Notes in Computer Science, 1446:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muggleton:1998:AIT

[2926] S. Muggleton. Advances in ILP theory and implementations. Lecture Notes in Computer Science, 1446:9–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivasan:1998:AIP

[2927] A. Srinivasan. Application of ILP to problems in chemistry and biology. Lecture Notes in Computer Science, 1446:10–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzeroski:1998:RRL

[2928] S. Dzeroski, L. De Raedt, and H. Blockeel. Relational reinforcement learning. Lecture Notes in Computer Science, 1446:11–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reddy:1998:LFO

[2929] C. Reddy and P. Tadepalli. Learning first-order acyclic Horn programs from entailment. Lecture Notes in Computer Science, 1446:23–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slattery:1998:CSR

[2930] S. Slattery and M. Craven. Combining statistical and relational methods for learning in hypertext domains. Lecture Notes in Computer Science, 1446: 38–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turcotte:1998:AIL

[2931] M. Turcotte, S. H. Muggleton, and M. J. E. Sternberg. Application of inductive logic programming to discover rules governing the three-dimensional topology of protein structure. Lecture Notes in Computer Science, 1446:53– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bohnebeck:1998:TCF

[2932] U. Bohnebeck, T. Horvath, and S. Wrobel. Term comparisons in first-order similarity measures. Lecture Notes in Computer Science, 1446:65–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1998:SPN

[2933] S. Kramer, B. Pfahringer, and C. Helma. Stochastic propositionalization of non-determinate background knowledge. Lecture Notes in Computer Science, 1446:80–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebag:1998:SSS

[2934] M. Sebag. A stochastic simple similarity. Lecture Notes in Computer Science, 1446:95–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cussens:1998:UPP

[2935] J. Cussens. Using prior probabilities and density estimation for relational classification. Lecture Notes in Computer Science, 1446:106–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eineborg:1998:ICG

[2936] M. Eineborg and N. Lindberg. Induction of constraint grammar-rules using progol. Lecture Notes in Computer Science, 1446:116–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kazakov:1998:HAW

[2937] D. Kazakov and S. Manandhar. A hybrid approach to word segmentation. Lecture Notes in Computer Science, 1446:125–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manandhar:1998:LMM

[2938] S. Manandhar, S. Dzeroski, and T. Erjavec. Learning multilingual morphology with CLOG. Lecture Notes in Computer Science, 1446:135–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1998:UIS

[2939] N. Jacobs, K. Driessens, and L. De Raedt. Using ILP systems for verification and validation of multi-agent systems. Lecture Notes in Computer Science, 1446:145–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakano:1998:ISH

[2940] T. Nakano, N. Inuzuka, H. Seki, and H. Itoh. Inducing shogi heuristics using inductive logic programming. Lecture Notes in Computer Science, 1446: 155–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khan:1998:RLU

[2941] K. Khan, S. Muggleton, and R. Parson. Repeat learning using predicate invention. Lecture Notes in Computer Science, 1446:165–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogel:1998:NPM

[2942] L. Fogel and G. Zaverucha. Normal programs and multiple predicate learning. Lecture Notes in Computer Science, 1446:175–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flach:1998:STI

[2943] P. A. Flach, C. Giraud-Carrier, and J. W. Lloyd. Strongly typed inductive concept learning. Lecture Notes in Computer Science, 1446:185–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nock:1998:FFH

[2944] R. Nock and P. Jappy. Functionfree Horn clauses are hard to approximate. Lecture Notes in Computer Science, 1446:195–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hekanaho:1998:DGB

[2945] J. Hekanaho. DOGMA: A GA-based relational learner. Lecture Notes in Computer Science, 1446:205–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markov:1998:GUI

[2946] Z. Markov. Generalization under implication by lambda-subsumption. Lecture Notes in Computer Science, 1446: 215–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sammut:1998:PRR

[2947] C. Sammut. Prolog, refinements and RLGG’s. Lecture Notes in Computer Science, 1446:225–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zucker:1998:LSI

[2948] J.-D. Zucker and J.-G Ganascia. Learning structurally indeterminate clauses. Lecture Notes in Computer Science, 1446:235–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muggleton:1998:CIE

[2949] S. Muggleton. Completing inverse entailment. Lecture Notes in Computer Science, 1446:245–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nienhuys-Cheng:1998:DLH

[2950] S.-H Nienhuys-Cheng. Distances and limits on Herbrand interpretations. Lecture Notes in Computer Science, 1446:250–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirsten:1998:RDB

[2951] M. Kirsten and S. Wrobel. Relational distance-based clustering. Lecture Notes in Computer Science, 1446: 261–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramon:1998:FDD

[2952] J. Ramon and M. Bruynooghe. A framework for defining distances between first-order logic objects. Lecture Notes in Computer Science, 1446:271– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzeroski:1998:DTP

[2953] S. Dzeroski, N. Jacobs, M. Molina, and C. Moure. Detecting traffic problems with ILP. Lecture Notes in Computer Science, 1446:281–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roberts:1998:CIP

[2954] S. Roberts, W. Van Laer, N. Jacobs, and S. Muggleton. A comparison of ILP and propositional systems on propositional traffic data. Lecture Notes in Computer Science, 1446:291–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dearholt:1998:CCC

[2955] D. W. Dearholt. The cumulative consensus of cognitive agents in scenarios: A framework for evolutionary processes in semantic memory. Lecture Notes in Computer Science, 1447:3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tesfatsion:1998:PPS

[2956] L. Tesfatsion. Preferential partner selection in evolutionary labor markets: A study in agent-based computational economics. Lecture Notes in Computer Science, 1447:15–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bankes:1998:SPE

[2957] S. Bankes. Subspace pursuit for exploratory modeling. Lecture Notes in Computer Science, 1447:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaufils:1998:CCS

[2958] B. Beaufils, J.-P. Delahaye, and P. Mathieu. Complete classes of strategies for the classical iterated prisoner’s dilemma. Lecture Notes in Computer Science, 1447:33–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bull:1998:ECM

[2959] L. Bull. Evolutionary computing in multi-agent environments: Operators. Lecture Notes in Computer Science, 1447:43–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olsson:1998:ESH

[2960] B. Olsson. Evalution of a simple hostparasite genetic algorithm. Lecture Notes in Computer Science, 1447:53– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walker:1998:TTP

[2961] J. F. Walker. Testing three paradigms for evolving groups of cooperative, simple agents. Lecture Notes in Computer Science, 1447:63–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ashlock:1998:AGA

[2962] D. Ashlock and J. E. Mayfield. Acquisition of general adaptive features by evolution. Lecture Notes in Computer Science, 1447:75–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myung:1998:HIL

[2963] H. Myung and J.-H. Kim. Hybrid interior-lagrangian penalty based evolutionary optimization. Lecture Notes in Computer Science, 1447:85– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breeden:1998:GOF

[2964] J. L. Breeden. GA-Optimal fitness functions. Lecture Notes in Computer Science, 1447:95–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1998:SEP

[2965] X. Yao and Y. Liu. Scaling up evolutionary programming algorithms. Lecture Notes in Computer Science, 1447: 103–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomon:1998:SNS

[2966] R. Salomon. Short notes on the schema theorem and the building block hypothesis in genetic algorithms. Lecture Notes in Computer Science, 1447:113– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeck:1998:SEA

[2967] T. Baeck, A. E. Eiben, and M. E.

Vink. A superior evolutionary algorithm for 3-SAT. Lecture Notes in Computer Science, 1447:125–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chau:1998:EHC

[2968] P. M. Chau, G. Clark, and A. V. Sebald. Evolvable hardware control for dynamic reconfigurable and adaptive computing. Lecture Notes in Computer Science, 1447:137–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirsh:1998:EPS

[2969] J. E. Hirsh and D. K. Young. Evolutionary programming strategies with self-adaptation applied to the design of rotorcraft using parallel processing. Lecture Notes in Computer Science, 1447:147–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chellapilla:1998:OTP

[2970] K. Chellapilla, R. Sathyanarayan, and A. Hoorfar. Optimization of thinned phased arrays using evolutionary programming. Lecture Notes in Computer Science, 1447:157–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosinski:1998:EDC

[2971] W. Kosinski, M. Weigl, and Z. Michalewicz. Evolutionary domain covering of an inference system for function approximation. Lecture Notes in Computer Science, 1447:167–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rychtyckyj:1998:LRE

[2972] N. Rychtyckyj and R. G. Reynolds. Learning to re-engineer semantic networks using cultural algorithms. Lecture Notes in Computer Science, 1447: 181–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ostrowski:1998:ISM

[2973] D. Ostrowski and R. G. Reynolds. Integration of slicing methods into a cultural algorithm in order to assist in large-scale engineering systems design. Lecture Notes in Computer Science, 1447:191–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Louis:1998:GSO

[2974] S. J. Louis, G. Bebis, S. Uthiram, and Y. Varol. Genetic search for object identification. Lecture Notes in Computer Science, 1447:199–??, 1998.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1998:FCA

[2975] S. Zhu and R. G. Reynolds. Fuzzy cultural algorithms with evolutionary programming. Lecture Notes in Computer Science, 1447:209–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chung:1998:CES

[2976] C.-J. Chung and R. G. Reynolds. Culturing evolution strategies to support the exploration of novel environments by an intelligent robotic agent. Lecture Notes in Computer Science, 1447: 219–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gessler:1998:SCA

[2977] N. Gessler. Skeuomorphs and cultural algorithms. Lecture Notes in Computer Science, 1447:229–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoenauer:1998:SOT

[2978] M. Schoenauer and Z. Michalewicz.

Sphere operators and their applicability for constrained parameter optimization problems. Lecture Notes in Computer Science, 1447:241–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1998:NMI

[2979] T. Fernandez and M. Evett. Numeric mutation as an improvement to symbolic regression in genetic programming. Lecture Notes in Computer Science, 1447:251–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sprave:1998:VDO

[2980] J. Sprave and S. Rolf. Variabledimensional optimization with evolutionary algorithms using fixed-length representations. Lecture Notes in Computer Science, 1447:261–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebald:1998:MPEa

[2981] A. V. Sebald and K. Chellapilla. On making problems evolutionarily friendly. part 1: Evolving the most convenient representations. Lecture Notes in Computer Science, 1447:271– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebald:1998:MPEb

[2982] A. V. Sebald and K. Chellapilla. On making problems evolutionarily friendly. part 2: Evolving the most convenient coordinate systems within which to pose (and solve) the given problem. Lecture Notes in Computer Science, 1447:281–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liang:1998:EIS

[2983] K.-H. Liang, X. Yao, Y. Liu, and C. Newton. An experimental investigation of self-adaptation in evolutionary programming. Lecture Notes in Computer Science, 1447:291–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hart:1998:AEP

[2984] W. E. Hart. On the application of evolutionary pattern search algorithms. Lecture Notes in Computer Science, 1447:303–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogel:1998:STM

[2985] D. B. Fogel and A. Ghozeil. The schema theorem and the misallocation of trials in the presence of stochastic effects. Lecture Notes in Computer Science, 1447:313–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beyer:1998:EPE

[2986] H.-G Beyer. On the “explorative power” of ES/EP-like algorithms. Lecture Notes in Computer Science, 1447: 323–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomon:1998:RAG

[2987] R. Salomon. Resampling and its avoidance in genetic algorithms. Lecture Notes in Computer Science, 1447:335– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rudolph:1998:ESM

[2988] G. Rudolph. Evolutionary search for minimal elements in partially ordered finite sets. Lecture Notes in Computer Science, 1447:345–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macready:1998:TML

[2989] W. G. Macready. Tailoring mutation to landscape properties. Lecture Notes in Computer Science, 1447:355–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moore:1998:GPM

[2990] F. W. Moore and O. N. Garcia. A genetic programming methodology for missile countermeasures optimization under uncertainty. Lecture Notes in Computer Science, 1447:367–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1998:EAV

[2991] I. K. Evans. Evolutionary algorithms for vertex cover. Lecture Notes in Computer Science, 1447:377–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thacore:1998:ESL

[2992] S. Thacore. An evolutionary selflearning methodology: Some preliminary results from a case study. Lecture Notes in Computer Science, 1447: 387–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sundaralingam:1998:EIF

[2993] S. Sundaralingam and K. Sharman. Evolving IIR filters in multipath environments. Lecture Notes in Computer Science, 1447:397–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murata:1998:FPI

[2994] T. Murata, H. Ishibuchi, T. Nakashima, and M. Gen. Fuzzy partition and input selection by genetic algorithms for designing fuzzy rule-based classification systems. Lecture Notes in Computer Science, 1447:407–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chellapilla:1998:ENC

[2995] K. Chellapilla. Evolving nonlinear controllers for backing up a truck-and-trailer using evolutionary programming. Lecture Notes in Computer Science, 1447:417–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogel:1998:RDS

[2996] G. B. Fogel, K. Chellapilla, and D. B.

Fogel. Reconstruction of DNA sequence information from a simulated DNA chip using evolutionary programming. Lecture Notes in Computer Science, 1447:429–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koza:1998:UPM

[2997] J. R. Koza, F. H. Bennett, and D. Andre. Using programmatic motifs and genetic programming to classify protein sequences as to cellular location. Lecture Notes in Computer Science, 1447:437–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehlhaar:1998:FAR

[2998] D. K. Gehlhaar, D. Bouzida, and P. A.

Rejto. Fully automated and rapid flexible docking of inhibitors covalently bound to serine proteases. Lecture Notes in Computer Science, 1447:449– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfaffmann:1998:MNM

[2999] J. O. Pfaffmann and M. Conrad. Microtubule networks as a medium for adaptive information processing. Lecture Notes in Computer Science, 1447: 463–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brewster:1998:EIM

[3000] J. Brewster and M. Conrad. Evolve IV: A metabolically-based artificial ecosystem model. Lecture Notes in Computer Science, 1447:473–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaffe:1998:SMS

[3001] K. Jaffe. Sex, mate selection, and evolution. Lecture Notes in Computer Science, 1447:483–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenwood:1998:FLE

[3002] G. W. Greenwood and Y.-P. Liu. Finding low energy conformations of atomic clusters using evolution strategies. Lecture Notes in Computer Science, 1447:493–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1998:EDN

[3003] R. W. Anderson, J. B. Badler, and E. L. Keller. Estimating the distribution of neural connections in the saccadic system using a biologically plausible learning rule — preliminary results. Lecture Notes in Computer Science, 1447:503–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamont:1998:EAC

[3004] G. B. Lamont, S. M. Brown, and G. H.

Gates. Evolutionary algorithms combined with deterministic search. Lecture Notes in Computer Science, 1447: 517–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ozcan:1998:SSM

[3005] E. Ozcan and C. K. Mohan. Steady state memetic algorithm for partial shape matching. Lecture Notes in Computer Science, 1447:527–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ashlock:1998:FCT

[3006] D. Ashlock and J. I. Lathrop. A fully characterized test suite for genetic programming. Lecture Notes in Computer Science, 1447:537–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mengshoel:1998:GAB

[3007] O. J. Mengshoel and D. C. Wilkins.

Genetic algorithms for belief network inference: The role of scaling and niching. Lecture Notes in Computer Science, 1447:547–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1998:BSF

[3008] M. A. Smith. Building software frameworks for evolutionary computation. Lecture Notes in Computer Science, 1447:557–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dunning:1998:RSD

[3009] T. Dunning. Recorded step directional mutation for faster convergence. Lecture Notes in Computer Science, 1447: 569–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennedy:1998:BP

[3010] J. Kennedy. The behavior of particles. Lecture Notes in Computer Science, 1447:581–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1998:PSP

[3011] Y. Shi and R. C. Eberhart. Parameter selection in particle swarm optimization. Lecture Notes in Computer Science, 1447:591–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angeline:1998:EOV

[3012] P. J. Angeline. Evolutionary optimization versus particle swarm optimization: Philosophy and performance differences. Lecture Notes in Computer Science, 1447:601–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eberhart:1998:CBG

[3013] R. C. Eberhart and Y. Shi. Comparison between genetic algorithms and particle swarm optimization. Lecture Notes in Computer Science, 1447:611– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wicker:1998:HEL

[3014] D. Wicker, M. M. Rizki, and L. A.

Tamburino. A hybrid evolutionary learning system for synthesizing neural network pattern recognition systems. Lecture Notes in Computer Science, 1447:619–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skourikhine:1998:EAD

[3015] A. N. Skourikhine. An evolutionary algorithm for designing feedforward neural networks. Lecture Notes in Computer Science, 1447:629–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FigueiraPujol:1998:DNR

[3016] J. C. Figueira Pujol and R. Poli. Dual network representation applied to the evolution of neural controllers. Lecture Notes in Computer Science, 1447:637– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imada:1998:WDL

[3017] A. Imada and K. Araki. What does the landscape of a Hopfield associative memory look like? Lecture Notes in Computer Science, 1447:647–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chorazyczewski:1998:VEA

[3018] A. Chorazyczewski and R. Galar. Visualization of evolutionary adaptation in R0n. Lecture Notes in Computer Science, 1447:659–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patton:1998:ACD

[3019] A. L. Patton, T. Dexter, E. D. Goodman, and W. F. Punch. On the application of cohort-driven operators to continuous optimization problems using evolutionary computation. Lecture Notes in Computer Science, 1447:671– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stanhope:1998:OMC

[3020] S. A. Stanhope and J. M. Daida. Optimal mutation and crossover rates for a genetic algorithm operating in a dynamic environment. Lecture Notes in Computer Science, 1447:693–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menczer:1998:LS

[3021] F. Menczer and R. K. Belew. Local selection. Lecture Notes in Computer Science, 1447:703–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McDonnell:1998:AMS

[3022] J. R. McDonnell. Asymmetric mutations for stochastic search. Lecture Notes in Computer Science, 1447:713– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brotherton:1998:ARE

[3023] T. W. Brotherton and G. Chadderdon.

Automated rule extraction for engine health monitoring. Lecture Notes in Computer Science, 1447:725–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stanhope:1998:GPA

[3024] S. A. Stanhope and J. M. Daida. Genetic programming for automatic target classification and recognition in synthetic aperture radar imagery. Lecture Notes in Computer Science, 1447: 735–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aler:1998:EHP

[3025] R. Aler, D. Borrajo, and P. Isasi.

Evolving heuristics for planning. Lecture Notes in Computer Science, 1447: 745–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liang:1998:SCS

[3026] K.-H. Liang, X. Yao, C. Newton, and D. Hoffman. Solving cutting stock problems by evolutionary programming. Lecture Notes in Computer Science, 1447:755–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kammeyer:1998:EIE

[3027] T. E. Kammeyer and R. K. Belew. An empirical investigation of an evolutionary algorithm’s ability to maintain a known good solution. Lecture Notes in Computer Science, 1447:767–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Porto:1998:EIL

[3028] V. W. Porto. Evolving integrated lowlevel behaviors into intelligently interactive simulated forces. Lecture Notes in Computer Science, 1447:777–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:1998:UOI

[3029] C. Davis and C. F. Eick. Using offset invariant crossover as a tool for discovering cycle lengths of a periodic function. Lecture Notes in Computer Science, 1447:789–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuscu:1998:EGB

[3030] I. Kuscu. Evolving a generalized behaviour: Artificial ant problem revisited. Lecture Notes in Computer Science, 1447:799–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahluwalia:1998:CEF

[3031] M. Ahluwalia and L. Bull. Coevolving functions in genetic programming: Dynamic ADF creation using GLiB. Lecture Notes in Computer Science, 1447:809–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waagen:1998:ESL

[3032] D. Waagen, J. Pecina, and R. Pickens. Evolving spatially-localized projection filters for SAR automatic target recognition. Lecture Notes in Computer Science, 1447:819–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1998:GPO

[3033] S.-H. Chen and C.-H Yeh. Genetic programming in the overlapping generations model: An illustration with the dynamics of the inflation rate. Lecture Notes in Computer Science, 1447: 829–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myers:1998:FBV

[3034] G. Myers. A fast bit-vector algorithm for approximate string matching based on dynamic programming. Lecture Notes in Computer Science, 1448:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Navarro:1998:BPA

[3035] G. Navarro and M. Raffinot. A bitparallel approach to suffix automata: Fast extended string matching. Lecture Notes in Computer Science, 1448: 14–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziv-Ukelson:1998:DMA

[3036] M. Ziv-Ukelson and A. Kershenbaum.

A dictionary matching algorithm fast on the avenge for terms of varying lengths. Lecture Notes in Computer Science, 1448:34–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charras:1998:VFS

[3037] C. Charras, T. Lecroq, and J. D. Pehoushek. A very fast string matching algorithm for small alphabets and long patterns. Lecture Notes in Computer Science, 1448:55–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Risvik:1998:AWS

[3038] K. M. Risvik. Approximate word sequence matching over spase suffix trees. Lecture Notes in Computer Science, 1448:65–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1998:EPA

[3039] K. Zhang. Efficient parallel algorithm for the editing distance between ordered trees. Lecture Notes in Computer Science, 1448:80–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myers:1998:REA

[3040] E. G. Myers, P. Oliva, and K. Guimaraes. Reporting exit and approximate regular expression matches. Lecture Notes in Computer Science, 1448:91– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1998:AOD

[3041] Y. Yang, K. Zhang, X. Wang, and J. T. L. Wang. An approximate oracle for distance in metric spaces. Lecture Notes in Computer Science, 1448: 104–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fredriksson:1998:RIF

[3042] K. Fredriksson and E. Ukkonen. A rotation invariant filter for twodimensional string matching. Lecture Notes in Computer Science, 1448:118– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1998:CSA

[3043] D. K. Kim, Y. A. Kim, and K. Park.

Constructing suffix arrays for multidimensional matrices. Lecture Notes in Computer Science, 1448:126–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoye:1998:SFD

[3044] J. Stoye and D. Gusfield. Simple and flexible detection of contiguous repates using a suffix tree. Lecture Notes in Computer Science, 1448:140–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedersen:1998:CCD

[3045] C. N. S. Pedersen, R. Lyngs, and J. Hein. Comparison and coding DNA. Lecture Notes in Computer Science, 1448:153–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ma:1998:FTA

[3046] B. Ma, L. Wang, and M. Li. Fixed topology alignment with recombination. Lecture Notes in Computer Science, 1448:174–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kececioglu:1998:AA

[3047] J. D. Kececioğlu and W. Zhang. Aligning alignments. Lecture Notes in Computer Science, 1448:189–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1998:ESC

[3048] A. Amir, G. M. Landau, M. Lewenstein, and N. Lewenstein. Efficient special cases of pattern matching with swaps. Lecture Notes in Computer Science, 1448:209–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hua:1998:ADS

[3049] Y. Hua, T. Jiang, and B. Wu. Aligning DNA sequences to minimize he changes in protein. Lecture Notes in Computer Science, 1448:221–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Mabrouk:1998:GH

[3050] N. El-Mabrouk, J. H. Nadeau, and D. Sankoff. Genome halving. Lecture Notes in Computer Science, 1448:235– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadimitriou:1998:AAI

[3051] C. H. Papadimitriou. Algorithmic approaches to information retrieval and data mining. Lecture Notes in Computer Science, 1449:1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1998:CPA

[3052] Fan Chung Graham. Combinatorial problems arising in massive data sets. Lecture Notes in Computer Science, 1449:2–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1998:EPM

[3053] M. J. Fischer. Estimating parameters of monotone Boolean functions. Lecture Notes in Computer Science, 1449:3–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosaraju:1998:AA

[3054] S. Rao Kosaraju and M. Pop. Deamortization of algorithms. Lecture Notes in Computer Science, 1449:4–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dai:1998:CNC

[3055] Y. Dai and N. Katoh. On computing new classes of optimal triangulations with angular constraints. Lecture Notes in Computer Science, 1449:15– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1998:MWTa

[3056] Cao An Wang, F. Y. Chin, and Bo Ting Yang. Maximum weight triangulation and its application on graph drawing. Lecture Notes in Computer Science, 1449:25–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1998:CSS

[3057] T. Graf, V. Kamakoti, N. S. Janaki Latha, and C. Pandu Rangan. The colored sector search tree: A dynamic data structure for efficient high dimensional nearest-foreign-neighbor queries. Lecture Notes in Computer Science, 1449:35–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1998:SEA

[3058] D. Z. Chen and O. Daescu. Spaceefficient algorithms for approximating polygonal curves in two dimensional space. Lecture Notes in Computer Science, 1449:45–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atallah:1998:PGA

[3059] M. J. Atallah and D. Z. Chen. Parallel geometric algorithms in coarsegrain network models. Lecture Notes in Computer Science, 1449:55–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleischer:1998:BP

[3060] R. Fleischer. On the bahncard problem. Lecture Notes in Computer Science, 1449:65–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez-Ortiz:1998:USS

[3061] A. Lopez-Ortiz and S. Schuierer. The ultimate strategy to search on m rays? Lecture Notes in Computer Science, 1449:75–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1998:BAD

[3062] M. Li and L. Zhang. Better approximation of diagonal-flip transformation and rotation transformation. Lecture Notes in Computer Science, 1449:85– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xue:1998:COP

[3063] G. Xue. A cost optimal parallel algorithm for computing force field in N-body simulations. Lecture Notes in Computer Science, 1449:95–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yagiura:1998:EFN

[3064] M. Yagiura and T. Ibaraki. Efficient 2 and 3-Flip neighborhood search algorithms for the MAX SAT. Lecture Notes in Computer Science, 1449:105– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egecioglu:1998:AAU

[3065] O. Egecioğlu and M. Peinado. Algorithms for almost-uniform generation with an unbiased binary source. Lecture Notes in Computer Science, 1449: 117–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBonis:1998:IAC

[3066] A. De Bonis, L. Gargano, and U. Vaccaro. Improved algorithms for chemical threshold testing problems. Lecture Notes in Computer Science, 1449: 127–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiwi:1998:MMB

[3067] M. Kiwi, D. A. Spielman, and S.- H Teng. Min-max-boundary domain decomposition. Lecture Notes in Computer Science, 1449:137–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reith:1998:BLB

[3068] S. Reith and K. W. Wagner. On Boolean lowness and Boolean highness. Lecture Notes in Computer Science, 1449:147–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hertrampf:1998:IDB

[3069] U. Hertrampf. The inherent dimension of bounded counting classes. Lecture Notes in Computer Science, 1449:157– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1998:ECS

[3070] C.-J. Lu. An exact characterization of symmetric functions in qAC00[2]. Lecture Notes in Computer Science, 1449: 167–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1998:RR

[3071] J.-Y. Cai, L. A. Hemaspaandra, and G. Wechsung. Robust reductions. Lecture Notes in Computer Science, 1449: 174–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brattka:1998:AES

[3072] V. Brattka, K. Weihrauch, and X. Zheng. Approaches to effective semicontinuity of real functions. Lecture Notes in Computer Science, 1449:184– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grebinski:1998:PAC

[3073] V. Grebinski. On the power of additive combinatorial search model. Lecture Notes in Computer Science, 1449: 194–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1998:NRS

[3074] F. K. Hwang and W.-D. Lin. The number of rearrangements in a 3stage Clos network using an auxiliary switch. Lecture Notes in Computer Science, 1449:204–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsai:1998:LBW

[3075] K.-H. Tsai and D.-W Wang. Lower bounds for wide-sense non-blocking Clos network. Lecture Notes in Computer Science, 1449:213–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1998:MMS

[3076] D. S. Kim and D.-Z Du. Multirate multicast switching networks. Lecture Notes in Computer Science, 1449:219– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwama:1998:ERR

[3077] K. Iwama, E. Miyano, S. Tajima, and H. Tamaki. Efficient randomized routing algorithms on the two-dimensional mesh of buses. Lecture Notes in Computer Science, 1449:229–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1998:EDJ

[3078] T.-S. Hsu and D. R. Lopez. Executing divisible jobs on a network with a fixed number of processors. Lecture Notes in Computer Science, 1449:241–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mans:1998:ACC

[3079] B. Mans, F. Pappalardi, and I. Shparlinski. On the Adam conjecture on circulant graphs. Lecture Notes in Computer Science, 1449:251–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zang:1998:PTC

[3080] W. Zang. Proof of Toft’s conjecture: Every graph containing no fully odd K4 is 3-colorable. Lecture Notes in Computer Science, 1449:261–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1998:NFO

[3081] J.-J. Wang, T.-Y. Sung, L.-H. Hsu, and M.-Y. Lin. A new family of optimal 1-Hamiltonian graphs with small diameter. Lecture Notes in Computer Science, 1449:269–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peng:1998:LTA

[3082] S.-L. Peng, C.-W. Ho, T.-S. Hsu, and M.-T. Ko. A linear-time algorithm for constructing an optimal node-search strategy of a tree. Lecture Notes in Computer Science, 1449:279–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1998:GOP

[3083] J. Wang and Y. Ge. A graph optimization problem in virtual colonoscopy. Lecture Notes in Computer Science, 1449:289–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1998:AEA

[3084] Bang Ye Wu, K.-M. Chao, and Chuan Yi Tang. Approximation and exact algorithms for constructing minimum ultrametric trees from distance matrices. Lecture Notes in Computer Science, 1449:299–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chao:1998:OAF

[3085] H. S. Chao, F. R. Hsu, and R. C. T.

Lee. An optimal algorithm for finding the minimum cardinality dominating set on permutation graphs. Lecture Notes in Computer Science, 1449: 309–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeza-Yates:1998:STD

[3086] R. A. Baeza-Yates. Similarity in twodimensional strings. Lecture Notes in Computer Science, 1449:319–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alber:1998:MDH

[3087] J. Alber and R. Niedermeier. On multidimensional Hilbert indexings. Lecture Notes in Computer Science, 1449:329– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernasconi:1998:CPC

[3088] A. Bernasconi. Combinatorial properties of classes of functions hard to compute in constant depth. Lecture Notes in Computer Science, 1449:339– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizuki:1998:ESK

[3089] T. Mizuki, H. Shizuya, and T. Nishizeki. Eulerian secret key exchange. Lecture Notes in Computer Science, 1449:349– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaver:1998:OTT

[3090] D. Beaver. One-time tables for twoparty computation. Lecture Notes in Computer Science, 1449:361–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ausiello:1998:HTR

[3091] Giorgio Ausiello, Giuseppe F. Italiano, and Umberto Nanni. Hypergraph traversal revisited: Cost measures and dynamic algorithms. Lecture Notes in Computer Science, 1450: 1–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500001.
                                        pdf.

Borger:1998:DJV

[3092] Egon Börger and Wolfram Schulte. Defining the Java Virtual Machine as platform for provably correct Java compilation. Lecture Notes in Computer Science, 1450:17–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500017.
pdf.

Harel:1998:TTR

[3093] David Harel. Towards a theory of recursive structures. Lecture Notes in Computer Science, 1450:36–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500036.
pdf.

Kesten:1998:MAK

[3094] Yonit Kesten and Amir Pnueli. Modularization and abstraction: The keys to practical formal verification. Lecture Notes in Computer Science, 1450: 54–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500054.
pdf.

Maass:1998:RTS

[3095] Wolfgang Maass. On the role of time and space in neural computation. Lecture Notes in Computer Science, 1450:

72–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500072.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500072.
                                        pdf.

Mehlhorn:1998:AWP

[3096] Kurt Mehlhorn and Stefan Näher. From algorithms to working programs: On the use of program checking in LEDA. Lecture Notes in Computer Science, 1450:84–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500084.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500084.
                                        pdf.

Micali:1998:CSC

[3097] Silvio Micali. Computationallysound checkers. Lecture Notes in Computer Science, 1450:94–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500094.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500094.
                                        pdf.

Nielsen:1998:RAP

[3098] Mogens Nielsen. Reasoning about the past. Lecture Notes in Computer Science, 1450:117–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1450/14500117.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1450/14500117.pdf.

Pudlak:1998:SAL

[3099] Pavel Pudlák. Satisfiability — algorithms and logic. Lecture Notes in Computer Science, 1450:129–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500129.
pdf.

Stirling:1998:JB

[3100] Colin Stirling. The joys of bisimulation.

Lecture Notes in Computer Science, 1450:142–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1450/14500142.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1450/14500142.
pdf.

Wiedermann:1998:TAE

[3101] Jirí Wiedermann. Towards algorithmic explanation of mind evolution and functioning. Lecture Notes in Computer Science, 1450:152–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500152.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1450/14500152.
pdf.

Aldaz:1998:CHP

[3102] Mikel Aldaz, Joos Heintz, Guillermo Matera, José L. Montaña, and Luis M. Pardo. Combinatorial hardness proofs for polynomial evaluation. Lecture Notes in Computer Science, 1450:167– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500167.
pdf.

Alekhnovich:1998:MPP

[3103] Michael Alekhnovich, Sam Buss, Shlomo Moran, and Toniann Pitassi. Minimum propositional proof length is NP-Hard to linearly approximate. Lecture Notes in Computer Science, 1450: 176–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500176.
pdf.

Chrobak:1998:RPS

[3104] Marek Chrobak and Christoph Dürr. Reconstructing polyatomic structures from discrete X-rays: NP-completeness proof for three atoms. Lecture Notes in Computer Science, 1450:185–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1450/14500185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500185.
pdf.

Kuzjurin:1998:LEC

[3105] Nikolai N. Kuzjurin. Locally explicit construction of Rödl’s asymptotically good packings. Lecture Notes in Computer Science, 1450:194–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500194.
pdf.

Baaz:1998:PTF

[3106] Matthias Baaz, Agata Ciabattoni, Christian Fermüller, and Helmut Veith. Proof theory of fuzzy logics: Urquhart’s C and related logics. Lecture Notes in Computer Science, 1450: 203–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500203.
pdf.

Bonner:1998:NLP

[3107] Richard Bonner, Rusins Freivalds, Janis Lapins, and Antra Lukjanska. Nonstochastic languages as projections of 2-Tape quasideterministic languages. Lecture Notes in Computer Science, 1450:213–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500213.
                                        pdf.

Nielson:1998:FLI

[3108] Flemming Nielson and Hanne Riis Nielson. Flow logic for imperative objects. Lecture Notes in Computer Science, 1450:220–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1450/14500220.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500220.
                                        pdf.

Rabinovich:1998:ECT

[3109] Alexander Rabinovich. Expressive completeness of temporal logic of action. Lecture Notes in Computer Science, 1450:229–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1450/14500229.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1450/14500229.pdf.

Ferreira:1998:RAT

[3110] Maria C. F. Ferreira, Delia Kesner, and Laurence Puel. Reducing ACtermination to termination. Lecture Notes in Computer Science, 1450:239– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500239.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1450/14500239.
pdf.

Fueloep:1998:OPT

[3111] Z. Fueloep, E. Jurvanen, M. Steinby, and S. Vagvoelgyi. On one-pass term rewriting. Lecture Notes in Computer Science, 1450:248–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fulop:1998:OPT

[3112] Zoltán Fülöp, Eija Jurvanen, Magnus Steinby, and Sándor Vágvölgyi. On one-pass term rewriting. Lecture Notes in Computer Science, 1450:248– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500248.
pdf.

Hermann:1998:WSC

[3113] Miki Hermann and Gernot Salzer.

On the word, subsumption, and complement problem for recurrent term schematizations. Lecture Notes in Computer Science, 1450:257–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500257.
pdf.

Touzet:1998:EHB

[3114] Hélène Touzet. Encoding the hydra battle as a rewrite system. Lecture Notes in Computer Science, 1450:267– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500267.
pdf.

Hagenah:1998:CFN

[3115] Christian Hagenah and Anca Muscholl. Computing ϵ-free NFA from regular expressions in O(n log^{2}(n)) time. Lecture Notes in Computer Science, 1450:277– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500277.
pdf.

Latteux:1998:ILP

[3116] Michel Latteux, David Simplot, and Alain Terlutte. Iterated lengthpreserving rational transductions. Lecture Notes in Computer Science, 1450: 286–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500286.
pdf.

Petersen:1998:HHO

[3117] Holger Petersen. The head hierarchy for oblivious finite automata with polynomial advice collapses. Lecture Notes in Computer Science, 1450:296– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500296.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500296.
pdf.

Senizergues:1998:EPD

[3118] Géraud Sénizergues. The equivalence problem for deterministic pushdown transducers into Abelian groups. Lecture Notes in Computer Science, 1450: 305–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500305.
pdf.

Barthe:1998:SFC

[3119] Gilles Barthe. The semi-full closure of pure type systems. Lecture Notes in Computer Science, 1450:316– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500316.
pdf.

Benke:1998:PPS

[3120] Marcin Benke. Predicative polymorphic subtyping. Lecture Notes in Computer Science, 1450:326–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500326.
                                        pdf.

Bierman:1998:CIC

[3121] G. M. Bierman. A computational interpretation of the λμ-calculus. Lecture Notes in Computer Science, 1450: 336–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500336.
                                        pdf.

Chrzaszcz:1998:PSD

[3122] Jacek Chrzaszcz. Polymorphic subtyping without distributivity. Lecture Notes in Computer Science, 1450:346– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500346.
                                        pdf.

Gastin:1998:NEM

[3123] Paul Gastin, Raphaël Meyer, and Antoine Petit. A (non-elementary) modular decision procedure for LTrL. Lecture Notes in Computer Science, 1450: 356–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500356.
pdf.

Giacobazzi:1998:CAI

[3124] Roberto Giacobazzi, Francesco Ranzato, and Francesca Scozzari. Complete abstract interpretations made constructive. Lecture Notes in Computer Science, 1450:366–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1450/14500366.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1450/14500366.pdf.

Hune:1998:TBO

[3125] Thomas Hune and Mogens Nielsen.

Timed bisimulation and open maps. Lecture Notes in Computer Science, 1450:378–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1450/14500378.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1450/14500378.
pdf.

Srba:1998:DSC

[3126] Jirí Srba. Deadlocking states in context-free process algebra. Lecture Notes in Computer Science, 1450:388– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500388.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500388.
                                        pdf.

Amano:1998:SLB

[3127] Kazuyuki Amano and Akira Maruoka. A superpolynomial lower bound for a circuit computing the clique function with at most (1/6) log log n negation gates. Lecture Notes in Computer Science, 1450:399–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500399.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500399.
                                        pdf.

Ambainis:1998:CAC

[3128] Andris Ambainis, David Mix Barrington, and Huong LêThanh. On counting AC^{0} circuits with negative constants. Lecture Notes in Computer Science, 1450:409–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500409.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500409.
                                        pdf.

Hemaspaandra:1998:SST

[3129] Lane A. Hemaspaandra and Jörg Rothe. A second step towards circuit complexity-theoretic analogs of Rice’s theorem. Lecture Notes in Computer Science, 1450:418–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500418.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500418.
pdf.

Emerson:1998:MCR

[3130] E. Allen Emerson and Richard J. Trefler. Model checking real-time properties of symmetric systems. Lecture Notes in Computer Science, 1450:427– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500427.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500427.
pdf.

Grohe:1998:LOI

[3131] Martin Grohe and Thomas Schwentick.

Locality of order-invariant firstorder formulas. Lecture Notes in Computer Science, 1450:437–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500437.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500437.
pdf.

DiPierro:1998:PCC

[3132] A. Di Pierro and H. Wiklicky. Probabilistic concurrent constraint programming: Towards a fully abstract model.

Lecture Notes in Computer Science, 1450:446–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierro:1998:PCC

[3133] Alessandra Di Pierro and Herbert Wiklicky. Probabilistic concurrent constraint programming: Towards a fully abstract model. Lecture Notes in Computer Science, 1450:446–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500446.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500446.
                                        pdf.

Simpson:1998:LFA

[3134] Alex K. Simpson. Lazy functional algorithms for exact real functionals. Lecture Notes in Computer Science, 1450: 456–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500456.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500456.
                                        pdf.

Ambos-Spies:1998:RVC

[3135] Klaus Ambos-Spies, Steffen Lempp, and Gunther Mainhardt. Randomness vs. completeness: On the diagonalization strength of resourcebounded random sets. Lecture Notes in Computer Science, 1450:465–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1450/14500465.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500465.
pdf.

Bentzien:1998:PTT

[3136] Levke Bentzien. Positive Turing and truth-table completeness for NEXP are incomparable. Lecture Notes in Computer Science, 1450:474–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500474.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500474.
pdf.

Goldsmith:1998:TNS

[3137] Judy Goldsmith, Mitsunori Ogihara, and Jörg Rothe. Tally NP sets and easy census functions. Lecture Notes in Computer Science, 1450:483– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500483.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500483.
pdf.

Kobler:1998:ACI

[3138] Johannes Köbler and Rainer Schuler.

Average-case intractability vs. worstcase intractability. Lecture Notes in Computer Science, 1450:493–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1450/14500493.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500493.
                                        pdf.

Koebler:1998:ACI

[3139] J. Koebler and R. Schuler. Averagecase intractability vs. worst-case intractability. Lecture Notes in Computer Science, 1450:493–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harju:1998:STS

[3140] Tero Harju, Alexandru Mateescu, and Arto Salomaa. Shuffle on trajectories: The Schützenberger product and related operations. Lecture Notes in Computer Science, 1450:503–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500503.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500503.
                                        pdf.

Kuich:1998:GEC

[3141] Werner Kuich. Gaussian elimination and a characterization of algebraic power series. Lecture Notes in Computer Science, 1450:512–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500512.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500512.
                                        pdf.

Lopez:1998:DSS

[3142] Luis-Miguel Lopez and Philippe Narbel. D0L-systems and surface automorphisms. Lecture Notes in Computer Science, 1450:522–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500522.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500522.
pdf.

Ryl:1998:ASL

[3143] Isabelle Ryl, Yves Roos, and Mireille Clerbout. About synchronization languages. Lecture Notes in Computer Science, 1450:533–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1450/14500533.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1450/14500533.pdf.

Barthelmann:1998:WCE

[3144] Klaus Barthelmann. When can an equational simple graph be generated by hyperedge replacement? Lecture Notes in Computer Science, 1450:543– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500543.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500543.
pdf.

Grosse-Rhode:1998:STR

[3145] Martin Große-Rhode, Francesco Parisi-Presicce, and Marta Simeoni. Spatial and temporal refinement of typed graph transformation systems. Lecture Notes in Computer Science, 1450:553– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500553.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500553.
                                        pdf.

Hofmeister:1998:AMI

[3146] Thomas Hofmeister and Hanno Lefmann. Approximating maximum independent sets in uniform hypergraphs. Lecture Notes in Computer Science, 1450:562–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1450/14500562.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500562.
                                        pdf.

LaTorre:1998:RHG

[3147] S. La Torre and M. Napoli. Representing hyper-graphs by regular languages. Lecture Notes in Computer Science, 1450:571–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torre:1998:RHG

[3148] Salvatore La Torre and Margherita Napoli. Representing hyper-graphs by regular languages. Lecture Notes in Computer Science, 1450:571–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500571.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500571.
pdf.

Iwama:1998:ITS

[3149] Kazuo Iwama and Chuzo Iwamoto.

Improved time and space hierarchies of one-tape off-line TMs. Lecture Notes in Computer Science, 1450:580– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500580.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500580.
pdf.

Mielniczuk:1998:TSC

[3150] Pawel Mielniczuk and Leszek Pacholski. Tarskian set constraints are in NEXPTIME. Lecture Notes in Computer Science, 1450:589–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500589.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500589.
pdf.

Vorobyov:1998:ETC

[3151] Sergei Vorobyov. ∀∃^{∗}-equational theory of context unification is Π_{1}^{0}-hard. Lecture Notes in Computer Science, 1450: 597–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500597.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500597.
                                        pdf.

Wiedermann:1998:SNS

[3152] Jirí Wiedermann. Speeding-up nondeterministic single-tape off-line computations by one alternation. Lecture Notes in Computer Science, 1450:607– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500607.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500607.
                                        pdf.

Courcelle:1998:FCP

[3153] Bruno Courcelle and Denis Lapoire. Facial circuits of planar graphs and context-free languages. Lecture Notes in Computer Science, 1450:616–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500616.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500616.
                                        pdf.

Iwama:1998:OOS

[3154] Kazuo Iwama, Mitsushi Nozoe, and Shuzo Yajima. Optimizing OB-DDs is still intractable for monotone functions. Lecture Notes in Computer Science, 1450:625–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500625.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500625.
pdf.

Preuss:1998:BVO

[3155] Harry Preuß and Anand Srivastav. Blockwise variable orderings for shared BDDs. Lecture Notes in Computer Science, 1450:636–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500636.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500636.
pdf.

Slobodova:1998:CPO

[3156] Anna Slobodová. On the composition problem for OBDDs with multiple variable orders. Lecture Notes in Computer Science, 1450:645–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500645.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500645.
pdf.

Choffrut:1998:ETS

[3157] Christian Choffrut and Sandor Horvath. Equations in transfinite strings. Lecture Notes in Computer Science, 1450:656–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/1450/14500656.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500656.
                                        pdf.

Crochemore:1998:MFW

[3158] M. Crochemore, F. Mignosi, and A. Restivo. Minimal forbidden words and factor automata. Lecture Notes in Computer Science, 1450:665–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500665.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500665.
                                        pdf.

Karhumaeki:1998:DEB

[3159] J. Karhumaeki, J. Manuch, and W. Plandowski. On defect effect of bi-infinite words. Lecture Notes in Computer Science, 1450:674–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karhumaki:1998:DEB

[3160] Juhani Karhumäki, Ján Manuch, and Wojciech Plandowski. On defect effect of bi-infinite words. Lecture Notes in Computer Science, 1450:674– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500674.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500674.
                                        pdf.

Kolpakov:1998:RFB

[3161] Roman Kolpakov, Gregory Kucherov, and Yuri Tarannikov. On repetitionfree binary words of minimal density. Lecture Notes in Computer Science, 1450:683–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1450/14500683.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1450/14500683.
pdf.

Bezrukov:1998:EHG

[3162] S. L. Bezrukov, J. D. Chavez, L. H. Harper, M. Röttger, and U.- P. Schroeder. Embedding of hypercubes into grids. Lecture Notes in Computer Science, 1450:693–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500693.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500693.
pdf.

Bodlaender:1998:TDS

[3163] Hans L. Bodlaender and Torben Hagerup. Tree decompositions of small diameter. Lecture Notes in Computer Science, 1450:702–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500702.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500702.
pdf.

Broersma:1998:DPF

[3164] Hajo Broersma, Andreas Huck, Ton Kloks, Otto Koppius, Dieter Kratsch, Haiko Müller, and Hilde Tuinstra. Degree-preserving forests. Lecture Notes in Computer Science, 1450:713– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500713.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500713.
pdf.

Crauser:1998:PDS

[3165] A. Crauser, K. Mehlhorn, U. Meyer, and P. Sanders. A parallelization of Dijkstra’s shortest path algorithm. Lecture Notes in Computer Science, 1450: 722–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500722.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500722.
pdf.

Durand:1998:CBC

[3166] Bruno Durand and Sylvain Porrot. Comparison between the complexity of a function and the complexity of its graph. Lecture Notes in Computer Science, 1450:732–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500732.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500732.
pdf.

Fernau:1998:ICL

[3167] Henning Fernau and Ludwig Staiger.

IFS and control languages. Lecture Notes in Computer Science, 1450:740– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500740.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500740.
pdf.

Matz:1998:OQW

[3168] Oliver Matz. One quantifier will do in existential monadic second-order logic over pictures. Lecture Notes in Computer Science, 1450:751–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500751.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500751.
pdf.

Reinhardt:1998:SRP

[3169] Klaus Reinhardt. On some recognizable picture-languages. Lecture Notes in Computer Science, 1450:760– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500760.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500760.
pdf.

Auletta:1998:CWC

[3170] Vincenzo Auletta, Ioannis Caragiannis, Christos Kaklamanis, and Pino Persiano. On the complexity of wavelength converters. Lecture Notes in Computer Science, 1450:771–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500771.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500771.
pdf.

Damm:1998:BVM

[3171] Carsten Damm. On Boolean vs. modular arithmetic for circuits and communication protocols. Lecture Notes in Computer Science, 1450:780–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500780.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500780.
pdf.

Hromkovic:1998:CCL

[3172] Juraj Hromkovic. Communication complexity and lower bounds on multilective computations. Lecture Notes in Computer Science, 1450:789–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500789.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500789.
pdf.

Weihrauch:1998:FHR

[3173] Klaus Weihrauch and Xizhong Zheng. A finite hierarchy of the recursively enumerable real numbers. Lecture Notes in Computer Science, 1450:798– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500798.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500798.
pdf.

Buchholz:1998:OGO

[3174] Thomas Buchholz, Andreas Klein, and Martin Kutrib. One guess oneway cellular arrays. Lecture Notes in Computer Science, 1450:807–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500807.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500807.
pdf.

Cattaneo:1998:TDC

[3175] Gianpiero Cattaneo and Luciano Margara. Topological definitions of chaos applied to cellular automata dynamics. Lecture Notes in Computer Science, 1450:816–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1450/14500816.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1450/14500816.
pdf.

Manzini:1998:CSL

[3176] Giovanni Manzini. Characterization of sensitive linear cellular automata with respect to the counting distance. Lecture Notes in Computer Science, 1450: 825–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500825.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500825.
                                        pdf.

Mazoyer:1998:ACA

[3177] Jacques Mazoyer and Ivan Rapaport. Additive cellular automata over Z_{p} and the bottom of (CA,≤ ). Lecture Notes in Computer Science, 1450:834– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1450/14500834.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1450/14500834.
                                        pdf.

Tiskin:1999:EBS

[3178] A. Tiskin. Erratum: Bulk-synchronous parallel multiplication of Boolean matrices. Lecture Notes in Computer Science, 1644:717–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). See [2842].